A Complete Bibliography of Publications in the Journal of Systems and Software (2010–2019)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

23 April 2025 Version 3.02

Title word cross-reference

(k, n) [YC11]. (n, t, n) [LHYZ12]. 1000 [ABJ^{+}17]. 2 [BMAH11]. 3 [GCLD13, JSL16, MKH^{+}12]. 3 + 1 [Fug12]. ^{2} [NJ17]. ^{th} [KT16]. α [TTL10]. HV^{2}M [CBZ^{+}16]. i^{∗} [DCG16, MNSA16]. K [HKS^{+}17, Cho13, LZ12, MLLK11, SHN14, Zha12b]. M [MMSD13]. M^{3} [DAG19]. n [SPSR17]. O(1) [PNY14]. p [hChSyCwL10]. q [GMS11].

* [TTC15].

-band [MMSD13]. -gram [SPSR17]. -hour [ABJ^{+}17]. -nearest [Cho13, LZ12]. -SDH [GMS11]. -SIP [hChSyCwL10]. -trimmed [TTL10].

.Net [QOLJG16].

1 [KJ10]. 103 [HST16]. 109c [Woh16]. 11 [KT16]. 128 [TSLL11]. 13-round [TSLL11]. 133 [YMM^{+}19]. 148 [WLL19a]. 192/256 [LGLL12]. 1H [JCYT16].

2.0 [BCG^{+}13, CIB^{+}19, GCC^{+}15, GLJ13, OGK13]. 2009 [CL11]. 2017 [WB19]. 2153 [TTT14]. 23rd [Bor12]. 256 [LGLL12].

35th [WC16]. 3E [ZGSH13]. 3G [Ski13]. 3GPP [EZOK14].

4.0 [DSGS17]. 4G [WCC13].

5 [WL15a]. 5W [JCYT16].

802.11 [WC11]. 83 [AAH12b]. 84 [YWEL^{+}13]. 85 [WZM12a, XTZX13]. 86 [BKSM14, TTT14, wZfG14a].

9th [LH12].

AADL [MBPM19, YHM^{+}14]. AAL [NAB^{+}13]. ABC [YAY13]. abilities [WS13]. ability [WS12, ZXL10]. abnormal [GHBD^{+}16]. absorptive [MRM16]. Abstract [YHM^{+}14, AR12, OMLB16, PC10]. abstraction [CS16, SKE10]. abstractions [OBS^{+}18]. academia [GK18]. Academic [BKW10, KBJZ15]. Accelerated [AN10, PS14]. accelerating [KMK17]. acceleration [MN19, XZZ^{+}16]. accelerator [GGK19, RBT11]. accelerator-based [RBT11]. acceptance [DLW^{+}13, OD17, PHR10, RBS19, SCC16, VHL14]. access [CH10b, FBB15, FNWL18, GAT15, KKL^{+}11, LLLK12, LH11b, MJ18, MGM10, NZM10, PCCB^{+}11, SM17a, WH15, WS12, ZML17, BDGP13]. accesses [EAH^{+}11]. accessing [MCV15]. accountable [ZZ12]. accounting [Al 12]. accuracy [CS15, RSB^{+}14]. Accurate [LLZW14, PPMM17, TAB^{+}16, BNS12, PSM12, ZCY^{+}16]. Achieving [ADET12, AGR19, NSDI16, PDL^{+}16, SLZ12, SPMG18]. ACL [PGRQVV12]. acquisition [Kel15, LMT16, NK15]. across [PAB^{+}17]. Action [MAR^{+}19, BP13, TTMI19]. Action-Oriented [MAR^{+}19]. actions [KHC16, SDB16, VvSvV16]. active [LCLP16, YTW^{+}13]. Activities [SSR18, Al 12, AAN11, ROR11, RDVC19, Xia13]. activity [BS12, ÇB16, DC17, LNY^{+}11, MS16, SGMHJ13]. actor [BSK^{+}18, CDRT13]. Actual [ETM10]. acyclic [LWLL12]. Ad [ACSC16, ACL13, BCLW11, hChSyCwL10, CWK10, Cho13, KSHC14, LLHY19, MLHL12, MDO^{+}10, WOC15]. ad-hoc [hChSyCwL10, MDO^{+}10]. Adam [XCM^{+}12]. adaptability [PPMM14]. Adaptable [EMSU11]. adaptation [APM^{+}14, ADET12, BGEP17, BBD18, BMLL14, CCdL^{+}16, CPYZ14, CG12, EGG^{+}11, FCB^{+}16, GBH^{+}16, GSP^{+}19, GDSB11, JS13, JS16, Pot13, SH17, VSS^{+}11, XCM^{+}12]. adapted [TPGdS13]. Adapter [XPBC11]. Adapting [MHB18, SH17, BJG11, HGP^{+}12, SBB^{+}16, PH13]. Adaptive [AR12, ABB15, CKMT10, CT11a, CKC15, HyLW^{+}12, KRD16, LCLF13, LXG10, MSHG18, PWLL13, Aki18, AG15, ARS17, BSK^{+}18, Bar15, CCdL^{+}16, CLH^{+}13, CHLW17, CZC^{+}18, CGPT14, EEAZ13, FS19, FGBC10, GZKL13, HWR17, JC15, KKG^{+}12, LXC13, LWW^{+}10, LZR16, LYC14, MLLK11, MPN^{+}17, MCS^{+}12, ND18, PCHW12, PPMM12, PPMM17, PZ15, QXYL16, SAA^{+}10, SB17a, SYBN12, SRWE10, TJT^{+}18, WDC12, WCX15, WMAS12, WKH11, YXP^{+}18, LLC17, SD16b]. adaptivity [ZHGL11]. Adding [KCR16, MTF14]. address [PN14]. addressing [GSN^{+}15]. adequate [DW11]. adjustable [WL17]. Adjusting [MG11, Oja16a, ST11, CV14]. adjustment [ANM15, MCC^{+}18]. ADL [PFF12, WRTP^{+}13]. ADLs [WB15]. admitted [MKS^{+}18, SSK19]. adopt [PWS^{+}15]. Adopter [RNR17]. Adopting [LPB19, TTMI19]. adoption [BdMSNO^{+}17, CCP18, Che17, EGHO16, FB18, GN15, JLL19, KCV^{+}19, KKA^{+}19, MRM16, NHH^{+}12, SG12, Wu11]. adoptive [SS12]. Advanced [LJA^{+}11]. Advancing [SW19]. adversary [dOCS13]. advertising [AM10b]. AES [BBBP13]. aesthetic [LQLC16]. AFChecker [LXC13]. affect [GGC16, NBM19, SMSH18, SKK^{+}18b]. affecting [LRD^{+}19, MP12, PWS^{+}15, Wu11, ZZP15, ZP17]. affiliations [ESM19a]. Affinity [SK13]. Affinity-aware [SK13]. affordable [CCG^{+}10]. after [KQ17]. against [BBBP13, HHH^{+}10a, KKHH11, OLV15, TLL13]. age [SSMvD16]. agenda [AS16, FS17]. Agent [CCG^{+}10, GGS15, LN13, ATHM17, AN16, BM17, CNKL12, GMPN16, GRBNA10, GCC^{+}15, GZKL13, GGM11, ISM11, JS16, KB16, LG17, MIBV14, SPTM15, SST16, ZMB14]. Agent-based [GGS15]. Agent-oriented [CCG^{+}10, LN13, ISM11]. agents [JSM10, JRO12, SÁMI17]. Aggregate [HCT^{+}15, LCC10, Shi10, YDGB^{+}12]. Aggregate-strength [HCT^{+}15]. aggregated [CCMOM19]. Aggregation [Bar15, AKB11, MT10, SGBCP12, YCWW15]. Agile [CP15, DvdVA^{+}13, GN15, KSM^{+}16, MB10, ASG17, AVGM19, AdB17, CCP18, CLL14, CBVF19, CNMR18, DPL16, DNBM12, DCT17, DGCA17, FFdRG^{+}14, FMRM15, GTF15, GTF17, HM16, KM14, LSD^{+}16, MHB18, MVSG18, NBF16, NBF^{+}19, OK18, nPHW^{+}16, SS12, SDG17, SNDC13, Sta14, SHHL12, WK15, WCC12, YLA16b, DCP12, HL10, JMML17, LGC17, OD17, RKK16]. Agile-developed [OD17]. agility [GTF15, JWA14]. aging [ACW10, PMMM11]. agnostic [ASMM18]. agreement [IB11, TLL12, OHJ10]. agreements [FSG^{+}11, IYS13]. Ahab [SCMS15]. ahead [YCF^{+}13]. aided [MB17]. AIOLOS [VSDD12]. air [HWHT11, MPTT14]. AJAX [YM13]. alarming [BRG^{+}12]. Alberta [GV10]. algebraic [BBA10, Pra18]. algorithm [AR18, AG15, BAI^{+}14, CL17a, CJ13, CWK^{+}13, CS12, DFJ19, EEAZ13, GWW^{+}11, HY11, HWL13a, JK13, JXLC15, JHYK10, KHS11, KSRD10, KSN17, KKHH11, KLB15, LPP15, LLZW14, LHY12, LZ13, LL14, Lin16, LLML13, MMZ^{+}16, MH11, MJZ^{+}10, NDS13, OCC12, PS13, PQBP16, PRN17, PS14, PM10, RFM10, RLL^{+}18, SLW^{+}15, TGKL19, jT12, TTL10, WGZ^{+}12, WX10, YCLY13, YH19, YH10, ZLW^{+}12, ZL17, ZWF^{+}18, ZL12b, ZGL^{+}10, FAM15].

Algorithms [FGLI15, BMAH11, DRCA^{+}19, FFV19, HBT16, LWOY16, PLHP^{+}15, PCCK18, RCCVB11, SA11, TdCAF16, VTZ^{+}17, wZfG13, wZfG14a]. Alibaba [DLW^{+}13]. aligned [WMW12]. Aligning [GGT^{+}19, VvSvV16, CRESF^{+}13]. alignment [LMR12, UGFK15, VLC^{+}17]. all-port [MV10]. Alleviating [MARD16]. allocation [BMAH11, BV15, BGLG13, CAG17, DM17b, GQ12, HNH15, KHSD10, Luk11, NK15, SWES16, TY18, ZWC^{+}19]. allowance [RXY^{+}19]. Alone [ESM19a]. alphabets [Kan15]. alternate [ZLW^{+}12]. alternating [WCB^{+}17]. alternatives [KK12]. Always [GFP11]. AM [KKP12]. Ambient [ARS10, ABB15, GMPN16, BHH^{+}12, RASL12]. Ambient-PRISMA [ARS10]. Ambients [ARS10]. among [GE15b, HG18]. amongst [RHRC13]. amorphous [BKSM13, BKSM14]. amount [EEAZ13]. amplification [DVPY^{+}19]. Analogy [ANC11, ANM15, IAA16]. Analogy-based [ANC11, ANM15, IAA16]. analyses [SNDD19]. Analysing [SB17a, SYBN12]. Analysis [AV12, CDPM17, SW10, SSP17, WCTK12, AAMS14, AAMS16, AMAY19, AHH^{+}10, AHW10, AS10, ACRD19, AHBA19, BL19, BLL^{+}18, BHH^{+}12, BLTY18, BGG10, BWH10, BRS10, BCL12, BSB12, CCGG14, CS15, CS16, CCN^{+}10, CMM15, CL17b, CVGP13, CCMOM19, DC17, DSGS17, DZT^{+}14, DH13, DOL^{+}16, EH19, EBJ17, FDN^{+}16, FBD^{+}18, FSGYP17, FP18, FMdAR16, GCDY16, GGT^{+}19, HGBS18, HBT16, HRB12, HSL14, HCL^{+}10, HCC10a, HWLM11, IZ18, IYS13, KK17a, KRDH12, KM17, KGW12, KR14, KRJ17, KCV^{+}19, Kim12, KKP12, LJH10, LKJR10a, LKJR10b, LS14, LGL^{+}10, LJM11, MGM10, MJ14, MTF14, MK17, MVSG18, MA10, MNM12, MMTS15, MMB10, NSAK10, dONTF^{+}19, NSM17, OHL17, OMLB16, PDS19, PH13, PPMM17, PDBD18, QBO^{+}14, RAS14, RPT19, RGH17, RITF^{+}11, RASL12, SG16, ŠK11, SS12, SCwY12, SGC^{+}17, dSSVV11, SRBT18]. analysis [SZS13, SLLL14, SLL^{+}15, TNK^{+}19, TXCX19, TCS18, VTZ^{+}17, VCMG17, VHFF^{+}17, WCC12, WLZ^{+}17a, WZY^{+}18, WGKW19, WKV11, WV11, WMOKY11, XYCL17, YAY13, YLXZ16, YFT^{+}15, ZYZ^{+}17, ZZP15, ZP17, dB12, JR15, MS17b]. analyst [SJ17]. analysts [RDPM19]. analytic [LZKW12, MB19]. Analytical [EK13, LJC16, VRPT18, MV11]. Analytics [UB19, CBVF19, KCR16, LLH^{+}16, VZT17]. analyze [SGMHJ13]. Analyzer [PÁC13, KH10]. Analyzing [CL15, CMR19, JLGM17, KG10, RSB^{+}16, BKRW19, CTKT13, KJS^{+}12, MBPM19, PB11, YAKK16, JLZ^{+}19, vAAJ16]. anatomy [TKP^{+}18]. anchors [LJ16]. and/or [HHKWB16]. Android [AAM^{+}17, BCA^{+}19, CCL^{+}19, HBM19, JLL19, KFLS18, LZL^{+}18, LRB^{+}19, LVVTP17, LWBH16, OBS^{+}18, SPCT18, TKZW17, YGN^{+}16]. Android-related [LRB^{+}19]. AndroidOff [CCL^{+}19]. angle [PKS18]. animation [KA14]. annealing [MK15b]. anniversary [WC16]. annotation [HA10, LGM^{+}18]. annotations [SM17b]. anomalies [ZXC^{+}17]. Anomaly [SKK^{+}18a, BLL^{+}18, WWZ^{+}14]. anonymity [MK15a, YL16]. answers [TLWS10]. ant [MDO^{+}10, TJH15]. Antecedents [GA11, LCCJ10]. anti [FHL^{+}18, MSK^{+}17, QZ14, Sta10]. anti-forensics [QZ14]. anti-patterns [FHL^{+}18, MSK^{+}17, Sta10]. antipatterns [KVGS11, TKCR14]. anytime [DRCA^{+}19]. AOSD [Ano13a]. Apache [MK17]. APDL [KSKP11]. aperiodic [OD10]. API [BHVR18, CKCK15, EZG15, KFLS18, NKZ17, SM17b, SPSR17, SCO13, XSL^{+}18]. APIs [Sal17, SSSA17]. app [GNA17, JED18, LLL^{+}17b, RM19a, vAAJ16]. appliances [ZDC^{+}11, ZDC^{+}11]. applicability [JWA14]. Application [FKWVH19, FCMJ12, AV12, AIE19, AR12, AYZI10, ANG^{+}19, ASMM18, AdAD17, AF16, BBBP13, BCF18, BCL^{+}18, BGG10, CS16, CDA11, CM15, CCL^{+}19, DBCG14, EAH^{+}11, FTSC12, HyLW^{+}12, HWLM11, HSS14, JS13, JRO12, KK17a, KSHC14, LWZ12, MJZ^{+}10, NHH^{+}12, PC15, PHR10, SCGL^{+}18, SCC16, SP14, TTMI19, VSDD12, WCC12, YLYL17, ZSG16, ZYZL12, ZS16, Zha16, DFCPSF15, FM11]. application-domain [SP14]. applications [AdB13, AAC16, AAB19, AHOP14, ABFM12, BPQP^{+}10, BZ14, BSDD14, BAAD17, BCS18, BK17, BPB19, CG15, CdCAdO18, CLR18, CRKH11, CCGdL16, CRESF^{+}13, CF12, CGPT14, DHC^{+}11, DS16b, DAG19, GE15a, GRBNA10, GBCI11, GZKL13, HGP^{+}12, HCY19, HVK11, HS15, KRJ17, KVH12, KQ17, LXJL10, LZJ^{+}19, LCJ10, LZHS11, LXC13, LASL14, MBD13, MGR^{+}13, MK15b, NOPF12, NK15, NBR^{+}14, OGK13, OD17, PDK^{+}16, PLHP^{+}15, PG15, PMMM11, QGZ^{+}15, RAS14, RHHT18, RLY^{+}13, RAJ15, RB16, RMD11, SRWE10, SC14, Shi17, SC19, SBB^{+}16, SLLL12, TKZW17, TJT^{+}18, TAF^{+}17, UIK17, VVA^{+}15, VSS^{+}11, WVT^{+}14, WWZ^{+}14, WXY^{+}17, WH15, YXH^{+}18, YXP^{+}18, YM13, YGN^{+}16, ZTZ^{+}11, CCCY17]. applied [FFV19, PPG^{+}13]. Applying [BRS^{+}18, FSGL12, KHMF13, PLHP^{+}15, AdB17, MGB16, PCFRP19, RSB^{+}16, RMCH^{+}14, ZFS15]. Appraisal [OKMD12]. Approach [CPDM16, HP16, SDB16, Aba13, AIE19, AdB13, AMKD13, AM15, AHBA19, AGC13, ACH19, AF16, AdB17, APS^{+}10, AGR19, BML^{+}13, BKH10, FGB^{+}19, CG15, CF13, CWK^{+}11, CCY11, CHLW17, CZC^{+}18, CBZ^{+}16, CJL11, CHL^{+}13, CAG17, CHCO11, CKL12, CLF^{+}13, CKS15, CGPT14, DBCdP11, DV10, DWC17, DAG19, EGHO16, EZRK16, FVHF^{+}15, FG15, GE15a, GN15, GMPN16, GGB19, GMLSF^{+}15, GCSSDP^{+}18, GPSS^{+}13, GEM15, HBM19, HdM17, HJ14, HNH15, HNS12, HPF16, HLLS13, IBM11, JS11, JG14, JC15, JCK^{+}17, JMML17, KCT12, KR14, KRJ17, KKH^{+}16, KVGS11, KY10, KKL^{+}11, KLB15, KD18, KTF^{+}16, KR16, KS16, KSS15, KHMF13, LLM^{+}17, LMN10, LMGHB17, LZJ^{+}19, LWXZ10, LT11, LLWL14, LLWL19, LASL14, MMP15, MPTT14, MFMCY12, Mer13, MdFD^{+}15, MA11, MCS^{+}12, MA17, NEM17, OZO^{+}14, PS13, PS15]. approach [PB11, PD16, PLGT10, PAR14, PCCK18, PAS^{+}10, RM19b, SCS15, SL10, SAMN12, SÁM^{+}16, SPTM15, SAKZ15, SJC13, SSP17, SHC^{+}11, SA18, SJH^{+}10, TBSvdW18, TB13, TGP11, TT13, TTT14, TTMI19, UIK17, VAM^{+}10, WDC12, WMW^{+}19, WV11, Wu11, WLD16, XYCL17, YSSaR14, YZC15, YJZ17, ZMB14, ZYZL12, ZJZ^{+}17, rBHM17, BBEM11]. Approaches [GMMGP15, VLC^{+}17, AJG^{+}15, ABCH13, AAGT16, ALRP16, BKS15, BS15, CNSG12, ESM^{+}19b, FDÁM12, GSM19, JSHW14, MKH^{+}12, PFG13, PMB15, RGV^{+}17, RM19a, SGMHJ13, SS14b, TAF^{+}17, WCC12, WNC17, dNPM18]. appropriate [LMPM18]. approval [ULS19]. Approximating [MMP15, OH15]. approximation [YH19]. apps [AAM^{+}17, JLL19, LVVTP17, PLVB^{+}18, QXYL16, SPC16]. AppSpear [LZL^{+}18]. April [KT16]. AQUArIUM [CdCAdO18]. arbitrary [ÁGBYB^{+}14, GBC16]. architect [HFLvV11, MTA^{+}16]. Architecting [DLM19, AdB17, PBM19, TZB19, PvV12]. Architectural [LL15, RAS14, UB19, BGS^{+}16, BBA10, BMB18, BWH10, CLS^{+}12, CH10c, FLRT19, GLZ15, GSP^{+}19, HZ15, KOS15, KG10, LJA^{+}11, LJDK10, MCV16, MFM10, MPLL18, PSEE12, PPMM14, SB17a, SAMN12, TKCR14, TGE17, Woo12, YLA16a, YLA^{+}17, ZMK12, dRSBA13]. architecturally [MSGM17]. Architecture [BCEF10, EB14b, KT16, LJH10, LH12, LLGZ13, MAG12, MOH16, SSR18, TKH^{+}11, ARS10, ARS17, APCS10, BKH10, BLTY18, BGG10, BCL12, BKRW19, CCdL^{+}16, CJT^{+}16, ÇT13, CHLW17, CG12, CD10, CFN10, CKS15, DYC19, EK13, ELHC13, FCB^{+}16, GAMW14, GBH^{+}16, GGvH^{+}18, GFP11, GKV14, GCLD13, GAKF13, GDSB11, GPL^{+}15, HNZ17, HBM19, HJN11, HA10, HN17, JRO12, KTT^{+}17, KGW12, KL10, KH14, LG17, LPXL10, LLX^{+}11, LLH^{+}16, LZR16, MCC^{+}18, MS16, MEB^{+}10, MKS10, MAH18, ME10, MCV15, NFSM11, NHH^{+}12, PGPC17, Pot13, RPT19, ŠK11, SA12, SLB14, SSS17, SKRB19, SHC^{+}11, SHGT16, TAJ^{+}10, TL14, TFS10, THWC10, UD10, VCB^{+}18, VHFF^{+}17, WB12, WMC17, YLA16b, ZK13, ZML10, ZMK12, dSB12, vHAH12, vHJPB^{+}17, vVT16, vdBSvS^{+}19, ÇT13, EMSU11]. Architecture-based [MOH16, CCdL^{+}16, CG12, EK13, FCB^{+}16, GDSB11, LZR16, MKS10]. Architecture-driven [MAG12, MEB^{+}10]. architectures [AB16, BD10, CBT^{+}14, CNSG12, CHL^{+}13, CV16b, CPDM16, Dut15, DAG19, GCC^{+}15, GA13, HTB12, KRD16, MCV16, MK11, MGvFGCB10, NCWK18, PN14, SRWE10, VZT17, WRTP^{+}13, WB10, BBEM11, MPRS14]. Area [MMTS15, DFG^{+}13, HBG^{+}13, LLS11, WCC13, ZÁ15]. areas [HWHT11]. argument [SGC^{+}17]. argument-fragments [SGC^{+}17]. argumentation [MOH16, YFT^{+}15]. ARIA [Kim12, TSL^{+}11]. arithmetic [LWC13]. arrays [SH17]. art [ACS13, ABL15, GAMW14, LDS^{+}19, MDP^{+}11, MRY17, PMR16, TJT^{+}18, WMAS12, CKMT10]. artery [CCWT13]. article [Ano19l]. Artifact [CFRPC^{+}18]. Artifact-based [CFRPC^{+}18]. artifacts [GE15b, SRBT18]. artificial [DRCG12, KCV11, KR16]. ask [BDDS11, vAAJ16]. aspect [ARS10, MGvFGCB10, NFSM11, NBR^{+}13, PFF12, SdSLS^{+}19, ZMB14]. aspect-mining [ZMB14]. aspect-oriented [ARS10, MGvFGCB10, NFSM11, NBR^{+}13, PFF12, SdSLS^{+}19]. AspectJ [FDN^{+}16]. aspects [CSF^{+}14, LY18, VM13]. assembly [JSM10]. assertion [DDF^{+}13]. assertion-based [DDF^{+}13]. assertions [ABS19]. assess [SFMB16, VVA^{+}15]. assessed [KM13]. Assessing [AKA^{+}15, BHH^{+}10, MPTT14, OLV15, UGFK15, VHL14, NSM17, dAGSdFS^{+}15, SM16, SJH^{+}10]. Assessment [CLL14, AD14, BP13, EFSJM17, FG15, JWA14, KCAS13, KLA^{+}19, LGM^{+}18, LMS12, LHLG^{+}15, MACB19, MGvFGCB10, PHBJ16, RDVC19, SL10, SED16, SLLL14, WTG^{+}11, Woo12, ZADA15]. asset [BWW^{+}18]. assets [LMN10, TTL^{+}13]. Assignment [KA17, BNS12, CdCMdMSNdA16, KHS11, LSE12, LCLS16, LMPM18, LZ13, LL14, MLHL12, MJ14, SAKZ15, Wen16, ZGL^{+}10]. assist [CHN19a, CKS15, SHGT16]. assistance [SPSR17]. Assisted [BHH^{+}12, RASL12, APS16, GMPN16, HHC12, WWL^{+}10, YCG^{+}14]. assisted-living [GMPN16]. Assisting [GXZ^{+}19]. assistive [MCV15, GMLSF^{+}15]. association [TL14]. assumption [YLA^{+}17, ZG10]. assumptions [YLA16a]. assurance [AS16, FFWE17, HNH15, OKMD12, WKD^{+}19]. Astor [MM19]. astronomy [DHC^{+}11]. Asymmetric [LTW16, CSS^{+}13]. Asymmetric-histogram [LTW16, CSS^{+}13]. Asynchrony [JLGM17]. ATAM [ZK13]. ATEMES [KSH^{+}12]. ATM [WMD^{+}10]. attack [CWK^{+}13, MBB11, TSL^{+}11, ZZ16, ZL12a]. attacks [BBBP13, BCR^{+}19, KPS10, KKHH11, KKP12, OLV15, RZMPM12, TLL13, jWLY^{+}13, YXH^{+}18, ZGZ^{+}13]. ATTEST [NC10]. attitudes [CLS^{+}12]. Attribute [FWCS12, BV15, FNWL18, KAM13, WGC^{+}14, ZML17].

Attribute-based

[FWCS12, WGC^{+}14, ZML17]. attributes [AAC^{+}17, HPF16, LMPM18, NC10]. ATTributes-based [NC10]. auction [BV15, KBRV17, KBRV18, TY18, ZSB19]. auction-based [KBRV18, TY18, ZSB19]. audio [BCA^{+}19, yWpNyL11]. audit [ZHAY12]. auditing [YYS^{+}16]. augmented [LGH^{+}17, SS13, VSS^{+}11]. Augmenting [ZYY^{+}19]. AUSTIN [LMH10]. authenticated [WZM12a, WZM12b, YC12, ZG10]. Authentication [MKH^{+}12, CTL12, CH10a, EA11, GJ13, GCSÁddP11, HCC10b, HS11b, IB11, Kan15, LH11a, LT13, NB13, SHBC19, TLL12, WKH11, WS13, YSL^{+}10, OHJ10]. authentication-chaining [EA11]. authoring [PSS11]. authority [CKCK15, ZZ12]. Auto [TSCB19, AIE19, DVV^{+}16]. auto-decomposition [AIE19]. Auto-scaling [TSCB19, DVV^{+}16]. automata [DPP^{+}18, WOLS12, WS13, ABCH13]. Automated [BPO^{+}16, BNS12, JSM10, JKL19, NBA^{+}17, PS13, PW18, TJH15, TSRC18, WL15b, WBS^{+}10, YFT^{+}15, ASdMGM14, AHBA19, ABC^{+}13, AS17, BSG^{+}18, CCM12, CdCMdMSNdA16, CWK^{+}11, CKS15, DW14, FKWVH19, FGLI15, HBM19, JSHW14, MG12, MGM16, SH17, SPLW17, Ski13, SWES16, VM13, VHFST15, VHFF^{+}17, YLCZ12]. Automatic [ÁGBYB^{+}14, ABL15, CL17b, CBSM16, FDÁM12, FMPS16, GLZ15, HAE^{+}15, JEEL16, KSH^{+}12, LPM15, LQLW12, LMPM18, LSLG17, WRTP^{+}13, AAM^{+}17, CCdR^{+}16, GGvH^{+}18, HZ15, HY11, HJ12, HPF16, KBHG17, LNW^{+}11, LHP^{+}09, LHP^{+}10, LDS^{+}19, MSK^{+}17, OGRJ^{+}18, PPS12, SZ11, SAKZ15, TAF^{+}17, WMW^{+}19]. Automatically [CHL^{+}19, LHG^{+}18, YFZ^{+}16, ATHM17, GPD^{+}19, RMCH^{+}14, HRRC16]. Automating [CNKL12, KKT17, LZL^{+}18, SKL10]. automation [BFLZ13, FVHF^{+}15, GCLD13, KAS18, KMK16, SJR^{+}11, WRR14]. Automotive [HBP^{+}17, DNSH13, ELHC13, MSS18, fLSN18]. autonomic [EGG^{+}11, MBT16, WTG^{+}15]. autonomous [BVV^{+}10, ETYL15, JSM10, KGG18, MSHG18, PBM19, YSDT11]. availability [BT17, OCC12, Pot13, SW10]. available [CSS10, JLQ^{+}10]. AVC [LCC^{+}13, LLML13, LW13c]. averse [Kel15]. avoid [FGBC10]. Avoiding [JSHW14, HST15, HST16]. aware [AAC16, AGBD14, AO16, BSDD14, BVV^{+}10, BCS18, CV16a, CDRT13, CYT16, CKC15, DBZ16, DFJ19, DHC^{+}11, EBJ17, FRGC10, GQ12, GDSB11, HGMB13, JLQ^{+}10, KC16, KRJ17, KSHC14, LJC16, LWL^{+}13, LZ13, LL14, LC11, LXC13, LVPMPCLS13, MOD^{+}19, MJ18, MRT17, MDO^{+}10, MV11, NK15, OB13, PCCB^{+}11, SRWE10, SGBCP12, SB19, SGEK19, SK13, TKJ15, TdCAF16, TC16b, WWL^{+}10, WWZ^{+}14, Wen16, WX10, WZJI14, XCM^{+}12, YZG^{+}13, wZfG13, wZfG14a, wZfG14b, ZWC^{+}19, ZADM10, ZW15]. Awareness [TKSRP11, AHOP14, CBC14, DM17a, EZRK16, FHY17, NBM19, NBR^{+}14, SSMvD16, TR18, UD10]. awareness- [SSMvD16].

backoff [MAAC17]. backup [CRSS14]. Backwards [CCGdL16]. Bad [KP10, WKbOS17]. bag [GGS15, PK10a, MK15b]. bag-of-tasks [GGS15, MK15b]. balance [Dan17, MB10]. balanced [MCC11, NNVD17]. balancer [CV16a]. Balancing [MSSMDC12, Woh16, BVV^{+}10, CCH14, DY15, FS19, HH17, WSM15, WOC15]. band [MMSD13]. bandwidth [NJ17]. banking [CDA11]. Barefoot [BS15]. barriers [PWS^{+}15, WRR14]. base [HCL12]. Based [AAH10, CVGP13, GMGTdFR14, HLS^{+}13, HC15, ASGJ13, ABB19, ABCH13, AN16, AAJD^{+}16, AQK11, AKAA18, AR18, AG15, AYZI10, AAH12b, AKL14, ARS17, AAN11, AGC13, AAB19, AM10a, AWSE19, AHC^{+}11, ANC11, ANM15, BRB14, BKLE18, BSK^{+}18, BM18, BSG12, BBBP13, BMLL14, BAAS13, BPGS13, BRG^{+}12, BCLW11, BDBLP15, BPSK18, CFRPC^{+}18, CCdL^{+}16, ÇT13, CFAP17, CBS16, CV14, CV16a, CCLL11, CTL12, CNL13, CPS11, CMK^{+}11, CWK10, CSS^{+}13, CW14, CXO^{+}15, CPX16, CSN^{+}17, CZC^{+}18, CCL^{+}19, CG12, CBZ^{+}16, CHL^{+}13, CFN10, CHCO11, CKL12, CBC^{+}15, CLF^{+}13, CPR13, DGBE18, DII^{+}17, DY15, DMSG11, DYC19, DPP^{+}18, DK15a, DFJ19, DH13, EAH^{+}11, EA14, EB14c, EUR^{+}13, EK13, FBB15, FWCS12, FYCL13, FVHF^{+}15, FKWVH19, FSPH^{+}16, FFV19, FCB^{+}16, FSS^{+}13, FNWL18, GKD13, GJ13, GVPPM18, GMR17, GKS18]. based [GPMI13, GS17, GFP11, GGB19, GKV14, GZY11, GHBD^{+}16, GCSÁddP11, GAKF13, DDF^{+}13, GDSB11, GMS11, GPL^{+}15, GLOM19, GGS15, HBG^{+}14, HP16, HSC15, HJBH10, HBT16, HZH^{+}16, HNH15, HSPD14, HRB12, HRRC16, HNS12, HWC^{+}10, HB13, HWL13a, HWL13b, HCC10a, HYWS11, HWLM11, HPH12, HLWS13, HKS^{+}17, HCC10b, HR10, IAA16, IB11, JS11, JK13, JC15, JXLC15, JS13, JS16, JH10, Jør10, JKL19, KAO13, KK11, KOL^{+}14, KVGS11, KU10, KY10, KKL^{+}11, KLL17, KKA^{+}19, KAM13, KB16, KTF^{+}16, KBRV18, KSS15, KKLB11, KTK19, LKR13, LCT10, LMN10, LC10, LLC10, LKL^{+}11, LESL11, LS14, LCLF13, LXJL10, LQLW12, LHZX12, LWL^{+}13, LG15, LQLC16, LGH^{+}17, LZCL19, LLHY19, LL10, LT11, LH11b, LHY12, Lin12b, LCC^{+}13, Lin14, LLWL14, LDZL15, LNY^{+}11, LBX12, LZKW12, LW13a, LZLC17, LASL14, LLL^{+}14, LHLG^{+}15, LWW^{+}10, LCL15, LTW16, LW13c, LZR16, LLGZ13]. based [MJF10, MOD^{+}19, MKS10, MLD^{+}14, MFTP18, MCV16, MBD13, MJ14, MK17, MLD16, MPN^{+}17, MOH16, MK15a, MBB11, MIUM12, MIBV14, MA11, MGM16, MCS^{+}12, MCKA18, MIKG13, NTT19, NC10, NKMM12, NCW^{+}19, ND18, NPC12, NDS13, OZO^{+}14, OD17, OLZN13, OSH^{+}18, OB13, Pal12, PEO11, PMDH13, PB15, PWY^{+}16, PAOC15, PWW10, Pen11, PCYZ12, PLGT10, PPB16, PPS12, PHR10, PWC12, PÁC13, QXYL16, RNC14, RBT11, RFM10, RRV19, RAK15, RZMPM12, RO13b, RG10, RLY^{+}13, RXY^{+}19, RSBA19, RCPZ19, ROFGFRM13, RMD11, SCMS15, SFMB16, SI12, ST13, SSMvD16, SNM14, SKE10, SRS15, SBZ^{+}17, SPLW17, Shi10, SAKZ15, SA11, SHBA^{+}16, SA16, SV12, SDB16, SHS16, SHC^{+}11, SGW^{+}15, SZW^{+}16, SHGT16, SLLL12, aSRS^{+}10, SHH^{+}15, aSRZ^{+}18, SHBC19, TY18, TG17, TKJL13, TBG13, TB13, TGKL19, TTC18, TSCB19, TPGdS13, TAB^{+}16]. based [TKCR14, TTL10, TPKT12, TXCX19, TMTB19, ULS19, UIK17, VCdA^{+}16, VKL16, VHFF^{+}17, WWYZ11, WWLG13, WCC^{+}14, WGC^{+}14, WXY^{+}17, WXZ^{+}17, WFY^{+}19, Wau19, WGZ^{+}12, WKD^{+}19, WDC10, WAWO12, WS12, WWY^{+}12, WOLS12, jWLY^{+}13, WS13, WZJI14, XJZ^{+}15, XLW18, XLM^{+}15, XZZ^{+}16, XB19b, XYZ^{+}19, YSG17, YWHL11, YCLY13, YTW^{+}13, YCC16, YCLC17, YLC18, YH19, YZL^{+}14, YBE17, YGN^{+}16, YKC^{+}12, YFT^{+}15, YZC15, YLZ^{+}16, YLYL17, YZY^{+}18, ZTZ^{+}11, ZLZ11, ZXTT11, ZLW^{+}12, ZM12, ZT14, ZML17, ZHH^{+}17, ZYZ^{+}17, ZWM^{+}18, ZCZZ11, ZZ12, ZGZ^{+}13, Zha16, ZWF^{+}18, ZL12b, ZLmLN14, ZSB19, ZYY^{+}19, dACM17, dL13, dCPV10, dNPM18, WL10, BLUH15, NBH19, TKSRP11]. basic [KP10]. basis [EVR11, SXYW14]. batch [AR18, SRS15]. Bayesian [DCT17, KVGS11, PRN17, RSB^{+}16, RBW18, SXYW14, YLYL17]. BDTEX [KVGS11]. Be [WLL17, FFdRG^{+}14, KM13, ZZ16, ZXC^{+}17]. beam [JC15]. bee [MCS^{+}12]. before [AS10]. beginning [GPD^{+}19]. behavior [Ala15, AAB19, BPGS13, CRSS14, LGH^{+}17, LLHY19, NJ17, OK11, dMCR19, KMWL12]. behavior-driven [NJ17]. Behavioral [LFW15, HJBH10, LZLC17]. behaviors [JKC19]. behaviour [ABJ^{+}17, BPQP^{+}10, HL10]. behavioural [BZ10]. belief [AC16]. benchmark [BGEP17, CDOP15]. benchmarking [FMdAR16, SA11, VVA^{+}15]. benchmarks [SPC16]. Benefits [BS12, BGG10, FDÁM12, LMT16, TTR^{+}13, ZGYS^{+}15]. Bessel [GJ13]. Best [GFP11, KT16, CL11, NI13, OZO^{+}14, PFL16, ZADA15]. better [ABL16, SRSC16, TC16a]. between [AJLS10, AC17, BDD^{+}15, BCD^{+}18, CFMRL11, CRC19, CKL12, EZOK14, GBDCR12, GKV14, HSM16, IF19, IBAH12, KBDGAW16, MER17, OBS^{+}18, PPMM14, SCGL^{+}18, SÁMI17, SBDB19, TR18, TGE17, VCMG17, WMW12, vdRBSvV10]. Beyond [dMSSS^{+}13, ZK13, AZX14, AT15, Bos12, GL14, MM19]. bi [DRCA^{+}19, VRPT18]. bi-objective [DRCA^{+}19]. bibliometric [KLA^{+}19]. bidding [GGC16]. bidirectional [DGWC16, SHC^{+}11]. bidirectional-transformation-based [SHC^{+}11]. Big [SKT17, TLK16b, UB19, GPL^{+}15, LDZL15, SGW^{+}15, VTZ^{+}17, XLM^{+}15, YF15, Dut15, FGD^{+}17]. bilateral [jT12]. binary [PQBP16, WCC^{+}14, WLZ^{+}17b, ZLmLN14]. Bioinformatics [PM10]. biometric [GCSÁddP11]. bit [PMDH13]. bit-rate [PMDH13]. bitstream [QZ12]. BitTorrent [KA14]. Black [AAB19, BAAD17, CF13, ZZ12]. Black-box [AAB19, CF13, ZZ12]. Blackboard [JRO12]. blame [DSGS17]. Blending [CSF^{+}14]. blind [yWpWyYpN13]. block [GCSSDP^{+}18, KM11, LCLF13, WCC^{+}14, ZL12a]. Blockchain [TMTB19]. Blockchain-based [TMTB19]. blocking [TGKL19, VGSN18]. blog [TPTV17]. blogs [DV10]. blood [HHC12]. BN [PSNB11]. Board [Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano11k, Ano11l, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano13j, Ano13k, Ano13l, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano17l, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano19a, Ano19b, Ano19c, Ano19d]. Board [Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k]. boolean [CHN19b, CW14, Kim17, YCC16]. Boolean-based [CW14, YCC16]. boost [CBZ^{+}16, LLC17]. Boosting [RNC14, ZHGL11, MRJD^{+}12, ROFGFRM13]. bootstrapping [AHH^{+}10]. Bord [BRG^{+}12]. Bord-and-Pillar [BRG^{+}12]. both [LWLL12, WCCL10, YLXZ16]. BotMosaic [HB13]. botnets [HB13]. Bottom [PK10b]. Bottom-up [PK10b]. bound [DAG19, SMZC12]. boundaries [Bos12]. box [AAB19, BBEM11, BAAD17, CF13, KCAS13, ZZ12]. BPEL [LQLW12, aSRZ^{+}18]. BPM [LGH^{+}17, THWC10]. BPM-oriented [THWC10]. branch [LMH10, PS13]. Brazil [CCP18, CVGP13, DFG^{+}13, Gar13, LCM^{+}13, NAB^{+}13, dMSSS^{+}13]. Brazilian [Bor12, Ano13a, WWSS13]. breaking [RvDV17]. breath [LSR13]. brick [SBAH17]. Bridging [CKL12, HS11a, TR18, LVPMPCLS13, SBDB19]. Bringing [BBEM11, BMKM15, NTdSX13]. Broadcast [RLL^{+}18, CLL10, HST15, HST16, MV11, NSAK10, SM17a, ZZ12]. broadcasting [MK11]. Broker [WZJI14, KAK^{+}13]. Broker-based [WZJI14]. brokering [BVV^{+}10]. brownout [XB19b]. BrownoutCon [XB19b]. BSN [HY11]. buffering [YZG^{+}13]. Bug [MRS18, ACB18, BNS12, CPZF19, HCY19, HK13, LYLC16, MRRS19, SAKZ15, ZCC^{+}19, ZM18]. bug-fixing [ACB18]. Bug-proneness [MRS18]. buggy [WMW^{+}19]. bugs [CPZF19, CPRT16, ECS15, VGSN18, WLL17, YLCZ12, ZCY^{+}16, ZFY^{+}19, IBAH12]. build [ABJ10, CdR^{+}14, SCC16]. Building [GZKL13, BSG12, GSM19, GCSSDP^{+}18, GTF17, KH14, TG17, LJDK10]. bulk [HSS10]. bullying [GGM11]. Bundle [ST11]. bursts [SAA^{+}10]. bursty [BP15, LJM11, PPMM17, WMOKY11]. bus [CTL10]. Business [ML18, PCFRP19, BGLG13, CFRPC^{+}18, CO12, CLF^{+}13, FDÁM12, FSG^{+}11, GBDCR12, LPM15, LMGHB17, MSGGL12, Oja16b, OFR^{+}12, PCCLdGP12, ŠK11, SJ17, SS14a, SDB18, SLR16, SK18, SSAS11, TAF^{+}17, VKL16, VvSvV16, RCL14]. Bypassing [TKJ16]. byte [Kim12].

c [KRDH12, dSACdLF17, CWK^{+}11, EBC10, KTK19, LMH10, MN19]. c-means [KRDH12]. C/C [KTK19]. cache [ÁRMC16, OB13, TDW^{+}14, nWsCqW12, CWK^{+}13]. cache-aware [OB13]. caching [GLJ13, NTT19]. CALA [HRRC16]. calibration [LHP^{+}09, LHP^{+}10, NBH19]. Call [QGZ^{+}15]. calls [TLZ^{+}16]. cam [PKS18]. Camellia [LGL^{+}10, LGLL12]. Camellia-192 [LGLL12]. Camellia-192/256 [LGLL12]. cameras [MKH^{+}12]. Can [vAAJ16, ZXC^{+}17]. Canada [GZ13]. Canadian [GV10]. cancelled [AS10]. capabilities [KCR16]. capability [LLM^{+}17, LT13]. capacity [AQK11, BK17, CAG17, LCT10, LBCL10, Lin12b, LCC^{+}13, MRM16, PWLL13, WLH13, WCC^{+}14]. Capitals [Woh16, WSM15]. capstone [VLL18]. Capturing [CBL^{+}15, MH11, PKS18, YAKK16]. card [TBSvdW18, ABFM12]. cards [HCC10b, KKP12, YSL^{+}10]. care [HWdS^{+}15]. carotid [CCWT13]. cascades [RNC14]. Case [RB16, SSP17, AAGT16, ABC^{+}13, AWSE19, BAM17, BS12, BAAD17, CKMT10, CXO^{+}15, CZC^{+}18, CFA^{+}19, EA12, EA14, EA19, EVR11, ELHC13, FMdAR16, GGK19, GPPT16, GSdS16, GEM15, HGBS18, Han12, HWC^{+}10, HCC10a, HPH12, IF10, JWA14, JCYT16, JC15, JAS19, JR15, Kan15, KOS15, KFN19, KJS^{+}12, KVH12, KSM^{+}16, KMG^{+}19, LQLW12, LWZ12, LPB19, MAH18, MVSG18, MCTM11, MPLL^{+}15, MMTS15, PPG^{+}13, PSS^{+}16, PAB^{+}17, PCCLdGP12, PWA^{+}19, RAS14, RASL12, SAA^{+}10, SS12, Shi12, SS14a, SDB18, SGC^{+}17, SCC16, TKP^{+}18, UGFK15, VTZ^{+}17, WGKW19, WKD^{+}19, WRR14, WLD16, YLA^{+}17, ZLL^{+}12, ZYA^{+}18, ZAY19, dB12, dSdMSNO^{+}14, vHAT13, TKSRP11]. cases [KSM^{+}16, LLWL19, TAS^{+}18, WZY^{+}18, ZYZ^{+}17, ZZC18]. Casper [CBSM16]. catalogs [dAGSdFS^{+}15]. catalogue [EL10]. catalogues [DV10]. categorization [ASMM18, BCL^{+}18]. category [YFZ^{+}16, ZA12, CPX16]. Category-choice [CPX16]. causality [CBSM16, CPV^{+}14]. CBRJS [KTK19]. CCA [SLZ12]. CCA2 [RG10, ZZ12]. CDH [ZG10]. CDL [WKZL10]. celebrate [WC16]. Cell [WCC13, AAMS14, GAT15]. Cell-related [WCC13]. cellular [WOLS12, WS13]. censorship [Shi12]. center [LZL^{+}15]. centered [KSKP11, ZÁ15]. centred [LSLG17]. centric [CCY11, OBS^{+}18, PCG^{+}14, RHHT18, WWY^{+}12]. Certificate [YLZ^{+}16, LHZX12]. Certificate-based [YLZ^{+}16, LHZX12]. certificateless [THS12, ZM12]. Certified [BDGP13]. CF [GSM19]. chain [AMAY19, CPS11, Aki18]. chaining [EA11, TSCB19]. chains [KMK17]. challenge [GGS^{+}19, Rya13]. Challenges [AZX14, CSMC19, DPL16, VHFST15, BGEP17, BCG^{+}13, Che17, Chr16, CDS19, DGCA17, GDLB16, HM16, JG14, KS19, MSS18, PMR16, Fug12]. challenging [MSHG18]. Chandra [TG10]. Change [HGBS18, SLLL14, AHBA19, AGR19, CS15, CPD^{+}18, CdCMdMSNdA16, DRELHE16, HJBH10, JLGM17, KWS^{+}17, MJZ^{+}10, MMB10, NKMM12, PB11, SVM19, SLL^{+}15, ZcKS17, dOFB^{+}19]. change-impact [CS15]. change-point [MJZ^{+}10]. change-proneness [HJBH10]. changeability [SLLL14]. changed [GV10, RBS19]. changes [BCD^{+}18, DNSH13, JLL19, RvDV17, SPC16, WMW12, WRS^{+}17, YFZ^{+}16]. Changing [CLS^{+}12, FS14a, HTB12, KGG18, XYCL17]. channel [KKP12, LZ13, LL14, MLHL12, ZGZ^{+}13]. channels [HSS10, TNK^{+}19, WGKW19]. Chaos [LW13c, jWLY^{+}13, CCLL11, LW13a, PPG^{+}13, jT12, ZLW^{+}12]. chaos-and-Hamming [CCLL11]. Chaos-based [LW13c, jWLY^{+}13, ZLW^{+}12]. chaotic [HRB12, LWC13, NES^{+}14, WGZ^{+}12, ZT14]. characteristic [CSW13, MA10]. characteristics [CRL^{+}12, CPRT16, DZT^{+}14, DGCA17, EED16, FVHF^{+}15, SJ17, SS15, SA18]. Characterization [dMCR19, BPQP^{+}10, DDD14]. Characterizing [MRT17, VGSN18, CPZF19]. characters [PWC12]. charging [LWOY16, LWL^{+}16, LWC^{+}18]. Chart [ZLG10]. cheat [WS12]. cheat-preventing [WS12]. checker [WKZL10]. Checking [ZTZ^{+}11, ABFM12, CHLW17, CSS10, EBEL18, LXC13, RSCB18, ZML10]. Checkpoint [SBZ^{+}17]. Checkpoint-based [SBZ^{+}17]. checkpointing [BM18, CLY14, KKH^{+}16, OD10]. Chen [LLLK10, YWEL^{+}13]. China [DLW^{+}13]. Chinese [JKC19]. chips [TC16b]. choice [CPX16]. Choquet [SNM14]. Chord [SBZ^{+}17]. choreographies [BMKM15]. chunking [SHGT16]. cipher [AMS^{+}10]. circuit [CCdR^{+}16, WMOKY11]. cities [AKA^{+}15, PCG^{+}14]. city [HWHT11, KLL^{+}11, HWdS^{+}15]. Class [NCS10, AR12, Al 12, CCR14, EVR11, MJ14, OWB11, QGZ^{+}15, RO13b, SS15, ZJZ^{+}17, BDO11, FTSC12]. class- [RO13b]. classes [AC17, CHN19a, ZXL10]. classification [DRCG12, KCT12, KU10, KKA^{+}19, LZ12, LLWL19, MRBN17, MRJD^{+}12, SZ11, SS14b, SLLY17, TCK14, VHL14, ZML10]. classifications [ALRP16]. classifiers [XHM^{+}11, Zha12a]. Classifying [dAGSdFS^{+}15, Ala15, CPZF19, GSM19, LHG^{+}18, YFZ^{+}16, HRRC16]. classroom [AAN11]. CLEFIA [TSLL11]. CLEFIA-128 [TSLL11]. client [BCF18, CWJK13]. clients [KNA11, OM13]. clone [MRS18, ND18, ZcKS17]. clones [BKSM13, BKSM14, MRS18]. cloning [MRRS19, ZYZ^{+}17]. Closed [WLC13a, NK15, NDS13, OH15]. closed-loop [NK15]. closeness [WKbOS17]. Cloud [AKAA18, FS14b, GDLB16, GGS15, HLS^{+}13, MT13, Rya13, AJG^{+}15, ALRP16, AO16, BMA^{+}13, BV15, BJK^{+}11, Bis13, CNM18, CZG^{+}15, CXO^{+}15, CHL^{+}13, CAG17, CDPM17, DKP^{+}19, DS16a, DFJ19, DEA^{+}14, DM17b, DS16b, EGHO16, FB18, FNWL18, GS17, GGS^{+}19, GGB19, GCSSDP^{+}18, GMMC13, GZS^{+}18, HS15, JCYT16, KSN17, KA18, KQ17, KBRV17, KBRV18, LMT16, LZCL19, LLHY19, LDZL15, LZY^{+}15, LZC14, LCL15, LY18, LZG15, MJ18, MGB16, MK17, MS17b, MBT16, MIKG13, MCV15, NK15, NB13, Oja16a, Oja16b, OSH^{+}18, PWS^{+}15, RQD^{+}17, SKK^{+}18a, SCO13, SBB^{+}16, Som13, SCC16, SS13, SWES16, TY18, TG17, VPMVM^{+}13, WDC12, WCX15, Wen16, WCB^{+}17, XZZ^{+}16, XB19b, YYS^{+}16, YL16, YCLC17, ZWC^{+}19, ZFY^{+}19, rBHM17, Cha17, LZO^{+}13, LZO^{+}16]. cloud-based [CXO^{+}15, CHL^{+}13, LDZL15, MK17, TG17, YCLC17]. cloud-native [KQ17]. Cloudera [MCL^{+}17]. clouds [DVV^{+}16, MK15b, SB19, TSCB19, ZSB19, ZHAY12, CdAM^{+}14, KKG^{+}12]. CLPL [CX10]. cluster [CDGJ10, MAS13, WGC^{+}14]. Clustering [CV14, LK13, LWOY16, XZZ^{+}16, BPGS13, CZC^{+}18, CL17b, DFJ19, HR10, KS16, LQC^{+}14, MK16, MJ14, NMM13, TZ12, YH19, ZCZZ11]. Clustering-based [XZZ^{+}16, MJ14]. clusters [AO16, BLM10, BHH^{+}10, IKBH14, RBT11, SBZ^{+}17, ZHGL11, dACM17]. CMMI [WL15a]. Co [DRELHE16, SVM19, BSG^{+}18, GGvH^{+}18, HyLW^{+}12, HNH15, KBHG17, SHHL12, WRS^{+}17, dOFB^{+}19]. Co-change [SVM19, dOFB^{+}19]. co-changes [WRS^{+}17]. Co-evolution [DRELHE16, BSG^{+}18, GGvH^{+}18, KBHG17]. co-fix [HNH15]. co-located [SHHL12]. co-scheduling [HyLW^{+}12]. coal [BRG^{+}12]. cocktail [OHJ10]. COCOMO [NBH19]. Code [KH10, SED16, WFF18, YC13, AMdLM17, CAHV15, CCLL11, CHL^{+}19, EAH^{+}11, FDN^{+}16, FLRT19, GK18, GE15b, GPD^{+}19, HNZ17, IKBH14, JKL19, KS19, KR14, LK13, LK16, LQLC16, LZL^{+}18, LGM^{+}18, LCL^{+}12, MRS18, MRRS19, MSGM17, ND18, NVPGMPSM17, OM13, OHL17, OKS^{+}15, PAR14, QBO^{+}14, SJR^{+}11, SRJL^{+}18, SPMG18, TAF^{+}17, WDC10, YXH^{+}18, YWHL11, ZTZ^{+}11]. code-based [JKL19]. code-smells [OKS^{+}15]. codebases [KTK19]. CodeCloud [CdAM^{+}14]. codependent [VvSvV16]. codes [Ala15, BMS11, WYCC13]. Coding [BV18, WAWO12, JXLC15, KM11, LWC13, PMDH13]. coding-based [JXLC15]. Coding-error [WAWO12]. cognitive [AMAY19, BPGS13, PV18, AAJD^{+}16]. cognizant [HPH12]. Coherent [IKBH14, PN14]. cohesion [Al 12, BDO11, MJF10, QGZ^{+}15]. coincidental [LLWL19]. collaborate [vAAJ16]. Collaboration [MdOBW^{+}15, ÇB16, CRSS14]. Collaborative [ESM^{+}19b, PSEE12, RF18, YSJ13, AAN11, AHOP14, BDG13, CX10, CC11, GLZ15, LLWL14, LWL^{+}16, LOFA17, NOPF12, PRS11, PQBP16, TT13, TTT14, WCB^{+}17, Xia13, XWZC14, HB13]. CollabRDL [LOFA17]. collection [AKA^{+}15, AN10, KKLB11]. collections [SH17]. Collective [MAR^{+}19, PGP^{+}19]. collinear [LXG10]. Collision [KHC16, ZL12b]. Collision-based [ZL12b]. collisions [RM19b]. collocations [WFF18]. Collusion [MMSD13]. colony [TJH15]. color [CPL13, SNM14, yWpWyYpN13, WGZ^{+}12]. CoM [LJDK10]. combination [BGG10, NSM17, PB15, YLA16b]. combinations [MRBN17]. Combinatorial [TY18, BV15, KBRV18, ZYZZ14]. Combining [CBVF19, DW14, SC19, TC16a, LY18]. come [DDMP14]. comment [IBAH12]. Commentary [WB10]. comments [CHL^{+}19, ZYY^{+}19]. commerce [DLW^{+}13]. commercial [HBR19, KKP12, LZO^{+}13, YSSaR14, vAAJ16]. commit [RCPZ19]. commitment [EBEL18, WKbOS17]. commits [SYXL17]. commodity [KMK17]. Common [AMKD13, CCWT13, HR10, LRB^{+}19, SCO13, TKZW17, dMCR19]. communicating [SK13]. Communication [LMS11, MV10, WGKW19, AN16, AHLH16, BML^{+}13, BCD^{+}18, GC13, HSS10, KA14, KKLC12, KM14, LT13, LyWSZ10, MRM16, NK14, SCMS15, TKSRP11, TNK^{+}19, TGKL19]. Communication-efficient [LMS11]. communications [BBA10, SS13]. Communities [SBGT13, ESM19a, GL14, TKH^{+}11]. community [Ano13a, ESRF19, KCV^{+}19, LWZ12, QGZ^{+}15]. Companies [ESWA18, BV16, GTF17, HBOS13, MVSG18, SNDC13, VHFF^{+}17]. company [AT18, DLW^{+}13, TTMI19, YJZ17]. comparative [LZO^{+}16, MVLJ18, MRS18, SMS11, SLL^{+}15, TAJ^{+}10, TdCAF16, vHAT13]. Comparing [BRB14, BV16, NBF^{+}19, RO13a, SGMHJ13, AAM^{+}17]. Comparison [DR12, CRC19, DC11, LMIV15, LFCL12, MA10, NBA^{+}17, OFR^{+}12, OSH^{+}18, RGV^{+}17, TLK16b, ZML10, ZZP15, ZP17]. compatibility [FCC^{+}10, KKT17]. competency [HJP15, PJK13]. competition [HSM16]. Compiler [WWL^{+}10]. Compiler-assisted [WWL^{+}10]. compilers [CWK^{+}11]. complete [WL17]. completely [SD16a]. Complex [CM12, AAA11, BM17, CX10, CHN19b, CL15, CL17b, Cic16, FS19, FGD^{+}17, KGG18, PRN17, SGK12, TS19]. complexity [DNSH13, EK12, LWW^{+}10, ZLT10, ZXL10]. compliance [MOH16]. Component [MPRS14, ADTZ12, ASGJ13, ARS17, BWP16, BM18, BKH10, BCS18, BKRW19, CHCO11, EL10, FM11, FCC^{+}10, GHBD^{+}16, HNS12, KM17, KAM13, LG15, LASL14, MBD13, MA11, PEO11, TAB^{+}16, VCdA^{+}16, YM13, ZLZ11, WL10]. component- [LASL14]. component-based [ASGJ13, ARS17, BM18, GHBD^{+}16, HNS12, MBD13, MA11, PEO11, TAB^{+}16, VCdA^{+}16, WL10]. components [BWW^{+}18, BDLM16, HJ14, ICSK14, JRO12, LCLP16, LLX^{+}11, MPAA15, OCC13, PDS19, RBT11, RITF^{+}11, SAMN12, SSSA17, SJ17, SS15, YSG17]. Composing [LLX^{+}11]. composite [Cic16, HS15, LQLW12, LASL14, MK15a, SYT^{+}17, WZJI14, YDGB^{+}12]. composite-metric [MK15a]. Composition [BWH10, BDBLP15, BBS10, BEK^{+}19, FYCL13, LKL^{+}11, LLZW14, MdOBW^{+}15, MS17b, TBG13, TKK^{+}19]. Compositional [SK18, TKJL13, MKS10, SGC^{+}17, TKJ15]. compositions [APM^{+}14, BBD18, Mer13, MSL12]. compound [KPS10, jT12]. Comprehensible [MdFD^{+}15]. Comprehensive [OD10, ABJ10, CS15, CPRT16, FBB15, FCC^{+}10, YZC15, rBHM17]. compressed [Lin12a]. compression [JEEL16, KPT13, QZ14, SI12, WCCL10]. compression-based [SI12]. computation [DEA^{+}14, TAB^{+}16, YDGB^{+}12, YZL^{+}14]. Computational [YGN^{+}16, RHHT18, SRS15, TdCAF16]. computations [Shi10, SK10]. Computer [BD10, LCJ10, HHC12, HLWS13, Ifi11, KBDGAW16, MB19, MCV15, SLW^{+}15]. computer/IT [Ifi11]. Computing [TMTB19, AJG^{+}15, ADMOK^{+}10, AR18, AHLH16, ALRP16, AAN11, AGBD14, BV15, CZG^{+}15, DKP^{+}19, EGHO16, FB18, FTC16, GGS^{+}19, GZKL13, HGP^{+}12, HH17, KHSD10, KHS11, KQ17, KBRV17, KBRV18, MT13, MCV15, Oja16a, PNJGF12, RQD^{+}17, Rya13, SY16b, Som13, SGEK19, TY18, TJT^{+}18, TLK16b, WCX15, WLZ^{+}17a, Wen16, XB19b, YL16, YSJ13, ZGSH13, rBHM17, vWSB13]. conceal [EEAZ13]. concealing [CPL13]. Concept [LBX12, MS17b, AACT13, CHN19a, DH13, LMGHB17, YF15]. concept-drift [YF15]. conception [BGS^{+}16]. concepts [CCD19, FM11, KSAR18, TKH^{+}11]. Conceptual [RKK16, SA14, AF16, ARH^{+}17, BDPRC18]. conceptualization [SOS^{+}16]. concern [ADTZ12, FSGL12, RDPM19]. concern-sensitive [FSGL12]. Concerns [SSR18, CHCO11, PSEE12, VM13]. concrete [PC10]. concurrency [HK13, SNDD19]. concurrent [ACRD19, dSACdLF17, BM18, CL18, LZR16, PTF^{+}15, WLL19a, WLL19b]. condition [CCWT13]. conditional [EBEL18, FSGW11]. conducting [CC11]. Conference [BKW10, KT16, LH12, SS17, BCL^{+}18, HL10]. conferences [LCM^{+}13]. CONFIDDENT [PGRQVV12]. confidence [LYC14]. Confidential [HS11b]. configurable [PSS^{+}16]. configuration [BLM10, ESM^{+}19b, HGBS18, JLL19, MSAH16, MAS13, OGRJ^{+}18, SP14, SHBC19, TBG13]. configurations [GBH^{+}16, WBS^{+}10, WGS^{+}14]. conflicts [EUR^{+}13, HST15, HST16]. conformance [ATHM17, LCLP16]. Confucian [WKbOS17]. congestion [XZP^{+}10]. congruence [ZCC^{+}19]. conjunctive [BL11]. connection [Cic16]. connections [Cic16, GBDCR12]. connectivity [TZB19]. connector [BKRW19, LASL14]. connector-based [LASL14]. connectors [EL10, NSDI16]. consecutive [AT18]. consequences [HTB12, SMB17]. considerations [ZW15]. Considering [BD16, SÁM^{+}16, WWSZ15]. consistency [TLGE18, ZcKS17]. Consistent [DEW^{+}16, TLWS10, DRELHE16, EA12, KH14, PGRQVV12]. console [BLL^{+}18]. consolidation [KCV11, LZY^{+}15, LN13]. constant [Shi10]. Constrained [LW13b, AR18, DVV^{+}16, HZG^{+}12, LZ13, PCCB^{+}11, Sko14, ZCC^{+}17, MGM10]. Constraint [LKR13, CCR14, EK12, TFS10]. Constraint-based [LKR13]. constRaints [LPP15, ACH19, CL17b, CF12, GLZ15, GWW^{+}11, KTT^{+}17, KBHG17, LGS^{+}19, NBF16, SÁM^{+}16, SRS15, ZTZ^{+}11]. construct [KTF^{+}16]. constructing [FYCL13, GMPN16, GCSSDP^{+}18, GPSS^{+}13, GAT15]. Construction [KK17b, BHM12, CX10, Luk11, RLL^{+}18, SKL10, SMK^{+}18, WWLG13, YKC^{+}12, SBGT13]. constructs [MNO18, PTF^{+}15]. consultants [CWJK13]. consumer [HTB12]. consumption [ÁRMC16, APS^{+}10, KA18, PAS^{+}10, SMSH18, XJZ^{+}15]. Container [dACM17]. Container-based [dACM17]. containers [SMSH18, XB19b]. contemporary [BGS^{+}16, TNK^{+}19]. contending [AAMS16]. content [CdR^{+}14, FMPS16, KM17, LHH10, LVPMPCLS13, PÁC13, Shi12, WWSZ15]. content-aware [LVPMPCLS13]. contention [BLS18, BL19, CYT16]. contention-free [BLS18, BL19]. contents [LLLK12, LAT10]. Context [BDV17, EZRK16, SGP12, AAC16, BD16, BSDD14, CBC14, DBZ16, FRGC10, GMR17, GDSB11, HGMB13, KOS15, KRJ17, KAK^{+}13, KSHC14, LC11, LXC13, LLL17a, MOD^{+}19, MRT17, MSK^{+}17, NK15, NBR^{+}14, PCCB^{+}11, RSBA19, SRWE10, SG16, VKL16, XCM^{+}12, ZYY^{+}19]. context-aware [AAC16, BSDD14, DBZ16, FRGC10, GDSB11, HGMB13, KRJ17, KSHC14, LC11, LXC13, MOD^{+}19, MRT17, PCCB^{+}11, SRWE10, XCM^{+}12]. Context-awareness [EZRK16]. context-based [VKL16]. Context-oriented [SGP12]. context-sensitive [SG16]. contexts [CCY11, KWS^{+}17, LK13, MER17, MHB18]. contextual [Aki18, WRS^{+}17]. continuity [SMB17]. Continuous [BK17, Che17, Cho13, FS17, RHL^{+}17, TGBF17, HMOK18, IBM11, LCC10, SB14, SMB17, YMM^{+}17, YMM^{+}19, FGMM17]. Continuously [BKRW19]. Contract [DGBE18, ASMN15, TKK^{+}19]. Contract-based [DGBE18]. contrast [DDD14, GLW13, MM14]. Contributions [LN13, CLL14, LMWM18]. control [ARS17, BSK^{+}18, BSKL10, BM17, CFL^{+}18, DMSG11, DYC19, EK12, FBB15, FNWL18, HVK11, KKL^{+}11, LBCL10, LH11b, LDS^{+}19, MGM10, NZM10, PCHW12, PCYZ12, PCCB^{+}11, SC19, WXY^{+}17, WLL19a, WLL19b, XZP^{+}10, ZML17, dRSBA13]. controlled [HC10, MNSA15]. controller [CV14, MMTS15]. Controlling [CWJK13, dSB12, CDGJ10]. conventions [HAE^{+}15]. convergence [KL11, TT10]. Convex [LSE12]. Cooperation [CRSS14, SSMvD16, dVRB13]. cooperation-based [SSMvD16]. Cooperative [CMR19, NMM13, ACSC16, BD10, HdM17, KSHC14, FH10]. Coopetitive [GD12]. coordinate [LOFA17]. Coordinated [BSK^{+}18]. Coordination [APCS10, SHHL12, NPC12, Sko14]. copyright [GJ13]. Coqcots [BDLM16]. core [CYT16, CKC15, FHL^{+}15, KSH^{+}12, LS14, PN14, PGPC17, WX10, ZCC^{+}17, fLSN18, CD10]. corporate [FG15]. correct [BHH^{+}12, LLWL19, LJDK10]. correcting [BMS11, CV16a]. correction [ABS19, LQLC16, OKS^{+}15, YLXZ16]. correctness [SMK^{+}18]. correctness-by-construction [SMK^{+}18]. correlated [HSC15]. correlation [LYLC16]. Correlations [SMB17, MC10]. Corrigendum [APS^{+}10, BKSM14, HST16, LHP^{+}10, TTT14, WZM12a, XTZX13, YMM^{+}19, YWEL^{+}13, wZfG14a]. corruption [WLZ^{+}17b]. cosine [Lin12b]. Cost [ALRP16, LMT16, RB16, SD16b, WAG15, ZGYS^{+}15, AN16, BCLW11, CCL^{+}19, DFJ19, HPH12, JRSN10, KGB11, LXG10, LNW^{+}11, LZG15, MK16, MBF12, MA10, MPAA15, PFO^{+}19, PvV12, Pot13, PACH15, WL15a, WQJZ10, WL17]. cost-cognizant [HPH12]. Cost-effective [RB16, SD16b, WAG15, DFJ19, LNW^{+}11, PACH15]. cost-efficient [LZG15]. cost-sensitive [WQJZ10]. Costs [GSdS16, HLWS13, Zha12a]. COTS [BWP16, RPK^{+}13, YSG17]. counterexample [YXP^{+}18]. Countering [YF15]. Countermeasure [BP13]. country [VBC^{+}14]. coupled [CDOP15, EZG15]. coupling [AC17, AAM16, CRC19, CCMOM19, DNSH13, FM11, MS16]. course [BRS^{+}18, KH10, RF18, RHM^{+}18]. courses [vWSB13]. cover [UUN13]. coverage [AMdLM17, GZY11, LMH10, LT11, LLK11, LCL^{+}12, MGM10, PAR14, SPMG18, WL17, WDC10]. coverage-based [WDC10]. covert [LT13, LyWSZ10]. CPSs [AM15]. CPU [BSKL10, CRK^{+}18, DAG19, SMZC12, SK13, YCF^{+}13]. CPU-bound [DAG19, SMZC12]. crash [GXZ^{+}19, LMS11, PNY14]. crash-recovery [LMS11]. crashing [GXZ^{+}19]. create [LK13]. created [KVH12]. Creating [SLLY17, Oja16b]. creation [CCdR^{+}16]. creativity [AVGM19]. credibility [SFMB16]. criteria [CCP18, EFSJM17, FMdAR16, MK15b, PB15, AKAA18]. Critical [DSGS17, CCN^{+}10, KHC16, SNDC13, Sta14]. Critical-blame [DSGS17]. criticality [CSMC19, LGHR16, PGPC17, XB19a]. Cross [CTHW12, GCDY16, HKS^{+}17, SRBT18, AdAD17, DAG19, GD12, GMMC13, KSHC14, MVSG18, NBR^{+}14, RM19a, RA16, SS12, XLL^{+}19, YFZ^{+}16, YJZ17]. cross-application [AdAD17]. cross-case [MVSG18, SS12]. cross-company [YJZ17]. Cross-factor [GCDY16]. cross-functional [GD12]. Cross-layer [CTHW12, KSHC14]. cross-platform [DAG19, NBR^{+}14, RM19a, RA16]. cross-project [YFZ^{+}16]. Cross-validation [HKS^{+}17]. crosscutting [Ano13a, CHCO11, VM13]. crossover [CV16b]. Crowdsourcing [PLVB^{+}18, BS15, GGC16, KA17, LHG^{+}18, MCHJ17, STA19, TT13, TTT14]. CRT [KKHH11]. CRT-RSA [KKHH11]. Cryptanalysis [WWYZ11, LLLK10, LGLL12, RITF^{+}11, SDM10, TSLL11]. Cryptanalyzing [ZLW^{+}12]. cryptographic [PSdO^{+}13]. cryptography [DDD14, LLLK10]. cryptosystem [IB11, NZM10]. cryptosystems [HRB12, SA16]. cube [ZGZ^{+}13]. CUP [VHL14]. curation [Bis13]. Current [BD16, Chr16, dONTF^{+}19]. curvature [GJ13]. curvature-feature [GJ13]. curve [IB11, NZM10]. curvelets [ZLmLN14]. curves [PSNB11]. custom [WLZ^{+}17a]. customer [AAMS14, AAMS16, GC13, HHK13, VLL18]. customer-developer [GC13]. customers [DLW^{+}13, OD17]. customization [MVLJ18, PD16, WVT^{+}14]. customized [AMGG14, GMPN16]. CUSUM [MJZ^{+}10]. CUTE [LMH10]. Cuts [CJ13]. CVSS [HFE10]. cyber [AZX14, AWSE19, CSMC19, DMA18, GVPPM18, GBH^{+}16, GSP^{+}19, LS17b, LL15, LDS^{+}19, MPLL18, VSDD12, ZAY19]. cyber-foraging [LL15, MPLL18]. cyber-physical [AWSE19, GBH^{+}16, GSP^{+}19, LS17b, LDS^{+}19, ZAY19]. cybernetics [Cha17, CHLW17, DWC17, LGH^{+}17, LZLC17, LLL17a, YCA17]. Cybersecurity [UB19]. cycle [Fei12, LMT16, SS15, WB12]. cyclic [LWLL12, OCC13].

D [BMAH11, GCLD13, JSL16, MKH^{+}12, SSP^{+}15]. D-P2P-Sim [SSP^{+}15]. DAG [LCLS16, SRS15]. daily [SSD16]. Data [CSS10, HNS12, HCL12, KC16, LZCL19, MJ18, MRBN17, SKT17, UB19, AQK11, AG15, ACL13, ACSC16, ÁGBYB^{+}14, AN10, Bis13, BTPLST15, CCGG14, CNL13, CPS11, CDOP15, CLL10, CBVF19, DII^{+}17, DM17a, DIB14, DS12, DK15a, DHC^{+}11, Dut15, EH19, FS14a, FF12, GZY11, GTY12, dGFDL16, GMGTdFR14, GP10b, GPL^{+}15, GZS^{+}18, HBG^{+}13, HSC15, HY11, HBT16, HST15, HST16, HC10, HCL^{+}10, HWL13b, HSS10, HTH13, IAA16, KRDH12, KCR16, KSAR18, KKLC12, LM13, LCT10, LC10, LZL^{+}15, LCLF13, LJL^{+}12, LBCL10, Lin12b, LCC^{+}13, LDZL15, LTK^{+}15, LWZ^{+}16, Lin16, LZ12, LW13b, LLML13, LZLC17, LKK14, LCL15, LLC17, MMP15, MQG^{+}17, MTF14, MDBC17, MG11, MC10, MIUM12, MT10, MdFD^{+}15, MSL12, MJZ^{+}10, NK15, NBH19, NDS13, OLZN13, PS13, PAOC15, PWC12, QZ12, RSB^{+}14, RLY^{+}13, RHRC13, RHRC15, SM17a, SD16a, SAA^{+}10]. data [SG16, SHN14, SAH12, Shi17, SJC13, SGBCP12, SBDB19, SGW^{+}15, TLWS10, TKJ16, TGKL19, TBC^{+}16, TLK16b, TPTV17, VTZ^{+}17, VZT17, WCCL10, WLH13, WYCC13, WLC13b, WCC^{+}14, WLZ^{+}17b, WWY^{+}12, XLM^{+}15, YWWS10, YWTW11, YWHL11, YCLY13, YTW^{+}13, YF15, YYS^{+}16, YZL^{+}14, YM13, ZJZ11, ZHH^{+}17, ZCZZ11, ZHAY12, FGD^{+}17, HBG^{+}14]. data-centric [WWY^{+}12]. data-hiding [WYCC13]. data-intensive [Shi17]. Data-locality-aware [KC16]. database [BL11, CH10d, EAH^{+}11, HyLW^{+}12, HNS12, KVT^{+}17, LKL^{+}11, LZG15, LGZ^{+}18, PDK^{+}16, RB16, SLLL12, ZTZ^{+}11]. database-driven [PDK^{+}16]. databases [DK15b, HHK13, JK13, KKR16, Lin12a, MLGA11, RVCM17, VGM13]. Dataclay [MQG^{+}17]. dataflow [LZJ^{+}19, SMM17, WLZ^{+}17a]. dataflow-driven [LZJ^{+}19]. datasets [HKS^{+}17, LXG10]. DCT [LCC^{+}13]. DCT-based [LCC^{+}13]. DDDAS [NK15]. DDS [PG15]. deadline [DVV^{+}16, LSE12]. deadlines [CBL^{+}15, HST15, HST16, SK10]. Deadlock [PRN17, AHW10]. Dealing [FRGC10, Sko14]. debt [BMB18, BMB19, CREH^{+}18, FKA16, FSGYP17, GSdS16, LAL15, MS16, MKS^{+}18, MGM16, SSK19, TAV13, YHMS16]. Debugging [KVH12, AZvG11, ASdMGM14, BND14, BLC^{+}18, DW14, MJ19, PGP^{+}19, PW18, WWSZ15, XST18, YLCZ12, CA14]. decade [DNBM12]. decaying [JLZ^{+}19]. decentralised [NPC12]. Decentralized [EMSU11, HSC15, JS13]. deception [DMA18]. decide [JK12]. Decision [MTA^{+}16, URG10, Zha12a, vVT16, BWP16, BWW^{+}18, DGCA17, DCP12, ETYL15, GLZ15, GPMI13, JLZ^{+}19, OWG19, PWS^{+}15, RPT19, SWA^{+}13, WQJZ10, WPL^{+}18, vHAT13, vHJPB^{+}17, AKAA18]. decision-making [BWP16, BWW^{+}18, ETYL15, SWA^{+}13, AKAA18]. decisional [LJC16]. decisions [CPS11, HGBS18, MAH18, MFM10, SHS16, VM12, vHAH12, vHAT13, vdBSvS^{+}19]. declarative [CCGdL10, CGPT14]. decoding [BMS11, LHY12]. decomposition [AIE19, MDBC17, SJ17, TC11]. decoupling [PC10]. decrease [MJ18]. decryption [FNWL18]. dedicated [ZLD13]. deduplication [XZZ^{+}16, ZHH^{+}17]. deep [DBL^{+}18, RCPZ19, XLX^{+}19, YXH^{+}18, ZLL^{+}12]. DeepLink [RCPZ19]. deExploit [WLZ^{+}17b]. defeasible [KB16]. Defect [XLW18, AC16, CSN^{+}17, HNH15, LAT10, MS16, NCW^{+}19, OCC13, ÖT18, PPB19, RSB^{+}16, SLLY17, XLX^{+}19, XLL^{+}19, YJZ17]. defect-related [MS16]. defectiveness [OY16]. defects [CPV^{+}14, WAWO12, XCM^{+}12]. defense [DMA18]. defined [FFdRG^{+}14, WTG^{+}15]. Defining [AAA11, CDGJ10, KBJZ15]. Definition [BGEP17, KSKP11, LFW15, MGR^{+}13, RRM17, YKC^{+}12, dSF12]. Definitions [ABL15]. definitive [RM19a]. degree [PD16]. delay [CSW10, LZ13, LGS^{+}19, NSAK10, TAB^{+}16]. delay-constrained [LZ13]. delay-tolerant [NSAK10]. DelDroid [HBM19].

Delegatable [WZ11]. delivering [SCO13]. Delivery [Che17, AN16, LHH10, VvSvV16]. Delphi [EGHO16]. Delta [LLL^{+}14, YLCZ12]. Delta-oriented [LLL^{+}14]. Demand [LS14, DR12, HST15, ZLC^{+}14, HST16]. Demand-based [LS14]. demand-driven [ZLC^{+}14]. demodulation [KKP12]. demographics [GCDY16]. deniable [HS11b]. Denial [OLV15]. density [ZCZZ11]. Dependability [FMdAR16, RASL12, HP16, MBPM19, SXYW14]. dependable [SJH^{+}10]. dependence [BHH^{+}10, CS16, YLYL17]. dependences [MH11, WLL19a, WLL19b]. dependencies [BRS10, MSL12, OCC13, SPLW17, dOFB^{+}19]. Dependency [ADTZ12, JLQ^{+}10, HJBH10, YZL^{+}14]. Dependency-aware [JLQ^{+}10]. dependency-based [YZL^{+}14]. dependent [FBD^{+}18, IBM11]. deployed [BZ14, MHLMG14]. deploying [DBL^{+}18]. deployment [AHH^{+}10, ABL15, ÇT13, CXO^{+}15, GDSB11, HS15, LLK11, MBAG11, PCCB^{+}11, RHL^{+}17, SMS11, VSS^{+}11, WL17]. deprecation [BHVR18]. depth [CJ13, KM17]. dereferences [CBSM16]. derivation [CL17b, CNKL12, LPM15, ROR11, RBW18]. Deriving [FSG^{+}11, PFF12]. describing [KT12]. description [BBA10, GGC16, GS17, LPXL10, XLM^{+}15]. descriptions [CHN19a, CHN19b, EVR11, LLL^{+}17b, NBA^{+}17, OFR^{+}12, RDVC19]. Design [AHH^{+}10, BDG13, hChSyCwL10, DS16a, HJ12, HCC10a, KLL17, MLGA11, MBD13, PZB10, RLY^{+}13, RAJ15, SM17b, AAN11, AKKS11, ACS13, ARH^{+}17, AAC^{+}17, BPO^{+}16, BD16, BZ10, BHH^{+}12, BPSK18, CSF^{+}14, CFL^{+}18, CNSG12, CKL12, DAR14, EZOK14, FM11, FVHF^{+}15, FSGL12, FMR11, FCRF16, GKD13, GVPPM18, GTA14, GA13, DDF^{+}13, HJBH10, HJP15, KRJ17, LKRYTS18, LJDK10, MCV16, MM19, MRY17, NBR^{+}13, NOPF12, PLGT10, PFF12, PGRQVV12, SCS15, SKRB19, SHS16, SMK^{+}18, SLLL12, TBSvdW18, TJH15, YZC15, YZY^{+}18, ZÁ15, ZFS15, ZADA15, ZLT10, vHJPB^{+}17]. design-based [AAN11]. design-time [AAC^{+}17]. designated [FWCS12, HYWS11, RPSL10]. designated-verifier [FWCS12]. designed [CFAP17]. designers [WK15, vHAT13]. Designing [AdB17, DFCPSF15, FHL^{+}18, TLK^{+}16a, AF16, BPB19, DBL^{+}18, GMLSF^{+}15, SJH^{+}10]. designs [ATHM17, RF14]. desk [ABL16]. destinations [WMOKY11]. detailed [PFF12]. detect [FCMJ12, KSS15, LTK^{+}15, TVMS18, YXH^{+}18]. detected [ZXC^{+}17]. Detecting [EUR^{+}13, AMdLM17, CHL^{+}19, YZY^{+}18]. detection [Aba13, ASMM18, BKLE18, BLL^{+}18, BRG^{+}12, CKCK15, CXO^{+}15, DPP^{+}18, FMR11, HK13, HAE^{+}15, HB13, KVGS11, KCV^{+}19, KHC16, LWB^{+}13, LG17, LYLC16, LTW16, MJZ^{+}10, ND18, PRN17, SG16, SKK^{+}18a, SKE10, SS14b, TLZ^{+}16, WWZ^{+}14, jWLY^{+}13, XTZX12, XTZX13, YWWS10, YLXZ16, ZFS15, ZLC^{+}14, AT18]. Detector [PÁC13]. determination [HBM19]. Determining [NBH19]. deterministic [DC11]. develop [AdB13]. developed [AT18, GN15, LMNA17, OD17, WK15]. developer [BCD^{+}18, BMB19, ÇB16, CPD^{+}18, GC13, HSM16, MSK^{+}17, SYXL17, YLCZ12, vAAJ16, LZHS11]. developer-related [CPD^{+}18]. developers [ABJ^{+}17, BDV17, CBVF19, GFWA18, HHKWB16, HAE^{+}15, LK16, LVVTP17, OBS^{+}18, WL15a]. Developing [Aki18, FGB^{+}19, CH11, KSAR18, CDS19, EA12, GMMC13, LMN10, PGPC17, SJR^{+}11, SÁM^{+}16, WRR14]. Development [AYZI10, AMGG14, LS17b, AJLS10, AKH12, ASG17, AVGM19, APS16, AB10, APCS10, AHC^{+}11, BCFP19, BCF18, BMB19, BBS10, Bos12, BS15, CM15, CNG16, CC11, CLL14, CHCO11, CBVF19, CNMR18, DC17, DNBM12, DCT17, DGCA17, DCP12, EL10, Fei12, FAI13, FFdRG^{+}14, FMRM15, FCRF16, GKD13, GRBNA10, GGC16, GD12, GC13, GTA14, GTF17, HGP^{+}12, HP16, HTB12, HVK11, IAA16, JED18, JK12, JST10, JR15, KFN19, Kel15, KRJ17, KSM^{+}16, KM14, KTK19, LCLP16, LGC17, LWSH19, LCCJ10, LSD^{+}16, LWZ12, LHCT19, LRD^{+}19, LASL14, LJ16, MWM12, MKS10, MDP^{+}11, MGB16, MT13, NCK^{+}15, NKZ17, OKS^{+}15, PJK13, PC15, PRS11, PFG13, PGRQVV12, PFL16, PM10, RF18, RM19a, RSGH12, SSMvD16, SKRB19, SWA^{+}13, SB14, SM16, SHHL12, SLLY17, SP14, VAM^{+}10, VM12, WK15, WCC12]. development [WWSS13, WBBK18, YLA16b, YHMS16, ZÁ15, ZGYS^{+}15, ZP17, BMKM15]. device [ASV^{+}16, OMLB16, SCL13, SHBC19]. device-related [SCL13]. devices [BJK^{+}11, CDA11, CCdR^{+}16, CTL12, CMK^{+}11, DS16a, GGB19, IB11, LZHS11, PCCB^{+}11, SKE10, SHBC19]. DevOps [LPB19]. diagnosability [BGLG13, KKH^{+}16]. diagnose [WLZ^{+}17b]. diagnosis [CBS16, LDZL15, MHLMG14, SKK^{+}18a, WBS^{+}10]. diagram [CTKT13]. Diagrams [LMIV15, BS12, CCR14, EA19, OWB11, SDB18, TLGE18, GC13]. diamond [CSW13, HCL12]. DIANNE [DBL^{+}18]. dictionary [MBB11]. did [DDMP14, SAR15]. difference [AQK11, JK13, LCT10]. differences [SB14]. differencing [HCL12, YWTW11]. different [CBAV16, GCC^{+}15, Kan15, LFCL12, MRS18]. Differential [Kim12, LGL^{+}10, EMBS17, LLLK10, LGLL12, SDM10, TSLL11]. diffuseness [GPD^{+}19]. Diffusion [jT12]. Digital [BEZ14, AM10b, CIB^{+}19, GMS11, KM11, KLP10, MM14, yWpNyL11]. dimensional [DGWC16, LWC^{+}18, LQC^{+}14, TC16b, ZCZZ11]. dimensions [FS14a]. directed [KPS10, SPC16, ZLL^{+}12]. Directing [KK11]. direction [WCB^{+}17, YCLY13]. directions [BGEP17, VHFST15]. disabled [HWdS^{+}15]. DISARM [KB16]. Disaster [MAEL19, HCL^{+}10]. discipline [FP19, PvV12]. Disciplined [HL10]. discourse [AT15]. discoverability [KTK19, SM17b]. Discovering [KVT^{+}17, Oja16b, SJC13]. discovery [CHL11, GLJ13, HHK13, KK11, KKR16, LGH^{+}17, LLWL14, SdSLS^{+}19, VPL^{+}10, WAWO12, ZWM^{+}18]. discrete [AMAY19, Lin12b, WS13]. discriminative [YFZ^{+}16]. disk [RFM10, SRT^{+}12]. disk-scheduling [RFM10]. disparities [WL16]. dispatching [OB13]. dissemination [ACSC16, HSS10, LKK14]. distinguisher [AMS^{+}10]. distortion [LBCL10]. distortions [MBF12]. Distributed [BND14, CSS10, KK17a, MLLK11, PNJGF12, PGPC17, XWZC14, AR12, ACRD19, AD14, ABL15, AM10b, ACW10, CDOP15, DBL^{+}18, DKP^{+}19, DK15a, DK15b, EH19, FVHF^{+}15, GSM15, HZG^{+}12, HN17, JRO12, KHSD10, KKH^{+}16, KSENM17, KA14, KM14, KB16, LPP^{+}10, LSE12, LC11, LZR16, MQG^{+}17, MARD16, NNVD17, NPC12, NBR^{+}14, PK10a, PDL^{+}16, SSP^{+}15, SB19, SK10, TR18, TAB^{+}16, THWC10, TLK16b, YCWW15, YCLC17, YZL^{+}14, ZK13, ZLC^{+}14, PD12]. Distributing [BGTC18]. Distribution [HBG^{+}14, HBG^{+}13, HSPD14, RSB^{+}16, WWSZ15]. dithered [UUN13]. diverged [MT13]. diverse [SNDD19]. diversity [CKMT10, SSMvD16]. division [MSAH16]. DiVM [RSCB18]. DNA [WGZ^{+}12]. Do [OBS^{+}18, RDPM19, BLLGSMB11, CPRT16, HA10, YHMS16]. Docker [SMSH18]. Document [CDS10, AF16, DII^{+}17]. document-driven [AF16]. documentation [HZ15, MTA^{+}16, VVA^{+}15, YLA^{+}17, ZGYS^{+}15, vHAH12, vHAT13]. documenting [AAA11]. documents [CH11, HR10]. Does [GXZ^{+}19, vHAT13, KS19, SMSH18]. Domain [PC10, ACG^{+}15, AMCC14, ARS17, AMK12, BML^{+}13, EMBS17, EZRK16, FH10, GJ13, HGMB13, IZ18, KMK16, LXCM11, LLL^{+}17b, MSS18, MPTT14, PWW10, SKL10, ST13, SHS16, SP14, yWpWyYpN13, YWWS10, KVH12, RASL12, VPdP13]. Domain-specific [PC10, ACG^{+}15, AMCC14, ARS17, EMBS17, HGMB13, KMK16, SKL10, SHS16, VPdP13]. domains [JHYK10, NES^{+}14, PAB^{+}17]. DoS-resistant [HCC10b]. DOTS [CL17a]. Double [NTRN11, BV15, KBRV17, KBRV18, TY18]. Double-layered [NTRN11]. doubly [AC16]. downlink [WC11]. Dr. [TG10]. drag [SDB16]. drag-and-drop [SDB16]. DRank [SPLW17]. DRE [TDW^{+}14]. drift [BGEP17, YF15]. Driven [PMR16, AdB13, Aki18, AF16, AC16, BCF18, BKRW19, FGB^{+}19, CWK^{+}13, CPYZ14, CHCO11, CV16b, ELHC13, FDÁM12, FAI13, GMPN16, GKS18, DDF^{+}13, GEM15, HP16, HVK11, HK13, JPGdL17, KBM18, LZJ^{+}19, LWZ^{+}16, MEB^{+}10, MGB16, MBAG11, MAG12, MB19, MCS^{+}12, MGR^{+}13, MD16, NK15, NJ17, PDK^{+}16, PG15, PBD^{+}12, PGRQVV12, PZ15, RRM17, SAMN12, SBDB19, TKJ16, TAF^{+}17, TTR^{+}13, TGP11, UIK17, VM12, Wau19, WWSS13, WLD16, ZLC^{+}14, BMKM15]. drivers [OMLB16]. drives [nWsCqW12]. DRM [LLLK12, LHH10]. DRM-protected [LLLK12]. DRMFS [LLLK12]. drop [SDB16]. DS [NJ17]. DSFMS [GPMI13]. DSL [MAGC^{+}17]. DSLs [BLC^{+}18]. DSP [WWL^{+}10]. DSS [GRR16]. DSVerifier [CIB^{+}19]. DTN [ST11, VT14]. due [BMB19]. dumb [MKRO14]. duplication [LYLC16]. duration [GGC16, LMA15, PCCK18]. durations [LNY^{+}11]. during [BRS^{+}18, BKRW19, FB18, MAAC17, RDVC19]. DyDAP [SGBCP12]. Dynamic [APM^{+}14, APT^{+}12, DVV^{+}16, EGG^{+}11, OCC12, PCCB^{+}11, VSS^{+}11, WCX15, WCTK12, ASV^{+}16, ADET12, AKA^{+}15, AR17, BRB14, BPQP^{+}10, BSKL10, BLM10, CS15, CFL19, CCdR^{+}16, CdSdSG^{+}18, CKCK15, CTL12, CYT16, CS12, DY15, DS12, DZT^{+}14, EA11, DDF^{+}13, JS16, LZL^{+}15, Li11, LJL^{+}12, LG15, LH11b, MCC^{+}18, MAAC17, NK15, OM13, PB15, PLHP^{+}15, PCYZ12, PAR14, PWA^{+}19, RO13b, RLL^{+}18, RMCH^{+}14, SGBCP12, SKF17, THWC10, VKL16, WL16, XJZ^{+}15, YLC18, CBT^{+}14]. dynamic-circuit [CCdR^{+}16]. dynamical [jT12, Wen16]. Dynamically [GAT15, CFN10, QOLJG16]. dynamics [ESRF19]. DYSCS [CdR^{+}14].

E-business [RCL14]. e-commerce [DLW^{+}13]. e-mail [MRJD^{+}12]. EA [vdRBSvV10]. Early [CM15, CREH^{+}18, ÖT18, SMK^{+}18, APS16, FSG^{+}11, NHC13, SG16, SS15, ZK13]. EAST [WRTP^{+}13]. EAST-ADL [WRTP^{+}13]. ECC [LH11b]. EClass [KCT12]. economical [GPL^{+}15]. economy [JK12]. ECOS [DFJ19]. ecosystem [CRL^{+}12, Han12, KHMA12, LZO^{+}16, PSZ17]. ecosystems [AS16, BBS10, Bos12, EB14b, HSS14, HSM16, MH13, Man16, SA12, VA17, VAJ18, vAAJ16, KJS^{+}12]. ECP [Cic16]. ECUs [fLSN18]. Edge [TMTB19, FF12, LTW16, SB19, TJT^{+}18, WCB^{+}17, YCLY13]. EDGES [MC10]. editor [CCM12]. Editorial [Ano11a, Ano11b, Ano11c, Ano11e, Ano11f, Ano11g, Ano12e, GP10a, JWT17, TBG17, Won10, WC16, SY16a, Ano11d, Ano11h, Ano11i, Ano11j, Ano11k, Ano11l, Ano12a, Ano12b, Ano12c, Ano12d, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano13j, Ano13k, Ano13l, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano17l, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g]. Editorial [Ano18h, Ano18i, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, BKW10]. editors [BDV17, CdS18, OPS11]. education [AM18, AdB17, CC11, CMR19, FHL^{+}18, GGT^{+}19, MSSMDC12, RZL^{+}18, SW19]. educational [vWSB13]. Effect [ÇB16, AL10, BDPRC18, CPYZ14, ETM10, HJN11, HNH15, JSL16, SRJL^{+}18, YAY13]. Effective [AKB11, CKCK15, HK13, JJC^{+}14, LCC10, ROFGFRM13, CX10, DFJ19, GPL^{+}15, KHS11, KLB15, LNW^{+}11, MQG^{+}17, PCCK18, PACH15, RB16, SD16b, WAG15, WDC10, LXC13]. effectively [KTF^{+}16, ZXC^{+}17]. effectiveness [JST10, RZL^{+}18, vdRBSvV10]. effects [JH10, Jør16, Kan15, KCV11, LJ16, MFM10, SSMvD16, SHBC19, SAN^{+}17, Xia13]. efficiency [CW12, DMSG11, GKS18, KCT12, LSNL^{+}19, PAR14, SB12, TDW^{+}14, WH15, WOC15, YM13]. Efficient [AMP12, ACSC16, BKLE18, BM18, CCG^{+}18, DRCA^{+}19, GLWY10, GGK19, HL11, JLY14, Kim17, KKR16, LRO19, LHJ10, LZL^{+}15, LWZ^{+}16, LHYZ12, MCKA18, NES^{+}14, NZM10, RO13b, RVCM17, SGO13, WVT^{+}14, WXZ^{+}17, YCLY13, YXP^{+}18, YZL^{+}14, ZGZ^{+}13, ZHAY12, ATvHJ18, ASV^{+}16, BCA^{+}19, Bar15, CDA11, CKCK15, CH11, CLY17, CHL11, CZG^{+}15, DFJ19, EMBS17, EZOK14, FNWL18, GQ12, GCSSDP^{+}18, HWL13a, HSS10, HS15, IB11, JXLC15, KKH^{+}16, KKHH11, KKLB11, LMS11, LH11a, LKL^{+}11, LHZX12, LZ13, LZG15, MPN^{+}17, MSAH16, MT10, NNVD17, OT17, Pen11, PPMM17, PFL16, SM17a, Shi17, TLL12, VT14, WMWZ12, WC11, XB19b, YH10, ZM12, ZGSH13, fLSN18, MC10]. Efficiently [YZY^{+}18, LBCL10, LGZ^{+}18]. effort [ASMN15, ABL16, ANC11, ANM15, CM15, DCT17, dGFDL16, IAA16, IHA16, JH10, Jør10, Jør16, KM13, LJ16, MdFD^{+}15, NBH19, PCCK18, SPCT18]. eight [GTF17, VCdA^{+}16]. elastic [HWR17, TSCB19, ZGSH13, dACM17]. elasticity [DM17a]. elasticizing [GE15a]. ElasticSFC [TSCB19]. elderly [HWdS^{+}15, TCCH12]. eLearning [JRO12]. election [LMS11]. electricity [LZL^{+}15]. electrocardiogram [SLW^{+}15]. electronic [WKV11]. electronics [HTB12]. elements [AMdLM17, FSGYP17, TKZW17]. Elicitation [LZLC17, AVGM19, GSM15, PG12]. eligibility [DMSG11]. elimination [LZ12, WAWO12]. elitism [PS13]. elliptic [BAAS13, IB11, NZM10, PSNB11]. embed [KPS10]. Embedded [ABCH13, LPXL10, WCTK12, ÁRMC16, CWK^{+}11, hChSyCwL10, EB14b, DDF^{+}13, HZG^{+}12, HNS12, KSM^{+}16, KSH^{+}12, LC11, LLS11, MFMCY12, MBAG11, NEM17, RAK15, SCwY12, SJH^{+}10, TC12, WWL^{+}10, WWSS13, YSSaR14, dRSBA13, fLSN18]. Embedding [LCT10, SÁMI17, AO16, EA11, HCL12, MKH^{+}12, PWLL13, YWWS10]. emergence [LN13]. emergency [HWdS^{+}15, MPLL^{+}15]. emergency-care [HWdS^{+}15]. emerging [BCG^{+}13, Han12, VA17, CA14]. Emotion [MPLL^{+}15]. Emotion-led [MPLL^{+}15]. emotions [CFA^{+}19]. Empirical [DDMP14, FAI13, FP19, MPLL18, PFL16, RPT19, RSGH12, SAN^{+}17, Wie14, WSJ14, ACG^{+}15, ANG^{+}19, AKKS11, ARH^{+}17, AB10, ANM15, BRB14, BCD^{+}18, BHVR18, CFL^{+}18, CH10c, CO12, DvdVA^{+}13, DOL^{+}16, EA14, EED16, EBC10, FB18, FLRT19, GTA14, HHKWB16, HP16, HJN11, HKS^{+}17, JSL16, JLL19, KFLS18, KY10, LMH10, LWSH19, LMS12, LCL15, MNS13, MDBC17, MGR^{+}13, MRRS19, MHLMG14, NCS10, dONTF^{+}19, NCW^{+}19, OWG19, PHR10, RDPM19, RNR17, STS^{+}19, SKRB19, SVM19, SLLL14, SKF17, TB13, VBC^{+}14, YC13, YHMS16, ZXC^{+}17, ZFY^{+}19, BWH10, MPTT14]. empirically [EA19, GN15]. empirically-developed [GN15]. Employing [PWA^{+}19, VTZ^{+}17]. Empowering [OD17]. Enable [MSB18, CdAM^{+}14, PACH15, VvSvV16]. enabled [AN10, EZRK16, GGB19, KR14]. enabler [LWZ12]. Enablers [ESWA18]. Enabling [BHH^{+}12, BLUH15, PC15, YYS^{+}16, TC12]. enactment [RRM17]. encoding [CNL13, CSW13, HCL12, MIUM12, WCCL10]. encountered [GSdS16]. encrypted [BTPLST15, BL11, CH11, GZS^{+}18]. encryption [BAAS13, FSGW11, LWC13, LW13a, LW13c, NES^{+}14, RG10, RPSL10, SNM14, SLZ12, jT12, WWYZ11, WHY^{+}12, WGZ^{+}12, YLZ^{+}16, ZLW^{+}12, ZT14, ZML17, ZZ12, ZL12b]. End [BCFP19, SP14, AKL14, CFL^{+}18, CTHW12, FGBC10, GCSSDP^{+}18, HBG^{+}13, HBG^{+}14, KY10, LKP13, LASL14, LSLG17, MNO18, SK10]. end-to-end [CTHW12, FGBC10, GCSSDP^{+}18, HBG^{+}13, HBG^{+}14, KY10, SK10]. End-user [BCFP19, SP14, AKL14, LASL14, LSLG17, MNO18]. Endpoint [AT18]. ends [LKJR10a, LKJR10b, PSS11]. endurance [nWsCqW12]. enemies [WLL17]. Energy [BCA^{+}19, CLY17, FHY17, LZL^{+}15, WH15, Wen16, ZWC^{+}19, ASV^{+}16, ÁRMC16, Bar15, CDA11, CZG^{+}15, DM17a, GQ12, HZG^{+}12, JXLC15, KCT12, KA18, LWL^{+}13, LGHR16, LSNL^{+}19, LZC14, MDO^{+}10, MT10, NNVD17, PPMM12, PFL16, SPC16, SMSH18, Sko14, TdCAF16, TC12, VT14, WMWZ12, WC11, XJZ^{+}15, XB19b, YZG^{+}13, ZGSH13]. Energy-aware [Wen16, ZWC^{+}19, GQ12, LWL^{+}13, MDO^{+}10, TdCAF16]. energy-efficiency [KCT12]. Energy-Efficient [LZL^{+}15, BCA^{+}19, CLY17, Bar15, CDA11, CZG^{+}15, JXLC15, MT10, PFL16, VT14, WMWZ12, WC11, XB19b, ZGSH13]. energy-saving [LZC14, YZG^{+}13]. Enforcement [SSR18, GLZ15, HBM19, ZTZ^{+}11]. engine [MSGM17]. Engineering [AAC16, CG15, CCCY17, CVGP13, Gar13, HC15, JWT17, LN13, NFSM11, PMR16, PSS11, TGBF17, Woh16, AC19, ADCO18, AS10, AM18, BMA^{+}13, BCFP19, BCR^{+}19, BBND^{+}18, BCL^{+}18, BDBLP15, BRS^{+}18, BDM^{+}19, BCG^{+}13, CdS18, CLR18, CC11, CMR19, CRESF^{+}13, CNMR18, DGRN10, ESM^{+}19b, ETM10, FDÁM12, FVHF^{+}15, FOR19, FKWVH19, FP19, FS17, FCC^{+}10, GPP^{+}17, GCBCD15, GCDY16, GGT^{+}19, GJ16, HBP^{+}17, HP16, HLS^{+}13, HS11a, HJP15, IF19, IZ18, JED18, JPGdL17, JDLS16, KLA^{+}19, KSIZ19, LLM^{+}17, LCM^{+}13, LFW15, LHG^{+}18, LSLG17, LHLG^{+}15, MCHJ17, MACB19, MAGC^{+}17, Mer13, MPLL^{+}15, MSSMDC12, dONTF^{+}19, NBM19, OK18, RF18, RRV19, RAK15, RSBA19, RHM^{+}18, SW19, SGP12, SNL16, San16, STA19, SSSA17, dMSSS^{+}13, SBDB19, Som13, TTL^{+}13, TKP^{+}18, UGFK15, UIK17, VCdA^{+}16]. engineering [VLL18, VBC^{+}14, VCMG17, WMAS12, WRdMSN^{+}13, WSM15, WBBK18, WTG^{+}11, WLD16, ZTCZ16, dSdMSNO^{+}14, Bor12, DDMP14, HLS^{+}13, VPMVM^{+}13]. engineers [HG18, dSF12]. engines [APT^{+}12]. English [Kan15]. enhance [OCC12, SC19]. Enhanced [FHL^{+}15, PPN^{+}15, YCC16, CdR^{+}14, LWC13, TKH^{+}11]. enhancement [ULS19]. Enhancements [LYLC16]. Enhancing [CPD^{+}18, FVHF^{+}15, KTK19, MKS10, SYXL17, ZCZZ11]. enrich [TCCH12]. Ensemble [LLC17, ANM15, IHA16]. ensembles [SH17]. ensure [CH10b]. ensuring [ATHM17]. Enterprise [ŠK11, BK17, JBSL12, JKC19, LJH10, LLX^{+}11, NHH^{+}12, NB13, RNR17, SS14a, SCC16, WAWO12, dSdMSNO^{+}14, vdBSvS^{+}19, FCMJ12]. enterprises [VA17]. Entity [DK15a, CTKT13, LWXZ10, MPN^{+}17, WWLG13, ZLZ11]. Entropy [WFY^{+}19]. environment [CZG^{+}15, FHL^{+}18, KSH^{+}12, LZR16, MJ18, SA11, SSAS11, TT13, TTT14, XZZ^{+}16, YH13, YLC18]. Environmental [ZP17, ZZP15, DFCPSF15]. environments [AR12, AHH^{+}10, AD14, AdAD17, AM10b, BSG12, CNM18, CFL19, CLL10, GGK19, HMOK18, HGP^{+}12, JS16, KSN17, KSENM17, KK17b, LVPMPCLS13, NK14, PLGT10, PM10, SLW^{+}15, WDC12, XYZ^{+}19, NFSM11]. EPR [UUN11]. equations [EMBS17]. equipment [AAMS16]. equipments [AAMS14]. equivalence [CHN19a, DPP^{+}18]. era [FGD^{+}17, Oja16a]. ERD [CTKT13]. Erlang [CF13]. erosion [dSB12]. ERP [CWJK13, Ifi11, MRM16, PD16, RPK^{+}13, SL10]. Erratum [AAH12b, LKJR10a, WLL19a, Woh16]. error [BMS11, CXO^{+}15, KBM18, LQLC16, LWBH16, MSGGL12, MA10, OBS^{+}18, WAWO12]. error-correcting [BMS11]. errors [CSS^{+}13, FCMJ12, JSHW14, LCLF13, ZW15]. ESPRET [TAS^{+}18]. estimate [SA18]. estimated [OGK13]. estimates [HFE10, Jør16, LJ16]. Estimating [LYC14, TTC18, CBAV16, KLB15, LXG10, MH12, WL15a]. estimation [ATvHJ18, ANC11, ANM15, CM15, CCL^{+}19, DW11, DCT17, dGFDL16, IAA16, IHA16, JH10, Jør10, Jør16, MBF12, MA10, NHC13, NBH19, PEO11, PD16, PCCK18, PD12, RPK^{+}13, TAS^{+}18]. estimations [CBVF19, MPAA15]. estimator [SD16a]. ethics [WKbOS17]. evading [YWWS10]. evaluate [ABJ10, CXO^{+}15, HLLS13, MNSA15, MNSA16, SSF15, dOCS13]. Evaluating [CBAV16, CPDM16, CFM^{+}16, EA19, Li11, OGK13, SdSLS^{+}19, dOSdAdSG17, FSGL12, LZO^{+}13, RZL^{+}18, SMSH18, YLCZ12]. Evaluation [AAH10, LCM^{+}13, ADMOK^{+}10, AKAA18, AK16, AAH12b, ANM15, BHM12, BMAH11, BAM17, BGG10, BT17, BK17, BS15, CdCAdO18, CMK^{+}11, CRC19, CREH^{+}18, CFA^{+}19, EB14a, EA14, EK13, FH10, GLWY10, GDLB16, HRD10, JS11, KGG18, LY18, LLGZ13, MK17, MACB19, MGAN18, MSHG18, OD10, PK10a, PCHW12, PZB10, PCFRP19, PFL16, RLY^{+}13, RM19a, RGH17, SA11, SXYW14, SM16, TB13, TPKT12, Wau19, WR10, WMD^{+}10, WSJ14, ZK13, ZJC^{+}10]. Evaluations [YLC18, KOS15]. even [HG18]. event [BRB14, CM12, FS19, FGD^{+}17, HSPD14, KK17a, LGH^{+}17, PG15, TKJ16, WLL15]. event-based [HSPD14]. event-driven [PG15, TKJ16]. event-extraction [BRB14]. EventHealer [TKJ16]. events [DM17b, KFN19, TS19, TLA18]. Everything [SST16]. Evidence [SdSGdMSN^{+}13, DLW^{+}13]. Evolution [NKMM12, dONTF^{+}19, PSZ17, VHFST15, ADTZ12, BCL12, BSG^{+}18, BKRW19, CCM12, CHLW17, DRELHE16, DGRN10, GGvH^{+}18, GPPT16, HNZ17, IF10, JLGM17, KBHG17, LGH^{+}17, MPTT14, MD16, NCS10, NBA^{+}15, NSM17, PLVB^{+}18, PS16, PBD^{+}12, RMCH^{+}14, SBT19, SA12, UD10, XYCL17, YAKK16, YLCZ12, ZWF^{+}18, dOSdAdSG17]. Evolutionary [GZY11, TCK14, AGR19, CV16b, GTY12, HJ14, PLHP^{+}15, XJZ^{+}15]. evolvability [BCL12]. evolved [GL14]. Evolving [PMR16, WGS^{+}14, HHKWB16, HGBS18, LWB^{+}13, RF14, URG10]. eVoting [Pen11]. exact [Kim17]. examination [PHR10, RNR17, Sta14]. Examining [DGCA17, Ifi11, BMB19]. exception [ECS15, KFLS18, SCL13, SHBA^{+}16, ZM18]. exception-related [ZM18]. exceptions [CF12, HdM17, OBS^{+}18]. exchange [RHRC13, RHRC15, WZM12a, WZM12b, YC12, YM13, ZG10]. exchanges [JS16]. executable [BLC^{+}18, ICSK14, KTT^{+}17, KH14]. Execution [BLL^{+}18, KMWL12, RXY^{+}19, ÁRMC16, ACH19, AAA11, CdAM^{+}14, DFJ19, EED16, FDÁM12, GGS15, HCB^{+}16, HSPD14, HS15, JJC^{+}14, KCT12, LWL^{+}13, NCK^{+}15, PH13, PPG^{+}10, SC19, SK18, TAS^{+}18]. Executions [MAR^{+}19, ASdMGM14]. Existence [MKRO14]. existing [MAGC^{+}17]. Exoneration [GLOM19]. Exoneration-based [GLOM19]. expansion [AQK11, JK13, LCT10, ZWM^{+}18]. expect [DOL^{+}16]. expectations [IF19]. expected [GGC16]. experience [AL10, FGB^{+}19, CMK^{+}11, CDS19, JS16, NBF^{+}19, SAH12, SJ17, WCC12, WKV11, WB15]. experiences [BDG13, HCY19, TCCH12, FH10]. experiment [CFRPC^{+}18, MNSA15, RZL^{+}18, SCMS15, WFF18, HWLM11]. Experimental [KOS15, BDD^{+}15, BDPRC18, LMIV15, LFCL12, OK11, OFR^{+}12, RM19b]. Experimentally [NSM17]. experimentation [YMM^{+}17, YMM^{+}19, FGMM17]. experiments [JDLS16, MNSA16, SKK^{+}18a]. Expert [SSR18, BDDS11]. expertise [Ifi11]. experts [RDVC19]. explaining [DNBM12]. explanation [CSN^{+}17]. Explicit [Cic16]. exploitations [SZ11]. Exploiting [ECRVMS11, GE15b, ILZ14, IZ18, TLZ^{+}16, TGE17, VT14, FDÁM12, FHL^{+}15]. exploits [WLZ^{+}17b]. exploration [JPGdL17, TAV13, vHJPB^{+}17]. exploratory [AMdLM17, BS12, CdSdSG^{+}18, ECS15, GCDY16, GW10, JR15, KNA11, MBF12, MAH18, MFM10, RASL12, SS12, TKZW17, WGKW19]. Exploring [BAM17, BGG10, IF19, KK12, MM19, OWB11, QGZ^{+}15, SPC16, SG16, ZZC18, JG14]. exponent [LCL15]. exporting [TTL^{+}13]. Expression [NTT19]. expressive [MMP15]. Extendable [NC10]. Extended [GMGTdFR14, CSW10, CH10b, EFSJM17, LMS12, OGRJ^{+}18]. Extending [MBT16]. extensible [LQLW12, Luk11, OAC11]. extension [BMB19, MLGA11, SdSLS^{+}19]. extensions [GCAH18]. external [Ifi11]. extra [TGE17]. extra-functional [TGE17]. extract [TC11, BDO11, FTSC12]. extracted [CCWT13]. Extracting [AK15, DCG16, SDB18, SHS16, JLGM17]. extraction [AACT13, BRB14, BKS15, BKSM13, BKSM14, CHN19a, EB14c, KKA^{+}19, LZL^{+}18, NBA^{+}17]. extreme [GJ13]. Eye [KWS^{+}17, GW10].

F [GMGTdFR14, PW18]. Face [ZLmLN14]. faceted [LAT10]. facets [KMG^{+}19]. facilitate [GSM15, HBR19, WWLG13]. facilitating [KCAS13, MDP^{+}11]. factor [GCDY16]. Factors [KNA11, MP12, VBC^{+}14, BPGS13, CPD^{+}18, DPL16, Jør14, LRD^{+}19, SNDC13, WR10, Wu11, ZZP15, ZP17, dSF12]. fail [AS10, AAB19, BAAD17]. fail-safe [AAB19, BAAD17]. failed [TTC15, ZYZ^{+}17]. Failure [FSS^{+}13, Jør14, DW11, DPVvV19, PD12]. failures [AD14, CLY14, WLL17]. fair [BV15, HH17]. fair-share [HH17]. fairness [TT10]. familiar [WLL17]. families [CGS19, CBAV16, CFAP17, KTF^{+}16, dMCR19]. family [MNSA16, PSNB11, PCCLdGP12, dAGSdFS^{+}15, SSS17, TFS10, WDC10]. far [DDMP14]. Fast [AAH10, PSM12, TT10, AAH12b, CL13, JHYK10, KAS18, LHY12, MBB11, PQBP16, VvSvV16]. Fault [BCS18, MGM10, aSRZ^{+}18, WL16, YSDT11, ZJC^{+}10, Al 12, AM15, ABJ10, BKLE18, BBBP13, CBS16, CCH14, CPR13, DW11, DW14, FP18, GGvH^{+}18, GPSS^{+}13, GXZ^{+}19, GLOM19, JJC^{+}14, KKHH11, Kim12, LGL^{+}10, LLH^{+}16, LLWL19, MLD^{+}14, MdFD^{+}15, MA17, PAR14, TXCX19, WL15b, WWSZ15, WMWZ12, WDC10, XYZ^{+}19, YLXZ16, YLYL17, ZCT^{+}11, ZS16, ZYZ^{+}17, ZZC18, ZXL10, ZHGL11, dCPV10]. Fault-aware [BCS18]. fault-prediction [dCPV10]. fault-prone [MA17, ZXL10]. fault-proneness [FP18]. Fault-tolerant [YSDT11, GPSS^{+}13, LLH^{+}16, WMWZ12, ZHGL11]. faultloads [CSM15]. faults [AZvG11, dSACdLF17, AMdLM17, MHLMG14, SRWE10, SPMG18, ZWF^{+}18]. Feature [BKS15, KCV^{+}19, TKK^{+}19, AGR19, BGEP17, BAM17, BLUH15, CHN19a, CFAP17, CV16b, GSM19, GJ13, GWW^{+}11, KKL^{+}11, KKA^{+}19, KMG^{+}19, LMN10, LHLG^{+}15, MRBN17, NCW^{+}19, PXT^{+}13, PBD^{+}12, PHBJ16, SdSGdMSN^{+}13, UIK17, WQJZ10, WBS^{+}10, WGS^{+}14, XLX^{+}19, YJZ17, dL13]. feature-based [KKL^{+}11, UIK17]. feature-driven [CV16b]. Feature-oriented [TKK^{+}19, LMN10]. features [AKL14, BZ10, BEK^{+}19, CCWT13, CRESF^{+}13, HHKWB16, KAU16, LYLC16, ZLmLN14, ZA12]. Federated [KAK^{+}13, AO16]. federation [NB13]. Feedback [XLW18, CBVF19, ILZ13, LGH^{+}17, LHCT19, NPC12, PCYZ12, RA16, ZJZ^{+}17]. Feedback-based [XLW18, NPC12]. FeGC [KKLB11]. Field [CRSS14, nQYD11, CVGP13, KL11, SCwY12, SCL13, CMK^{+}11]. fields [dGFDL16]. File [ZCC^{+}19, CCH14, JLZ^{+}19, KA14, LLLK12, LZCL19, Luk11, MCC11, MK17, PNY14, TXLC12, YCLC17]. File-level [ZCC^{+}19]. files [FSS^{+}13, RBS19].

Filling [SCGL^{+}18]. filter [AG15]. filter-based [AG15]. filtering [CCdR^{+}16, KK17a, LLWL14, ND18, PQBP16, ROFGFRM13, Shi12, XWZC14]. find [HG18]. Finding [dOFB^{+}19, JSHW14, MSGM17, SHGT16]. findings [RSGH12]. Fine [PPB19, ZML17, FSGW11]. Fine-grained [PPB19, ZML17]. finite [BKRW19, DCG16, EFSJM17]. fire [BRG^{+}12]. firewall [PGRQVV12]. Firm [CFMRL11]. firms [JKC19]. fission [HWR17]. Fitness [HBT16]. fitting [WQJZ10]. five [AS10, UGFK15]. fix [BCD^{+}18, HNH15]. fix-inducing [BCD^{+}18]. fixed [FHL^{+}15, Kim17, RXY^{+}19, wZfG14b, dOCS13]. fixed-priority [FHL^{+}15, dOCS13]. fixer [ZCY^{+}16]. fixing [ACB18, JSHW14]. FL [FP18]. FLANN [DRCG12]. Flash [PNY14, CH10d, KKLB11]. flexibility [LCCJ10, LWZ12]. Flexible [ES14, GBDCR12, KTF^{+}16, GCSSDP^{+}18, ILZ14, KLP10, LMT16, VRG^{+}16, ZL12b]. FlexIQ [ILZ14]. Flipping [CCGG14]. flocking [YSDT11]. FLOSS [HBR19]. Flow [BCF18, AM10a, ABFM12, CCdR^{+}16, DC17, LQLW12, SG16, APS16]. FLOW-assisted [APS16]. FM [GSM19]. FM-CF [GSM19]. focus [AHLH16, BPSK18]. focused [WSJ14]. Fog [MAR^{+}19, TMTB19, GGB19]. FogBus [TMTB19]. Follow [SSF15]. Follow-The-Sun [SSF15]. foraging [LL15, MPLL18, MCS^{+}12, VSDD12]. forecasting [LNY^{+}11, SKF17]. forensics [QZ14]. foreword [SY16a]. fork [GL14, OH15]. fork-join [OH15]. form [BHM12, OH15, Xia13]. Formal [BZ10, BCF18, KSN17, WKV11, BHH^{+}12, CTKT13, DAR14, DBZ16, DH13, GKV14, HdM17, MGM10, MA11, AHH^{+}10, MS17b]. formalism [KU10, SSF15]. Formally [PPS12]. Format [CDS10]. formation [LLHY19, OCC12]. formats [JH10, ZT14]. formed [VA17]. former [SNDC13]. forming [LS17a]. formulas [SGK12]. FORTUNA [GKD13]. forward [dONTF^{+}19, WLL17]. Foundations [DPVvV19, VPMVM^{+}13]. four [VBC^{+}14]. Fourier [GJ13, yWpWyYpN13]. FPA [FP18]. FPA-FL [FP18]. FPGA [MM14]. FPT [YH19]. FPT-approximation [YH19]. FPZL [dOCS13]. fractal [KM11]. fractional [MIUM12]. fragile [CCLL11]. fragility [CIB^{+}19]. fragmentation [HSPD14]. fragments [SGC^{+}17]. frame [LWL^{+}13]. frame-based [LWL^{+}13]. frames [LCC^{+}13, CKL12]. Framework [ILZ14, JS11, TMTB19, AV12, AM13, ATHM17, AK16, AAM^{+}17, BKLE18, BSG^{+}18, BSB12, CNM18, ÇT13, CPX16, CBC14, CMR19, CBC^{+}15, DBL^{+}18, DS16a, DBZ16, DM17b, ETYL15, FBB15, FTC16, FCC^{+}10, FMRM15, GSM19, GKD13, GN15, GPP^{+}17, GPMI13, GSN^{+}15, GDLB16, GMMC13, GZKL13, HGP^{+}12, HZH^{+}16, HSL14, ILZ13, KC16, KH14, KSAR18, KT12, KTK19, LCLP16, LSE12, LHH10, LDZL15, LC11, LNW^{+}11, LLC17, LZR16, MEB^{+}10, MIKG13, MAAC17, NK15, NBR^{+}14, OAC11, OCC12, PPG^{+}13, PWY^{+}16, PSdO^{+}13, PPMM12, PGRQVV12, RGV^{+}17, RAS14, RLL^{+}18, RM19a, RGH17, SJR^{+}11, SA16, SSP^{+}15, SBDB19, SWES16, TKJL13, TPGdS13, TTL^{+}13, TC16b, VM12, VPdP13, VRG^{+}16, VvSvV16, XLL^{+}19, YLA^{+}17, YAKK16, ZLC^{+}14, dRSBA13, fLSN18, rBHM17, vHAH12, CV14]. framework-intensive [RAS14]. frameworks [CdL18, GAKF13, MDP^{+}11, OLV15, PHR10, ROFGFRM13, SKL10, SPCT18, TJT^{+}18, TKJ15, RCL14]. Free [Shi12, BLS18, BL19, DFCPSF15, GW10, HL10, Kan15, RBW18, WDC12, Xia13, YAY13]. free-spirited [HL10]. FreeRTOS [GPPT16]. frequency [CS12, HFE10]. frequent [DS12, KKR16, KVT^{+}17, LJL^{+}12, LW13b, MSB18, NDS13, Sal17, ZJL10]. friendly [MCV15, PSNB11, WOLS12]. friends [EBC10, RNC14]. front [PSS11]. frontiers [WMC17]. FSA [LMS12]. FTM [AHH^{+}10]. full [JJC^{+}14]. Fully [ZZ12, ZML17]. fun [GCMB17]. Function [EAH^{+}11, CSW13, HBT16, LC10, TSCB19, WWSZ15, LWSH19, SB19]. Function-as-a-Service [LWSH19]. Functional [ABB15, DRCG12, EGM^{+}11, GD12, GEM15, HPF16, KR16, NSDI16, OMLB16, SA14, TTM13, TGE17, YCG^{+}14]. functionalities [RAJ15]. functions [LWBH16, MRBN17, TSCB19, TC12, MG11]. Further [WHY^{+}12]. fusion [TXLC12, YCF^{+}13]. Future [CG15, BMA^{+}13, BGEP17, CJT^{+}16, Chr16, DFG^{+}13, Fug12, PMR16, WTG^{+}15]. fuzzing [ZLL^{+}12]. fuzzy [BSKL10, BMLL14, KRDH12, LLWL19, MMSD13, SFMB16, SNM14, ANC11, MG11]. fuzzy-based [SFMB16].

GA [MMSD13, MA17]. GA-fuzzy

[MMSD13]. game

[LWL^{+}16, LHCT19, XJZ^{+}15, vWSB13].

gameplay [Dan17]. games [Dan17, GSM15]. Gamification [AM18, GPP^{+}17]. gangs [PK10a]. gap [CKL12, IF19, PFG13, SBDB19, TR18]. garbage [KKLB11]. GASR [FDN^{+}16]. Gateway [WZM12a, WZM12b]. Gateway-oriented [WZM12a, WZM12b]. gathering [CFA^{+}19, MC10]. Gaussian [ZL17]. gaze [KWS^{+}17]. General [Woh16, AAM^{+}17, KL10, LNW^{+}11, SNDD19, WSM15, WS12, YC11]. Generalized [CCGG14, YDGB^{+}12]. Generalizing [SED16]. generate [MM19, SGC^{+}17]. generate-and-validate [MM19]. generated [GPD^{+}19, LW13a, SCL13]. Generating [DV10, KTT^{+}17, ZYZZ14, CL18, Cic16, UIK17]. Generation [GKV14, AZ11, AG15, ÁGBYB^{+}14, ABC^{+}13, CLS^{+}12, EVR11, EGM^{+}11, FAM15, GZY11, GTY12, GEM15, HY11, HBT16, HZH^{+}16, HWC^{+}10, KL10, KL11, KD18, MSHG18, PS13, PAOC15, PW18, TAF^{+}17, TGKL19, VRPT18, VPMVM^{+}13, ZAY19, ZYY^{+}19]. generator [GP10b]. generic [BMS11, DK15b, KD18, XPBC11]. Genetic [JK13, TGKL19, AR18, AG15, EEAZ13, GWW^{+}11, KSN17, KLB15, LHJ10, PS13, RCCVB11]. GenProg [MM19]. geographic [DBCdP11]. geographically [CdR^{+}14]. geolocation [PWY^{+}16]. geometrical [TLL13]. gesture [SHBC19]. gesture-based [SHBC19]. gestures [GCSÁddP11]. GeX [MMP15]. GitHub [BV18, JR15, TNK^{+}19, TLA18]. gives [Jør16]. GLBM [ZADM10]. global [APCS10, BHH^{+}10, BBS10, CL18, GGS^{+}19, GBC16, Jør14, KK11, KR14, LH11a, LCLS16, SKRB19, ZGL^{+}10, dOCS13]. globally [KM14, TR18]. Goal [LMR12, PZ15, CPYZ14, CHL^{+}13, GPMI13, MTF14, PNJGF12, SCS15, ZWM^{+}18]. goal-based [GPMI13]. Goal-driven [PZ15, CPYZ14]. goal-oriented [CHL^{+}13, MTF14, PNJGF12, SCS15, ZWM^{+}18]. goals [CFAP17, GBH^{+}16, KGG18, MPS^{+}12]. GoF [ACS13]. Going [DC17]. good [CHL^{+}13]. governance [HBR19, VvSvV16, Wau19]. GPU [BAI^{+}14, HCB^{+}16, MBB11, PS14]. GPU-SAM [HCB^{+}16]. GQM [GPMI13, KVGS11]. GQM-based [KVGS11]. GQM-DSFMS [GPMI13]. gradient [YCLY13]. gradient-based [YCLY13]. grain [FSGW11]. grained [PPB19, ZML17]. gram [SPSR17]. grammar [HWC^{+}10]. grammar-based [HWC^{+}10]. grammars [PACH15]. grammatical [RMCH^{+}14]. granules [IBM11]. Graph [QGZ^{+}15, WWLG13, ÁGBYB^{+}14, BKLE18, CL17b, HWR17, LQLW12, MMP15, PXT^{+}13, PRN17, RSCB18, YLYL17, CJ13].

Graph-based [WWLG13]. graph-modeled [MMP15]. graphical [CTL12, LK16, OFR^{+}12]. graphical-based [CTL12]. graphs [AR12, BP13, BNS12, SK10]. gray [Che13, JBSL12, UUN13]. gray-level [Che13]. greedy [KHS11]. Greek [KKA^{+}19]. Green [LZL^{+}15]. grid [ALRP16, DHC^{+}11, JLQ^{+}10, LT11, SRS15, Sko14, WS12, YWEL^{+}13, ZCZZ11, GQ12, KK11, PM10, SLLL12, XPBC11]. grid-based [LT11, WS12]. grid-density [ZCZZ11]. grids [CT11b, CLH^{+}13, GLW13, HSC15, TdCAF16]. Grindstone4Spam [MRJD^{+}12]. ground [KA17]. Grounded [GN15, WLD16, AKH12, JG14, JMML17, SSD16]. groundwater [LHP^{+}09, LHP^{+}10]. Group [GTF17, SZS13, BPSK18, IF19, LCC10, YSDT11, ZcKS17]. group-based [BPSK18]. group-by [LCC10]. grouped [SD16a]. Grouping [GTY12, GZY11]. groupware [BDG13, MGR^{+}13, PLGT10]. Growing [HHKWB16, EZG15, KHMA12]. growth [KLB15, RSB^{+}14]. GSR [RFM10]. GTCharge [LWL^{+}16]. guaranteed [LWL^{+}13, LGHR16, LLK11]. guaranteeing [FCC^{+}10]. guarantees [AMP12, CGS19, LGZ^{+}18]. Guest [BDV17, CCM12, SY16a, CdS18, OPS11]. GUI [BRB14, HCC10a, YCG^{+}14]. guide [PPG^{+}10, dSF12]. guidelines [CPDM16]. guiding [LK13]. Guilt [TKCR14]. Guilt-based [TKCR14].

H.264 [LCC^{+}13, LLML13, LW13c].

H.264/AVC [LCC^{+}13, LLML13, LW13c]. hack [SCwY12]. Hadoop [LZCL19, MK17, SGW^{+}15]. Hadoop-based [LZCL19]. Hamming [CCLL11, ZGZ^{+}13]. handheld [CTL12]. Handling [BBA10, CF12, FS14a, WQJZ10, CPYZ14, ECS15, IYS13, KFLS18, LNW^{+}11, OBS^{+}18, TKCR14]. handoff [PZB10]. Handover [AAH10, AAH12b, CL13, EZOK14, LRD^{+}19]. handshake [WZ11]. HaoLap [SGW^{+}15]. happens [GFWA18]. happy [GFWA18]. hard [LSE12, LWL^{+}13, WMWZ12, wZfG14b]. hardening [AMKD13]. hardware [GKD13]. hardware-based [GKD13]. harmonization [PPG^{+}13]. harmony [ZGL^{+}10]. hash [PPB16]. hash-based [PPB16]. Haskell [LSNL^{+}19]. haystack [dOFB^{+}19]. hazard [KHC16]. HB [NJ17]. HCI [CdCAdO18, JST10]. HDD [CH10d]. HDFS [DZT^{+}14, LLGZ13]. healing [TTC15]. Health [HWdS^{+}15, LZHS11]. healthcare [PPN^{+}15, VPL^{+}10]. heap [HHH^{+}10a]. Heart [VPL^{+}10]. help [Aki18, ABL16, LK16, RNC14, RDPM19, vHAT13]. heterogeneity [CDGJ10]. heterogeneous [AR18, AYZI10, BLM10, CLY17, CTHW12, DBL^{+}18, DK15a, DFJ19, GPL^{+}15, JRO12, KHS11, MMZ^{+}16, MK15b, NEM17, NTRN11, OZO^{+}14, PK10a, TBC^{+}16, WH15, Zha12a, ZLD13, ZCC^{+}17, ZSB19, ZGSH13]. Heuristic [dNPM18, KS16, TVMS18, TPGdS13]. Heuristic-based [dNPM18, TPGdS13]. heuristics [DHC^{+}11, FSGL12, WDC10]. hidden [LZL^{+}18]. hidden-code [LZL^{+}18]. hiding [AQK11, CNL13, FF12, HC10, HWL13b, HTH13, LCT10, LC10, LCLF13, LBCL10, Lin12b, LCC^{+}13, LLML13, LTW16, OLZN13, PMDH13, PWC12, QZ12, UUN11, WCCL10, WLH13, WYCC13, WLC13b, WCC^{+}14, YWTW11, YWHL11, YCLY13]. Hierarchical [ABB19, BLLGSMB11, GBC16, KKG^{+}12, LH11b, NZM10, RG10, SS13, WWYZ11, WL15b]. hierarchy [CCD19, LZKW12]. High [AQK11, CS12, Lin12b, AdB13, AHLH16, AKA^{+}15, BML^{+}13, CS19, FF12, FTC16, HTH13, LCC^{+}13, NJ17, PN14, PC15, RLY^{+}13, RQD^{+}17, SPMG18, Shi17, SS13, TBC^{+}16, TC12, WLH13, WYCC13, WCC^{+}14, WKH11, XZP^{+}10, ZHH^{+}17, ZCZZ11, ÇT13, NK14]. high-bandwidth [NJ17]. high-level [PN14, PC15, TC12]. high-performance [FTC16, RLY^{+}13, Shi17, WYCC13, NK14]. high-speed [XZP^{+}10]. higher [LHJ10, dPLV19, nQYD11]. higher-order [nQYD11]. highly [BNS12, CSS10, JLQ^{+}10, PSS^{+}16, PDBD18]. highly-accurate [BNS12]. Hindering [BTPLST15]. HiP [MBPM19]. HiP-HOPS [MBPM19]. HIPAA [HL11]. Histogram [WLC13b, CSS^{+}13, HC10, HTH13, Lin14, LTW16]. histogram-shifting [HC10]. Histogram-shifting-imitated [WLC13b]. Historical [JRSN10, RSB^{+}14, SYXL17]. history [HPH12, KM17, OKS^{+}15, PDBD18]. history-based [HPH12]. hitting [TXCX19]. Hoc [ACSC16, ACL13, BCLW11, hChSyCwL10, CWK10, Cho13, KSHC14, LLHY19, MLHL12, MDO^{+}10, WOC15]. holes [NNVD17]. Holistic [GGS^{+}19]. Home [LDZL15, CFL^{+}18, GGB19, KLP10, SJR^{+}11]. Home-diagnosis [LDZL15]. homeostasis [GSP^{+}19]. homing [HSM16]. honeybee [KHSD10]. hop [CW12, JXLC15]. HOPS [MBPM19]. horizon [HZG^{+}12]. hospital [TKSRP11]. hosting [RQD^{+}17]. hosts [Wen16]. Hot [WLZ^{+}17a, WMOKY11]. hot-spot [WMOKY11]. hour [ABJ^{+}17]. hours [Jør16]. house [BWP16, Ifi11]. HPC [CNM18]. HPobSAM [KJS^{+}12]. HSFal [JJC^{+}14]. HSP [HHH^{+}10a]. Huffman [LHY12, YWHL11]. Huffman-code [YWHL11]. Human [Woh16, CFRPC^{+}18, LWW^{+}10, WSM15, YCG^{+}14]. human-perceived [CFRPC^{+}18]. Hurst [LCL15]. HVMs [CBZ^{+}16]. Hybrid [KAM13, KR16, LWC^{+}18, PN14, BDBLP15, BT17, CCdR^{+}16, CNL13, CDOP15, DBCdP11, DAG19, EEAZ13, JS11, JJC^{+}14, KHMF13, LMT16, LG17, LZCL19, LT11, LQW^{+}12, MLHL12, QOLJG16, SBZ^{+}17, SLW^{+}15, YXH^{+}18, YH10]. hybridization [MMSD13]. Hyper [KS16, TZB19, WGZ^{+}12]. hyper-chaotic [WGZ^{+}12]. Hyper-heuristic [KS16]. hypervisor [PWY^{+}16]. hypervisor-based [PWY^{+}16].

i* [MNSA15]. I-Cache [CWK^{+}13]. I-star [MTF14]. I/O [FTC16, SMZC12, SC19]. IaaS [DVV^{+}16, DR12]. IBM [XPBC11]. IBUPROFEN [PCFRP19]. ICSR [WB19]. ID [HCC10b, IB11, Shi10, SV12]. ID-based [HCC10b, IB11, Shi10, SV12]. IDE [ÇT13, GMR17]. IDE-based [GMR17]. ideal [BMLL14]. Identification [FSGYP17, FTSC12, Sal17, TC10, TC11, CKS15, HZ15, KM14, PG12, RO13a, TLGE18, CPX16]. identifier [AACT13]. identifiers [CAHV15]. identify [HJ14, LLWL19, TTC15]. Identifying [BDO11, FBB^{+}12, WLZ^{+}17b, WRR14, CBVF19, LZJ^{+}19, XCM^{+}12]. Identity [HYWS11, RG10, SA16, WWYZ11, YYS^{+}16, YKC^{+}12, ZZ12]. Identity-based [HYWS11, RG10, SA16, YKC^{+}12, ZZ12]. IDF [LCLP16]. idle [SRS15]. IDTV [BPB19]. IEEE [KT16, LH12, AAMS14, PZB10, WC11]. IEEE/IFIP [KT16, LH12]. IFIP [KT16, LH12]. IFML [BCF18]. IKAROS [FTC16]. illustration [AB10]. Image [BAAS13, Che13, KPS10, PWW10, CPL13, CT11a, CJ13, CW14, EA11, HRB12, HHH10b, KRDH12, KM11, KAS18, LXCM11, LW13a, NES^{+}14, SNM14, jT12, TTL10, TLL13, UUN11, UUN13, WCCL10, yWpWyYpN13, WGZ^{+}12, WKH11, WOLS12, WS13, XZZ^{+}16, YC11, YCC16, ZLW^{+}12, ZT14, ZL12b]. images [AQK11, AMK12, CCWT13, Che13, HSL14, HWL13b, HHC12, HTH13, KSRD10, LW13a, MM14, MKH^{+}12, UUN13, WLH13, WCC^{+}14, YWTW11]. imbalanced [LLC17]. imitated [WLC13b]. Impact [JAS19, Al 12, AHBA19, Ano13a, BHH^{+}10, BBS10, CS15, CS16, CCP18, CREH^{+}18, CFA^{+}19, DNSH13, HGBS18, HFE10, HWLM11, IYS13, KA18, LRB^{+}19, MS16, MT13, PB11, PSZ17, RvDV17, SLLL14, SLL^{+}15, Wau19, dL13]. impacted [AMdLM17]. Impacts [WKbOS17, Li11, NBF16, SPC16]. Impala [MCL^{+}17]. imperative [BSB12]. Imperceptible [Lin14]. imperfect [WWSZ15]. implement [FCRF16, HdM17]. Implementation [HN17, BBA10, BAI^{+}14, CdSdSG^{+}18, CNSG12, DS16a, HJP15, LLGZ13, MM14, NES^{+}14, PSNB11, PPS12, SDB16, ZADA15]. implementation-friendly [PSNB11]. implementations [dB12]. Implementing [AAN11, CMK^{+}11, FHL^{+}18, JKC19, PN14, RAJ15, SA16, SV19]. Implications [LHCT19, APCS10, CFL^{+}18, MVLJ18, Han12]. implicit [OWB11]. implied [dMCR19]. importance [LMPM18, OK18]. Impossible [TSLL11, LGLL12, SDM10]. imprecise [SK10]. impressions [BCG^{+}14]. improve [BLLGSMB11, BGLG13, JDLS16, MTF14, SM17b, VPdP13, YWHL11, YM13, ZYZ^{+}17]. Improved [BL19, LKP13, LGLL12, QZ14, ZL12a, BLUH15, DRCG12, GLW13, HWL13b, IZ18, KSN17, LCC^{+}13, PS13, PWLL13, SDM10]. Improvement [CWK^{+}13, AAGT16, BD16, CSW13, GMMGP15, GC13, GLJ13, KSAR18, LMR12, PW10, PPG^{+}10, VLC^{+}17, MMB10]. improves [vdBSvS^{+}19]. Improving [CFAP17, CSW10, CJ13, FCB^{+}16, GMS11, KAO13, KA14, LGC17, LZKW12, LZR16, OKS^{+}15, PB15, PXT^{+}13, SOS^{+}18, SB12, TBSvdW18, XSL^{+}18, DY15, EA19, HJBH10, KCT12, KM14, LMNA17, RSB^{+}14, VSDD12]. imputation [HKS^{+}17, TC16a, ZJZ11, Zha12b]. in-depth [KM17]. In-house [BWP16, Ifi11]. in-vehicle [BKLE18]. inaccurate [CBVF19]. incentive [ZSB19]. incentives [LLW12, dVRB13]. inception [CBSM16]. incidents [ABL16]. incomplete [ZJZ11]. inconsistencies [EA14, EUR^{+}13, FKWVH19]. Incorporating [CCdL^{+}16, YLXZ16, FP18]. Incorrect [JDLS16]. Increasing [GKS18, PKS18]. Incremental [IYS13, CLY14, LCLP16, dNPM18]. incrementally [YF15]. independent [BEK^{+}19, CF13, DDD14, FLRT19, ZGSH13]. index [JRSN10, Lin12a, SLLL12]. indexes [YWHL11]. indexing [TBC^{+}16, ZXTT11, ZHH^{+}17, FSS^{+}13]. indicators [YC13]. indirect [AAM16, GMGTdFR14]. indoor [DC11]. inducing [BCD^{+}18]. induction [BBBP13]. Industrial [BKW10, Woo12, AAGT16, APW14, ABJ^{+}17, AHC^{+}11, CCdL^{+}16, GGT^{+}19, HVK11, KBJZ15, KSM^{+}16, LWSH19, MSSMDC12, PCFRP19, SCwY12, SCL13, SCC16, SM16, Sta14, SAN^{+}17, VHFF^{+}17, WGKW19, WB15, YLA^{+}17, dSdMSNO^{+}14, dOSdAdSG17, ELHC13]. Industry [Bis13, HBR19, CBT^{+}14, EB14a, EbAT13, ETM10, GK18, HTB12, MSB18, MTA^{+}16, MFTP18, SB14, TTR^{+}13, WRR14]. infeasible [KSS15]. inference [RSB^{+}16, TSRC18]. infinite [ASdMGM14]. inflow [RSB^{+}16]. influence [ARH^{+}17, BRS^{+}18, CO12, EED16, HSM16, KS19, SJ17, SS15]. influences [Ifi11]. influencing [KFN19, KNA11]. informal [NBA^{+}17]. Information [AAH10, CMM15, PCG^{+}14, ABFM12, BPO^{+}16, BDBLP15, CPL13, CSW10, CH10b, JAS19, KCV^{+}19, KHW19, LS17a, LK16, LTW16, MCC11, MKH^{+}12, PMDH13, PDS19, nQYD11, RNC14, ST13, SHGT16, SYXL17, VM12, WCCL10, WRS^{+}17, WB15, YAY13, YAT11, ZLZ11, ZJZ11, ZYY^{+}19, ZMK12, BDGP13]. informations [AAH12b]. infrastructure [AO16, CX10, CMM15, DBL^{+}18, TG17, WC11]. infrastructures [GQ12, SGEK19]. inheritance [NCS10, RO13b]. Inhibitors [ESWA18]. initial [MAAC17]. initiatives [GMMGP15]. injection [PDK^{+}16, RNC14, YXH^{+}18]. innate [BDD^{+}15]. Innovation [ESWA18, EbAT13, ZÁ15, LMWM18]. innovative [GGS^{+}19]. Input [JC15, LXJL10, RHHT18, SRT^{+}12, SED16, WLZ^{+}17b]. Input-based [JC15]. input-centric [RHHT18]. input/output [SED16]. InRob [MFMCY12]. insertion [JRSN10]. insider [dMSSS^{+}13]. insight [KTF15]. insourcing [SWA^{+}13]. inspection [FAI13, SdSGdMSN^{+}13]. inspections [CTKT13]. inspired [MDO^{+}10, NEM17]. installations [CMK^{+}11]. instance [LTK^{+}15, TCK14, TC16a]. instances [YZC15, YZY^{+}18, ZJZ11]. instantiation [VPdP13]. institutions [CLL14, KLA^{+}19, WTG^{+}11]. instrument [JC10]. instrumented [OM13]. insulated [RG10, WWYZ11]. integer [AMK12, CAG17, Lin16]. integral [DAR14, SNM14]. integrate [JRO12]. integrated [KSAR18, LJM11, TLWS10, TTMI19, XLW18]. Integrating [KAU16, KFN19, MPAA15, MMTS15, WK15, ZTCZ16, DK15a]. Integration [FSPH^{+}16, VCMG17, AT15, BBS10, FCRF16, DDF^{+}13, HMOK18, HLW^{+}15, ICSK14, JST10, KM17, LLX^{+}11, LLL^{+}14, MSB18, NTRN11, RPK^{+}13, SB14, SMB17, SJH^{+}10, WBBK18, ZJZ^{+}17, FCMJ12]. integrator [JLZ^{+}19]. integrators [LMPM18]. integrity [ZTZ^{+}11, ZHAY12]. Intel [DSGS17]. intelligence [PGP^{+}19]. Intelligent [AMK12, BD16, GGB19, LPP^{+}10, MKH^{+}12]. intensive [AAA11, FOR19, GBH^{+}16, MAH18, RAS14, RHL^{+}17, SCL13, SMM17, Shi17, dSSVV11, YMM^{+}17, YMM^{+}19]. intentional [MBF12]. intentions [GA11]. inter [AHLH16, BML^{+}13, CBZ^{+}16, FKWVH19, MQG^{+}17, SHBC19, WK15, WLC13a, ESM19a]. inter-domain [BML^{+}13]. inter-model [FKWVH19]. inter-organisational [WK15]. Inter-organizational [ESM19a]. inter-player [MQG^{+}17]. inter-process [AHLH16]. inter-sequence [WLC13a]. inter-session [SHBC19]. inter-VM [CBZ^{+}16]. interact [HA10]. interaction [AZ11, GBDCR12, HSPD14, HLWS13, HCT^{+}15, KWS^{+}17, MOD^{+}19, BCF18]. Interactions [MAR^{+}19, SÁMI17, SKK^{+}18b]. Interactive [FSGW11, AM10b, Hoo14, ILZ14, MFTP18, MGR^{+}13, PGP^{+}19, QXYL16]. interconnection [BMAH11, WMOKY11]. interdiscipline [FP19]. interest [TZ12]. interesting [ZZ16]. Interface [GC13, AYZI10, MCV15, TKZW17]. interfaces [Aki18, AK15]. interference [AdAD17]. interleaving [BP15]. interlinked [MK15b]. intermediate [LSE12]. Internal [ESWA18, GAKF13]. International [SS17]. Internet [CG15, CRL^{+}12, FGBC10, JSM10, MOD^{+}19, Pal12, PC15, RLL^{+}18, SST16, Shi12, SXYW14, WTG^{+}15, ZXG10]. Internet-scale [JSM10, SXYW14]. internetworking [VT14]. Interoperability [RCL14, MFMCY12, NSDI16, GMGTdFR14]. Interoperable [MIBV14]. interpersonal [WKbOS17]. Interplay [AJLS10, AC17]. Interpretation [JK12, ADET12, OMLB16]. interpreted [AMCC14]. interruptions [FGBC10]. interval [LNY^{+}11, LYC14]. intervention [APT^{+}12, VvSvV16]. interventions [SSMvD16]. interview [AHC^{+}11]. intra [LCC^{+}13]. introduced [HHKWB16]. Introducing [Ano19l, Ski13, WBBK18, YMM^{+}17, YMM^{+}19, TC10]. Introduction [BCG^{+}13, DIB14, FKA16, FOR19, GBG10, ML18, NBM19, OPS11, PBM19, SS17, WMAS12, WMC17, XST18, ZTPT18, dAK18, BDV17, CCM12, CdS18, NBF^{+}19, PS16]. Intrusion [SKE10, YKC^{+}12, LG17]. Intrusion-resilient [YKC^{+}12]. invalid [SLLY17]. invariance [KAS18, yWpNyL11]. invariant [LXCM11]. invariants [CCGdL10, TLL13, WL16]. inversion [SYT^{+}17]. inverted [Luk11]. Invertible [UUN13]. investigate [ASGJ13]. Investigating [HNH15, JWA14, KOS15, MBF12, SPMG18]. investigation [AKKS11, ARH^{+}17, ABJ10, BDD^{+}15, CFL^{+}18, CH10c, GTA14, GTF17, KKA^{+}19, LMH10, LZO^{+}16, MNS13, ZADA15]. investigations [LRB^{+}19]. investment [vdBSvS^{+}19]. involvement [CFMRL11]. involving [JSM10]. iOS [LZHS11]. IoT [CDS19, DS16a, GGB19, WNC17]. IP [BP15, OSH^{+}18]. IP-based [OSH^{+}18]. IPAC [KVH12]. IPv6 [AAH12b, AAH10, CL13]. IR [BLUH15]. IR-based [BLUH15]. IRC [HB13]. IRC-based [HB13]. ISBSG [dGFDL16]. ISO [DRCG12]. ISO-FLANN [DRCG12]. ISODAC [TBC^{+}16]. Issue [BCEF10, CCCY17, OPS11, SS17, WCTK12, AC19, ADMOK^{+}10, BEZ14, Bor12, BCG^{+}13, CCM12, CdS18, CLR18, CA14, CL11, DIB14, Dut15, FKA16, FOR19, GBG10, JWT17, LH12, ML18, MJ19, MACB19, MS17a, NBM19, PBM19, PS16, RCPZ19, TZB19, TB13, VZT17, WB19, WMAS12, WMC17, WC16, XST18, YAT11, ZTPT18, dAK18, GP10a, Won10]. issue-based [TB13]. issue-commit [RCPZ19]. Issues [FGD^{+}17, ALRP16, CDS10, EGHO16, JR15]. iStar [GCAH18]. IT-outsourcing [DPVvV19]. Italian [ETM10, TTR^{+}13]. item [CLL10]. items [ACL13]. itemset [DS12, NDS13]. itemsets [ZJL10, CCGG14]. Iterated [LM15, KHS11]. iterative [PXT^{+}13]. iteratively [Zha12b]. iTravel [YH13].

J [AAH12b, APS^{+}10, BKSM14, LKJR10a, LHP^{+}10, WZM12a, XTZX13, YWEL^{+}13, wZfG14a]. Jaccard [LQC^{+}14]. Jacobian [BAAS13]. JaguarCode [YLC18]. Java [WLL19a, ASdMGM14, AYZI10, ABFM12, BD17, FGB^{+}19, ECS15, ES14, EED16, HdM17, HWLM11, LRO19, MLGA11, NCS10, PTF^{+}15, QLBS17, SS14b, WLL19b, ZYY^{+}19]. JavaScript [HHKWB16]. JCSI [ABFM12]. JMove [TVMS18]. Job [GQ12, PRS11]. jobs [AR18, LZY^{+}15]. join [OH15]. Journal [Ano19l, HST16, TTT14, WLL19a, Woh16, YMM^{+}19, vV13]. JPEG [HWL13b, QZ12, QZ14, WLH13]. JSS [BCG^{+}13, GP10a, MJ19, Won10, WC16]. judgment [JH10, Jør10]. judgment-based [JH10, Jør10]. junior [vHAT13]. just [PPB19]. just-in-time [PPB19].

Kalman [AG15]. Kanban [ADCO18]. KAOS [MNSA15]. Kernel [CYT16, Fei12, IF10, LWBH16, OY16, SCwY12]. Key [ROR11, HL11, IB11, LH11b, LW13a, RG10, RPSL10, SLZ12, Shi17, TLL12, WWYZ11, WZM12a, WZM12b, YC12, YLZ^{+}16, ZG10, OHJ10]. key-insulated [RG10, WWYZ11]. key-value [Shi17]. keyrings [MBB11]. keystroke [CTL12, Kan15]. Keyword [TZ12, BL11, GZS^{+}18, LWXZ10, WHY^{+}12]. kill [LGC17]. Kintala [TG10]. kit [FCRF16]. Kmeans [LQC^{+}14]. Knowledge [HHK13, AAH12a, BSG^{+}18, CJT^{+}16, CHL11, FB18, FH10, GK18, GLJ13, Ifi11, JS11, Kel15, KK17b, LJA^{+}11, LLL^{+}17b, NTdSX13, SdSLS^{+}19, SKE10, SBDB19, TAJ^{+}10, TNK^{+}19, TBG13, WBBK18, YCG^{+}14, ZMK12]. knowledge-based [SKE10, TBG13]. Korean [Kan15]. Kuali [LWZ12].

label [CTHW12]. labeling [KA17]. labelling [XYZ^{+}19]. landscape [KS19].

Landscaping [KSIZ19]. Language [ASMM18, BCF18, ARS17, BKS15, BWH10, BEK^{+}19, CAHV15, CF13, CG12, DPP^{+}18, EMBS17, HHKWB16, HGMB13, KMWL12, KMK16, LPXL10, LOFA17, MAGC^{+}17, Mer13, MGR^{+}13, NBR^{+}13, PC10, SA14, SPSR17, SAH12, KVH12]. Language-agnostic [ASMM18]. language-independent [CF13]. languages [ACG^{+}15, AMKD13, BBA10, BSB12, QOLJG16, RO13b, SKL10, SHS16, TFS10, VPdP13]. Large [ESWA18, FNWL18, TPTV17, AHLH16, APS16, AAA11, BLL^{+}18, BV16, ÇB16, CSM15, DvdVA^{+}13, DPL16, EEAZ13, EH19, GTF17, KLL^{+}11, KKR16, LCL^{+}12, Lin12a, LTK^{+}15, LWOY16, LLL^{+}14, nPHW^{+}16, PFG13, PTF^{+}15, RSB^{+}16, RVCM17, SAH12, Shi17, SVM19, SM16, SGO13, SYT^{+}17, SAN^{+}17, TTL^{+}13, TTMI19, WK15, WFF18, WB15, XWZC14, YMM^{+}17, YMM^{+}19, ZK13]. large-scale [APS16, BLL^{+}18, CSM15, DvdVA^{+}13, DPL16, EH19, KLL^{+}11, LLL^{+}14, nPHW^{+}16, PFG13, PTF^{+}15, SAH12, Shi17, SAN^{+}17, TTMI19, WFF18, WB15, ZK13]. Lascad [ASMM18]. late [MRS18]. Latency [SB19, BCA^{+}19, KY10, LJC16, MJ18]. Latency-aware [SB19]. lattices [CHN19a]. Laws [dOSdAdSG17]. laxity [LESL11]. Layer [AAH10, KCAS13, AAH12b, CTL10, CTHW12, KSHC14, LLLK12, PZB10, PGRQVV12, SRWE10, ZYZL12, LZG15]. layer-3 [PGRQVV12]. layered [BPO^{+}16, MGB16, NTRN11, LLW12]. layers [ST11]. layout [DS12]. LBS [JRSN10]. LCBM [PQBP16]. LDFR [XLX^{+}19]. LDoS [jWLY^{+}13]. lead [JAS19]. lead-time [JAS19]. leader [LMS11]. leading [DLW^{+}13]. Leagile [WCC12]. leak [CCN^{+}10, SS14b]. leakage [WWL^{+}10, YLZ^{+}16]. leakage-aware [WWL^{+}10]. LEAM [PW10]. Lean [ESWA18, NBF16, PFG13, WCC12, PW10]. learned [RB16, SKK^{+}18a, VM12]. Learning [LTK^{+}15, LMS12, XLX^{+}19, AIE19, AAN11, BNS12, CX10, EB14c, FMPS16, GJ13, JRO12, KCT12, KHMF13, LG17, LLC17, MDP^{+}11, NI13, PCCK18, RCPZ19, SRSC16, SZW^{+}16, SGEK19, VTZ^{+}17, VLL18, WQJZ10, WCX15, Xia13, XHM^{+}11, YXH^{+}18, ZFS15]. learning-based [EB14c]. least [HBM19, MT10]. least-privilege [HBM19]. least-time [MT10]. leave [GA11, KM13]. leave-one-out [KM13]. lecture [ZXG10]. led [MPLL^{+}15]. legacy [CRESF^{+}13, KTK19, SV19, YM13]. Lego [SBAH17]. legs [SCMS15]. Lehman [dOSdAdSG17]. length [GGC16, LWC13, XTZX12, XTZX13]. less [CLY14]. LESSON [ZXG10]. lessons [RB16, SKK^{+}18a, VM12]. Let [BDDS11]. leukocyte [HHC12]. Level [SRS15, WL15a, ÁRMC16, BLS18, CS16, Che13, FSG^{+}11, GE15a, GBC16, HM16, HFE10, IYS13, LCLS16, MT10, OKMD12, PN14, PC15, PPMM14, PK10b, PBD^{+}12, Sal17, ST13, TC12, UUN13, WLZ^{+}17b, WL10, YC13, ZJZ^{+}17, ZCC^{+}19, ÇT13]. levels [AL10]. Leveraging [AVGM19, HCB^{+}16, dOFB^{+}19]. libraries [LRB^{+}19, SOS^{+}18]. LibreOffice [GL14]. license [KTF15, KKT17, LGC17]. licenses [SA12]. life [Fei12, SS15, WB12]. lifecycle [SRBT18, TTMI19]. Lifetime [YZG^{+}13]. Lightweight [SCwY12, TMTB19, KT12, PSdO^{+}13, PQBP16, RQD^{+}17, ZADA15]. like [ZLZ11]. likelihood [RBW18]. likelihood-free [RBW18]. likely [TSRC18]. limitations [HLWS13, MMTS15, SBAH17]. Limiting [CPYZ14]. line [AC19, BWW^{+}18, BBD18, CHN19b, CV16b, DGRN10, DWC17, FHY17, HGBS18, HPF16, JKL19, KCV^{+}19, LMN10, LNTS19, MCV16, MB10, PBD^{+}12, SSS17, UIK17, dSdMSNO^{+}14, CBT^{+}14]. linear [ACH19, CAG17, MPAA15, NHC13]. lines [AWSE19, BKS15, BdMSNO^{+}17, BBS10, CdSdSG^{+}18, CNKL12, ESM^{+}19b, FFV19, GWW^{+}11, HBOS13, KTF^{+}16, LGS^{+}19, MAGC^{+}17, MD16, NBA^{+}15, OGRJ^{+}18, PLHP^{+}15, RTM19, ROR11, SBT19, SdSGdMSN^{+}13, TBG13, WVT^{+}14, WAG15, WGS^{+}14, dOSdAdSG17]. Link [AAH10, AAH12b, DRCG12, KR16, PSM12, RNC14, SCGL^{+}18, SZS13]. Link-Layer [AAH10]. linkage [dNPM18]. Links [HRRC16, KBDGAW16, RCPZ19]. Linux [ABB19, Fei12, IF10, SMZC12]. Linux-based [ABB19]. list [SBZ^{+}17]. lists [LMIV15]. Literature [GCAH18, SKT17, AAGT16, AKAA18, AVGM19, APW14, AT15, AS16, BWP16, BKS15, BMB18, CFL^{+}18, CP15, CS19, DVPY^{+}19, DPL16, DBCG14, GJ16, GNA17, GA11, IHA16, JED18, KGB11, KNA11, LFW15, LL15, MH13, Man16, MRT17, MRY17, MGAN18, MMB10, OGRJ^{+}18, PG12, PMB15, PFO^{+}19, RAK15, STA19, TKP^{+}18, TCS18, VLC^{+}17, VCMG17, ZADA15]. little [RNC14]. live [FGLI15]. lives [TLK^{+}16a]. Living [RASL12, CFAP17, GMPN16, BHH^{+}12]. LLVM [RSCB18]. Load [MCC11, SLW^{+}15, BVV^{+}10, CV16a, CCH14, CS12, DY15, FS19, LJL^{+}12, NNVD17, WOC15, YCF^{+}13]. load-balancing [DY15]. Load-prediction [SLW^{+}15]. local [CL18, FF12, HC10, JC15, KAU16, LM15, LWW^{+}10, ZLmLN14]. localisation [aSRZ^{+}18, XYZ^{+}19]. locality [KC16]. localization [DC11, DW14, FP18, GXZ^{+}19, GLOM19, JJC^{+}14, LLWL19, MLD^{+}14, PAR14, TXCX19, WL15b, WL16, WDC10, YLYL17, ZJC^{+}10, ZCT^{+}11, ZS16, ZYZ^{+}17, ZZC18]. localize [dSACdLF17]. Localizing [ZWF^{+}18]. located [SHHL12]. location [AACT13, AL10, BLUH15, CFAP17, IBM11, NCS10, PXT^{+}13, WCC13, XSL^{+}18, dL13]. location-dependent [IBM11]. lock [PMWC12]. Log [XPBC11, BLL^{+}18, FSS^{+}13, MK17, NHC13, WWSZ15]. log-linear [NHC13]. log-logistic [WWSZ15]. logic [BMLL14, EBEL18, KAO13, KB16]. logic-based [BMLL14, KAO13]. Logical [MCL^{+}17, AC17, HJ14]. logistic [WWSZ15]. logistics [Hoo14, TTL^{+}13]. logs [LGH^{+}17]. lonesome [HFLvV11]. long [CFAP17, SB17b, UD10]. long-living [CFAP17]. long-term [UD10]. longevity [NCWK18]. longitudinal [Han12, Man16, vHJPB^{+}17]. look [Dan17]. Loop [TXLC12, NK15, PCYZ12, WWL^{+}10]. LoopFix [WMW^{+}19]. loops [WMW^{+}19]. Loosely [EZG15, CDOP15]. loosely-coupled [CDOP15]. losing [SLZ12]. loss [BMB19, ZYZL12]. losses [BP15]. LossEstimate [PD12]. Lossless [QZ12, GJ13, HWL13b, TTL10, WCCL10, WLH13]. Low [PMDH13, BCA^{+}19, LKP13]. low-end [LKP13]. low-latency [BCA^{+}19]. lower [Jør16, ST11]. LSB [YWWS10]. LTE [EZOK14]. LTL [PW18, SGK12]. LZW [WYCC13].

m [CDA11]. m-banking [CDA11]. M2M [SDB16]. MAC [Bar15, PZB10, WC11]. machine [AO16, BNS12, EFSJM17, GJ13, KCT12, KCV11, MCV15, PCCK18, SZW^{+}16, TTL10, VTZ^{+}17, XHM^{+}11, XZZ^{+}16, ZFS15, ZWC^{+}19, Zha16]. machines [ABB19, BML^{+}13, FGLI15, LQW^{+}12, PWW10, SK13, WXZ^{+}17, FSS^{+}13, KMWL12, YHM^{+}14]. macro [GAT15]. Mahtab [MN19]. mail [MRJD^{+}12]. main [SLLL12]. main-memory [SLLL12]. mainstream [AMKD13]. maintainability [AAM16, CL15, CREH^{+}18, KR16, SAN^{+}17, YC13, dAGSdFS^{+}15]. maintainable [FCRF16]. maintainers [dL13]. Maintaining [CSS10, GAT15]. maintenance [ACG^{+}15, CW12, DFCPSF15, JLQ^{+}10, KBHG17, MG12, OBS^{+}18, PGRQVV12, SL10]. maintenance-centric [OBS^{+}18]. maintenance-free [DFCPSF15]. major [PWS^{+}15]. make [JLZ^{+}19]. makespan [ZCC^{+}17]. Making [OFR^{+}12, BWP16, BWW^{+}18, DCP12, ETYL15, GLZ15, OWG19, RPT19, SWA^{+}13, vVT16, AKAA18]. malicious [Ala15]. malleability [SC19]. Malware [CRL^{+}12, LZL^{+}18]. man [MCV15]. manage [AKH12, AMCC14, FSGYP17, KTF^{+}16, YHMS16]. Management [TKSRP11, ADTZ12, ASMN15, ADET12, BCS18, CJT^{+}16, CS19, CLY17, CBC^{+}15, CS12, CDPM17, DS16a, DIB14, EB14a, EH19, EGG^{+}11, EBJ17, FOR19, GGS^{+}19, GGB19, GSdS16, HNS12, HM16, HCL^{+}10, HL11, JG14, KLP10, KGG18, KH10, LJL^{+}12, LAL15, LH11b, MPTT14, MGB16, MBT16, NKMM12, PCHW12, PPMM12, PvV12, SBGT13, Sta10, TAJ^{+}10, THWC10, TC12, TTMI19, Uzz13, WDC12, YAY13, ZDC^{+}11, ZFY^{+}19, ZMK12]. managers [MNS13, RMCH^{+}14]. Managing [BMB18, CHCO11, FKWVH19, MSL12, NCK^{+}15, PCHW12, aSRS^{+}10, VAJ18]. MANET [LJC16]. manifestation [CPRT16]. manifold [LYLC16]. manipulate [LLWL19]. manipulating [MCTM11, MCTM11]. manipulation [OHL17]. manual [TAS^{+}18]. Manufacturing [NCK^{+}15, AHW10, FVHF^{+}15, TTL^{+}13]. many [GZY11, PN14, RRV19, ZZ16]. many-core [PN14]. many-objective [RRV19]. Map [KPT13, CJ13, ÖT18, ZT14]. Map-matched [KPT13]. Mapping [KBM18, ASGJ13, AJG^{+}15, AB16, ADCO18, AM18, APS16, ACS13, AAC^{+}17, AS16, BCFP19, BBND^{+}18, BLTY18, BDM^{+}19, CS19, CNMR18, DLM19, ESM^{+}19b, FSGYP17, GMMGP15, GRR16, dGFDL16, HBP^{+}17, JCYT16, KBJZ15, KSIZ19, Kit10, KQ17, LAL15, dPLV19, MM14, MRY17, MAEL19, MD16, NVPGMPSM17, PXT^{+}13, PMB15, RHL^{+}17, TAF^{+}17, WNC17, WRdMSN^{+}13, YLA16b, ZSG16, ZGYS^{+}15]. mapreduce [KC16, BGTC18, TLK16b]. maps [BAAS13, DEA^{+}14, KOS15, LWC13, OD17]. MARBLE [PCCLdGP12]. March [WZM12a]. margin [KBM18]. market [TY18]. marketplace [Jør14, KBRV18]. Markov [WFY^{+}19, AMAY19, LCL15, WCC13]. MAS [GCC^{+}15, dVRB13]. MAS-ML [GCC^{+}15]. mashup [LLX^{+}11]. Masquerade [RZMPM12, XTZX12, XTZX13]. massive [LWZ^{+}16]. massively [Dan17]. matched [KPT13]. matching [BKLE18, CJL11, KOL^{+}14, MPN^{+}17, WWLG13, YJZ17]. materialized [GLWY10]. mating [KHSD10]. matrices [NBA^{+}17]. matrix [TK14]. Matthew [YAY13]. maturing [FMRM15]. maturity [BRS^{+}18, CS19, FFdRG^{+}14, GTF15, GTF17, KKA^{+}19, NBF^{+}19]. Maude [ABS19]. Maven [RvDV17]. maximization [YXP^{+}18]. maximize [PAR14]. maximizing [KHSD10]. maximum [LGS^{+}19, WFY^{+}19]. may [AS10, EED16]. maze [LLC10]. MCPs [CD10]. McTorrent [HSS10]. MD4 [ZL12a]. MDD [PFF12]. MDE [DV10, GDFFPG^{+}10, HZH^{+}16, HBOS13]. me [BDDS11]. mean [TTL10]. Means [LKJR10a, LKJR10b, CCGG14, FSGYP17, KRDH12, KHW19, KM14]. Means-ends [LKJR10a, LKJR10b]. Measure [DDD14, BLLGSMB11, CCMOM19, GD12, WL10]. Measure-independent [DDD14]. Measurement [PW10, Al 12, DAG19, EAH^{+}11, EbAT13, GPMI13, GTF15, KMK16, LAT10, QGZ^{+}15, SM16]. measurement-modeling [DAG19]. measures [BDO11, BAM17, CdCAdO18, FMdAR16, KOL^{+}14, MSGGL12]. Measuring [BPSK18, DNSH13, HJBH10, HWHT11, JST10, LS17a, PDS19, WLL15, ZLT10, DPVvV19, nPHW^{+}16, PEO11, TPGdS13]. mechanism [BDBLP15, CL13, KKG^{+}12, KBRV18, LQLC16, LZG15, SM17a, TY18, WCX15, YL16, YGN^{+}16, ZDC^{+}11, ZSB19]. mechanisms [CYT16, KBRV17, SdSLS^{+}19, SC19]. median [LCLF13]. mediating [JC10, KP10]. Mediation [BDBLP15]. Medical [UUN11, AQK11, AMK12, FM11, HTH13, KSRD10, KRDH12, LDZL15]. medium [VA17, dSdMSNO^{+}14]. medium-sized [dSdMSNO^{+}14]. meet [TSL^{+}11]. meet-in-the-middle [TSL^{+}11]. meeting [SSD16]. Meetings [BJK^{+}11, SKK^{+}18b]. members [JLY14]. Memetic [FAM15]. Memoriam [TG10]. Memory [CCN^{+}10, SS14b, ÁRMC16, BPQP^{+}10, KKLB11, PS13, RMCH^{+}14, RSCB18, Shi17, SLLL12, WLZ^{+}17b]. memory-corruption [WLZ^{+}17b]. memory-efficient [Shi17]. merge [HCB^{+}16]. mergence [ZHH^{+}17]. merging [DEW^{+}16, MKL^{+}15]. mesh [BMAH11, LZ13, LL14, WMD^{+}10, YCLC17, ZADM10]. MeshFS [YCLC17]. MeSRAM [SM16]. message [AN16, BKRW19, CL18, EEAZ13, JEEL16, SV12]. message-driven [BKRW19]. message-passing [CL18]. messages [BHVR18, KPS10]. messaging [RA16]. Meta [rBHM17, BCR^{+}19, CdL18, GGT^{+}19, KBHG17, AM13]. meta-analysis [GGT^{+}19]. meta-model [BCR^{+}19]. meta-modelling [CdL18]. meta-object [CdL18]. Meta-Protocol [AM13]. Meta-synthesis [rBHM17]. metadata [BPQP^{+}10, GAKF13, VRPT18]. metadata-based [GAKF13]. metaheuristic [HBT16]. Metamodel [MGR^{+}13, KOL^{+}14, KTF^{+}16, SdSLS^{+}19, WKD^{+}19, ZMK12]. metamodel-based [KTF^{+}16]. Metamodel-driven [MGR^{+}13]. metamodels [DRELHE16, HS11a]. metamorphic [JCK^{+}17, TSRC18, XHM^{+}11, CPX16]. metaphorical [MMB10]. metaprogramming [LRO19]. metasearching [AKB11]. Method [CS16, CCGdL16, AKAA18, ATvHJ18, AKL14, dSACdLF17, AS17, AB10, BKSM13, BKSM14, Che13, CSS^{+}13, EB14c, FAI13, HJ12, HC10, HHC12, HTH13, KKLC12, LLC10, LT13, LWSH19, LGH^{+}17, LWBH16, MRBN17, PMDH13, PWW10, PW10, PWC12, RFM10, RSB^{+}16, RBW18, SI12, SCwY12, SKE10, SPLW17, SSP17, SGC^{+}17, SHS16, SBB^{+}16, SZW^{+}16, SM16, Sta14, SNDD19, TVMS18, TB13, TTC18, TC11, WCCL10, WCB^{+}17, XSL^{+}18, YXH^{+}18, YTW^{+}13, YZC15, ZK13, ZLZ11, ZYY^{+}19]. method-based [AKAA18]. Method-level [CS16]. Methodological [BHM12, WV11]. methodologies [ABC^{+}13, DDMP14, DNBM12, ISM11, SDG17, TLK16b]. methodology [DAR14, HGP^{+}12, PN14, RMCH^{+}14, WSJ14, ZÁ15]. Methods [ACW10, Al 12, ABJ10, ANM15, CP15, CBAV16, DC11, HCC10a, KSENM17, KCV^{+}19, KSM^{+}16, MRT17, PG12, TC11, TPKT12, Wie14, ZADA15, ZXTT11]. metric [CRC19, KCAS13, MK15a, dAGSdFS^{+}15, TDW^{+}14, CPX16]. Metrics [PSZ17, ASGJ13, AAM16, AAC^{+}17, CPR13, DMSG11, FBB^{+}12, GPMI13, Kit10, MJF10, NVPGMPSM17, PDS19, PFO^{+}19, dAGSdFS^{+}15, ZXL10, LGM^{+}18]. MFCVQ [YWHL11]. MFCVQ-based [YWHL11]. micro [FMR11, LVVTP17]. micro-optimize [LVVTP17]. micro-structures [FMR11]. microaggregation [LM13, LM15]. microcontrollers [LKP13]. microgrids [AMCC14]. microservices [AIE19, DLM19, LZJ^{+}19]. middle [TSL^{+}11]. middleware [AGBD14, BMLL14, CCN^{+}10, CDRT13, DKP^{+}19, DIB14, HBG^{+}13, HWLM11, KSHC14, LC11, MKS10, MDP^{+}11, RLY^{+}13, RMD11, TLK^{+}16a, FS14b, KLL^{+}11, VSDD12]. migratable [GMMC13]. migrating [FGB^{+}19]. migration [FGLI15, GDLB16, KKA^{+}19, RRM17, YTW^{+}13, rBHM17]. MIH [OZO^{+}14]. MIH-based [OZO^{+}14]. mind [OD17]. mindfulness [BDPRC18]. Miner [LWBH16]. mines [BRG^{+}12]. minimal [CCY11]. minimisation [YH10]. minimization [CTL10, FHY17, LGHR16, MB17, WAG15, ZAY19, ZCC^{+}17]. minimized [PWY^{+}16]. Minimizing [BGLG13]. Minimum [LGS^{+}19, WL17, BCLW11, CHL11, HWL13a]. Minimum-cost [WL17]. Minimum/maximum [LGS^{+}19]. Mining [CHN19b, LJL^{+}12, LLL^{+}17b, MKHLB16, Wu11, ZM18, DS12, GNA17, HSC15, HWL13a, KSAR18, Lin16, LZ12, LW13b, LWBH16, MG11, MdFD^{+}15, NJ17, NDS13, PCCLdGP12, PWA^{+}19, RDPM19, SOS^{+}18, SAH12, SJC13, SLLY17, SYXL17, TLK16b, TPTV17, WLC13a, YF15, ZMB14, ZJL10]. mismatch [MARD16]. miss [BKSM13, BKSM14]. missed [SPMG18]. Missing [IAA16, ZJZ11, SCGL^{+}18, TC16a]. mission [CCN^{+}10]. mission-critical [CCN^{+}10]. mistakes [SCMS15]. misuse [EA12, EA14, KOS15]. misuses [WLZ^{+}17b]. Mitigating [SMZC12, LMT16]. mitigation [WAWO12]. mixed [CSMC19, CAG17, LWSH19, LGHR16, MMTS15, NI13, PGPC17, TB13, XB19a]. mixed-criticality [LGHR16, PGPC17]. mixed-method [LWSH19, TB13]. Mixing [GDFFPG^{+}10]. ML [GCC^{+}15]. MMPP [LJM11]. MMU [CLY14]. MMU-less [CLY14]. Mobiiscape [KLL^{+}11]. Mobile [ASV^{+}16, AAH10, CBS16, CL13, GBCI11, LLHY19, PMMM11, AR12, AN16, ARS10, AAH12b, AAN11, AAM^{+}17, AGBD14, BGS^{+}16, BCF18, BGG10, BJK^{+}11, BND14, BSDD14, BDM^{+}19, CDA11, CdCAdO18, CTL12, CMK^{+}11, CWK10, CRKH11, Chr16, CGPT14, DKP^{+}19, DIB14, FRGC10, GRBNA10, GNA17, GFP11, IB11, JED18, KAU16, KSHC14, LL14, LZHS11, LKK14, MDP^{+}11, MT13, NOPF12, PLVB^{+}18, PPN^{+}15, PLHP^{+}15, RHHT18, SM17a, SRWE10, SHN14, SKE10, SHBC19, TG17, TKH^{+}11, UIK17, VSS^{+}11, VSDD12, VPL^{+}10, WCB^{+}17, YSDT11, YH13, YL16, YGN^{+}16, ZSG16, ZSB19, FS14b]. mobile-cloud [DKP^{+}19]. mobile-health [LZHS11]. mobiles [GCSÁddP11]. Mobility [BCEF10, AN10, BD10, KLL^{+}11, LH11a, MEB^{+}10, ME10, WB10]. mobility-enabled [AN10]. Model [AHH^{+}10, DPP^{+}18, EBEL18, HVK11, MAR^{+}19, MKL^{+}15, MBPM19, OB13, PMR16, PBD^{+}12, SDB16, WKD^{+}19, AdB13, Aki18, AK16, AdAD17, AF16, AAB19, BRB14, BSK^{+}18, BV15, BCR^{+}19, BGTC18, BCF18, BDDS11, BMB18, BRS^{+}18, BPB19, BL11, CCGdL10, FGB^{+}19, CFAP17, CV14, CHLW17, CD10, CH10b, CCGdL16, CHCO11, DEW^{+}16, DK15b, DGWC16, DCT17, DM17b, EVR11, EUR^{+}13, FDÁM12, FGMM17, FVHF^{+}15, Fei12, FKWVH19, FAI13, GSM19, GMPN16, GMR17, GKS18, GD12, GRT13, GTF15, DDF^{+}13, GEM15, HP16, HA10, HZH^{+}16, HLWS13, JBSL12, JPGdL17, KP10, KBHG17, KLL17, KD18, KB16, KSS15, LJC16, LKR13, LKRYTS18, LS17b, LPM15, LJA^{+}11, LXC13, LLL^{+}14, LPB19, MJ14, MGB16, MAG12, MLD16, MPRS14, MV11, NHC13, NPC12, NB13, PB15, PCHW12, PRS11, PCFRP19, PHR10, PGRQVV12, RAK15]. model [RHHT18, RKK16, RRM17, SAMN12, SÁMI17, SFMB16, ST13, SDG17, SXYW14, SSP17, SS14a, SZW^{+}16, SGEK19, SLLY17, SXYM11, SS13, TS19, TKJL13, TAF^{+}17, TTR^{+}13, TGP11, TSRC18, UIK17, Uzz13, VM12, WKZL10, WDC12, WWSZ15, WTG^{+}15, Wau19, WZM12a, WZM12b, WBS^{+}10, WGS^{+}14, WWSS13, WCC13, XTZX12, XTZX13, XB19a, YFZ^{+}16, YC12, YCF^{+}13, YHM^{+}14, YF15, ZML10, ZM18, dCPV10, nQYD11, RSCB18, WFY^{+}19]. Model-based [DPP^{+}18, OB13, SDB16, AAB19, BRB14, CFAP17, EUR^{+}13, FVHF^{+}15, FKWVH19, GKS18, KSS15, LLL^{+}14]. Model-Driven [PMR16, HVK11, PBD^{+}12, AdB13, Aki18, FGB^{+}19, FDÁM12, FAI13, GMPN16, DDF^{+}13, GEM15, HP16, JPGdL17, MGB16, PGRQVV12, RRM17, SAMN12, TAF^{+}17, TGP11, UIK17, VM12, Wau19, WWSS13]. model-free [WDC12]. model-to-model [CCGdL10]. modeled [MMP15]. Modeling [AAMS14, ABB15, BPGS13, BCF18, CCD19, PG15, SB14, aSRS^{+}10, TAF^{+}17, VPdP13, YLXZ16, AHW10, AGC13, APS^{+}10, BM18, BKH10, BDPRC18, BT17, DMA18, DGRN10, DAG19, EZRK16, FBB15, FCB^{+}16, HGMB13, JC10, KJS^{+}12, KKL^{+}11, LGH^{+}17, MB19, MV10, MGR^{+}13, MNSA15, MNSA16, PPMM17, PAS^{+}10, RTM19, SA14, SÁM^{+}16, SKL10, SJ17, dSSVV11, SK13, SWES16, TB13, TGP11, VRPT18, WSJ14, YAKK16, ZMK12, BBA10]. Modelling [CHN19a, WB15, WMOKY11, AD14, BRS10, CHN19b, Cic16, CdL18, ETYL15, GCC^{+}15, KMK16, MPS^{+}12, MPLL^{+}15, PC10, SB17a, SS15, TTR^{+}13, VKL16]. Models [PS16, APM^{+}14, ASV^{+}16, AHBA19, AMCC14, ADET12, AKA^{+}15, AF16, AGR19, ABJ10, AMGG14, AK15, BAM17, BGLG13, BWH10, BSG^{+}18, CdAM^{+}14, CCG^{+}18, CHN19a, CLS^{+}12, CPD^{+}18, CRK^{+}18, CFM^{+}16, DCG16, DRELHE16, DZT^{+}14, EA12, EA14, EGG^{+}11, FDÁM12, GVPPM18, GGvH^{+}18, GTA14, HGBS18, HJBH10, HBT16, JED18, JK12, Kim12, KLB15, KM13, KTF^{+}16, LWB^{+}13, LHP^{+}09, LHP^{+}10, LHLG^{+}15, MWM12, MSGGL12, MA10, MPAA15, NTT19, NHH^{+}12, NBF^{+}19, PMR16, PN14, PPG^{+}13, PFF12, RSB^{+}14, SPSR17, SNDD19, SGO13, TTL^{+}13, TKCR14, TGE17, WMW12, YAKK16, ZYA^{+}18, LJH10]. Modern [YCA17, VAM^{+}10]. modernization [CRESF^{+}13]. modernizations [SdSLS^{+}19]. modifiability [CFRPC^{+}18, LJH10]. modification [LCLF13]. modified [EEAZ13]. Modular [BRS10, DBL^{+}18, SNDD19]. Modularity [VHFF^{+}17, dB12]. modularization [NMM13]. module [EB14c, KS16]. modules [MTF14, MA17]. modulo [SYT^{+}17]. modulus [CSW13, LC10]. Moitree [DKP^{+}19]. Mojave [BCBZ14]. moment [GJ13, TPKT12, yWpNyL11]. moment-based [TPKT12]. MOMM [MKL^{+}15]. Monetary [AB10]. Monitoring [DFCPSF15, TJT^{+}18, BRG^{+}12, CLY17, CLF^{+}13, ES14, JR15, KKG^{+}12, KLL^{+}11, MLLK11, MB10, OM13, PZ15, RGV^{+}17, SYBN12, SZ11, VRG^{+}16, WWY^{+}12]. monolithic [LZJ^{+}19]. monotone [SD16a]. MOO [dRSBA13]. mortem [AS10]. Mosco [AGBD14]. MostoDE [RHRC13]. MostoDEx [RHRC15]. motifs [WFY^{+}19]. motifs-based [WFY^{+}19]. motivate [VBC^{+}14]. Motivating [LMWM18]. motivational [MPS^{+}12, dSF12]. move [TVMS18]. movement [NCS10]. moving [IBM11, KLL^{+}11, Lin12a, URG10]. MPI [TGKL19]. MPLS [CTHW12, LL10]. MPSoCs [NEM17]. MULAPI [XSL^{+}18]. Multi [ÁRMC16, BLS18, CTL10, DMA18, FMdAR16, GRT13, HM16, LAT10, LyWSZ10, MS17b, MK15b, PSS^{+}16, SRWE10, ACL13, ÁGBYB^{+}14, BV15, BPO^{+}16, BM17, BCS18, CLL10, CW12, CW14, CYT16, CKC15, CAG17, CNKL12, CV16b, FHL^{+}15, FFV19, GMPN16, GCC^{+}15, GZS^{+}18, GGM11, HCB^{+}16, HSM16, JLY14, JXLC15, JS16, KBJZ15, KM11, Kim12, KSH^{+}12, KAM13, KS16, LHJ10, LKP13, LS14, LXG10, LQC^{+}14, LZG15, MLHL12, MVLJ18, MMZ^{+}16, MGB16, MIBV14, NCW^{+}19, OKS^{+}15, PB15, PCHW12, PK10b, PGPC17, PWA^{+}19, PHBJ16, Sal17, SPTM15, SCO13, SHS16, SA18, SZS13, TKP^{+}18, TLL12, WVT^{+}14, WDC12, WCX15, WX10, YFZ^{+}16, YCF^{+}13, YCC16, YLXZ16, YH10, ZMB14, ZJZ^{+}17, ZCC^{+}17, ZP17, fLSN18, AKAA18, CD10, MKL^{+}15, ACSC16]. multi-agent [BM17, CNKL12, GMPN16, GCC^{+}15, GGM11, JS16, MIBV14, SPTM15, ZMB14]. multi-attribute [BV15, KAM13]. multi-byte [Kim12]. multi-category [YFZ^{+}16]. multi-channel [MLHL12]. Multi-cloud [MS17b, CAG17, WCX15]. multi-collinear [LXG10]. multi-component [BCS18]. multi-core [CYT16, CKC15, FHL^{+}15, KSH^{+}12, LS14, PGPC17, WX10, ZCC^{+}17, fLSN18, CD10]. Multi-criteria [FMdAR16, MK15b, PB15, AKAA18]. multi-dimensional [LQC^{+}14]. Multi-faceted [LAT10]. multi-GPU [HCB^{+}16]. multi-homing [HSM16]. multi-hop [CW12, JXLC15]. multi-item [CLL10]. multi-keyword [GZS^{+}18]. Multi-layer [CTL10, SRWE10]. multi-layered [BPO^{+}16, MGB16]. Multi-level [BLS18, HM16, PK10b, Sal17, ZJZ^{+}17]. multi-members [JLY14]. multi-method [SHS16]. multi-model [PCHW12]. Multi-objective [ÁRMC16, PSS^{+}16, CV16b, FFV19, KS16, LHJ10, NCW^{+}19, OKS^{+}15, PWA^{+}19, YH10, MKL^{+}15]. Multi-paradigm [DMA18]. multi-partite [ÁGBYB^{+}14]. Multi-party [LyWSZ10]. multi-precision [LKP13]. multi-purpose [KM11]. multi-rat [MMZ^{+}16]. multi-relational [SZS13]. multi-release [YLXZ16, ZP17]. multi-RSU [ACL13, ACSC16]. multi-secret [CW14]. multi-server [TLL12]. Multi-sprint [GRT13]. multi-step-ahead [YCF^{+}13]. multi-target [SA18]. multi-tenancy [KBJZ15]. multi-tenant [LZG15, MVLJ18, PHBJ16, WVT^{+}14]. multi-tier [WDC12]. multi-vendor [SCO13]. multi-vocal [TKP^{+}18]. multiagent [VAM^{+}10]. multibit [KPS10]. multicast [JXLC15, LZ13, LL14, MV10, ZYZL12, ZADM10]. multicluster [ZLD13]. multicore [HH17, LFCL12, OB13, TC16b]. multicriteria [SL10]. multidimensional [ZXTT11]. multilayer [NHC13]. multimedia [GFP11, GPL^{+}15, LLLK10, YWTW11]. multiobjective [PLHP^{+}15, dCPV10]. multiplayer [CMK^{+}11, Dan17]. Multiple [KSM^{+}16, ACL13, BAI^{+}14, CHL11, GZS^{+}18, HCL12, HWL13a, HSS10, JKC19, NSM17, PPG^{+}13, SC19, SK10, WL15b, WH15, ZWF^{+}18, dRSBA13, vHAT13]. multiple-base [HCL12]. Multiple-case [KSM^{+}16, vHAT13]. multiple-GPU [BAI^{+}14]. multiplexed [MSAH16]. multipliers [WCB^{+}17]. multiprocessor [BLS18, CBL^{+}15, LESL11, LCLS16, WXZ^{+}17]. multiprocessors [dOCS13]. multithreaded [AR17, TLZ^{+}16]. multivariate [AdAD17]. multiwavelet [PWW10]. municipalities [KKA^{+}19]. music [LHH10]. mutant [DPP^{+}18, MHLMG14]. mutants [MK16]. Mutation [CNM18, MK16, DW14, HBT16, LHJ10, dPLV19, PFO^{+}19]. Mutomvo [CNM18]. mutual [IB11, LZ12]. my [BDDS11, KMG^{+}19].

narratives [BS12]. native [KQ17, PDK^{+}16, YGN^{+}16]. natural [BKS15, CAHV15, KMWL12, SA14]. navigation [BPGS13, LK13]. near [BKSM13, BKSM14, CL17a]. near-miss [BKSM13, BKSM14]. near-optimal [CL17a]. Nearest [Zha12b, Cho13, HKS^{+}17, LZ12]. Nebo [EMBS17]. need [JCK^{+}17]. needles [dOFB^{+}19]. needs [GGT^{+}19, MSSMDC12]. negative [CPYZ14]. neglect [OBS^{+}18]. neighbor [Cho13, HKS^{+}17, LZ12, Zha12b]. nested [RQD^{+}17]. nested-virtualization [RQD^{+}17]. net [AHW10, FYCL13]. nets [BHM12, OH15, PPMM12]. Network [AO16, GMGTdFR14, JXLC15, SB19, BKLE18, BRG^{+}12, CCH14, CBZ^{+}16, CL15, CL17b, DY15, DRCG12, DAR14, DCT17, HY11, HST16, HB13, JLY14, KY10, KR16, KSHC14, LH11a, LG17, LL10, LLH^{+}16, MB19, MK15a, MJZ^{+}10, NJ17, OK11, OZO^{+}14, SMZC12, Ski13, TJH15, TAF^{+}17, TSCB19, TAB^{+}16, TCCH12, TT10, YCLC17, YLYL17, MMTS15]. Network-aware [AO16]. network-based [BKLE18, BRG^{+}12]. networked [ADMOK^{+}10]. networking [PSM12, SCGL^{+}18, YAT11]. Networks [ACSC16, AAMS16, AMAY19, ACL13, AC16, BMAH11, Bar15, BCLW11, BND14, CCdR^{+}16, CBS16, CSW10, CWK10, CTHW12, CW12, Cho13, CL13, DBL^{+}18, DBCdP11, EEAZ13, ECRVMS11, HBG^{+}13, HST15, HWHT11, HSS10, JXLC15, KLP10, KCV11, LS17a, LCC10, LT11, LZ13, LL14, LWOY16, LWL^{+}16, LWC^{+}18, LMA15, LKK14, MLHL12, MCC^{+}18, MLLK11, MMZ^{+}16, MLD16, MC10, MDO^{+}10, MT10, MKRO14, MARD16, MAAC17, NSAK10, NNVD17, OZO^{+}14, OH15, PZB10, PD12, RNC14, SM17a, SHN14, SMS11, SGBCP12, SZS13, SHH^{+}15, WMD^{+}10, WMOKY11, WCC13, WOC15, XZP^{+}10, YH19, ZADM10, ZYY^{+}19, AAJD^{+}16, CDRT13, DFCPSF15, GMGTdFR14, SXYW14]. Neural [LMA15, DBL^{+}18, DRCG12, EEAZ13, KCV11, KR16, ZYY^{+}19]. newcomer [TTC18]. news [KP10, LQC^{+}14, TPTV17, WKbOS17]. next [CRK^{+}18, PS15, TLK^{+}16a, VRPT18, VPMVM^{+}13, DRCA^{+}19]. NFV [KMK17]. NN [Zha12b]. No [SBAH17]. NoCs [MV10, MV11]. node [CLY14, TLK^{+}16a]. Node.js [KTK19]. Node/ [TLK^{+}16a]. nodes [BK11, GAT15, MKRO14]. noise [CKS15, DEA^{+}14]. Noisy [LZ12]. Non [ABB15, ZCT^{+}11, ACH19, CTKT13, CHN19b, MPAA15, MJZ^{+}10, NSDI16, PN14, PGRQVV12, ZLD13, ZL17]. non-boolean [CHN19b]. non-coherent [PN14]. non-dedicated [ZLD13]. non-formal [CTKT13]. Non-Functional [ABB15]. non-linear [ACH19]. Non-parametric [ZCT^{+}11, MPAA15]. non-redundant [PGRQVV12]. non-stationary [MJZ^{+}10, ZL17]. noncontiguous [BMAH11]. nonlinear [GSN^{+}15]. nonparametric [SD16a]. NoSQL [DII^{+}17, DK15b]. Nosv [RQD^{+}17]. Notation [AHBA19]. notational [HCL12]. notations [HRD10, OFR^{+}12]. note [Ano11m, Ano17m]. notes [ZXG10]. Novel [CNSG12, HBT16, KSRD10, CNL13, CH10a, CBZ^{+}16, DS16a, GSN^{+}15, HLLS13, KRJ17, LC10, LHH10, LXCM11, LGH^{+}17, LWXZ10, LNW^{+}11, LNY^{+}11, LWW^{+}10, LLC17, LY18, MCC^{+}18, MRBN17, OY16, PZB10, ST13, SSP^{+}15, TVMS18, jT12, TT13, TTT14, WGZ^{+}12, ZGL^{+}10]. novice [CCP18, CDS19]. nucleus [HHC12]. null [CBSM16]. NUMA [CYT16]. number [MIUM12, SYT^{+}17]. numbering [Ano19l]. numbers [ANC11]. Numerical [LJ16]. numerically [EMBS17].

O [FTC16, SMZC12, SC19]. O1FS [PNY14]. object [Al 12, BPSK18, CZC^{+}18, CL15, CL17b, CdL18, CCMOM19, EVR11, EB14c, FBB^{+}12, FTSC12, KCAS13, KCR16, KS19, KR16, MJ14, Mer13, OAC11, OB13, SKL10, SH17, SSSA17, SSS17, VTZ^{+}17, YLC18, ZXL10, dAGSdFS^{+}15]. object-oriented [Al 12, BPSK18, CZC^{+}18, CL15, CL17b, CCMOM19, EVR11, EB14c, FBB^{+}12, FTSC12, KCAS13, KS19, KR16, MJ14, Mer13, OAC11, OB13, SKL10, SSSA17, SSS17, YLC18, ZXL10, dAGSdFS^{+}15]. objective [ÁRMC16, CV16b, DRCA^{+}19, FFV19, KS16, LHJ10, LY18, NCW^{+}19, OKS^{+}15, PSS^{+}16, PWA^{+}19, RRV19, YH10, MKL^{+}15]. objectives [dRSBA13]. objects [CRC19, IBM11, KLL^{+}11, Lin12a, SJ17]. Oblivious [MXZ11]. Obscured [DM17b]. observation [CV16a, WHY^{+}12]. Observational [YBE17]. Observations [CBT^{+}14, KWS^{+}17]. Obstacles [DCP12, GSdS16]. obtaining [CHL^{+}13]. OCCI [MBT16]. OCL [CCR14, KBHG17, OT17]. Octopus [BSG12]. ODC [CPR13]. off [AHC^{+}11, CFMRL11]. off-the-shelf-based [AHC^{+}11]. Offloading [CCL^{+}19, AR12, ASV^{+}16, DSGS17, RHHT18, YGN^{+}16]. offs [SPCT18]. Offshore [SWA^{+}13, KNA11]. OLAP [SGW^{+}15]. OMG [BCF18, HBG^{+}13]. omnipresent [AHH^{+}10]. Omniscient [BLC^{+}18]. On-demand [HST15, DR12, HST16]. on-line [DWC17, FHY17]. One [BMS11, AAN11, KM13, LW13a, MT10, ZL12a]. one-block [ZL12a]. one-level [MT10]. One-time [BMS11, LW13a]. one-to-one [AAN11]. Online [SGEK19, VPL^{+}10, CL17a, Dan17, GSM15, KH10, MCS^{+}12, YCWW15]. only [HRB12]. onto [AO16]. ontologies [HS11a, LPP^{+}10, RHRC13, ZLT10]. Ontology [MCS^{+}12, YSG17, AACT13, BLLGSMB11, KSAR18, MJF10]. Ontology-based [YSG17, MJF10]. Open [CdL18, GPPT16, ACB18, ALRP16, BCG^{+}14, DFCPSF15, ESM19a, EB14b, GDLB16, GW10, HBR19, JBSL12, KTF15, KKT17, KHMA12, KKA^{+}19, KK17b, LRD^{+}19, NPC12, PAB^{+}17, PPS12, RA16, RNR17, SA12, SG12, VGSN18, WFF18, YLXZ16, ZFY^{+}19, CFMRL11, GL14, LMWM18, LLS11, MP12, Shi12]. open-source [RA16, WFF18]. OpenFlow [CCdR^{+}16]. Opening [JBSL12]. OpenMP [DSGS17, NEM17]. OpenPGP [MBB11]. OpenStack [ZFY^{+}19]. Operating [ESRF19, GPPT16, HK13, SRT^{+}12]. operation [LWB^{+}13, Lin14, WGZ^{+}12]. operational [OD10, OKMD12]. operations [DZT^{+}14, KD18, PWY^{+}16]. Operator [ILZ14, CV16b, HWR17]. Opinion [GNA17, TPTV17]. opportunistic [BCLW11, NSAK10, SCGL^{+}18]. Opportunities [SBAH17, AZX14, ACW10, BDO11, CDPM17, LAH^{+}16, Oja16b, TVMS18, TC10, TC11]. Optical [AT18]. Optimal [LM13, WXY^{+}17, AM10a, CL17a, DDD14]. optimisation [GA13, PACH15, RRV19, WRTP^{+}13]. Optimization [Pot13, ADMOK^{+}10, ALRP16, ÁRMC16, BLM10, BZ14, BAI^{+}14, CPYZ14, CAG17, CV16b, ELHC13, GRT13, GCSÁddP11, KHSD10, KAM13, LSE12, LZCL19, LLHY19, LLZW14, LCL^{+}12, MCL^{+}17, MdOBW^{+}15, MBAG11, MAG12, MRJD^{+}12, PS15, PRN17, RCCVB11, RGH17, San16, Ski13, SGO13, SWES16, TJH15, TXLC12, TDW^{+}14, XJZ^{+}15, YTW^{+}13, ZYZZ14, Zha16, dCPV10, dRSBA13, AZ11]. optimize [AN16, AKL14, LVVTP17, MAS13, RMCH^{+}14]. Optimized [DHC^{+}11, DRCG12, GWW^{+}11, KCV11, YF15, ZDC^{+}11]. Optimizing [LQW^{+}12, QOLJG16, BM18, ÇT13]. oracle [JCK^{+}17, KAS18]. oracles [CL18, PW18, RG10, ZTPT18]. orchestrated [ABC^{+}13]. orchestrations [TTC15, ZTCZ16]. Order [LPP15, LHJ10, dPLV19, MCKA18, NBF^{+}19, PMDH13, nQYD11, WCC13, ZJZ^{+}17]. ordered [JHYK10, MLD16, YZY^{+}18]. Ordering [ZA12, PS13]. organisational [WK15]. organisations [YMM^{+}17, YMM^{+}19]. organization [Bos12, JBSL12]. Organizational [ISM11, SG12, Woh16, JMML17, MMB10, WKbOS17, WRR14, WSM15, ESM19a]. organizations [ASG17, AK16, BdMSNO^{+}17, BCG^{+}14, KK11, KFN19, PPG^{+}10, SST16, SM16]. organizing [BM17, GAKF13, HM16, XLM^{+}15]. Oriented [BBEM11, EMSU11, MAR^{+}19, Al 12, AM15, ARS10, AK15, BPSK18, CCG^{+}10, CZC^{+}18, CL15, CL17b, CHL^{+}13, CGPT14, CCMOM19, EVR11, EB14c, FBB^{+}12, FTSC12, GMMC13, ISM11, JLQ^{+}10, KCAS13, KS19, KKH^{+}16, KR16, LMN10, LMGHB17, LLL^{+}14, LN13, MJF10, MJ14, MTF14, Mer13, MPS^{+}12, MPLL^{+}15, MGvFGCB10, NFSM11, NBR^{+}13, OAC11, OB13, PNJGF12, PFF12, Pot13, PHBJ16, SCS15, SGP12, SKL10, SdSLS^{+}19, SSSA17, SSS17, SWES16, TKK^{+}19, THWC10, WXY^{+}17, WZM12a, WZM12b, YLC18, ZWM^{+}18, ZXL10, dVRB13, dAGSdFS^{+}15]. origins [BWW^{+}18]. OSS [BWP16, KTK19, ZCC^{+}19]. other [SC14]. our [WLL17]. outage [DM17b]. outcomes [CBAV16, FMRM15]. outlook [DFG^{+}13]. output [KAS18, SRT^{+}12, SED16]. outsource [SYT^{+}17]. outsourced [DvdVA^{+}13, ZML17]. outsourcing [AV12, AK16, BWP16, DPVvV19, Jør14, KNA11, ZHAY12]. over-fitting [WQJZ10]. Overcoming [Che17]. Overflow [CCD19]. overlapped [MK16]. overlay [DY15, MARD16]. overload [JEEL16]. overview [CBT^{+}14, EGM^{+}11]. owners [GZS^{+}18].

P [LJDK10]. P/S [LJDK10]. P/S-CoM [LJDK10]. P2P [SSP^{+}15, LLW12, OK11, SSP^{+}15, ZXTT11]. P2P-based [ZXTT11]. Package [ESRF19, AAM16]. packages [AKKS11, DYC19, JS11]. packed [LZL^{+}18]. Packet [GFP11, BP15, FGBC10]. pages [DH13]. pains [EZG15]. pair [CRSS14]. pair-programming [CRSS14]. paired [LWBH16]. pairing [RZL^{+}18, Shi10]. Pairwise [LPP15]. palette [CPL13].

Palladio [FSPH^{+}16, TKCR14]. palm [WLL^{+}13]. PAM [TKSRP11]. paper [TZ12]. Papers [KT16, BCL^{+}18, Bor12, CL11, LH12]. paradigm [DMA18]. Parallel [JWT17, AHW10, BAI^{+}14, EMBS17, GE15a, GTY12, LZY^{+}15, MCC11, PDBD18, RG10, TS19, TGKL19, TLK16b, WWYZ11]. parallelism [CBL^{+}15, OWB11]. parallelization [LAH^{+}16, MDBC17, NEM17]. parallelize [XPBC11]. parameterized [NTT19]. parameters [KLB15, MAG12, PG15]. parametric [GVPPM18, LGS^{+}19, MPAA15, ZCT^{+}11]. Pareto [NCW^{+}19, YH10]. parity [CSS10]. Parkinson [GMPN16]. PARS [TS19]. Parsed [EVR11]. parsing [AACT13]. part [LKJR10a, LKJR10b, BKW10]. partial [CHL11, EMBS17, KVT^{+}17, MCKA18]. Partially [SBT19, HRB12, JHYK10]. partially-ordered [JHYK10]. participant [AL10]. participating [CH10a]. participatory [CRKH11, Chr16, DEA^{+}14]. particle [LLZW14, dCPV10, AZ11]. partite [ÁGBYB^{+}14]. partitioned [XB19a]. partitioning [CH10d, KSENM17, MCC11, YZL^{+}14]. partnership [AK16]. party [AHC^{+}11, LyWSZ10, YC12]. passing [CL18]. password [CTL12, HCC10b, WZM12a, WZM12b, YC12]. password-authenticated [WZM12a, WZM12b]. path [GP10b, MK15a]. paths [GZY11, GTY12, KSS15, LWLL12, LWBH16]. patients [GMPN16]. pattern [BKLE18, FM11, HZ15, HP16, HK13, JCYT16, KPS10, KY10, KLL^{+}11, KLL17, LNY^{+}11, LW13b, NKZ17, SOS^{+}18, WLC13a, YCF^{+}13, YZC15, YZY^{+}18, ZFS15, ZLmLN14]. pattern-based [HP16, KY10, KLL17]. pattern-driven [HK13]. Patterns [ABJ^{+}17, FHL^{+}18, AKKS11, ACS13, BZ10, CSF^{+}14, CCG^{+}10, CHL11, Cic16, CRESF^{+}13, FVHF^{+}15, FMR11, GGM11, HSC15, HA10, HJ12, HWL13a, HHK13, JLGM17, KA18, KKR16, KVT^{+}17, LKRYTS18, LJL^{+}12, LLX^{+}11, MRY17, MSK^{+}17, MKHLB16, RAJ15, SCS15, Sal17, ŠK11, SB17b, SJC13, SVM19, Sta10, VPL^{+}10, WCC^{+}14, YZC15, ZTZ^{+}11]. patterns-based [HSC15]. pave [WLL17]. payload [FF12]. payment [YL16]. PCS [WCC13]. PDE [OLZN13]. Peak [ACRD19]. Peer [BCG^{+}14, KSHC14, LHH10, LKK14, MLD16, OK11, SHN14, SS13, YH13]. peer-to-peer [KSHC14, LHH10, LKK14, MLD16, YH13]. people [AKH12, FFdRG^{+}14, HWdS^{+}15, MPLL^{+}15]. people-oriented [MPLL^{+}15]. Per-flow [AM10a]. Perceived [OK18, CFRPC^{+}18, ETM10, OGK13, VLL18]. percentage [LLK11]. perceptions [GW10, LLS11]. perceptron [NHC13]. perceptual [MK11]. Percolation [Pal12]. Percolation-based [Pal12]. perfect [LLC10]. perfecting [RHM^{+}18]. performability [EBJ17, EK13]. Performance [AAMS16, AMAY19, AAH10, AAH12b, BMAH11, BZ14, DZT^{+}14, GLJ13, LJM11, MK17, MNM12, NSAK10, PK10a, SPCT18, TPKT12, Zha16, AdB13, AHLH16, ATvHJ18, AL10, BML^{+}13, BHM12, BDPRC18, BRS^{+}18, BT17, BK17, CSW10, CLL10, CBZ^{+}16, CS12, DWC17, FTC16, HCY19, HZH^{+}16, KBDGAW16, KA18, KA14, KCV11, LZCL19, LZR16, NK14, PCHW12, PH13, Pot13, QOLJG16, RHHT18, RLY^{+}13, RQD^{+}17, RVCM17, ROFGFRM13, SPC16, SCwY12, Shi17, SA11, SSP^{+}15, SS13, TBC^{+}16, TKCR14, TDW^{+}14, VSDD12, WYCC13, WMD^{+}10, ZHH^{+}17, dL13, ADMOK^{+}10, OSH^{+}18]. performance-directed [SPC16]. Performing [CTKT13]. period [FHL^{+}15]. Periodic [HLW^{+}15, CHL11, HSC15, HyLW^{+}12, KPS10, KKR16, KVT^{+}17, TKJL13]. periodic-frequent [KKR16, KVT^{+}17]. Periphery [LLK11]. permutation [HRB12]. Perpetual [Fei12]. persistence [SPCT18, TGP11]. persona [ARH^{+}17]. personal [BSG12]. personalities [KBDGAW16]. personality [ARH^{+}17]. personalization [FMPS16]. personalized [AM10b, ECRVMS11, MOD^{+}19]. personalizing [SAA^{+}10]. personnel [GA11, PPN^{+}15]. perspective [BDM^{+}19, EED16, EUR^{+}13, HM16, JCYT16, KBJZ15, LS17b, LMWM18, LZHS11, LDS^{+}19, SGP12, SAR15, SOS^{+}16, Som13, GRR16]. perspectives [AKH12, CSMC19, NAB^{+}13, SW19, YLCZ12]. Perturbation [LXJL10, LCC^{+}13]. Perturbation-based [LXJL10]. perturbations [XYZ^{+}19]. pervasive [AHH^{+}10, BSG12, CMK^{+}11, KAK^{+}13]. Petri [AHW10, BHM12, FYCL13, OH15, PPMM12]. Petri-nets [OH15]. PF [LWBH16]. PF-Miner [LWBH16]. Phase [MN19, HJ12, LZC14]. Phase-wise [MN19]. phases [APS16]. PHash [Shi17]. Phi [DSGS17, GGK19]. PHP [DGBE18]. physical [AWSE19, CSMC19, GVPPM18, GBH^{+}16, GSP^{+}19, LS17b, LDS^{+}19, NI13, ZAY19]. physical-task [NI13]. Pig [SAH12]. Pillar [BRG^{+}12]. pipeline [KAS18]. pipelined [WMOKY11]. pivoting [PS14]. pixel [HCL12, YWTW11]. pixel-value [YWTW11]. PL [FGB^{+}19]. PL/SQL [FGB^{+}19]. placement [DHC^{+}11, OCC12, XJZ^{+}15]. plagiarism [TLZ^{+}16]. plain [LW13a]. plan [GLWY10]. planning [ABL16, BMLL14, BRS^{+}18, BK17, FCB^{+}16, GRT13, LHCT19, MH12, Ski13, WAWO12, WCC13]. plans [Pot13]. plasticity [dVRB13]. Platform [BV18, vHJPB^{+}17, AKL14, APS^{+}10, AM10b, BD17, CdAM^{+}14, CdR^{+}14, CMM15, CDPM17, DAG19, HS15, HWdS^{+}15, MOD^{+}19, MIBV14, NBR^{+}14, PAS^{+}10, QOLJG16, RM19a, RA16, ZLD13, vAAJ16]. Platforms [SKT17, FHL^{+}15, PGPC17, ZCC^{+}17]. played [WLL17]. player [MQG^{+}17]. playing [Dan17]. playout [FGBC10]. PLC [VHFF^{+}17]. PLC-based [VHFF^{+}17]. PMIPv6 [CL13]. point [BGLG13, EAH^{+}11, MJZ^{+}10, WL10]. points [GAT15, ZZ16, SSP17]. Poisson [AC16]. poker [MH12]. policies [FBB15, PCCB^{+}11, SW10]. policy [BLS18, BL19, FSGW11, RMD11, ST11, TKSRP11]. policy-based [RMD11, TKSRP11]. polymorphism [TC10]. popularity [CFMRL11]. port [BKH10, MV10]. portability [TLK^{+}16a]. portable [EMBS17, CDS10]. portals [HYA11]. portfolio [CBC^{+}15, URG10]. positioning [CSW13]. Positive [SKK^{+}18b]. post [SS12, AS10]. post-adoptive [SS12]. Post-mortem [AS10]. postcamera [Lin14]. posteriori [LWB^{+}13]. posting [MCC11]. posture [SHBC19]. potential [KWS^{+}17, VT14]. Power [BLM10, TKJ15, wZfG13, wZfG14a, wZfG14b, ASV^{+}16, APS^{+}10, CWK10, CW12, CRK^{+}18, EBJ17, KKHH11, MAS13, PAS^{+}10, RITF^{+}11, TC12, VA17, VAJ18, WDC12]. Power-aware [TKJ15, wZfG13, wZfG14a, wZfG14b]. Practical [CSM15, SKK^{+}18a, AMS^{+}10, FGB^{+}19, FG15, LWBH16, MMTS15, SOS^{+}16]. Practice [BKW10, CJT^{+}16, BGS^{+}16, BDPRC18, GTA14, Han12, KTF^{+}16, LKRYTS18, LWSH19, NBF^{+}19, RZL^{+}18, SB14, VCB^{+}18, Wie14]. Practices [BV18, BD16, BV16, BCG^{+}13, CCP18, CWJK13, ETM10, GV10, GZ13, GCBCD15, GCDY16, IBAH12, JDLS16, KFN19, KT12, MHB18, NBF16, OK18, PFL16, SW19, WWSS13, ZADA15]. Practitioner [LLS11, GCDY16, LMNA17, NBF^{+}19]. practitioners [AHC^{+}11, CCP18, PV18]. Pragmatic [GVPPM18]. Praspel [DGBE18]. precedence [AR18]. precious [vV10]. precision [LKP13]. predicate [aSRZ^{+}18, WL15b]. predicates [DOL^{+}16, GLOM19]. predict [AAM16, KY10, MER17, NHH^{+}12, OY16, RBS19, WRS^{+}17, XYCL17, ZXL10]. Predictable [ICSK14]. Predicting [ACB18, ABL16, CPV^{+}14, SD16a, ZcKS17, AdAD17, GXZ^{+}19, KR16, LMA15, VGSN18]. prediction [Al 12, AGC13, AC16, ABJ10, CPD^{+}18, CBAV16, CMM15, CSS^{+}13, FF12, FSS^{+}13, HJBH10, HC10, JLZ^{+}19, KY10, KCV11, LCT10, LCLF13, LG15, LJA^{+}11, MdFD^{+}15, NCW^{+}19, ÖT18, PEO11, PSM12, PB15, PPB19, RSB^{+}16, SLW^{+}15, SHBA^{+}16, TAB^{+}16, ULS19, VTZ^{+}17, WFY^{+}19, XLW18, XLX^{+}19, XLL^{+}19, YCLY13, YCF^{+}13, YJZ17, ZCY^{+}16, ZL17, dCPV10]. Predictive [PJT^{+}17, CS15, RSB^{+}14]. predictor [OLZN13]. preemption [Kim17]. preemptive [FSPH^{+}16]. Preface [MS17a, SLR16]. Preference [FFV19]. preferences [MLD16, SPLW17]. preferences-based [MLD16]. preferTrust [MLD16]. Prefetching [MJ18, LZCL19]. Prefetching-aware [MJ18]. prefix [ND18]. preimage [ZL12a]. preliminary [Kit10]. premier [LCM^{+}13]. premise [AAMS14]. preparation [SAH12]. presence [LJM11, PS15, PJK13, SMZC12]. PRESENT [ZGZ^{+}13]. presentation [ZLZ11]. preservation [LCLF13, ZLmLN14]. Preserving [AAH12a, MCV16, BKSM13, BKSM14, BJK^{+}11, DEA^{+}14, HL11, Lin16]. preventing [WS12]. prevention [Aba13, BRG^{+}12, KHC16, WAWO12]. price [LZL^{+}15]. pricing [AB10, LZO^{+}16, Oja16a]. Prime [CG15]. primitive [LCLP16]. primitives [HZ15]. Principles [LMGHB17, BGS^{+}16, BPSK18, GDFFPG^{+}10, KFN19, ZMK12]. print [KPS10, PKS18]. print-cam [PKS18]. print-scan [KPS10]. prioritization [AWSE19, CZC^{+}18, DvdVA^{+}13, HMOK18, HCC10a, HPH12, HLLS13, HCT^{+}15, JC15, LZKW12, MCTM11, MKS^{+}18, MB17, PSS^{+}16, PSEE12, PMB15, PWA^{+}19, SPLW17, SB12, SCC16, SD16b]. prioritize [WZY^{+}18]. Prioritized [ZS16, PD16]. Priority [HYA11, AKA^{+}15, FHL^{+}15, FSPH^{+}16, KSN17, Kim17, LCLS16, LZ13, RXY^{+}19, wZfG14b, dOCS13]. priority-aware [LZ13]. PRISMA [ARS10]. Privacy [Chr16, DEA^{+}14, SY16b, AGBD14, CDS10, CRKH11, ECRVMS11, Lin16, MXZ11, MIKG13, SLZ12, SGBCP12, TKH^{+}11, WSJ14, YYS^{+}16, BJK^{+}11]. privacy-aware [AGBD14]. privacy-enhanced [TKH^{+}11]. privacy-focused [WSJ14]. Privacy-preserving [DEA^{+}14, Lin16, BJK^{+}11]. private [CHL^{+}13, GPSS^{+}13]. privilege [HBM19]. Pro [BPB19]. Pro-IDTV [BPB19]. Proactive [DM17a, SKK^{+}18b]. probabilistic [AMP12, DC11, DK15a, HN17, NTT19, PACH15, RXY^{+}19, SGO13]. probability [MSGGL12, RCCVB11, YXP^{+}18]. problem [EK12, KK17b, MJ14, MARD16, PS15, RSBA19, SS15, XJZ^{+}15, ZJZ^{+}17, ZGL^{+}10, CKL12, DRCA^{+}19]. problem-solving [KK17b]. problems [AR17, JK12, KSENM17, LCL^{+}12, SYBN12, TTR^{+}13, VLL18, VHL14, YF15]. Procedural [WV11]. procedure [AK15, BKSM13, BKSM14]. procedure-oriented [AK15]. procedures [Ski13]. Process [HSPD14, MMB10, AKH12, AAMS14, AAGT16, AHLH16, APW14, AAN11, AF16, AMGG14, BWW^{+}18, BBA10, BGLG13, BPB19, CNG16, CNKL12, FDÁM12, GMMGP15, GDLB16, HVK11, HPF16, HBOS13, IBM11, JED18, JAS19, JMML17, KKT17, KSKP11, KSAR18, KTF^{+}16, LPM15, LMR12, LMGHB17, LZKW12, LRD^{+}19, LMNA17, MdOBW^{+}15, MSGGL12, MAAC17, OFR^{+}12, PB11, PCCLdGP12, PCFRP19, PW10, PPG^{+}10, RDPM19, ŠK11, SS14a, SWA^{+}13, SZW^{+}16, SK18, TAF^{+}17, TTC15, Uzz13, VLC^{+}17, VKL16, VVA^{+}15, Wau19, WMW12, XLW18, ZADA15, ZL17]. process-centered [KSKP11]. process-line [BWW^{+}18]. Processes [FFdRG^{+}14, AHW10, AC16, AM10a, BRS^{+}18, CFRPC^{+}18, CXO^{+}15, CLF^{+}13, FSG^{+}11, JST10, JR15, PRS11, RRM17, SMZC12, YLXZ16]. processing [CM12, DM17a, DWC17, FS19, FGD^{+}17, HWR17, KAS18, LCC10, LPP^{+}10, Lin12a, RVCM17, SHN14, SMM17, TS19]. processor [TC12]. Processors [CD10, TXLC12, WWL^{+}10]. produced [RM19b]. producing [BV16, JBSL12, MPAA15]. Product [CBT^{+}14, ESWA18, AC19, AKL14, AWSE19, BKS15, BHM12, BBD18, BdMSNO^{+}17, BBS10, CHN19a, CHN19b, CdSdSG^{+}18, CFAP17, CNKL12, CV16b, DGRN10, DV10, EB14a, ESM^{+}19b, FFV19, GMMGP15, GWW^{+}11, HGBS18, HJN11, HPF16, JG14, JKL19, KCV^{+}19, LMN10, LNTS19, LGS^{+}19, MNS13, MCV16, MAGC^{+}17, MD16, NBA^{+}17, NBA^{+}15, OGRJ^{+}18, OH15, PLHP^{+}15, PBD^{+}12, RTM19, ROR11, SBT19, SSS17, SdSGdMSN^{+}13, SSAS11, TBG13, URG10, UD10, UIK17, WAG15, WGS^{+}14, WR10, WBBK18, YMM^{+}17, YMM^{+}19, dSdMSNO^{+}14, dOSdAdSG17]. product-form [BHM12, OH15]. product-line [LNTS19, UIK17]. production [FKWVH19, VHFST15, VHFF^{+}17]. productivity [BMB19, RSGH12]. products [HBR19, RHL^{+}17]. profile [Cic16, OCC13, RZMPM12]. profiles [BK17, PC10]. Profiling [Ala15, KMK17, LWLL12, TZ12, TC12, WLZ^{+}17a, WLL19a, WLL19b]. profit [GCMB17]. program [BHH^{+}10, CS16, DW14, EK12, JCK^{+}17, JRO12, LLL17a, MM19, MJ19, SZ11, aSRZ^{+}18, XST18, YLYL17]. Programmable [CBC14, AYZI10]. programmers [CDS19]. programmes [LLM^{+}17]. Programming [BSDD14, MAR^{+}19, AR17, BCFP19, BDG13, BSB12, CdAM^{+}14, CCR14, CCG^{+}10, CAG17, CRSS14, GE15a, KBDGAW16, LHJ10, MNO18, NBR^{+}13, PN14, PTF^{+}15, RBS19, RAJ15, SGP12, SPCT18, YLC18]. programming-level [GE15a]. programs [ASdMGM14, ABS19, dSACdLF17, CL18, DOL^{+}16, ECS15, ES14, EED16, GPMI13, LMH10, LAH^{+}16, MNM12, PJK13, SÁMI17, SM16, aSRZ^{+}18, TKJ16, TLZ^{+}16, TGKL19, YWWS10, dSF12]. progress [HH17]. Progressive [HHH10b, YCWW15, JHYK10, FMRM15]. project [ASMN15, ACB18, APCS10, CS19, CBAV16, CC11, GL14, GGC16, HM16, JSL16, Jør16, JR15, LS17a, LMIV15, LSD^{+}16, NBF16, PCCK18, RSBA19, RKK16, SSF15, SAR15, Sta10, VLL18, WK15, YAY13, YFZ^{+}16]. projects [AS10, AAH12a, AdB17, ÇB16, CFMRL11, DvdVA^{+}13, GC13, IF19, JKC19, Jør14, KP10, LHCT19, LMA15, LMNA17, MSB18, MAH18, MVSG18, PD16, RSB^{+}14, RSB^{+}16, RCCVB11, SNDC13, SHHL12, TNK^{+}19, Uzz13, VGSN18, WKbOS17, ZCC^{+}19, dOSdAdSG17]. Prolog [AR17]. ProMARTES [TAB^{+}16]. promised [HS11b]. promote [GMMGP15]. Promoting [GRBNA10, dVRB13]. prone [MA17, ZXL10]. proneness [FP18, HJBH10, MRS18, ZCC^{+}19]. proof [LMGHB17]. proof-of-concept [LMGHB17]. Propagating [WMW12]. propagation [ATvHJ18, DRELHE16, MRS18, MRRS19]. properties [DNSH13, HBG^{+}13, OMLB16, PH13, PJT^{+}17, PDBD18, WWY^{+}12]. Property [ZLG10, KU10, ZLmLN14]. proposal [FCMJ12, SLLL14]. proprietary [KHMA12]. prospective [Gar13]. prospects [FGD^{+}17, GTF15, KKA^{+}19]. protected [LLLK12, RF14]. protection [GMGTdFR14, JEEL16, KJ10, SY16b, TLL13]. protections [BCR^{+}19]. protocol [AN16, CSW10, CTHW12, CW12, DMSG11, MDO^{+}10, MT10, OHJ10, PZB10, PPS12, SC14, WZM12a, WZM12b, WC11, XZP^{+}10, YC12, ZG10, AM13, GMGTdFR14, ST11]. protocols [BCS18, CdL18, DAR14, KKLC12, SSP^{+}15, YSL^{+}10]. prototype [RO13b]. prototype-based [RO13b]. prototyping [CH10c]. Provably [LH11a, YC12, ZG10]. provide [ECRVMS11, LLW12, WL17]. provider [CWJK13]. providers [MIKG13]. Providing [HH17, MCV15, CX10]. province [GV10]. provisioning [KAK^{+}13, MPLL18, SB19, THWC10, WZJI14, ZDC^{+}11]. PROW [LPP15]. Proxy [TLK^{+}16a, FSGW11, HNS12, SLZ12, SV12, SXYM11, WHY^{+}12, CL13]. pruning [WQJZ10]. PS [CDRT13]. PS-QUASAR [CDRT13]. Pseudo [JC10]. PSO [MA17, TLL13]. PSO-GA [MA17]. PSP [RZL^{+}18]. public [CMK^{+}11, RPSL10, YYS^{+}16]. public-key [RPSL10]. publish [CDRT13, HBG^{+}13, LJC16, LVPMPCLS13, RMD11, LJDK10]. publish/subscribe [CDRT13, HBG^{+}13, LJC16, LVPMPCLS13, RMD11, LJDK10]. Publisher [Ano11m, Ano17m]. Publishing [VGM13]. pull [JLZ^{+}19, LMPM18]. pure [OK11]. purpose [KL10, KM11]. pursuit [RHM^{+}18]. PVD [YWWS10]. Pycots [BDLM16].

QA [JSHW14]. QoS [BMLL14, BVV^{+}10, CV16a, CDRT13, DHC^{+}11, EGG^{+}11, HBG^{+}14, KAM13, Li11, LG15, LLWL14, LGZ^{+}18, MLHL12, MV11, PPMM12, PPMM14, PPMM17, PG15, SWES16, WTG^{+}15, YZG^{+}13, ZADM10]. QoS-aware [BVV^{+}10, CV16a, DHC^{+}11, MV11, YZG^{+}13]. QoS-based [LLWL14]. QoS-oriented [SWES16]. QR [LQLC16]. QSIC [CL11]. quad [LBCL10]. Qualitative [San16, GTF17, PV18, RM19b]. qualities [PSZ17]. Quality [AJG^{+}15, AS16, CLH^{+}13, ELHC13, FGBC10, GA13, MVLJ18, MVSG18, WNC17, AHOP14, AAC^{+}17, BDD^{+}15, BLLGSMB11, BAM17, ÇB16, CFMRL11, CdCAdO18, FOR19, FFWE17, FG15, GMMGP15, HBG^{+}13, HJN11, HNH15, HPF16, HWHT11, HTH13, HKS^{+}17, JLGM17, KGB11, KS19, LAT10, LJA^{+}11, LWZ^{+}16, LSD^{+}16, LHCT19, LMNA17, OWG19, OGK13, PCYZ12, SKF17, VSS^{+}11, WKH11, ZTCZ16, ZGYS^{+}15, vdBSvS^{+}19, YDGB^{+}12]. Quality-adaptive [CLH^{+}13]. Quality-driven [ELHC13, LWZ^{+}16]. QualityScan [WOC15]. quantify [EED16]. Quantifying [ACG^{+}15, HFE10]. quantitative [AdAD17, CSF^{+}14, CGS19, GTF15, GTF17, KGG18, MGvFGCB10]. Quantitatively [nPHW^{+}16]. Quantum [AR18, LyWSZ10]. QUASAR [CDRT13]. Quasi [WMWZ12, CBL^{+}15, KKH^{+}16, MWM12, MRT17]. quasi-deadlines [CBL^{+}15]. Quasi-static [WMWZ12]. quasi-synchronous [KKH^{+}16]. quasi-systematic [MWM12, MRT17]. quaternion [yWpWyYpN13]. queries [Cho13, GSN^{+}15, IBM11, JHYK10, LCC10, MMP15, SED16, ZJL10]. query [ACL13, CH11, CJL11, DII^{+}17, GLWY10, ILZ13, LPP^{+}10, LWXZ10, MCL^{+}17, Pra18, RVCM17, SHN14, TLWS10, ZWM^{+}18]. Querying [ILZ14, CNG16, MIUM12]. question [PMWC12]. queueing [OH15]. queues [KSN17]. queuing [MB19]. QuickFuzz [GCMB17]. QVT [KLL17].

R [PSZ17, PDBD18]. R-SHT [PDBD18]. R-Tree [PDBD18]. races [TKJ16]. Radigost [MIBV14]. Radio [AAJD^{+}16, AMAY19]. RAID [YTW^{+}13]. RAID-structured [YTW^{+}13]. raising [LCT10]. RAMCloud [LLGZ13]. Random [CKMT10, MGM10, WS12, YCG^{+}14, CT11b, CLH^{+}13, CZC^{+}18, GP10b, GLW13, HZH^{+}16, MRBN17, RG10, YWEL^{+}13, FSS^{+}13]. randomized [JC15]. range [Cho13, GSN^{+}15]. ranging [MAAC17]. rank [AKB11]. ranked [GZS^{+}18]. ranking [DH13, SED16]. rankings [RM19b]. ranks [AN10]. Rapid [FFWE17, TGBF17, CCG^{+}10]. rat [MMZ^{+}16]. Rate [VPL^{+}10, FBD^{+}18, PMDH13, PDS19]. rate-dependent [FBD^{+}18]. rates [DW11]. ratings [PQBP16, XWZC14]. rationale [Xia13]. rationalize [vHAT13]. ray [BAI^{+}14]. RCDA [PvV12]. RDF [RHRC15]. RDMA [RLY^{+}13]. RDMA-based [RLY^{+}13]. RDOTE [VGM13]. Re [CRESF^{+}13, SV19, FSGW11, NCS10, SLZ12, WHY^{+}12, GRR16]. re-encryption [FSGW11, SLZ12, WHY^{+}12]. Re-engineering [CRESF^{+}13]. Re-implementing [SV19]. re-location [NCS10]. react [RMD11]. reactive [MNSA16, SÁM^{+}16, MNSA15, SAMN12, SÁMI17]. reactor [KJ10]. read [DZT^{+}14]. readers [WL17]. ready [OSH^{+}18]. Real [CLF^{+}13, DYC19, GBC16, LKK14, MK11, AMP12, ABB19, ACL13, ACH19, BLS18, BL19, CPS11, CLL10, CZG^{+}15, CBL^{+}15, CS12, CF12, DOL^{+}16, EBEL18, EK13, FHL^{+}15, FHY17, GLZ15, GPPT16, HyLW^{+}12, HCB^{+}16, HZG^{+}12, HNS12, ICSK14, KC16, KLB15, LESL11, LSE12, LS14, LS17b, LFCL12, LWL^{+}13, LC11, LHP^{+}09, LHP^{+}10, LGS^{+}19, LPB19, MBD13, MFMCY12, NPC12, PG15, RFM10, RXY^{+}19, RGH17, SPC16, SMS11, SK10, TKJL13, TKJ15, TC16b, VZT17, WMWZ12, WX10, YLCZ12, wZfG13, wZfG14a, wZfG14b, ZW15, ZHGL11, ABCH13]. Real-time [CLF^{+}13, DYC19, GBC16, LKK14, MK11, AMP12, ABB19, ACL13, BLS18, BL19, CPS11, CLL10, CZG^{+}15, CBL^{+}15, CS12, CF12, EBEL18, EK13, FHL^{+}15, FHY17, GLZ15, GPPT16, HyLW^{+}12, HCB^{+}16, HZG^{+}12, HNS12, ICSK14, KC16, LESL11, LSE12, LS14, LS17b, LFCL12, LWL^{+}13, LC11, LHP^{+}09, LHP^{+}10, LGS^{+}19, MBD13, MFMCY12, NPC12, PG15, RXY^{+}19, RGH17, SK10, TKJL13, TKJ15, TC16b, WMWZ12, WX10, wZfG13, wZfG14a, wZfG14b, ZW15, ZHGL11, ABCH13]. real-valued [KLB15]. reality [IF19, NI13, VSS^{+}11]. realization [hChSyCwL10]. really [RDPM19]. realtime [WFY^{+}19]. reasoning [CCGdL16, OT17, TBSvdW18]. reasons [CBVF19]. reassembling [LZL^{+}18]. Reassessing [KP10]. receiver [MXZ11]. recently [HHKWB16]. recently-evolving [HHKWB16]. recently-introduced [HHKWB16]. rechargeable [LWOY16, LWL^{+}16, LWC^{+}18]. recognition [CCWT13, HHC12, WLL^{+}13, ZLmLN14]. recommend [GJ16, dOFB^{+}19]. recommendation [GJ16, GMR17, HSL14, LQC^{+}14, MCS^{+}12, NKZ17, SZW^{+}16, SYXL17, XWZC14, XSL^{+}18, ZCY^{+}16]. Recommender [SHH^{+}15, CCY11, LK16, NTdSX13, TZ12, YSG17, YH13, GMLSF^{+}15]. Recommending [BCBZ14]. Reconciling [AKH12, HNZ17, MWM12]. reconfigurable [CFN10, GHBD^{+}16, KRD16]. reconfiguration [BJG11, BBD18, BDLM16, DS16b, Li11, LG15, LJDK10, PDL^{+}16]. reconfiguring [PLHP^{+}15]. reconstruction [BAI^{+}14]. record [dNPM18]. recoverable [LNW^{+}11]. Recovering [QBO^{+}14, RCPZ19, SSS17, KMG^{+}19]. recovery [CKS15, LMS11, LL10, MAEL19, PNY14, SV12, YZC15, ZYZL12]. recursion [LHY12]. Recursive [ZL17]. Redirection [LL10]. redistributed [LXCM11]. reduce [CYT16, EA14, PFO^{+}19]. reduced [LGLL12, TSL^{+}11]. reduced-round [LGLL12, TSL^{+}11]. Reducing [SSMvD16, CWK^{+}11, JRSN10]. reduction [MK16, MGM16, MCKA18, SRS15]. redundant [PGRQVV12]. reengineering [FGB^{+}19]. refactored [CFRPC^{+}18]. Refactoring [YM13, Al 12, AMdLM17, BDO11, BDD^{+}15, KS19, MGM16, MSK^{+}17, MCKA18, PCFRP19, SGMHJ13, SAN^{+}17, TVMS18, TC10, TC11, VM13]. refactorings [CFM^{+}16, FTSC12, dOFB^{+}19]. Reference [ZMK12, AF16, GAKF13, KSKP11, NFSM11, PPG^{+}13, WWLG13]. refined [EBC10]. refinement [APT^{+}12, ILZ13, KHW19, TZ12]. Refining [SDG17]. reflective [LC11]. register [TXLC12]. registration [AAMS14]. registries [SBGT13]. regression [AAB19, HPH12, JK12, JKL19, LXG10, LQLW12, LNTS19, MA10, MN19, NHC13, RB16, SD16b, SSP17, TTC18, WXY^{+}17, YLCZ12]. regression-based [TTC18]. regulations [HL11]. regulatory [MOH16]. reinforcement [FMPS16]. rejuvenation [ACW10, OD10, SW10, SPTM15]. rekeying [SA11]. related [CPD^{+}18, JK12, LRB^{+}19, MS16, SCL13, TLZ^{+}16, WCC13, ZM18]. relatedness [LBX12]. Relation [CPX16, BCD^{+}18, HSL14, vdRBSvV10]. relational [BGTC18, BL11, CDOP15, JK13, LKL^{+}11, MLGA11, SZS13, VGM13]. relations [HN17, TSRC18]. relationship [BDD^{+}15, CTKT13, FHL^{+}15, IBAH12, OBS^{+}18, VLL18, ZCC^{+}19]. relationships [CHN19b, GD12, GMGTdFR14, JLZ^{+}19, LLL17a, MER17, PPMM14, PSZ17, VAJ18, vAAJ16]. Release [DRCA^{+}19, MXZ11, PS15, YLXZ16, ZP17]. releases [AT18]. Relevance [NAB^{+}13, TTR^{+}13, FMR11]. relevant [NBH19, TTC15]. Reliability [AT18, LWL^{+}13, LGHR16, MBAG11, ZCC^{+}17, AGC13, CL15, FHY17, HWLM11, KHSD10, KLB15, LZR16, LYC14, MAG12, MPRS14, PEO11, PB15, Pot13, RAS14, RSB^{+}14, SD16a, VHL14, WZY^{+}18, WFY^{+}19, WRdMSN^{+}13, YTW^{+}13, YLXZ16, ZZP15, ZP17]. Reliability-driven [MBAG11]. reliable [DS16b, FYCL13, GGS^{+}19, SJC13, ZYZL12]. remedy [WS13]. reminder [TBSvdW18]. ReMinds [VRG^{+}16]. remote [HSL14, IB11, YSL^{+}10]. rendering [KA14, SNDD19]. Rendex [AS17]. reordering [TXLC12]. repackaging [KTK19]. repair [JCK^{+}17, MM19, MJ19, WMW^{+}19, ZM18]. Repeating [SB17b]. replacement [BHVR18]. replanning [GRT13]. replay [WXZ^{+}17]. replica [DHC^{+}11]. replicated [EBC10, GV10, RZL^{+}18, SHN14]. Replication [MJ18, ACB18, BDPRC18, BMB19, CdS18, HSC15, OCC12, Zha16]. report [SAH12, SAKZ15, WCC12, WKbOS17, WKV11, WB15]. Reported [ASMN15]. reporting [KP10, OKMD12]. reports [HCY19, LYLC16, ULS19]. repositories [LPM15, SAH12, SGMHJ13, TLA18]. Repository [BV18, CBC^{+}15, RvDV17]. repository-based [CBC^{+}15]. representation [SB17b, XLX^{+}19]. representative [CSM15, LTK^{+}15, OSH^{+}18]. representing [SCS15, XLM^{+}15]. Reproducing [HCY19]. reputation [KB16, TTC18]. request [CLL10, JLZ^{+}19, JH10]. requests [CdCMdMSNdA16, DR12, HYA11, KK11, LHG^{+}18, LMPM18]. required [ABL16, FSGYP17]. Requirement [MD16, PLGT10]. Requirement-based [PLGT10]. Requirement-driven [MD16].

Requirements [ABB15, AHBA19, CNMR18, LZLC17, WPL^{+}18, AVGM19, AS17, AGR19, BKS15, CKL12, CRESF^{+}13, CFA^{+}19, DvdVA^{+}13, Dan17, EGM^{+}11, EUR^{+}13, FSG^{+}11, GSM15, GKV14, HBR19, HJP15, JC10, KMWL12, LKJR10a, LKJR10b, LHG^{+}18, MPTT14, MVSG18, MFM10, MPLL^{+}15, MIKG13, OK18, OWG19, PG12, PD16, PMB15, RDPM19, RO13a, RM19b, RHM^{+}18, SCMS15, SA14, SJR^{+}11, SPLW17, dSSVV11, SMK^{+}18, SRBT18, TKP^{+}18, UGFK15, VVA^{+}15, VCMG17, WLD16, XYCL17, YFT^{+}15, ZTCZ16, ZHGL11, dSdMSNO^{+}14, DDMP14, FFWE17]. Research [ACS13, BKW10, Ano13a, AS16, BP13, BPSK18, CBT^{+}14, DDMP14, DFG^{+}13, JDLS16, KGB11, LCM^{+}13, Man16, dONTF^{+}19, RPT19, TTMI19, VHFST15, Wie14, KS19, VCB^{+}18]. researcher [HCY19]. ReSeer [WXZ^{+}17]. Resemblance [ZHH^{+}17]. reservation [ZWC^{+}19]. reside [GXZ^{+}19]. residence [GXZ^{+}19]. resilience [PDL^{+}16]. resilient [KPS10, MMSD13, YKC^{+}12, YLZ^{+}16]. resistant [HCC10b]. resolution [DBCdP11, DK15a, KHC16]. resolving [MKS^{+}18]. Resource [AD14, CDPM17, KK11, ACRD19, AK15, BV15, BK17, CLY17, CYT16, DM17b, ES14, GGS^{+}19, GGB19, GHBD^{+}16, GWW^{+}11, HNH15, HLW^{+}15, HLWS13, MCC^{+}18, MAS13, NEM17, NK15, SWES16, TY18, THWC10, WDC12, WAWO12, ZWC^{+}19, fLSN18, vV10]. resource-restricted [NEM17]. resources [AHW10, JSL16, MSAH16, SCO13, Sko14]. response [EGG^{+}11, MMTS15]. response-time [EGG^{+}11, MMTS15]. responsibility [KP10, MJ14]. REST [AK15, CPDM16]. restricted [NEM17]. Results [JKC19, LL15, MRT17, APT^{+}12, BPSK18, JDLS16, LGLL12, TGE17]. retargeted [CWK^{+}11]. Retest [LNTS19]. retrieval [GPL^{+}15, KCV^{+}19, MCC11, nQYD11, ST13]. Retrospective [Gar13]. retrospectives [LMIV15]. Reusability [PAB^{+}17, AKKS11, GMGTdFR14, PDS19, SOS^{+}18]. Reusable [RBT11, KTT^{+}17, LMN10, NOPF12, SSSA17, SGC^{+}17, SHS16]. Reuse [SS17, dAK18, BKS15, BV16, EL10, ESRF19, ICSK14, KCAS13, LOFA17, MB17, OAC11, PDS19, PK10b]. ReuseTool [OAC11]. Reusing [FB18, SJ17, MAH18, MSGM17, SBB^{+}16]. Revealing [GGM11]. revenue [Oja16a]. Reverse [MAGC^{+}17, SSSA17, BCR^{+}19, LHLG^{+}15]. Reversible [CSS^{+}13, FF12, OLZN13, AMK12, CNL13, CT11a, HC10, HTH13, JK13, LCT10, LCLF13, LBCL10, Lin12b, Lin16, LTW16, MM14, PWLL13, TK14, WYCC13, WLC13b, WOLS12, YWHL11, YCLY13]. Reversibly [MKH^{+}12]. Review [CVGP13, GCAH18, SKT17, UB19, AAGT16, AKAA18, AVGM19, APW14, ALRP16, BWP16, BKS15, BMB18, CFL^{+}18, CP15, DPL16, DBCG14, GJ16, GNA17, GA11, IHA16, JED18, KGB11, KNA11, LFW15, LL15, LZO^{+}13, MWM12, MH13, MRT17, MGAN18, OGRJ^{+}18, ÖT18, PG12, PFG13, PMB15, PFO^{+}19, RAK15, RCL14, RSBA19, SNL16, SRJL^{+}18, STA19, SLB14, TJT^{+}18, TTM13, TL14, TPKT12, TCS18, VLC^{+}17, VCMG17, WPL^{+}18, WLL17, ZADA15, Zha16]. Reviewers [vV10]. Reviewing [AHOP14]. reviews [AS17, GNA17, PLVB^{+}18]. revisited [HYWS11]. Revisiting [LRB^{+}19, Man16]. reward [TKJL13]. reward-based [TKJL13]. rewriting [GLWY10]. REX [CM12]. RFID [Aba13, AYZI10, CPS11, KSKP11, SLLL12, WL17]. RFM [HHK13]. RGB [SNM14]. RIA [CRESF^{+}13]. RIAs [CRESF^{+}13]. rich [BCF18]. rich-client [BCF18]. riddle [KFLS18]. ridge [LXG10]. right [BDDS11, FGMM17]. rights [KLP10]. RISC [LKP13]. risk [AV12, AD14, BP13, BRS10, CMM15, CLF^{+}13, GKS18, HFE10, Kel15, PvV12, RO13a, WZY^{+}18, YFT^{+}15]. risk- [PvV12]. risk-averse [Kel15]. risk-based [YFT^{+}15]. risk-driven [GKS18]. risks [AB10, SL10]. RO [Jua10]. road [DII^{+}17]. Roadmap [BD10, FS17, ME10, SST16, WB10]. roadmapping [SSAS11]. roads [MT13]. ROAR [SWES16]. RoboCup [KHMF13]. robot [CCG^{+}10, DYC19, MOD^{+}19, ESRF19]. Robotic [MOD^{+}19, AB16, BDM^{+}19]. robots [YSDT11]. Robust [BSKL10, DS16b, LSR13, TK14, TTL10, AMP12, ATvHJ18, CCG^{+}18, KLP10, LWZ12, LLML13, PS15, PKS18, yWpNyL11, yWpWyYpN13, YSL^{+}10]. Robustness [XYZ^{+}19, MFMCY12, SM16]. role [Bis13, Dan17, FBB15, FM11, KP10, KKL^{+}11, LHCT19, MRM16, SA12, SKK^{+}18b]. role-based [FBB15, KKL^{+}11]. role-playing [Dan17]. roles [JMML17, MNS13, MPS^{+}12, WLL17]. Rolling [HZG^{+}12]. Rolling-horizon [HZG^{+}12]. ROS [DYC19]. rostering [PPN^{+}15]. rough [Wu11]. round [LGLL12, TSLL11, TSL^{+}11]. routed [MV10, MV11]. ROUTER [GGB19]. routing [BCLW11, CSW10, CWK10, CW12, DBCdP11, JXLC15, MDO^{+}10, MT10, NNVD17, Pal12, YH19]. RSA [BBBP13, CWK^{+}13, KKHH11, ZM12]. RSA-based [ZM12]. RSU [ACL13, ACSC16]. Rule [VKL16, NBR^{+}13, PWA^{+}19, QLBS17, ROFGFRM13]. Rule- [VKL16]. rule-based [ROFGFRM13]. rules [CCdR^{+}16, PS14, SDB18, TLGE18]. Run [SHBA^{+}16, fLSN18]. Run-based [SHBA^{+}16]. running [Li11]. runs [LZY^{+}15]. Runtime [KGG18, ASV^{+}16, ADET12, LRO19, NTT19, OM13, PJT^{+}17, QOLJG16, RGV^{+}17, SB17a, SHC^{+}11, VRG^{+}16, WLL19a, WLL19b, YGN^{+}16, dRSBA13]. Rust [KTK19]. Rust/Node.js/WebAssembly [KTK19].

S [ÇT13]. S-CoM [LJDK10]. S-IDE

[ÇT13]. SAAD [PÁC13]. SaaS [HS15, MVLJ18, Wu11]. Sacbe [GCSSDP^{+}18]. Safe [BDLM16, NBA^{+}15, TBG13, AAB19, BAAD17, LJDK10, SBT19]. Safety [BDM^{+}19, KHC16, LKJR10a, LKJR10b, RO13a, SGC^{+}17, VCMG17, WGKW19]. Sakai [LWZ12]. SALSA [BVV^{+}10]. SAM [HCB^{+}16]. same [CPZF19]. samples [RHRC15]. Sampling [QXYL16]. Sampling-based [QXYL16]. SAN [SSF15]. SAND [LLH^{+}16]. satisfaction [EK12, PRS11, VLL18, vdRBSvV10]. saving [CWK10, LZC14, YZG^{+}13]. SBSE [HC15, PMB15]. SBVR [SDB18]. scalability [PHBJ16, YC11]. Scalable [CCH14, FS19, JPGdL17, LKL^{+}11, LQC^{+}14, Luk11, MSAH16, ND18, DK15b, FTC16, KLL^{+}11, PN14, PPMM17, SST16, YC11, CSS10]. scale [APS16, BLL^{+}18, ÇB16, CSM15, DvdVA^{+}13, DPL16, EH19, JSM10, JK12, KLL^{+}11, LTK^{+}15, LLL^{+}14, nPHW^{+}16, PFG13, PTF^{+}15, SAH12, Shi17, SXYW14, SVM19, SAN^{+}17, TPTV17, TTMI19, WFF18, WB15, XWZC14, YAY13, ZK13]. scale-free [YAY13]. scaled [KCR16, LWOY16, TS19]. scaled-out [KCR16]. Scaling [KGW12, AR17, CS12, DVV^{+}16, KAS18, LCL15, TSCB19, Wie14]. scan [KPS10]. ScapeGoat [GHBD^{+}16]. SCC [KMK17]. scenario [DK15b, HRD10, SCMS15]. scenario-based [SCMS15]. scenarios [BRS10, JS13, KCV11, MSHG18, SSF15, dMCR19]. Scented [GPD^{+}19]. schedulability [BL19, FBD^{+}18, Kim17, LS14]. scheduler [AR18, FSPH^{+}16]. schedulers [HN17, LFCL12]. Scheduling [LZL^{+}15, LZY^{+}15, SKT17, SK10, ZLD13, ABB19, ALRP16, BLS18, BJK^{+}11, CSMC19, CLL10, CZG^{+}15, CYT16, CKC15, CBL^{+}15, DVV^{+}16, DR12, DFJ19, FHL^{+}15, FHY17, FGBC10, GBC16, HyLW^{+}12, HZG^{+}12, HYA11, HH17, KC16, KSN17, Kim17, KCV11, LESL11, LS14, LCLS16, LS17b, LWC^{+}18, LJM11, MMZ^{+}16, MCKA18, MK15b, PK10a, RFM10, RXY^{+}19, RSBA19, ROFGFRM13, SRS15, SBZ^{+}17, SLW^{+}15, TKJL13, TKJ15, TdCAF16, TC16b, WWL^{+}10, WMWZ12, WX10, WC11, WCB^{+}17, wZfG13, wZfG14a, wZfG14b, ZW15, ZCC^{+}17, ZHGL11, ZGSH13, dOCS13]. schema [KSKP11, NTRN11]. schemas [OT17]. scheme [BMS11, BCL^{+}18, CBS16, CCLL11, CNL13, CH10a, CT11a, CW14, FWCS12, GJ13, HSPD14, HWL13b, HCC10b, IB11, KKLB11, LC10, LSR13, LHZX12, LH11b, Lin12b, LWC13, LCC^{+}13, LWL^{+}16, LJM11, LW13a, LTW16, LY18, MKS^{+}18, MIUM12, NNVD17, Pen11, RPSL10, Shi10, SGBCP12, SV12, SXYM11, TK14, TLL13, TLL12, UUN11, VHL14, WWYZ11, yWpNyL11, WLH13, WYCC13, WCC^{+}14, WZ11, WKH11, WOLS12, WS13, WOC15, YWTW11, YC11, YCC16, ZM12, ZADM10]. schemes [AQK11, DDD14, DR12, LWC^{+}18, LHYZ12, OD10, PCHW12, WMWZ12, YZG^{+}13, ZT14]. Scholar [Won10]. scholars [CLL14, KLA^{+}19, WTG^{+}11]. science [HG18]. Scientific [Kel15, ALRP16, DFJ19, GE15a, LNW^{+}11, Rya13, SZS13, ZLD13]. scientists [HG18]. SCOOP [MNM12]. Scope [MB17, AKL14]. Scope-aided [MB17]. scopes [CHL^{+}19]. scoping [DFG^{+}13, dSdMSNO^{+}14]. Score [GCSÁddP11]. scores [SA18]. screen [CTL12, EAH^{+}11]. Scrum [RKK16, vWSB13, PPG^{+}10, SRSC16, SBAH17, VvSvV16]. SCRUMIA [vWSB13]. SDH [GMS11]. SDN [MCC^{+}18].

SE [IZ18]. Search

[AWSE19, CVGP13, KOL^{+}14, APT^{+}12, BL11, CCY11, ECRVMS11, GZS^{+}18, HNH15, HG18, JC15, JRSN10, KAU16, LM15, LHLG^{+}15, MFTP18, MCV16, MGM16, MSGM17, ND18, PMDH13, PWA^{+}19, RRV19, RSBA19, SS15, SED16, WHY^{+}12, WAG15, WXZ^{+}17, ZGL^{+}10, HLS^{+}13, HC15]. Search-based [AWSE19, KOL^{+}14, HNH15, LHLG^{+}15, MFTP18, RRV19, RSBA19, WXZ^{+}17]. search-centric [CCY11]. search-order-coding [PMDH13]. searchable [KTT^{+}17, RPSL10]. searching [TBC^{+}16, ZXG10]. Secret [EA11, WS13, CT11b, CLH^{+}13, CW14, EEAZ13, GLW13, HHH10b, LT13, LyWSZ10, LHYZ12, MBB11, UUN11, UUN13, WZ11, WS12, WOLS12, YWEL^{+}13, YC11, YCC16]. Section [BKW10, BFLZ13, SLR16]. Secure [GZS^{+}18, LH11b, ABFM12, CDA11, CCLL11, CW14, CH10b, CL13, EZOK14, GKD13, GRBNA10, IB11, KKHH11, LH11a, LSR13, PSdO^{+}13, RG10, RITF^{+}11, SM17a, SC14, SXYM11, SS13, TLL12, THS12, WLL^{+}13, YC12, ZG10, ZZ12]. securely [SYT^{+}17]. SecureSMS [SC14]. Securing [CPL13, OM13, PDK^{+}16]. Security [CDS10, HRB12, AMKD13, ANG^{+}19, BP13, BSG^{+}18, BL11, DAR14, GMS11, HFE10, KOS15, LDS^{+}19, MIKG13, OLV15, OKMD12, PPS12, PCCB^{+}11, RO13a, RPSL10, Rya13, SZ11, SLZ12, UUN11, WV11, YFT^{+}15, YKC^{+}12]. SEED [KKP12]. Seeing [GW10]. segmentation [HHC12, KSRD10, ST11]. Segmenting [AHLH16, KSRD10]. Selected [LH12, Bor12]. Selecting [RSB^{+}14, LMPM18]. Selection [AHC^{+}11, Jør10, AM10a, BWW^{+}18, CPR13, EFSJM17, GPMI13, GWW^{+}11, HJ12, JS11, JKL19, KNA11, LQLW12, LTK^{+}15, LWZ^{+}16, LNTS19, MK15a, MB17, MIKG13, MAAC17, NCW^{+}19, OZO^{+}14, PB15, PMB15, RAK15, SSP17, TCK14, TC16a, WQJZ10, WGC^{+}14, WCX15, WXY^{+}17, WH15, XLL^{+}19, Zha12b, MGM10]. selective [LW13c]. Self [ABB15, BJG11, BBD18, BM17, CHLW17, EK12, GBH^{+}16, HWR17, JS16, PCYZ12, SRT^{+}12, ARS17, BSK^{+}18, CCdL^{+}16, CV16a, CPYZ14, CG12, DWC17, FCB^{+}16, GSP^{+}19, HGP^{+}12, HM16, KKG^{+}12, LT13, LZR16, MKS^{+}18, MCS^{+}12, MAS13, PCHW12, PPMM12, PDL^{+}16, QXYL16, SB17a, SSK19, SGEK19, TJT^{+}18, WMAS12, YXP^{+}18, CV14]. Self-adaptation [BBD18, GBH^{+}16, JS16, CCdL^{+}16, CG12, FCB^{+}16, GSP^{+}19]. Self-adapting [BJG11, HGP^{+}12]. Self-Adaptive [ABB15, CHLW17, HWR17, ARS17, BSK^{+}18, KKG^{+}12, LZR16, PPMM12, QXYL16, SB17a, TJT^{+}18, WMAS12, YXP^{+}18]. Self-Adjusting [CV14]. self-admitted [MKS^{+}18, SSK19]. self-authentication [LT13]. self-aware [SGEK19]. self-configuration [MAS13]. Self-control [EK12]. self-correcting [CV16a]. self-managing [PCHW12]. self-optimization [CPYZ14]. Self-organizing [BM17, HM16]. self-reconfiguration [PDL^{+}16]. Self-tuning [PCYZ12, SRT^{+}12, DWC17]. selfish [GAT15]. SelfMotion [CGPT14]. Semantic [DH13, MJF10, RvDV17, XLM^{+}15, BDO11, BKSM13, BKSM14, GMLSF^{+}15, GPL^{+}15, KKLC12, LPM15, LBX12, MTF14, RHRC13, ST13, TJH15, EZRK16, KM17, KR14, TTM13, VGM13, ZLT10]. semantic-based [GPL^{+}15]. semantic-preserving [BKSM13, BKSM14]. semantic-web [RHRC13]. semantically [CdR^{+}14]. Semantics [MOD^{+}19, BCF18, Cic16, GKV14, LLLK12, SK18, YBE17, Zha16]. Semantics-based [MOD^{+}19]. Semi [HZ15, BSG^{+}18, CdCMdMSNdA16, GGvH^{+}18, KBHG17, OGRJ^{+}18, PPS12, SPLW17, XB19a]. semi-automated [BSG^{+}18, CdCMdMSNdA16, SPLW17]. Semi-automatic [HZ15, GGvH^{+}18, KBHG17, OGRJ^{+}18, PPS12]. semi-partitioned [XB19a]. semiconductor [AT18]. senior [CC11]. sense [OFR^{+}12, RMD11]. sense-and-react [RMD11]. sensing [CMK^{+}11, CRKH11, Chr16, FF12, HSL14]. sensitive [FSGL12, SG16, WQJZ10, Zha12a]. sensitivity [LWW^{+}10, LTW16]. Sensor [DFCPSF15, AN10, Bar15, BRG^{+}12, BK11, CBS16, CLY14, CLF^{+}13, DBCdP11, HWHT11, HSS10, LCC10, LT11, LLK11, LWOY16, LWL^{+}16, LWC^{+}18, LHP^{+}09, LHP^{+}10, MLLK11, MC10, MT10, MKRO14, NSAK10, NNVD17, SMS11, SGBCP12, TAF^{+}17, YH19, CDRT13]. sensor-based [CLF^{+}13]. Sentiment [ULS19, IZ18, JR15]. SentiStrength [IZ18]. SentiStrength-SE [IZ18]. separate [ADTZ12]. SEProf [TC12]. Sequence [ZLG10, CJ13, CZC^{+}18, HHK13, WLC13a, WGZ^{+}12]. sequences [LK13, MJZ^{+}10, Pra18, YZY^{+}18, ZJC^{+}10]. sequential [HWL13a, HHK13, LAH^{+}16, SJC13]. series [AGC13, LNY^{+}11, SB17b, SKF17]. serious [GSM15]. server [BLM10, hChSyCwL10, EBJ17, HCY19, HWLM11, MAS13, SLLY17, THWC10, TC16b, TLL12]. Service [AM15, CNG16, CBC^{+}15, EMSU11, HBG^{+}14, HS15, AJG^{+}15, AKAA18, APM^{+}14, AM10a, AK15, BBD18, BMLL14, BMKM15, BZ14, BDBLP15, BVV^{+}10, CFN10, CDPM17, CGPT14, DVV^{+}16, DS16a, DYC19, DLW^{+}13, FYCL13, FSG^{+}11, GS17, GCLD13, GMMC13, HBG^{+}13, HWLM11, IYS13, JLQ^{+}10, KMK17, LMN10, LPM15, LQLW12, LLZW14, LLWL14, LDZL15, LVPMPCLS13, LY18, LZG15, LGZ^{+}18, MS17b, MSL12, OLV15, OCC12, PSS11, Pot13, RAS14, RBW18, SW10, SBGT13, aSRS^{+}10, TG17, TSCB19, WVT^{+}14, WCX15, WXY^{+}17, WNC17, WBBK18, WWY^{+}12, WZJI14, XYCL17, YMM^{+}17, YMM^{+}19, ZTCZ16, ZWM^{+}18, ZHGL11, ZHAY12, dVRB13, BBEM11, LWSH19, MPRS14, OLV15, WVT^{+}14, YDGB^{+}12]. service-based [CFN10, LMN10, aSRS^{+}10, WWY^{+}12]. Service-oriented [AM15, CGPT14, GMMC13, JLQ^{+}10, Pot13, WXY^{+}17, dVRB13]. services [AM10a, CCH14, CH10b, GFP11, GPSS^{+}13, KTT^{+}17, LRO19, LKL^{+}11, LZO^{+}13, LLX^{+}11, MGB16, MCTM11, Oja16b, PWS^{+}15, PCG^{+}14, PHBJ16, RHL^{+}17, SFMB16, SKK^{+}18a, SCO13, SKF17, TTM13, TTC18, VPL^{+}10, Wau19, XPBC11, YDGB^{+}12, YAT11, ZMK12]. session [SHBC19]. set [DW11, LWL^{+}13, VvSvV16, Wu11]. sets [TXCX19]. Setting [NI13]. SETZ [TTL^{+}13]. several [ZT14]. severity [SA18, ZCY^{+}16]. SF-PMIPv6 [CL13]. SGEESS [LZL^{+}15]. Shades [JBSL12]. Shamir [UUN11]. Shannon [AMS^{+}10]. Shape [RITF^{+}11]. share [HH17, LMWM18]. shared [AHW10, SBZ^{+}17, Xia13, YYS^{+}16]. shared-resources [AHW10]. Sharetouch [TCCH12]. sharing [CT11b, Che13, CLH^{+}13, CW14, EA11, GGK19, GLW13, HHH10b, LT13, LJA^{+}11, LyWSZ10, LHYZ12, MQG^{+}17, TNK^{+}19, UUN11, UUN13, WKH11, WS12, WOLS12, WS13, YWEL^{+}13, YC11, YCC16, ZXG10]. shelf [AHC^{+}11]. shifting [CSS^{+}13, HC10, HTH13, WLC13b]. short [LHZX12, San16, THS12]. Shorter [PPB16, LMT16]. should [FFdRG^{+}14, JLZ^{+}19, KM13, ZZ16]. showcase [CMK^{+}11]. SHT [PDBD18]. shuffling [Pen11]. Side [KKP12, ZGZ^{+}13]. side-channel [ZGZ^{+}13]. Signal [CWK10, RITF^{+}11, RA16]. signature [BMS11, FWCS12, HYWS11, LHZX12, Shi10, SV12, SLLL12, SXYM11, YKC^{+}12, ZM12]. signature-based [SLLL12]. signatures [GMS11, PPB16, THS12, YZC15]. signcryption [HS11b]. significant [MSGM17, Wu11]. Signs [vV13]. silver [SBAH17]. Sim [SSP^{+}15]. SimFuzz [ZLL^{+}12]. similar [ASMM18]. similarities [JLZ^{+}19]. Similarity [MG11, DII^{+}17, LBX12, LQC^{+}14, MER17, PXT^{+}13, ZLL^{+}12]. simple [MT10]. simplex [PS14]. Simplification [OT17, CL17a]. simulated [CNM18, MK15b]. Simulation [HWLM11, APW14, AWSE19, ÇT13, CXO^{+}15, CHL^{+}13, KSN17, PB11, SCGL^{+}18, SMS11, SLW^{+}15, Uzz13, VKL16]. Simulation-based [HWLM11, AWSE19]. simulations [DPP^{+}18]. simulators [dOCS13]. Simulink [HBT16]. Simultaneous [AZvG11]. since [GPD^{+}19]. single [ÁRMC16, URG10]. singular [XWZC14]. sink [CBS16]. SIP [hChSyCwL10, GFP11, HBG^{+}14]. SIP-based [GFP11, HBG^{+}14]. Sirius [TPGdS13]. SIT [QXYL16]. sites [CdR^{+}14, FG15]. situational [LK16]. situations [HCL^{+}10]. six [MVSG18]. size [ASMN15, DW11, RSGH12, SHBC19, WL10]. sized [dSdMSNO^{+}14]. Skyline [ILZ14, JHYK10, YZL^{+}14]. SLA [WZJI14]. SLA-aware [WZJI14]. Slantlet [TK14]. Slice [MLD^{+}14]. Slice-based [MLD^{+}14]. slices [JJC^{+}14]. slicing [PB11, QBO^{+}14, aSRZ^{+}18, YBE17, ZS16]. sliding [DS12, NDS13]. slot [SRS15]. slowly [FS14a]. SMACK [TDW^{+}14]. small [AT18, BdMSNO^{+}17, DY15, HBOS13, Jør14, PPG^{+}10, VA17, dSdMSNO^{+}14]. small-to-medium [VA17]. smart [AMCC14, AKA^{+}15, CFL^{+}18, GGB19, HCC10b, KKP12, PBM19, Sko14, YSL^{+}10, GSN^{+}15, HWdS^{+}15, LZL^{+}15, PCG^{+}14]. SMCD [EA14]. smear [HHC12]. smell [SRJL^{+}18]. Smells [GK18, FLRT19, GPD^{+}19, KHW19, OKS^{+}15, SS18, WFF18, YC13, FFWE17]. SMEs [CO12]. smooth [GRT13, YC11]. smoothing [WQJZ10]. smoothness [LBCL10]. SMS [PSdO^{+}13, SC14]. SMSCrypto [PSdO^{+}13]. snowballing [DVPY^{+}19]. SOA [PZ15]. SoC [CTL10]. Social [AZX14, BV18, GMGTdFR14, Woh16, AGBD14, CdR^{+}14, ECRVMS11, HY11, JLY14, KAU16, KB16, LS17a, PSM12, RNC14, Sko14, SZS13, SHH^{+}15, TCCH12, TPTV17, WSM15, dVRB13, Cha17]. socially [MPS^{+}12]. Society [BEZ14, PMMM11]. socio [ZCC^{+}19]. socio-technical [ZCC^{+}19]. sociotechnical [BPB19]. Soft [HJP15, CF12, LSE12, WX10, ZW15]. SoftProcessors [WLZ^{+}17a]. Softw [AAH12b, WZM12a, XTZX13, wZfG14a, YWEL^{+}13]. Softw. [BKSM14]. Software [AAGT16, AB16, AS10, AK16, AC16, APS^{+}10, BWP16, BGS^{+}16, BCEF10, BEZ14, BPQP^{+}10, BdMSNO^{+}17, BBND^{+}18, BMB19, BD10, Bor12, Bos12, BCL12, CBT^{+}14, CFL19, CS19, CC11, CA14, CVGP13, EB14a, ESWA18, Gar13, GMLSF^{+}15, HLS^{+}13, HC15, HST16, HG18, JED18, KM17, KR14, KT16, KM13, KJ10, LH12, LLM^{+}17, LKJR10a, LCCJ10, LGH^{+}17, LHP^{+}10, LN13, ML18, MH13, MSS18, ME10, PW10, PMB15, RTM19, RSBA19, SPTM15, dAGSdFS^{+}15, SS17, SCL13, SKRB19, Sta10, SSAS11, TL14, TGBF17, UD10, VLL18, VZT17, VCB^{+}18, WL15a, WTG^{+}15, WLL19a, WL10, Woh16, WCTK12, dSdMSNO^{+}14, dAK18, ASGJ13, AJLS10, AZvG11, AT18, AC19, AKH12, ADCO18, AC17, ACB18, ASG17, AAH12a, ACG^{+}15, AVGM19, ATvHJ18, AMKD13, AM18, ARS10, APW14, APS16, AKL14, ASMM18, AGC13, AKKS11]. software [ABC^{+}13, AdB17, AB10, ABL15, AAA11, APCS10, ACW10, AS16, AHC^{+}11, ANC11, BP13, BCBZ14, BWW^{+}18, BMA^{+}13, BM18, BKS15, BCFP19, BCR^{+}19, BV16, BKH10, BFLZ13, BCL^{+}18, BZ14, BLTY18, BGG10, Bis13, BDV17, BRS^{+}18, BBS10, BDM^{+}19, BS15, BK11, CX10, ÇB16, CCdL^{+}16, CGS19, CJT^{+}16, CCM12, CdS18, CFMRL11, CdCAdO18, CdSdSG^{+}18, CdCMdMSNdA16, CBAV16, CKCK15, CLR18, CCG^{+}10, CHLW17, CSN^{+}17, CZC^{+}18, CD10, CL15, CL17b, CLL14, CNSG12, CMR19, CO12, CBVF19, CKS15, CSM15, CPR13, CPRT16, CPV^{+}14, CNMR18, CFA^{+}19, CCMOM19, DC17, DGRN10, DWC17, DNBM12, DCT17, DGCA17, DCP12, DLW^{+}13, DFG^{+}13, DNSH13, Dut15, ESM^{+}19b, EbAT13, ETM10, EB14b, EBC10, EZRK16, EK13, FHL^{+}18, FKA16, FM11, FOR19, FP19, FBB^{+}12, FFV19, FS17, FFdRG^{+}14, FMRM15, FRGC10, FCB^{+}16, FCRF16]. software [GAMW14, GL14, GMMGP15, GPP^{+}17, GV10, GZ13, GCBCD15, GCDY16, GK18, GGT^{+}19, GJ16, GGC16, GBH^{+}16, GD12, GC13, GHBD^{+}16, dGFDL16, GTA14, GFWA18, GA13, DDF^{+}13, GMMC13, GWW^{+}11, GW10, HBP^{+}17, HNZ17, HJN11, Han12, HTB12, HBR19, HS11a, HJP15, HPF16, HLLS13, HKS^{+}17, HBOS13, HSM16, IF19, IAA16, IZ18, IF10, JLGM17, JS11, JBSL12, JG14, JCYT16, JSM10, JS13, JK12, Jør14, JDLS16, Jør16, JST10, JKL19, JR15, JC10, KCAS13, KRD16, KCT12, KTF15, KKT17, KLA^{+}19, KGB11, KFN19, KS19, KGW12, Kel15, KNA11, KSIZ19, KLB15, KT12, Kit10, KK17b, KSH^{+}12, KM14, KAM13, KTF^{+}16, KR16, KS16, LWB^{+}13, LCM^{+}13, LMIV15, LWSH19, LFW15, LXG10, LAT10, LG15, LJA^{+}11, LSD^{+}16, LCJ10, LLL17a, LSLG17, LMS12, LJ16, LMA15, LMNA17]. software [LLS11, LZR16, LYC14, MSB18, MBF12, MWM12, MdOBW^{+}15, MNS13, MEB^{+}10, Man16, MAH18, MCHJ17, MFTP18, MCV16, MRT17, MFMCY12, MDBC17, MVSG18, MACB19, MS17a, MGAN18, Mer13, MT13, MB17, MA10, MPAA15, MdFD^{+}15, MGvFGCB10, MD16, MSK^{+}17, MSSMDC12, MA17, MKHLB16, NCK^{+}15, NMM13, NHC13, dONTF^{+}19, NBA^{+}15, NCW^{+}19, NKZ17, NSM17, NBM19, Oja16a, OD10, OY16, OB13, OCC13, ÖT18, PEO11, PDS19, PB11, PB15, PAB^{+}17, PLHP^{+}15, PCHW12, PH13, PCYZ12, PPMM14, PFG13, PPG^{+}10, PRN17, PSZ17, PCCK18, PDL^{+}16, PFL16, QGZ^{+}15, ROR11, RBT11, RF18, RRV19, RSB^{+}14, RSB^{+}16, RPT19, RCCVB11, RSGH12, RHL^{+}17, RF14, SOS^{+}18, SW19, SD16a, SB17a, SGP12, SBT19, SNL16, San16, STA19, SAR15, SA12, SSMvD16, SCwY12, SLB14, SAH12, SS18, SSSA17, SSS17]. software [SXYW14, dMSSS^{+}13, SWA^{+}13, SLR16, dSSVV11, SGMHJ13, SBDB19, Som13, SHC^{+}11, SZW^{+}16, SdSGdMSN^{+}13, SA18, SG12, SB14, SNDC13, SM16, SHHL12, SHGT16, SJH^{+}10, TBSvdW18, TJH15, TBG13, TTR^{+}13, TR18, TKP^{+}18, TKCR14, TGE17, TC12, TCS18, TTMI19, URG10, UGFK15, Uzz13, VCdA^{+}16, VA17, VAJ18, VLC^{+}17, VVA^{+}15, VBC^{+}14, VHFST15, VHFF^{+}17, WCC12, WWSZ15, WKbOS17, WB12, WMC17, WGS^{+}14, WWSS13, WRR14, WRdMSN^{+}13, WSM15, WTG^{+}11, WLL17, WAWO12, XLW18, XYCL17, XB19b, XLX^{+}19, YMM^{+}17, YMM^{+}19, YFZ^{+}16, YLXZ16, YLA16b, YLA16a, YCA17, YAKK16, YHMS16, YLCZ12, ZÁ15, ZADA15, ZML10, ZLC^{+}14, ZCY^{+}16, ZZC18, ZWF^{+}18, ZFY^{+}19, ZGYS^{+}15, ZZP15, ZP17, dSF12, dL13, dOSdAdSG17, dRSBA13, dSB12, fLSN18, vVT16, CCCY17, IBAH12, JWT17, LMWM18, MP12, MMB10, NFSM11, Shi12, TTT14, VPMVM^{+}13, WVT^{+}14, WB10, Ano19l]. software-as-a-service [BZ14, WVT^{+}14]. software-intensive [AAA11, FOR19, GBH^{+}16, MAH18, dSSVV11, YMM^{+}17, YMM^{+}19]. software-producing [BV16]. solid [nWsCqW12]. Solidifying [VPMVM^{+}13]. solution [GGS^{+}19, HHH^{+}10a, LQLC16, PPN^{+}15, TBC^{+}16, nWsCqW12, XJZ^{+}15]. solutions [FCMJ12, FCRF16, KSKP11, MSS18, MAEL19, Rya13]. solve [DRCA^{+}19]. solver [EK12]. solving [ADTZ12, ACH19, BRS^{+}18, EMBS17, KK17b]. some [HHKWB16, MNO18]. Sonata [GBDCR12]. sound [LSR13]. Source [LMWM18, MP12, NVPGMPSM17, OHL17, Shi12, ACB18, BCG^{+}14, CAHV15, CHL^{+}19, DFCPSF15, ESM19a, EAH^{+}11, GPPT16, GW10, HNZ17, HBR19, IKBH14, KTF15, KKT17, KR14, KHMA12, KKA^{+}19, KK17b, LAT10, LWZ^{+}16, LWZ12, LRD^{+}19, PAB^{+}17, RA16, RNR17, SG12, VGSN18, WFF18, YLXZ16, ZFY^{+}19, CFMRL11, GL14, LLS11]. Sources [HSS14, CDOP15, LWZ^{+}16, NTRN11]. space [BAI^{+}14, DGRN10, MM19, PWC12, RKK16, VVA^{+}15, Xia13, vHJPB^{+}17]. spaces [CGS19, GBDCR12, PN14]. Spam [PÁC13, ROFGFRM13]. SPAPE [BKSM14, BKSM13]. spark [MPN^{+}17, MK17]. spark-based [MPN^{+}17]. sparse [vV10]. spatial [HSL14, MLGA11, MC10, RVCM17, YWWS10]. spatio [Lin12a]. spatio-temporal [Lin12a]. SPDX [KKT17]. Special [AC19, ADMOK^{+}10, BCEF10, BEZ14, BFLZ13, Bor12, BKW10, CCCY17, CLR18, CA14, CL11, Dut15, GP10a, LH12, MACB19, MS17a, OPS11, TZB19, VZT17, WB19, Won10, WCTK12, YAT11, Al 12, BCG^{+}13, CCM12, CdS18, DIB14, FKA16, FOR19, GBG10, JWT17, ML18, MJ19, NBM19, PBM19, PS16, SLR16, WMAS12, WMC17, WC16, XST18, ZTPT18, dAK18, SS17]. specialization [LMGHB17]. Specific [KVH12, ACG^{+}15, AMCC14, ARS17, EMBS17, HAE^{+}15, HGMB13, KMK16, PC10, SKL10, SHS16, VPdP13]. Specification [WWY^{+}12, BZ10, CF13, DBZ16, KU10, LKR13, MVSG18, MA11, SGK12, SdSGdMSN^{+}13, TFS10]. specifications [GA13, SAMN12]. specificity [IZ18]. specified [PRN17]. Specifying [ZYA^{+}18, MGR^{+}13]. spectrum [AMAY19, CCWT13, JJC^{+}14, MMSD13, TXCX19, XYZ^{+}19, ZYZ^{+}17]. spectrum-based [XYZ^{+}19, ZYZ^{+}17]. speed [XZP^{+}10]. speeding [dNPM18]. SPI [PW10, CO12, WR10]. SPI-LEAM [PW10]. SPIN [ASdMGM14]. spirited [HL10]. SPLC [GP10a]. split [HCB^{+}16]. split-and-merge [HCB^{+}16]. splitting [LWOY16]. sporadic [wZfG13, wZfG14a, wZfG14b]. Spot [LZO^{+}16, WMOKY11]. spots [WLZ^{+}17a]. Spotting [GHBD^{+}16]. sprays [HHH^{+}10a]. spread [MMSD13]. spreading [HLWS13]. spreadsheet [CFM^{+}16, JSHW14, KHW19, ZXC^{+}17]. spreadsheets [LT13]. sprint [GRT13, LHCT19]. SQL [FGB^{+}19, GLOM19]. SQLIA [Aba13]. squaring [LKP13]. Stability [MGvFGCB10, SB17a]. stack [GXZ^{+}19, ZFY^{+}19, CCD19]. stage [XLL^{+}19]. stages [MAAC17]. Stakeholder [Hoo14, PG12, vdRBSvV10]. stand [SSD16]. stand-up [SSD16]. standard [MG11, SXYM11, WZM12a, WZM12b, YC12]. standardized [GS17]. Star [BV18, MTF14]. Starring [BV18]. start [FHL^{+}18, SAR15]. start-up [FHL^{+}18]. started [AS10]. startup [BBND^{+}18]. Startups [ESWA18, TKP^{+}18]. starvation [SMZC12]. State [GAMW14, LDS^{+}19, MDP^{+}11, PMR16, YHM^{+}14, ACS13, ABL15, DCG16, EFSJM17, MRY17, dONTF^{+}19, PDBD18, TJT^{+}18, nWsCqW12, WMAS12, Zha16, KMWL12]. State-of-the-art [PMR16, TJT^{+}18]. statecharts [SÁMI17]. Stateless [CL18]. statements [SKK^{+}18b]. Static [ABS19, OMLB16, SLL^{+}15, ANG^{+}19, BSB12, FP18, PDS19, SNDD19, WMWZ12, YLC18, ZS16]. statically [QOLJG16]. stationary [MKRO14, MJZ^{+}10, ZL17]. statistical [FP18, KSN17, LNY^{+}11, Luk11, MLD^{+}14, dONTF^{+}19, XYCL17, ZCT^{+}11]. statistically [YAKK16]. Status [FGD^{+}17, LZHS11, GBCI11, PMMM11]. steganographic [LLC10]. steganography [EEAZ13, LyWSZ10, LWW^{+}10, SI12, WKH11, WOLS12]. step [SH17, YCF^{+}13]. steps [dONTF^{+}19]. Stepwise [SPSR17, SSP17]. stereo [CJ13]. stereotypes [SSMvD16]. Stitch [CG12]. Stochastic [BT17, ZW15, AC16, BHM12, CCG^{+}18, FCB^{+}16, OH15, PACH15]. Storage [LLGZ13, BT17, FNWL18, GCSSDP^{+}18, GPSS^{+}13, KKLB11, LMT16, LZC14, Luk11, MCC11, OSH^{+}18, WCB^{+}17, YTW^{+}13, YYS^{+}16, NC10]. store [DII^{+}17, GNA17, KCR16, MQG^{+}17, Shi17]. stories [MH12]. strategic [Uzz13, VLC^{+}17, Wau19]. Strategies [VAJ18, CNL13, CXO^{+}15, GQ12, HS15, Jør10, KA17, MVLJ18, NSM17, Oja16a, RB16, ROFGFRM13, SD16b, YWHL11]. strategy [AZ11, HSC15, KHMF13, LWL^{+}13, LNY^{+}11, LZC14, LYC14, MLHL12, SRS15, WGC^{+}14, WC11, LZL^{+}15]. stream [APS16, DM17a, HWR17, LW13a, MRBN17, TXLC12, VZT17, YF15, YCWW15]. streaming [FGBC10, LLW12, LLH^{+}16, MLHL12]. streams [CPS11, CJL11, DS12, KK17a, LJL^{+}12, LLML13, NDS13, VTZ^{+}17]. strength [AZ11, CWK10, HCT^{+}15]. stress [AL10]. Striving [Dan17]. Strong [KRDH12, FWCS12, HYWS11, LJ16, CCGG14]. Strongly [SXYM11, THS12, EZG15]. structural [AC17, BDO11, CFMRL11, KOL^{+}14, KCV^{+}19, LMIV15, NOPF12, PXT^{+}13, PACH15, XLM^{+}15, YZY^{+}18]. structure [GAKF13, HTB12, HR10, JRSN10, KHW19, LBX12, QGZ^{+}15, SM17a, ZLW^{+}12, dSF12]. structured [SMM17, WKD^{+}19, YTW^{+}13]. structures [Cic16, FMR11, ISM11, Lin12a, SAA^{+}10, WS12, CSS10]. Structuring [DGRN10, SWA^{+}13]. student [IF19]. studies [CdS18, CRSS14, DDMP14, GNA17, HWC^{+}10, JCYT16, KSIZ19, LCM^{+}13, MPTT14, PPG^{+}13, PCCLdGP12, SAH12, UGFK15, WRdMSN^{+}13]. Study [KBM18, SSR18, ASGJ13, AJG^{+}15, AAGT16, AB16, ADCO18, ACB18, ASG17, ACG^{+}15, ANG^{+}19, AMdLM17, ACS13, AAC^{+}17, ABJ^{+}17, AHC^{+}11, BRB14, BCFP19, BBND^{+}18, BCD^{+}18, BMB19, BAM17, BLTY18, BS12, BAAD17, BDM^{+}19, BHVR18, CSF^{+}14, CS15, CdSdSG^{+}18, CC11, CXO^{+}15, CO12, CPRT16, CNMR18, CFA^{+}19, DvdVA^{+}13, DVPY^{+}19, DLM19, DFG^{+}13, ECS15, ESM^{+}19b, EGHO16, EED16, EBC10, ELHC13, FSGYP17, FLRT19, FMdAR16, GMMGP15, GCDY16, dGFDL16, GTF15, GPPT16, GSdS16, GW10, HHKWB16, HBP^{+}17, HJN11, Han12, HKS^{+}17, IF10, JSL16, JWA14, JLL19, JCYT16, JAS19, JR15, KBJZ15, Kan15, KFN19, KFLS18, KJS^{+}12, KNA11, KSIZ19, Kit10, KSM^{+}16, KQ17, KMG^{+}19, KBRV17, LWSH19, LAL15, dPLV19, LPB19, MBF12, MVLJ18, Man16, MAH18, MDBC17, MAEL19, MFM10, MPLL^{+}15, MGvFGCB10, MGR^{+}13, MRS18, MRRS19]. study [MD16, MHLMG14, NCS10, NCW^{+}19, NVPGMPSM17, NBF16, OK11, OBS^{+}18, OWG19, OCC13, PSS^{+}16, PAB^{+}17, PWS^{+}15, PTF^{+}15, PV18, RTM19, RDPM19, RAS14, RASL12, RHL^{+}17, RB16, RM19b, RVCM17, SCMS15, STS^{+}19, SCwY12, SCL13, SMS11, Shi12, SKRB19, SVM19, SHS16, SCC16, SNDC13, SSD16, SLLL14, SLL^{+}15, SKF17, SAN^{+}17, TKZW17, TKSRP11, TAJ^{+}10, TAF^{+}17, TdCAF16, VBC^{+}14, WGKW19, WRR14, YC13, YLA16b, YLA^{+}17, YHMS16, ZSG16, ZXC^{+}17, ZFY^{+}19, dSdMSNO^{+}14, vHAT13, vHJPB^{+}17]. styles [KBDGAW16, KG10, LJDK10, MKS10, MCV16, SRSC16]. sub [ELHC13, LLZW14, YZC15]. sub-patterns [YZC15]. sub-swarms [LLZW14]. sub-system [ELHC13]. Subdomain [PAOC15]. Subdomain-based [PAOC15]. subject [EA14]. subject-based [EA14]. subjective [AL10, LY18]. subscribe [CDRT13, HBG^{+}13, LJC16, LVPMPCLS13, RMD11, LJDK10]. subset [XLL^{+}19]. substitutes [TTC15]. subtree [LWXZ10]. success [CO12, DPVvV19, DPL16, GGC16, Ifi11, LSD^{+}16, MP12, PHR10, RCCVB11, SNDC13, WR10]. successful [ZADA15]. suggestions [BD16]. suite [CdCAdO18, FAM15, HCT^{+}15, WAG15, YH10, ZYZZ14, LGM^{+}18]. suites [AZ11, CWK^{+}11, MH11, SPMG18]. Summarizing [RDVC19]. Summary [ZJL10]. Sun [SSF15]. super [ZLZ11]. supercomputer [SMM17]. supercomputing [RGH17]. superscalar [CD10]. supplementary [SYXL17]. supplementing [BS12]. supplier [SAR15]. supply [CPS11]. support [AHOP14, BWH10, BDG13, CNG16, EL10, EH19, GPMI13, HNZ17, HP16, HCB^{+}16, IBM11, KLL^{+}11, LZG15, MLHL12, MKS10, MIKG13, NI13, OAC11, PLVB^{+}18, PH13, PWW10, PBD^{+}12, PV18, RO13b, RA16, RRM17, RDVC19, ŠK11, TJH15, TTL10, URG10, WPL^{+}18, FSS^{+}13]. supported [AAN11, BD10, ISM11, KLL17]. Supporting [AACT13, ACL13, CFL^{+}18, CPS11, GGvH^{+}18, HBG^{+}13, HBG^{+}14, HP16, JS13, SHC^{+}11, YFZ^{+}16, DGRN10, JCK^{+}17, LGZ^{+}18, THWC10, WB12, WBBK18, YYS^{+}16, GCC^{+}15]. supportive [SKK^{+}18b]. supports [CHL11, HWL13a]. suppression [LM13]. surface [CPRT16]. Surfing [BAI^{+}14]. surrogate [MPLL18]. surveillance [MJZ^{+}10, XLM^{+}15]. Survey [CDS19, HCL^{+}10, KKA^{+}19, AAC16, ABC^{+}13, AT15, BGEP17, BCG^{+}14, CCP18, CRKH11, FBD^{+}18, GV10, GZ13, GCBCD15, GK18, GDLB16, HG18, JSHW14, LKRYTS18, LCM^{+}13, LSD^{+}16, MCHJ17, MARD16, OK18, PWS^{+}15, RRV19, Rya13, San16, SS18, SSK19, SNDC13, TTR^{+}13, TKP^{+}18, WWSS13, YLA16a, ZXTT11, dSB12]. Survey-based [KKA^{+}19]. surveys [JWA14, Sta14]. Survivable [WMD^{+}10]. Surviving [CLY14]. Sustainability [GL14, NCWK18, CFAP17, VCB^{+}18]. sustainable [GGS^{+}19]. sustained [SDG17]. SVM [TLL13]. Swarm [AZ11, PGP^{+}19, DRCG12, LLZW14, MDO^{+}10, dCPV10]. swarm-inspired [MDO^{+}10]. swarms [LLZW14]. Switching [GFP11, CCdR^{+}16, CTHW12, SYBN12, aSRZ^{+}18, WL15b, WMOKY11]. symbolic [ACH19, BSB12, EED16, dCPV10]. symbols [SB17b]. Symposium [Bor12]. synchronization [DGWC16]. synchronizing [LZCL19]. synchronous [BKRW19, KKH^{+}16]. syndrome [BMS11]. synergistic [TGP11]. synergy [ST11]. syntactic [CJL11, KOL^{+}14, QLBS17]. syntax [EA19]. syntaxes [PC10]. synthesis [CCG^{+}18, rBHM17]. synthesised [KMWL12]. synthesized [NSDI16]. Synthesizing [AMCC14, CGS19]. SysML [CKL12]. Syst [AAH12b, APS^{+}10, BKSM14, LKJR10a, LHP^{+}10, WZM12a, XTZX13, YWEL^{+}13, wZfG14a]. system [ASGJ13, AT18, AYZI10, AR17, AAA11, AWSE19, BSG12, BRG^{+}12, BDBLP15, BDG13, BSG^{+}18, CCdL^{+}16, CH11, CTL12, CNSG12, CHL^{+}13, DII^{+}17, DvdVA^{+}13, EH19, ELHC13, FNWL18, GBH^{+}16, GPSS^{+}13, GPPT16, HCB^{+}16, HCL12, Hoo14, HAE^{+}15, HWLM11, Ifi11, JS11, KK11, KAK^{+}13, KA14, KLP10, KGG18, KSM^{+}16, KMK16, KJ10, KH10, LS17a, LLLK12, LXG10, LLLK10, LHP^{+}09, LHP^{+}10, MCL^{+}17, MS16, MAH18, MCS^{+}12, MCV15, NI13, NJ17, OD10, PK10a, PNY14, PH13, PDBD18, RAK15, RA16, ST13, SV19, SGW^{+}15, SB12, SMK^{+}18, TKSRP11, TG17, TLZ^{+}16, TTL^{+}13, TCCH12, TDW^{+}14, WRTP^{+}13, WGZ^{+}12, WKD^{+}19, WKV11, WL10, WLL^{+}13, XB19b, YC13, YSG17, YH13, YCWW15, YCLC17, ZML17, ZXG10, dRSBA13, ESRF19, LLGZ13, WFY^{+}19]. system-level [WL10, YC13]. system-specific [HAE^{+}15]. system-wide [HCB^{+}16]. system/software [CNSG12]. Systematic [BEK^{+}19, GCAH18, IHA16, KBM18, PHBJ16, SKT17, AJG^{+}15, AAGT16, AB16, ADCO18, AKAA18, AVGM19, AM18, APW14, ABJ10, AS16, BWP16, BKS15, BCFP19, BBND^{+}18, BMB18, BLTY18, BDM^{+}19, CX10, CP15, CS19, CNMR18, DLM19, DPL16, DBCG14, DZT^{+}14, ESM^{+}19b, FSGYP17, GRR16, GJ16, GNA17, GA11, dGFDL16, HBP^{+}17, JED18, JCYT16, KBJZ15, KGB11, KNA11, KSIZ19, KQ17, KBRV17, LFW15, LL15, LZO^{+}13, LAL15, dPLV19, MWM12, MH13, MRT17, MRY17, MAEL19, MGAN18, MD16, NVPGMPSM17, OGRJ^{+}18, ÖT18, PG12, PPG^{+}13, PMB15, PFO^{+}19, RAK15, RSBA19, RHL^{+}17, SNL16, SRJL^{+}18, STA19, SLB14, TTM13, TAF^{+}17, TLGE18, TCS18, VLC^{+}17, VCMG17, WPL^{+}18, WNC17, YLA16b, ZADA15, ZSG16, ZGYS^{+}15, BPQP^{+}10]. Systems [ABB15, Ano19l, BEZ14, HST16, TTT14, UB19, WLL19a, Woh16, YMM^{+}19, Aba13, AC19, AZX14, AB16, ADMOK^{+}10, AR18, AHLH16, ACRD19, AAC16, ÁRMC16, ABL15, ACW10, BL19, BSK^{+}18, BM18, BLL^{+}18, BPO^{+}16, BD16, BHH^{+}12, BM17, BT17, BDM^{+}19, BKRW19, CX10, CCG^{+}18, CGS19, CSMC19, ÇT13, CWK^{+}11, CCY11, CCH14, CLY17, CIB^{+}19, CYT16, CKC15, CNKL12, CHCO11, CH10d, CDS19, CSM15, CDPM17, CCMOM19, DWC17, DBZ16, DFJ19, DNSH13, Dut15, EZOK14, EGG^{+}11, EB14c, EBJ17, EK13, ETYL15, FKA16, FVHF^{+}15, FOR19, FKWVH19, FTC16, FTSC12, FGBC10, GKD13, GMPN16, GVPPM18, GJ16, GMR17, GBH^{+}16, GSP^{+}19, GMLSF^{+}15, GCC^{+}15, GHBD^{+}16, GBC16, GGM11, HyLW^{+}12, HdM17, HZG^{+}12, HNS12, HK13, HGMB13, HH17, JKC19, JSM10, JS16, JAS19, JRO12, KRD16, KHSD10, KHS11, KTF15, KH14]. systems [KKL^{+}11, KAS18, KGG18, KAM13, KKLB11, KHC16, LJH10, LJC16, LLM^{+}17, LMS11, LMN10, LSE12, LS17b, LFCL12, Li11, LLW12, LWL^{+}13, LG15, LGHR16, LL10, LLK11, LC11, LNW^{+}11, LNY^{+}11, LLL^{+}14, LZC14, LDS^{+}19, LLS11, MJF10, MRT17, MBAG11, MB19, MAEL19, MPLL^{+}15, MA11, MNSA15, MNSA16, MHLMG14, NCK^{+}15, NCWK18, NKMM12, NK15, NPC12, NTdSX13, OSH^{+}18, OKMD12, OB13, PSM12, PSS^{+}16, PCHW12, PBM19, PCYZ12, PPMM17, PGPC17, PFG13, PRN17, PDL^{+}16, PZ15, RAK15, RXY^{+}19, SYBN12, SJR^{+}11, SÁM^{+}16, SNL16, SRT^{+}12, SY16b, SJ17, SS14a, dSSVV11, SK10, aSRS^{+}10, SHH^{+}15, SJH^{+}10, SAN^{+}17, TZ12, TAB^{+}16, THWC10, TCS18, URG10, VM12, VRPT18, VZT17, VRG^{+}16, VHFST15, VHFF^{+}17, WFF18, WL17, WMWZ12, WMAS12, WBBK18, WTG^{+}11, WB15, WX10, WWY^{+}12, XB19a, YAY13, YWWS10, YTW^{+}13, YSJ13, YSSaR14, ZK13, ZMB14]. systems [wZfG13, wZfG14a, wZfG14b, ZXL10, ZGSH13, ABCH13, JWT17, WFY^{+}19, WL10, ZAY19].

T [CM12]. T-REX [CM12]. ‘T. [YWEL^{+}13]. table [WWLG13]. tabling [AR17]. tactical [ETYL15, STS^{+}19]. Tactics [UB19, WK15, HA10, LL15, MPLL18]. Tactile [RLL^{+}18]. tags [CCD19, WL17]. TAIC [BKW10]. TAIC-PART [BKW10]. tailoring [CP15, CCP18]. Taking [Bos12]. TALISMAN [GDFFPG^{+}10]. TARA [Woo12]. target [GTY12, LT11, SA18]. targeting [MA17]. Task [KHSD10, KHS11, SKT17, BRS^{+}18, CSMC19, CYT16, CKC15, DFJ19, DCT17, FHL^{+}15, KSN17, LS17b, LWL^{+}13, NI13, SK10, TKJL13, TC16b, WX10, ZW15, ZCC^{+}17, ZGL^{+}10]. task-aware [CYT16]. task-clustering [DFJ19]. task-solving [BRS^{+}18]. tasks [AMP12, ABB19, CZG^{+}15, FBD^{+}18, GGS15, JSL16, KWS^{+}17, KA17, LCLS16, MER17, RBS19, Wen16, wZfG13, wZfG14a, wZfG14b, ZHGL11, ZGSH13, MK15b]. Taxation [LLW12]. Taxonomy [KSENM17, SZ11, SS14b, DGWC16]. TD [SOS^{+}16]. TDQN [MB19]. teacher [NI13]. Teaching [Som13, SBAH17, vWSB13]. team [HM16, LCCJ10, OCC12, RSGH12, RKK16]. teams [DCP12, GD12, GTF17, LS17a, LSD^{+}16, TR18, VBC^{+}14, VvSvV16, YHMS16]. Teamwork [JKC19, LSD^{+}16, IF19, RF18]. Technical [MS16, MGM16, BMB18, BMB19, CREH^{+}18, FKA16, FSGYP17, GSdS16, LAL15, MKS^{+}18, PWS^{+}15, SSK19, TAV13, YHMS16, ZCC^{+}19, VM12]. technique [BBBP13, EZOK14, GCSÁddP11, HPH12, HR10, LWLL12, MK15a, PMDH13, SAA^{+}10]. Techniques [BKW10, AAM^{+}17, BRB14, BPGS13, BLTY18, CBT^{+}14, CL18, CKS15, CPR13, DC17, FYCL13, FBD^{+}18, IAA16, KSAR18, LH11a, LHLG^{+}15, PFO^{+}19, RO13a, RGH17, SLB14, SLL^{+}15, TSCB19, TTR^{+}13, WAG15, ZFS15, ZSG16, ZML10, WMD^{+}10]. technologies [PPN^{+}15, WPL^{+}18, YSJ13]. Technology [CCCY17, LWZ12, AT15, CLR18, CCWT13, LPM15, MCV15, NHH^{+}12, PHR10, SMM17, WPL^{+}18, Wie14, XLM^{+}15, ZMK12]. teleo [MNSA16, SÁM^{+}16, MNSA15, SAMN12, SÁMI17]. teleo-reactive [MNSA16, SÁM^{+}16, MNSA15, SAMN12, SÁMI17]. TelosB [APS^{+}10, PAS^{+}10]. Temperature [WX10, ZCC^{+}17]. Temperature-aware [WX10]. template [GCSÁddP11, ZZ16]. templates [NBA^{+}15, SGK12]. Temporal [Pra18, Lin12a, LNW^{+}11, MC10, SKE10, WWY^{+}12]. tenancy [KBJZ15]. tenant [LZG15, MCC^{+}18, MVLJ18, PHBJ16, WVT^{+}14]. tendency [MRS18]. Tensor [nQYD11]. tentative [LZY^{+}15]. term [UD10]. terms [CAHV15]. tertiary [KSIZ19, NBF16, RTM19]. Test [AG15, AMdLM17, CZC^{+}18, HMOK18, KMK16, LCL^{+}12, MGM10, SD16b, SCC16, AAGT16, AZ11, ABC^{+}13, AWSE19, BFLZ13, BGLG13, CL18, CF13, CWK^{+}11, CKMT10, DVPY^{+}19, DW11, EFSJM17, EGM^{+}11, FAM15, GK18, GKS18, GZY11, GTY12, GP10b, GPD^{+}19, GEM15, HBT16, HN17, HWC^{+}10, HCC10a, HPH12, HCT^{+}15, JC15, JCK^{+}17, JKL19, KAS18, LQLW12, LNTS19, LLWL19, MH11, MCTM11, MB17, PS13, PSS^{+}16, PAOC15, PWA^{+}19, Pra18, QBO^{+}14, SPMG18, SB12, TAS^{+}18, TGKL19, UGFK15, WQJZ10, WGC^{+}14, WAG15, WXY^{+}17, WZY^{+}18, YH10, ZYZZ14, ZJZ^{+}17, ZYZ^{+}17, ZZC18, ZAY19, ZTPT18, BMKM15, ZLL^{+}12]. test-case [HCC10a]. Test-Driven [BMKM15]. test-point [BGLG13]. test-to-code [QBO^{+}14]. testability [AAM16]. testbed [RLY^{+}13]. tester [RPSL10]. Testing [ABCH13, BKW10, CKMT10, CPR13, WCTK12, XHM^{+}11, ANG^{+}19, AAM^{+}17, AAB19, AL10, AWSE19, BRB14, BBEM11, BAAD17, CNM18, CF13, CWK^{+}11, DGBE18, DBCG14, DFG^{+}13, EED16, FFV19, GV10, GZ13, GKS18, GP10b, GCMB17, HZH^{+}16, HJP15, HPH12, JCYT16, JCK^{+}17, JKL19, KAO13, KSN17, KSH^{+}12, LHJ10, LCM^{+}13, LXJL10, LQLW12, dPLV19, LNTS19, LLL^{+}14, LGS^{+}19, LYC14, MK16, MFTP18, MRT17, MFMCY12, MS17a, MN19, MSHG18, OD17, PS13, PK10b, PW18, PFO^{+}19, PACH15, PV18, QXYL16, RB16, SD16b, SCL13, SSP^{+}15, SRBT18, TTM13, TG17, TT13, TTT14, XLW18, YCG^{+}14, YSSaR14, ZSG16, FH10]. tests [CPV^{+}14, Kim17, RBS19, SCC16]. text [BLTY18, IZ18, Kan15, PWC12, SI12, SLLY17, TCK14]. text-based [PWC12]. textual [AS17, OFR^{+}12, QBO^{+}14]. theft [BTPLST15, CKCK15]. their [CFL^{+}18, CHN19a, HRRC16, KCV11, LJ16, MNO18, PSZ17, RSB^{+}14, VLL18, WFF18, vHAT13]. Theoretical [SOS^{+}16, LWL^{+}16, ZYZ^{+}17]. Theory [GN15, Woh16, AKH12, CL17b, DC17, Han12, JG14, JMML17, LPB19, SSD16, VA17, WSM15, XJZ^{+}15]. there [KL11]. Thermal [TC16b, CKC15, ZW15]. thermal-aware [CKC15, TC16b, ZW15]. Thermal-throttling [TC16b]. Things [MOD^{+}19, PC15]. Thinking [CMR19]. third [AHC^{+}11]. Thread [LCLS16, TLZ^{+}16, WLL19a, WLL19b]. Thread-level [LCLS16]. thread-related [TLZ^{+}16]. threads [WLL19a, WLL19b]. Threat [TCS18, WSJ14]. threats [CRL^{+}12, KOS15]. three [DGWC16, KSM^{+}16, LWC^{+}18, LZC14, ST13, TC16b, YC12]. three-dimensional [DGWC16, LWC^{+}18, TC16b]. three-level [ST13]. three-party [YC12]. three-phase [LZC14]. Threshold [CT11b, GLW13, YWEL^{+}13, Kim17]. Thresholds [MSGGL12, FBB^{+}12]. thriving [vV13]. throttling [TC16b]. tied [EZG15]. tier [WDC12]. Time [SKF17, AMP12, ABB19, AMAY19, ACB18, ACL13, ÁRMC16, AGC13, AAC^{+}17, BLS18, BL19, BMS11, BKRW19, CPS11, CLL10, CZG^{+}15, CBL^{+}15, CLF^{+}13, CS12, CF12, DYC19, EBEL18, EGG^{+}11, EK12, EK13, FHL^{+}15, FHY17, GLZ15, GPPT16, GBC16, HyLW^{+}12, HCB^{+}16, HZG^{+}12, HNS12, ICSK14, IYS13, JLZ^{+}19, JAS19, KC16, LESL11, LSE12, LS14, LS17b, LFCL12, LWL^{+}13, LC11, LNY^{+}11, LW13a, LHP^{+}09, LHP^{+}10, LKK14, LGS^{+}19, MBD13, MFMCY12, MSAH16, MT10, MK11, MMTS15, NPC12, PNY14, PPB19, PG15, RFM10, RXY^{+}19, RGH17, SW10, SB17b, SMS11, SAKZ15, SK10, TAS^{+}18, TKJL13, TKJ15, TC16b, VZT17, WMWZ12, WX10, wZfG13, wZfG14a, wZfG14b, ZW15, ZHGL11, ABCH13, HL10]. time-based [SAKZ15]. time-decaying [JLZ^{+}19]. time-division [MSAH16]. ‘Time-out [HL10]. time-series [LNY^{+}11]. time-synchronous [BKRW19]. time-triggered [SW10]. timed [HRD10, MXZ11, PJT^{+}17, ABCH13, YHM^{+}14, ZLG10]. timed-release [MXZ11]. timing [CWK^{+}13, CF12, SÁM^{+}16]. tiny [PWY^{+}16]. TinyOS [OMLB16]. together [ESM19a]. tolerance [AM15, CCH14]. tolerant [CSW10, GPSS^{+}13, LLH^{+}16, NSAK10, WMWZ12, YSDT11, ZHGL11]. tomography [BAI^{+}14]. too [HLS^{+}13]. tool [AT15, ABFM12, CDGJ10, ÇT13, HP16, MTA^{+}16, OAC11, RHRC13, RHRC15, RRM17, TAS^{+}18, TVMS18, TC12, WBBK18]. tool-support [HP16]. tooling [CPDM16]. toolkit [MRJD^{+}12]. tools [ANG^{+}19, HBR19, KTF15, MG11, RDPM19, RAK15, TAJ^{+}10]. Top [SHN14, Won10, MLLK11]. Top- [SHN14, MLLK11]. Topic [CSN^{+}17, YFZ^{+}16]. Topic-based [CSN^{+}17]. Topics [CA14, CCD19]. topological [TNK^{+}19]. topology [AN10, DMSG11, LLHY19, MARD16]. TOPSIS [LY18]. Tor [MK15a]. ToscaMart [SBB^{+}16]. tossing [BNS12]. totally [JHYK10]. totally- [JHYK10]. TOTAM [BSDD14]. Touch [SHBC19, CTL12]. TPM [PWY^{+}16]. Trace [MB19, CWK^{+}13, GKV14, GXZ^{+}19, dL13]. trace-based [dL13]. Trace-driven [MB19, CWK^{+}13]. traceability [GE15b, LKJR10a, LKJR10b, MG12, MSS18, QBO^{+}14, TGE17, WPL^{+}18, WBBK18, YYS^{+}16]. traces [AHLH16, CBSM16, GKV14, MHLMG14, PH13, PDBD18]. track [BSK^{+}18]. track-based [BSK^{+}18]. Tracking [WLL19a, WLL19b, CBSM16, LT11]. trade [CFMRL11, SPCT18]. trade-off [CFMRL11]. trade-offs [SPCT18]. tradeoff [CGS19, PCYZ12, Pot13]. traditional [VTZ^{+}17]. traffic [BSK^{+}18, CCdR^{+}16, DII^{+}17, LJM11, LLH^{+}16, MPTT14, WC11, WMOKY11]. TRAILS [WBBK18]. training [DBL^{+}18, KJ10, LMNA17, NBH19, XLL^{+}19]. traits [ARH^{+}17, BD17]. trajectories [GSN^{+}15]. trajectory [CL17a, KPT13, VTZ^{+}17]. transaction [DK15b, DWC17]. transactional [KVT^{+}17]. transactions [HyLW^{+}12]. transfer [MXZ11, RLY^{+}13, WPL^{+}18, YJZ17]. Transferring [MFTP18]. transform [CJ13, Lin16, NES^{+}14, TK14, yWpWyYpN13, WS13]. transformation [BGTC18, EBEL18, JMML17, KLL17, LKRYTS18, Lin12b, MBPM19, nPHW^{+}16, PRN17, SHC^{+}11, YHM^{+}14]. transformations [CCGdL10, CCGdL16, DPL16, LKR13, SDB16, TSRC18]. Transforming [SS14a]. Transition [GC13, JMML17, GN15, KK12, LCL15]. transitions [EAH^{+}11]. translation [KKLC12, KAS18]. transmission [MMTS15, PSdO^{+}13]. transparent [CCdR^{+}16, LLLK12]. transport [LHP^{+}09, LHP^{+}10, ST11, XZP^{+}10]. Trapdoor [RPSL10]. tree [GGvH^{+}18, HWL13a, LHY12, PDBD18, RLL^{+}18, SA11, WQJZ10, ZLZ11, Zha12a, PDBD18]. tree-based [HWL13a, SA11]. tree-like [ZLZ11]. trees [BTPLST15, KMWL12]. trends [Chr16, GBCI11, LZHS11, MKHLB16, PMMM11, YCA17]. trials [TKH^{+}11]. triggered [SW10]. triggers [FGB^{+}19]. trimmed [TTL10]. Triple [LW13a]. Triple-image [LW13a]. TRiStar [MNSA16]. troubled [KP10]. Trust [AHH^{+}10, BCLW11, KK11, MRM16, MLD16, RNR17, SFMB16, TR18]. Trust-based [BCLW11]. trusted [PWY^{+}16]. trustworthiness [KR14, LY18, SXYW14]. Trustworthy [BEZ14, KK11, LLWL14, MA11]. truthful [KBRV18]. truths [KA17]. Tsao [YWEL^{+}13]. TSTSS [XLL^{+}19]. Tuning [GSP^{+}19, DWC17, PCYZ12, SRT^{+}12]. Turkey [GCBCD15, GCDY16]. TV [AM10b]. Twenty [VCdA^{+}16]. Twenty-eight [VCdA^{+}16]. twig [CJL11]. twig-query [CJL11]. Twitter [CCGG14]. Two [HWC^{+}10, MT13, YSL^{+}10, BV16, ÇB16, HJ12, LCM^{+}13, PPG^{+}13, PFL16, TLK^{+}16a, XLL^{+}19]. two-phase [HJ12]. two-stage [XLL^{+}19]. type [ASMN15, KCV11, TPGdS13]. typed [QOLJG16]. types [CPZF19, CPR13, MRS18, WH15]. Typical [ZDC^{+}11]. typing [SY16b].

U [GMGTdFR14]. ubiquitous [ADMOK^{+}10, CdCAdO18, GZKL13, HGP^{+}12, MDP^{+}11, SNL16, SY16b]. UCSD [LGC17]. UDP [BP15]. UI [KL10, KL11]. ultrasound [CCWT13]. UML [CCR14, Cic16, HJBH10, KSS15, LCLP16, OT17, PC10, SDB18, SHS16, TLGE18, TGP11, WWSS13]. UML-based [HJBH10, SHS16]. UML/OCL [CCR14, OT17]. UMTS [OHJ10]. unbounded [LGS^{+}19]. uncaught [OBS^{+}18]. uncertain [CZG^{+}15, LW13b, MAG12]. uncertainties [CIB^{+}19, PS15]. Uncertainty [CPYZ14, ZAY19, ATvHJ18, GE15b, SFMB16, TGE17, WLL15, ZYA^{+}18]. Uncertainty-wise [ZAY19]. underfeeding [BBBP13]. undergraduate [IF19, RF18, RHM^{+}18]. underlying [dSF12]. Undersampling [LLC17]. Undersampling-Boost [LLC17]. understandability [CFRPC^{+}18, MNO18, MNSA15, MNSA16]. Understanding [AC17, CPZF19, CFA^{+}19, EGHO16, FMR11, KA18, KQ17, MPS^{+}12, NBF^{+}19, QLBS17, SA12, SS12, Kel15, ZFY^{+}19, dSF12, BV18]. underwater [LWC^{+}18]. unforgeable [SXYM11]. Unicode [PWC12]. unified [BMB18, GPSS^{+}13]. uniform [GP10b, MGB16]. UniSpaCh [PWC12]. Unit [Jør16, EED16, PV18]. univariate [LM13, LW13b]. universal [RA16]. universe [FNWL18]. universities [Fug12]. university [MHB18]. unknown [HAE^{+}15]. unlabelled [ZZC18]. Unreliable [XZP^{+}10]. Unsupervised [AIE19]. Unusual [TLA18]. Unveiling [LAH^{+}16, JLY14]. update [HyLW^{+}12, IBAH12]. updating [GCSÁddP11, MIUM12]. upgrades [BCBZ14]. upon [Lin12b]. Upperware [BSG12]. urgency [CBL^{+}15]. URL [HRRC16]. URLs [CCY11]. usability [AJLS10, ACG^{+}15, AL10, BGG10, BS15, FAI13, FH10, JAS19, RAJ15, TPGdS13, VHL14, WK15, WR10, WRR14]. usable [PSS11]. Usage [SHGT16, GHBD^{+}16, dGFDL16, NHH^{+}12, NKZ17, PTF^{+}15, QLBS17, SOS^{+}18, Sal17, SPSR17, SS12, SDG17, Sta14, SK13, TKZW17, XSL^{+}18]. Usage-based [SHGT16]. USDL [GS17]. use [APW14, BS12, BHVR18, EA19, EVR11, EBC10, FG15, GGK19, GTA14, HHKWB16, HGBS18, HBR19, JK12, LS17b, MCHJ17, MG11, MAS13, MSK^{+}17, MHLMG14, OGK13, SS14a, SDB18, WLD16, ZYA^{+}18, dB12, SSP17]. use-case [GGK19]. used [ZZ16]. usefulness [ZZC18]. User [DLW^{+}13, OD17, dSSVV11, Aki18, AKL14, APT^{+}12, BCFP19, CCY11, CMK^{+}11, CH10a, GNA17, GW10, ILZ13, LXJL10, LHG^{+}18, LLHY19, LZLC17, LASL14, LSLG17, MNO18, MH12, MCV15, PLVB^{+}18, RZMPM12, SP14, SHBC19, TZ12, TKH^{+}11, WOLS12, XYCL17, YSL^{+}10, ZÁ15, AHBA19, GC13]. user-centered [ZÁ15]. user-friendly [MCV15, WOLS12]. user-input-validation [LXJL10]. user-participating [CH10a]. users [BPGS13, CFL^{+}18, Kan15]. Using [ASdMGM14, AAM16, CCdR^{+}16, CXO^{+}15, EA14, JG14, LBCL10, Lin12a, MER17, NHH^{+}12, PPG^{+}10, RBS19, SSF15, SAH12, WZY^{+}18, WRS^{+}17, XPBC11, YH10, AR12, ABCH13, AMAY19, AZ11, ACB18, AC16, ANC11, BDO11, BNS12, BK17, CCR14, CF13, CPD^{+}18, CNL13, CIB^{+}19, CSW10, CCWT13, CSW13, CL15, CL17b, CBL^{+}15, CHCO11, DW11, DYC19, DPP^{+}18, EEAZ13, EBEL18, FS19, FF12, FSS^{+}13, GZKL13, HZ15, HSPD14, HC10, HCL12, HPF16, HCC10a, HCC10b, HS11b, JJC^{+}14, KHSD10, KHS11, KSN17, KNA11, KSAR18, KM11, KCV^{+}19, KA14, KMWL12, KKP12, KLB15, KMK16, LMH10, LK16, LWXZ10, LQW^{+}12, LWC13, Lin16, LLX^{+}11, LZKW12, LZ12, LQC^{+}14, LTW16, MH12, MMSD13, MM14, MKH^{+}12, MRBN17, MTF14, MS17b, MBPM19, MC10, MB10, MGM16, NHC13, NBH19, OCC12, OH15, OKS^{+}15, PS13, PDS19, PB11, PD16, PPN^{+}15, PXT^{+}13]. using [PCCLdGP12, PFF12, PRN17, PMB15, PWC12, QBO^{+}14, RSB^{+}14, RCCVB11, RHRC15, SCS15, SAA^{+}10, SPSR17, SRSC16, SKE10, SBZ^{+}17, SSP17, SLLY17, TJH15, TAF^{+}17, TK14, TTC18, TKP^{+}18, TLL13, TXCX19, UUN11, WRTP^{+}13, yWpNyL11, WAG15, WCX15, WL16, WLL19a, WLL19b, WKD^{+}19, Woo12, WB15, Wu11, WCB^{+}17, XZP^{+}10, XLM^{+}15, YWTW11, YSL^{+}10, YZL^{+}14, ZLW^{+}12, ZYZZ14, ZLmLN14, dOCS13, rBHM17, vHJPB^{+}17, HSS10]. utilization [BSKL10, KK17b, NZM10]. Utilizing [GSM15, LLW12, APT^{+}12, SK10, ZJZ11, SRBT18]. UX [KFN19].

v [CIB^{+}19]. validate [MM19]. Validated [SGK12]. validating [XHM^{+}11]. validation [AMGG14, CCGdL10, EZOK14, FAI13, GKV14, GTF15, HP16, HKS^{+}17, KKH^{+}16, KM13, LJH10, LMGHB17, LXJL10, LSLG17, LHP^{+}09, LHP^{+}10, MPLL18, SCMS15, SMK^{+}18, Wie14, YXP^{+}18]. validity [VHL14]. value [ASG17, APS16, CSW13, HCL12, HSS14, LMGHB17, MKS^{+}18, PCYZ12, Shi17, TC16a, VvSvV16, YWTW11]. value-based [PCYZ12]. value-oriented [LMGHB17]. valued [KLB15]. values [KK11]. VANETs [ACL13, ACSC16, WOC15]. Variability [GAMW14, APM^{+}14, CFL19, CHN19b, FFV19, FRGC10, RTM19, SRBT18, aSRS^{+}10, SHBC19, TB13, VPL^{+}10]. variable [AZ11, LWC13, WCC13, XTZX12, XTZX13]. variable-length [LWC13, XTZX12, XTZX13]. variable-order [WCC13]. variables [BHH^{+}10, MH11]. variance [HC10]. variance-controlled [HC10]. variant [CNG16]. variants [BZ10, CRC19, LNTS19, MAGC^{+}17, RBW18, SSS17]. variation [LMT16]. variations [RF14]. Varied [YWWS10]. varieties [YWTW11]. VAS [SC14]. Vector [FSS^{+}13, LBX12, PWW10, TTL10]. vectors [CKCK15]. vehicle [BKLE18]. vehicles [MSHG18]. Vehicular [ACSC16, ACL13, Cho13, HWHT11, WOC15]. vein [WLL^{+}13]. vendor [AK16, SCO13]. vendors [KNA11, RNR17]. veracity [WLL15]. Verification [CCGdL10, ABB15, BBA10, BK11, CCR14, DAR14, DBZ16, FDN^{+}16, DDF^{+}13, KK12, KSN17, LCLP16, LLL17a, LSLG17, MS17a, MA11, NTT19, PJT^{+}17, SK18]. verified [YHM^{+}14]. verifier [FWCS12, HYWS11]. Verifying [CIB^{+}19]. VERSAG [GZKL13]. Versatile [EBJ17]. version [XLL^{+}19]. versioning [RvDV17]. versions [DEW^{+}16, LNTS19]. versus [FFdRG^{+}14, GCDY16, Shi12]. Vertical [CH10d]. very [KKR16]. vetting [RDPM19]. via [ADET12, CKCK15, CdL18, GLJ13, HP16, HH17, KCT12, KM13, LT13, LPM15, PDL^{+}16, RLL^{+}18, SD16a, SPTM15, SYXL17, WL15b, WLL15, YFZ^{+}16, ZLmLN14]. video [FGBC10, LLML13, MLHL12, MK11, XLM^{+}15]. video-streaming [MLHL12]. view [CV14, DZT^{+}14, GLWY10, NI13, dMSSS^{+}13]. viewed [Kel15]. viewpoint [VCB^{+}18]. viewpoints [AAA11, GCLD13, vHJPB^{+}17]. views [Gar13]. violation [IYS13]. violations [CF12, LNW^{+}11]. Virtual [ZDC^{+}11, ABB19, AdAD17, AO16, BML^{+}13, FGLI15, GAT15, KK11, KCV11, LQW^{+}12, MCC^{+}18, NI13, SK13, WXZ^{+}17, XZZ^{+}16, ZWC^{+}19, dACM17]. Virtualization [AAJD^{+}16, LQW^{+}12, RQD^{+}17, TZB19, TSCB19]. Virtualization-based [AAJD^{+}16, TSCB19]. Virtualized [MAS13, EBJ17, GGK19, NK14, SB19]. virtually [TLWS10]. virus [HLWS13]. visible [Lin14]. vision [LWW^{+}10, NCK^{+}15]. visits [SAA^{+}10]. VISOR [KAS18]. Visual [MA10, CT11b, CLH^{+}13, DDD14, EA19, GLW13, KAS18, MGR^{+}13, WS12, YWEL^{+}13, YBE17]. visualisation [WBBK18]. visualization [JSL16, MGAN18, NSM17, PDBD18, SLB14, YLC18]. Visualizing [RF14]. VLC [HWL13b]. VLC-based [HWL13b]. VLIW [WWL^{+}10]. VM [CBZ^{+}16, LCL15]. VMM [RQD^{+}17]. VMs [XJZ^{+}15]. vocabularies [SDB18]. vocal [TKP^{+}18]. void [DBCdP11]. VoIP [hChSyCwL10]. volatile [SSAS11]. voltage [BBBP13, CS12]. voltage/frequency [CS12]. volume [LMT16]. voting [WKV11]. VQ [CNL13, YWHL11]. VQ-based [CNL13]. VR [KJ10]. VR-1 [KJ10]. VRSS [LZKW12]. VRSS-based [LZKW12]. vs [CFRPC^{+}18, ETM10, SMS11, YCG^{+}14]. vulnerabilitie [RDVC19]. vulnerabilities [MKHLB16, PDK^{+}16, STS^{+}19]. vulnerability [CMM15, HLLS13, LZKW12, RDVC19, SG16, SZ11, SA18, ZLC^{+}14]. waiting [SBZ^{+}17]. WANs [HBG^{+}14]. warehouses [FS14a, MTF14]. warnings [ANG^{+}19]. WAS [WGC^{+}14]. Waste [KM14, AKA^{+}15]. watermark [HB13, TLL13]. watermarking [AMK12, CCLL11, CT11a, CSS^{+}13, JK13, KPS10, KM11, LSR13, LXCM11, Lin14, MMSD13, MM14, MK11, PWLL13, PWW10, PKS18, TK14, TTL10, TPKT12, yWpNyL11, yWpWyYpN13]. waveform [CCWT13]. wavelet [AMK12, LXCM11, yWpNyL11, WS13]. wavelets [MMSD13]. way [WLL17]. weaving [AMKD13, HPF16, MKS10]. web [AIE19, AdB13, AAB19, BPO^{+}16, BMKM15, BAAD17, CM15, CCY11, CCH14, DH13, FMPS16, FG15, GLJ13, HYA11, LXJL10, LASL14, LSLG17, OM13, OLV15, OD17, PDK^{+}16, RAS14, RHRC13, RAJ15, SAA^{+}10, SKF17, TTC18, TPGdS13, WLL15, YLC18, ZTCZ16, AM10a, BPGS13, BLM10, BCG^{+}13, CH10b, CRESF^{+}13, DBCG14, EAH^{+}11, EZRK16, ECRVMS11, EUR^{+}13, EZG15, FAI13, GMGTdFR14, GLJ13, KM17, KR14, LKL^{+}11, LAT10, LLWL14, MIBV14, MAS13, OGK13, PÁC13, SFMB16, SBGT13, aSRS^{+}10, TTM13, VGM13, WWZ^{+}14, ZLT10, ZWM^{+}18]. web-based [OD17, YLC18, MIBV14]. web-centred [LSLG17]. web-clients [OM13]. WebAssembly [KTK19]. website [TPGdS13]. weight [HCC10a, LL14, ZGZ^{+}13, LPP15]. weight-aware [LL14]. weight-based [HCC10a, ZGZ^{+}13]. weighted [CL15, CL17b, HHK13, HR10, LLWL19, WGC^{+}14]. Where [KMG^{+}19]. whistleblowing [KP10]. white [BBEM11, KCAS13]. white-box [BBEM11, KCAS13]. Who [JLZ^{+}19]. whole [FAM15, LKJR10a, LKJR10b]. whole-part [LKJR10a, LKJR10b]. WICSA [KT16, LH12]. Wide [BPGS13, HBG^{+}13, HCB^{+}16]. wikis [OD17]. willingness [WKbOS17]. WiMAX [CTHW12]. WiMAX-MPLS [CTHW12]. window [DS12, NDS13, NBH19]. Window-based [NBH19]. windowing [MPN^{+}17]. Wireless [BRG^{+}12, AAMS16, AMAY19, Bar15, CBS16, CLY17, CW12, DBCdP11, HST15, HST16, HWHT11, HSS10, JXLC15, LLK11, LZ13, LL14, LWOY16, LWL^{+}16, LWC^{+}18, LHP^{+}09, LHP^{+}10, MLLK11, MMZ^{+}16, MDO^{+}10, MT10, MKRO14, MAAC17, NSAK10, NNVD17, OZO^{+}14, PZB10, PD12, SM17a, SMS11, SGBCP12, TKSRP11, TAF^{+}17, WH15, YCLC17, YH19, ZADM10, CDRT13, DFCPSF15]. wise [MN19, ZAY19]. within [AVGM19, GMMGP15, TJT^{+}18, TZ12, TTMI19, WR10, ZJZ11, dOSdAdSG17]. without [JCK^{+}17, RG10, SLZ12]. wizard [LSLG17]. WLANs [EZOK14, WC11]. WordNet [LBX12]. Work [FH10, Jør16, BMB19, BD10, WK15, WRR14]. Work-domain [FH10]. Work-hours [Jør16]. Workbench [FSPH^{+}16]. workdays [Jør16]. workflow [ALRP16, BJG11, KSENM17, LNW^{+}11, LNY^{+}11, MVLJ18, THWC10, WXY^{+}17, ZLD13]. workflows [CFN10, DFJ19, SHBA^{+}16]. Working [KT16, LH12, RKK16]. Workload [WWZ^{+}14, EBJ17, KCV11, PPMM17]. Workload-aware [WWZ^{+}14, EBJ17]. workloads [BSKL10, DVV^{+}16, SMSH18]. workshops [SBAH17]. world [DY15, DOL^{+}16, LCM^{+}13, LPB19, BPGS13]. wormhole [MV10, MV11]. wormhole-routed [MV10, MV11]. worst [MMTS15]. worst-case [MMTS15]. would [LK16]. WRAN [AAMS14]. Wrekavoc [CDGJ10]. write [DZT^{+}14]. WS [aSRZ^{+}18, TTC15, WKZL10]. WS-* [TTC15]. WS-BPEL [aSRZ^{+}18]. WS-CDL [WKZL10]. WSAN [TLK^{+}16a]. WSN [APS^{+}10, PAS^{+}10]. Wu [LLLK10].

X [BAI^{+}14]. X-ray [BAI^{+}14]. XACML [CH10b]. Xen [CBZ^{+}16]. Xen-based [CBZ^{+}16]. Xen2MX [NK14]. Xenomai [DYC19]. Xeon [DSGS17, GGK19]. XML [CH11, CDOP15, CJL11, GLWY10, HR10, KSKP11, LWXZ10, MCTM11, MIUM12, NKMM12, NTRN11, PDK^{+}16, SM17a, TLWS10]. XML-based [NKMM12]. XML-manipulating [MCTM11, MCTM11]. XQuery [PDK^{+}16]. Xtraitj [BD17].

years [CJT^{+}16, DFG^{+}13, KQ17,

dMSSS^{+}13, VCdA^{+}16]. Yen [LLLK10].

Yugoslavia [SNDC13].

Zero [LESL11, TLL13]. Zero-laxity [LESL11]. zero-watermark [TLL13]. Zodiac [SDM10].

References

Arias:2011:DDE

[AAA11] Trosky B. Callo Arias, Pierre America, and Paris Avgeriou. Defining and documenting execution viewpoints for a large and complex software-intensive system. The Journal of Systems and Software, 84 (9):1447–1461, September 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                                         //www.sciencedirect.
                                                         com/science/article/
                                                         pii/S016412121000316X.

Andrews:2019:BBM

[AAB19] Anneliese Andrews, Ahmed Alhaddad, and Salah Boukhris. Black-box model-based regression testing of fail-safe behavior in web applications. The Journal of Systems and Software, 149(??):318–339, March 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302541.

Alegre:2016:ECA

[AAC16] Unai Alegre, Juan Carlos Augusto, and Tony Clark. Engineering context-aware systems and applications: a survey. The Journal of Systems and Software, 117(??):55–83, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000467.

Arvanitou:2017:MSD

[AAC^{+}17] Elvira Maria Arvanitou, Apostolos Ampatzoglou, Alexander Chatzigeorgiou, Matthias Galster, and Paris Avgeriou. A mapping study on design-time quality attributes and metrics. The Journal of Systems and Software, 127(??):52–77, May 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121730016X.

Abebe:2013:SCL

[AACT13] Surafel Lemma Abebe, Anita Alicante, Anna Corazza, and Paolo Tonella. Supporting concept location through identifier parsing and ontology extraction. The Journal of Systems and Software, 86(11):2919– 2938, November 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213001702.

Afzal:2016:STP

[AAGT16] Wasif Afzal, Snehal Alone, Kerstin Glocksien, and Richard Torkar. Software test process improvement approaches: a systematic literature review and an industrial case study. The Journal of Systems and Software, 111(??):1–33, January 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0164121215001910.

Alnas:2010:PEF

[AAH10] Mohamed Alnas, Irfan Awan, and R. D. W. Holton. Performance evaluation of fast handover in mobile IPv6 based on link-layer information. The Journal of Systems and Software, 83(10):1644–1650, October 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). See erratum [AAH12b].

Alam:2012:PKS

[AAH12a] Omar Alam, Bram Adams, and Ahmed E. Hassan. Preserving knowledge in software projects. The Journal of Systems and Software, 85 (10):2318–2330, October 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000787.

Alnas:2012:EPE

[AAH12b] Mohamed Alnas, Irfan Awan, and R. D. W. Holton. Erratum to “Performance evaluation of fast handover in mobile IPv6 based on link layer informations” [J.

Syst. Softw. 83 (2010) 1644–1650]. The Journal of Systems and Software, 85(1):61, January 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001816.

See [AAH10].

Al-Ayyoub:2016:VBC

[AAJD^{+}16] Mahmoud Al-Ayyoub, Yaser Jararweh, Ahmad Doulat, Haythem A. Bany Salameh, Ahmad Al Abed Al Aziz, Mohammad Alsmirat, and Abdallah A. Khreishah. Virtualization-based Cognitive Radio Networks. The Journal of Systems and Software, 117(??): 15–29, July 2016. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121216000479.

Almugrin:2016:UIC

[AAM16] Saleh Almugrin, Waleed Albattah, and Austin Melton. Using indirect coupling metrics to predict package maintainability and testability. The Journal of Systems and Software, 121(??):298–310, November 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121600056X.

Amalfitano:2017:GFC

[AAM^{+}17] Domenico Amalfitano, Nicola Amatucci, Atif M. Memon, Porfirio Tramontana, and Anna Rita Fasolino. A general framework for comparing automatic testing techniques of Android mobile apps. The Journal of Systems and Software, 125(??):322–343, March 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL /

/www.sciencedirect.
com/science/article/
pii/S016412121630259X.

Afzal:2014:MAC

[AAMS14] Humaira Afzal, Irfan Awan, Mohammed Rafiq Mufti, and Ray E. Sheriff. Modeling and analysis of customer premise equipments registration process in IEEE 802.22 WRAN cell. The Journal of Systems and Software, 98(??):107–116, December 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214001873.

Afzal:2016:PAC

[AAMS16] Humaira Afzal, Irfan Awan, Mohammed Rafiq Mufti, and Ray E. Sheriff. Performance analysis of contending customer equipment in wireless networks. The Journal of Systems and Software, 117(??):357–365, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216300115.

Alvarez:2011:ICL

[AAN11] Claudio Alvarez, Rosa Alarcon, and Miguel Nussbaum. Implementing collaborative learning activities in the classroom supported by oneto-one mobile computing: a design-based process. The Journal of Systems and Software, 84 (11):1961–1976, November 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211001865.

Appari:2010:MPS

[AB10] Ajit Appari and Michel Benaroch. Monetary pricing of software development risks: a method and empirical illustration. The Journal of Systems and Software, 83 (11):2098–2107, November 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ahmad:2016:SAR

[AB16] Aakash Ahmad and Muhammad Ali Babar. Software architectures for robotic systems: a systematic mapping study. The Journal of Systems and Software, 122(??):16–39, December 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301479.

Abawajy:2013:SDP

[Aba13] Jemal Abawajy. SQLIA detection and prevention approach for RFID systems. The Journal of Systems and Software, 86(3):751–758, March 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121212003159.

Ahmad:2015:MVF

[ABB15] Manzoor Ahmad, Nicolas Belloir, and Jean-Michel Bruel. Modeling and verification of functional and non-functional requirements of ambient self-adaptive systems. The Journal of Systems and Software, 107(??):50–70, September 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215001065.

Abeni:2019:HSR

[ABB19] Luca Abeni, Alessandro Biondi, and Enrico Bini. Hierarchical scheduling of real-time tasks over Linux-based virtual machines. The Journal of Systems and Software, 149(??):234–249, March 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121830270X.

Anand:2013:OSM

[ABC^{+}13] Saswat Anand, Edmund K. Burke, Tsong Yueh Chen, John Clark, Myra B.

Cohen, Wolfgang Grieskamp, Mark Harman, Mary Jean Harrold, Phil McMinn, Orchestrators, and Editors. An orchestrated survey of methodologies for automated software test case generation. The Journal of Systems and Software, 86(8):1978–2001, August 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000563.

AbouTrab:2013:TRT

[ABCH13] M. Saeed AbouTrab, Michael Brockway, Steve Counsell, and Robert M. Hierons. Testing Real-Time Embedded Systems using Timed Automata based approaches. The Journal of Systems and Software, 86(5):1209– 1223, May 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212003391.

Avvenuti:2012:JTC

[ABFM12] Marco Avvenuti, Cinzia Bernardeschi, Nicoletta De Francesco, and Paolo Masci. JCSI: a tool for checking secure information flow in Java Card applications. The Journal of Systems and Software, 85(11):2479–2493, November 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212001513.

Arisholm:2010:SCI

[ABJ10] Erik Arisholm, Lionel C. Briand, and Eivind B. Johannessen. A systematic and comprehensive investigation of methods to build and evaluate fault prediction models. The Journal of Systems and Software, 83(1):2–17, January 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Astromskis:2017:PDB

[ABJ^{+}17] Saulius Astromskis, Gabriele Bavota, Andrea Janes, Barbara Russo, and Massimiliano Di Penta. Patterns of developers behaviour: a 1000-hour industrial study. The Journal of Systems and Software, 132(??):85–97, October 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S016412121730136X.

Arcangeli:2015:ADD

[ABL15] Jean-Paul Arcangeli, Raja Boujbel, and Sébastien Leriche. Automatic deployment of distributed software systems: Definitions and state of the art. The Journal of Systems and Software, 103(??):198–218, May 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000308.

Andrews:2016:TBH

[ABL16] Anneliese Amschler Andrews, Philip Beaver, and Joseph Lucente. Towards better help desk planning: Predicting incidents and required effort. The Journal of Systems and Software, 117(??):426–449, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300127.

Alpuente:2019:SCM

[ABS19] M. Alpuente, D. Ballis, and J. Sapiña. Static correction of Maude programs with assertions. The Journal of Systems and Software, 153(??):64–85, July 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300640.

Andreou:2016:SDP

[AC16] Andreas S. Andreou and Sotirios P. Chatzis. Software defect prediction using doubly stochastic Poisson processes driven by stochastic belief networks. The Journal of Systems and Software, 122(??):72–82, December 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301601.

Ajienka:2017:UIB

[AC17] Nemitari Ajienka and Andrea Capiluppi. Understanding the interplay between the logical and structural coupling of software classes. The Journal of Systems and Software, 134 (??):120–137, December 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121730184X.

Acher:2019:SIS

[AC19] Mathieu Acher and Myra B. Cohen. Special issue on systems and software product line engineering. The Journal of Systems and Software, 154(??):110–111, August 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300950.

Akbarinasaji:2018:PBF

[ACB18] Shirin Akbarinasaji, Bora Caglayan, and Ayse Bener. Predicting bugfixing time: a replication study using an open source software project. The Journal of Systems and Software, 136(??):173–186, February 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217300365.

Albuquerque:2015:QUD

[ACG^{+}15] Diego Albuquerque, Bruno Cafeo, Alessandro Garcia, Simone Barbosa, Silvia Abrahão, and António Ribeiro. Quantifying usability of domainspecific languages: an empirical study on software maintenance. The Journal of Systems and Software, 101(??):245– 259, March 2015. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121214002799.

Amiri-Chimeh:2019:ASN

[ACH19] Saeed Amiri-Chimeh and Hassan Haghighi. An approach to solving nonlinear real constraints for symbolic execution. The Journal of Systems and Software, 157(??):??, November 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301505.

Ali:2013:SRT

[ACL13] G. G. Md. Nawaz Ali, Edward Chan, and Wenzhong Li. Supporting real-time multiple data items query in multi-RSU vehicular ad hoc networks (VANETs). The Journal of Systems and Software, 86 (8):2127–2142, August 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000769.

Albert:2019:PRA

[ACRD19] Elvira Albert, Jesús Correas, and Guillermo Román-Díez. Peak resource analysis of concurrent distributed systems. The Journal of Systems and Software, 149(??):35–62, March 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302528.

Ampatzoglou:2013:RSA

[ACS13] Apostolos Ampatzoglou, Sofia Charalampidou, and Ioannis Stamelos. Research state of the art on GoF design patterns: a mapping study. The Journal of Systems and Software, 86(7):1945– 1964, July 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000757.

Ali:2016:EDD

[ACSC16] G. G. Md. Nawaz Ali, Peter Han Joo Chong, Syeda Khairunnesa Samantha, and Edward Chan. Efficient data dissemination in cooperative multi-RSU Vehicular Ad Hoc Networks (VANETs). The Journal of Systems and Software, 117(??): 508–527, July 2016. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216300218.

Avritzer:2010:MOR

[ACW10] Alberto Avritzer, Robert G. Cole, and Elaine J. Weyuker. Methods and opportunities for rejuvenation in aging distributed software systems. The Journal of Systems and Software, 83 (9):1568–1578, September 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Alsoghayer:2014:RFR

[AD14] Raid Alsoghayer and Karim Djemame. Resource failures risk assessment modelling in distributed environments. The Journal of Systems and Software, 88 (??):42–53, February 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002331.

Alves:2017:MQM

[AdAD17] Maicon Melo Alves and Lúcia Maria de Assumpção Drummond. A multivariate and quantitative model for predicting cross-application interference in virtual environments. The Journal of Systems and Software, 128(??):150–163, June 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217300699.

Agustin:2013:MDA

[AdB13] José Luis Herrero Agustin and Pablo Carmona del Barco. A model-driven approach to develop high performance web applications. The Journal of Systems and Software, 86 (12):3013–3023, December 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001751.

Angelov:2017:DAA

[AdB17] S. Angelov and P. de Beer. Designing and applying an approach to software architecting in agile projects in education. The Journal of Systems and Software, 127(??):78–90, May 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217300195.

Ahmad:2018:KSE

[ADCO18] Muhammad Ovais Ahmad, Denis Dennehy, Kieran Conboy, and Markku Oivo. Kanban in software engineering: a systematic mapping study. The Journal of Systems and Software, 137(??):96–113, March 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217302820.

Amoui:2012:ADA

[ADET12] Mehdi Amoui, Mahdi Derakhshanmanesh, Jürgen Ebert, and Ladan Tahvildari. Achieving dynamic adaptation via management and interpretation of runtime models. The Journal of Systems and Software, 85 (12):2720–2737, December 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001458.

Al-Dubai:2010:SIP

[ADMOK^{+}10] Ahmed Y. Al-Dubai, Geyong Min, Mohamed Ould-Khaoua, Xiaolong Jin, and William Buchanan. Special issue: Performance evaluation and optimization of ubiquitous computing and networked systems. The Journal of Systems and Software, 83(8):1299– 1300, August 2010. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Abate:2012:DSS

[ADTZ12] Pietro Abate, Roberto Di Cosmo, Ralf Treinen, and Stefano Zacchiroli. Dependency solving: a separate concern in component evolution management. The Journal of Systems and Software, 85(10):2228–2240, October 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121212000477.

Andreas:2016:TDD

[AF16] Jede Andreas and Teuteberg Frank. Towards a document-driven approach for designing reference models: From a conceptual process model to its application. The Journal of Systems and Software, 111(??):254–269, January 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215002137.

Aleti:2015:TDG

[AG15] Aldeida Aleti and Lars Grunske. Test data generation with a Kalman filter-based adaptive genetic algorithm. The Journal of Systems and Software, 103(??):343– 352, May 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214002660.

Anh:2014:MPA

[AGBD14] Dinh Tien Tuan Anh, Milind Ganjoo, Stefano Braghin, and Anwitaman Datta. Mosco: a privacyaware middleware for mobile social computing. The Journal of Systems and Software, 92(??):20– 31, June 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002835.

Alvarez-Garcia:2014:AMP

[ÁGBYB^{+}14] Sandra Álvarez-García, Ricardo Baeza-Yates, Nieves R. Brisaboa, Josep-Lluis Larriba-Pey, and Oscar Pedreira. Automatic multi-partite graph generation from arbitrary data. The Journal of Systems and Software, 94(??):72–86, August 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214000715.

Amin:2013:ASR

[AGC13] Ayman Amin, Lars Grunske, and Alan Colman. An approach to software reliability prediction based on time series modeling. The Journal of Systems and Software, 86(7):1923–1932, July 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213000617.

Arcaini:2019:ACR

[AGR19] Paolo Arcaini, Angelo Gargantini, and Marco Radavelli. Achieving change requirements of feature models by an evolutionary approach. The Journal of Systems and Software, 150(??):64–76, April 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219300056.

Alkaf:2019:ACI

[AHBA19] Hasan Alkaf, Jameleddine Hassine, Taha Binalialhag, and Daniel Amyot. An automated change impact analysis approach for User Requirements Notation models. The Journal of Systems and Software, 157(??):??, November 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301724.

Ayala:2011:STP

[AHC^{+}11] Claudia Ayala, Øyvind Hauge, Reidar Conradi, Xavier Franch, and Jingyue Li. Selection of third party software in off-the-shelf-based software development — an interview study with industrial practitioners. The Journal of Systems and Software, 84(4):620– 637, April 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ahamed:2010:DAD

[AHH^{+}10] Sheikh I. Ahamed, Munirul M. Haque, Md. Endadul Hoque, Farzana Rahman, and Nilothpal Talukder. Design, analysis, and deployment of omnipresent Formal Trust Model (FTM) with trust bootstrapping for pervasive environments. The Journal of Systems and Software, 83(2):253– 270, February 2010. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Alawneh:2016:SLT

[AHLH16] Luay Alawneh, Abdelwahab Hamou-Lhadj, and Jameleddine Hassine. Segmenting large traces of inter-process communication with a focus on high performance computing systems. The Journal of Systems and Software, 120(??):1–16, October 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216300954.

Antunes:2014:RQA

[AHOP14] Pedro Antunes, Valeria Herskovic, Sergio F. Ochoa, and José A. Pino. Reviewing the quality of awareness support in collaborative applications. The Journal of Systems and Software, 89(??):146–169, March 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213002756.

Ahmad:2010:PNM

[AHW10] Farooq Ahmad, Hejiao Huang, and Xiao-Long Wang. Petri net modeling and deadlock analysis of parallel manufacturing processes with shared-resources. The Journal of Systems and Software, 83(4):675–688, April 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Abdullah:2019:ULA

[AIE19] Muhammad Abdullah, Waheed Iqbal, and Abdelkarim Erradi. Unsupervised learning approach for web application auto-decomposition into microservices. The Journal of Systems and Software, 151(??):243– 257, May 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300408.

Abdelmaboud:2015:QSA

[AJG^{+}15] Abdelzahir Abdelmaboud, Dayang N. A. Jawawi, Imran Ghani, Abubakar Elsafi, and Barbara Kitchenham. Quality of service approaches in cloud computing: a systematic mapping study. The Journal of Systems and Software, 101(??):159–179, March 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002830.

Abrahao:2010:IBU

[AJLS10] Silvia Abrahão, Natalia Juristo, Effie L.-C. Law, and Jan Stage. Interplay between usability and software development. The Journal of Systems and Software, 83 (11):2015–2018, November 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Athanasopoulos:2015:ERR

[AK15] Michael Athanasopoulos and Kostas Kontogiannis. Extracting REST resource models from procedure-oriented service interfaces. The Journal of Systems and Software, 100(??):149–166, February 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214002362.

Ali:2016:SOP

[AK16] Sikandar Ali and Siffat Ullah Khan. Software outsourcing partnership model: an evaluation framework for vendor organizations. The Journal of Systems and Software, 117(??):402– 425, July 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121630019X.

Anagnostopoulos:2015:ADM

[AKA^{+}15] Theodoros Anagnostopoulos, Kostas Kolomvatsos, Christos Anagnostopoulos, Arkady Zaslavsky, and Stathes Hadjiefthymiades. Assessing dynamic models for high priority waste collection in smart cities. The Journal of Systems and Software, 110(??):178–192, December 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001922.

Alabool:2018:CSE

[AKAA18] Hamzeh Alabool, Ahmad Kamil, Noreen Arshad, and Deemah Alarabiat. Cloud service evaluation method-based Multi-Criteria Decision-Making: A systematic literature review. The Journal of Systems and Software, 139(??):161– 188, May 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300244.

Akritidis:2011:ERA

[AKB11] Leonidas Akritidis, Dimitrios Katsaros, and Panayiotis Bozanis. Effective rank aggregation for metasearching. The Journal of Systems and Software, 84(1):130–143, January 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Adolph:2012:RPG

[AKH12] Steve Adolph, Philippe Kruchten, and Wendy Hall. Reconciling perspectives: a grounded theory of how people manage the process of software development. The Journal of Systems and Software, 85(6):1269–1286, June 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212000386.

Akiki:2018:CDM

[Aki18] Pierre A. Akiki. CHAIN: Developing model-driven contextual help for adaptive user interfaces. The Journal of Systems and Software, 135(??):165– 190, January 2018. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217302509.

Ampatzoglou:2011:EIR

[AKKS11] Apostolos Ampatzoglou, Apostolos Kritikos, George Kakarontzas, and Ioannis Stamelos. An empirical investigation on the reusability of design patterns and software packages. The Journal of Systems and Software, 84 (12):2265–2283, December 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001592.

Alsawalqah:2014:MOS

[AKL14] Hamad I. Alsawalqah, Sungwon Kang, and Jihyun Lee. A method to optimize the scope of a software product platform based on enduser features. The Journal of Systems and Software, 98(??):79–106, December 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001861.

AlDallal:2012:IAS

[Al 12] Jehad Al Dallal. The impact of accounting for special methods in the measurement of objectoriented class cohesion on refactoring and fault prediction activities. The Journal of Systems and Software, 85(5):1042– 1057, May 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211003001.

Andrzejczak:2010:ETL

[AL10] Chris Andrzejczak and Dahai Liu. The effect of testing location on usability testing performance, participant stress levels, and subjective testing experience. The Journal of Systems and Software, 83(7):1258– 1266, July 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Alazab:2015:PCB

[Ala15] Mamoun Alazab. Profiling and classifying the behavior of malicious codes. The Journal of Systems and Software, 100(??):91–102, February 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214002283.

Alkhanak:2016:COA

[ALRP16] Ehab Nabiel Alkhanak, Sai Peck Lee, Reza Rezaei, and Reza Meimandi Parizi. Cost optimization approaches for scientific workflow scheduling in cloud and grid computing: a review, classifications, and open issues. The Journal of Systems and Software, 113(??):1– 26, March 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002484.

Ardagna:2010:PFO

[AM10a] Danilo Ardagna and Raffaela Mirandola. Per-flow optimal service selection for Web services based processes. The Journal of Systems and Software, 83(8):1512–1523, August 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Athanasiadis:2010:DPP

[AM10b] Erast Athanasiadis and Sarandis Mitropoulos. A distributed platform for personalized advertising in digital interactive TV environments. The Journal of Systems and Software, 83(8):1453–1469, August 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Abdullah:2013:MPF

[AM13] Ibrahim S. Abdullah and Daniel A. Menascé. The Meta-Protocol framework. The Journal of Systems and Software, 86 (11):2711–2724, November 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213001386.

Alho:2015:SOA

[AM15] Pekka Alho and Jouni Mattila. Service-oriented approach to fault tolerance in CPSs. The Journal of Systems and Software, 105(??):1–17, July 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215000643.

Alhammad:2018:GSE

[AM18] Manal M. Alhammad and Ana M. Moreno. Gamification in software engineering education: a systematic mapping. The Journal of Systems and Software, 141(??):131– 150, July 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300645.

Afzal:2019:PAR

[AMAY19] Humaira Afzal, Muhammad Rafiq Mufti, Irfan Awan, and Muhammad Yousaf. Performance analysis of radio spectrum for cognitive radio wireless networks using discrete time Markov chain. The Journal of Systems and Software, 151(??):1–7, May 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300068.

Allison:2014:SID

[AMCC14] Mark Allison, Karl A.

Morris, Fábio M. Costa, and Peter J. Clarke. Synthesizing interpreted domain-specific models to manage smart microgrids. The Journal of Systems and Software, 96(??):172–193, October 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121214001435.

Alves:2017:TCI

[AMdLM17] Everton L. G. Alves, Tiago Massoni, and Patrícia Duarte de Lima Machado. Test coverage of impacted code elements for detecting refactoring faults: an exploratory study. The Journal of Systems and Software, 123(??):223–238, January 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216000388.

Asadi:2014:DVC

[AMGG14] Mohsen Asadi, Bardia Mohabbati, Gerd Gröner, and Dragan Gasevic. Development and validation of customized process models. The Journal of Systems and Software, 96(??):73–92, October 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214001344.

Arsalan:2012:IRW

[AMK12]  Muhammad  Arsalan,

Sana Ambreen Malik, and Asifullah Khan. Intelligent reversible watermarking in integer wavelet domain for medical images. The Journal of Systems and Software, 85(4):883–894, April 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002858.

Alhadidi:2013:CWA

[AMKD13] Dima Alhadidi, Azzam Mourad, Hakim Idrissi Kaitouni, and Mourad Debbabi. Common weaving approach in mainstream languages for software security hardening. The Journal of Systems and Software, 86(10):2654–2674, October 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001325.

Abeni:2012:ERP

[AMP12] Luca Abeni, Nicola Manica, and Luigi Palopoli. Efficient and robust probabilistic guarantees for real-time tasks. The Journal of Systems and Software, 85(5):1147– 1156, May 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211003232.

Ahmadian:2010:PDS

[AMS^{+}10] Zahra Ahmadian, Javad Mohajeri, Mahmoud Salmasizadeh, Risto M. Hakala, and Kaisa Nyberg. A practical distinguisher for the Shannon cipher. The Journal of Systems and Software, 83 (4):543–547, April 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Angelopoulos:2010:ACS

[AN10] Constantinos Marios Angelopoulos and Sotiris Nikoletseas. Accelerated collection of sensor data by mobility-enabled topology ranks. The Journal of Systems and Software, 83(12):2471– 2477, December 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ahmed:2016:MAB

[AN16] Salman Ahmed and Aamer Nadeem. A mobile agent based communication protocol to optimize message delivery cost. The Journal of Systems and Software, 121(??):40–48, November 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301388.

Azzeh:2011:ABS

[ANC11] Mohammad Azzeh, Daniel Neagu, and Peter I. Cowling. Analogy-based software effort estimation using Fuzzy numbers. The Journal of Systems and Software, 84(2): 270–284, February 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Aloraini:2019:ESS

[ANG^{+}19] Bushra Aloraini, Meiyappan Nagappan, Daniel M. German, Shinpei Hayashi, and Yoshiki Higo. An empirical study of security warnings from static application security testing tools. The Journal of Systems and Software, 158(??):??, December 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219302018.

Azzeh:2015:EEE

[ANM15] Mohammad Azzeh, Ali Bou Nassif, and Leandro L.

Minku. An empirical evaluation of ensemble adjustment methods for analogy-based effort estimation. The Journal of Systems and Software, 103(??):36–52, May 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215000199.

Anonymous:2011:EBa

[Ano11a] Anonymous. Editorial Board. The Journal of Systems and Software, 84(1):??, January 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2011:EBb

[Ano11b] Anonymous. Editorial Board. The Journal of Systems and Software, 84(2):??, February 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2011:EBc

[Ano11c] Anonymous. Editorial Board. The Journal of Systems and Software, 84 (3):??, March 2011. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2011:EBd

[Ano11d] Anonymous. Editorial Board. The Journal of Systems and Software, 84 (4):??, April 2011. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2011:EBe

[Ano11e] Anonymous. Editorial Board. The Journal of Systems and Software, 84 (5):??, May 2011. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2011:EBf

[Ano11f] Anonymous. Editorial Board. The Journal of Systems and Software, 84 (6):??, June 2011. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2011:EBg

[Ano11g] Anonymous. Editorial Board. The Journal of Systems and Software, 84 (7):??, July 2011. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2011:EBh

[Ano11h] Anonymous. Editorial Board. The Journal of Systems and Software, 84 (8):??, August 2011. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2011:EBi

[Ano11i] Anonymous. Editorial Board. The Journal of Systems and Software, 84(9):??, September 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001725.

Anonymous:2011:EBj

[Ano11j] Anonymous. Editorial Board. The Journal of Systems and Software, 84(10):??, October 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001890.

Anonymous:2011:EBk

[Ano11k] Anonymous. Editorial Board. The Journal of Systems and Software, 84(11):??, November 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002196.

Anonymous:2011:EBl

[Ano11l] Anonymous. Editorial Board. The Journal of Systems and Software, 84(12):??, December 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002469.

Anonymous:2011:PN

[Ano11m] Anonymous. Publisher’s note. The Journal of Systems and Software, 84(3): 525, March 2011. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2012:EBa

[Ano12a] Anonymous. Editorial Board. The Journal of Systems and Software, 85(1):??, January 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002639.

Anonymous:2012:EBb

[Ano12b] Anonymous. Editorial Board. The Journal of Systems and Software, 85(2):??, February 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211003074.

Anonymous:2012:EBc

[Ano12c] Anonymous. Editorial Board. The Journal of Systems and Software, 85(3):??, March 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000118.

Anonymous:2012:EBd

[Ano12d] Anonymous. Editorial Board. The Journal of Systems and Software, 85(4):??, April 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000283.

Anonymous:2012:EBe

[Ano12e] Anonymous. Editorial Board. The Journal of Systems and Software, 85(5):??, May 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S016412121200057X.

Anonymous:2012:EBf

[Ano12f] Anonymous. Editorial Board. The Journal of Systems and Software, 85(6):??, June 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000878.

Anonymous:2012:EBg

[Ano12g] Anonymous. Editorial Board. The Journal of Systems and Software, 85(7):??, July 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001124.

Anonymous:2012:EBh

[Ano12h] Anonymous. Editorial Board. The Journal of Systems and Software, 85(8):??, August 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001392.

Anonymous:2012:EBi

[Ano12i] Anonymous. Editorial Board. The Journal of Systems and Software, 85(9):??, September 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212001793.

Anonymous:2012:EBj

[Ano12j] Anonymous. Editorial Board. The Journal of Systems and Software, 85(10):??, October 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212001963.

Anonymous:2012:EBk

[Ano12k] Anonymous. Editorial Board. The Journal of Systems and Software, 85(11):??, November 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121200235X.

Anonymous:2012:EBl

[Ano12l] Anonymous. Editorial Board. The Journal of Systems and Software, 85(12):??, December 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002580.

Anonymous:2013:CIA

[Ano13a] Anonymous. The crosscutting impact of the AOSD Brazilian research community. The Journal of Systems and Software, 86(4):905–933, April 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002427.

Anonymous:2013:EBa

[Ano13b] Anonymous. Editorial Board. The Journal of Systems and Software, 86(1):??, January 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002919.

Anonymous:2013:EBb

[Ano13c] Anonymous. Editorial Board. The Journal of Systems and Software, 86(2):??, February 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212003469.

Anonymous:2013:EBc

[Ano13d] Anonymous. Editorial Board. The Journal of Systems and Software, 86(3):??, March 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213000071.

Anonymous:2013:EBd

[Ano13e] Anonymous. Editorial Board. The Journal of Systems and Software, 86(4):??, April 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121300040X.

Anonymous:2013:EBe

[Ano13f] Anonymous. Editorial Board. The Journal of Systems and Software, 86(5):??, May 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000630.

Anonymous:2013:EBf

[Ano13g] Anonymous. Editorial Board. The Journal of Systems and Software, 86(6):??, June 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000939.

Anonymous:2013:EBg

[Ano13h] Anonymous. Editorial Board. The Journal of Systems and Software, 86(7):??, July 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001106.

Anonymous:2013:EBh

[Ano13i] Anonymous. Editorial Board. The Journal of Systems and Software, 86(8):??, August 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001520.

Anonymous:2013:EBi

[Ano13j] Anonymous. Editorial Board. The Journal of Systems and Software, 86(9):??, September 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213001829.

Anonymous:2013:EBj

[Ano13k] Anonymous. Editorial Board. The Journal of Systems and Software, 86(10):??, October 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213002057.

Anonymous:2013:EBk

[Ano13l] Anonymous. Editorial Board. The Journal of Systems and Software, 86(11):??, November 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213002203.

Anonymous:2014:EBa

[Ano14a] Anonymous. Editorial Board. The Journal of Systems and Software, 94(??):IFC, August 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001253.

Anonymous:2014:EBb

[Ano14b] Anonymous. Editorial Board. The Journal of Systems and Software, 95(??):IFC, September 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001617.

Anonymous:2014:EBc

[Ano14c] Anonymous. Editorial Board. The Journal of Systems and Software, 96(??):IFC, October 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001770.

Anonymous:2014:EBd

[Ano14d] Anonymous. Editorial Board. The Journal of Systems and Software, 97(??):IFC, November 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121400199X.

Anonymous:2014:EBe

[Ano14e] Anonymous. Editorial Board. The Journal of Systems and Software, 98(??):IFC, December 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214002209.

Anonymous:2015:EBa

[Ano15a] Anonymous. Editorial Board. The Journal of Systems and Software, 99(??):IFC, January 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214002453.

Anonymous:2015:EBb

[Ano15b] Anonymous. Editorial Board. The Journal of Systems and Software, 100(??):IFC, February 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S016412121400257X.

Anonymous:2015:EBc

[Ano15c]  Anonymous.  Edito-

rial Board. The Journal of Systems and Software, 101(??):IFC, March 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000059.

Anonymous:2015:EBd

[Ano15d] Anonymous. Editorial Board. The Journal of Systems and Software, 102(??):IFC, April 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000230.

Anonymous:2015:EBe

[Ano15e] Anonymous. Editorial Board. The Journal of Systems and Software, 103(??):ifc, May 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000515.

Anonymous:2015:EBf

[Ano15f] Anonymous. Editorial Board. The Journal of Systems and Software, 104(??):ifc, June 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215000734.

Anonymous:2015:EBg

[Ano15g] Anonymous. Editorial Board. The Journal of Systems and Software, 105(??):ifc, July 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215000904.

Anonymous:2015:EBh

[Ano15h] Anonymous. Editorial Board. The Journal of Systems and Software, 106(??):ifc, August 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215001107.

Anonymous:2015:EBi

[Ano15i] Anonymous. Editorial Board. The Journal of Systems and Software, 107(??):ifc, September 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001363.

Anonymous:2015:EBj

[Ano15j] Anonymous. Editorial Board. The Journal of Systems and Software, 109(??):ifc, November 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002046.

Anonymous:2015:EBk

[Ano15k] Anonymous. Editorial Board. The Journal of Systems and Software, 110(??):ifc, December 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002186.

Anonymous:2016:EBa

[Ano16a] Anonymous. Editorial Board. The Journal of Systems and Software, 111(??):ifc, January 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215002526.

Anonymous:2016:EBb

[Ano16b] Anonymous. Editorial Board. The Journal of Systems and Software, 112(??):ifc, February 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215002733.

Anonymous:2016:EBc

[Ano16c] Anonymous. Editorial Board. The Journal of Systems and Software, 113(??):ifc, March 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216000108.

Anonymous:2016:EBd

[Ano16d] Anonymous. Editorial Board. The Journal of Systems and Software, 114(??):ifc, April 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S016412121600025X.

Anonymous:2016:EBe

[Ano16e] Anonymous. Editorial Board. The Journal of Systems and Software, 115(??):ifc, May 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000686.

Anonymous:2016:EBf

[Ano16f] Anonymous. Editorial Board. The Journal of Systems and Software, 116(??):ifc, June 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000625.

Anonymous:2016:EBg

[Ano16g] Anonymous. Editorial Board. The Journal of Systems and Software, 117(??):ifc, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300474.

Anonymous:2016:EBh

[Ano16h] Anonymous. Editorial Board. The Journal of Systems and Software, 118(??):ifc, August 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121630067X.

Anonymous:2016:EBi

[Ano16i] Anonymous. Editorial Board. The Journal of Systems and Software, 119(??):ifc, September 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216301108.

Anonymous:2016:EBj

[Ano16j] Anonymous. Editorial Board. The Journal of Systems and Software, 120(??):ifc, October 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216301406.

Anonymous:2016:EBk

[Ano16k] Anonymous. Editorial Board. The Journal of Systems and Software, 121(??):ifc, November 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301650.

Anonymous:2016:EBl

[Ano16l] Anonymous. Editorial Board. The Journal of Systems and Software, 122(??):ifc, December 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216302047.

Anonymous:2017:EBa

[Ano17a] Anonymous. Editorial Board. The Journal of Systems and Software, 123(??):ifc, January 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216302187.

Anonymous:2017:EBb

[Ano17b] Anonymous. Editorial Board. The Journal of Systems and Software, 124(??):ifc, February 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216302436.

Anonymous:2017:EBc

[Ano17c] Anonymous. Editorial Board. The Journal of Systems and Software, 125(??):ifc, March 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217300079.

Anonymous:2017:EBd

[Ano17d] Anonymous. Editorial Board. The Journal of Systems and Software, 126(??):ifc, April 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217300249.

Anonymous:2017:EBe

[Ano17e] Anonymous. Editorial Board. The Journal of Systems and Software, 127(??):ifc, May 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217300481.

Anonymous:2017:EBf

[Ano17f] Anonymous. Editorial Board. The Journal of Systems and Software, 128(??):ifc, June 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217300730.

Anonymous:2017:EBg

[Ano17g] Anonymous. Editorial Board. The Journal of Systems and Software, 129(??):ifc, July 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217300894.

Anonymous:2017:EBh

[Ano17h] Anonymous. Editorial Board. The Journal of Systems and Software, 130(??):ifc, August 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301139.

Anonymous:2017:EBi

[Ano17i] Anonymous. Editorial Board. The Journal of Systems and Software, 131(??):ifc, September 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301425.

Anonymous:2017:EBj

[Ano17j] Anonymous. Editorial Board. The Journal of Systems and Software, 132(??):ifc, October 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301693.

Anonymous:2017:EBk

[Ano17k] Anonymous. Editorial Board. The Journal of Systems and Software, 133(??):ifc, November 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301905.

Anonymous:2017:EB

[Ano17l] Anonymous. Editorial Board. The Journal of Systems and Software, 134(??):ifc, December 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302431.

Anonymous:2017:PN

[Ano17m] Anonymous. Publisher’s note. The Journal of Systems and Software, 130(??):113, August 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301115.

Anonymous:2018:EBa

[Ano18a] Anonymous. Editorial Board. The Journal of Systems and Software, 135(??):ifc, January 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302637.

Anonymous:2018:EBb

[Ano18b] Anonymous. Editorial Board. The Journal of Systems and Software, 136(??):ifc, February 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121217302868.

Anonymous:2018:EBc

[Ano18c] Anonymous. Editorial Board. The Journal of Systems and Software, 138(??):ii, April 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218300190.

Anonymous:2018:EBd

[Ano18d] Anonymous. Editorial Board. The Journal of Systems and Software, 139(??):ii, May 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218300384.

Anonymous:2018:EBe

[Ano18e] Anonymous. Editorial Board. The Journal of Systems and Software, 140(??):ii, June 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218300578.

Anonymous:2018:EBf

[Ano18f] Anonymous. Editorial Board. The Journal of Systems and Software, 141(??):ii, July 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300785.

Anonymous:2018:EBg

[Ano18g] Anonymous. Editorial Board. The Journal of Systems and Software, 142(??):ii, August 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301006.

Anonymous:2018:EBh

[Ano18h] Anonymous. Editorial Board. The Journal of Systems and Software, 143(??):ii, September 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301171.

Anonymous:2018:EBi

[Ano18i] Anonymous. Editorial Board. The Journal of Systems and Software, 144(??):ii, October 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301560.

Anonymous:2019:EBa

[Ano19a] Anonymous. Editorial Board. The Journal of Systems and Software, 148(??):ii, February 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302620.

Anonymous:2019:EBb

[Ano19b] Anonymous. Editorial Board. The Journal of Systems and Software, 149(??):ii, March 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121930010X.

Anonymous:2019:EBc

[Ano19c] Anonymous. Editorial Board. The Journal of Systems and Software, 150(??):ii, April 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300251.

Anonymous:2019:EBd

[Ano19d] Anonymous. Editorial Board. The Journal of Systems and Software, 151(??):ii, May 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300494.

Anonymous:2019:EBe

[Ano19e] Anonymous. Editorial Board. The Journal of Systems and Software, 152(??):ii, June 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300688.

Anonymous:2019:EBf

[Ano19f] Anonymous. Editorial Board. The Journal of Systems and Software, 153(??):ii, July 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121219300871.

Anonymous:2019:EBg

[Ano19g] Anonymous. Editorial Board. The Journal of Systems and Software, 154(??):ii, August 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301141.

Anonymous:2019:EBh

[Ano19h] Anonymous. Editorial Board. The Journal of Systems and Software, 155(??):ii, September 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301244.

Anonymous:2019:EBi

[Ano19i] Anonymous. Editorial Board. The Journal of Systems and Software, 156(??):ii, October 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121930161X.

Anonymous:2019:EBj

[Ano19j] Anonymous. Editorial Board. The Journal of Systems and Software, 157(??):??, November 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301852.

Anonymous:2019:EBk

[Ano19k] Anonymous. Editorial Board. The Journal of Systems and Software, 158(??):??, December 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219302122.

Anonymous:2019:IAN

[Ano19l] Anonymous. Introducing article numbering to The Journal of Systems Software. The Journal of Systems and Software, 157(??):??, November 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301888.

Aral:2016:NAE

[AO16] Atakan Aral and Tolga Ovatman. Networkaware embedding of virtual machine clusters onto federated cloud infrastructure. The Journal of Systems and Software, 120(??):89–104, October 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301078.

Avritzer:2010:CIS

[APCS10] Alberto Avritzer, Daniel Paulish, Yuanfang Cai, and Kanwarpreet Sethi. Coordination implications of software architecture in a global software development project. The Journal of Systems and Software, 83 (10):1881–1895, October 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Alferez:2014:DAS

[APM^{+}14] G. H. Alférez, V. Pelechano, R. Mazo, C. Salinesi, and D. Diaz. Dynamic adaptation of service compositions with variability models. The Journal of Systems and Software, 91(??):24–47, May 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001465.

Antonopoulos:2010:CMA

[APS^{+}10] Ch. Antonopoulos, A. Prayati, T. Stoyanova, C. Koulamas, and G. Papadopoulos. Corrigendum to “A modeling approach on the TelosB WSN platform power consumption” [J. Syst. Software 83 (2010) 1355–1363]. The Journal of Systems and Software, 83(12): 2665, December 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). See [PAS^{+}10].

Ali:2016:FAV

[APS16] Nauman Bin Ali, Kai Petersen, and Kurt Schneider. FLOW-assisted value stream mapping in the early phases of large-scale software development. The Journal of Systems and Software, 111(??):213–227, January 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002253.

Antoniou:2012:DRS

[APT^{+}12] Dimitris Antoniou, Yannis Plegas, Athanasios Tsakalidis, Giannis Tzimas, and Emmanouil Viennas. Dynamic refinement of search engines results utilizing the user intervention. The Journal of Systems and Software, 85(7):1577–1587, July 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212000350.

Ali:2014:SLR

[APW14] Nauman Bin Ali, Kai Petersen, and Claes Wohlin. A systematic literature review on the industrial use of software process simulation. The Journal of Systems and Software, 97(??):65–85, November 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214001502.

Al-Qershi:2011:HCD

[AQK11] Osamah M. Al-Qershi and Bee Ee Khoo. High capacity data hiding schemes for medical images based on difference expansion. The Journal of Systems and Software, 84(1):105–112, January 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Abebe:2012:AAO

[AR12] Ermyas Abebe and Caspar Ryan. Adaptive application offloading using distributed abstract class graphs in mobile environments. The Journal of Systems and Software, 85(12):2755–2769, December 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001653.

Areias:2017:SDP

[AR17] Miguel Areias and Ricardo Rocha. On scaling dynamic programming problems with a multithreaded tabling Prolog system. The Journal of Systems and Software, 125(??):417– 426, March 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

//www.sciencedirect.
com/science/article/
pii/S0164121216300929.

Alam:2018:QGA

[AR18] Taj Alam and Zahid Raza. Quantum genetic algorithm based scheduler for batch of precedence constrained jobs on heterogeneous computing systems. The Journal of Systems and Software, 135(??):126–142, January 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302194.

Anvari:2017:EII

[ARH^{+}17] Farshid Anvari, Deborah Richards, Michael Hitchens, Muhammad Ali Babar, Hien Minh Thi Tran, and Peter Busch. An empirical investigation of the influence of persona with personality traits on conceptual design. The Journal of Systems and Software, 134(??):324–339, December 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302078.

Alvarez:2016:MOO

[ÁRMC16] Josefa Díaz Álvarez, José L. Risco-Martín, and J. Manuel Colmenar. Multi-objective optimization of energy consumption and execution time in a single level cache memory for embedded systems. The Journal of Systems and Software, 111(??):200–212, January 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002241.

Ali:2010:APA

[ARS10] Nour Ali, Isidro Ramos, and Carlos Solís. Ambient-PRISMA: Ambients in mobile aspect-oriented software architecture. The Journal of Systems and Software, 83(6):937– 958, June 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Alvares:2017:DSL

[ARS17] Frederico Alvares, Eric Rutten, and Lionel Seinturier. A domainspecific language for the control of self-adaptive component-based architecture. The Journal of Systems and Software, 130(??):94–112, August 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217300201.

Ahonen:2010:SEP

[AS10] Jarmo J. Ahonen and Paula Savolainen. Software engineering projects may fail before they are started: Post-mortem analysis of five cancelled projects. The Journal of Systems and Software, 83 (11):2175–2187, November 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Axelsson:2016:QAS

[AS16] Jakob Axelsson and Mats Skoglund. Quality assurance in software ecosystems: a systematic literature mapping and research agenda. The Journal of Systems and Software, 114(??):69–81, April 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215002861.

Antinyan:2017:RMA

[AS17] Vard Antinyan and Mirosław Staron. Rendex: a method for automated reviews of textual requirements. The Journal of Systems and Software, 131(??):63–77, September 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301000.

Adalid:2014:USA

[ASdMGM14] Damián Adalid, Alberto Salmerón, María del Mar Gallardo, and Pedro Merino. Using SPIN for automated debugging of infinite executions of Java programs. The Journal of Systems and Software, 90(??):61– 75, April 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002641.

Alahyari:2017:SVA

[ASG17] Hiva Alahyari, Richard Berntsson Svensson, and Tony Gorschek. A study of value in agile software development organizations. The Journal of Systems and Software, 125(??): 271–288, March 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

//www.sciencedirect.
com/science/article/
pii/S0164121216302539.

Abdellatief:2013:MSI

[ASGJ13] Majdi Abdellatief, Abu Bakar Md Sultan, Abdul Azim Abdul Ghani, and Marzanah A. Jabar. A mapping study to investigate componentbased software system metrics. The Journal of Systems and Software, 86(3):587–603, March 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002798.

Altarawy:2018:LLA

[ASMM18] Doaa Altarawy, Hossameldin Shahin, Ayat Mohammed, and Na Meng. Lascad: Language-agnostic software categorization and similar application detection. The Journal of Systems and Software, 142(??):21–34, August 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300682.

Ahonen:2015:RPM

[ASMN15] Jarmo J. Ahonen, Paula Savolainen, Helena Merikoski, and Jaana Nevalainen. Reported project management effort, project size, and contract type. The Journal of Systems and Software, 109 (??):205–213, November 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001739.

Sun:2010:MMV

[aSRS^{+}10] Chang ai Sun, Rowan Rossing, Marco Sinnema, Pavel Bulanov, and Marco Aiello. Modeling and managing the variability of Web servicebased systems. The Journal of Systems and Software, 83(3):502–516, March 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sun:2018:FLW

[aSRZ^{+}18] Chang ai Sun, Yufeng Ran, Caiyun Zheng, Huai Liu, Dave Towey, and Xiangyu Zhang. Fault localisation for WS-BPEL programs based on predicate switching and program slicing. The Journal of Systems and Software, 135(??):191–204, January 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302534.

Ali:2016:MDP

[ASV^{+}16] Farhan Azmat Ali, Pieter Simoens, Tim Verbelen, Piet Demeester, and Bart Dhoedt. Mobile device power models for energy efficient dynamic offloading at runtime. The Journal of Systems and Software, 113(??):173–187, March 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215002666.

Asplund:2015:DTI

[AT15] Fredrik Asplund and Martin Törngren. The discourse on tool integration beyond technology, a literature survey. The Journal of Systems and Software, 106(??):117–131, August 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215000886.

Abuta:2018:RCR

[AT18] Eric Abuta and Jeff Tian. Reliability over consecutive releases of a semiconductor Optical Endpoint Detection software system developed in a small company. The Journal of Systems and Software, 137(??):355–365, March 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302996.

Abushark:2017:FAE

[ATHM17] Yoosef Abushark, John Thangarajah, James Harland, and Tim Miller. A framework for automatically ensuring the conformance of agent designs. The Journal of Systems and Software, 131 (??):266–310, September 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301048.

Aleti:2018:EMU

[ATvHJ18] Aldeida Aleti, Catia Trubiani, André van Hoorn, and Pooyan Jamshidi. An efficient method for uncertainty propagation in robust software performance estimation. The Journal of Systems and Software, 138(??):222– 235, April 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218300013.

Abdullah:2012:AAO

[AV12] Lili Marziana Abdullah and June M. Verner. Analysis and application of an outsourcing risk framework. The Journal of Systems and Software, 85(8):1930–1952, August 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212000647.

Aldave:2019:LCR

[AVGM19] Ainhoa Aldave, Juan M. Vara, David Granada, and Esperanza Marcos. Leveraging creativity in requirements elicitation within agile software development: a systematic literature review. The Journal of Systems and Software, 157(??):??, November 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301712.

Arrieta:2019:SBT

[AWSE19] Aitor Arrieta, Shuai Wang, Goiuria Sagardui, and Leire Etxeberria. Search-based test case prioritization for simulation-based testing of cyber-physical system product lines. The Journal of Systems and Software, 149(??):1–34, March 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302085.

Ali:2010:DJB

[AYZI10] Mohammed F. M. Ali, Mohammed I. Younis, Kamal Z. Zamli, and Widad Ismail. Development of Java based RFID application programmable interface for heterogeneous RFID system. The Journal of Systems and Software, 83 (11):2322–2331, November 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ahmed:2011:VSI

[AZ11] Bestoun S. Ahmed and Kamal Z. Zamli. A variable strength interaction test suites generation strategy using Particle Swarm Optimization. The Journal of Systems and Software, 84 (12):2171–2185, December 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211001464.

Abreu:2011:SDS

[AZvG11] Rui Abreu, Peter Zoeteweij, and Arjan J. C. van Gemund. Simultaneous debugging of software faults. The Journal of Systems and Software, 84 (4):573–586, April 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Agarwal:2014:SCS

[AZX14] Nitin Agarwal, Aoying Zhou, and Guandong Xu. Social cyber systems — challenges, opportunities, and beyond. The Journal of Systems and Software, 94(??):1– 3, August 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214001204.

Boukhris:2017:CSB

[BAAD17] Salah Boukhris, Anneliese Andrews, Ahmed Alhaddad, and Rinku Dewri. A case study of black box fail-safe testing in web applications. The Journal of Systems and Software, 131 (??):146–167, September 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121630190X.

Behnia:2013:IEB

[BAAS13] S. Behnia, A. Akhavan, A. Akhshani, and A. Samsudin. Image encryption based on the Jacobian elliptic maps. The Journal of Systems and Software, 86 (9):2429–2438, September 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001283.

Blas:2014:SOS

[BAI^{+}14] Javier Garcia Blas, Monica Abella, Florin Isaila, Jesus Carretero, and Manuel Desco. Surfing the optimization space of a multiple-GPU parallel implementation of a X-ray tomography reconstruction algorithm. The Journal of Systems and Software, 95 (??):166–175, September 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214001034.

Bezerra:2017:EQM

[BAM17] Carla I. M. Bezerra, Rossana M. C. Andrade, and Jose Maria Monteiro. Exploring quality measures for the evaluation of feature models: a case study. The Journal of Systems and Software, 131 (??):366–385, September 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216301340.

Barnawi:2015:AAE

[Bar15] Abdulaziz Y. Barnawi. Aggregation for adaptive and energy-efficient MAC in wireless sensor networks. The Journal of Systems and Software, 109(??):161–176, November 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215001636.

Bernardo:2010:HCP

[BBA10] Marco Bernardo, Edoardo Bontà, and Alessandro Aldini. Handling communications in process algebraic architectural description languages: Modeling, verification, and implementation. The Journal of Systems and Software, 83(8): 1404–1429, August 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Barenghi:2013:FIT

[BBBP13] Alessandro Barenghi, Guido M. Bertoni, Luca Breveglieri, and Gerardo Pelosi. A fault induction technique based on voltage underfeeding with application to attacks against AES and RSA. The Journal of Systems and Software, 86(7):1864–1878, July 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000320.

Bashari:2018:SAS

[BBD18] Mahdi Bashari, Ebrahim Bagheri, and Weichang Du. Self-adaptation of service compositions through product line reconfiguration. The Journal of Systems and Software, 144(??):84–105, October 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218301134.

Bartolini:2011:BWB

[BBEM11] Cesare Bartolini, Antonia Bertolino, Sebastian Elbaum, and Eda Marchetti. Bringing white-box testing to Service Oriented Architectures through a Service Oriented Approach. The Journal of Systems and Software, 84(4):655–668, April 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Berg:2018:SSE

[BBND^{+}18] Vebjørn Berg, Jørgen Birkeland, Anh Nguyen-Duc, Ilias O. Pappas, and Letizia Jaccheri. Software startup engineering: a systematic mapping study. The Journal of Systems and Software, 144(??):255–274, October 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218301286.

Bosch:2010:ICI

[BBS10] Jan Bosch and Petra Bosch-Sijtsema. From integration to composition: On the impact of software product lines, global development and ecosystems. The Journal of Systems and Software, 83(1):67–76, January 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Balsini:2019:EEL

[BCA^{+}19] Alessio Balsini, Tommaso Cucinotta, Luca Abeni, Joel Fernandes, Phil Burk, Patrick Bellasi, and Morten Rasmussen. Energy-efficient low-latency audio on Android. The Journal of Systems and Software, 152(??):182–195, June 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300585.

Bachwani:2014:RSU

[BCBZ14] Rekha Bachwani, Olivier Crameri, Ricardo Bianchini, and Willy Zwaenepoel. Recommending software upgrades with Mojave. The Journal of Systems and Software, 96(??):10– 23, October 2014. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214001198.

Bernardi:2018:RBD

[BCD^{+}18] Mario Luca Bernardi, Gerardo Canfora, Giuseppe A. Di Lucca, Massimiliano Di Penta, and Damiano Distante. The relation between developers’ communication and fix-inducing changes: An empirical study. The Journal of Systems and Software, 140(??):111– 125, June 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218300426.

Bahsoon:2010:SIS

[BCEF10] Rami Bahsoon, Licia Capra, Wolfgang Emmerich, and Mohamed E. Fayad. Special issue on software architecture and mobility. The Journal of Systems and Software, 83 (6):883–884, June 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bernaschina:2018:FSO

[BCF18] Carlo Bernaschina, Sara Comai, and Piero Fraternali. Formal semantics of OMG’s Interaction Flow Modeling Language (IFML) for mobile and rich-client application model driven development. The Journal of Systems and Software, 137(??):239–260, March 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302959.

Barricelli:2019:EUD

[BCFP19] Barbara Rita Barricelli, Fabio Cassano, Daniela Fogli, and Antonio Piccinno. End-user development, end-user programming and enduser software engineering: a systematic mapping study. The Journal of Systems and Software, 149(??):101–137, March 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302577.

Brambilla:2013:IJS

[BCG^{+}13] Marco Brambilla, Cinzia Cappiello, Irene Garrigós, Jose-Norberto Mazón, and Santiago Meliá. Introduction to the JSS special issue of Web 2.0 engineering: New practices and emerging challenges. The Journal of Systems and Software, 86 (12):2967–2969, December 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121300215X.

Bosu:2014:PIO

[BCG^{+}14] Amiangshu Bosu, Jeffrey Carver, Rosanna Guadagno, Blake Bassett, Debra McCallum, and Lorin Hochstein. Peer impressions in open source organizations: a survey. The Journal of Systems and Software, 94(??):4–15, August 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214000818.

Breivold:2012:SAE

[BCL12] Hongyu Pei Breivold, Ivica Crnkovic, and Magnus Larsson. Software architecture evolution through evolvability analysis. The Journal of Systems and Software, 85(11):2574–2592, November 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121200163X.

Bertolino:2018:CSS

[BCL^{+}18] Antonia Bertolino, Antonello Calabrò, Francesca Lonetti, Eda Marchetti, and Breno Miranda. A categorization scheme for software engineering conference papers and its application. The Journal of Systems and Software, 137(??):114–129, March 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302844.

Bo:2011:TBM

[BCLW11] Wang Bo, Huang Chuanhe, Li Layuan, and Yang Wenzhong. Trust-based minimum cost opportunistic routing for ad hoc networks. The Journal of Systems and Software, 84(12):2107–2122, December 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001373.

Basile:2019:MMS

[BCR^{+}19] Cataldo Basile, Daniele Canavese, Leonardo Regano, Paolo Falcarin, and Bjorn De Sutter. A metamodel for software protections and reverse engineering attacks. The Journal of Systems and Software, 150(??):3–21, April 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302838.

Brogi:2018:FAM

[BCS18] Antonio Brogi, Andrea Canciani, and Jacopo Soldani. Fault-aware management protocols for multi-component applications. The Journal of Systems and Software, 139(??):189–210, May 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121830030X.

Bishop:2010:CSC

[BD10] Judith Bishop and Theo Danzfuss. Computer supported cooperative work and “Software architectures and mobility: A Roadmap”. The Journal of Systems and Software, 83(6):902–905, June 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bauer:2016:CCD

[BD16] Christine Bauer and Anind K. Dey. Considering context in the design of intelligent systems: Current practices and suggestions for improvement. The Journal of Systems and Software, 112(??):26–47, February 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002320.

Bettini:2017:XTJ

[BD17] Lorenzo Bettini and Ferruccio Damiani. Xtraitj: Traits for the Java platform. The Journal of Systems and Software, 131(??):419–441, September 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121216301297.

Boissel-Dallier:2015:MIS

[BDBLP15] Nicolas Boissel-Dallier, Fréderick Benaben, Jean-Pierre Lorré, and Hervé Pingaud. Mediation information system engineering based on hybrid service composition mechanism. The Journal of Systems and Software, 108(??):39–59, October 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001168.

Bavota:2015:EII

[BDD^{+}15] Gabriele Bavota, Andrea De Lucia, Massimiliano Di Penta, Rocco Oliveto, and Fabio Palomba. An experimental investigation on the innate relationship between quality and refactoring. The Journal of Systems and Software, 107(??):1–14, September 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001053.

Bertolino:2011:MMR

[BDDS11] Antonia Bertolino, Guglielmo De Angelis, Alessio Di Sandro, and Antonino Sabetta. Is my model right? Let me ask the expert. The Journal of Systems and Software, 84(7): 1089–1099, July 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bravo:2013:GSS

[BDG13] Crescencio Bravo, Rafael Duque, and Jesús Gallardo. A groupware system to support collaborative programming: Design and experiences. The Journal of Systems and Software, 86(7): 1759–1771, July 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121212002439.

Blundo:2013:CIA

[BDGP13] Carlo Blundo, Angelo De Caro, Clemente Galdi, and Giuseppe Persiano. Certified Information Access. The Journal of Systems and Software, 86 (9):2439–2450, September 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001313.

Buisson:2016:SRC

[BDLM16] Jérémy Buisson, Fabien Dagnat, Elena Leroux, and Sébastien Martinez. Safe reconfiguration of Coqcots and Pycots components. The Journal of Systems and Software, 122(??):430–444, December 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002630.

Bozhinoski:2019:SMR

[BDM^{+}19] Darko Bozhinoski, Davide Di Ruscio, Ivano Malavolta, Patrizio Pelliccione, and Ivica Crnkovic. Safety for mobile robotic systems: a systematic mapping study from a software engineering perspective. The Journal of Systems and Software, 151(??):150–179, May 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300317.

Bastos:2017:SPL

[BdMSNO^{+}17] Jonatas Ferreira Bastos, Paulo Anselmo da Mota Silveira Neto, Pádraig O’Leary, Eduardo Santana de Almeida, and Silvio Romero de Lemos Meira. Software product lines adoption in small organizations. The Journal of Systems and Software, 131(??):112–128, September 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121217300997.

Bavota:2011:IEC

[BDO11] Gabriele Bavota, Andrea De Lucia, and Rocco Oliveto. Identifying Extract Class refactoring opportunities using structural and semantic cohesion measures. The Journal of Systems and Software, 84(3):397–414, March 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bernardez:2018:ERE

[BDPRC18] Beatriz Bernárdez, Amador Durán, José A. Parejo, and Antonio Ruiz-Cortés. An experimental replication on the effect of the practice of mindfulness in conceptual modeling performance. The Journal of Systems and Software, 136(??):153–172, February 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301005.

Blincoe:2017:GEI

[BDV17] Kelly Blincoe, Daniela Damian, and Giuseppe Valetto. Guest editors introduction: Context for software developers. The Journal of Systems and Software, 128(??):198– 199, June 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217300468.

Butting:2019:SCI

[BEK^{+}19] Arvid Butting, Robert Eikermann, Oliver Kautz, Bernhard Rumpe, and Andreas Wortmann. Systematic composition of independent language features. The Journal of Systems and Software, 152(??):50–69, June 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300366.

Bai:2014:SIS

[BEZ14] Xiaoying Bai, Atilla Elci, and Mohammad Zulkernine. Special issue on “Trustworthy Software Systems for the Digital Society”. The Journal of Systems and Software, 89(??):1–2, March 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214000144.

Bertolino:2013:SSA

[BFLZ13] Antonia Bertolino, Howard Foster, J. Jenny Li, and Hong Zhu. Special section on automation of software test. The Journal of Systems and Software, 86(8):1977, August 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000344.

Barddal:2017:SFD

[BGEP17] Jean Paul Barddal, Heitor Murilo Gomes, Fabrício Enembreck, and Bernhard Pfahringer. A survey on feature drift adaptation: Definition, benchmark, challenges and future directions. The Journal of Systems and Software, 127(??):278– 294, May 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301030.

Biel:2010:EBC

[BGG10] Bettina Biel, Thomas Grill, and Volker Gruhn. Exploring the benefits of the combination of a software architecture analysis and a usability evaluation of a mobile application. The Journal of Systems and Software, 83 (11):2031–2044, November 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Borrego:2013:MTP

[BGLG13] Diana Borrego, María Teresa Gómez-López, and Rafael M. Gasca. Minimizing testpoint allocation to improve diagnosability in business process models. The Journal of Systems and Software, 86 (11):2725–2741, November 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001398.

Bagheri:2016:SAP

[BGS^{+}16] Hamid Bagheri, Joshua Garcia, Alireza Sadeghi, Sam Malek, and Nenad Medvidovic. Software architectural principles in contemporary mobile software: from conception to practice. The Journal of Systems and Software, 119 (??):31–44, September 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300607.

Benelallam:2018:DRM

[BGTC18] Amine Benelallam, Abel Gómez, Massimo Tisi, and Jordi Cabot. Distributing relational model transformation on MapReduce. The Journal of Systems and Software, 142(??):1–20, August 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300669.

Binkley:2010:AIG

[BHH^{+}10] David Binkley, Mark Harman, Youssef Hassoun, Syed Islam, and Zheng Li. Assessing the impact of global variables on program dependence and dependence clusters. The Journal of Systems and Software, 83(1):96– 107, January 2010. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Benghazi:2012:ECD

[BHH^{+}12] Kawtar Benghazi, María V.

Hurtado, Miguel J. Hornos, María L. Rodríguez, Carlos Rodríguez-Domínguez, Ana B. Pelegrina, and María J. Rodríguez-Fórtiz. Enabling correct design and formal analysis of Ambient Assisted Living systems. The Journal of Systems and Software, 85(3):498–510, March 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001221.

Balsamo:2012:MCP

[BHM12] Simonetta Balsamo, Peter G. Harrison, and Andrea Marin. Methodological construction of product-form stochastic Petri nets for performance evaluation. The Journal of Systems and Software, 85(7):1520– 1539, July 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002986.

Brito:2018:URM

[BHVR18] Gleison Brito, Andre Hora, Marco Tulio Valente, and Romain Robbes. On the use of replacement messages in API deprecation: an empirical study. The Journal of Systems and Software, 137(??):306–321, March 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121730300X.

Bishop:2013:IRD

[Bis13] Judith Bishop. Industry’s role in data and software curation in the cloud. The Journal of Systems and Software, 86 (9):2327–2329, September 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000204.

Baird:2011:SAW

[BJG11] R. Baird, N. Jorgenson, and R. Gamble. Self-adapting workflow reconfiguration. The Journal of Systems and Software, 84(3):510–524, March 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bilogrevic:2011:MTC

[BJK^{+}11] Igor Bilogrevic, Murtuza Jadliwala, Praveen Kumar, Sudeep Singh Walia, Jean-Pierre Hubaux, Imad Aad, and Valtteri Niemi. Meetings through the cloud: Privacypreserving scheduling on mobile devices. The Journal of Systems and Software, 84(11):1910– 1927, November 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121211001026.

Bucur:2011:SVS

[BK11] Doina Bucur and Marta Kwiatkowska. On software verification for sensor nodes. The Journal of Systems and Software, 84(10):1693–1707, October 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001051.

Brunnert:2017:CPE

[BK17] Andreas Brunnert and Helmut Krcmar. Continuous performance evaluation and capacity planning using resource profiles for enterprise applications. The Journal of Systems and Software, 123(??):239–262, January 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001831.

Bennouar:2010:NAC

[BKH10] D. Bennouar, T. Khammaci, and A. Henni. A new approach for component’s port modeling in software architecture. The Journal of Systems and Software, 83(8): 1430–1442, August 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Baek:2018:EGP

[BKLE18] Sun Geol Baek, Dong Hyun Kang, Sungkil Lee, and Young Ik Eom. Efficient graph pattern matching framework for network-based in-vehicle fault detection. The Journal of Systems and Software, 140(??):17–31, June 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300347.

Butting:2019:CAF

[BKRW19] Arvid Butting, Oliver Kautz, Bernhard Rumpe, and Andreas Wortmann. Continuously analyzing finite, message-driven, time-synchronous component and connector systems during architecture evolution. The Journal of Systems and Software, 149(??):437–461, March 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302760.

Bakar:2015:FEA

[BKS15] Noor Hasrina Bakar, Zarinah M. Kasirun, and Norsaremah Salleh. Feature extraction approaches from natural language requirements for reuse in software product lines: a systematic literature review. The Journal of Systems and Software, 106(??):132–149, August 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001004.

Bian:2013:SSP

[BKSM13] Yixin Bian, Gunes Koru, Xiaohong Su, and Peijun Ma. SPAPE: a semantic-preserving amorphous procedure extraction method for nearmiss clones. The Journal of Systems and Software, 86(8):2077–2093, August 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000733.

See corrigendum [BKSM14].

Bian:2014:CSS

[BKSM14] Yixin Bian, Gunes Koru, Xiaohong Su, and Peijun Ma. Corrigendum to: “SPAPE: a semantic-preserving amorphous procedure extraction method for nearmiss clones”: [J. Syst. Softw. 86 (2013) 2077– 2093]. The Journal of Systems and Software, 88(??):250, February 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002732.

See [BKSM13].

Bottaci:2010:TPT

[BKW10] Leonardo Bottaci, Gregory M. Kapfhammer, and Neil Walkinshaw. TAIC-PART 2009 — Testing: Academic & Industrial Conference — Practice And Research Techniques: Special Section Editorial. The Journal of Systems and Software, 83(12):2367–2368, December 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Byun:2011:SMC

[BL11] Jin Wook Byun and Dong Hoon Lee. On a security model of conjunctive keyword search over encrypted relational database. The Journal of Systems and Software, 84(8):1364–1372, August 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Baek:2019:ISA

[BL19] Hyeongboo Baek and Jinkyu Lee. Improved schedulability analysis of the contention-free policy for real-time systems. The Journal of Systems and Software, 154(??):112–124, August 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301025.

Bousse:2018:ODE

[BLC^{+}18] Erwan Bousse, Dorian Leroy, Benoit Combemale, Manuel Wimmer, and Benoit Baudry. Omniscient debugging for executable DSLs. The Journal of Systems and Software, 137(??):261– 288, March 2018. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121217302765.

Bao:2018:EAD

[BLL^{+}18] Liang Bao, Qian Li, Peiyao Lu, Jie Lu, Tongxiao Ruan, and Ke Zhang. Execution anomaly detection in large-scale systems through console log analysis. The Journal of Systems and Software, 143 (??):172–186, September 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301031.

Beydoun:2011:HDW

[BLLGSMB11] Ghassan Beydoun, Antonio A. Lopez-Lorca, Francisco García-Sánchez, and Rodrigo Martínez-Béjar. How do we measure and improve the quality of a hierarchical ontology? The Journal of Systems and Software, 84(12):2363–2373, December 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211001853.

Bertini:2010:POD

[BLM10] Luciano Bertini, Julius C. B. Leite, and Daniel Mossé. Power optimization for dynamic configuration in heterogeneous Web server clusters. The Journal of Systems and Software, 83(4):585–598, April 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Baek:2018:MLC

[BLS18] Hyeongboo Baek, Jinkyu Lee, and Insik Shin. Multi-level contentionfree policy for real-time multiprocessor scheduling. The Journal of Systems and Software, 137(??):36–49, March 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217302777.

Bi:2018:SMS

[BLTY18] Tingting Bi, Peng Liang, Antony Tang, and Chen Yang. A systematic mapping study on text analysis techniques in software architecture. The Journal of Systems and Software, 144(??):533–558, October 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301493.

Binkley:2015:EII

[BLUH15] Dave Binkley, Dawn Lawrie, Christopher Uehlinger, and Daniel Heinz. Enabling improved IRbased feature location. The Journal of Systems and Software, 101(??): 30–42, March 2015. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121214002428.

Boes:2017:SOM

[BM17] Jérémy Boes and Frédéric Migeon. Self-organizing multi-agent systems for the control of complex systems. The Journal of Systems and Software, 134(??):12–28, December 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301838.

Bajunaid:2018:EMO

[BM18] Noor Bajunaid and Daniel A. Menascé. Efficient modeling and optimizing of checkpointing in concurrent component-based software systems. The Journal of Systems and Software, 139(??):1–13, May 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218300116.

Bahsoon:2013:FSE

[BMA^{+}13] Rami Bahsoon, Ivan Mistrík, Nour Ali, T. S. Mohan, and Nenad Medvidović. The future of software engineering IN and FOR the cloud. The Journal of Systems and Software, 86 (9):2221–2224, September 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213001349.

Bani-Mohammad:2011:PEN

[BMAH11] Saad Bani-Mohammad, Ismail Ababneh, and Mazen Hamdan. Performance evaluation of noncontiguous allocation algorithms for 2D mesh interconnection networks.

The Journal of Systems and Software, 84 (12):2156–2170, December 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001452.

Besker:2018:MAT

[BMB18] Terese Besker, Antonio Martini, and Jan Bosch. Managing architectural technical debt: a unified model and systematic literature review. The Journal of Systems and Software, 135(??):1–16, January 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302121.

Besker:2019:SDP

[BMB19] Terese Besker, Antonio Martini, and Jan Bosch. Software developer productivity loss due to technical debt — a replication and extension study examining developers’ development work. The Journal of Systems and Software, 156(??):41–61, October 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301335.

Besson:2015:BTD

[BMKM15] Felipe Besson, Paulo Moura, Fabio Kon, and Dejan Milojicic. Bringing Test-Driven Development to web service choreographies. The Journal of Systems and Software, 99 (??):135–154, January 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121400209X.

Bai:2013:HPI

[BML^{+}13] Yuebin Bai, Yao Ma, Cheng Luo, Duo Lv, and Yuanfeng Peng. A high performance interdomain communication approach for virtual machines. The Journal of Systems and Software, 86(2):367–376, February 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212002476.

Beggas:2014:TIS

[BMLL14] Mounir Beggas, Lionel Médini, Frederique Laforest, and Mohamed Tayeb Laskri. Towards an ideal service QoS in fuzzy logic-based adaptation planning middleware. The Journal of Systems and Software, 92(??):71– 81, June 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001738.

Barreto:2011:OTS

[BMS11] Paulo S. L. M. Barreto, Rafael Misoczki, and Marcos A. Simplicio, Jr. One-time signature scheme from syndrome decoding over generic error-correcting codes. The Journal of Systems and Software, 84(2):198– 204, February 2011. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Boix:2014:DDM

[BND14] Elisa Gonzalez Boix, Carlos Noguera, and Wolfgang De Meuter. Distributed debugging for mobile networks. The Journal of Systems and Software, 90(??):76–90, April 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213002811.

Bhattacharya:2012:AHA

[BNS12] Pamela Bhattacharya, Iulian Neamtiu, and Christian R. Shelton. Automated, highly-accurate, bug assignment using machine learning and tossing graphs. The Journal of Systems and Software, 85(10):2275– 2292, October 2012. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212001240.

Borba:2012:SIS

[Bor12] Paulo Borba. Special issue with selected papers from the 23rd Brazilian Symposium on Software Engineering. The Journal of Systems and Software, 85(2):215, February 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211002767.

Bosch:2012:SET

[Bos12] Jan Bosch. Software ecosystems: Taking software development beyond the boundaries of the organization. The Journal of Systems and Software, 85(7):1453– 1454, July 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000829.

Baca:2013:CGS

[BP13] Dejan Baca and Kai Petersen. Countermeasure graphs for software security risk assessment: an action research. The Journal of Systems and Software, 86 (9):2411–2428, September 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001027.

Botta:2015:IPI

[BP15] Alessio Botta and Antonio Pescapé. IP packet interleaving for UDP bursty losses. The Journal of Systems and Software, 109(??):177–191, November 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001673.

Buchdid:2019:PIS

[BPB19] Samuel B. Buchdid, Roberto Pereira, and M. Cecília C. Baranauskas. Pro-IDTV: a sociotechnical process model for designing IDTV applications. The Journal of Systems and Software, 154(??):234–254, August 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301037.

Belk:2013:MUW

[BPGS13] Marios Belk, Efi Papatheocharous, Panagiotis Germanakos, and George Samaras. Modeling users on the World Wide Web based on cognitive factors, navigation behavior and clustering techniques. The Journal of Systems and Software, 86 (12):2995–3012, December 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213001040.

Basso:2016:ADM

[BPO^{+}16] Fábio Paulo Basso, Raquel Mainardi Pillat, Toacy Cavalcante Oliveira, Fabricia Roos-Frantz, and Rafael Z.

Frantz. Automated design of multi-layered web information systems. The Journal of Systems and Software, 117(??):612– 637, July 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300358.

Bartzas:2010:SMS

[BPQP^{+}10] Alexandros Bartzas, Miguel Peon-Quiros, Christophe Poucet, Christos Baloukas, Stylianos Mamagkakis, Francky Catthoor, Dimitrios Soudris, and Jose M. Mendias. Software metadata: Systematic characterization of the memory behaviour of dynamic applications. The Journal of Systems and Software, 83(6):1051–1075, June 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Brauer:2018:MOO

[BPSK18] Johannes Bräuer, Reinhold Plösch, Matthias Saft, and Christian Körner. Measuring object-oriented design principles: The results of focus groupbased research. The Journal of Systems and Software, 140(??):74–90, June 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121830044X.

Bae:2014:CMB

[BRB14] Gigon Bae, Gregg Rothermel, and Doo-Hwan Bae. Comparing model-based and dynamic eventextraction based GUI testing techniques: an empirical study. The Journal of Systems and Software, 97(??):15–46, November 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214001472.

Bhattacharjee:2012:WSN

[BRG^{+}12] Sudipta Bhattacharjee, Pramit Roy, Soumalya Ghosh, Sudip Misra, and Mohammad S. Obaidat. Wireless sensor network-based fire detection, alarming, monitoring and prevention system for Bord-and-Pillar coal mines. The Journal of Systems and Software, 85(3):571–581, March 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002378.

Braendeland:2010:MAM

[BRS10] Gyrd Brændeland, Atle Refsdal, and Ketil Stølen. Modular analysis and modelling of risk scenarios with dependencies. The Journal of Systems and Software, 83 (10):1995–2013, October 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bollin:2018:AMM

[BRS^{+}18] Andreas Bollin, Elisa Reçi, Csaba Szabó, Veronika Szabóová, and Rudolf Siebenhofer. Applying a maturity model during a software engineering course — how planning and task-solving processes influence the course performance. The Journal of Systems and Software, 144(??):397– 408, October 2018. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121218301407.

Bolloju:2012:BSU

[BS12] Narasimha Bolloju and Sherry X. Y. Sun. Benefits of supplementing use case narratives with activity diagrams — an exploratory study. The Journal of Systems and Software, 85 (9):2182–2191, September 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121200129X.

Bruun:2015:NAU

[BS15] Anders Bruun and Jan Stage. New approaches to usability evaluation in software development: Barefoot and crowdsourcing. The Journal of Systems and Software, 105(??):40–53, July 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000655.

Burgstaller:2012:SAF

[BSB12] Bernd Burgstaller, Bernhard Scholz, and Johann Blieberger. A symbolic analysis framework for static analysis of imperative programming languages. The Journal of Systems and Software, 85(6):1418–1439, June 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002974.

Boix:2014:PMC

[BSDD14] Elisa Gonzalez Boix, Christophe Scholliers, Wolfgang De Meuter, and Theo D’Hondt. Programming mobile context-aware applications with TOTAM. The Journal of Systems and Software, 92(??):3–19, June 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001799.

Ballesteros:2012:OUB

[BSG12] Francisco J. Ballesteros, Enrique Soriano, and Gorka Guardiola. Octopus: an Upperware based system for building personal pervasive environments. The Journal of Systems and Software, 85(7):1637–1649, July 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121200043X.

Burger:2018:FSA

[BSG^{+}18] Jens Bürger, Daniel Strüber, Stefan Gärtner, Thomas Ruhroth, Jan Jürjens, and Kurt Schneider. A framework for semi-automated coevolution of security knowledge and system models. The Journal of Systems and Software, 139(??):142–160, May 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121830027X.

Bagheri:2018:CAM

[BSK^{+}18] Maryam Bagheri, Marjan Sirjani, Ehsan Khamespanah, Narges Khakpour, Ilge Akkaya, Ali Movaghar, and Edward A. Lee. Coordinated actor model of self-adaptive track-based traffic control systems. The Journal of Systems and Software, 143 (??):116–139, September 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301092.

Basaran:2010:RFC

[BSKL10] Can Basaran, Mehmet H. Suzer, Kyoung-Don Kang, and Xue Liu. Robust fuzzy CPU utilization control for dynamic workloads. The Journal of Systems and Software, 83(7):1192–1204, July 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Borba:2017:SMP

[BT17] Eric Borba and Eduardo Tavares. Stochastic modeling for performance and availability evaluation of hybrid storage systems. The Journal of Systems and Software, 134(??):1–11, December 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301851.

Blasco:2015:HDT

[BTPLST15] Jorge Blasco, Juan E.

Tapiador, Pedro Peris-Lopez, and Guillermo Suarez-Tangil. Hindering data theft with encrypted data trees. The Journal of Systems and Software, 101(??):147– 158, March 2015. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121214002775.

Baranwal:2015:FMA

[BV15] Gaurav Baranwal and Deo Prakash Vidyarthi. A fair multi-attribute combinatorial double auction model for resource allocation in cloud computing. The Journal of Systems and Software, 108(??):60–76, October 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215001272.

Bauer:2016:CRP

[BV16] Veronika Bauer and Antonio Vetro’. Comparing reuse practices in two large software-producing companies. The Journal of Systems and Software, 117(??):545–582, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216300176.

Borges:2018:WGS

[BV18] Hudson Borges and Marco Tulio Valente. What’s in a GitHub star? Understanding repository starring practices in a social coding platform. The Journal of Systems and Software, 146(??):112–129, December 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301961.

Boone:2010:SQA

[BVV^{+}10] Bas Boone, Sofie Van Hoecke, Gregory Van Seghbroeck, Niels Joncheere, Viviane Jonckers, Filip De Turck, Chris Develder, and Bart Dhoedt. SALSA: QoS-aware load balancing for autonomous service brokering. The Journal of Systems and Software, 83(3):446–456, March 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Boucke:2010:CAM

[BWH10] Nelis Boucké, Danny Weyns, and Tom Holvoet. Composition of architectural models: Empirical analysis and language support. The Journal of Systems and Software, 83 (11):2108–2127, November 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Badampudi:2016:SCD

[BWP16] Deepika Badampudi, Claes Wohlin, and Kai Petersen. Software component decision-making: In-house, OSS, COTS or outsourcing — a systematic literature review. The Journal of Systems and Software, 121(??):105–124, November 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301212.

Badampudi:2018:DMP

[BWW^{+}18] Deepika Badampudi, Krzysztof Wnuk, Claes Wohlin, Ulrik Franke, Darja Smite, and Antonio Cicchetti. A decision-making processline for selection of software asset origins and components. The Journal of Systems and Software, 135(??):88–104, January 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302182.

Bayley:2010:FSV

[BZ10] Ian Bayley and Hong Zhu. Formal specification of the variants and behavioural features of design patterns. The Journal of Systems and Software, 83(2):209–221, February 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bezemer:2014:POD

[BZ14] Cor-Paul Bezemer and Andy Zaidman. Performance optimization of deployed software-as-a-service applications. The Journal of Systems and Software, 87(??):87–103, January 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121300232X.

Chen:2014:SIE

[CA14] Zhenyu Chen and Hira Agrawal. Special issue on Emerging Topics on Software Debugging. The Journal of Systems and Software, 90(??):1– 2, April 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121400034X.

Ciavotta:2017:MIL

[CAG17] Michele Ciavotta, Danilo Ardagna, and Giovanni Paolo Gibilisco. A mixed integer linear programming optimization approach for multi-cloud capacity allocation. The Journal of Systems and Software, 123(??):64–78, January 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301996.

Carvalho:2015:SCI

[CAHV15] Nuno Ramos Carvalho, José João Almeida, Pedro Rangel Henriques, and Maria João Varanda. From source code identifiers to natural language terms. The Journal of Systems and Software, 100(??):117–128, February 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002179.

Caglayan:2016:EDC

[ÇB16] Bora Çaglayan and Ayse Basar Bener. Effect of developer collaboration activity on software quality in two large scale projects. The Journal of Systems and Software, 118(??):288–296, August 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0164121216300097.

Cerpa:2016:EDF

[CBAV16] Narciso Cerpa, Matthew Bardeen, César A. Astudillo, and June Verner. Evaluating different families of prediction methods for estimating software project outcomes. The Journal of Systems and Software, 112 (??):48–64, February 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121500223X.

Chihani:2014:PCA

[CBC14] Bachir Chihani, Emmanuel Bertin, and Noël Crespi. Programmable context awareness framework. The Journal of Systems and Software, 92(??):59–70, June 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001908.

Comerio:2015:SPM

[CBC^{+}15] Marco Comerio, Carlo Batini, Marco Castelli, Simone Grega, Marco Rossetti, and Gianluigi Viscusi. Service portfolio management: a repository-based framework. The Journal of Systems and Software, 104(??):112–125, June 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000333.

Chwa:2015:CUP

[CBL^{+}15] Hoon Sung Chwa, Hyoungbu Back, Jinkyu Lee, Kieu-My Phan, and Insik Shin. Capturing urgency and parallelism using quasi-deadlines for real-time multiprocessor scheduling. The Journal of Systems and Software, 101(??):15–29, March 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002544.

Chanak:2016:MSB

[CBS16] Prasenjit Chanak, Indrajit Banerjee, and R. Simon Sherratt. Mobile sink based fault diagnosis scheme for wireless sensor networks. The Journal of Systems and Software, 119 (??):45–57, September 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300620.

Cornu:2016:CAT

[CBSM16] Benoit Cornu, Earl T.

Barr, Lionel Seinturier, and Martin Monperrus. Casper: Automatic tracking of null dereferences to inception with causality traces. The Journal of Systems and Software, 122(??):52–62, December 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301522.

Capilla:2014:ODS

[CBT^{+}14] Rafael Capilla, Jan Bosch, Pablo Trinidad, Antonio Ruiz-Cortés, and Mike Hinchey. An overview of Dynamic Software Product Line architectures and techniques: Observations from research and industry. The Journal of Systems and Software, 91(??):3–23, May 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214000119.

Conoscenti:2019:CDA

[CBVF19] Marco Conoscenti, Veronika Besner, Antonio Vetrò, and Daniel Méndez Fernández. Combining data analytics and developers feedback for identifying reasons of inaccurate estimations in agile software development. The Journal of Systems and Software, 156(??):126–135, October 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301372.

Cheng:2016:VMN

[CBZ^{+}16] Kun Cheng, Yuebin Bai, Yongwang Zhao, Yao Ma, Duo Lu, Yuanfeng Peng, and Minxuan Zhou. HV^{2}M: a novel approach to boost inter-VM network performance for Xenbased HVMs. The Journal of Systems and Software, 114(??):54–68, April 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215002782.

Chen:2011:SEE

[CC11] Chung-Yang Chen and P. Pete Chong. Software engineering education: a study on conducting collaborative senior project development. The Journal of Systems and Software, 84 (3):479–491, March 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Carver:2017:SIS

[CCCY17] Doris Carver, W. K.

Chan, Carl K. Chang, and Hongji Yang. Special issue on Software Engineering Technology and Applications. The Journal of Systems and Software, 126(??):85–86, April 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216302552.

Chen:2019:MSO

[CCD19] Hui Chen, John Coogle, and Kostadin Damevski. Modeling Stack Overflow tags and topics as a hierarchy of concepts. The Journal of Systems and Software, 156(??):283–299, October 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301499.

Camara:2016:IAB

[CCdL^{+}16] Javier Cámara, Pedro Correia, Rogério de Lemos, David Garlan, Pedro Gomes, Bradley Schmerl, and Rafael Ventura. Incorporating architecture-based selfadaptation into an adaptive industrial software system. The Journal of Systems and Software, 122(??):507–523, December 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002113.

Cardoso:2016:UTF

[CCdR^{+}16] K. V. Cardoso, S. L. Correa, J. F. de Rezende, B. S. da Silva, M. O. M. C. de Mello, and M. A. Cruz. Using traffic filtering rules and Open-Flow devices for transparent flow switching and automatic dynamiccircuit creation in hybrid networks. The Journal of Systems and Software, 117(??):113–128, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000522.

Chella:2010:AOS

[CCG^{+}10] Antonio Chella, Massimo Cossentino, Salvatore Gaglio, Luca Sabatucci, and Valeria Seidita. Agentoriented software patterns for rapid and affordable robot programming. The Journal of Systems and Software, 83 (4):557–573, April 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Calinescu:2018:ESR

[CCG^{+}18] Radu Calinescu, Milan Ceska, Simos Gerasimou, Marta Kwiatkowska, and Nicola Paoletti. Efficient synthesis of robust models for stochastic systems. The Journal of Systems and Software, 143 (??):140–158, September 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300967.

Cabot:2010:VVD

[CCGdL10] Jordi Cabot, Robert Clarisó, Esther Guerra, and Juan de Lara. Verification and validation of declarative model-to-model transformations through invariants. The Journal of Systems and Software, 83(2):283–302, February 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Clariso:2016:BRM

[CCGdL16] Robert Clarisó, Jordi Cabot, Esther Guerra, and Juan de Lara. Backwards reasoning for model transformations: Method and applications. The Journal of Systems and Software, 116(??):113–132, June 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001788.

Cagliero:2014:TDA

[CCGG14] Luca Cagliero, Tania Cerquitelli, Paolo Garza, and Luigi Grimaudo. Twitter data analysis by means of Strong Flipping Generalized Itemsets. The Journal of Systems and Software, 94(??):16–29, August 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214000806.

Chang:2014:SNF

[CCH14] Hsien-Tsung Chang, Yi-Min Chang, and Sheng-Yuan Hsiao. Scalable network file systems with load balancing and fault tolerance for web services. The Journal of Systems and Software, 93(??):102–109, July 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214000685.

Chen:2019:AOA

[CCL^{+}19] Xing Chen, Jiaqing Chen, Bichun Liu, Yun Ma, Ying Zhang, and Hao Zhong. AndroidOff: Offloading Android application based on cost estimation. The Journal of Systems and Software, 158(??):??, December 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121930192X.

Chang:2011:SFW

[CCLL11] Chin-Chen Chang, Kuo-Nan Chen, Chin-Feng Lee, and Li-Jen Liu. A secure fragile watermarking scheme based on chaos-and-Hamming code. The Journal of Systems and Software, 84 (9):1462–1470, September 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211000549.

Capiluppi:2012:GEI

[CCM12] Andrea Capiluppi, Anthony Cleve, and Naouel Moha. Guest editors’ introduction to the special issue on automated software evolution. The Journal of Systems and Software, 85(10):2193–2194, October 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001598.

Czibula:2019:ACM

[CCMOM19] Istvan Gergely Czibula, Gabriela Czibula, Diana-Lucia Miholca, and Zsuzsanna Onet-Marian. An aggregated coupling measure for the analysis of object-oriented software systems. The Journal of Systems and Software, 148(??):1–20, February 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302371.

Carrozza:2010:MLA

[CCN^{+}10] G. Carrozza, D. Cotroneo, R. Natella, A. Pecchia, and S. Russo. Memory leak analysis of mission-critical middleware. The Journal of Systems and Software, 83 (9):1556–1567, September 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Campanelli:2018:ITC

[CCP18] Amadeu Silveira Campanelli, Ronaldo Darwich Camilo, and Fernando Silva Parreiras. The impact of tailoring criteria on agile practices adoption: a survey with novice agile practitioners in Brazil. The Journal of Systems and Software, 137(??):366–379, March 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217303059.

Cabot:2014:VUO

[CCR14] J. Cabot, R. Clarisó, and D. Riera. On the verification of UML/OCL class diagrams using constraint programming. The Journal of Systems and Software, 93(??):1– 23, July 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214000739.

Chen:2013:CCA

[CCWT13] Chiu-Mei Chen, Chiao-Min Chen, Hsien-Chu Wu, and Chwei-Shyong Tsai. Common carotid artery condition recognition technology using waveform features extracted from ultrasound spectrum images. The Journal of Systems and Software, 86(1):38–46, January 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212001926.

Chan:2011:WSC

[CCY11] W. K. Chan, Yuen Yau Chiu, and Yuen Tak Yu. A web search-centric approach to recommender systems with URLs as minimal user contexts.

The Journal of Systems and Software, 84(6):930– 941, June 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Choi:2010:SSA

[CD10] Gyu Sang Choi and Chita R. Das. A superscalar software architecture model for Multi-Core Processors (MCPs). The Journal of Systems and Software, 83 (10):1823–1837, October 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cano:2011:SEE

[CDA11] Maria-Dolores Cano and Gines Domenech-Asensi. A secure energy-efficient m-banking application for mobile devices. The Journal of Systems and Software, 84(11):1899– 1909, November 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S016412121100149X.

Caballer:2014:CPE

[CdAM^{+}14] Miguel Caballer, Carlos de Alfonso, Germán Moltó, Eloy Romero, Ignacio Blanquer, and Andrés García. Code-Cloud: a platform to enable execution of programming models on the Clouds. The Journal of Systems and Software, 93(??):187–198, July 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214000521.

Carvalho:2018:ASS

[CdCAdO18] Rainara Maia Carvalho, Rossana Maria de Castro Andrade, and Káthia Marçal de Oliveira. AQUArIUM

— a suite of software measures for HCI quality evaluation of ubiquitous mobile applications. The Journal of Systems and Software, 136(??):101–136, February 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302728.

Cavalcanti:2016:TSA

[CdCMdMSNdA16] Yguaratã Cerqueira Cavalcanti, Ivan do Carmo Machado, Paulo Anselmo da Motal S. Neto, and Eduardo Santana de Almeida. Towards semi-automated assignment of software change requests. The Journal of Systems and Software, 115(??):82– 101, May 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000352.

Canon:2010:DCH

[CDGJ10]  Louis-Claude  Canon,

Olivier Dubuisson, Jens Gustedt, and Emmanuel Jeannot. Defining and controlling the heterogeneity of a cluster: The Wrekavoc tool. The Journal of Systems and Software, 83(5):786–802, May 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cuadrado:2018:OMM

[CdL18] Jesús Sánchez Cuadrado and Juan de Lara. Open meta-modelling frameworks via meta-object protocols. The Journal of Systems and Software, 145(??):1–24, November 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301432.

Chaudhari:2015:THR

[CDOP15] Mahesh B. Chaudhari, Suzanne W. Dietrich, Jennifer Ortiz, and Spencer Pearson. Towards a hybrid relational and XML benchmark for loosely-coupled distributed data sources. The Journal of Systems and Software, 109 (??):78–87, November 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001557.

Costache:2017:RMC

[CDPM17] Stefania Costache, Djawida Dib, Nikos Parlavantzas, and Christine Morin. Resource management in cloud platform as a service systems: Analysis and opportunities. The Journal of Systems and Software, 132(??):98–118, October 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217300845.

CavalcantedeMenezes:2014:DPB

[CdR^{+}14] Luciana Cavalcante de Menezes, Cláudio de Souza Baptista, Ana Gabrielle Ramos Falcão, Maxwell Guimarães de Oliveira, and Leandro Balby Marinho. DYSCS: a platform to build geographically and semantically enhanced social content sites. The Journal of Systems and Software, 94(??):39–49, August 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121400079X.

Chen:2013:PQP

[CDRT13] Jaime Chen, Manuel Díaz, Bartolomé Rubio, and José M. Troya. PS-QUASAR: a publish/subscribe QoS aware middleware for Wireless Sensor and Actor Networks. The Journal of Systems and Software, 86(6):1650–1662, June 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000502.

Castiglione:2010:SPI

[CDS10] Aniello Castiglione, Alfredo De Santis, and Claudio Soriente. Security and privacy issues in the Portable Document Format. The Journal of Systems and Software, 83 (10):1813–1822, October 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Capiluppi:2018:GEI

[CdS18] Andrea Capiluppi and Fabio Queda Bueno da Silva. Guest Editors’ introduction to the special issue on replication studies in software engineering. The Journal of Systems and Software, 136(??):137–138, February 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217302145.

Corno:2019:CNP

[CDS19] Fulvio Corno, Luigi De Russis, and Juan Pablo Sáenz. On the challenges novice programmers experience in developing IoT systems: a survey. The Journal of Systems and Software, 157(??):??, November 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301566.

Carvalho:2018:IDS

[CdSdSG^{+}18] Michelle Larissa Luciano Carvalho, Matheus Lessa Gonçalves da Silva, Gecynalda Soares da Silva Gomes, Alcemir Rodrigues Santos, Ivan do Carmo Machado, Magno Luã de Jesus Souza, and Eduardo Santana de Almeida. On the implementation of dynamic software product lines: an exploratory study. The Journal of Systems and Software, 136(??):74– 100, February 2018. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121217302601.

Cucinotta:2012:HTC

[CF12] Tommaso Cucinotta and Dario Faggioli. Handling timing constraints violations in soft realtime applications as exceptions. The Journal of Systems and Software, 85(4):995–1011, April 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002949.

Castro:2013:LIA

[CF13] Laura M. Castro and Miguel A. Francisco. A language-independent approach to black-box testing using Erlang as test specification language. The Journal of Systems and Software, 86 (12):3109–3122, December 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213001714.

Curumsing:2019:UIE

[CFA^{+}19] Maheswaree Kissoon Curumsing, Niroshinie Fernando, Mohamed Abdelrazek, Rajesh Vasa, Kon Mouzakis, and John Grundy. Understanding the impact of emotions on software: a case study in requirements gathering and evaluation. The Journal of Systems and Software, 147(??):215–229, January 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218301341.

Cetina:2017:IFL

[CFAP17] Carlos Cetina, Jaime Font, Lorena Arcega, and Francisca Pérez. Improving feature location in long-living modelbased product families designed with sustainability goals. The Journal of Systems and Software, 134(??):261–278, December 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121730211X.

Caivano:2018:SEU

[CFL^{+}18] Danilo Caivano, Daniela Fogli, Rosa Lanzilotti, Antonio Piccinno, and Fabio Cassano. Supporting end users to control their smart home: design implications from a literature review and an empirical investigation. The Journal of Systems and Software, 144(??):295–313, October 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301262.

Capilla:2019:SVD

[CFL19] Rafael Capilla, Lidia Fuentes, and Malte Lochau. Software variability in dynamic environments. The Journal of Systems and Software, 156(??):62–64, October 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301323.

Cunha:2016:ERS

[CFM^{+}16] Jácome Cunha, João Paulo Fernandes, Pedro Martins, Jorge Mendes, Rui Pereira, and João Saraiva. Evaluating refactorings for spreadsheet models. The Journal of Systems and Software, 118(??):234–250, August 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216300280.

Capra:2011:FIO

[CFMRL11] Eugenio Capra, Chiara Francalanci, Francesco Merlo, and Cristina Rossi-Lamastra. Firms’ involvement in Open Source projects: a tradeoff between software structural quality and popularity. The Journal of Systems and Software, 84(1):144–161, January 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cicirelli:2010:SBA

[CFN10] Franco Cicirelli, Angelo Furfaro, and Libero Nigro. A servicebased architecture for dynamically reconfigurable workflows. The Journal of Systems and Software, 83(7):1148–1164, July 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Caivano:2018:ABV

[CFRPC^{+}18] Danilo Caivano, María Fernández-Ropero, Ricardo Pérez-Castillo, Mario Piattini, and Michele Scalera. Artifact-based vs. human-perceived understandability and modifiability of refactored business processes: an experiment. The Journal of Systems and Software, 144(??):143–164, October 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301225.

Cheng:2012:SLA

[CG12] Shang-Wen Cheng and David Garlan. Stitch: a language for architecturebased self-adaptation. The Journal of Systems and Software, 85 (12):2860–2875, December 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121212000714.

Caporuscio:2015:EFI

[CG15] Mauro Caporuscio and Carlo Ghezzi. Engineering future Internet applications: the Prime approach. The Journal of Systems and Software, 106(??):9–27, August 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215000783.

Cugola:2014:SDA

[CGPT14] Gianpaolo Cugola, Carlo Ghezzi, Leandro Sales Pinto, and Giordano Tamburrelli. SelfMotion: a declarative approach for adaptive serviceoriented mobile applications. The Journal of Systems and Software, 92(??):32–44, June 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213002653.

Camara:2019:STS

[CGS19] Javier Cámara, David Garlan, and Bradley Schmerl. Synthesizing tradeoff spaces with quantitative guarantees for families of software systems. The Journal of Systems and Software, 152(??):33–49, June 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300445.

Chen:2010:NUP

[CH10a] Tzung-Her Chen and Jyun-Ci Huang. A novel user-participating authentication scheme. The Journal of Systems and Software, 83(5):861– 867, May 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chou:2010:EXM

[CH10b] Shih-Chien Chou and Chun-Hao Huang. An extended XACML model to ensure secure information access for Web services. The Journal of Systems and Software, 83 (1):77–84, January 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Christensen:2010:EIA

[CH10c] Henrik Bærbak Christensen and Klaus Marius Hansen. An empirical investigation of architectural prototyping. The Journal of Systems and Software, 83(1):133–142, January 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Clementsen:2010:VPF

[CH10d] Davur S. Clementsen and Zhen He. Vertical partitioning for flash and HDD database systems. The Journal of Systems and Software, 83 (11):2237–2250, November 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:2011:DEQ

[CH11] Tao-Ku Chang and Gwan-Hwan Hwang. Developing an efficient query system for encrypted XML documents. The Journal of Systems and Software, 84(8):1292–1305, August 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:2017:CSC

[Cha17] Victor Chang. A cybernetics Social Cloud. The Journal of Systems and Software, 124 (??):195–211, February 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002939.

Clemente:2011:MCC

[CHCO11] Pedro J. Clemente, Juan Hernández, José M. Conejero, and Guadalupe Ortiz. Managing crosscutting concerns in component based systems using a model driven development approach. The Journal of Systems and Software, 84(6):1032– 1053, June 2011. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:2013:ISM

[Che13] Wei-Kuei Chen. Image sharing method for gray-level images. The Journal of Systems and Software, 86(2):581–585, February 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002774.

Chen:2017:CDO

[Che17] Lianping Chen. Continuous delivery: Overcoming adoption challenges. The Journal of Systems and Software, 128(??):72–86, June 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217300353.

Chen:2011:NEK

[CHL11] Shih-Sheng Chen, Tony Cheng-Kui Huang, and Zhe-Min Lin. New and efficient knowledge discovery of partial periodic patterns with multiple minimum supports. The Journal of Systems and Software, 84 (10):1638–1651, October 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211000975.

Chung:2013:GOS

[CHL^{+}13] Lawrence Chung, Tom Hill, Owolabi Legunsen, Zhenzhou Sun, Adip Dsouza, and Sam Supakkul. A goaloriented simulation approach for obtaining good private cloud-based system architectures. The Journal of Systems and Software, 86 (9):2242–2262, September 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0164121212002865.

Chen:2019:ADS

[CHL^{+}19] Huanchao Chen, Yuan Huang, Zhiyong Liu, Xiangping Chen, Fan Zhou, and Xiaonan Luo. Automatically detecting the scopes of source code comments. The Journal of Systems and Software, 153(??):45–63, July 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121930055X.

Chen:2017:SAA

[CHLW17] Luxi Chen, Linpeng Huang, Chen Li, and Xiwen Wu. Self-adaptive architecture evolution with model checking: a software cybernetics approach. The Journal of Systems and Software, 124(??):228–246, February 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000820.

Carbonnel:2019:MEC

[CHN19a] Jessie Carbonnel, Marianne Huchard, and Clémentine Nebut. Modelling equivalence classes of feature models with concept lattices to assist their extraction from product descriptions. The Journal of Systems and Software, 152(??):1– 23, June 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300378.

Carbonnel:2019:TCP

[CHN19b] Jessie Carbonnel, Marianne Huchard, and Clémentine Nebut. Towards complex product line variability modelling: Mining relationships from non-boolean descriptions. The Journal of Systems and Software, 156(??):341–360, October 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301311.

Cho:2013:CRN

[Cho13] Hyung-Ju Cho. Continuous range k-nearest neighbor queries in vehicular ad hoc networks. The Journal of Systems and Software, 86(5):1323–1332, May 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212003433.

Christin:2016:PMP

[Chr16] Delphine Christin. Privacy in mobile participatory sensing: Current trends and future challenges. The Journal of Systems and Software, 116(??):57–68, June 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000692.

Chaves:2019:VFD

[CIB^{+}19] Lennon C. Chaves, Hussama I. Ismail, Iury V. Bessa, Lucas C. Cordeiro, and Eddie B. de Lima Filho. Verifying fragility in digital systems with uncertainties using DSVerifier v 2.0. The Journal of Systems and Software, 153(??):22–43, July 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300603.

Ciccozzi:2016:ECP

[Cic16] Federico Ciccozzi. Explicit connection patterns (ECP) profile and semantics for modelling and generating explicit connections in complex UML composite structures. The Journal of Systems and Software, 121(??):329–344, November 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216000571.

Chen:2013:IGC

[CJ13] Wei-Ming Chen and Sheng-Hao Jhang. Improving Graph Cuts algorithm to transform sequence of stereo image to depth map. The Journal of Systems and Software, 86(1):198–210, January 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212002191.

Chou:2011:SAT

[CJL11] Chien-Ping Chou, Kuen-Fang Jea, and Heng-Hsun Liao. A syntactic approach to twigquery matching on XML streams. The Journal of Systems and Software, 84 (6):993–1007, June 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Capilla:2016:YSA

[CJT^{+}16] Rafael Capilla, Anton Jansen, Antony Tang, Paris Avgeriou, and Muhammad Ali Babar. 10 years of software architecture knowledge management: Practice and future. The Journal of Systems and Software, 116(??):191–205, June 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002034.

Chu:2015:ATA

[CKC15] Hsin-Hao Chu, Yu-Chon Kao, and Ya-Shu Chen. Adaptive thermal-aware task scheduling for multicore systems. The Journal of Systems and Software, 99(??):155–174, January 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002106.

Chae:2015:EED

[CKCK15] Dong-Kyu Chae, Sang-Wook Kim, Seong-Je Cho, and Yesol Kim. Effective and efficient detection of software theft via dynamic API authority vectors. The Journal of Systems and Software, 110(??):1–9, December 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121500179X.

Colombo:2012:BGB

[CKL12] Pietro Colombo, Ferhat Khendek, and Luigi Lavazza. Bridging the gap between requirements and design: an approach based on Problem Frames and SysML. The Journal of Systems and Software, 85(3):717–745, March 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002561.

Chen:2010:ART

[CKMT10] Tsong Yueh Chen, Fei-Ching Kuo, Robert G. Merkel, and T. H. Tse. Adaptive random testing: The ART of test case diversity. The Journal of Systems and Software, 83 (1):60–66, January 2010.

CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Constantinou:2015:AAN

[CKS15] Eleni Constantinou, George Kakarontzas, and Ioannis Stamelos. An automated approach for noise identification to assist software architecture recovery techniques. The Journal of Systems and Software, 107 (??):142–157, September 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121500117X.

Choi:2011:SIB

[CL11] Byoungju Choi and Richard Lai. Special issue on the best papers of QSIC 2009. The Journal of Systems and Software, 84(4):527, April 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chuang:2013:SPS

[CL13] Ming-Chin Chuang and Jeng-Farn Lee. SF-PMIPv6: a secure fast handover mechanism for Proxy Mobile IPv6 networks. The Journal of Systems and Software, 86(2):437–448, February 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212002555.

Chong:2015:AMR

[CL15] Chun Yong Chong and Sai Peck Lee. Analyzing maintainability and reliability of object-oriented software using weighted complex network. The Journal of Systems and Software, 110(??):28–53, December 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215001752.

Cao:2017:DON

[CL17a] Weiquan Cao and Yunzhao Li. DOTS: an online and near-optimal trajectory simplification algorithm. The Journal of Systems and Software, 126(??):34–44, April 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217300031.

Chong:2017:ACC

[CL17b] Chun Yong Chong and Sai Peck Lee. Automatic clustering constraints derivation from object-oriented software using weighted complex network with graph theory analysis. The Journal of Systems and Software, 133(??):28–53, November 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301772.

Carver:2018:STG

[CL18] R. Carver and Yu Lei.

Stateless techniques for generating global and local test oracles for message-passing concurrent programs. The Journal of Systems and Software, 136(??):237–265, February 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302753.

Conforti:2013:RTR

[CLF^{+}13] Raffaele Conforti, Marcello La Rosa, Giancarlo Fortino, Arthur H. M. ter Hofstede, Jan Recker, and Michael Adams. Real-time risk monitoring in business processes: a sensor-based approach. The Journal of Systems and Software, 86(11):2939–2965, November 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121300174X.

Chen:2013:QAV

[CLH^{+}13] Tzung-Her Chen, Yao-Sheng Lee, Wei-Lun Huang, Justie Su-Tzu Juan, Ying-Yu Chen, and Ming-Jheng Li. Quality-adaptive visual secret sharing by random grids. The Journal of Systems and Software, 86(5):1267–1274, May 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212003342.

Chen:2010:PRT

[CLL10] Jun Chen, Victor C. S. Lee, and Kai Liu. On the performance of realtime multi-item request scheduling in data broadcast environments. The Journal of Systems and Software, 83(8):1337– 1345, August 2010. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chuang:2014:AIS

[CLL14] Sun-Wen Chuang, Tainyi Luor, and Hsi-Peng Lu. Assessment of institutions, scholars, and contributions on agile software development (2001– 2012). The Journal of Systems and Software, 93(??):84–101, July 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214000697.

Chan:2018:SIS

[CLR18] W. K. Chan, Xiaodong Liu, and Hridesh Rajan. Special issue on software engineering technology and applications. The Journal of Systems and Software, 137(??):34–35, March 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302789.

Castro:2012:CAT

[CLS^{+}12] Jaelson Castro, Marcia Lucena, Carla Silva, Fernanda Alencar, Emanuel Santos, and João Pimentel. Changing attitudes towards the generation of architectural models. The Journal of Systems and Software, 85(3):463–479, March 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001415.

Chang:2014:SSN

[CLY14] Hsung-Pin Chang, Yen-Ting Liu, and Shang-Sheng Yang. Surviving sensor node failures by MMU-less incremental checkpointing. The Journal of Systems and Software, 87(??):74–86, January 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121300229X.

Chang:2017:EEH

[CLY17] Che-Wei Chang, Chun-Yi Liu, and Chuan-Yue Yang. Energy-efficient heterogeneous resource management for wireless monitoring systems. The Journal of Systems and Software, 131 (??):168–180, September 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0164121216301984.

Cugola:2012:CEP

[CM12] Gianpaolo Cugola and Alessandro Margara. Complex event processing with T-REX. The Journal of Systems and Software, 85(8):1709–1728, August 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000842.

Ceke:2015:EEE

[CM15] Denis Ceke and Boris Milasinović. Early effort estimation in web application development. The Journal of Systems and Software, 103(??):219– 237, May 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000394.

Chatzigiannakis:2011:IMP

[CMK^{+}11] Ioannis Chatzigiannakis, Georgios Mylonas, Panagiotis Kokkinos, Orestis Akribopoulos, Marios Logaras, and Irene Mavrommati. Implementing multiplayer pervasive installations based on mobile sensing devices: Field experience and user evaluation from a public showcase. The Journal of Systems and Software, 84 (11):1989–2004, November 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001658.

Chatzipoulidis:2015:IIR

[CMM15] Aristeidis Chatzipoulidis, Dimitrios Michalopoulos, and Ioannis Mavridis. Information infrastructure risk prediction through platform vulnerability analysis. The Journal of Systems and Software, 106(??):28–41, August 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000837.

Ciancarini:2019:CTA

[CMR19] Paolo Ciancarini, Marcello Missiroli, and Daniel Russo. Cooperative thinking: Analyzing a new framework for software engineering education. The Journal of Systems and Software, 157(??):??, November 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121930175X.

Chan:2016:SQS

[CNG16] Nguyen Ngoc Chan, Nattawat Nonsung, and Walid Gaaloul. Service querying to support process variant development. The Journal of Systems and Software, 122(??):538–552, December 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001697.

Cirilo:2012:APD

[CNKL12] Elder Cirilo, Ingrid Nunes, Uirá Kulesza, and Carlos Lucena. Automating the product derivation process of multiagent systems product lines. The Journal of Systems and Software, 85(2):258–276, February 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001075.

Chang:2013:NVB

[CNL13] Chin-Chen Chang, Thai Son Nguyen, and Chia-Chen Lin. A novel VQ-based reversible data hiding scheme by using hybrid encoding strategies. The Journal of Systems and Software, 86(2):389–402, February 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002518.

Canizares:2018:MMT

[CNM18] Pablo C. Cañizares, Alberto Núñez, and Mercedes G. Merayo. Mutomvo: Mutation testing framework for simulated cloud and HPC environments. The Journal of Systems and Software, 143(??):187–207, September 2018. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121218300931.

Curcio:2018:RES

[CNMR18] Karina Curcio, Tiago Navarro, Andreia Malucelli, and Sheila Reinehr. Requirements engineering: A systematic mapping study in agile software development. The Journal of Systems and Software, 139(??):32–50, May 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300141.

Chung:2012:NAD

[CNSG12] Lawrence Chung, Manuel Noguera, Nary Subramanian, and José Luis Garrido. Novel approaches in the design and implementation of system/software architectures. The Journal of Systems and Software, 85(3):459–462, March 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002962.

Clarke:2012:ISB

[CO12] Paul Clarke and Rory V.

O’Connor. The influence of SPI on business success in software SMEs: an empirical study. The Journal of Systems and Software, 85(10):2356–2367, October 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001355.

Campanelli:2015:AMT

[CP15] Amadeu Silveira Campanelli and Fernando Silva Parreiras. Agile methods tailoring — a systematic literature review. The Journal of Systems and Software, 110(??):85–100, December 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215001843.

Catolino:2018:ECP

[CPD^{+}18] Gemma Catolino, Fabio Palomba, Andrea De Lucia, Filomena Ferrucci, and Andy Zaidman. Enhancing change prediction models using developer-related factors. The Journal of Systems and Software, 143(??):14–28, September 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218300918.

Costa:2016:ERA

[CPDM16] Bruno Costa, Paulo F. Pires, Flávia C. Delicato, and Paulo Merson. Evaluating REST architectures — approach, tooling and guidelines.

The Journal of Systems and Software, 112 (??):156–180, February 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002150.

Chaumont:2013:SCI

[CPL13] M. Chaumont, W. Puech, and C. Lahanier. Securing color information of an image by concealing the color palette. The Journal of Systems and Software, 86(3):809–825, March 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121200324X.

Cotroneo:2013:TTS

[CPR13] Domenico Cotroneo, Roberto Pietrantuono, and Stefano Russo. Testing techniques selection based on ODC fault types and software metrics. The Journal of Systems and Software, 86(6):1613–1637, June 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121213000319.

Cotroneo:2016:HDB

[CPRT16] Domenico Cotroneo, Roberto Pietrantuono, Stefano Russo, and Kishor Trivedi. How do bugs surface? A comprehensive study on the characteristics of software bugs manifestation. The Journal of Systems and Software, 113(??):27–43, March 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215002460.

Chatziantoniou:2011:SRT

[CPS11] Damianos Chatziantoniou, Katerina Pramatari, and Yannis Sotiropoulos. Supporting real-time supply chain decisions based on RFID data streams. The Journal of Systems and Software, 84(4):700–710, April 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Couto:2014:PSD

[CPV^{+}14] Cesar Couto, Pedro Pires, Marco Tulio Valente, Roberto S. Bigonha, and Nicolas Anquetil. Predicting software defects with causality tests. The Journal of Systems and Software, 93(??):24– 41, July 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214000351.

Chen:2016:MMR

[CPX16] Tsong Yueh Chen, Pak-Lok Poon, and Xiaoyuan Xie. METRIC: METamorphic Relation Identification based on the Category-choice framework. The Journal of Systems and Software, 116(??):177–190, June 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001624.

Chen:2014:UHG

[CPYZ14] Bihuan Chen, Xin Peng, Yijun Yu, and Wenyun Zhao. Uncertainty handling in goal-driven selfoptimization — limiting the negative effect on adaptation. The Journal of Systems and Software, 90(??):114–127, April 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214000065.

Catolino:2019:ABS

[CPZF19] Gemma Catolino, Fabio Palomba, Andy Zaidman, and Filomena Ferrucci. Not all bugs are the same: Understanding, characterizing, and classifying bug types. The Journal of Systems and Software, 152(??):165–181, June 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219300536.

Child:2019:CEV

[CRC19] Mike Child, Peter Rosner, and Steve Counsell. A comparison and evaluation of variants in the coupling between objects metric. The Journal of Systems and Software, 151(??):120–132, May 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219300305.

Conejero:2018:EET

[CREH^{+}18] José M. Conejero, Roberto Rodríguez-Echeverría, Juan Hernández, Pedro J. Clemente, Carmen Ortiz-Caraballo, Elena Jurado, and Fernando Sánchez-Figueroa. Early evaluation of technical debt impact on maintainability. The Journal of Systems and Software, 142(??):92–114, August 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300736.

Conejero:2013:REL

[CRESF^{+}13] José M. Conejero, Roberto Rodríguez-Echeverría, Fernando Sánchez-Figueroa, Marino Linaje, Juan C. Preciado, and Pedro J. Clemente. Re-engineering legacy Web applications into RIAs by aligning modernization requirements, patterns and RIA features. The Journal of Systems and Software, 86 (12):2981–2994, December 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001076.

Colmant:2018:NCP

[CRK^{+}18] Maxime Colmant, Romain Rouvoy, Mascha Kurpicz, Anita Sobe, Pascal Felber, and Lionel Seinturier. The next 700 CPU power models. The Journal of Systems and Software, 144(??):382–396, October 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218301377.

Christin:2011:SPM

[CRKH11] Delphine Christin, Andreas Reinhardt, Salil S. Kanhere, and Matthias Hollick. A survey on privacy in mobile participatory sensing applications. The Journal of Systems and Software, 84 (11):1928–1946, November 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211001701.

Chen:2012:MCT

[CRL^{+}12] Zhongqiang Chen, Mema Roussopoulos, Zhanyan Liang, Yuan Zhang, Zhongrong Chen, and Alex Delis. Malware characteristics and threats on the Internet ecosystem. The Journal of Systems and Software, 85(7):1650–1672, July 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000441.

Coman:2014:CCP

[CRSS14] Irina D. Coman, Pierre N.

Robillard, Alberto Sillitti, and Giancarlo Succi. Cooperation, collaboration and pairprogramming: Field studies on backup behavior. The Journal of Systems and Software, 91(??):124–134, May 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214000107.

Coronel:2012:HPD

[CS12] J. O. Coronel and J. E.

Simó. High performance dynamic voltage/ frequency scaling algorithm for real-time dynamic load management. The Journal of Systems and Software, 85(4):906– 919, April 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121211002871.

Cai:2015:CSP

[CS15] Haipeng Cai and Raul Santelices. A comprehensive study of the predictive accuracy of dynamic change-impact analysis. The Journal of Systems and Software, 103(??):248–265, May 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215000424.

Cai:2016:MLP

[CS16] Haipeng Cai and Raul Santelices. Method-level program dependence abstraction and its application to impact analysis. The Journal of Systems and Software, 122(??):311–326, December 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216301960.

Cerdeiral:2019:SPM

[CS19] Cristina T. Cerdeiral and Gleison Santos. Software project management in high maturity: a systematic literature mapping. The Journal of Systems and Software, 148(??):56–87, February 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302218.

Cacho:2014:BDP

[CSF^{+}14] Nelio Cacho, Claudio Sant’anna, Eduardo Figueiredo, Francisco Dantas, Alessandro Garcia, and Thais Batista. Blending design patterns with aspects: a quantitative study. The Journal of Systems and Software, 98(??):117–139, December 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001885.

Costa:2015:PRF

[CSM15] Pedro Costa, João Gabriel Silva, and Henrique Madeira. Practical and representative faultloads for large-scale software systems. The Journal of Systems and Software, 103(??):182–197, May 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215000357.

Capota:2019:TMC

[CSMC19] Eugenia Ana Capota, Cristina Sorina Stangaciu, Mihai Victor Micea, and Daniel-Ioan Curiac. Towards mixed criticality task scheduling in cyber physical systems: Challenges and perspectives. The Journal of Systems and Software, 156(??):204–216, October 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301426.

Chen:2017:TBS

[CSN^{+}17] Tse-Hsun Chen, Weiyi Shang, Meiyappan Nagappan, Ahmed E. Hassan, and Stephen W. Thomas. Topic-based software defect explanation. The Journal of Systems and Software, 129(??):79–106, July 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216300528.

Cieslicki:2010:MCP

[CSS10] Damian Cieslicki, Stefan Schaeckeler, and Thomas Schwarz. Maintaining and checking parity in highly available Scalable Distributed Data Structures. The Journal of Systems and Software, 83 (4):529–542, April 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:2013:RWM

[CSS^{+}13] Xianyi Chen, Xingming Sun, Huiyu Sun, Zhili Zhou, and Jianjun Zhang. Reversible watermarking method based on asymmetrichistogram shifting of prediction errors. The Journal of Systems and Software, 86(10):2620–2626, October 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121300126X.

Chen:2010:IRP

[CSW10] Xiao Chen, Jian Shen, and Jie Wu. Improving routing protocol performance in delay tolerant networks using extended information. The Journal of Systems and Software, 83(8):1301–1309, August 2010. CODEN JSSODM.

ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:2013:IDE

[CSW13] Jeanne Chen, Chih-Wei Shiu, and Mei-Chen Wu. An improvement of diamond encoding using characteristic value positioning and modulus function. The Journal of Systems and Software, 86(5):1377–1383, May 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213000034.

Chen:2011:ARI

[CT11a] Chien-Chang Chen and Yao-Hong Tsai. Adaptive reversible image watermarking scheme. The Journal of Systems and Software, 84(3):428–434, March 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:2011:TVS

[CT11b] Tzung-Her Chen and Kai-Hsiang Tsao. Threshold visual secret sharing by random grids. The Journal of Systems and Software, 84(7):1197– 1208, July 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). See corrigendum [YWEL^{+}13].

Celik:2013:ITF

[ÇT13] Turgay Çelik and Bedir Tekinerdogan. S-IDE: a tool framework for optimizing deployment architecture of High Level Architecture based simulation systems. The Journal of Systems and Software, 86(10):2520–2541, October 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000599.

Chen:2012:CLE

[CTHW12] Jenhui Chen, Woei-Hwa Tarn, Wu-Hsiao Hsu, and Chih-Chieh Wang. Cross-layer end-to-end label switching protocol for WiMAX-MPLS heterogeneous networks. The Journal of Systems and Software, 85 (11):2459–2469, November 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001483.

Cagiltay:2013:PAN

[CTKT13] Nergiz Ercil Cagiltay, Gul Tokdemir, Ozkan Kilic, and Damla Topalli. Performing and analyzing non-formal inspections of entity relationship diagram (ERD). The Journal of Systems and Software, 86 (8):2184–2195, August 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001015.

Chen:2010:MLB

[CTL10] Ya-Shu Chen, Hsin-Liang Tsai, and Shi-Wu Lo. Multi-layer bus minimization for SoC. The Journal of Systems and Software, 83(1):121–132, January 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:2012:GBP

[CTL12] Ting-Yi Chang, Cheng-Jung Tsai, and Jyun-Hao Lin. A graphical-based password keystroke dynamic authentication system for touch screen handheld mobile devices. The Journal of Systems and Software, 85(5):1157–1165, May 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211003256.

Chandakanna:2014:MVC

[CV14] Veerabhadra Rao Chandakanna and Valli Kumari Vatsavayi. A model view controller based Self-Adjusting Clustering Framework. The Journal of Systems and Software, 89(??):193–206, March 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002859.

Chandakanna:2016:QAS

[CV16a] Veerabhadra Rao Chandakanna and Valli Kumari Vatsavayi. A QoSaware self-correcting observation based load balancer. The Journal of Systems and Software, 115(??):111–129, May 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000376.

Colanzi:2016:FDC

[CV16b] Thelma Elita Colanzi and Silvia Regina Vergilio. A feature-driven crossover operator for multi-objective and evolutionary optimization of product line architectures. The Journal of Systems and Software, 121(??):126–143, November 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000583.

Colanzi:2013:SBS

[CVGP13] Thelma Elita Colanzi, Silvia Regina Vergilio, Wesley Klewerton Guez Assunção, and Aurora Pozo. Search based software engineering: Review and analysis of the field in Brazil. The Journal of Systems and Software, 86(4):970–984, April 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002166.

Chen:2012:PER

[CW12] Ching-Wen Chen and Chuan-Chi Weng. A power efficiency routing and maintenance protocol in wireless multi-hop networks. The Journal of Systems and Software, 85(1):62–76, January 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001877.

Chen:2014:SBB

[CW14] Chien-Chang Chen and Wei-Jie Wu. A secure Boolean-based multisecret image sharing scheme. The Journal of Systems and Software, 92(??):107–114, June 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214000181.

Chang:2013:CEC

[CWJK13] Jamie Y. T. Chang, Eric T. G. Wang, James J. Jiang, and Gary Klein. Controlling ERP consultants: Client and provider practices. The Journal of Systems and Software, 86(5):1453– 1461, May 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000149.

Chen:2010:SSB

[CWK10] Ching-Wen Chen, Chuan-Chi Weng, and Yu-Chen Kuo. Signal strength based routing for power saving in mobile ad hoc networks. The Journal of Systems and Software, 83(8):1373–1386, August 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chae:2011:AAR

[CWK^{+}11] Heung Seok Chae, Gyun Woo, Tae Yeon Kim, Jung Ho Bae, and Won-Young Kim. An automated approach to reducing test suites for testing retargeted C compilers for embedded systems. The Journal of Systems and Software, 84 (12):2053–2064, December 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211000987.

Chen:2013:ITD

[CWK^{+}13] CaiSen Chen, Tao Wang, YingZhan Kou, Xiao-Cen Chen, and Xiong Li. Improvement of tracedriven I-Cache timing attack on the RSA algorithm. The Journal of Systems and Software, 86(1):100–107, January 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002130.

Caballe:2010:CPS

[CX10] Santi Caballé and Fatos Xhafa. CLPL: Providing software infrastructure for the systematic and effective construction of complex collaborative learning systems. The Journal of Systems and Software, 83 (11):2083–2097, November 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:2015:USE

[CXO^{+}15] Jie Chen, Xiwei Xu, Leon J. Osterweil, Liming Zhu, Yuriy Brun, Len Bass, Junchao Xiao, Mingshu Li, and Qing Wang. Using simulation to evaluate error detection strategies: a case study of cloudbased deployment processes. The Journal of Systems and Software, 110(??):205–221, December 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001879.

Chiang:2016:KMD

[CYT16] Mei-Ling Chiang, Chieh-Jui Yang, and Shu-Wei Tu. Kernel mechanisms with dynamic taskaware scheduling to reduce resource contention in NUMA multi-core systems. The Journal of Systems and Software, 121(??):72–87, November 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216301376.

Chen:2018:TCP

[CZC^{+}18] Jinfu Chen, Lili Zhu, Tsong Yueh Chen, Dave Towey, Fei-Ching Kuo, Rubing Huang, and Yuchi Guo. Test case prioritization for objectoriented software: an adaptive random sequence approach based on clustering. The Journal of Systems and Software, 135(??):107–125, January 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217302170.

Chen:2015:TEE

[CZG^{+}15] Huangke Chen, Xiaomin Zhu, Hui Guo, Jianghan Zhu, Xiao Qin, and Jianhong Wu. Towards energy-efficient scheduling for real-time tasks under uncertain cloud computing environment. The Journal of Systems and Software, 99(??):20–35, January 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001903.

deAlfonso:2017:CBV

[dACM17] Carlos de Alfonso, Amanda Calatrava, and Germán Moltó. Container-based virtual elastic clusters. The Journal of Systems and Software, 127(??):1– 11, May 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217300146.

Duttagupta:2019:HMM

[DAG19] Subhasri Duttagupta, Varsha Apte, and Devidas Gawali. M^{3} — a hybrid measurementmodeling approach for CPU-bound applications on cross-platform architectures. The Journal of Systems and Software, 156(??):232–245, October 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121930144X.

Saraiva:2015:CMA

[dAGSdFS^{+}15] Juliana de A. G. Saraiva, Micael S. de França, Sérgio C. B. Soares, Fernando J. C. L. Filho, and Renata M. C. R. de Souza. Classifying metrics for assessing Object-Oriented Software Maintainability: a family of metrics’ catalogs. The Journal of Systems and Software, 103(??):85–101, May 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215000126.

deAlmeida:2018:ISI

[dAK18] Eduardo Santana de Almeida and Georgia M. Kapitsaki. Introduction to the special issue on “Software Reuse”. The Journal of Systems and Software, 137(??):216, March 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0164121217302935.

Daneva:2017:SBL

[Dan17] Maya Daneva. Striving for balance: a look at gameplay requirements of massively multiplayer online role-playing games. The Journal of Systems and Software, 134(??):54–75, December 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301644.

Diaz:2014:FMI

[DAR14] Jesus Diaz, David Arroyo, and Francisco B. Rodriguez. A formal methodology for integral security design and verification of network protocols. The Journal of Systems and Software, 89(??):87–98, March 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002355.

dAmorim:2012:MAU

[dB12] Fernanda d’Amorim and Paulo Borba. Modularity analysis of use case implementations. The Journal of Systems and Software, 85(4):1012–1027, April 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002950.

Denardin:2011:GRH

[DBCdP11] Gustavo Weber Denardin, Carlos Henrique Barriquello, Alexandre Campos, and Ricardo Nederson do Prado. A geographic routing hybrid approach for void resolution in wireless sensor networks. The Journal of Systems and Software, 84(10):1577–1590, October 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211000732.

Dogan:2014:WAT

[DBCG14] Serdar Dogan, Aysu Betin-Can, and Vahid Garousi. Web application testing: a systematic literature review. The Journal of Systems and Software, 91(??):174– 201, May 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0164121214000223.

Coninck:2018:DMF

[DBL^{+}18] Elias De Coninck, Steven Bohez, Sam Leroux, Tim Verbelen, Bert Vankeirsbilck, Pieter Simoens, and Bart Dhoedt. DI-ANNE: a modular framework for designing, training and deploying deep neural networks on heterogeneous distributed infrastructure. The Journal of Systems and Software, 141(??):52–65, July 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300487.

Djoudi:2016:FFC

[DBZ16] Brahim Djoudi, Chafia Bouanaka, and Nadia Zeghib. A formal framework for context-aware systems specification and verification. The Journal of Systems and Software, 122(??):445–462, December 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002599.

Dawes:2011:CDP

[DC11] Brett Dawes and Kwan-Wu Chin. A comparison of deterministic and probabilistic methods for indoor localization. The Journal of Systems and Software, 84(3):442–451, March 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Dennehy:2017:GFA

[DC17] Denis Dennehy and Kieran Conboy. Going with the flow: an activity theory analysis of flow techniques in software development. The Journal of Systems and Software, 133(??):160–173, November 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216302011.

Deb:2016:EFS

[DCG16] Novarun Deb, Nabendu Chaki, and Aditya Ghose. Extracting finite state models from i^{∗} models. The Journal of Systems and Software, 121(??):265–280, November 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0164121216300048.

Drury:2012:ODM

[DCP12] Meghann Drury, Kieran Conboy, and Ken Power. Obstacles to decision making in Agile software development teams. The Journal of Systems and Software, 85(6): 1239–1254, June 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121212000374.

deCarvalho:2010:SFP

[dCPV10] André B. de Carvalho, Aurora Pozo, and Silvia Regina Vergilio. A symbolic fault-prediction model based on multiobjective particle swarm optimization. The Journal of Systems and Software, 83(5):868–882, May 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Dragicevic:2017:BNM

[DCT17] Srdjana Dragicevic, Stipe Celar, and Mili Turic. Bayesian network model for task effort estimation in agile software development. The Journal of Systems and Software, 127(??):109–119, May 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217300171.

DArco:2014:MIC

[DDD14] P. D’Arco, R. De Prisco, and A. De Santis. Measureindependent characterization of contrast optimal visual cryptography schemes. The Journal of Systems and Software, 95(??):89–99, September 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214000995.

Guglielmo:2013:IMD

[DDF^{+}13] Giuseppe Di Guglielmo, Luigi Di Guglielmo, Andreas Foltinek, Masahiro Fujita, Franco Fummi, Cristina Marconcini, and Graziano Pravadelli. On the integration of modeldriven design and dynamic assertion-based verification for embedded software. The Journal of Systems and Software, 86(8):2013–2033, August 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002506.

Daneva:2014:ERM

[DDMP14] Maya Daneva, Daniela Damian, Alessandro Marchetto, and Oscar Pastor. Empirical research methodologies and studies in Requirements Engineering: How far did we come? The Journal of Systems and Software, 95(??):1–9, September 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001460.

Drosatos:2014:PPC

[DEA^{+}14] George Drosatos, Pavlos S. Efraimidis, Ioannis N. Athanasiadis, Matthias Stevens, and Ellie D’Hondt. Privacypreserving computation of participatory noise maps in the cloud. The Journal of Systems and Software, 92(??):170– 183, June 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214000430.

Dam:2016:CMM

[DEW^{+}16] Hoa Khanh Dam, Alexander Egyed, Michael Winikoff, Alexander Reder, and Roberto E. Lopez-Herrejon. Consistent merging of model versions. The Journal of Systems and Software, 112(??):137–155, February 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121500134X.

Delamo:2015:DOS

[DFCPSF15] Manuel Delamo, Santiago Felici-Castell, Juan J. Pérez-Solano, and Andrew Foster. Designing an open source maintenance-free Environmental Monitoring Application for Wireless Sensor Networks. The Journal of Systems and Software, 103(??):238– 247, May 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215000400.

Durelli:2013:SSY

[DFG^{+}13] Vinicius Humberto Serapilha Durelli, Rodrigo Fraxino Araujo, Marco Aurelio Graciotto Silva, Rafael Alves Paes de Oliveira, Jose Carlos Maldonado, and Marcio Eduardo Delamaro. A scoping study on the 25 years of research into software testing in Brazil and an outlook on the future of the area. The Journal of Systems and Software, 86(4):934–950, April 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121200283X.

Dong:2019:EET

[DFJ19] Minggang Dong, Lili Fan, and Chao Jing. ECOS: an efficient taskclustering based costeffective aware scheduling algorithm for scientific workflows execution on heterogeneous cloud systems. The Journal of Systems and Software, 158(??):??, December 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301797.

Dadeau:2018:CBT

[DGBE18] Frederic Dadeau, Alain Giorgetti, Fabrice Bouquet, and Ivan Enderlin. Contract-based testing for PHP with Praspel. The Journal of Systems and Software, 136(??):209–222, February 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121730119X.

Drury-Grogan:2017:EDC

[DGCA17] Meghann L. Drury-Grogan, Kieran Conboy, and Tom Acton. Examining decision characteristics & challenges for agile software development. The Journal of Systems and Software, 131 (??):248–265, September 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217301103.

Gonzalez-Ladron-de-Guevara:2016:UID

[dGFDL16] Fernando González-Ladrón de Guevara, Marta Fernández-Diego, and Chris Lokan. The usage of ISBSG data fields in software effort estimation: a systematic mapping study. The Journal of Systems and Software, 113(??):188–215, March 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002642.

Dhungana:2010:SMS

[DGRN10] Deepak Dhungana, Paul Grünbacher, Rick Rabiser, and Thomas Neumayer. Structuring the modeling space and supporting evolution in software product line engineering. The Journal of Systems and Software, 83(7):1108–1122, July 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Diskin:2016:TDT

[DGWC16] Zinovy Diskin, Hamid Gholizadeh, Arif Wider, and Krzysztof Czarnecki. A three-dimensional taxonomy for bidirectional model synchronization. The Journal of Systems and Software, 111 (??):298–322, January 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121500120X.

Du:2013:SRW

[DH13] YaJun Du and YuFeng Hai. Semantic ranking of web pages based on formal concept analysis. The Journal of Systems and Software, 86(1):187–197, January 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002154.

Du:2011:OQA

[DHC^{+}11] Zhihui Du, Jingkun Hu, Yinong Chen, Zhili Cheng, and Xiaoying Wang. Optimized QoSaware replica placement heuristics and applications in astronomy data grid. The Journal of Systems and Software, 84(7): 1224–1232, July 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Delot:2014:ISI

[DIB14] Thierry Delot, Sergio Ilarri, and Cristian Borcea. Introduction to the special issue on middleware for mobile data management. The Journal of Systems and Software, 92(??):1– 2, June 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0164121214000673.

Damaiyanti:2017:SQS

[DII^{+}17] Titus Irma Damaiyanti, Ardi Imawan, Fitri Indra Indikawati, Yoon-Ho Choi, and Joonho Kwon. A similarity query system for road traffic data based on a NoSQL document store. The Journal of Systems and Software, 127(??):28–51, May 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217300158.

Dharavath:2015:ERB

[DK15a] Ramesh Dharavath and Chiranjeev Kumar. Entity resolution based EM for integrating heterogeneous distributed probabilistic data. The Journal of Systems and Software, 107(??):93–109, September 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001077.

Dharavath:2015:SGT

[DK15b] Ramesh Dharavath and Chiranjeev Kumar. A scalable generic transaction model scenario for distributed NoSQL databases. The Journal of Systems and Software, 101(??):43–58, March 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002684.

Debnath:2019:MMD

[DKP^{+}19] Hillol Debnath, Mohammad A. Khan, Nafize R. Paiker, Xiaoning Ding, Narain Gehani, Reza Curtmola, and Cristian Borcea. The Moitree middleware for distributed mobile-cloud computing. The Journal of Systems and Software, 157(??):??, November 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301542.

deAlmeidaMaia:2013:ITB

[dL13] Marcelo de Almeida Maia and Raquel Fialho Lafetá. On the impact of trace-based feature location in the performance of software maintainers. The Journal of Systems and Software, 86(4):1023–1037, April 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121200341X.

DiFrancesco:2019:AMS

[DLM19] Paolo Di Francesco, Patricia Lago, and Ivano Malavolta. Architecting with microservices: a systematic mapping study. The Journal of Systems and Software, 150(??):77–97, April 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300019.

Du:2013:UAS

[DLW^{+}13] Jian Du, Jing Lu, Dong Wu, Huiping Li, and Jie Li. User acceptance of software as a service: Evidence from customers of China’s leading e-commerce company, Alibaba. The Journal of Systems and Software, 86(8):2034–2044, August 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000587.

DeMatteis:2017:PEE

[DM17a] Tiziano De Matteis and Gabriele Mencagli. Proactive elasticity and energy awareness in data stream processing. The Journal of Systems and Software, 127(??):302– 319, May 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216301467.

Dunne:2017:OCR

[DM17b] Jonathan Dunne and David Malone. Obscured by the cloud: a resource allocation framework to model cloud outage events. The Journal of Systems and Software, 131 (??):218–229, September 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217301231.

DeFaveri:2018:MPD

[DMA18] Cristiano De Faveri, Ana Moreira, and Vasco Amaral. Multi-paradigm deception modeling for cyber defense. The Journal of Systems and Software, 141(??):32–51, July 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300499.

deMelo:2019:CIS

[dMCR19] Caio Batista de Melo, André Luiz Fernandes Cançado, and Genaína Nunes Rodrigues. Characterization of implied scenarios as families of common behavior. The Journal of Systems and Software, 158(??):??, December 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301992.

Dargie:2011:TCP

[DMSG11] Waltenegus Dargie, Rami Mochaourab, Alexander Schill, and Lin Guan. A topology control protocol based on eligibility and efficiency metrics. The Journal of Systems and Software, 84(1):2–11, January 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

SilveiraNeto:2013:YSE

[dMSSS^{+}13] Paulo Anselmo da Mota Silveira Neto, Joás Sousa Gomes, Eduardo Santana de Almeida, Jair Cavalcanti Leite, Thais Vasconcelos Batista, and Larissa Leite. 25 years of software engineering in Brazil: Beyond an insider’s view. The Journal of Systems and Software, 86(4):872–889, April 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002981.

Dingsoyr:2012:DAM

[DNBM12] Torgeir Dingsøyr, Sridhar Nerur, VenuGopal Balijepally, and Nils Brede Moe. A decade of agile methodologies: Towards explaining agile software development. The Journal of Systems and Software, 85(6):1213–1221, June 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000532.

doNascimento:2018:HBA

[dNPM18] Dimas Cassimiro do Nascimento, Carlos Eduardo Santos Pires, and Demetrio Gomes Mestre. Heuristic-based approaches for speeding up incremental record linkage. The Journal of Systems and Software, 137(??):335–354, March 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302972.

Durisic:2013:MIC

[DNSH13] Darko Durisic, Martin Nilsson, Mirosław Staron, and Jörgen Hansson. Measuring the impact of changes to the complexity and coupling properties of automotive software systems. The Journal of Systems and Software, 86(5):1275– 1293, May 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212003330.

deOliveira:2013:UAS

[dOCS13] Romulo Silva de Oliveira, Andreu Carminati, and Renan Augusto Starke. On using adversary simulators to evaluate global fixed-priority and FPZL scheduling of multiprocessors. The Journal of Systems and Software, 86(2):403–411, February 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121200252X.

deOliveira:2019:FNH

[dOFB^{+}19] Marcos César de Oliveira, Davi Freitas, Rodrigo Bonifácio, Gustavo Pinto, and David Lo. Finding needles in a haystack: Leveraging co-change dependencies to recommend refactorings. The Journal of Systems and Software, 158(??):??, December 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301943.

Durelli:2016:WEP

[DOL^{+}16] Vinicius H. S. Durelli, Jeff Offutt, Nan Li, Marcio E. Delamaro, Jin Guo, Zengshu Shi, and Xinge Ai. What to expect of predicates: an empirical analysis of predicates in real world programs. The Journal of Systems and Software, 113(??):324–336, March 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0164121215002885.

Neto:2019:ESA

[dONTF^{+}19] Francisco Gomes de Oliveira Neto, Richard Torkar, Robert Feldt, Lucas Gren, Carlo A. Furia, and Ziwei Huang. Evolution of statistical analysis in empirical software engineering research: Current state and steps forward. The Journal of Systems and Software, 156(??):246–267, October 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301451.

deOliveira:2017:ELL

[dOSdAdSG17] Raphael Pereira de Oliveira, Alcemir Rodrigues Santos, Eduardo Santana de Almeida, and Gecynalda Soares da Silva Gomes. Evaluating Lehman’s Laws of software evolution within software product lines industrial projects. The Journal of Systems and Software, 131(??):347–365, September 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121216301339.

Dikert:2016:CSF

[DPL16] Kim Dikert, Maria Paasivaara, and Casper Lassenius. Challenges and success factors for largescale agile transformations: a systematic literature review. The Journal of Systems and Software, 119(??):87–108, September 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216300826.

Lima:2019:SMS

[dPLV19] Jackson Antonio do Prado Lima and Silvia Regina Vergilio. A systematic mapping study on higher order mutation testing. The Journal of Systems and Software, 154(??):92– 109, August 2019. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219300809.

Devroey:2018:MBM

[DPP^{+}18] Xavier Devroey, Gilles Perrouin, Mike Papadakis, Axel Legay, Pierre-Yves Schobbens, and Patrick Heymans.

Model-based mutant equivalence detection using automata language equivalence and simulations. The Journal of Systems and Software, 141(??):1–15, July 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300475.

Delen:2019:FMI

[DPVvV19] G. P. A. J. Delen, R. J. Peters, C. Verhoef, and S. F. M. van Vlijmen. Foundations for measuring IToutsourcing success and failure. The Journal of Systems and Software, 156(??):113–125, October 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301360.

Do:2012:CSS

[DR12] Tien Van Do and Csaba Rotter. Comparison of scheduling schemes for on-demand IaaS requests. The Journal of Systems and Software, 85(6):1400–1408, June 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212000052.

Dominguez-Rios:2019:EAA

[DRCA^{+}19] Miguel Ángel Domínguez-Ríos, Francisco Chicano, Enrique Alba, Isabel del Águila, and José del Sagrado. Efficient anytime algorithms to solve the bi-objective Next Release Problem. The Journal of Systems and Software, 156(??):217–231, October 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301414.

Dehuri:2012:ISO

[DRCG12] Satchidananda Dehuri, Rahul Roy, Sung-Bae Cho, and Ashish Ghosh. An improved swarm optimized functional link artificial neural network (ISO-FLANN) for classification. The Journal of Systems and Software, 85(6):1333–1345, June 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212000210.

Demuth:2016:CEM

[DRELHE16] Andreas Demuth, Markus Riedl-Ehrenleitner, Roberto E. Lopez-Herrejon, and Alexander Egyed. Co-evolution of metamodels and models through consistent change propagation. The Journal of Systems and Software, 111(??):281– 297, January 2016. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121215000564.

deRoo:2013:MAF

[dRSBA13] Arjan de Roo, Hasan Sözer, Lodewijk Bergmans, and Mehmet Aksit. MOO: an architectural framework for runtime optimization of multiple system objectives in embedded control software. The Journal of Systems and Software, 86(10):2502–2519, October 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000897.

Deypir:2012:DLS

[DS12] Mahmood Deypir and Mohammad Hadi Sadreddini. A dynamic layout of sliding window for frequent itemset mining over data streams. The Journal of Systems and Software, 85(3):746–759, March 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002573.

Dehury:2016:DIN

[DS16a] Chinmaya Kumar Dehury and Prasan Kumar Sahoo. Design and implementation of a novel service management framework for IoT devices in cloud. The Journal of Systems and Software, 119 (??):149–161, September 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300887.

Duran:2016:RRR

[DS16b] Francisco Durán and Gwen Salaün. Robust and reliable reconfiguration of cloud applications. The Journal of Systems and Software, 122(??):524–537, December 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002101.

Alves:2017:MLF

[dSACdLF17] Erickson H. da S. Alves, Lucas C. Cordeiro, and Eddie B. de L. Filho. A method to localize faults in concurrent C programs. The Journal of Systems and Software, 132(??):336–352, October 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217300614.

deSilva:2012:CSA

[dSB12] Lakshitha de Silva and Dharini Balasubramaniam. Controlling software architecture erosion: a survey. The Journal of Systems and Software, 85(1):132–151, January 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002044.

daSilva:2014:SPL

[dSdMSNO^{+}14] Ivonei Freitas da Silva, Paulo Anselmo da Mota Silveira Neto, Pádraig O’Leary, Eduardo Santana de Almeida, and Silvio Romero de Lemos Meira. Software product line scoping and requirements engineering in a small and medium-sized enterprise: an industrial case study. The Journal of Systems and Software, 88(??):189–206, February 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002598.

daSilva:2012:TUU

[dSF12] Fabio Q. B. da Silva and A. César C. França. Towards understanding the underlying structure of motivational factors for software engineers to guide the definition of motivational programs. The Journal of Systems and Software, 85(2):216–226, February 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121210003390.

Dietrich:2017:CBA

[DSGS17] Robert Dietrich, Felix Schmitt, Alexander Grund, and Jonas Stolle. Critical-blame analysis for OpenMP 4.0 offloading on Intel Xeon Phi. The Journal of Systems and Software, 125(??): 381–388, March 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

//www.sciencedirect.
com/science/article/
pii/S0164121215002940.

Soares:2011:URM

[dSSVV11] Michel dos Santos Soares, Jos Vrancken, and Alexander Verbraeck. User requirements modeling and analysis of softwareintensive systems. The Journal of Systems and Software, 84(2):328–339, February 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Dutta:2015:SIS

[Dut15] Kaushik Dutta. Special issue on software architectures and systems for Big data. The Journal of Systems and Software, 102(??):145, April 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000175.

Diaz:2010:GBP

[DV10] Oscar Diaz and Felipe M.

Villoria. Generating blogs out of product catalogues: an MDE approach. The Journal of Systems and Software, 83 (10):1970–1982, October 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Daneva:2013:ARP

[DvdVA^{+}13] Maya Daneva, Egbert van der Veen, Chintan Amrit, Smita Ghaisas, Klaas Sikkel, Ramesh Kumar, Nirav Ajmeri, Uday Ramteerthkar, and Roel Wieringa. Agile requirements prioritization in large-scale outsourced system projects: an empirical study. The Journal of Systems and Software, 86(5):1333–1353, May 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212003536.

Danglot:2019:SLS

[DVPY^{+}19] Benjamin Danglot, Oscar Vera-Perez, Zhongxing Yu, Andy Zaidman, Martin Monperrus, and Benoit Baudry. A snowballing literature study on test amplification. The Journal of Systems and Software, 157(??):??, November 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301736.

delVal:2013:PCS

[dVRB13] E. del Val, M. Rebollo, and V. Botti. Promoting cooperation in serviceoriented MAS through social plasticity and incentives. The Journal of Systems and Software, 86(2):520–537, February 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002725.

DeConinck:2016:DAS

[DVV^{+}16] Elias De Coninck, Tim Verbelen, Bert Vankeirsbilck, Steven Bohez, Pieter Simoens, and Bart Dhoedt. Dynamic autoscaling and scheduling of deadline constrained service workloads on IaaS clouds. The Journal of Systems and Software, 118(??):101–114, August 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300449.

Debroy:2011:EAT

[DW11] Vidroha Debroy and W. Eric Wong. On the estimation of adequate test set size using fault failure rates. The Journal of Systems and Software, 84 (4):587–602, April 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Debroy:2014:CMF

[DW14] Vidroha Debroy and W. Eric Wong. Combining mutation and fault localization for automated program debugging. The Journal of Systems and Software, 90(??):45–60, April 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213002616.

Ding:2017:SCA

[DWC17] Zuohua Ding, Zhijie Wei, and Haibo Chen. A software cybernetics approach to self-tuning performance of on-line transaction processing systems. The Journal of Systems and Software, 124(??):247–259, February 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000832.

Daraghmi:2015:SWB

[DY15] Eman Yasser Daraghmi and Shyan-Ming Yuan. A small world based overlay network for improving dynamic load-balancing. The Journal of Systems and Software, 107 (??):187–203, September 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001181.

Delgado:2019:RTC

[DYC19] Raimarius Delgado, Bum-Jae You, and Byoung Wook Choi. Real-time control architecture based on Xenomai using ROS packages for a service robot. The Journal of Systems and Software, 151(??):8–19, May 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300160.

Dong:2014:PMD

[DZT^{+}14] Bo Dong, Qinghua Zheng, Feng Tian, Kuo-Ming Chao, Nick Godwin, Tian Ma, and Haipeng Xu. Performance models and dynamic characteristics analysis for HDFS write and read operations: a systematic view. The Journal of Systems and Software, 93(??):132–151, July 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121400065X.

Eslami:2011:SIS

[EA11] Z. Eslami and J. Zarepour Ahmadabadi. Secret image sharing with authentication-chaining and dynamic embedding. The Journal of Systems and Software, 84(5):803– 809, May 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

El-Attar:2012:TDC

[EA12] Mohamed El-Attar. Towards developing consistent misuse case models. The Journal of Systems and Software, 85(2):323–339, February 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0164121211002160.

El-Attar:2014:USR

[EA14] Mohamed El-Attar. Using SMCD to reduce inconsistencies in misuse case models: a subjectbased empirical evaluation. The Journal of Systems and Software, 87(??):104–118, January 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002458.

El-Attar:2019:EEI

[EA19] Mohamed El-Attar. Evaluating and empirically improving the visual syntax of use case diagrams. The Journal of Systems and Software, 156(??):136–163, October 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301402.

Edagawa:2011:FPM

[EAH^{+}11] T. Edagawa, T. Akaike, Y. Higo, S. Kusumoto, S. Hanabusa, and T. Shibamoto. Function point measurement from Web application source code based on screen transitions and database accesses. The Journal of Systems and Software, 84(6):976–984, June 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ebert:2014:SPM

[EB14a] Christof Ebert and Sjaak Brinkkemper. Software product management — an industry evaluation. The Journal of Systems and Software, 95(??):10–18, September 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214000156.

Eklund:2014:AEO

[EB14b] Ulrik Eklund and Jan Bosch. Architecture for embedded open software ecosystems. The Journal of Systems and Software, 92(??):128–142, June 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214000211.

Erdemir:2014:LBM

[EB14c] Ural Erdemir and Feza Buzluca. A learningbased module extraction method for objectoriented systems. The Journal of Systems and Software, 97(??):156– 177, November 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121214001599.

Edison:2013:TIM

[EbAT13] Henry Edison, Nauman bin Ali, and Richard Torkar. Towards innovation measurement in the software industry. The Journal of Systems and Software, 86(5):1390– 1407, May 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000058.

English:2010:RRE

[EBC10] Michael English, Jim Buckley, and Tony Cahill. A replicated and refined empirical study of the use of friends in C++ software. The Journal of Systems and Software, 83 (11):2275–2286, November 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

ElMenshawy:2018:MCR

[EBEL18] Mohamed El Menshawy, Jamal Bentahar, Warda El Kholy, and Amine Laarej. Model checking real-time conditional commitment logic using transformation. The Journal of Systems and Software, 138(??):189– 205, April 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217303126.

Escheikh:2017:VWA

[EBJ17] Mohamed Escheikh, Kamel Barkaoui, and Hana Jouini. Versatile workloadaware power management performability analysis of server virtualized systems. The Journal of Systems and Software, 125(??):365– 379, March 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216302588.

Erola:2011:ESN

[ECRVMS11] Arnau Erola, Jordi Castellà-Roca, Alexandre Viejo, and Josep M. Mateo-Sanz. Exploiting social networks to provide privacy in personalized Web search. The Journal of Systems and Software, 84(10):1734–1745, October 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001117.

Ebert:2015:ESE

[ECS15] Felipe Ebert, Fernando Castor, and Alexander Serebrenik. An exploratory study on exception handling bugs in Java programs. The Journal of Systems and Software, 106(??):82– 101, August 2015. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121215000862.

El-Emam:2013:NSA

[EEAZ13] Nameer N. El-Emam and Rasheed Abdul Shaheed Al-Zubidy. New steganography algorithm to conceal a large amount of secret message using hybrid adaptive neural networks with modified adaptive genetic algorithm. The Journal of Systems and Software, 86(6):1465–1481, June 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212003317.

Eler:2016:ESQ

[EED16] Marcelo M. Eler, Andre T. Endo, and Vinicius H. S. Durelli. An empirical study to quantify the characteristics of Java programs that may influence symbolic execution from a unit testing perspective. The Journal of Systems and Software, 121(??):281–297, November 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216000868.

El-Fakih:2017:AEF

[EFSJM17] Khaled El-Fakih, Adenilso Simao, Noshad Jadoon, and Jose Carlos Maldonado. An assessment of extended finite state machine test selection criteria. The Journal of Systems and Software, 123(??):106–118, January 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216301923.

Entrialgo:2011:DAR

[EGG^{+}11] Joaquín Entrialgo, Daniel F.

García, Javier García, Manuel García, Pablo Valledor, and Mohammad S. Obaidat. Dynamic adaptation of response-time models for QoS management in autonomic systems. The Journal of Systems and Software, 84(5):810–820, May 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

El-Gazzar:2016:UCC

[EGHO16] Rania El-Gazzar, Eli Hustad, and Dag H. Olsen. Understanding cloud computing adoption issues: a Delphi study approach. The Journal of Systems and Software, 118(??):64–84, August 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121630036X.

Escalona:2011:OTG

[EGM^{+}11] M. J. Escalona, J. J.

Gutierrez, M. Mejías, G. Aragón, I. Ramos, J. Torres, and F. J. Domínguez. An overview on test generation from functional requirements.

The Journal of Systems and Software, 84(8): 1379–1393, August 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Emara:2019:DDM

[EH19] Tamer Z. Emara and Joshua Zhexue Huang. A distributed data management system to support large-scale data analysis. The Journal of Systems and Software, 148(??):105–115, February 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302437.

Eracar:2012:SCT

[EK12] Yönet A. Eracar and Mieczysław M. Kokar. Self-control of the time complexity of a constraint satisfaction problem solver program. The Journal of Systems and Software, 85 (12):2697–2706, December 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001501.

Eshragh:2013:AAB

[EK13] Faeze Eshragh and Mehdi Kargahi. Analytical architecture-based performability evaluation of real-time software systems. The Journal of Systems and Software, 86(1):233–246, January 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002300.

Elizondo:2010:CCC

[EL10] Perla Velasco Elizondo and Kung-Kiu Lau. A catalogue of component connectors to support development with reuse. The Journal of Systems and Software, 83(7): 1165–1178, July 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Etemaadi:2013:QDO

[ELHC13] Ramin Etemaadi, Kenneth Lind, Rogardt Heldal, and Michel R. V. Chaudron. Qualitydriven optimization of system architecture: Industrial case study on an automotive sub-system. The Journal of Systems and Software, 86 (10):2559–2573, October 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001404.

Earl:2017:NEP

[EMBS17] Christopher Earl, Matthew Might, Abhishek Bagusetty, and James C. Sutherland. Nebo: an efficient, parallel, and portable domain-specific language for numerically solving partial differential equations. The Journal of Systems and Software, 125(??):389– 400, March 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

//www.sciencedirect.
com/science/article/
pii/S0164121216000182.

Esfahani:2011:ADS

[EMSU11] Faramarz Safi Esfahani, Masrah Azrifah Azmi Murad, Md. Nasir B. Sulaiman, and Nur Izura Udzir. Adaptable Decentralized Service Oriented Architecture. The Journal of Systems and Software, 84(10):1591–1617, October 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0164121211000744.

Eichelberger:2014:FRM

[ES14] Holger Eichelberger and Klaus Schmid. Flexible resource monitoring of Java programs. The Journal of Systems and Software, 93(??):163– 186, July 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214000533.

Eckert:2019:ATI

[ESM19a] Remo Eckert, Matthias Stuermer, and Thomas Myrach. Alone or together? Inter-organizational affiliations of open source communities. The Journal of Systems and Software, 149(??):250–262, March 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302693.

Edded:2019:CCA

[ESM^{+}19b] Sabrine Edded, Sihem Ben Sassi, Raúl Mazo, Camille Salinesi, and Henda Ben Ghezala. Collaborative configuration approaches in software product lines engineering: a systematic mapping study. The Journal of Systems and Software, 158(??):??, December 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301967.

Estefo:2019:ROS

[ESRF19] Pablo Estefo, Jocelyn Simmonds, Romain Robbes, and Johan Fabry. The Robot Operating System: Package reuse and community dynamics. The Journal of Systems and Software, 151(??):226–242, May 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300342.

Edison:2018:LIS

[ESWA18] Henry Edison, Nina M. Smørsgård, Xiaofeng Wang, and Pekka Abrahamsson. Lean internal startups for software product innovation in large companies: Enablers and inhibitors. The Journal of Systems and Software, 135(??):69–87, January 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302157.

Egorova:2010:AVP

[ETM10] Evgenia Egorova, Marco Torchiano, and Maurizio Morisio. Actual vs. perceived effect of software engineering practices in the Italian industry. The Journal of Systems and Software, 83 (10):1907–1916, October 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Evertsz:2015:FMT

[ETYL15] Rick Evertsz, John Thangarajah, Nitin Yadav, and Thanh Ly. A framework for modelling tactical decision-making in autonomous systems. The Journal of Systems and Software, 110 (??):222–238, December 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001892.

Escalona:2013:DWR

[EUR^{+}13] M. J. Escalona, M. Urbieta, G. Rossi, J. A. Garcia-Garcia, and E. Robles Luna. Detecting Web requirements conflicts and inconsistencies under a model-based perspective. The Journal of Systems and Software, 86 (12):3024–3038, December 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001337.

Elbendak:2011:PUC

[EVR11] Mosa Elbendak, Paul Vickers, and Nick Rossiter. Parsed use case descriptions as a basis for object-oriented class model generation. The Journal of Systems and Software, 84(7):1209– 1223, July 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Espinha:2015:WAG

[EZG15] Tiago Espinha, Andy Zaidman, and Hans-Gerhard Gross. Web API growing pains: Loosely coupled yet strongly tied. The Journal of Systems and Software, 100(??):27–43, February 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0164121214002180.

Elbouabidi:2014:EDV

[EZOK14] Imen Elbouabidi, Faouzi Zarai, Mohammad S. Obaidat, and Lotfi Kamoun. An efficient design and validation technique for secure handover between 3GPP LTE and WLANs systems. The Journal of Systems and Software, 91(??):163–173, May 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214000193.

Erfani:2016:CAS

[EZRK16] Mostafa Erfani, Mohammadnaser Zandi, Juergen Rilling, and Iman Keivanloo. Contextawareness in the software domain — a Semantic Web enabled modeling approach. The Journal of Systems and Software, 121(??):345–357, November 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000558.

Fernandez:2013:EVU

[FAI13] Adrian Fernandez, Silvia Abrahão, and Emilio Insfran. Empirical validation of a usability inspection method for model-driven Web development. The Journal of Systems and Software, 86(1):161–186, January 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121200218X.

Fraser:2015:MAW

[FAM15] Gordon Fraser, Andrea Arcuri, and Phil McMinn. A Memetic Algorithm for whole test suite generation. The Journal of Systems and Software, 103(??):311– 327, May 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001216.

Fahmideh:2018:REK

[FB18] Mahdi Fahmideh and Ghassan Beydoun. Reusing empirical knowledge during cloud computing adoption. The Journal of Systems and Software, 138(??):124–157, April 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217303047.

Ferreira:2012:ITO

[FBB^{+}12] Kecia A. M. Ferreira, Mariza A. S. Bigonha, Roberto S. Bigonha, Luiz F. O. Mendes, and Heitor C. Almeida. Identifying thresholds for object-oriented software metrics. The Journal of Systems and Software, 85(2):244–257, February 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001385.

Fadhel:2015:CMF

[FBB15] Ameni Ben Fadhel, Domenico Bianculli, and Lionel Briand. A comprehensive modeling framework for role-based access control policies. The Journal of Systems and Software, 107 (??):110–126, September 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001041.

Feld:2018:SSA

[FBD^{+}18] Timo Feld, Alessandro Biondi, Robert I. Davis, Giorgio Buttazzo, and Frank Slomka. A survey of schedulability analysis techniques for ratedependent tasks. The Journal of Systems and Software, 138(??):100– 107, April 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217303102.

Franco:2016:ISA

[FCB^{+}16] João M. Franco, Francisco Correia, Raul Barbosa, Mário Zenha-Rela, Bradley Schmerl, and David Garlan. Improving self-adaptation planning through software architecture-based stochastic modeling. The Journal of Systems and Software, 115(??):42–60, May 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216000212.

Floch:2010:CEF

[FCC^{+}10] J. Floch, C. Carrez, P. Cieślak, M. Rój, R. T. Sanders, and M. M.

Shiaa. A comprehensive engineering framework for guaranteeing component compatibility. The Journal of Systems and Software, 83 (10):1759–1779, October 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Frantz:2012:PDE

[FCMJ12] Rafael Z. Frantz, Rafael Corchuelo, and Carlos Molina-Jiménez. A proposal to detect errors in Enterprise Application Integration solutions. The Journal of Systems and Software, 85(3):480–497, March 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002809.

Frantz:2016:DMS

[FCRF16] Rafael Z. Frantz, Rafael Corchuelo, and Fabricia Roos-Frantz. On the design of a maintainable software development kit to implement integration solutions. The Journal of Systems and Software, 111(??):89–104, January 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215001880.

Fabra:2012:AEB

[FDÁM12] J. Fabra, V. De Castro, P. Álvarez, and E. Marcos. Automatic execution of business process models: exploiting the benefits of model-driven engineering approaches. The Journal of Systems and Software, 85(3):607–625, March 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211002391.

Fabry:2016:ACA

[FDN^{+}16] Johan Fabry, Coen De Roover, Carlos Noguera, Steffen Zschaler, Awais Rashid, and Viviane Jonckers. AspectJ code analysis and verification with GASR. The Journal of Systems and Software, 117(??):528–544, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216300279.

Feitelson:2012:PDM

[Fei12] Dror G. Feitelson. Perpetual development: a model of the Linux kernel life cycle. The Journal of Systems and Software, 85(4):859–875, April 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002822.

Feng:2012:RDH

[FF12] Guorui Feng and Lingyan Fan. Reversible data hiding of high payload using local edge sensing prediction. The Journal of Systems and Software, 85(2):392–399, February 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002330.

Fontana:2014:PVP

[FFdRG^{+}14] Rafaela Mantovani Fontana, Isabela Mantovani Fontana, Paula Andrea da Rosa Garbuio, Sheila Reinehr, and Andreia Malucelli. Processes versus people: How should agile software development maturity be defined? The Journal of Systems and Software, 97(??):140–155, November 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214001587.

Filho:2019:PBM

[FFV19] Helson Luiz Jakubovski Filho, Thiago Nascimento Ferreira, and Silvia Regina Vergilio. Preference based multiobjective algorithms applied to the variability testing of software product lines. The Journal of Systems and Software, 151(??):194–209, May 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121930038X.

Femmer:2017:RQA

[FFWE17] Henning Femmer, Daniel Méndez Fernández, Stefan Wagner, and Sebastian Eder. Rapid quality assurance with Requirements Smells. The Journal of Systems and Software, 123(??):190–213, January 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216000789.

Fogli:2015:PAA

[FG15] Daniela Fogli and Giovanni Guida. A practical approach to the assessment of quality in use of corporate web sites. The Journal of Systems and Software, 99(??):52–65, January 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001964.

Candel:2019:DMD

[FGB^{+}19] Carlos Javier Fernández Candel, Jesús García Molina, Francisco Javier Bermúdez Ruiz, Jose Ramón Hoyos Barceló, Diego Sevilla Ruiz, and Benito José Cuesta Viera. Developing a modeldriven reengineering approach for migrating PL/ SQL triggers to Java: a practical experience. The Journal of Systems and Software, 151(??):38–64, May 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300214.

Fortuna:2010:QAE

[FGBC10] Rossella Fortuna, Luigi Alfredo Grieco, Gennaro Boggia, and Pietro Camarda. Quality adaptive end-to-end packet scheduling to avoid playout interruptions in Internet video streaming systems. The Journal of Systems and Software, 83(8):1489–1499, August 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Flouris:2017:ICE

[FGD^{+}17] Ioannis Flouris, Nikos Giatrakos, Antonios Deligiannakis, Minos Garofalakis, Michael Kamp, and Michael Mock. Issues in complex event processing: Status and prospects in the Big Data era. The Journal of Systems and Software, 127(??):217– 236, May 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300802.

Forsman:2015:AAL

[FGLI15] Mattias Forsman, Andreas Glad, Lars Lundberg, and Dragos Ilie. Algorithms for automated live migration of virtual machines. The Journal of Systems and Software, 101(??):110–126, March 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002751.

Fagerholm:2017:RMC

[FGMM17] Fabian Fagerholm, Alejandro Sanchez Guinea, Hanna Mäenpää, and Jürgen Münch. The RIGHT model for Continuous Experimentation. The Journal of Systems and Software, 123(??):292–305, January 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300024.

Folstad:2010:WDK

[FH10] Asbjørn Følstad and Kasper Hornbæk. Workdomain knowledge in usability evaluation: Experiences with Cooperative Usability Testing. The Journal of Systems and Software, 83(11):2019– 2030, November 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fan:2015:EFP

[FHL^{+}15] Ming Fan, Qiushi Han, Shuo Liu, Shaolei Ren, Gang Quan, and Shangping Ren. Enhanced fixed-priority real-time scheduling on multi-core platforms by exploiting task period relationship. The Journal of Systems and Software, 99(??):85–96, January 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001976.

Fagerholm:2018:DIE

[FHL^{+}18] Fabian Fagerholm, Arto Hellas, Matti Luukkainen, Kati Kyllönen, Sezin Yaman, and Hanna Mäenpää. Designing and implementing an environment for software start-up education: Patterns and antipatterns. The Journal of Systems and Software, 146(??):1–13, December 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301742.

Fan:2017:EML

[FHY17] Ming Fan, Qiushi Han, and Xiaokun Yang. Energy minimization for online real-time scheduling with reliability awareness. The Journal of Systems and Software, 127(??):168–176, May 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121730033X.

Falessi:2016:ISI

[FKA16] Davide Falessi, Philippe Kruchten, and Paris Avgeriou. Introduction to the special issue on technical debt in software systems. The Journal of Systems and Software, 120(??):154–155, October 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300978.

Feldmann:2019:MIM

[FKWVH19] S. Feldmann, K. Kernschmidt, M. Wimmer, and B. Vogel-Heuser. Managing inter-model inconsistencies in modelbased systems engineering: Application in automated production systems engineering. The Journal of Systems and Software, 153(??):105– 134, July 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219300639.

Fontana:2019:ASI

[FLRT19] Francesca Arcelli Fontana, Valentina Lenarduzzi, Riccardo Roveda, and Davide Taibi. Are architectural smells independent from code smells? An empirical study. The Journal of Systems and Software, 154(??):139–156, August 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301013.

faragardi:2018:REF

[fLSN18] Hamid Reza faragardi, Björn Lisper, Kristian Sandström, and Thomas Nolte. A resource efficient framework to run automotive embedded software on multicore ECUs. The Journal of Systems and Software, 139(??):64–83, May 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218300153.

Fasquel:2011:DPC

[FM11] Jean-Baptiste Fasquel and Johan Moreau. A design pattern coupling role and component concepts: Application to medical software. The Journal of Systems and Software, 84 (5):847–863, May 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Friginal:2016:MCA

[FMdAR16] Jesús Friginal, Miquel Martínez, David de Andrés, and Juan-Carlos Ruiz. Multi-criteria analysis of measures in benchmarking: Dependability benchmarking as a case study. The Journal of Systems and Software, 111(??):105–118, January 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002009.

Ferretti:2016:AWC

[FMPS16] Stefano Ferretti, Silvia Mirri, Catia Prandi, and Paola Salomoni. Automatic web content personalization through reinforcement learning. The Journal of Systems and Software, 121 (??):157–169, November 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000443.

Fontana:2011:URM

[FMR11] F. Arcelli Fontana, S. Maggioni, and C. Raibulet. Understanding the relevance of micro-structures for design patterns detection. The Journal of Systems and Software, 84 (12):2334–2347, December 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121100183X.

Fontana:2015:POF

[FMRM15] Rafaela Mantovani Fontana, Victor Meyer, Jr., Sheila Reinehr, and Andreia Malucelli. Progressive Outcomes: a framework for maturing in agile software development. The Journal of Systems and Software, 102(??):88– 108, April 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002908.

Fu:2018:LUA

[FNWL18] Xingbing Fu, Xuyun Nie, Ting Wu, and Fagen Li. Large universe attribute based access control with efficient decryption in cloud storage system. The Journal of Systems and Software, 135(??):157–164, January 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302510.

Felderer:2019:ISI

[FOR19] Michael Felderer, Helena Holmström Olsson, and Rick Rabiser. Introduction to the special issue on quality engineering and management of software-intensive systems. The Journal of Systems and Software, 149(??):533–534, March 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302796.

Feyzi:2018:FFI

[FP18] Farid Feyzi and Saeed Parsa. FPA-FL: Incorporating static faultproneness analysis into statistical fault localization. The Journal of Systems and Software, 136(??):39–58, February 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302583.

Fernandez:2019:ESE

[FP19] Daniel Méndez Fernández and Jan-Hendrik Passoth. Empirical software engineering: From discipline to interdiscipline. The Journal of Systems and Software, 148(??):170–179, February 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121830253X.

Fortier:2010:DVC

[FRGC10] Andrés Fortier, Gustavo Rossi, Silvia E. Gordillo, and Cecilia Challiol. Dealing with variability in context-aware mobile software. The Journal of Systems and Software, 83 (6):915–936, June 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Faisal:2014:HSC

[FS14a] Sidra Faisal and Mansoor Sarwar. Handling slowly changing dimensions in data warehouses. The Journal of Systems and Software, 94(??):151–160, August 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214000843.

Flores:2014:MCM

[FS14b] Huber Flores and Satish Narayana Srirama. Mobile Cloud Middleware. The Journal of Systems and Software, 92(??):82–94, June 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002318.

Fitzgerald:2017:CSE

[FS17] Brian Fitzgerald and Klaas-Jan Stol. Continuous software engineering: a roadmap and agenda. The Journal of Systems and Software, 123(??):176–189, January 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121215001430.

Fardbastani:2019:SCE

[FS19] Mohammad Ali Fardbastani and Mohsen Sharifi. Scalable complex event processing using adaptive load balancing. The Journal of Systems and Software, 149(??):305–317, March 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218302723.

Frankova:2011:DBP

[FSG^{+}11] Ganna Frankova, Magali Séguran, Florian Gilcher, Slim Trabelsi, Jörg Dörflinger, and Marco Aiello. Deriving business processes with service level agreements from early requirements. The Journal of Systems and Software, 84(8):1351–1363, August 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Figueiredo:2012:AEC

[FSGL12] Eduardo Figueiredo, Claudio Sant’Anna, Alessandro Garcia, and Carlos Lucena. Applying and evaluating concernsensitive design heuristics. The Journal of Systems and Software, 85(2):227–243, February 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002585.

Fang:2011:ICP

[FSGW11] Liming Fang, Willy Susilo, Chunpeng Ge, and Jiandong Wang. Interactive conditional proxy re-encryption with fine grain policy. The Journal of Systems and Software, 84(12):2293– 2302, December 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121211001579.

Fernandez-Sanchez:2017:IAE

[FSGYP17] Carlos Fernández-Sánchez, Juan Garbajosa, Agustín Yagüe, and Jennifer Perez. Identification and analysis of the elements required to manage technical debt by means of a systematic mapping study. The Journal of Systems and Software, 124(??):22–38, February 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216302138.

Fernandez-Salgado:2016:IPP

[FSPH^{+}16] Javier Fernández-Salgado, Pablo Parra, Michael Hauck, Agustín M. Hellín, Sebastián Sánchez-Prieto, Klaus Krogmann, and Óscar R. Polo. Integration of a preemptive priority based scheduler in the Palladio Workbench. The Journal of Systems and Software, 114(??):20–37, April 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215002915.

Fronza:2013:FPB

[FSS^{+}13] Ilenia Fronza, Alberto Sillitti, Giancarlo Succi, Mikko Terho, and Jelena Vlasenko. Failure prediction based on log files using Random Indexing and Support Vector Machines. The Journal of Systems and Software, 86(1):2–11, January 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0164121212001732.

Filippidis:2016:ISF

[FTC16] Christos Filippidis, Panayiotis Tsanakas, and Yiannis Cotronis. IKAROS: a scalable I/O framework for high-performance computing systems. The Journal of Systems and Software, 118(??):277– 287, August 2016. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121216300589.

Fokaefs:2012:IAE

[FTSC12] Marios Fokaefs, Nikolaos Tsantalis, Eleni Stroulia, and Alexander Chatzigeorgiou. Identification and application of Extract Class refactorings in object-oriented systems. The Journal of Systems and Software, 85(10):2241–2260, October 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001057.

Fuggetta:2012:CFU

[Fug12] Alfonso Fuggetta. 3 + 1 Challenges for the future of universities. The Journal of Systems and Software, 85(10):2417–2424, October 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001525.

Fay:2015:EMB

[FVHF^{+}15] Alexander Fay, Birgit Vogel-Heuser, Timo Frank, Karin Eckert, Thomas Hadlich, and Christian Diedrich. Enhancing a model-based engineering approach for distributed manufacturing automation systems with characteristics and design patterns. The Journal of Systems and Software, 101(??):221–235, March 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002866.

Fan:2012:ABS

[FWCS12] Chun-I Fan, Chien-Nan Wu, Wei-Kuei Chen, and Wei-Zhe Sun. Attributebased strong designatedverifier signature scheme. The Journal of Systems and Software, 85(4):944– 959, April 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002895.

Fan:2013:PNB

[FYCL13] Guisheng Fan, Huiqun Yu, Liqiong Chen, and Dongmei Liu. Petri net based techniques for constructing reliable service composition. The Journal of Systems and Software, 86(4):1089–1106, April 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212003214.

Ghapanchi:2011:AIP

[GA11] Amir Hossein Ghapanchi and Aybuke Aurum. Antecedents to IT personnel’s intentions to leave: a systematic literature review. The Journal of Systems and Software, 84(2):238–249, February 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Grunske:2013:QOS

[GA13] Lars Grunske and Aldeida Aleti. Quality optimisation of software architectures and design specifications. The Journal of Systems and Software, 86(10):2465–2466, October 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001416.

Guerra:2013:RAO

[GAKF13] Eduardo Guerra, Felipe Alves, Uirá Kulesza, and Clovis Fernandes. A reference architecture for organizing the internal structure of metadatabased frameworks. The Journal of Systems and Software, 86(5):1239– 1256, May 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212003366.

Galster:2014:VSA

[GAMW14] Matthias Galster, Paris Avgeriou, Tomi Männistö, and Danny Weyns. Variability in software architecture — state of the art. The Journal of Systems and Software, 91(??):1–2, May 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214000454.

Garcia:2013:SEB

[Gar13] Alessandro Garcia. Software engineering in Brazil: Retrospective and prospective views. The Journal of Systems and Software, 86(4):869– 871, April 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121300023X.

Gui:2015:DCM

[GAT15] Jinsong Gui, Maryam Ahmadi, and Fei Tong. Dynamically constructing and maintaining virtual access points in a macro cell with selfish nodes. The Journal of Systems and Software, 108(??):1–22, October 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001193.

Guasque:2016:RTH

[GBC16] Ana Guasque, Patricia Balbastre, and Alfons Crespo. Real-time hierarchical systems with arbitrary scheduling at global level. The Journal of Systems and Software, 119(??):70–86, September 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216300619.

Gavalas:2011:MAS

[GBCI11] Damianos Gavalas, Paolo Bellavista, Jiannong Cao, and Valérie Issarny. Mobile applications: Status and trends. The Journal of Systems and Software, 84(11):1823–1826, November 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211002019.

Godet-Bar:2012:SFC

[GBDCR12] Guillaume Godet-Bar, Sophie Dupuy-Chessa, and Dominique Rieu. Sonata: Flexible connections between interaction and business spaces. The Journal of Systems and Software, 85(5):1105– 1118, May 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211003219.

Grottke:2010:ISI

[GBG10] Michael Grottke, Doo-Hwan Bae, and Bjørn Axel Gran. Introduction to the special issue. The Journal of Systems and Software, 83(9):1553– 1554, September 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gerostathopoulos:2016:SAS

[GBH^{+}16] Ilias Gerostathopoulos, Tomas Bures, Petr Hnetynka, Jaroslav Keznikl, Michal Kit, Frantisek Plasil, and Noël Plouzeau. Self-adaptation in softwareintensive cyber-physical systems: From system goals to architecture configurations. The Journal of Systems and Software, 122(??):378–397, December 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000601.

Gomez:2013:UIT

[GC13] M. Gómez and J. Cervantes. User Interface Transition Diagrams for customer-developer communication improvement in software development projects. The Journal of Systems and Software, 86 (9):2394–2410, September 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213001003.

Goncalves:2018:SLR

[GCAH18] Enyo Gonçalves, Jaelson Castro, João Araújo, and Tiago Heineck. A systematic literature review of iStar extensions. The Journal of Systems and Software, 137(??):1–33, March 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217302741.

Garousi:2015:SSE

[GCBCD15] Vahid Garousi, Ahmet Coskunçay, Aysu Betin-Can, and Onur Demirörs. A survey of software engineering practices in Turkey. The Journal of Systems and Software, 108(??):148–177, October 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215001314.

Goncalves:2015:MMS

[GCC^{+}15] Enyo José Tavares Gonçalves, Mariela I. Cortés, Gustavo Augusto Lima Campos, Yrleyjander S.

Lopes, Emmanuel S. S. Freire, Viviane Torres da Silva, Kleinner Silva Farias de Oliveira, and Marcos Antonio de Oliveira. MAS-ML 2.0: Supporting the modelling of multi-agent systems with different agent architectures. The Journal of Systems and Software, 108(??):77–109, October 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001247.

Garousi:2016:CFA

[GCDY16] Vahid Garousi, Ahmet Coskunçay, Onur Demirörs, and Ali Yazici. Cross-factor analysis of software engineering practices versus practitioner demographics: an exploratory study in Turkey. The Journal of Systems and Software, 111(??): 49–73, January 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121215002071.

Gu:2013:AVS

[GCLD13] Qing Gu, Félix Cuadrado, Patricia Lago, and Juan C. Dueñas. 3D architecture viewpoints on service automation. The Journal of Systems and Software, 86(5):1307–1322, May 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212003445.

Grieco:2017:QTF

[GCMB17] Gustavo Grieco, Martín Ceresa, Agustín Mista, and Pablo Buiras. Quick-Fuzz testing for fun and profit. The Journal of Systems and Software, 134(??):340–354, December 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217302066.

Guerra-Casanova:2011:SOT

[GCSÁddP11] J. Guerra-Casanova, C. Sánchez-Ávila, A. de Santos Sierra, and G. Bailador del Pozo. Score optimization and template updating in a biometric technique for authentication in mobiles based on gestures. The Journal of Systems and Software, 84(11):2013–2021, November 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001427.

Gonzalez-Compean:2018:SBB

[GCSSDP^{+}18] J. L. Gonzalez-Compean, Victor Sosa-Sosa, Arturo Diaz-Perez, Jesus Carretero, and Jedidiah Yanez-Sierra. Sacbe: a building block approach for constructing efficient and flexible end-to-end cloud storage. The Journal of Systems and Software, 135(??):143–156, January 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302212.

Ghobadi:2012:CRC

[GD12] Shahla Ghobadi and John D’Ambra. Coopetitive relationships in cross-functional software development teams: How to model and measure? The Journal of Systems and Software, 85(5):1096–1104, May 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211003190.

Garcia-Diaz:2010:TMM

[GDFFPG^{+}10] Vicente García-Díaz, Héctor Fernández-Fernández, Elías Palacios-González, B. Cristina Pelayo G-Bustelo, Oscar Sanjuán-Martínez, and Juan Manuel Cueva Lovelle. TALISMAN MDE: Mixing MDE principles. The Journal of Systems and Software, 83(7):1179– 1191, July 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gholami:2016:CMP

[GDLB16] Mahdi Fahmideh Gholami, Farhad Daneshgar, Graham Low, and Ghassan Beydoun. Cloud migration process — a survey, evaluation framework, and open challenges. The Journal of Systems and Software, 120(??):31–69, October 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216300966.

Gui:2011:TAB

[GDSB11] Ning Gui, Vincenzo De Florio, Hong Sun, and Chris Blondia. Toward architecture-based context-aware deployment and adaptation. The Journal of Systems and Software, 84(2):185– 197, February 2011. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Galante:2015:PLA

[GE15a] Guilherme Galante and Luis Carlos Erpen De Bona. A programminglevel approach for elasticizing parallel scientific applications. The Journal of Systems and Software, 110(??):239–252, December 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001946.

Ghabi:2015:ETU

[GE15b] Achraf Ghabi and Alexander Egyed. Exploiting traceability uncertainty among artifacts and code. The Journal of Systems and Software, 108(??):178–192, October 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001338.

Gutierrez:2015:MDA

[GEM15] J. J. Gutiérrez, M. J.

Escalona, and M. Mejías. A model-driven approach for functional test case generation. The Journal of Systems and Software, 109(??):214–228, November 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001703.

Ghini:2011:ABP

[GFP11] Vittorio Ghini, Stefano Ferretti, and Fabio Panzieri. The “Always Best Packet Switching” architecture for SIPbased mobile multimedia services. The Journal of Systems and Software, 84(11):1827–1851, November 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001506.

Graziotin:2018:WHW

[GFWA18] Daniel Graziotin, Fabian Fagerholm, Xiaofeng Wang, and Pekka Abrahamsson. What happens when software developers are (un)happy. The Journal of Systems and Software, 140(??):32–47, June 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300323.

Gill:2019:RFE

[GGB19] Sukhpal Singh Gill, Peter Garraghan, and Rajkumar Buyya. ROUTER: Fog enabled cloud based intelligent resource management approach for smart home IoT devices. The Journal of Systems and Software, 154(??):125–138, August 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300986.

Gefen:2016:HPD

[GGC16] David Gefen, Gavriel Gefen, and Erran Carmel. How project description length and expected duration affect bidding and project success in crowdsourcing software development. The Journal of Systems and Software, 116(??):75–84, June 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121500062X.

Gerangelos:2019:EAS

[GGK19] Stefanos Gerangelos, Georgios Goumas, and Nectarios Koziris. Efficient accelerator sharing in virtualized environments: a Xeon Phi use-case. The Journal of Systems and Software, 150(??):37–50, April 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218302863.

Gutierrez:2011:RBP

[GGM11] Celia Gutierrez and Iván García-Magariño. Revealing bullying patterns in multi-agent systems. The Journal of Systems and Software, 84 (9):1563–1575, September 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211000720.

Gutierrez-Garcia:2015:ABC

[GGS15] J. Octavio Gutierrez-Garcia and Kwang Mong Sim. Agent-based cloud bag-of-tasks execution. The Journal of Systems and Software, 104(??): 17–31, June 2015. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S016412121500045X.

Gill:2019:HRM

[GGS^{+}19] Sukhpal Singh Gill, Peter Garraghan, Vlado Stankovski, Giuliano Casale, Ruppa K. Thulasiram, Soumya K. Ghosh, Kotagiri Ramamohanarao, and Rajkumar Buyya. Holistic resource management for sustainable and reliable cloud computing: an innovative solution to global challenge. The Journal of Systems and Software, 155 (??):104–129, September 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301098.

Garousi:2019:ASE

[GGT^{+}19] Vahid Garousi, Görkem Giray, Eray Tüzün, Cagatay Catal, and Michael Felderer. Aligning software engineering education with industrial needs: a metaanalysis. The Journal of Systems and Software, 156(??):65–83, October 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301347.

Getir:2018:SSA

[GGvH^{+}18] Sinem Getir, Lars Grunske, André van Hoorn, Timo Kehrer, Yannic Noller, and Matthias Tichy. Supporting semi-automatic co-evolution of architecture and fault tree models. The Journal of Systems and Software, 142(??):115–135, August 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218300657.

Gonzalez-Herrera:2016:SSA

[GHBD^{+}16] I. Gonzalez-Herrera, J. Bourcier, E. Daubert, W. Rudametkin, O. Barais, F. Fouquet, J. M. Jézéquel, and B. Baudry. Scape-Goat: Spotting abnormal resource usage in component-based reconfigurable software systems. The Journal of Systems and Software, 122(??):398–415, December 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000595.

Gao:2013:LCA

[GJ13] Guangyong Gao and Guoping Jiang. A lossless copyright authentication scheme based on Bessel–Fourier moment and extreme learning machine in curvaturefeature domain. The Journal of Systems and Software, 86(1):222–232, January 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002270.

Gasparic:2016:WRS

[GJ16] Marko Gasparic and Andrea Janes. What recommendation systems for software engineering recommend: a systematic literature review. The Journal of Systems and Software, 113(??):101–113, March 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002605.

Garousi:2018:SST

[GK18] Vahid Garousi and Baris Küçük. Smells in software test code: a survey of knowledge in industry and academia. The Journal of Systems and Software, 138(??):52–81, April 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217303060.

Gallo:2013:FFD

[GKD13] Roberto Gallo, Henrique Kawakami, and Ricardo Dahab. FORTUNA — a framework for the design and development of hardware-based secure systems. The Journal of Systems and Software, 86(8):2063–2076, August 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121300071X.

Gebizli:2018:ITE

[GKS18] Ceren Sahin Gebizli, Abdulhadi Kirkici, and Hasan Sözer. Increasing test efficiency by riskdriven model-based testing. The Journal of Systems and Software, 144(??):356–365, October 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301328.

Goknil:2014:GVT

[GKV14] Arda Goknil, Ivan Kurtev, and Klaas Van Den Berg. Generation and validation of traces between requirements and architecture based on formal trace semantics. The Journal of Systems and Software, 88 (??):112–137, February 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002410.

Gamalielsson:2014:SOS

[GL14] Jonas Gamalielsson and Björn Lundell. Sustainability of Open Source software communities beyond a fork: How and why has the LibreOffice project evolved? The Journal of Systems and Software, 89(??):128– 145, March 2014. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121213002744.

Guerrero:2013:PIW

[GLJ13] Carlos Guerrero, Isaac Lera, and Carlos Juiz. Performance improvement of web caching in Web 2.0 via knowledge discovery. The Journal of Systems and Software, 86(12):2970–2980, December 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213001209.

Guo:2019:EBF

[GLOM19] Yun Guo, Nan Li, Jeff Offutt, and Amihai Motro. Exonerationbased fault localization for SQL predicates. The Journal of Systems and Software, 147(??):230–245, January 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218302358.

Guo:2013:TVS

[GLW13] Teng Guo, Feng Liu, and ChuanKun Wu. Threshold visual secret sharing by random grids with improved contrast. The Journal of Systems and Software, 86 (8):2094–2109, August 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000745.

Gao:2010:EEQ

[GLWY10] Jun Gao, Jiaheng Lu, Tengjiao Wang, and Dongqing Yang. Efficient evaluation of query rewriting plan over materialized XML view. The Journal of Systems and Software, 83(6):1029– 1038, June 2010. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gaubatz:2015:AEC

[GLZ15] Patrick Gaubatz, Ioanna Lytra, and Uwe Zdun. Automatic enforcement of constraints in realtime collaborative architectural decision making. The Journal of Systems and Software, 103(??):128–149, May 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000345.

Gonzalez-Manzano:2014:EUS

[GMGTdFR14] Lorena González-Manzano, Ana I. González-Tablas, José M. de Fuentes, and Arturo Ribagorda. Extended U+F Social Network Protocol: Interoperability, reusability, data protection and indirect relationships in Web Based Social Networks. The Journal of Systems and Software, 94(??):50–71, August 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001083.

Gomez-Martinez:2015:SAD

[GMLSF^{+}15] Elena Gómez-Martínez, Marino Linaje, Fernando Sánchez-Figueroa, Andrés Iglesias-Pérez, Juan Carlos Preciado, Rafael González-Cabero, and José Merseguer. A semantic approach for designing Assistive Software Recommender systems. The Journal of Systems and Software, 104(??):166–178, June 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000606.

Guillen:2013:SOF

[GMMC13] Joaquín Guillén, Javier Miranda, Juan Manuel Murillo, and Carlos Canal. A service-oriented framework for developing cross cloud migratable software. The Journal of Systems and Software, 86 (9):2294–2308, September 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212003421.

Garcia-Mireles:2015:APP

[GMMGP15] Gabriel Alberto García-Mireles, Ma. Ángeles Moraga, Félix García, and Mario Piattini. Approaches to promote product quality within software process improvement initiatives: a mapping study. The Journal of Systems and Software, 103(??):150– 166, May 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000369.

Garcia-Magarino:2016:MDA

[GMPN16] Iván García-Magariño and Guillermo Palacios-Navarro. A model-driven approach for constructing ambient assistedliving multi-agent systems customized for Parkinson patients. The Journal of Systems and Software, 111(??):34–48, January 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215002083.

Gasparic:2017:CMI

[GMR17] Marko Gasparic, Gail C. Murphy, and Francesco Ricci. A context model for IDE-based recommendation systems. The Journal of Systems and Software, 128(??):200– 219, June 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216301807.

Guo:2011:ISS

[GMS11] Fuchun Guo, Yi Mu, and Willy Susilo. Improving security of q-SDH based digital signatures. The Journal of Systems and Software, 84(10):1783–1790, October 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211001233.

Gandomani:2015:EDF

[GN15] Taghi Javdani Gandomani and Mina Ziaei Nafchi. An empiricallydeveloped framework for agile transition and adoption: a grounded theory approach. The Journal of Systems and Software, 107(??):204– 219, September 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121215001223.

Genc-Nayebi:2017:SLR

[GNA17] Necmiye Genc-Nayebi and Alain Abran. A systematic literature review: Opinion mining studies from mobile app store user reviews. The Journal of Systems and Software, 125(??):207– 219, March 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

//www.sciencedirect.
com/science/article/
pii/S0164121216302291.

Geppert:2010:EJS

[GP10a] Birgit Geppert and Klaus Pohl. Editorial for the JSS SPLC 2008 Special Issue. The Journal of Systems and Software, 83 (7):1077, July 2010. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gotlieb:2010:URT

[GP10b] Arnaud Gotlieb and Matthieu Petit. A uniform random test data generator for path testing. The Journal of Systems and Software, 83 (12):2618–2626, December 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Grano:2019:SSB

[GPD^{+}19] Giovanni Grano, Fabio Palomba, Dario Di Nucci, Andrea De Lucia, and Harald C. Gall. Scented since the beginning: On the diffuseness of test smells in automatically generated test code. The Journal of Systems and Software, 156 (??):312–327, October 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301487.

Guo:2015:EEA

[GPL^{+}15] Kehua Guo, Wei Pan, Mingming Lu, Xiaoke Zhou, and Jianhua Ma. An effective and economical architecture for semantic-based heterogeneous multimedia big data retrieval. The Journal of Systems and Software, 102(??):207–216, April 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002040.

Gencel:2013:DSF

[GPMI13] Cigdem Gencel, Kai Petersen, Aftab Ahmad Mughal, and Muhammad Imran Iqbal. A decision support framework for metrics selection in goal-based measurement programs: GQM-DSFMS. The Journal of Systems and Software, 86 (12):3091–3108, December 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001726.

Garcia:2017:FGS

[GPP^{+}17] Félix García, Oscar Pedreira, Mario Piattini, Ana Cerdeira-Pena, and Miguel Penabad. A framework for gamification in software engineering. The Journal of Systems and Software, 132(??):21–40, October 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217301218.

Guan:2016:OSF

[GPPT16] Fei Guan, Long Peng, Luc Perneel, and Martin Timmerman. Open source FreeRTOS as a case study in real-time operating system evolution. The Journal of Systems and Software, 118(??):19–35, August 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216300383.

Gonzalez:2013:ACP

[GPSS^{+}13] J. L. Gonzalez, Jesus Carretero Perez, Victor Sosa-Sosa, Juan F. Rodriguez Cardoso, and Ricardo Marcelin-Jimenez. An approach for constructing private storage services as a unified fault-tolerant system. The Journal of Systems and Software, 86(7):1907–1922, July 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S016412121300054X.

Galizia:2012:JAS

[GQ12] Antonella Galizia and Alfonso Quarati. Job allocation strategies for energy-aware and efficient Grid infrastructures. The Journal of Systems and Software, 85(7):1588–1606, July 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000362.

Garrigues:2010:PDS

[GRBNA10] Carles Garrigues, Sergi Robles, Joan Borrell, and Guillermo Navarro-Arribas. Promoting the development of secure mobile agent applications. The Journal of Systems and Software, 83 (6):959–971, June 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Garcia:2016:DRP

[GRR16] Stephany García, Oscar Romero, and Ruth Raventós. DSS from an RE Perspective: a systematic mapping. The Journal of Systems and Software, 117(??):488– 507, July 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216300061.

Golfarelli:2013:MSP

[GRT13] Matteo Golfarelli, Stefano Rizzi, and Elisa Turricchia. Multi-sprint planning and smooth replanning: an optimization model. The Journal of Systems and Software, 86(9):2357–2370, September 2013. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213001039.

Ghazouani:2017:TSC

[GS17] Souad Ghazouani and Yahya Slimani. Towards a standardized cloud service description based on USDL. The Journal of Systems and Software, 132(??):1–20, October 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217301322.

Guo:2016:COE

[GSdS16] Yuepu Guo, Carolyn Seaman, and Fabio Q. B. da Silva. Costs and obstacles encountered in technical debt management — a case study. The Journal of Systems and Software, 120 (??):156–169, October 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121630108X.

Ghanbari:2015:UOS

[GSM15] Hadi Ghanbari, Jouni Similä, and Jouni Markkula. Utilizing online serious games to facilitate distributed requirements elicitation. The Journal of Systems and Software, 109(??):32–49, November 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001491.

Gacitua:2019:FCF

[GSM19] Ricardo Gacitúa, Samuel Sepúlveda, and Raúl Mazo. FM-CF: a framework for classifying feature model building approaches. The Journal of Systems and Software, 154(??):1–21, August 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219300767.

Gerolymatos:2015:SNF

[GSN^{+}15] Panagiotis Gerolymatos, Spyros Sioutas, Nikolaos Nodarakis, Alexandros Panaretos, and Konstantinos Tsakalidis. SMaRT: a novel framework for addressing range queries over nonlinear trajectories. The Journal of Systems and Software, 105(??):79–90, July 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215000709.

Gerostathopoulos:2019:TSA

[GSP^{+}19] Ilias Gerostathopoulos, Dominik Skoda, Frantisek Plasil, Tomas Bures, and Alessia Knauss. Tuning self-adaptation in cyber-physical systems through architectural homeostasis. The Journal of Systems and Software, 148(??):37–55, February 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S016412121830236X.

Gorschek:2014:USD

[GTA14] Tony Gorschek, Ewan Tempero, and Lefteris Angelis. On the use of software design models in software development practice: an empirical investigation. The Journal of Systems and Software, 95(??):176–193, September 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001022.

Gren:2015:PQM

[GTF15] Lucas Gren, Richard Torkar, and Robert Feldt. The prospects of a quantitative measurement of agility: a validation study on an agile maturity model. The Journal of Systems and Software, 107 (??):38–49, September 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001028.

Gren:2017:GDG

[GTF17] Lucas Gren, Richard Torkar, and Robert Feldt. Group development and group maturity when building agile teams: a qualitative and quantitative investigation at eight large companies. The Journal of Systems and Software, 124(??):104–119, February 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216302266.

Gong:2012:GTP

[GTY12] Dunwei Gong, Tian Tian, and Xiangjuan Yao. Grouping target paths for evolutionary generation of test data in parallel. The Journal of Systems and Software, 85(11):2531–2540, November 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001562.

Garousi:2010:RSS

[GV10] Vahid Garousi and Tan Varma. A replicated survey of software testing practices in the Canadian province of Alberta: What has changed from 2004 to 2009? The Journal of Systems and Software, 83(11):2251–2262, November 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Garcia-Valls:2018:PCP

[GVPPM18]  Marisol  García-Valls,

Diego Perez-Palacin, and Raffaela Mirandola. Pragmatic cyber physical systems design based on parametric models. The Journal of Systems and Software, 144 (??):559–572, October 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301298.

Gwebu:2010:SEE

[GW10] Kholekile L. Gwebu and Jing Wang. Seeing eye to eye? An exploratory study of free open source software users’ perceptions. The Journal of Systems and Software, 83 (11):2287–2296, November 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Guo:2011:GAO

[GWW^{+}11] Jianmei Guo, Jules White, Guangxin Wang, Jian Li, and Yinglin Wang. A genetic algorithm for optimized feature selection with resource constraints in software product lines. The Journal of Systems and Software, 84 (12):2208–2221, December 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001518.

Gu:2019:DFR

[GXZ^{+}19] Yongfeng Gu, Jifeng Xuan, Hongyu Zhang, Lanxin Zhang, Qingna Fan, Xiaoyuan Xie, and Tieyun Qian. Does the fault reside in a stack trace? Assisting crash localization by predicting crashing fault residence. The Journal of Systems and Software, 148(??):88–104, February 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302401.

Garousi:2013:SST

[GZ13] Vahid Garousi and Junji Zhi. A survey of software testing practices in Canada. The Journal of Systems and Software, 86(5):1354–1376, May 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212003561.

Gunasekera:2013:BUC

[GZKL13] Kutila Gunasekera, Arkady Zaslavsky, Shonali Krishnaswamy, and Seng Wai Loke. Building ubiquitous computing applications using the VERSAG adaptive agent framework. The Journal of Systems and Software, 86(2):501–519, February 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002695.

Guo:2018:SMK

[GZS^{+}18] Ziqing Guo, Hua Zhang, Caijun Sun, Qiaoyan Wen, and Wenmin Li. Secure multi-keyword ranked search over encrypted cloud data for multiple data owners. The Journal of Systems and Software, 137(??): 380–395, March 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121217303011.

Gong:2011:EGT

[GZY11] Dunwei Gong, Wanqiu Zhang, and Xiangjuan Yao. Evolutionary generation of test data for many paths coverage based on grouping. The Journal of Systems and Software, 84 (12):2222–2233, December 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121100152X.

Harrison:2010:HDA

[HA10] Neil B. Harrison and Paris Avgeriou. How do architecture patterns and tactics interact? A model and annotation. The Journal of Systems and Software, 83(10):1735– 1758, October 2010. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hora:2015:ADS

[HAE^{+}15] André Hora, Nicolas Anquetil, Anne Etien, Stéphane Ducasse, and Marco Túlio Valente. Automatic detection of system-specific conventions unknown to developers. The Journal of Systems and Software, 109(??):192–204, November 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001727.

Hanssen:2012:LCS

[Han12] Geir K. Hanssen. A longitudinal case study of an emerging software ecosystem: Implications for practice and theory. The Journal of Systems and Software, 85(7):1455–1466, July 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211000963.

Houmansadr:2013:BCN

[HB13] Amir Houmansadr and Nikita Borisov. Bot-Mosaic: Collaborative network watermark for the detection of IRCbased botnets. The Journal of Systems and Software, 86(3):707–715, March 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212003068.

Hakiri:2013:SEE

[HBG^{+}13] Akram Hakiri, Pascal Berthou, Aniruddha Gokhale, Douglas C. Schmidt, and Thierry Gayraud. Supporting end-to-end quality of service properties in OMG data distribution service publish/subscribe middleware over wide area networks. The Journal of Systems and Software, 86(10):2574–2593, October 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213001210.

Hakiri:2014:SSB

[HBG^{+}14] Akram Hakiri, Pascal Berthou, Aniruddha Gokhale, Douglas C. Schmidt, and Gayraud Thierry. Supporting SIPbased end-to-end Data Distribution Service QoS in WANs. The Journal of Systems and Software, 95(??):100–121, September 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214000983.

Hammad:2019:DAA

[HBM19] Mahmoud Hammad, Hamid Bagheri, and Sam Malek. DelDroid: an automated approach for determination and enforcement of least-privilege architecture in Android. The Journal of Systems and Software, 149(??):83– 100, March 2019. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121218302589.

Hurtado:2013:MSP

[HBOS13] Julio Ariel Hurtado, María Cecilia Bastarrica, Sergio F. Ochoa, and Jocelyn Simmonds. MDE software process lines in small companies. The Journal of Systems and Software, 86(5):1153– 1171, May 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002749.

Haghighatkhah:2017:ASE

[HBP^{+}17] Alireza Haghighatkhah, Ahmad Banijamali, Olli-Pekka Pakanen, Markku Oivo, and Pasi Kuvaja. Automotive software engineering: a systematic mapping study. The Journal of Systems and Software, 128(??):25–55, June 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217300560.

Harutyunyan:2019:IRF

[HBR19] Nikolay Harutyunyan, Andreas Bauer, and Dirk Riehle. Industry requirements for FLOSS governance tools to facilitate the use of open source software in commercial products. The Journal of Systems and Software, 158(??):??, December 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301578.

Hanh:2016:NFF

[HBT16] Le Thi My Hanh, Nguyen Thanh Binh, and Khuat Thanh Tung. A novel fitness function of metaheuristic algorithms for test data generation for Simulink models based on mutation analysis. The Journal of Systems and Software, 120(??):17–30, October 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0164121216301017.

Hong:2010:LVC

[HC10] Wien Hong and Tung-Shou Chen. A local variance-controlled reversible data hiding method using prediction and histogram-shifting. The Journal of Systems and Software, 83 (12):2653–2663, December 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Harman:2015:SBS

[HC15] Mark Harman and Francisco Chicano. Search Based Software Engineering (SBSE). The Journal of Systems and Software, 103(??):266, May 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121500031X.

Han:2016:GSL

[HCB^{+}16] Wookhyun Han, Hoon Sung Chwa, Hwidong Bae, Hyosu Kim, and Insik Shin. GPU-SAM: Leveraging multi-GPU split-and-merge execution for system-wide realtime support. The Journal of Systems and Software, 117(??):1–14, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216000455.

Huang:2010:DAG

[HCC10a] Chin-Yu Huang, Jun-Ru Chang, and Yung-Hsin Chang. Design and analysis of GUI test-case prioritization using weightbased methods. The Journal of Systems and Software, 83(4):646–659, April 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hwang:2010:RIB

[HCC10b] Min-Shiang Hwang, Song-Kong Chong, and Te-Yu Chen. DoS-resistant ID-based password authentication scheme using smart cards. The Journal of Systems and Software, 83(1):163–172, January 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:2010:DRA

[hChSyCwL10] Lin huang Chang, Chun hui Sung, Shih yi Chiu, and Yen wen Lin. Design and realization of ad-hoc VoIP with embedded p-SIP server. The Journal of Systems and Software, 83(12):2536–2555, December 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hristidis:2010:SDM

[HCL^{+}10] Vagelis Hristidis, Shu-Ching Chen, Tao Li, Steven Luis, and Yi Deng. Survey of data management and analysis in disaster situations. The Journal of Systems and Software, 83(10):1701– 1714, October 2010. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hong:2012:DEU

[HCL12] Wien Hong, Tung-Shou Chen, and Chih-Wei Luo. Data embedding using pixel value differencing and diamond encoding with multiple-base notational system. The Journal of Systems and Software, 85(5):1166–1175, May 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211003268.

Huang:2015:ASI

[HCT^{+}15] Rubing Huang, Jinfu Chen, Dave Towey, Alvin T. S. Chan, and Yansheng Lu. Aggregatestrength interaction test suite prioritization. The Journal of Systems and Software, 99(??):36–51, January 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001940.

Han:2019:RPB

[HCY19] Xue Han, Daniel Carroll, and Tingting Yu. Reproducing performance bug reports in server applications: the researchers’ experiences. The Journal of Systems and Software, 156(??):268–282, October 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301438.

Hanazumi:2017:FAI

[HdM17] Simone Hanazumi and Ana C. V. de Melo. A formal approach to implement Java exceptions in cooperative systems. The Journal of Systems and Software, 131 (??):475–490, September 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301285.

Houmb:2010:QSR

[HFE10] Siv Hilde Houmb, Virginia N. L. Franqueira, and Erlend A. Engum. Quantifying security risk level from CVSS estimates of frequency and impact. The Journal of Systems and Software, 83 (9):1622–1634, September 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hoorn:2011:LA

[HFLvV11] Johan F. Hoorn, Rik Farenhorst, Patricia Lago, and Hans van Vliet. The lonesome architect. The Journal of Systems and Software, 84 (9):1424–1435, September 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121210003171.

Hucka:2018:SSS

[HG18] M. Hucka and M. J. Graham. Software search is not a science, even among scientists: A survey of how scientists and engineers find software. The Journal of Systems and Software, 141(??):171–191, July 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300517.

Hajri:2018:CIA

[HGBS18] Ines Hajri, Arda Goknil, Lionel C. Briand, and Thierry Stephany. Change impact analysis for evolving configuration decisions in product line use case models. The Journal of Systems and Software, 139(??):211–237, May 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300293.

Hoyos:2013:DSL

[HGMB13] José R. Hoyos, Jesús García-Molina, and Juan A. Botía. A domainspecific language for context modeling in contextaware systems. The Journal of Systems and Software, 86(11):2890–2905, November 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0164121213001696.

Hallsteinsen:2012:DFM

[HGP^{+}12] S. Hallsteinsen, K. Geihs, N. Paspallis, F. Eliassen, G. Horn, J. Lorenzo, A. Mamelli, and G. A. Papadopoulos. A development framework and methodology for selfadapting applications in ubiquitous computing environments. The Journal of Systems and Software, 85(12):2840–2859, December 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002245.

Huh:2017:PFS

[HH17] Sungju Huh and Seongsoo Hong. Providing fairshare scheduling on multicore computing systems via progress balancing. The Journal of Systems and Software, 125(??): 183–196, March 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

//www.sciencedirect.
com/science/article/
pii/S0164121216302412.

Huang:2012:CAM

[HHC12] Der-Chen Huang, Kun-Ding Hung, and Yung-Kuan Chan. A computer assisted method for leukocyte nucleus segmentation and recognition in blood smear images. The Journal of Systems and Software, 85 (9):2104–2118, September 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001045.

Hsu:2010:HSA

[HHH^{+}10a] Fu-Hau Hsu, Cheng-Hsien Huang, Chi-Hsien Hsu, Chih-Wen Ou, Li-Han Chen, and Ping-Cheng Chiu. HSP: a solution against heap sprays. The Journal of Systems and Software, 83 (11):2227–2236, November 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Huang:2010:PSS

[HHH10b] Chin-Pan Huang, Chaur-Heh Hsieh, and Ping Sheng Huang. Progressive sharing for a secret image. The Journal of Systems and Software, 83(3):517– 527, March 2010. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hu:2013:KDW

[HHK13] Ya-Han Hu, Tony Cheng-Kui Huang, and Yu-Hua Kao. Knowledge discovery of weighted RFM sequential patterns from customer sequence databases. The Journal of Systems and Software, 86(3):779–788, March 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121200310X.

Hafiz:2016:GLE

[HHKWB16] Munawar Hafiz, Samir Hasan, Zachary King, and Allen Wirfs-Brock. Growing a language: an empirical study on how (and why) developers use some recently-introduced and/or recently-evolving JavaScript features. The Journal of Systems and Software, 121(??):191– 208, November 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121216300309.

Hasheminejad:2012:DPS

[HJ12] Seyed Mohammad Hossein Hasheminejad and Saeed Jalili. Design patterns selection: an automatic two-phase method. The Journal of Systems and Software, 85 (2):408–424, February 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002317.

Hasheminejad:2014:EAI

[HJ14] Seyed Mohammad Hossein Hasheminejad and Saeed Jalili. An evolutionary approach to identify logical components. The Journal of Systems and Software, 96(??):24–50, October 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001228.

Han:2010:MBD

[HJBH10] Ah-Rim Han, Sang-Uk Jeon, Doo-Hwan Bae, and Jang-Eui Hong. Measuring behavioral dependency for improving change-proneness prediction in UML-based design models. The Journal of Systems and Software, 83(2):222–234, February 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hansen:2011:ESS

[HJN11] Klaus Marius Hansen, Kristjan Jonasson, and Helmut Neukirchen. An empirical study of software architectures’ effect on product quality. The Journal of Systems and Software, 84(7): 1233–1243, July 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Holtkamp:2015:SCR

[HJP15] Philipp Holtkamp, Jussi P. P. Jokinen, and Jan M. Pawlowski. Soft competency requirements in requirements engineering, software design, implementation, and testing. The Journal of Systems and Software, 101(??):136–146, March 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002829.

Hong:2013:EPD

[HK13] Shin Hong and Moonzoo Kim. Effective patterndriven concurrency bug detection for operating systems. The Journal of Systems and Software, 86(2):377–388, February 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212002671.

Huang:2017:CVB

[HKS^{+}17] Jianglin Huang, Jacky Wai Keung, Federica Sarro, Yan-Fu Li, Y. T. Yu, W. K. Chan, and Hongyi Sun. Cross-validation based K nearest neighbor imputation for software quality datasets: an empirical study. The Journal of Systems and Software, 132(??):226–252, October 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217301516.

Hazzan:2010:DFS

[HL10] Orit Hazzan and Uri Leron. Disciplined and free-spirited: ‘Time-out behaviour’ at the Agile conference. The Journal of Systems and Software, 83(11):2363–2365, November 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Huang:2011:EKM

[HL11] Hui-Feng Huang and Kuo-Ching Liu. Efficient key management for preserving HIPAA regulations. The Journal of Systems and Software, 84(1):113–119, January 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Huang:2013:NAE

[HLLS13]  Chien-Cheng  Huang,

Feng-Yu Lin, Frank Yeong-Sung Lin, and Yeali S. Sun. A novel approach to evaluate software vulnerability prioritization. The Journal of Systems and Software, 86(11):2822–2840, November 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001489.

Harman:2013:CES

[HLS^{+}13] Mark Harman, Kiran Lakhotia, Jeremy Singer, David R. White, and Shin Yoo. Cloud engineering is Search Based Software Engineering too. The Journal of Systems and Software, 86 (9):2225–2241, September 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002853.

Hua:2015:PRI

[HLW^{+}15] Xiayu Hua, Zheng Li, Hao Wu, Chunhui Guo, and Shangping Ren. Periodic resource integration. The Journal of Systems and Software, 110(??):193–204, December 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215001934.

Huang:2013:CVS

[HLWS13] Chung-Yuan Huang, Chun-Liang Lee, Tzai-Hung Wen, and Chuen-Tsai Sun. A computer virus spreading model based on resource limitations and interaction costs. The Journal of Systems and Software, 86(3):801–808, March 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212003202.

Hoda:2016:MLA

[HM16] Rashina Hoda and Latha K. Murugesan. Multi-level agile project management challenges: a selforganizing team perspective. The Journal of Systems and Software, 117(??):245–257, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000807.

Haghighatkhah:2018:TPC

[HMOK18] Alireza Haghighatkhah, Mika Mäntylä, Markku Oivo, and Pasi Kuvaja. Test prioritization in continuous integration environments. The Journal of Systems and Software, 146(??):80–98, December 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301730.

Hierons:2017:IRP

[HN17] Robert M. Hierons and Manuel Núñez. Implementation relations and probabilistic schedulers in the distributed test architecture. The Journal of Systems and Software, 132(??):319–335, October 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217300602.

Hemmati:2015:IED

[HNH15] Hadi Hemmati, Meiyappan Nagappan, and Ahmed E. Hassan. Investigating the effect of “defect co-fix” on quality assurance resource allocation: a search-based approach. The Journal of Systems and Software, 103(??):412–422, May 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214002714.

Hjertstrom:2012:DMC

[HNS12] Andreas Hjertström, Dag Nyström, and Mikael Sjödin. Data management for componentbased embedded realtime systems: The database proxy approach. The Journal of Systems and Software, 85(4):821– 834, April 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211002792.

Haitzer:2017:RSA

[HNZ17] Thomas Haitzer, Elena Navarro, and Uwe Zdun. Reconciling software architecture and source code in support of software evolution. The Journal of Systems and Software, 123(??):119– 144, January 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121216302114.

Hoorn:2014:SLI

[Hoo14] Johan F. Hoorn. Stakeholder logistics of an interactive system. The Journal of Systems and Software, 95(??):52–69, September 2014. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S016412121300277X.

Hamid:2016:SPB

[HP16] Brahim Hamid and Jon Perez. Supporting pattern-based dependability engineering via model-driven development: Approach, toolsupport and empirical validation. The Journal of Systems and Software, 122(??):239–273, December 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121216301868.

Horcas:2016:APW

[HPF16] Jose-Miguel Horcas, Mónica Pinto, and Lidia Fuentes. An automatic process for weaving functional quality attributes using a software product line approach. The Journal of Systems and Software, 112(??):78–95, February 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121500240X.

Huang:2012:HBC

[HPH12] Yu-Chi Huang, Kuan-Li Peng, and Chin-Yu Huang. A history-based cost-cognizant test case prioritization technique in regression testing. The Journal of Systems and Software, 85(3):626–637, March 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211002780.

Hwang:2010:WCS

[HR10] Jeong Hee Hwang and Keun Ho Ryu. A weighted common structure based clustering technique for XML documents. The Journal of Systems and Software, 83(7):1267–1274, July 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hermassi:2012:SAI

[HRB12] Houcemeddine Hermassi, Rhouma Rhouma, and Safya Belghith. Security analysis of image cryptosystems only or partially based on a chaotic permutation. The Journal of Systems and Software, 85 (9):2133–2144, September 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001070.

Hassine:2010:ETS

[HRD10] Jameleddine Hassine, Juergen Rilling, and Rachida Dssouli. An evaluation of timed scenario notations. The Journal of Systems and Software, 83(2):326–350, February 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hernandez:2016:CCL

[HRRC16] Inma Hernández, Carlos R. Rivero, David Ruiz, and Rafael Corchuelo. CALA: ClAssifying Links Automatically based on their URL. The Journal of Systems and Software, 115(??):130–143, May 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121600042X.

Henderson-Sellers:2011:BMO

[HS11a] B. Henderson-Sellers. Bridging metamodels and ontologies in software engineering. The Journal of Systems and Software, 84(2):301–313, February 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hwang:2011:CDA

[HS11b] Shin-Jia Hwang and Yun-Hao Sung. Confidential deniable authentication using promised signcryption. The Journal of Systems and Software, 84(10):1652–1659, October 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211000999.

Huang:2015:SDS

[HS15] Kuo-Chan Huang and Bo-Jun Shen. Service deployment strategies for efficient execution of composite SaaS applications on cloud platform. The Journal of Systems and Software, 107 (??):127–141, September 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001156.

Hamrouni:2015:DMC

[HSC15] Tarek Hamrouni, Sarra Slimani, and Faouzi Ben Charrada. A data mining correlated patternsbased periodic decentralized replication strategy for data grids. The Journal of Systems and Software, 110(??):10–27, December 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001806.

Hong:2014:RFR

[HSL14] Jung-Hong Hong, Zeal Li-Tse Su, and Eric Hsueh-Chan Lu. A recommendation framework for remote sensing images by spatial relation analysis. The Journal of Systems and Software, 90(??):151–166, April 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121400003X.

Hyrynsalmi:2016:IDM

[HSM16] Sami Hyrynsalmi, Arho Suominen, and Matti Mäntymäki. The influence of developer multihoming on competition between software ecosystems. The Journal of Systems and Software, 111(??):119–127, January 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002010.

Hens:2014:PFD

[HSPD14] Pieter Hens, Monique Snoeck, Geert Poels, and Manu De Backer. Process fragmentation, distribution and execution using an event-based interaction scheme. The Journal of Systems and Software, 89(??):170–192, March 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002847.

Huang:2010:MUM

[HSS10] Leijun Huang, Sanjeev Setia, and Robert Simon. Mctorrent: Using multiple communication channels for efficient bulk data dissemination in wireless sensor networks. The Journal of Systems and Software, 83(1):108–120, January 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hyrynsalmi:2014:SVA

[HSS14] Sami Hyrynsalmi, Marko Seppänen, and Arho Suominen. Sources of value in application ecosystems. The Journal of Systems and Software, 96(??):61–72, October 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001356.

He:2015:DDB

[HST15] Ping He, Hong Shen, and Hui Tian. Ondemand data broadcast with deadlines for avoiding conflicts in wireless networks. The Journal of Systems and Software, 103(??):118–127, May 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000163.

See corrigendum [HST16].

He:2016:CDD

[HST16] Ping He, Hong Shen, and Hui Tian. Corrigendum to ’On-demand data broadcast with deadlines for avoiding conflicts in wireless networks’ [The Journal of Systems and Software 103 (2015) 118–127]. The Journal of Systems and Software, 111(??):323, January 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000825.

See [HST15].

Hartmann:2012:CIS

[HTB12] Herman Hartmann, Tim Trew, and Jan Bosch. The changing industry structure of software development for consumer electronics and its consequences for software architectures. The Journal of Systems and Software, 85(1):178–192, January 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002081.

Huang:2013:RDH

[HTH13] Li-Chin Huang, Lin-Yu Tseng, and Min-Shiang Hwang. A reversible data hiding method by histogram shifting in high quality medical images. The Journal of Systems and Software, 86(3):716–727, March 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212003172.

Hastbacka:2011:MDD

[HVK11] David Hästbacka, Timo Vepsäläinen, and Seppo Kuikka. Model-driven development of industrial process control applications. The Journal of Systems and Software, 84(7):1100–1113, July 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hoffman:2010:TCS

[HWC^{+}10] Daniel Hoffman, Hong-Yi Wang, Mitch Chang, David Ly-Gagnon, Lewis Sobotkiewicz, and Paul Strooper. Two case studies in grammar-based test generation. The Journal of Systems and Software, 83(12):2369– 2378, December 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hussain:2015:HEC

[HWdS^{+}15] Aamir Hussain, Rao Wenbi, Aristides Lopes da Silva, Muhammad Nadher, and Muhammad Mudhish. Health and emergency-care platform for the elderly and disabled people in the Smart City. The Journal of Systems and Software, 110(??):253–263, December 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001867.

Hu:2011:MAQ

[HWHT11] Shu-Chiung Hu, You-Chiun Wang, Chiuan-Yu Huang, and Yu-Chee Tseng. Measuring air quality in city areas by vehicular wireless sensor networks. The Journal of Systems and Software, 84(11):2005–2012, November 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001555.

Hu:2013:ETB

[HWL13a] Ya-Han Hu, Fan Wu, and Yi-Jiun Liao. An efficient tree-based algorithm for mining sequential patterns with multiple minimum supports. The Journal of Systems and Software, 86(5):1224– 1238, May 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212003329.

Hu:2013:IVB

[HWL13b] Yongjian Hu, Kan Wang, and Zhe-Ming Lu. An improved VLC-based lossless data hiding scheme for JPEG images. The Journal of Systems and Software, 86(8):2166– 2173, August 2013. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121213000848.

Huang:2011:SBA

[HWLM11] Gang Huang, Weihu Wang, Tiancheng Liu, and Hong Mei. Simulationbased analysis of middleware service impact on system reliability: Experiment on Java application server. The Journal of Systems and Software, 84(7):1160–1170, July 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hidalgo:2017:SAP

[HWR17] Nicolas Hidalgo, Daniel Wladdimiro, and Erika Rosas. Self-adaptive processing graph with operator fission for elastic stream processing. The Journal of Systems and Software, 127(??):205– 216, May 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216300796.

Han:2011:BAG

[HY11] Lixin Han and Hong Yan. BSN: an automatic generation algorithm of social network data. The Journal of Systems and Software, 84(8): 1261–1269, August 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Holton:2011:PSR

[HYA11] D. R. W. Holton, M. Younas, and I. U. Awan. Priority scheduling of requests to web portals. The Journal of Systems and Software, 84(8):1373–1378, August 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Han:2012:ACS

[HyLW^{+}12] Song Han, Kam yiu Lam, Jiantao Wang, Sang H. Son, and Aloysius K. Mok. Adaptive co-scheduling for periodic application and update transactions in real-time database systems. The Journal of Systems and Software, 85(8):1729–1743, August 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000830.

Huang:2011:IBS

[HYWS11] Qiong Huang, Guomin Yang, Duncan S. Wong, and Willy Susilo. Identitybased strong designated verifier signature revisited. The Journal of Systems and Software, 84(1): 120–129, January 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Haitzer:2015:SAA

[HZ15] Thomas Haitzer and Uwe Zdun. Semiautomatic architectural pattern identification and documentation using architectural primitives. The Journal of Systems and Software, 102(??):35–57, April 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214002921.

He:2012:RHS

[HZG^{+}12] Chuan He, Xiaomin Zhu, Hui Guo, Dishan Qiu, and Jianqing Jiang. Rolling-horizon scheduling for energy constrained distributed realtime embedded systems. The Journal of Systems and Software, 85(4):780– 794, April 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211002615.

He:2016:MPT

[HZH^{+}16] Xiao He, Tian Zhang, Chang-Jun Hu, Zhiyi Ma, and Weizhong Shao. An MDE performance testing framework based on random model generation. The Journal of Systems and Software, 121(??):247–264, November 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300292.

Idri:2016:MDT

[IAA16] Ali Idri, Ibtissam Abnane, and Alain Abran. Missing data techniques in analogy-based software development effort estimation. The Journal of Systems and Software, 117(??):595–611, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300334.

Islam:2011:MES

[IB11] Sk. Hafizul Islam and G. P. Biswas. A more efficient and secure IDbased remote mutual authentication with key agreement scheme for mobile devices on elliptic curve cryptosystem. The Journal of Systems and Software, 84 (11):1892–1898, November 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211001646.

Ibrahim:2012:RBC

[IBAH12] Walid M. Ibrahim, Nicolas Bettenburg, Bram Adams, and Ahmed E. Hassan. On the relationship between comment update practices and Software Bugs. The Journal of Systems and Software, 85 (10):2293–2304, October 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121100238X.

Ilarri:2011:APC

[IBM11] Sergio Ilarri, Carlos Bobed, and Eduardo Mena. An approach to process continuous location-dependent queries on moving objects with support for location granules. The Journal of Systems and Software, 84(8):1327–1350, August 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Inam:2014:PIR

[ICSK14] Rafia Inam, Jan Carlson, Mikael Sjödin, and Jirí Kuncar. Predictable integration and reuse of executable real-time components. The Journal of Systems and Software, 91(??):147–162, May 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214000132.

Israeli:2010:LKC

[IF10] Ayelet Israeli and Dror G.

Feitelson. The Linux kernel as a case study in software evolution. The Journal of Systems and Software, 83(3):485–501, March 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Iacob:2019:EGB

[IF19] Claudia Iacob and Shamal Faily. Exploring the gap between the student expectations and the reality of teamwork in undergraduate software engineering group projects. The Journal of Systems and Software, 157(??):??, November 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301682.

Ifinedo:2011:EIE

[Ifi11] Princely Ifinedo. Examining the influences of external expertise and in-house computer/IT knowledge on ERP system success. The Journal of Systems and Software, 84(12):2065–2078, December 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211001208.

Idri:2016:SLR

[IHA16] Ali Idri, Mohamed Hosni, and Alain Abran. Systematic literature review of ensemble effort estimation. The Journal of Systems and Software, 118(??):151–175, August 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216300450.

Islam:2014:CCS

[IKBH14] Syed Islam, Jens Krinke, David Binkley, and Mark Harman. Coherent clusters in source code. The Journal of Systems and Software, 88(??):1–24, February 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121300188X.

Islam:2013:FQR

[ILZ13] Md. Saiful Islam, Chengfei Liu, and Rui Zhou. A framework for query refinement with user feedback. The Journal of Systems and Software, 86(6):1580–1595, June 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000265.

Islam:2014:FFI

[ILZ14] Md. Saiful Islam, Chengfei Liu, and Rui Zhou. FlexIQ: a flexible interactive querying framework by exploiting the skyline operator. The Journal of Systems and Software, 97(??):97–117, November 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001538.

Isern:2011:OSS

[ISM11] David Isern, David Sánchez, and Antonio Moreno. Organizational structures supported by agentoriented methodologies. The Journal of Systems and Software, 84(2):169– 184, February 2011. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ismail:2013:ISL

[IYS13] Azlan Ismail, Jun Yan, and Jun Shen. Incremental service level agreements violation handling with time impact analysis. The Journal of Systems and Software, 86(6):1530–1544, June 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000216.

Islam:2018:SSE

[IZ18] Md Rakibul Islam and Minhaz F. Zibran. SentiStrength-SE: Exploiting domain specificity for improved sentiment analysis in software engineering text. The Journal of Systems and Software, 145 (??):125–146, November 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301675.

Joshi:2019:IUP

[JAS19] Poornima Joshi, Azin Akbari, and Richard Berntsson Svensson. Impact of usability on process lead-time in information systems: a case study. The Journal of Systems and Software, 148(??):148–169, February 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302383.

Jansen:2012:SGO

[JBSL12] Slinger Jansen, Sjaak Brinkkemper, Jurriaan Souer, and Lutzen Luinenburg. Shades of gray: Opening up a software producing organization with the open software enterprise model. The Journal of Systems and Software, 85(7):1495– 1510, July 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211003013.

Jwo:2010:PSM

[JC10] Jung-Sing Jwo and Yu Chin Cheng. Pseudo software: a mediating instrument for modeling software requirements. The Journal of Systems and Software, 83(4):599–608, April 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jiang:2015:IBA

[JC15] Bo Jiang and W. K. Chan. Input-based adaptive randomized test case prioritization: a local beam search approach. The Journal of Systems and Software, 105(??):91–106, July 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000680.

Jiang:2017:MTA

[JCK^{+}17] Mingyue Jiang, Tsong Yueh Chen, Fei-Ching Kuo, Dave Towey, and Zuohua Ding. A metamorphic testing approach for supporting program repair without the need for a test oracle. The Journal of Systems and Software, 126(??):127–140, April 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300206.

Jia:2016:PPS

[JCYT16] Changjiang Jia, Yan Cai, Yuen Tak Yu, and T. H. Tse. 5W+1H pattern: a perspective of systematic mapping studies and a case study on cloud software testing. The Journal of Systems and Software, 116(??):206– 219, June 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000370.

Jorgensen:2016:IRS

[JDLS16] Magne Jørgensen, Tore Dybå, Knut Liestøl, and Dag I. K. Sjøberg. Incorrect results in software engineering experiments: How to improve research practices. The Journal of Systems and Software, 116(??):133– 145, June 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000679.

Jabangwe:2018:SEP

[JED18] Ronald Jabangwe, Henry Edison, and Anh Nguyen Duc. Software engineering process models for mobile app development: a systematic literature review. The Journal of Systems and Software, 145(??):98–111, November 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301638.

Jagemar:2016:AMC

[JEEL16] Marcus Jägemar, Sigrid Eldh, Andreas Ermedahl, and Björn Lisper. Automatic message compression with overload protection. The Journal of Systems and Software, 121(??):209–222, November 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300267.

Jantunen:2014:UGT

[JG14] Sami Jantunen and Donald C. Gause. Using a grounded theory approach for exploring software product management challenges. The Journal of Systems and Software, 95(??):32–51, September 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0164121214000776.

Jorgensen:2010:ERF

[JH10] Magne Jørgensen and Torleif Halkjelsvik. The effects of request formats on judgment-based effort estimation. The Journal of Systems and Software, 83(1):29–36, January 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jung:2010:FPA

[JHYK10] Hyungsoo Jung, Hyuck Han, Heon Y. Yeom, and Sooyong Kang. A fast and progressive algorithm for skyline queries with totallyand partially-ordered domains. The Journal of Systems and Software, 83 (3):429–445, March 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ju:2014:HEF

[JJC^{+}14] Xiaolin Ju, Shujuan Jiang, Xiang Chen, Xingya Wang, Yanmei Zhang, and Heling Cao. HSFal: Effective fault localization using hybrid spectrum of full slices and execution slices. The Journal of Systems and Software, 90(??):3–17, April 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002823.

Jorgensen:2012:IPR

[JK12] Magne Jørgensen and Barbara Kitchenham. Interpretation problems related to the use of regression models to decide on economy of scale in software development. The Journal of Systems and Software, 85 (11):2494–2503, November 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001549.

Jawad:2013:GAD

[JK13] Khurram Jawad and Asifullah Khan. Genetic algorithm and difference expansion based reversible watermarking for relational databases. The Journal of Systems and Software, 86 (11):2742–2753, November 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001428.

Jiang:2019:TBI

[JKC19] James J. Jiang, Gary Klein, and Jamie Y. T. Chang. Teamwork behaviors in implementing enterprise systems with multiple projects: Results from Chinese firms. The Journal of Systems and Software, 157(??):??, November 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301670.

Jung:2019:ACB

[JKL19] Pilsu Jung, Sungwon Kang, and Jihyun Lee. Automated code-based test selection for software product line regression testing. The Journal of Systems and Software, 158(??):??, December 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301931.

Jaafar:2017:ASE

[JLGM17] Fehmi Jaafar, Angela Lozano, Yann-Gaël Guéhéneuc, and Kim Mens. Analyzing software evolution and quality by extracting Asynchrony change patterns. The Journal of Systems and Software, 131(??):311–322, September 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217300948.

Jha:2019:ESC

[JLL19] Ajay Kumar Jha, Sunghee Lee, and Woo Jin Lee. An empirical study of configuration changes and adoption in Android apps. The Journal of Systems and Software, 156(??):164–180, October 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301396.

Jin:2010:DAM

[JLQ^{+}10] Hai Jin, Yaqin Luo, Li Qi, Jie Dai, and Song Wu. Dependency-aware maintenance for highly available service-oriented grid. The Journal of Systems and Software, 83 (10):1983–1994, October 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jelassi:2014:EUM

[JLY14] M. Nidhal Jelassi, Christine Largeron, and Sadok Ben Yahia. Efficient unveiling of multi-members in a social network. The Journal of Systems and Software, 94(??):30–38, August 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001647.

Jiang:2019:WSM

[JLZ^{+}19] Jing Jiang, David Lo, Jiateng Zheng, Xin Xia, Yun Yang, and Li Zhang. Who should make decision on this pull request? Analyzing time-decaying relationships and file similarities for integrator prediction. The Journal of Systems and Software, 154(??):196–210, August 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300962.

Jovanovic:2017:TOR

[JMML17] Milos Jovanović, Antonia Mas, Antoni-Lluís Mesquida, and Bojan Lalić. Transition of organizational roles in Agile transformation process: a grounded theory approach. The Journal of Systems and Software, 133(??):174–194, November 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217301486.

Jorgensen:2010:SSJ

[Jør10] Magne Jørgensen. Selection of strategies in judgment-based effort estimation. The Journal of Systems and Software, 83(6):1039–1050, June 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jorgensen:2014:FFS

[Jør14] Magne Jørgensen. Failure factors of small software projects at a global outsourcing marketplace. The Journal of Systems and Software, 92(??):157–169, June 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214000429.

Jorgensen:2016:UES

[Jør16] Magne Jørgensen. Unit effects in software project effort estimation: Workhours gives lower effort estimates than workdays. The Journal of Systems and Software, 117(??):274–281, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300085.

Jimenez-Pastor:2017:SME

[JPGdL17] Antonio Jiménez-Pastor, Antonio Garmendia, and Juan de Lara. Scalable model exploration for model-driven engineering. The Journal of Systems and Software, 132(??):204–225, October 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301504.

Jurado:2015:SAM

[JR15] Francisco Jurado and Pilar Rodriguez. Sentiment Analysis in monitoring software development processes: an exploratory case study on GitHub’s project issues. The Journal of Systems and Software, 104(??):82–89, June 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215000485.

Jurado:2012:BAI

[JRO12] Francisco Jurado, Miguel A. Redondo, and Manuel Ortega. Blackboard architecture to integrate components and agents in heterogeneous distributed eLearning systems: an application for learning to program. The Journal of Systems and Software, 85(7):1621– 1636, July 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212000416.

Jung:2010:HIS

[JRSN10] Young Jin Jung, Keun Ho Ryu, Moon Sun Shin, and Silvia Nittel. Historical index structure for reducing insertion and search cost in LBS. The Journal of Systems and Software, 83(8):1500– 1511, August 2010. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jadhav:2011:FES

[JS11] Anil S. Jadhav and Rajendra M. Sonar. Framework for evaluation and selection of the software packages: a hybrid knowledge based system approach. The Journal of Systems and Software, 84(8):1394–1407, August 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jiao:2013:SAD

[JS13] Wenpin Jiao and Yanchun Sun. Supporting adaptation of decentralized software based on application scenarios. The Journal of Systems and Software, 86(7):1891–1906, July 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000551.

Jiao:2016:SAM

[JS16] Wenpin Jiao and Yanchun Sun. Self-adaptation of multi-agent systems in dynamic environments based on experience exchanges. The Journal of Systems and Software, 122(??):165–179, December 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301844.

Jannach:2014:AFF

[JSHW14] Dietmar Jannach, Thomas Schmitz, Birgit Hofer, and Franz Wotawa. Avoiding, finding and fixing spreadsheet errors — a survey of automated approaches for spreadsheet QA. The Journal of Systems and Software, 94(??):129–150, August 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214000788.

Jaber:2016:ESE

[JSL16] Khaled Jaber, Bonita Sharif, and Chang Liu. An empirical study on the effect of 3D visualization for project tasks and resources. The Journal of Systems and Software, 115(??):1–17, May 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121600008X.

Jiao:2010:AAI

[JSM10] Wenpin Jiao, Yanchun Sun, and Hong Mei. Automated assembly of Internet-scale software systems involving autonomous agents. The Journal of Systems and Software, 83(10):1838– 1850, October 2010. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Joshi:2010:MEH

[JST10] Anirudha Joshi, N. L.

Sarda, and Sanjay Tripathi. Measuring effectiveness of HCI integration in software development processes. The Journal of Systems and Software, 83(11):2045– 2058, November 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tong:2012:NBD

[jT12] Xiao jun Tong. The novel bilateral — Diffusion image encryption algorithm with dynamical compound chaos. The Journal of Systems and Software, 85(4):850–858, April 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002834.

Juang:2010:R

[Jua10] Wen-Shenq Juang. RO.

The Journal of Systems and Software, 83(4):638– 645, April 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jalali:2014:IAA

[JWA14] Samireh Jalali, Claes Wohlin, and Lefteris Angelis. Investigating the applicability of Agility assessment surveys: a case study. The Journal of Systems and Software, 98(??):172–190, December 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214001927.

Wu:2013:CBD

[jWLY^{+}13] Zhi jun Wu, Jin Lei, Di Yao, Ming hua Wang, and Sarhan M. Musa. Chaos-based detection of LDoS attacks. The Journal of Systems and Software, 86(1):211–221, January 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212002269.

Jannesari:2017:ESI

[JWT17] Ali Jannesari, Felix Wolf, and Walter F. Tichy. Editorial of special issue on Software Engineering for Parallel Systems. The Journal of Systems and Software, 125(??): 380, March 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

//www.sciencedirect.
com/science/article/
pii/S0164121216301881.

Jiang:2015:NCB

[JXLC15] Dingde Jiang, Zhengzheng Xu, Wenpan Li, and Zhenhua Chen. Network coding-based energyefficient multicast routing algorithm for multihop wireless networks. The Journal of Systems and Software, 104(??): 152–165, June 2015. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121215000576.

Kijsipongse:2014:ICP

[KA14] Ekasit Kijsipongse and Namfon Assawamekin. Improving the communication performance of distributed animation rendering using BitTorrent file system. The Journal of Systems and Software, 97 (??):178–191, November 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S016412121400171X.

Kubota:2017:ASG

[KA17] Takuya Kubota and Masayoshi Aritsugi. Assignment strategies for ground truths in the crowdsourcing of labeling tasks. The Journal of Systems and Software, 126(??):113–126, April 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216300930.

Khomh:2018:UIC

[KA18] Foutse Khomh and S. Amirhossein Abtahizadeh. Understanding the impact of cloud patterns on performance and energy consumption. The Journal of Systems and Software, 141(??):151–170, July 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218300621.

Kiani:2013:FBS

[KAK^{+}13] Saad Liaquat Kiani, Ashiq Anjum, Michael Knappmeyer, Nik Bessis, and Nikolaos Antonopoulos. Federated broker system for pervasive context provisioning. The Journal of Systems and Software, 86(4):1107–1123, April 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212003275.

Koziolek:2013:HMA

[KAM13] Anne Koziolek, Danilo Ardagna, and Raffaela Mirandola. Hybrid multi-attribute QoS optimization in component based software systems. The Journal of Systems and Software, 86(10):2542–2558, October 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000800.

Kang:2015:EDA

[Kan15]  Pilsung Kang.  The

effects of different alphabets on free text keystroke authentication: a case study on the Korean–English users. The Journal of Systems and Software, 102(??):1– 11, April 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214002842.

Kaminski:2013:ILB

[KAO13] Gary Kaminski, Paul Ammann, and Jeff Offutt. Improving logicbased testing. The Journal of Systems and Software, 86(8):2002–2012, August 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212002403.

Kirac:2018:VFI

[KAS18] M. Furkan Kiraç, Baris Aktemur, and Hasan Sözer. VISOR: a fast image processing pipeline with scaling and translation invariance for test oracle automation of visual output systems. The Journal of Systems and Software, 136(??):266–277, February 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217301243.

Kahveci:2016:ISF

[KAU16] Basri Kahveci, Ismail Sengör Altingövde, and Özgür Ulusoy. Integrating social features into mobile local search. The Journal of Systems and Software, 122(??):155–164, December 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301819.

Kravari:2016:DSD

[KB16] Kalliopi Kravari and Nick Bassiliades. DIS-ARM: a social distributed agent reputation model based on defeasible logic. The Journal of Systems and Software, 117(??):130–152, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000492.

Karimi:2016:LBP

[KBDGAW16] Zahra Karimi, Ahmad Baraani-Dastjerdi, Nasser Ghasem-Aghaee, and Stefan Wagner. Links between the personalities, styles and performance in computer programming. The Journal of Systems and Software, 111(??):228–241, January 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121500206X.

Khelladi:2017:SAM

[KBHG17] Djamel Eddine Khelladi, Reda Bendraou, Regina Hebig, and Marie-Pierre Gervais. A semiautomatic maintenance and co-evolution of OCL constraints with (meta)model evolution. The Journal of Systems and Software, 134(??):242–260, December 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121730198X.

Kabbedijk:2015:DMT

[KBJZ15] Jaap Kabbedijk, Cor-Paul Bezemer, Slinger Jansen, and Andy Zaidman. Defining multitenancy: a systematic mapping study on the academic and the industrial perspective. The Journal of Systems and Software, 100 (??):139–148, February 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214002313.

Kosar:2018:SMS

[KBM18] Tomaz Kosar, Sudev Bohra, and Marjan Mernik. A systematic mapping study driven by the margin of error. The Journal of Systems and Software, 144(??):439–449, October 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301353.

Kumar:2017:SSD

[KBRV17] Dinesh Kumar, Gaurav Baranwal, Zahid Raza, and Deo Prakash Vidyarthi. A systematic study of double auction mechanisms in cloud computing. The Journal of Systems and Software, 125(??):234– 255, March 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

//www.sciencedirect.
com/science/article/
pii/S0164121216302540.

Kumar:2018:TCD

[KBRV18] Dinesh Kumar, Gaurav Baranwal, Zahid Raza, and Deo Prakash Vidyarthi. A truthful combinatorial double auctionbased marketplace mechanism for cloud computing. The Journal of Systems and Software, 140(??):91–108, June 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300451.

Kao:2016:DLA

[KC16] Yu-Chon Kao and Ya-Shu Chen. Data-locality-aware mapreduce realtime scheduling framework. The Journal of Systems and Software, 112(??):65–77, February 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002344.

Kakarontzas:2013:LAO

[KCAS13] George Kakarontzas, Eleni Constantinou, Apostolos Ampatzoglou, and Ioannis Stamelos. Layer assessment of objectoriented software: a metric facilitating white-box reuse. The Journal of Systems and Software, 86(2):349–366, February 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0164121212002452.

Karakoyunlu:2016:ADA

[KCR16]  Cengiz  Karakoyunlu,

John A. Chandy, and Alma Riska. Adding data analytics capabilities to scaled-out object store. The Journal of Systems and Software, 121(??):16–27, November 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121630125X.

Kan:2012:EEC

[KCT12] Edward Y. Y. Kan, W. K. Chan, and T. H. Tse. EClass: an execution classification approach to improving the energy-efficiency of software via machine learning. The Journal of Systems and Software, 85(4):960–973, April 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002913.

Kousiouris:2011:ESW

[KCV11] George Kousiouris, Tommaso Cucinotta, and Theodora Varvarigou.

The effects of scheduling, workload type and consolidation scenarios on virtual machine performance and their prediction through optimized artificial neural networks. The Journal of Systems and Software, 84(8): 1270–1291, August 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kicsi:2019:FAU

[KCV^{+}19] András Kicsi, Viktor Csuvik, László Vidács, Ferenc Horváth, Árpád Beszédes, Tibor Gyimóthy, and Ferenc Kocsis. Feature analysis using information retrieval, community detection and structural analysis methods in product line adoption. The Journal of Systems and Software, 155(??):70–90, September 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301049.

Kleiner:2018:GAM

[KD18] Mathias Kleiner and Marcos Didonet Del Fabro. A generic approach to model generation operations. The Journal of Systems and Software, 142(??):136– 155, August 2018. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121218300839.

Kelly:2015:SSD

[Kel15] Diane Kelly. Scientific software development viewed as knowledge acquisition: Towards understanding the development of risk-averse scientific software. The Journal of Systems and Software, 109(??):50–61, November 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001533.

Kechagia:2018:EHR

[KFLS18] Maria Kechagia, Marios Fragkoulis, Panos Louridas, and Diomidis Spinellis. The exception handling riddle: An empirical study on the Android API. The Journal of Systems and Software, 142(??):248–270, August 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121218300724.

Kashfi:2019:IUP

[KFN19] Pariya Kashfi, Robert Feldt, and Agneta Nilsson. Integrating UX principles and practices into software development organizations: a case study of influencing events. The Journal of Systems and Software, 154(??):37–58, August 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219300731.

Kim:2010:AAS

[KG10] Jung Soo Kim and David Garlan. Analyzing architectural styles. The Journal of Systems and Software, 83(7):1216–1235, July 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Karg:2011:SLR

[KGB11] Lars M. Karg, Michael Grottke, and Arne Beckhaus. A systematic literature review of software quality cost research. The Journal of Systems and Software, 84 (3):415–427, March 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Klos:2018:RMQ

[KGG18] Verena Klös, Thomas Göthel, and Sabine Glesner. Runtime management and quantitative evaluation of changing system goals in complex autonomous systems. The Journal of Systems and Software, 144(??):314–327, October 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121830133X.

Kazman:2012:SSA

[KGW12] Rick Kazman, Michael Gagliardi, and William Wood. Scaling up software architecture analysis. The Journal of Systems and Software, 85(7):1511–1519, July 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211000793.

Kuo:2010:CAO

[KH10] Jong Yih Kuo and Fu Chu Huang. Code analyzer for an online course management system. The Journal of Systems and Software, 83 (12):2478–2486, December 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Khan:2014:BCF

[KH14] Imran Khan and Sajjad Haider. On building a consistent framework for executable systems architecture. The Journal of Systems and Software, 98(??):155–171, December 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001897.

Kwon:2016:CDR

[KHC16] Jang-Jin Kwon, Jang-Eui Hong, and Lawrence Chung. Collision detection and resolution of hazard prevention actions in safety critical systems. The Journal of Systems and Software, 118(??):1–18, August 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300322.

Kilamo:2012:POS

[KHMA12] Terhi Kilamo, Imed Hammouda, Tommi Mikkonen, and Timo Aaltonen. From proprietary to open source — growing an open source ecosystem. The Journal of Systems and Software, 85(7):1467–1478, July 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001683.

Kuo:2013:AHL

[KHMF13] Jong Yih Kuo, Fu-Chu Huang, Shang-Pin Ma, and Yong-Yi Fanjiang. Applying hybrid learning approach to RoboCup’s strategy. The Journal of Systems and Software, 86(7):1933–1944, July 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000605.

Kang:2011:TAH

[KHS11] Qinma Kang, Hong He, and Huimin Song. Task assignment in heterogeneous computing systems using an effective iterated greedy algorithm. The Journal of Systems and Software, 84(6):985–992, June 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kang:2010:TAM

[KHSD10] Qin-Ma Kang, Hong He, Hui-Min Song, and Rong Deng. Task allocation for maximizing reliability of distributed computing systems using honeybee mating optimization. The Journal of Systems and Software, 83 (11):2165–2174, November 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Koch:2019:RSS

[KHW19] Patrick Koch, Birgit Hofer, and Franz Wotawa. On the refinement of spreadsheet smells by means of structure information. The Journal of Systems and Software, 147(??):64–85, January 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121830219X.

Kim:2012:DFA

[Kim12] Chong Hee Kim. Differential fault analysis of ARIA in multi-byte fault models. The Journal of Systems and Software, 85 (9):2096–2103, September 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001033.

Kim:2017:EEB

[Kim17] Saehwa Kim. Efficient exact Boolean schedulability tests for fixed priority preemption threshold scheduling. The Journal of Systems and Software, 134(??):153–169, December 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301978.

Kitchenham:2010:WSM

[Kit10] Barbara Kitchenham.

What’s up with software metrics? — a preliminary mapping study. The Journal of Systems and Software, 83(1):37– 51, January 2010. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kropik:2010:SPS

[KJ10] Martin Kropik and Monika Jurickova. Software for protection system of VR-1 training reactor. The Journal of Systems and Software, 83(9):1605– 1611, September 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Khakpour:2012:HMA

[KJS^{+}12] Narges Khakpour, Saeed Jalili, Marjan Sirjani, Ursula Goltz, and Bahareh Abolhasanzadeh. HPobSAM for modeling and analyzing IT Ecosystems — Through a case study. The Journal of Systems and Software, 85 (12):2770–2784, December 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000738.

Karaoglanoglou:2011:RDG

[KK11] K. Karaoglanoglou and H. Karatza. Resource discovery in a Grid system: Directing requests to trustworthy virtual organizations based on global trust values. The Journal of Systems and Software, 84(3):465–478, March 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kapus-Kolar:2012:EAT

[KK12] Monika Kapus-Kolar. On “Exploring alternatives for transition verification”. The Journal of Systems and Software, 85(8):1744–1748, August 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000817.

Kalamatianos:2017:DAF

[KK17a] Theodoros Kalamatianos and Kostas Kontogiannis. Distributed analysis and filtering of application event streams. The Journal of Systems and Software, 129(??):1– 25, July 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217300687.

Koo:2017:CUP

[KK17b] Hyung-Min Koo and In-Young Ko. Construction and utilization of problem-solving knowledge in open source software environments. The Journal of Systems and Software, 131 (??):402–418, September 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300899.

Koloniaris:2019:SBI

[KKA^{+}19] Stavros Koloniaris, George Kousiouris, Dimosthenis Anagnostopoulos, Mara Nikolaidou, and Konstantinos Tserpes. Survey-based investigation, feature extraction and classification of Greek municipalities maturity for open source adoption and migration prospects. The Journal of Systems and Software, 158(??):??, December 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219302055.

Katsaros:2012:SAH

[KKG^{+}12] Gregory Katsaros, George Kousiouris, Spyridon V. Gogouvitis, Dimosthenis Kyriazis, Andreas Menychtas, and Theodora Varvarigou. A self-adaptive hierarchical monitoring mechanism for Clouds. The Journal of Systems and Software, 85(5): 1029–1041, May 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121211002998.

Khlif:2016:EVA

[KKH^{+}16] Houda Khlif, Hatem Hadj Kacem, Saúl E. Pomares Hernandez, Ahmed Hadj Kacem, Cédric Eichler, and Alberto Calixto Simón. An efficient validation approach for quasi-synchronous checkpointing oriented to distributed diagnosability. The Journal of Systems and Software, 122 (??):364–377, December 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300395.

Kim:2011:ECR

[KKHH11] Sung-Kyoung Kim, Tae Hyun Kim, Dong-Guk Han, and Seokhie Hong. An efficient CRT-RSA algorithm secure against power and fault attacks. The Journal of Systems and Software, 84 (10):1660–1669, October 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121211001014.

Kim:2011:FBA

[KKL^{+}11] Sangsig Kim, Dae-Kyoo Kim, Lunjin Lu, Suntae Kim, and Sooyong Park. A feature-based approach for modeling role-based access control systems. The Journal of Systems and Software, 84(12):2035–2052, December 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211000926.

Kwon:2011:FEG

[KKLB11] Ohhoon Kwon, Kern Koh, Jaewoo Lee, and Hyokyung Bahn. FeGC: an efficient garbage collection scheme for flash memory based storage systems. The Journal of Systems and Software, 84 (9):1507–1523, September 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211000665.

Kim:2012:STM

[KKLC12] Jingyu Kim, Sungwon Kang, Jihyun Lee, and Bong Wan Choi. A semantic translation method for data communication protocols. The Journal of Systems and Software, 85(12):2876–2898, December 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001720.

Kim:2012:SCA

[KKP12] Tae Hyun Kim, ChangKyun Kim, and IlHwan Park. Side channel analysis attacks using AM demodulation on commercial smart cards with SEED. The Journal of Systems and Software, 85 (12):2899–2908, December 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121200194X.

Kiran:2016:EDP

[KKR16] R. Uday Kiran, Masaru Kitsuregawa, and P. Krishna Reddy. Efficient discovery of periodicfrequent patterns in very large databases. The Journal of Systems and Software, 112 (??):110–121, February 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215002307.

Kapitsaki:2017:ALC

[KKT17] Georgia M. Kapitsaki, Frederik Kramer, and Nikolaos D. Tselikas. Automating the license compatibility process in open source software with SPDX. The Journal of Systems and Software, 131(??):386–401, September 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216300905.

Kennard:2010:TGP

[KL10] Richard Kennard and John Leaney. Towards a general purpose architecture for UI generation. The Journal of Systems and Software, 83 (10):1896–1906, October 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kennard:2011:TCF

[KL11] Richard Kennard and John Leaney. Is there convergence in the field of UI generation? The Journal of Systems and Software, 84(12):2079– 2087, December 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121211001348.

Karanatsiou:2019:BAS

[KLA^{+}19] Dimitra Karanatsiou, Yihao Li, Elvira-Maria Arvanitou, Nikolaos Misirlis, and W. Eric Wong. A bibliometric assessment of software engineering scholars and institutions (2010–2017). The Journal of Systems and Software, 147(??):246–261, January 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302334.

Kim:2015:EAE

[KLB15] Taehyoun Kim, Kwangkyu Lee, and Jongmoon Baik. An effective approach to estimating the parameters of software reliability growth models using a real-valued genetic algorithm. The Journal of Systems and Software, 102(??):134–144, April 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121215000023.

Kim:2011:MMS

[KLL^{+}11] Byoungjip Kim, Sang-Jeong Lee, Youngki Lee, Inseok Hwang, Yunseok Rhee, and Junehwa Song. Mobiiscape: Middleware support for scalable mobility pattern monitoring of moving objects in a largescale city. The Journal of Systems and Software, 84(11):1852–1870, November 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211001671.

Kim:2017:DPB

[KLL17] Dae-Kyoo Kim, Lunjin Lu, and Byunghun Lee. Design patternbased model transformation supported by QVT. The Journal of Systems and Software, 125(??): 289–308, March 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216302564.

Kim:2010:RFD

[KLP10] Heeyoul Kim, Younho Lee, and Yongsu Park. A robust and flexible digital rights management system for home networks. The Journal of Systems and Software, 83 (12):2431–2440, December 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kiani:2011:MPD

[KM11] Soheila Kiani and Mohsen Ebrahimi Moghaddam. A multipurpose digital image watermarking using fractal block coding. The Journal of Systems and Software, 84(9):1550– 1562, September 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121211000707.

Kocaguneli:2013:SEM

[KM13] Ekrem Kocaguneli and Tim Menzies. Software effort models should be assessed via leave-one-out validation. The Journal of Systems and Software, 86(7):1879–1890, July 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000538.

Korkala:2014:WIM

[KM14] Mikko Korkala and Frank Maurer. Waste identification as the means for improving communication in globally distributed agile software development. The Journal of Systems and Software, 95(??):122–140, September 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214001009.

Kaur:2017:SCS

[KM17] Loveleen Kaur and Ashutosh Mishra. Software component and the Semantic Web: an in-depth content analysis and integration history. The Journal of Systems and Software, 125(??):152– 169, March 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216302308.

Kruger:2019:WMF

[KMG^{+}19] Jacob Krüger, Mukelabai Mukelabai, Wanzi Gu, Hui Shen, Regina Hebig, and Thorsten Berger. Where is my feature and what is it about? A case study on recovering feature facets. The Journal of Systems and Software, 152(??):239– 253, June 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300184.

Kos:2016:TAM

[KMK16] Tomaz Kos, Marjan Mernik, and Tomaz Kosar. Test automation of a measurement system using a domain-specific modelling language. The Journal of Systems and Software, 111(??):74–88, January 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002058.

Katsikas:2017:PAC

[KMK17] Georgios P. Katsikas, Gerald Q. Maguire, Jr., and Dejan Kostić. Profiling and accelerating commodity NFV service chains with SCC. The Journal of Systems and Software, 127(??):12–27, May 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121217300055.

Kim:2012:ENL

[KMWL12] Soon-Kyeong Kim, Toby Myers, Marc-Florian Wendland, and Peter A. Lindsay. Execution of natural language requirements using State Machines synthesised from Behavior Trees. The Journal of Systems and Software, 85 (11):2652–2664, November 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212001690.

Khan:2011:FIC

[KNA11] Siffat Ullah Khan, Mahmood Niazi, and Rashid Ahmad. Factors influencing clients in the selection of offshore software outsourcing vendors: an exploratory study using a systematic literature review. The Journal of Systems and Software, 84(4): 686–699, April 2011. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kessentini:2014:SBM

[KOL^{+}14] Marouane Kessentini, Ali Ouni, Philip Langer, Manuel Wimmer, and Slim Bechikh. Searchbased metamodel matching with structural and syntactic measures. The Journal of Systems and Software, 97(??):1–14, November 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001484.

Karpati:2015:IST

[KOS15] Peter Karpati, Andreas L.

Opdahl, and Guttorm Sindre. Investigating security threats in architectural context: Experimental evaluations of misuse case maps. The Journal of Systems and Software, 104(??):90– 111, June 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000461.

Keil:2010:BNR

[KP10] Mark Keil and Chong-Woo Park. Bad news reporting on troubled IT projects: Reassessing the mediating role of responsibility in the basic whistleblowing model. The Journal of Systems and Software, 83 (11):2305–2316, November 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Keskinarkaus:2010:IWD

[KPS10] A. Keskinarkaus, A. Pramila, and T. Seppänen. Image watermarking with a directed periodic pattern to embed multibit messages resilient to printscan and compound attacks. The Journal of Systems and Software, 83 (10):1715–1725, October 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kellaris:2013:MMT

[KPT13] Georgios Kellaris, Nikos Pelekis, and Yannis Theodoridis. Mapmatched trajectory compression. The Journal of Systems and Software, 86(6):1566–1579, June 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213000289.

Kratzke:2017:UCN

[KQ17] Nane Kratzke and Peter-Christian Quint. Understanding cloud-native applications after 10 years of cloud computing — a systematic mapping study. The Journal of Systems and Software, 126(??):1–16, April 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217300018.

Keivanloo:2014:STS

[KR14] Iman Keivanloo and Juergen Rilling. Software trustworthiness 2.0 — a Semantic Web enabled global source code analysis approach. The Journal of Systems and Software, 89(??):33–50, March 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002173.

Kumar:2016:HFL

[KR16] Lov Kumar and Santanu Ku. Rath. Hybrid functional link artificial neural network approach for predicting maintainability of object-oriented software. The Journal of Systems and Software, 121(??):170–190, November 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121216000054.

Kallel:2016:ARS

[KRD16] Slim Kallel, Ismail Bouassida Rodruigez, and Khalil Drira. Adaptive and reconfigurable software systems and architectures. The Journal of Systems and Software, 122(??):342–343, December 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216301790.

Kannan:2012:SFC

[KRDH12] S. R. Kannan, S. Ramathilagam, R. Devi, and E. Hines. Strong fuzzy c-means in medical image data analysis. The Journal of Systems and Software, 85 (11):2425–2438, November 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211003050.

Khabou:2017:NAA

[KRJ17] Nesrine Khabou, Ismael Bouassida Rodriguez, and Mohamed Jmaiel. A novel analysis approach for the design and the development of context-aware applications. The Journal of Systems and Software, 133(??):113–125, November 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301528.

Kumari:2016:HHA

[KS16] A. Charan Kumari and K. Srinivas. Hyperheuristic approach for multi-objective software module clustering. The Journal of Systems and Software, 117(??):384– 401, July 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300231.

Kaur:2019:HDO

[KS19] Satnam Kaur and Paramvir Singh. How does objectoriented code refactoring influence software quality? Research landscape and challenges. The Journal of Systems and Software, 157(??):??, November 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121219301694.

Khanbabaei:2018:DIF

[KSAR18] Mohammad Khanbabaei, Farzad Movahedi Sobhani, Mahmood Alborzi, and Reza Radfar. Developing an integrated framework for using data mining techniques and ontology concepts for process improvement. The Journal of Systems and Software, 137(??):78–95, March 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217302613.

Khorsand:2017:TWP

[KSENM17] Reihaneh Khorsand, Faramarz Safi-Esfahani, Naser Nematbakhsh, and Mehran Mohsenzade. Taxonomy of workflow partitioning problems and methods in distributed environments. The Journal of Systems and Software, 132(??):253–271, October 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217300821.

Koong:2012:ATE

[KSH^{+}12] Chorng-Shiuh Koong, Chihhsiong Shih, Pao-Ann Hsiung, Hung-Jui Lai, Chih-Hung Chang, William C. Chu, Nien-Lin Hsueh, and Chao-Tung Yang. Automatic testing environment for multi-core embedded software — ATEMES. The Journal of Systems and Software, 85(1):43–60, January 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002305.

Kuo:2014:CLM

[KSHC14] Jun-Li Kuo, Chen-Hua Shih, Cheng-Yuan Ho, and Yaw-Chung Chen. A cross-layer middleware for context-aware cooperative application on mobile ad hoc peerto-peer network. The Journal of Systems and Software, 92(??):95–106, June 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002422.

Khan:2019:LSM

[KSIZ19] Muhammad Uzair Khan, Salman Sherin, Muhammad Zohaib Iqbal, and Rubab Zahid. Landscaping systematic mapping studies in software engineering: a tertiary study. The Journal of Systems and Software, 149(??):396–436, March 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302784.

Kefalakis:2011:ARX

[KSKP11] Nikos Kefalakis, John Soldatos, Nikolaos Konstantinou, and Neeli R. Prasad. APDL: a reference XML schema for process-centered definition of RFID solutions. The Journal of Systems and Software, 84(7): 1244–1259, July 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Konnola:2016:AME

[KSM^{+}16] Kaisa Könnölä, Samuli Suomi, Tuomas Mäkilä, Tero Jokela, Ville Rantala, and Teijo Lehtonen. Agile methods in embedded system development: Multiple-case study of three industrial cases. The Journal of Systems and Software, 118(??):134–150, August 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300413.

Keshanchi:2017:IGA

[KSN17] Bahman Keshanchi, Alireza Souri, and Nima Jafari Navimipour. An improved genetic algorithm for task scheduling in the cloud environments using the priority queues: Formal verification, simulation, and statistical testing. The Journal of Systems and Software, 124(??):1–21, February 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301066.

Kannan:2010:NSA

[KSRD10] S. R. Kannan, A. Sathya, S. Ramathilagam, and R. Devi. Novel segmentation algorithm in segmenting medical images. The Journal of Systems and Software, 83 (12):2487–2495, December 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kundu:2015:UMB

[KSS15] Debasish Kundu, Monalisa Sarma, and Debasis Samanta. A UML model-based approach to detect infeasible paths. The Journal of Systems and Software, 107 (??):71–92, September 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215001016.

Kirk:2012:LFD

[KT12] Diana Kirk and Ewan Tempero. A lightweight framework for describing software practices. The Journal of Systems and Software, 85(3):582–595, March 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121100241X.

Klein:2016:BPW

[KT16] John Klein and Antony Tang. Best papers from the 11^{th} Working IEEE/IFIP Conference on Software Architecture (WICSA 2014 7^{th}–11^{th} April 2014). The Journal of Systems and Software, 112(??):122, February 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002472.

Kapitsaki:2015:ILT

[KTF15] Georgia M. Kapitsaki, Nikolaos D. Tselikas, and Ioannis E. Foukarakis. An insight into license tools for open source software systems. The Journal of Systems and Software, 102(??):72–87, April 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002945.

Kuhrmann:2016:FSP

[KTF^{+}16] Marco Kuhrmann, Thomas Ternité, Jan Friedrich, Andreas Rausch, and Manfred Broy. Flexible software process lines in practice: a metamodelbased approach to effectively construct and manage families of software process models. The Journal of Systems and Software, 121(??):49–71, November 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121216301236.

Kyriakou:2019:ECC

[KTK19] Kyriakos-Ioannis D. Kyriakou, Nikolaos D. Tselikas, and Georgia M. Kapitsaki. Enhancing C/C++ based OSS development and discoverability with CBRJS: a

                                        Rust/Node.js/WebAssembly

framework for repackaging legacy codebases. The Journal of Systems and Software, 157(??):??, November 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301700.

Kallel:2017:GRS

[KTT^{+}17] Sahar Kallel, Bastien Tramoni, Chouki Tibermacine, Christophe Dony, and Ahmed Hadj Kacem. Generating reusable, searchable and executable “architecture constraints as services”. The Journal of Systems and Software, 127(??):91–108, May 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217300213.

Khwaja:2010:PBS

[KU10] Amir A. Khwaja and Joseph E. Urban. A property based specification formalism classification. The Journal of Systems and Software, 83 (11):2344–2362, November 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Khomh:2011:BGB

[KVGS11] Foutse Khomh, Stephane Vaucher, Yann-Gaël Guéhéneuc, and Houari Sahraoui. BDTEX: a GQM-based Bayesian approach for the detection of antipatterns. The Journal of Systems and Software, 84 (4):559–572, April 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kolomvatsos:2012:DAC

[KVH12] Kostas Kolomvatsos, George Valkanas, and Stathes Hadjiefthymiades. Debugging applications created by a Domain Specific Language: The IPAC case. The Journal of Systems and Software, 85(4):932–943, April 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121211002901.

Kiran:2017:DPP

[KVT^{+}17] R. Uday Kiran, J. N. Venkatesh, Masashi Toyoda, Masaru Kitsuregawa, and P. Krishna Reddy. Discovering partial periodic-frequent patterns in a transactional database. The Journal of Systems and Software, 125(??):170– 182, March 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216302382.

Kevic:2017:EGI

[KWS^{+}17] K. Kevic, B. M. Walters, T. R. Shaffer, B. Sharif, D. C. Shepherd, and T. Fritz. Eye gaze and interaction contexts for change tasks — Observations and potential. The Journal of Systems and Software, 128(??):252– 266, June 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216000881.

Kim:2010:PBP

[KY10] JunSeong Kim and Jongsu Yi. A pattern-based prediction: an empirical approach to predict end-to-end network latency. The Journal of Systems and Software, 83(11):2317– 2321, November 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Li:2016:UPO

[LAH^{+}16] Zhen Li, Rohit Atre, Zia Huda, Ali Jannesari, and Felix Wolf. Unveiling parallelization opportunities in sequential programs. The Journal of Systems and Software, 117(??):282–295, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121630005X.

Li:2015:SMS

[LAL15] Zengyang Li, Paris Avgeriou, and Peng Liang. A systematic mapping study on technical debt and its management. The Journal of Systems and Software, 101(??):193– 220, March 2015. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121214002854.

Lizcano:2014:CCB

[LASL14] D. Lizcano, F. Alonso, J. Soriano, and G. López. A component- and connectorbased approach for enduser composite web applications development. The Journal of Systems and Software, 94 (??):108–128, August 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214000740.

Li:2010:MFQ

[LAT10] Zhao Li, Nasser Alaeddine, and Jeff Tian. Multi-faceted quality and defect measurement for Web software and source contents. The Journal of Systems and Software, 83 (1):18–28, January 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lin:2010:UQS

[LBCL10] Chi-Nan Lin, Daniel J. Buehrer, Chin-Chen Chang, and Tzu-Chuen Lu. Using quad smoothness to efficiently control capacity–distortion of reversible data hiding. The Journal of Systems and Software, 83(10):1805– 1812, October 2010. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liu:2012:CVS

[LBX12] Hongzhe Liu, Hong Bao, and De Xu. Concept vector for semantic similarity and relatedness based on WordNet structure. The Journal of Systems and Software, 85(2):370–381, February 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002299.

Lee:2010:NDH

[LC10] Chin-Feng Lee and Hsing-Ling Chen. A novel data hiding scheme based on modulus function. The Journal of Systems and Software, 83(5):832–843, May 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liu:2011:CAR

[LC11] Shengpu Liu and Liang Cheng. A contextaware reflective middleware framework for distributed real-time and embedded systems. The Journal of Systems and Software, 84(2):205–218, February 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:2010:EPC

[LCC10] Chun-Hee Lee, Chin-Wan Chung, and Seok-Ju Chun. Effective processing of continuous groupby aggregate queries in sensor networks. The Journal of Systems and Software, 83(12):2627– 2641, December 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lin:2013:IDB

[LCC^{+}13] Tseng-Jung Lin, Kuo-Liang Chung, Po-Chun Chang, Yong-Huai Huang, Hong-Yuan Mark Liao, and Chiung-Yao Fang. An improved DCT-based perturbation scheme for high capacity data hiding in H.264/AVC intra frames. The Journal of Systems and Software, 86(3):604–614, March 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212003020.

Li:2010:SDT

[LCCJ10] Yuzhu Li, Kuo-Chung Chang, Houn-Gee Chen, and James J. Jiang. Software development team flexibility antecedents. The Journal of Systems and Software, 83 (10):1726–1734, October 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liu:2010:CSA

[LCJ10] Xiaoqing (Frank) Liu, Carl K. Chang, and T. Ming Jiang. Computer software and applications. The Journal of Systems and Software, 83 (2):173, February 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lin:2012:TCO

[LCL^{+}12] Ying-Dar Lin, Chi-Heng Chou, Yuan-Cheng Lai, Tse-Yau Huang, Simon Chung, Jui-Tsun Hung, and Frank C. Lin. Test coverage optimization for large code problems. The Journal of Systems and Software, 85(1):16–27, January 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121100121X.

Lu:2015:VSB

[LCL15] Chien-Tung Lu, Chia-Wei Chang, and Jung-Shian Li. VM scaling based on Hurst exponent and Markov transition with empirical cloud data. The Journal of Systems and Software, 99(??):199–207, January 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002155.

Leung:2013:ARD

[LCLF13] H. Y. Leung, L. M. Cheng, F. Liu, and Q. K. Fu. Adaptive reversible data hiding based on block median preservation and modification of prediction errors. The Journal of Systems and Software, 86(8):2204–2219, August 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000903.

Lambolais:2016:IFI

[LCLP16] Thomas Lambolais, Anne-Lise Courbis, Hong-Viet Luong, and Christian Percebois. IDF: a framework for the incremental development and conformance verification of UML active primitive components. The Journal of Systems and Software, 113(??):275–295, March 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002459.

Lee:2016:TLP

[LCLS16] Jiyeon Lee, Hoon Sung Chwa, Jinkyu Lee, and Insik Shin. Threadlevel priority assignment in global multiprocessor scheduling for DAG tasks. The Journal of Systems and Software, 113(??):246–256, March 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002800.

LazzariniLemos:2013:ESS

[LCM^{+}13] Otávio Augusto Lazzarini Lemos, Fabiano Cutigi Ferrari, Marcelo Medeiros Eler, José Carlos Maldonado, and Paulo Cesar Masiero. Evaluation studies of software testing research in Brazil and in the world: a survey of two premier software engineering conferences. The Journal of Systems and Software, 86(4):951–969, April 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212003238.

Lee:2010:ECR

[LCT10] Chin-Feng Lee, Hsing-Ling Chen, and Hao-Kuan Tso. Embedding capacity raising in reversible data hiding based on prediction of difference expansion. The Journal of Systems and Software, 83(10):1864– 1872, October 2010. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lun:2019:SAC

[LDS^{+}19] Yuriy Zacchia Lun, Alessandro D’Innocenzo, Francesco Smarra, Ivano Malavolta, and Maria Domenica Di Benedetto. State of the art of cyber-physical systems security: an automatic control perspective. The Journal of Systems and Software, 149(??):174–216, March 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218302681.

Lin:2015:CBF

[LDZL15] Wenmin Lin, Wanchun Dou, Zuojian Zhou, and Chang Liu. A cloud-based framework for home-diagnosis service over big medical data. The Journal of Systems and Software, 102(??):192–206, April 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001368.

Lee:2011:ZLB

[LESL11] Jinkyu Lee, Arvind Easwaran, Insik Shin, and Insup Lee. Zerolaxity based real-time multiprocessor scheduling. The Journal of Systems and Software, 84 (12):2324–2333, December 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001828.

Lelli:2012:ECD

[LFCL12] Juri Lelli, Dario Faggioli, Tommaso Cucinotta, and Giuseppe Lipari. An experimental comparison of different real-time schedulers on multicore systems. The Journal of Systems and Software, 85(10):2405–2416, October 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121200146X.

Lenberg:2015:BSE

[LFW15] Per Lenberg, Robert Feldt, and Lars Göran Wallgren. Behavioral software engineering: a definition and systematic literature review. The Journal of Systems and Software, 107 (??):15–37, September 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000989.

Li:2015:QPD

[LG15] Wei Li and William Guo. QoS prediction for dynamic reconfiguration of component based software systems. The Journal of Systems and Software, 102(??):12–34, April 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002787.

Leite:2017:HLA

[LG17] Adriana Leite and Rosario Girardi. A hybrid and learning agent architecture for network intrusion detection. The Journal of Systems and Software, 130(??):59–80, August 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217300183.

Larusdottir:2017:LKI

[LGC17] Marta Larusdottir, Jan Gulliksen, and Åsa Cajander. A license to kill — improving UCSD in Agile development. The Journal of Systems and Software, 123 (??):214–222, January 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000194.

Li:2017:SCB

[LGH^{+}17] Chuanyi Li, Jidong Ge, Liguo Huang, Haiyang Hu, Budan Wu, Hao Hu, and Bin Luo. Software cybernetics in BPM: modeling software behavior as feedback for evolution by a novel discovery method based on augmented event logs. The Journal of Systems and Software, 124 (??):260–273, February 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216000844.

Li:2016:RGE

[LGHR16] Zheng Li, Chunhui Guo, Xiayu Hua, and Shangping Ren. Reliability guaranteed energy minimization on mixedcriticality systems. The Journal of Systems and Software, 112(??):1–10, February 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215002289.

Li:2010:DFA

[LGL^{+}10] Wei Li, Dawu Gu, Juanru Li, Zhiqiang Liu, and Ya Liu. Differential fault analysis on Camellia. The Journal of Systems and Software, 83(5): 844–851, May 2010. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liu:2012:IRI

[LGLL12] Ya Liu, Dawu Gu, Zhiqiang Liu, and Wei Li. Improved results on impossible differential cryptanalysis of reducedround Camellia-192/256. The Journal of Systems and Software, 85 (11):2451–2458, November 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001495.

Lima:2018:MSC

[LGM^{+}18] Phyllipe Lima, Eduardo Guerra, Paulo Meirelles, Lucas Kanashiro, Hélio Silva, and Fábio Fagundes Silveira. A Metrics Suite for code annotation assessment. The Journal of Systems and Software, 137(??):163–183, March 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121730273X.

Luthmann:2019:MMD

[LGS^{+}19] Lars Luthmann, Timo Gerecht, Andreas Stephan, Johannes Bürdek, and Malte Lochau. Minimum/maximum delay testing of product lines with unbounded parametric real-time constraints. The Journal of Systems and Software, 149(??):535–553, March 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218302851.

Luo:2018:TES

[LGZ^{+}18] Yifeng Luo, Junshi Guo, Jiaye Zhu, Jihong Guan, and Shuigeng Zhou. Towards efficiently supporting database as a service with QoS guarantees. The Journal of Systems and Software, 139(??):51–63, May 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121830013X.

Lee:2011:PSE

[LH11a] Tian-Fu Lee and Tzonelih Hwang. Provably secure and efficient authentication techniques for the global mobility network. The Journal of Systems and Software, 84(10):1717–1725, October 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0164121211001099.

Lin:2011:SKM

[LH11b] Yu-Li Lin and Chien-Lung Hsu. Secure key management scheme for dynamic hierarchical access control based on ECC. The Journal of Systems and Software, 84 (4):679–685, April 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lago:2012:SIS

[LH12] Patricia Lago and Rich Hilliard. Special issue: Selected papers from the 9th Working IEEE/IFIP Conference on Software Architecture (WICSA 2011). The Journal of Systems and Software, 85 (9):1969–1970, September 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001331.

Liu:2019:RSP

[LHCT19] Jing-Wei Liu, Chia-Yu Ho, Jamie Y. T. Chang, and Jacob Chia-An Tsai. The role of Sprint planning and feedback in game development projects: Implications for game quality. The Journal of Systems and Software, 154(??):79–91, August 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300974.

Li:2018:ACU

[LHG^{+}18] Chuanyi Li, Liguo Huang, Jidong Ge, Bin Luo, and Vincent Ng. Automatically classifying user requests in crowdsourcing requirements engineering. The Journal of Systems and Software, 138(??):108–123, April 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217303096.

Li:2010:NDF

[LHH10] Jung-Shian Li, Che-Jen Hsieh, and Cheng-Fu Hung. A novel DRM framework for peer-to-peer music content delivery. The Journal of Systems and Software, 83 (10):1689–1700, October 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Langdon:2010:EMO

[LHJ10] William B. Langdon, Mark Harman, and Yue Jia. Efficient multiobjective higher order mutation testing with genetic programming. The Journal of Systems and Software, 83(12):2416– 2430, December 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lopez-Herrejon:2015:ASB

[LHLG^{+}15] Roberto E. Lopez-Herrejon, Lukas Linsbauer, José A. Galindo, José A. Parejo, David Benavides, Sergio Segura, and Alexander Egyed. An assessment of search-based techniques for reverse engineering feature models. The Journal of Systems and Software, 103(??):353– 369, May 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002349.

Loden:2009:WSS

[LHP^{+}09] Philip Loden, Qi Han, Lisa Porta, Tissa Illangasekare, and Anura P. Jayasumana. A wireless sensor system for validation of real-time automatic calibration of groundwater transport models. The Journal of Systems and Software, 82 (11):1859–1868, November 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). See corrigendum [LHP^{+}10].

Loden:2010:CWS

[LHP^{+}10] Philip Loden, Qi Han, Lisa Porta, Tissa Illangasekare, and Anura P. Jayasumana. Corrigendum to “A wireless sensor system for validation of real-time automatic calibration of groundwater transport models” [J. Syst. Software 82 (2009) 1859–1868]. The Journal of Systems and Software, 83(4):710, April 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). See [LHP^{+}09].

Lin:2012:FAH

[LHY12] Yih-Kai Lin, Shu-Chien Huang, and Cheng-Hsing Yang. A fast algorithm for Huffman decoding based on a recursion Huffman tree. The Journal of Systems and Software, 85(4):974–980, April 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0164121211002925.

Liu:2012:ESS

[LHYZ12] Yan-Xiao Liu, Lein Harn, Ching-Nung Yang, and Yu-Qing Zhang. Efficient (n, t, n) secret sharing schemes. The Journal of Systems and Software, 85(6):1325–1332, June 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000234.

Li:2012:ESC

[LHZX12] Jiguo Li, Xinyi Huang, Yichen Zhang, and Lizhong Xu. An efficient short certificate-based signature scheme. The Journal of Systems and Software, 85(2):314–322, February 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002111.

Li:2011:EID

[Li11]  Wei Li.  Evaluating

the impacts of dynamic reconfiguration on the QoS of running systems. The Journal of Systems and Software, 84 (12):2123–2138, December 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211001439.

Lin:2012:UCI

[Lin12a] Hung-Yi Lin. Using compressed index structures for processing moving objects in large spatiotemporal databases. The Journal of Systems and Software, 85(1):167–177, January 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121100207X.

Lin:2012:HCR

[Lin12b] Yih-Kai Lin. High capacity reversible data hiding scheme based upon discrete cosine transformation. The Journal of Systems and Software, 85(10):2395–2404, October 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212001446.

Lin:2014:IVW

[Lin14]  Pei-Yu Lin.  Imper-

ceptible visible watermarking based on postcamera histogram operation. The Journal of Systems and Software, 95(??):194–208, September 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001071.

Lin:2016:RDT

[Lin16] Chen-Yi Lin. A reversible data transform algorithm using integer transform for privacypreserving data mining. The Journal of Systems and Software, 117(??):104–112, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000418.

Lohre:2016:NAT

[LJ16] Erik Løhre and Magne Jørgensen. Numerical anchors and their strong effects on software development effort estimates. The Journal of Systems and Software, 116(??):49–56, June 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215000618.

Liang:2011:AQP

[LJA^{+}11] Peng Liang, Anton Jansen, Paris Avgeriou, Antony Tang, and Lai Xu. Advanced quality prediction model for software architectural knowledge sharing. The Journal of Systems and Software, 84(5): 786–802, May 2011. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lahyani:2016:ADM

[LJC16] Imene Lahyani, Mohamed Jmaiel, and Christophe Chassot. Analytical decisional model for latency aware publish/ subscribe systems on MANET. The Journal of Systems and Software, 122(??):484–495, December 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215002381.

Loulou:2010:PCB

[LJDK10] Imen Loulou, Mohamed Jmaiel, Khalil Drira, and Ahmed Hadj Kacem. P/S-CoM: Building correct by design Publish/ Subscribe architectural styles with safe reconfiguration. The Journal of Systems and Software, 83 (3):412–428, March 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lagerstrom:2010:AAE

[LJH10] Robert Lagerström, Pontus Johnson, and David Höök. Architecture analysis of enterprise systems modifiability — Models, analysis, and validation. The Journal of Systems and Software, 83(8): 1387–1403, August 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Li:2012:MFP

[LJL^{+}12] Chao-Wei Li, Kuen-Fang Jea, Ru-Ping Lin, Ssu-Fan Yen, and Chih-Wei Hsu. Mining frequent patterns from dynamic data streams with data load management. The Journal of Systems and Software, 85(6):1346–1362, June 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000209.

Liu:2011:PAI

[LJM11] Lei Liu, Xiaolong Jin, and Geyong Min. Performance analysis of an integrated scheduling scheme in the presence of bursty MMPP traffic. The Journal of Systems and Software, 84(1):37–44, January 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:2013:CNS

[LK13] Seonah Lee and Sungwon Kang. Clustering navigation sequences to create contexts for guiding code navigation. The Journal of Systems and Software, 86(8):2154–2165, August 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121300085X.

Lee:2016:WSI

[LK16] Seonah Lee and Sungwon Kang. What situational information would help developers when using a graphical code recommender? The Journal of Systems and Software, 117(??):199–217, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0164121216000819.

Lee:2010:EME

[LKJR10a] Jang-Soo Lee, Vikash Katta, Eun-Kyoung Jee, and Christian Raspotnig. Erratum to “Means-ends and whole-part traceability analysis of safety requirements” [J. Syst. Software 83 (2010) 1612– 1621]. The Journal of Systems and Software, 83(12):2664, December 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). See [LKJR10b].

Lee:2010:MEW

[LKJR10b] Jang-Soo Lee, Vikash Katta, Eun-Kyoung Jee, and Christian Raspotnig. Means-ends and wholepart traceability analysis of safety requirements. The Journal of Systems and Software, 83 (9):1612–1621, September 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). See erratum [LKJR10a].

Loukos:2014:RTD

[LKK14] Fotis Loukos, Helen Karatza, and Vana Kalogeraki. Real-time data dissemination in mobile peer-to-peer networks. The Journal of Systems and Software, 90(??):179–190, April 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002884.

Lee:2011:SEW

[LKL^{+}11] Daewook Lee, Joonho Kwon, Sangjun Lee, Seog Park, and Bonghee Hong. Scalable and efficient Web services composition based on a relational database. The Journal of Systems and Software, 84(12):2139–2155, December 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001440.

Lee:2013:IMP

[LKP13] Younho Lee, Ill-Hee Kim, and Yongsu Park. Improved multi-precision squaring for low-end RISC microcontrollers. The Journal of Systems and Software, 86(1):60– 71, January 2013. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121212002105.

Lano:2013:CBS

[LKR13] K. Lano and S. Kolahdouz-Rahimi. Constraintbased specification of model transformations. The Journal of Systems and Software, 86 (2):412–436, February 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002543.

Lano:2018:SMT

[LKRYTS18] Kevin Lano, Shekoufeh Kolahdouz-Rahimi, Sobhan Yassipour-Tehrani, and Mohammadreza Sharbaf. A survey of model transformation design patterns in practice. The Journal of Systems and Software, 140(??):48–73, June 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300438.

Lin:2010:RBR

[LL10] Jenn-Wei Lin and Huang-Yu Liu. Redirection based recovery for MPLS network systems. The Journal of Systems and Software, 83(4):609–620, April 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lin:2014:WAC

[LL14] Jenn-Wei Lin and Shian-Min Lin. A weight-aware channel assignment algorithm for mobile multicast in wireless mesh networks. The Journal of Systems and Software, 94(??):98–107, August 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214000752.

Lewis:2015:ATC

[LL15] Grace Lewis and Patricia Lago. Architectural tactics for cyberforaging: Results of a systematic literature review. The Journal of Systems and Software, 107 (??):158–186, September 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215001211.

Lee:2010:PMB

[LLC10] Hui-Lung Lee, Chia-Feng Lee, and Ling-Hwei Chen. A perfect maze based steganographic method. The Journal of Systems and Software, 83(12):2528– 2535, December 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lu:2017:AEU

[LLC17] Wei Lu, Zhe Li, and Jinghui Chu. Adaptive Ensemble Undersampling-Boost: a novel learning framework for imbalanced data. The Journal of Systems and Software, 132(??):272–282, October 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301395.

Luo:2013:RSS

[LLGZ13] Yifeng Luo, Siqiang Luo, Jihong Guan, and Shuigeng Zhou. A RAM-Cloud Storage System based on HDFS: Architecture, implementation and evaluation. The Journal of Systems and Software, 86(3):744–750, March 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212003184.

Liu:2016:SFT

[LLH^{+}16] Qin Liu, John C. S. Lui, Cheng He, Lujia Pan, Wei Fan, and Yunlong Shi. SAND: a faulttolerant streaming architecture for network traffic analytics. The Journal of Systems and Software, 122(??):553–563, December 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215001685.

Li:2019:MUB

[LLHY19] Chunlin Li, Zhu Liye, Tang Hengliang, and Luo Youlong. Mobile user behavior based topology formation and optimization in ad hoc mobile cloud. The Journal of Systems and Software, 148(??):132–147, February 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218302413.

Lin:2011:PDW

[LLK11] Chun-Han Lin, Huang-Chen Lee, and Chung-Ta King. Periphery deployment for wireless sensor systems with guaranteed coverage percentage. The Journal of Systems and Software, 84(5):763–774, May 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lochau:2014:DOM

[LLL^{+}14] Malte Lochau, Sascha Lity, Remo Lachmann, Ina Schaefer, and Ursula Goltz. Delta-oriented model-based integration testing of large-scale systems. The Journal of Systems and Software, 91(??):63–84, May 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002781.

Liu:2017:VPR

[LLL17a] Huaxiao Liu, Yuzhou Liu, and Lei Liu. The verification of program relationships in the context of software cybernetics. The Journal of Systems and Software, 124(??):212–227, February 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000224.

Liu:2017:MDK

[LLL^{+}17b] Yuzhou Liu, Lei Liu, Huaxiao Liu, Xiaoyu Wang, and Hongji Yang. Mining domain knowledge from app descriptions. The Journal of Systems and Software, 133(??):126–144, November 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217301784.

Li:2010:DCY

[LLLK10] Chengqing Li, Shujun Li, Kwok-Tung Lo, and Kyandoghere Kyamakya. A differential cryptanalysis of Yen–Chen–Wu multimedia cryptography system. The Journal of Systems and Software, 83(8):1443–1452, August 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:2012:DFS

[LLLK12] Sangho Lee, Hay-Rim Lee, Seungkwang Lee, and Jong Kim. DRMFS: a file system layer for transparent access semantics of DRM-protected contents. The Journal of Systems and Software, 85(5):1058–1066, May 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211003025.

Landwehr:2017:SSE

[LLM^{+}17] Carl Landwehr, Jochen Ludewig, Robert Meersman, David Lorge Parnas, Peretz Shoval, Yair Wand, David Weiss, and Elaine Weyuker. Software systems engineering programmes: a capability approach. The Journal of Systems and Software, 125(??):354– 364, March 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

//www.sciencedirect.
com/science/article/
pii/S0164121216302576.

Liu:2013:RDH

[LLML13] Yunxia Liu, Zhitang Li, Xiaojing Ma, and Jian Liu. A robust data hiding algorithm for H.264/AVC video streams. The Journal of Systems and Software, 86(8):2174–2183, August 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000836.

Lundell:2011:PPO

[LLS11] Björn Lundell, Brian Lings, and Anna Syberfeldt. Practitioner perceptions of Open Source software in the embedded systems area. The Journal of Systems and Software, 84 (9):1540–1549, September 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211000719.

Li:2012:ULT

[LLW12] Aikun Li, Yi Liang, and Di Wu. Utilizing Layered Taxation to provide incentives in P2P streaming systems. The Journal of Systems and Software, 85(8):1749–1756, August 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212000799.

Lin:2014:TQB

[LLWL14] Szu-Yin Lin, Chin-Hui Lai, Chih-Heng Wu, and Chi-Chun Lo. A trustworthy QoS-based collaborative filtering approach for Web service discovery. The Journal of Systems and Software, 93(??):217–228, July 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214000442.

Liu:2019:WFC

[LLWL19] Yong Liu, Meiying Li, Yonghao Wu, and Zheng Li. A weighted fuzzy classification approach to identify and manipulate coincidental correct test cases for fault localization. The Journal of Systems and Software, 151(??):20–37, May 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300196.

Liu:2011:CEM

[LLX^{+}11] Yan Liu, Xin Liang, Lingzhi Xu, Mark Staples, and Liming Zhu. Composing enterprise mashup components and services using architecture integration patterns. The Journal of Systems and Software, 84 (9):1436–1446, September 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211000227.

Liao:2014:ASS

[LLZW14] Jianxin Liao, Yang Liu, Xiaomin Zhu, and Jingyu Wang. Accurate subswarms particle swarm optimization algorithm for service composition. The Journal of Systems and Software, 90(??):191–203, April 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213002860.

Laszlo:2013:OUM

[LM13] Michael Laszlo and Sumitra Mukherjee. Optimal univariate microaggregation with data suppression. The Journal of Systems and Software, 86(3):677–682, March 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212003019.

Laszlo:2015:ILS

[LM15] Michael Laszlo and Sumitra Mukherjee. Iterated local search for microaggregation. The Journal of Systems and Software, 100(??):15–26, February 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002167.

Lopez-Martin:2015:NNP

[LMA15] Cuauhtémoc López-Martín and Alain Abran. Neural networks for predicting the duration of new software projects. The Journal of Systems and Software, 101(??):127–135, March 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002805.

Leshob:2017:VOA

[LMGHB17] Abderrahmane Leshob, Hafedh Mili, Javier Gonzalez-Huerta, and Anis Boubaker. A value-oriented approach to business process specialization: Principles, proof-of-concept, and validation. The Journal of Systems and Software, 127(??):120–149, May 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121217300316.

Lakhotia:2010:EIB

[LMH10] Kiran Lakhotia, Phil McMinn, and Mark Harman. An empirical investigation into branch coverage for C programs using CUTE and AUSTIN. The Journal of Systems and Software, 83 (12):2379–2391, December 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lehtinen:2015:DSL

[LMIV15] Timo O. A. Lehtinen, Mika V. Mäntylä, Juha Itkonen, and Jari Vanhanen. Diagrams or structural lists in software project retrospectives — an experimental comparison. The Journal of Systems and Software, 103(??):17–35, May 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215000151.

Lee:2010:FOA

[LMN10] Jaejoon Lee, Dirk Muthig, and Matthias Naab. A feature-oriented approach for developing reusable product line assets of service-based systems. The Journal of Systems and Software, 83(7):1123–1136, July 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lopez-Martin:2017:TPI

[LMNA17] Cuauhtémoc López-Martín, Ali Bou Nassif, and Alain Abran. A training process for improving the quality of software projects developed by a practitioner. The Journal of Systems and Software, 131(??):98–111, September 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217300973.

LimeiradeLimaJunior:2018:AAI

[LMPM18] Manoel Limeira de Lima Júnior, Daricélio Moreira Soares, Alexandre Plastino, and Leonardo Murta. Automatic assignment of integrators to pull requests: the importance of selecting appropriate attributes. The Journal of Systems and Software, 144 (??):181–196, October 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218301122.

Lepmets:2012:GAP

[LMR12] Marion Lepmets, Tom McBride, and Eric Ras. Goal alignment in process improvement. The Journal of Systems and Software, 85(6):1440– 1452, June 2012. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121200026X.

Larrea:2011:CEL

[LMS11] Mikel Larrea, Cristian Martín, and Iratxe Soraluze. Communicationefficient leader election in crash-recovery systems. The Journal of Systems and Software, 84 (12):2186–2195, December 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211001476.

Lo:2012:LEF

[LMS12] David Lo, Leonardo Mariani, and Mauro Santoro. Learning extended FSA from software: an empirical assessment. The Journal of Systems and Software, 85 (9):2063–2076, September 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001008.

Laatikainen:2016:CBF

[LMT16] Gabriella Laatikainen, Oleksiy Mazhelis, and Pasi Tyrvainen. Cost benefits of flexible hybrid cloud storage: Mitigating volume variation with shorter acquisition cycle. The Journal of Systems and Software, 122(??):180–201, December 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301613.

Linaaker:2018:MCO

[LMWM18] J. Linåker, H. Munir, K. Wnuk, and C. E. Mols. Motivating the contributions: an Open Innovation perspective on what to share as Open Source Software. The Journal of Systems and Software, 135(??):17–36, January 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217302169.

Lucena:2013:CEC

[LN13] Carlos Lucena and Ingrid Nunes. Contributions to the emergence and consolidation of agent-oriented software engineering. The Journal of Systems and Software, 86(4):890–904, April 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212002567.

Lity:2019:RTS

[LNTS19] Sascha Lity, Manuel Nieke, Thomas Thüm, and Ina Schaefer. Retest test selection for productline regression testing of variants and versions of variants. The Journal of Systems and Software, 147(??):46–63, January 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218302176.

Liu:2011:NGF

[LNW^{+}11] Xiao Liu, Zhiwei Ni, Zhangjun Wu, Dong Yuan, Jinjun Chen, and Yun Yang. A novel general framework for automatic and cost-effective handling of recoverable temporal violations in scientific workflow systems. The Journal of Systems and Software, 84 (3):492–509, March 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liu:2011:NST

[LNY^{+}11] Xiao Liu, Zhiwei Ni, Dong Yuan, Yuanchun Jiang, Zhangjun Wu, Jinjun Chen, and Yun Yang. A novel statistical time-series pattern based interval forecasting strategy for activity durations in workflow systems. The Journal of Systems and Software, 84(3):354–376, March 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lucas:2017:CLC

[LOFA17] Edson M. Lucas, Toacy C.

Oliveira, Kleinner Farias, and Paulo S. C. Alencar. CollabRDL: a language to coordinate collaborative reuse. The Journal of Systems and Software, 131(??):505–527, September 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217300225.

Luz:2019:ADR

[LPB19] Welder Pinheiro Luz, Gustavo Pinto, and Rodrigo Bonifácio. Adopting DevOps in the real world: a theory, a model, and a case study. The Journal of Systems and Software, 157(??):??, November 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301517.

Leopold:2015:ASD

[LPM15] Henrik Leopold, Fabian Pittke, and Jan Mendling. Automatic service derivation from business process model repositories via semantic technology. The Journal of Systems and Software, 108(??):134–147, October 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215001235.

Lee:2010:IQP

[LPP^{+}10] Jihyun Lee, Jeong-Hoon Park, Myung-Jae Park, Chin-Wan Chung, and Jun-Ki Min. An intelligent query processing for distributed ontologies. The Journal of Systems and Software, 83 (1):85–95, January 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lamancha:2015:PPA

[LPP15] Beatriz Pérez Lamancha, Macario Polo, and Mario Piattini. PROW: a Pairwise algorithm with constRaints, Order and Weight. The Journal of Systems and Software, 99(??):1–19, January 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001733.

Li:2010:EAD

[LPXL10] Juncao Li, Nicholas T.

Pilkington, Fei Xie, and Qiang Liu. Embedded architecture description language. The Journal of Systems and Software, 83(2):235–252, February 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lu:2014:SNR

[LQC^{+}14] Meilian Lu, Zhen Qin, Yiming Cao, Zhichao Liu, and Mengxing Wang.

Scalable news recommendation using multidimensional similarity and Jaccard–Kmeans clustering. The Journal of Systems and Software, 95(??):242–251, September 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001162.

Li:2016:AQC

[LQLC16] Li Li, Jinxia Qiu, Jianfeng Lu, and Chin-Chen Chang. An aesthetic QR code solution based on error correction mechanism. The Journal of Systems and Software, 116(??):85–94, June 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121500148X.

Li:2012:ATC

[LQLW12] Bixin Li, Dong Qiu, Hareton Leung, and Di Wang. Automatic test case selection for regression testing of composite service based on extensible BPEL flow graph. The Journal of Systems and Software, 85(6):1300–1324, June 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000258.

Lin:2012:OVM

[LQW^{+}12] Qian Lin, Zhengwei Qi, Jiewei Wu, Yaozu Dong, and Haibing Guan. Optimizing virtual machines using hybrid virtualization. The Journal of Systems and Software, 85 (11):2593–2603, November 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001677.

Li:2019:RIC

[LRB^{+}19] Li Li, Timothée Riom, Tegawendé F. Bissyandé, Haoyu Wang, Jacques Klein, and Le Traon Yves. Revisiting the impact of common libraries for Android-related investigations. The Journal of Systems and Software, 154(??):157–175, August 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301001.

Liu:2019:WFA

[LRD^{+}19] Bohan Liu, Guoping Rong, Liming Dong, He Zhang, Danni Chen, Tiange Chen, Yuyan Chen, and Tiantian Zhang. What are the factors affecting the handover process in open source development? The Journal of Systems and Software, 153(??):238–254, July 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219300718.

Lagartos:2019:ERM

[LRO19] Ignacio Lagartos, Jose Manuel Redondo, and Francisco Ortin. Efficient runtime metaprogramming services for Java. The Journal of Systems and Software, 153(??):220– 237, July 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219300792.

Lee:2014:DBS

[LS14] Jinkyu Lee and Insik Shin. Demand-based schedulability analysis for real-time multi-core scheduling. The Journal of Systems and Software, 89(??):99–108, March 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002367.

Latorre:2017:MSN

[LS17a] Roberto Latorre and Javier Suárez. Measuring social networks when forming information system project teams. The Journal of Systems and Software, 134(??):304–323, December 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121730208X.

Lee:2017:DUN

[LS17b] Jinkyu Lee and Kang G.

Shin. Development and use of a new task model for cyberphysical systems: a real-time scheduling perspective. The Journal of Systems and Software, 126(??):45–56, April 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121217300043.

Lindsjorn:2016:TQP

[LSD^{+}16] Yngve Lindsjørn, Dag I. K. Sjøberg, Torgeir Dingsøyr, Gunnar R. Bergersen, and Tore Dybå. Teamwork quality and project success in software development: a survey of agile development teams. The Journal of Systems and Software, 122(??):274–286, December 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121630187X.

Lee:2012:COF

[LSE12] Jinkyu Lee, Insik Shin, and Arvind Easwaran. Convex optimization framework for intermediate deadline assignment in soft and hard realtime distributed systems. The Journal of Systems and Software, 85(10):2331–2339, October 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212001100.

Lizcano:2017:AVV

[LSLG17] David Lizcano, Javier Soriano, Genoveva López, and Javier J. Gutiérrez. Automatic verification and validation wizard in web-centred enduser software engineering. The Journal of Systems and Software, 125 (??):47–67, March 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

//www.sciencedirect.
com/science/article/
pii/S0164121216302278.

Lima:2019:HEE

[LSNL^{+}19] Luís Gabriel Lima, Francisco Soares-Neto, Paulo Lieuthier, Fernando Castor, Gilberto Melfe, and João Paulo Fernandes. On Haskell and energy efficiency. The Journal of Systems and Software, 149(??):554–580, March 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302747.

Lei:2013:RSW

[LSR13] Baiying Lei, Insu Song, and Shah Atiqur Rahman. Robust and secure watermarking scheme for breath sound. The Journal of Systems and Software, 86(6):1638–1649, June 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000332.

Lin:2011:GBC

[LT11] Jenn-Wei Lin and Shih-Chieh Tang. A gridbased coverage approach for target tracking in hybrid sensor networks. The Journal of Systems and Software, 84 (10):1746–1756, October 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001129.

Lee:2013:CCM

[LT13] Che-Wei Lee and Wen-Hsiang Tsai. A covert communication method via spreadsheets by secret sharing with a selfauthentication capability. The Journal of Systems and Software, 86(2):324–334, February 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002464.

Lin:2015:LDR

[LTK^{+}15] Wei-Chao Lin, Chih-Fong Tsai, Shih-Wen Ke, Chia-Wen Hung, and William Eberle. Learning to detect representative data for large scale instance selection. The Journal of Systems and Software, 106(??):1–8, August 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000813.

Lu:2016:AHB

[LTW16] Tzu-Chuen Lu, Chun-Ya Tseng, and Jhih-Huei Wu. Asymmetrichistogram based reversible information hiding scheme using edge sensitivity detection. The Journal of Systems and Software, 116(??):2–21, June 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000990.

Luk:2011:SSS

[Luk11] Robert W. P. Luk. Scalable, statistical storage allocation for extensible inverted file construction. The Journal of Systems and Software, 84(7):

1082–1088, July 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lopez-Vega:2013:CAB

[LVPMPCLS13] Jose M. Lopez-Vega, Javier Povedano-Molina, Gerardo Pardo-Castellote, and Juan M. Lopez-Soler. A content-aware bridging service for publish/subscribe environments. The Journal of Systems and Software, 86(1):108–124, January 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212002142.

Linares-Vasquez:2017:HDM

[LVVTP17] Mario Linares-Vásquez, Christopher Vendome, Michele Tufano, and Denys Poshyvanyk. How developers micro-optimize Android apps. The Journal of Systems and Software, 130(??):1–23, August 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121730081X.

Liu:2013:TIE

[LW13a] Hongjun Liu and Xingyuan Wang. Triple-image encryption scheme based on one-time key stream generated by chaos and plain images. The Journal of Systems and Software, 86(3):826–834, March 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212003196.

Liu:2013:CFP

[LW13b] Ying-Ho Liu and Chun-Sheng Wang. Constrained frequent pattern mining on univariate uncertain data. The Journal of Systems and Software, 86(3):759–778, March 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212003135.

Lui:2013:CBS

[LW13c] Oi-Yan Lui and Kwok-Wo Wong. Chaos-based selective encryption for H.264/AVC. The Journal of Systems and Software, 86(12):3183–3192, December 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121213001994.

Langer:2013:POD

[LWB^{+}13] Philip Langer, Manuel Wimmer, Petra Brosch, Markus Herrmannsdörfer, Martina Seidl, Konrad Wieland, and Gerti Kappel. A posteriori operation detection in evolving software models. The Journal of Systems and Software, 86(2):551–566, February 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212002762.

Liu:2016:PMP

[LWBH16] Hu-Qiu Liu, Yu-Ping Wang, Jia-Ju Bai, and Shi-Min Hu. PF-Miner: a practical paired functions mining method for Android kernel in error paths. The Journal of Systems and Software, 121(??):234–246, November 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216000431.

Lin:2013:EVL

[LWC13] Qiuzhen Lin, Kwok-Wo Wong, and Jianyong Chen. An enhanced variable-length arithmetic coding and encryption scheme using chaotic maps. The Journal of Systems and Software, 86(5):1384–1389, May 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000046.

Lin:2018:HCS

[LWC^{+}18] Chi Lin, Kang Wang, Zihao Chu, Kai Wang, Jing Deng, Mohammad S. Obaidat, and Guowei Wu. Hybrid charging scheduling schemes for three-dimensional underwater wireless rechargeable sensor networks. The Journal of Systems and Software, 146 (??):42–58, December 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301766.

Li:2013:RGE

[LWL^{+}13] Zheng Li, Li Wang, Shuhui Li, Shangping Ren, and Gang Quan. Reliability guaranteed energy-aware frame-based task set execution strategy for hard real-time systems. The Journal of Systems and Software, 86 (12):3060–3070, December 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213001787.

Lin:2016:GGT

[LWL^{+}16] Chi Lin, Youkun Wu, Zhicheng Liu, Mohammad S. Obaidat, Chang Wu Yu, and Guowei Wu. GTCharge: a game theoretical collaborative charging scheme for wireless rechargeable sensor networks. The Journal of Systems and Software, 121(??):88–104, November 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216301480.

Li:2012:PAP

[LWLL12] Bixin Li, Lulu Wang, Hareton Leung, and Fei Liu. Profiling all paths: a new profiling technique for both cyclic and acyclic paths. The Journal of Systems and Software, 85(7):1558–1576, July 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000349.

Lin:2016:CSC

[LWOY16] Chi Lin, Guowei Wu, Mohammad S. Obaidat, and Chang Wu Yu. Clustering and splitting charging algorithms for large scaled wireless rechargeable sensor networks. The Journal of Systems and Software, 113(??):381–394, March 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002836.

Leitner:2019:MME

[LWSH19] Philipp Leitner, Erik Wittern, Josef Spillner, and Waldemar Hummer. A mixed-method empirical study of Functionas-a-Service software development in industrial practice. The Journal of Systems and Software, 149(??):340–359, March 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302735.

Lou:2010:NAS

[LWW^{+}10] Der-Chyuan Lou, Nan-I Wu, Chung-Ming Wang, Zong-Han Lin, and Chwei-Shyong Tsai. A novel adaptive steganography based on local complexity and human vision sensitivity. The Journal of Systems and Software, 83(7):1236– 1248, July 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lin:2010:NXK

[LWXZ10] Xudong Lin, Ning Wang, De Xu, and Xiaoning Zeng. A novel XML keyword query approach using entity subtree. The Journal of Systems and Software, 83(6):990– 1003, June 2010. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liu:2012:TFE

[LWZ12] Manlu Liu, Harry Jiannan Wang, and J. Leon Zhao. Technology flexibility as enabler of robust application development in community source: the case of Kuali and Sakai. The Journal of Systems and Software, 85(12):2921–2928, December 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001744.

Lin:2016:EQD

[LWZ^{+}16] Yiming Lin, Hongzhi Wang, Shuo Zhang, Jianzhong Li, and Hong Gao. Efficient qualitydriven source selection from massive data sources. The Journal of Systems and Software, 118 (??):221–233, August 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300577.

Liu:2013:AEM

[LXC13] Yepang Liu, Chang Xu, and S. C. Cheung. AFChecker: Effective model checking for context-aware adaptive applications. The Journal of Systems and Software, 86(3):854–867, March 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212003287.

Li:2011:NIW

[LXCM11] Li Li, He-Huan Xu, Chin-Chen Chang, and Ying-Ying Ma. A novel image watermarking in redistributed invariant wavelet domain. The Journal of Systems and Software, 84(6):923–929, June 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Li:2010:ARR

[LXG10] Yan-Fu Li, Min Xie, and Thong-Ngee Goh. Adaptive ridge regression system for software cost estimating on multi-collinear datasets. The Journal of Systems and Software, 83 (11):2332–2343, November 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Li:2010:PBU

[LXJL10] Nuo Li, Tao Xie, Maozhong Jin, and Chao Liu. Perturbation-based userinput-validation testing of web applications. The Journal of Systems and Software, 83(11):2263– 2274, November 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lu:2018:NTE

[LY18] Lilei Lu and Yuyu Yuan. A novel TOP-SIS evaluation scheme for cloud service trustworthiness combining objective and subjective aspects. The Journal of Systems and Software, 143(??):71–86, September 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121830092X.

Lv:2014:ECI

[LYC14] Junpeng Lv, Bei-Bei Yin, and Kai-Yuan Cai. Estimating confidence interval of software reliability with adaptive testing strategy. The Journal of Systems and Software, 97(??):192–206, November 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001721.

Lin:2016:EDD

[LYLC16] Meng-Jie Lin, Cheng-Zen Yang, Chao-Yuan Lee, and Chun-Chang Chen. Enhancements for duplication detection in bug reports with manifold correlation features. The Journal of Systems and Software, 121(??):223–233, November 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216000546.

Liao:2010:MPC

[LyWSZ10] Xin Liao, Qiao yan Wen, Ying Sun, and Jie Zhang. Multi-party covert communication with steganography and quantum secret sharing. The Journal of Systems and Software, 83 (10):1801–1804, October 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liu:2012:NDE

[LZ12] Huawen Liu and Shichao Zhang. Noisy data elimination using mutual knearest neighbor for classification mining. The Journal of Systems and Software, 85(5):1067– 1074, May 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211003049.

Lin:2013:DCP

[LZ13] Jenn-Wei Lin and Jian-Yan Zhuang. A delayconstrained and priorityaware channel assignment algorithm for efficient multicast in wireless mesh networks. The Journal of Systems and Software, 86(3):789–800, March 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212003111.

Long:2014:TPE

[LZC14] Saiqin Long, Yuelong Zhao, and Wei Chen. A three-phase energysaving strategy for cloud storage systems. The Journal of Systems and Software, 87(??):38–47, January 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002148.

Li:2019:DPF

[LZCL19] Chunlin Li, Jing Zhang, Yi Chen, and Youlong Luo. Data prefetching and file synchronizing for performance optimization in Hadoopbased hybrid cloud. The Journal of Systems and Software, 151(??):133– 149, May 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121219300287.

Luo:2015:LCE

[LZG15] Yifeng Luo, Shuigeng Zhou, and Jihong Guan. LAYER: a cost-efficient mechanism to support multi-tenant database as a service in cloud. The Journal of Systems and Software, 101(??):86–96, March 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214002696.

Liu:2011:STM

[LZHS11] Chang Liu, Qing Zhu, Kenneth A. Holroyd, and Elizabeth K. Seng. Status and trends of mobile-health applications for iOS devices: a developer’s perspective. The Journal of Systems and Software, 84 (11):2022–2033, November 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211001610.

Li:2019:DDA

[LZJ^{+}19] Shanshan Li, He Zhang, Zijia Jia, Zheng Li, Cheng Zhang, Jiaqi Li, Qiuya Gao, Jidong Ge, and Zhihao Shan. A dataflow-driven approach to identifying microservices from monolithic applications. The Journal of Systems and Software, 157(??):??, November 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301475.

Liu:2012:IVB

[LZKW12] Qixu Liu, Yuqing Zhang, Ying Kong, and Qianru Wu. Improving VRSSbased vulnerability prioritization using analytic hierarchy process. The Journal of Systems and Software, 85 (8):1699–1708, August 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000854.

Lei:2015:SSG

[LZL^{+}15] Hongtao Lei, Tao Zhang, Yajie Liu, Yabing Zha, and Xiaomin Zhu. SGEESS: Smart Green Energy-Efficient Scheduling Strategy with dynamic electricity price for data center. The Journal of Systems and Software, 108(??):23–38, October 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215001284.

Li:2018:AAH

[LZL^{+}18] Bodong Li, Yuanyuan Zhang, Juanru Li, Wenbo Yang, and Dawu Gu. AppSpear: Automating the hidden-code extraction and reassembling of packed Android malware. The Journal of Systems and Software, 140(??):3– 16, June 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218300311.

Liu:2017:RCE

[LZLC17] Lin Liu, Qing Zhou, Jilei Liu, and Zhanqiang Cao. Requirements cybernetics: Elicitation based on user behavioral data. The Journal of Systems and Software, 124(??):187–194, February 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0164121215002927.

Li:2013:ECC

[LZO^{+}13] Zheng Li, He Zhang, Liam O’Brien, Rainbow Cai, and Shayne Flint. On evaluating commercial Cloud services: a systematic review. The Journal of Systems and Software, 86 (9):2371–2393, September 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000915.

Li:2016:SPC

[LZO^{+}16] Zheng Li, He Zhang, Liam O’Brien, Shu Jiang, You Zhou, Maria Kihl, and Rajiv Ranjan. Spot pricing in the Cloud ecosystem: a comparative investigation. The Journal of Systems and Software, 114(??):1–19, April 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002332.

Lung:2016:ISP

[LZR16] Chung-Horng Lung, Xu Zhang, and Pragash Rajeswaran. Improving software performance and reliability in a distributed and concurrent environment with an architecturebased self-adaptive framework. The Journal of Systems and Software, 121(??):311–328, November 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300991.

Liu:2015:SPJ

[LZY^{+}15] Xiaocheng Liu, Yabing Zha, Quanjun Yin, Yong Peng, and Long Qin. Scheduling parallel jobs with tentative runs and consolidation in the cloud. The Journal of Systems and Software, 104(??):141–151, June 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000588.

Mittas:2010:VCS

[MA10] Nikolaos Mittas and Lefteris Angelis. Visual comparison of software cost estimation models by regression error characteristic analysis. The Journal of Systems and Software, 83(4):621–637, April 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mohammad:2011:FAS

[MA11] Mubarak Mohammad and Vangalur Alagar. A formal approach for the specification and verification of trustworthy component-based systems. The Journal of Systems and Software, 84(1):77–104, January 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Moussa:2017:PGA

[MA17] Rebecca Moussa and Danielle Azar. A PSO-GA approach targeting fault-prone software modules. The Journal of Systems and Software, 132(??):41–49, October 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301309.

Mufti:2017:FDS

[MAAC17] Muhammad Rafiq Mufti, Humaira Afzal, Irfan Awan, and Andrea Cullen. A framework for dynamic selection of backoff stages during initial ranging process in wireless networks. The Journal of Systems and Software, 133(??):17–27, November 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121730167X.

Mendes:2019:SIE

[MACB19] Emilia Mendes, Nauman Bin Ali, Steve Counsell, and Maria Teresa Baldassare. Special issue on evaluation and assessment in software engineering. The Journal of Systems and Software, 151(??):224–225, May 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300202.

Mendonca:2019:DRS

[MAEL19] Júlio Mendonça, Ermeson Andrade, Patricia Takako Endo, and Ricardo Lima. Disaster recovery solutions for IT systems: a systematic mapping study. The Journal of Systems and Software, 149(??):511– 530, March 2019. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121218302814.

Meedeniya:2012:ADR

[MAG12] Indika Meedeniya, Aldeida Aleti, and Lars Grunske. Architecture-driven reliability optimization with uncertain model parameters. The Journal of Systems and Software, 85(10):2340–2355, October 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001276.

Mendez-Acuna:2017:REL

[MAGC^{+}17] David Méndez-Acuña, José A. Galindo, Benoît Combemale, Arnaud Blouin, and Benoît Baudry. Reverse engineering language product lines from existing DSL variants. The Journal of Systems and Software, 133(??):145–158, November 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217300857.

Manteuffel:2018:ECS

[MAH18] Christian Manteuffel, Paris Avgeriou, and Roelof Hamberg. An exploratory case study on reusing architecture decisions in softwareintensive system projects. The Journal of Systems and Software, 144(??): 60–83, October 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121218301110.

Manikas:2016:RSE

[Man16] Konstantinos Manikas. Revisiting software ecosystems research: a longitudinal literature study. The Journal of Systems and Software, 117(??): 84–103, July 2016. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121216000406.

Makitalo:2019:AOP

[MAR^{+}19] Niko Mäkitalo, Timo Aaltonen, Mikko Raatikainen, Aleksandr Ometov, Sergey Andreev, Yevgeni Koucheryavy, and Tommi Mikkonen. Actionoriented programming model: Collective executions and interactions in the fog. The Journal of Systems and Software, 157(??):??, November 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301669.

Moustakas:2016:ATM

[MARD16] Vassilis Moustakas, Hüseyin Akcan, Mema Roussopoulos, and Alex Delis. Alleviating the topology mismatch problem in distributed overlay networks: a survey. The Journal of Systems and Software, 113(??):216–245, March 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002629.

Monteiro:2013:VWS

[MAS13] Andre Felipe Monteiro, Marcus Vinicius Azevedo, and Alexandre Sztajnberg. Virtualized Web server cluster selfconfiguration to optimize resource and power use. The Journal of Systems and Software, 86 (11):2779–2796, November 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121213001453.

Miranda:2010:AMU

[MB10] Eduardo Miranda and Pierre Bourque. Agile monitoring using the line of balance. The Journal of Systems and Software, 83(7):1205–1215, July 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Miranda:2017:SAT

[MB17] Breno Miranda and Antonia Bertolino. Scopeaided test prioritization, selection and minimization for software reuse. The Journal of Systems and Software, 131 (??):528–549, September 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216300875.

Menasce:2019:TTD

[MB19] Daniel A. Menascé and Shouvik Bardhan. TDQN: Trace-driven analytic queuing network modeling of computer systems. The Journal of Systems and Software, 147(??):162–171, January 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302346.

Meedeniya:2011:RDD

[MBAG11] Indika Meedeniya, Barbora Buhnova, Aldeida Aleti, and Lars Grunske. Reliabilitydriven deployment optimization for embedded systems. The Journal of Systems and Software, 84 (5):835–846, May 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Milo:2011:FGB

[MBB11] Fabrizio Milo, Massimo Bernaschi, and Mauro Bisson. A fast, GPU based, dictionary attack to OpenPGP secret keyrings. The Journal of Systems and Software, 84(12):2088–2096, December 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001270.

Martinez:2013:DCB

[MBD13] Patricia López Martínez, Laura Barros, and José M. Drake. Design of component-based realtime applications. The Journal of Systems and Software, 86(2):449–467, February 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002750.

Magazinius:2012:IID

[MBF12] Ana Magazinius, Sofia Börjesson, and Robert Feldt. Investigating intentional distortions in software cost estimation — an exploratory study. The Journal of Systems and Software, 85(8):1770–1781, August 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000763.

Mian:2019:MTA

[MBPM19] Zhibao Mian, Leonardo Bottaci, Yiannis Papadopoulos, and Nidhal Mahmud. Model transformation for analyzing dependability of AADL model by using HiP-HOPS. The Journal of Systems and Software, 151(??):258–282, May 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0164121219300299.

Mohamed:2016:EOA

[MBT16]  Mohamed  Mohamed,

Djamel Belaïd, and Samir Tata. Extending OCCI for autonomic management in the cloud. The Journal of Systems and Software, 122(??):416–429, December 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000030.

Min:2010:EED

[MC10] Jun-Ki Min and Chin-Wan Chung. EDGES: Efficient data gathering in sensor networks using temporal and spatial correlations. The Journal of Systems and Software, 83(2):271–282, February 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ma:2011:LSB

[MCC11] Yung-Cheng Ma, Chung-Ping Chung, and Tien-Fu Chen. Load and storage balanced posting file partitioning for parallel information retrieval. The Journal of Systems and Software, 84(5):864–884, May 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ma:2018:NDR

[MCC^{+}18] Yi-Wei Ma, Jiann-Liang Chen, Chen-Chia Chang, Akihiro Nakao, and Shu Yamamoto. A novel dynamic resource adjustment architecture for virtual tenant networks in SDN. The Journal of Systems and Software, 143(??):100–115, September 2018. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218300712.

Mao:2017:SUC

[MCHJ17] Ke Mao, Licia Capra, Mark Harman, and Yue Jia. A survey of the use of crowdsourcing in software engineering. The Journal of Systems and Software, 126(??):57–84, April 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216301832.

Morales:2018:ERS

[MCKA18] Rodrigo Morales, Francisco Chicano, Foutse Khomh, and Giuliano Antoniol. Efficient refactoring scheduling based on partial order reduction. The Journal of Systems and Software, 145(??):25–51, November 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301523.

Ma:2017:LQO

[MCL^{+}17] Jiaoyang Ma, Ling Chen, Mingqi Lv, Yi Yang, Yuliang Zhao, Yong Wu, and Jingchang Wang. Logical query optimization for Cloudera Impala system. The Journal of Systems and Software, 125(??):35– 46, March 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

//www.sciencedirect.
com/science/article/
pii/S0164121216302400.

Mohanraj:2012:ODB

[MCS^{+}12] V. Mohanraj, M. Chandrasekaran, J. Senthilkumar, S. Arumugam, and Y. Suresh. Ontology driven bee’s foraging approach based self adaptive online recommendation system. The Journal of Systems and Software, 85(11):2439–2450, November 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211003037.

Mei:2011:XMT

[MCTM11] Lijun Mei, W. K. Chan, T. H. Tse, and Robert G. Merkel. XML-manipulating test case prioritization for XML-manipulating services. The Journal of Systems and Software, 84 (4):603–619, April 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mulfari:2015:CSA

[MCV15] Davide Mulfari, Antonio Celesti, and Massimo Villari. A computer system architecture providing a user-friendly man machine interface for accessing assistive technology in cloud computing. The Journal of Systems and Software, 100(??):129–138, February 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214002325.

Mariani:2016:PAS

[MCV16] Thainá Mariani, Thelma Elita Colanzi, and Silvia Regina Vergilio. Preserving architectural styles in the search based design of software product line architectures. The Journal of Systems and Software, 115(??):157–173, May 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000364.

Montalvillo:2016:RDE

[MD16] Leticia Montalvillo and Oscar Díaz. Requirementdriven evolution in software product lines: a systematic mapping study. The Journal of Systems and Software, 122 (??):110–143, December 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301510.

Meade:2017:ESD

[MDBC17] Anne Meade, Deva Kumar Deeptimahanti, Jim Buckley, and J. J. Collins. An empirical study of data decomposition for software parallelization. The Journal of Systems and Software, 125(??):401– 416, March 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121600039X.

Moeyersoms:2015:CSF

[MdFD^{+}15] Julie Moeyersoms, Enric Junqué de Fortuny, Karel Dejaeger, Bart Baesens, and David Martens. Comprehensible software fault and effort prediction: a data mining approach. The Journal of Systems and Software, 100(??):80–90, February 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214002295.

Misra:2010:ASI

[MDO^{+}10] Sudip Misra, Sanjay K. Dhurandher, Mohammad S. Obaidat, Pushkar Gupta, Karan Verma, and Prayag Narula. An ant swarm-inspired energy-aware routing protocol for wireless adhoc networks. The Journal of Systems and Software, 83(11):2188–2199, November 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Magdaleno:2015:COS

[MdOBW^{+}15] Andréa Magalhães Magdaleno, Marcio de Oliveira Barros, Cláudia Maria Lima Werner, Renata Mendes de Araujo, and Carlos Freud Alves Batista. Collaboration optimization in software process composition. The Journal of Systems and Software, 103(??):452–466, May 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002672.

Martin:2011:SAF

[MDP^{+}11] Sergio Martin, Gabriel Diaz, Inmaculada Plaza, Elena Ruiz, Manuel Castro, and Juan Peire. State of the art of frameworks and middleware for facilitating mobile and ubiquitous learning development. The Journal of Systems and Software, 84(11):1883–1891, November 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001543.

Medvidovic:2010:SAM

[ME10] Nenad Medvidovic and George Edwards. Software architecture and mobility: a roadmap. The Journal of Systems and Software, 83(6):885– 898, June 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Malek:2010:ADS

[MEB^{+}10] Sam Malek, George Edwards, Yuriy Brun, Hossein Tajalli, Joshua Garcia, Ivo Krka, Nenad Medvidovic, Marija Mikic-Rakic, and Gaurav S. Sukhatme. An architecture-driven software mobility framework. The Journal of Systems and Software, 83 (6):972–989, June 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mernik:2013:OOA

[Mer13]  Marjan Mernik.  An

object-oriented approach to language compositions for software language engineering. The Journal of Systems and Software, 86 (9):2451–2464, September 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213001271.

Maalej:2017:UCS

[MER17] Walid Maalej, Mathias Ellmann, and Romain Robbes. Using contexts similarity to predict relationships between tasks. The Journal of Systems and Software, 128(??):267–284, June 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216302357.

Miller:2010:ESA

[MFM10] James A. Miller, Remo Ferrari, and Nazim H. Madhavji. An exploratory study of architectural effects on requirements decisions. The Journal of Systems and Software, 83 (12):2441–2455, December 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mattiello-Francisco:2012:IAT

[MFMCY12] Fátima Mattiello-Francisco, Eliane Martins, Ana Rosa Cavalli, and Edgar Toshiro Yano. InRob: an approach for testing interoperability and robustness of real-time embedded software. The Journal of Systems and Software, 85(1):3–15, January 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211000550.

Marculescu:2018:TIS

[MFTP18] Bogdan Marculescu, Robert Feldt, Richard Torkar, and Simon Poulding. Transferring interactive search-based software testing to industry. The Journal of Systems and Software, 142(??):156– 170, August 2018. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218300852.

Meged:2011:AFS

[MG11] Avichai Meged and Roy Gelbard. Adjusting Fuzzy Similarity Functions for use with standard data mining tools. The Journal of Systems and Software, 84 (12):2374–2383, December 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211001841.

Mader:2012:TAT

[MG12] Patrick Mäder and Orlena Gotel. Towards automated traceability maintenance. The Journal of Systems and Software, 85(10):2205–2227, October 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002779.

Merino:2018:SLR

[MGAN18] L. Merino, M. Ghafari, C. Anslow, and O. Nierstrasz. A systematic literature review of software visualization evaluation. The Journal of Systems and Software, 144(??):165–180, October 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301237.

Mastelic:2016:TUM

[MGB16] Toni Mastelić, Andrés García García, and Ivona Brandić. Towards uniform management of multi-layered cloud services by applying model-driven development. The Journal of Systems and Software, 121(??):358–371, November 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216000029.

Masood:2010:FCC

[MGM10] Ammar Masood, Arif Ghafoor, and Aditya P. Mathur. Fault coverage of Constrained Random Test Selection for access control: a formal analysis. The Journal of Systems and Software, 83 (12):2607–2617, December 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mohan:2016:TDR

[MGM16] Michael Mohan, Des Greer, and Paul Mc-Mullan. Technical debt reduction using search based automated refactoring. The Journal of Systems and Software, 120(??):183–194, October 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216300541.

Molina:2013:MDD

[MGR^{+}13] Ana I. Molina, Jesús Gallardo, Miguel A. Redondo, Manuel Ortega, and William J. Giraldo. Metamodel-driven definition of a visual modeling language for specifying interactive groupware applications: an empirical study. The Journal of Systems and Software, 86(7):1772–1789, July 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121200221X.

Molesini:2010:SAA

[MGvFGCB10] Ambra Molesini, Alessandro Garcia, Christina von Flach Garcia Chavez, and Thais Vasconcelos Batista. Stability assessment of aspect-oriented software architectures: a quantitative study. The Journal of Systems and Software, 83(5):711–722, May 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Masri:2011:ACV

[MH11] Wes Masri and Hiba Halabi. An algorithm for capturing variables dependences in test suites. The Journal of Systems and Software, 84(7): 1171–1190, July 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mahnic:2012:UPP

[MH12] Viljan Mahnic and Tomaz Hovelja. On using planning poker for estimating user stories. The Journal of Systems and Software, 85 (9):2086–2095, September 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212001021.

Manikas:2013:SES

[MH13] Konstantinos Manikas and Klaus Marius Hansen. Software ecosystems — a systematic literature review. The Journal of Systems and Software, 86(5):1294–1306, May 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121200338X.

Masood:2018:AAP

[MHB18] Zainab Masood, Rashina Hoda, and Kelly Blincoe. Adapting agile practices in university contexts. The Journal of Systems and Software, 144(??):501–510, October 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301419.

Murtaza:2014:ESU

[MHLMG14] Syed Shariyar Murtaza, Abdelwahab Hamou-Lhadj, Nazim H. Madhavji, and Mechelle Gittens. An empirical study on the use of mutant traces for diagnosis of faults in deployed systems. The Journal of Systems and Software, 90(??):29–44, April 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002768.

Mitrovic:2014:RIW

[MIBV14] Dejan Mitrović, Mirjana Ivanović, Zoran Budimac, and Milan Vidaković. Radigost: Interoperable Web-based multi-agent platform. The Journal of Systems and Software, 90(??): 167–178, April 2014. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121214000028.

Mouratidis:2013:FSS

[MIKG13] Haralambos Mouratidis, Shareeful Islam, Christos Kalloniatis, and Stefanos Gritzalis. A framework to support selection of cloud providers based on security and privacy requirements. The Journal of Systems and Software, 86 (9):2276–2293, September 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213000575.

Mirabi:2012:ESB

[MIUM12] Meghdad Mirabi, Hamidah Ibrahim, Nur Izura Udzir, and Ali Mamat. An encoding scheme based on fractional number for querying and updating XML data. The Journal of Systems and Software, 85(8):1831–1851, August 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212000702.

Masoud:2014:CBM

[MJ14] Hamid Masoud and Saeed Jalili. A clusteringbased model for class responsibility assignment problem in object-oriented analysis. The Journal of Systems and Software, 93(??):110–131, July 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214000661.

Mansouri:2018:NPA

[MJ18] N. Mansouri and M. M.

Javidi. A new prefetchingaware data replication to decrease access latency in cloud environment. The Journal of Systems and Software, 144(??):197–215, October 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301055.

Mayer:2019:JSI

[MJ19] Wolfgang Mayer and Bo Jiang. JSS special issue program debugging and repair. The Journal of Systems and Software, 158(??):??, December 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301785.

Ma:2010:SOO

[MJF10] Yinglong Ma, Beihong Jin, and Yulin Feng. Semantic oriented ontology cohesion metrics for ontology-based systems. The Journal of Systems and Software, 83(1):143– 152, January 2010. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

MontesDeOca:2010:CCP

[MJZ^{+}10] Veronica Montes De Oca, Daniel R. Jeske, Qi Zhang, Carlos Rendon, and Mazda Marvasti. A CUSUM changepoint detection algorithm for non-stationary sequences with application to data network surveillance. The Journal of Systems and Software, 83(7):1288–1297, July 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mohanty:2011:RTP

[MK11] Saraju P. Mohanty and Elias Kougianos. Realtime perceptual watermarking architectures for video broadcasting. The Journal of Systems and Software, 84(5):724–738, May 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Milajerdi:2015:CMB

[MK15a] Sadegh Momeni Milajerdi and Mehdi Kharrazi. A compositemetric based path selection technique for the Tor anonymity network. The Journal of Systems and Software, 103(??):53–61, May 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000035.

Moschakis:2015:MCS

[MK15b] Ioannis A. Moschakis and Helen D. Karatza. Multi-criteria scheduling of Bag-of-Tasks applications on heterogeneous interlinked clouds with simulated annealing. The Journal of Systems and Software, 101(??):1–14, March 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121400243X.

Ma:2016:MTC

[MK16] Yu-Seung Ma and Sang-Woon Kim. Mutation testing cost reduction by clustering overlapped mutants. The Journal of Systems and Software, 115(??):18–30, May 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216000078.

Mavridis:2017:PEC

[MK17] Ilias Mavridis and Helen Karatza. Performance evaluation of cloud-based log file analysis with Apache Hadoop and Apache Spark. The Journal of Systems and Software, 125(??):133– 151, March 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216302370.

Malik:2012:AIC

[MKH^{+}12] Sana Ambreen Malik, Asifullah Khan, Mutawarra Hussain, Khurram Jawad, Rafiullah Chamlawi, and Abdul Jalil. Authentication of images for 3D cameras: Reversibly embedding information using intelligent approaches. The Journal of Systems and Software, 85 (11):2665–2673, November 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001719.

Murtaza:2016:MTP

[MKHLB16] Syed Shariyar Murtaza, Wael Khreich, Abdelwahab Hamou-Lhadj, and Ayse Basar Bener. Mining trends and patterns of software vulnerabilities. The Journal of Systems and Software, 117(??):218–228, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000790.

Mansoor:2015:MMO

[MKL^{+}15] Usman Mansoor, Marouane Kessentini, Philip Langer, Manuel Wimmer, Slim Bechikh, and Kalyanmoy Deb. MOMM: Multi-Objective Model Merging. The Journal of Systems and Software, 103(??):423–439, May 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121400274X.

Misra:2014:EDN

[MKRO14] Sudip Misra, Pushpendu Kar, Arijit Roy, and Mohammad S. Obaidat. Existence of dumb nodes in stationary wireless sensor networks. The Journal of Systems and Software, 91(??):135–146, May 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214000120.

Malek:2010:EMS

[MKS10] Sam Malek, Harshini Ramnath Krishnan, and Jayalakshmi Srinivasan. Enhancing middleware support for architecturebased development through compositional weaving of styles. The Journal of Systems and Software, 83 (12):2513–2527, December 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mensah:2018:VPS

[MKS^{+}18] Solomon Mensah, Jacky Keung, Jeffery Svajlenko, Kwabena Ebo Bennin, and Qing Mi. On the value of a prioritization scheme for resolving self-admitted technical debt. The Journal of Systems and Software, 135(??):37–54, January 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302133.

Maglyas:2018:ISI

[ML18] Andrey Maglyas and Anna-Lena Lamprecht. Introduction to the special issue on “Software Business”. The Journal of Systems and Software, 135(??):105–106, January 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302200.

Mao:2014:SBS

[MLD^{+}14] Xiaoguang Mao, Yan Lei, Ziying Dai, Yuhua Qi, and Chengsong Wang. Slice-based statistical fault localization. The Journal of Systems and Software, 89(??):51–62, March 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121213002185.

Meng:2016:POP

[MLD16] Xianfu Meng, Tianjiao Li, and Yu Deng. preferTrust: an ordered preferences-based trust model in peer-to-peer networks. The Journal of Systems and Software, 113(??):309–323, March 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215002848.

Martinez-Llario:2011:DJS

[MLGA11] J. Martinez-Llario and M. Gonzalez-Alcaide. Design of a Java spatial extension for relational databases. The Journal of Systems and Software, 84(12):2314–2323, December 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211001695.

Ma:2012:HCA

[MLHL12] Xiaolin Ma, Fangmin Li, Fei Hu, and Xinhua Liu. A hybrid channel assignment strategy to QoS support of videostreaming over multichannel ad hoc networks. The Journal of Systems and Software, 85(2):300–308, February 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002135.

Mai:2011:DAT

[MLLK11] Hai Thanh Mai, Yu Won Lee, Ki Yong Lee, and Myoung Ho Kim. Distributed adaptive topk monitoring in wireless sensor networks. The Journal of Systems and Software, 84(2):314–327, February 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Maity:2014:FIR

[MM14] Hirak Kumar Maity and Santi P. Maity. FPGA implementation of reversible watermarking in digital images using reversible contrast mapping. The Journal of Systems and Software, 96(??):93–104, October 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001393.

Martinez:2019:AED

[MM19] Matias Martinez and Martin Monperrus. Astor: Exploring the design space of generate-and-validate program repair beyond GenProg. The Journal of Systems and Software, 151(??):65–80, May 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219300159.

Muller:2010:SPI

[MMB10] Sune Dueholm Müller, Lars Mathiassen, and Hans Henrik Balshøj. Software Process Improvement as organizational change: a metaphorical analysis of the literature. The Journal of Systems and Software, 83(11):2128–2146, November 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mandreoli:2015:AEQ

[MMP15] Federica Mandreoli, Riccardo Martoglia, and Wilma Penzo. Approximating expressive queries on graph-modeled data: the GeX approach. The Journal of Systems and Software, 109(??):106– 123, November 2015.

CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121215001545.

Maity:2013:CRS

[MMSD13] Santi P. Maity, Seba Maity, Jaya Sil, and Claude Delpha. Collusion resilient spread spectrum watermarking in M-band wavelets using GA-fuzzy hybridization. The Journal of Systems and Software, 86(1):47–59, January 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001938.

Mubeen:2015:IMT

[MMTS15] Saad Mubeen, Jukka Mäki-Turja, and Mikael Sjödin. Integrating mixed transmission and practical limitations with the worst-case responsetime analysis for Controller Area Network. The Journal of Systems and Software, 99(??):66– 84, January 2015. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121214001952.

Mansouri:2016:NMR

[MMZ^{+}16] Wahida Mansouri, Kais Mnif, Faouzi Zarai, Mohammad S. Obaidat, and Lotfi Kamoun. A new multi-rat scheduling algorithm for heterogeneous wireless networks. The Journal of Systems and Software, 115(??):174–184, May 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215000497.

Mondal:2019:MPW

[MN19] Shouvick Mondal and Rupesh Nasre. Mahtab: Phase-wise acceleration of regression testing for C. The Journal of Systems and Software, 158(??):??, December 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301773.

Morandi:2012:PAS

[MNM12] Benjamin Morandi, Sebastian Nanz, and Bertrand Meyer. Performance analysis of SCOOP programs. The Journal of Systems and Software, 85 (11):2519–2530, November 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001616.

Mackowiak:2018:SEU

[MNO18] Michał Maćkowiak, Jerzy Nawrocki, and Mirosław Ochodek. On some enduser programming constructs and their understandability. The Journal of Systems and Software, 142(??):206–222, August 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300633.

Maglyas:2013:WRS

[MNS13] Andrey Maglyas, Uolevi Nikula, and Kari Smolander. What are the roles of software product managers? An empirical investigation. The Journal of Systems and Software, 86(12):3071–3090, December 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121213001891.

Morales:2015:CEE

[MNSA15] José Miguel Morales, Elena Navarro, Pedro Sánchez, and Diego Alonso. A controlled experiment to evaluate the understandability of KAOS and i* for modeling Teleo-Reactive systems. The Journal of Systems and Software, 100(??):1–14, February 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214002143.

Morales:2016:FEE

[MNSA16] José Miguel Morales, Elena Navarro, Pedro Sánchez, and Diego Alonso. A family of experiments to evaluate the understandability of TRiStar and i^{∗} for modeling teleo-reactive systems. The Journal of Systems and Software, 114(??):82–100, April 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216000042.

Mahieu:2019:SBP

[MOD^{+}19] Christof Mahieu, Femke Ongenae, Femke De Backere, Pieter Bonte, Filip De Turck, and Pieter Simoens. Semanticsbased platform for contextaware and personalized robot interaction in the Internet of Robotic Things. The Journal of Systems and Software, 149(??):138–157, March 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302553.

Mihaylov:2016:ABR

[MOH16] Boyan Mihaylov, Lucian Onea, and Klaus Marius Hansen. Architecturebased regulatory compliance argumentation. The Journal of Systems and Software, 119 (??):1–30, September 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300310.

Midha:2012:FAS

[MP12] Vishal Midha and Prashant Palvia. Factors affecting the success of Open Source Software. The Journal of Systems and Software, 85(4):895–905, April 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121100286X.

Mittas:2015:INP

[MPAA15] Nikolaos Mittas, Efi Papatheocharous, Lefteris Angelis, and Andreas S. Andreou. Integrating non-parametric models with linear components for producing software cost estimations. The Journal of Systems and Software, 99(??):120–134, January 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002088.

Miller:2015:ELM

[MPLL^{+}15] Tim Miller, Sonja Pedell, Antonio A. Lopez-Lorca, Antonette Mendoza, Leon Sterling, and Alen Keirnan. Emotion-led modelling for people-oriented requirements engineering: the case study of emergency systems. The Journal of Systems and Software, 105(??):54–71, July 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000667.

Moghaddam:2018:EVC

[MPLL18] Fahimeh Alizadeh Moghaddam, Giuseppe Procaccianti, Grace A. Lewis, and Patricia Lago. Empirical validation of cyber-foraging architectural tactics for surrogate provisioning. The Journal of Systems and Software, 138(??):37–51, April 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302832.

Mestre:2017:ESB

[MPN^{+}17] Demetrio Gomes Mestre, Carlos Eduardo Santos Pires, Dimas Cassimiro Nascimento, Andreza Raquel Monteiro de Queiroz, Veruska Borges Santos, and Tiago Brasileiro Araujo. An efficient spark-based adaptive windowing for entity matching. The Journal of Systems and Software, 128(??):1–10, June 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217300559.

Mirandola:2014:RMS

[MPRS14] R. Mirandola, P. Potena, E. Riccobene, and P. Scandurra. A reliability model for Service Component Architectures. The Journal of Systems and Software, 89(??):109–127, March 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213002665.

Miller:2012:USO

[MPS^{+}12] Tim Miller, Sonja Pedell, Leon Sterling, Frank Vetere, and Steve Howard. Understanding socially oriented roles and goals through motivational modelling. The Journal of Systems and Software, 85(9):2160–2170, September 2012. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212001094.

Massacci:2014:ARE

[MPTT14] Fabio Massacci, Federica Paci, Le Minh Sang Tran, and Alessandra Tedeschi.

Assessing a requirements evolution approach: Empirical studies in the air traffic management domain. The Journal of Systems and Software, 95(??):70–88, September 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121300280X.

Marti:2017:DDD

[MQG^{+}17] Jonathan Martí, Anna Queralt, Daniel Gasull, Alex Barceló, Juan José Costa, and Toni Cortes. Dataclay: a distributed data store for effective inter-player data sharing. The Journal of Systems and Software, 131 (??):129–145, September 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301012.

Marron:2017:DSC

[MRBN17] Diego Marrón, Jesse Read, Albert Bifet, and Nacho Navarro. Data stream classification using random feature functions and novel method combinations. The Journal of Systems and Software, 127(??):195–204, May 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216300784.

Mendez:2012:GOT

[MRJD^{+}12] José R. Méndez, M. Reboiro-Jato, Fernando Díaz, Eduardo Díaz, and Florentino Fdez-Riverola. Grindstone4Spam: an optimization toolkit for boosting e-mail classification. The Journal of Systems and Software, 85 (12):2909–2920, December 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212001756.

Mayeh:2016:RAC

[MRM16] Maral Mayeh, T. Ramayah, and Alok Mishra. The role of absorptive capacity, communication and trust in ERP adoption. The Journal of Systems and Software, 119(??):58–69, September 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0164121216300565.

Mondal:2019:ESB

[MRRS19] Manishankar Mondal, Banani Roy, Chanchal K. Roy, and Kevin A. Schneider. An empirical study on bug propagation through code cloning. The Journal of Systems and Software, 158(??):??, December 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301815.

Mondal:2018:BPL

[MRS18] Manishankar Mondal, Chanchal K. Roy, and Kevin A. Schneider. Bugproneness and late propagation tendency of code clones: a comparative study on different clone types. The Journal of Systems and Software, 144(??):41–59, October 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301079.

Matalonga:2017:CTM

[MRT17] Santiago Matalonga, Felyppe Rodrigues, and Guilherme Horta Travassos. Characterizing testing methods for contextaware software systems: Results from a quasisystematic literature review. The Journal of Systems and Software, 131(??):1–21, September 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121730095X.

Mayvan:2017:SAD

[MRY17] B. Bafandeh Mayvan, A. Rasoolzadegan, and Z. Ghavidel Yazdi. The state of the art on design patterns: a systematic mapping of the literature. The Journal of Systems and Software, 125 (??):93–118, March 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

//www.sciencedirect.
com/science/article/
pii/S0164121216302321.

MacCormack:2016:TDS

[MS16] Alan MacCormack and Daniel J. Sturtevant. Technical debt and system architecture: the impact of coupling on defect-related activity. The Journal of Systems and Software, 120 (??):170–182, October 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300760.

Merayo:2017:PSI

[MS17a] Mercedes G. Merayo and Gwen Salaün. Preface: Special issue on software verification and testing. The Journal of Systems and Software, 132(??):317–318, October 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301577.

Mezni:2017:MCS

[MS17b] Haithem Mezni and Mokhtar Sellami. Multicloud service composition using Formal Concept Analysis. The Journal of Systems and Software, 134(??):138–152, December 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301760.

Minaeva:2016:SEC

[MSAH16] Anna Minaeva, Premysl Sucha, Benny Akesson, and Zdenek Hanzálek. Scalable and efficient configuration of timedivision multiplexed resources. The Journal of Systems and Software, 113(??):44–58, March 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002447.

Maartensson:2018:EMF

[MSB18] Torvald Mårtensson, Daniel Ståhl, and Jan Bosch. Enable more frequent integration of software in industry projects. The Journal of Systems and Software, 142 (??):223–236, August 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300906.

Mendling:2012:TEP

[MSGGL12] Jan Mendling, Laura Sánchez-González, Félix García, and Marcello La Rosa. Thresholds for error probability measures of business process models. The Journal of Systems and Software, 85(5):1188–1197, May 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000040.

Mujhid:2017:SEF

[MSGM17] Ibrahim Jameel Mujhid, Joanna C. S. Santos, Raghuram Gopalakrishnan, and Mehdi Mirakhorli. A search engine for finding and reusing architecturally significant code. The Journal of Systems and Software, 130(??):81–93, August 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216302345.

Mullins:2018:AGC

[MSHG18] Galen E. Mullins, Paul G.

Stankiewicz, R. Chad Hawthorne, and Satyandra K. Gupta. Adaptive generation of challenging scenarios for testing and evaluation of autonomous vehicles. The Journal of Systems and Software, 137(??):197–215, March 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217302546.

Morales:2017:UDC

[MSK^{+}17] Rodrigo Morales, Zéphyrin Soh, Foutse Khomh, Giuliano Antoniol, and Francisco Chicano. On the use of developers’ context for automatic refactoring of software antipatterns. The Journal of Systems and Software, 128(??):236–251, June 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216300632.

Monsieur:2012:MDD

[MSL12] Geert Monsieur, Monique Snoeck, and Wilfried Lemahieu. Managing data dependencies in service compositions. The Journal of Systems and Software, 85 (11):2604–2628, November 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212001665.

Maro:2018:STA

[MSS18] Salome Maro, Jan-Philipp Steghöfer, and Miroslaw Staron. Software traceability in the automotive domain: challenges and solutions. The Journal of Systems and Software, 141(??):85–110, July 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300608.

Moreno:2012:BSE

[MSSMDC12] Ana M. Moreno, Maria-Isabel Sanchez-Segura, Fuensanta Medina-Dominguez, and Laura Carvajal. Balancing software engineering education and industrial needs. The Journal of Systems and Software, 85(7):1607–1620, July 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000398.

Misra:2010:SLT

[MT10] Sudip Misra and P. Dias Thomasinous. A simple, least-time, and energyefficient routing protocol with one-level data aggregation for wireless sensor networks. The Journal of Systems and Software, 83(5):852–860, May 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mikkonen:2013:CCI

[MT13] Tommi Mikkonen and Antero Taivalsaari. Cloud computing and its impact on mobile software development: Two roads diverged. The Journal of Systems and Software, 86 (9):2318–2320, September 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213000241.

Manteuffel:2016:DAD

[MTA^{+}16] Christian Manteuffel, Dan Tofan, Paris Avgeriou, Heiko Koziolek, and Thomas Goldschmidt. Decision architect — a decision documentation tool for industry. The Journal of Systems and Software, 112 (??):181–198, February 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215002290.

Mate:2014:ASM

[MTF14] Alejandro Maté, Juan Trujillo, and Xavier Franch. Adding semantic modules to improve goal-oriented analysis of data warehouses using Istar. The Journal of Systems and Software, 88(??):102–111, February 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002446.

Moadeli:2010:CMM

[MV10] Mahmoud Moadeli and Wim Vanderbauwhede. Communication modeling of multicast in allport wormhole-routed NoCs. The Journal of Systems and Software, 83(8):1327–1336, August 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Moadeli:2011:AMB

[MV11] Mahmoud Moadeli and Wim Vanderbauwhede. An analytical model of broadcast in QoS-aware wormhole-routed NoCs. The Journal of Systems and Software, 84(1):12– 20, January 2011. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Makki:2018:CSW

[MVLJ18] Majid Makki, Dimitri Van Landuyt, Bert Lagaisse, and Wouter Joosen. A comparative study of workflow customization strategies: Quality implications for multi-tenant SaaS. The Journal of Systems and Software, 144(??):423–438, October 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301420.

Medeiros:2018:QSR

[MVSG18] Juliana Medeiros, Alexandre Vasconcelos, Carla Silva, and Miguel Goulão. Quality of software requirements specification in agile projects: A cross-case analysis of six companies. The Journal of Systems and Software, 142(??):171–194, August 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300888.

Magdaleno:2012:RSD

[MWM12] Andréa Magalhães Magdaleno, Cláudia Maria Lima Werner, and Renata Mendes de Araujo. Reconciling software development models: a quasisystematic review. The Journal of Systems and Software, 85(2):351–369, February 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002287.

Ma:2011:OTT

[MXZ11] Xu Ma, Lingling Xu, and Fangguo Zhang. Oblivious transfer with timedrelease receiver’s privacy. The Journal of Systems and Software, 84(3):460– 464, March 2011. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Nakagawa:2013:RPA

[NAB^{+}13] Elisa Y. Nakagawa, Pablo O. Antonino, Martin Becker, José C. Maldonado, Holger Storf, Karina B. Villela, and Dieter Rombach. Relevance and perspectives of AAL in Brazil. The Journal of Systems and Software, 86(4):985–996, April 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002841.

Noureddine:2013:AMT

[NB13] M. Noureddine and R. Bashroush. An authentication model towards cloud federation in the enterprise. The Journal of Systems and Software, 86 (9):2269–2275, September 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212003408.

Neves:2015:SET

[NBA^{+}15] L. Neves, P. Borba, V. Alves, L. Turnes, L. Teixeira, D. Sena, and U. Kulesza. Safe evolution templates for software product lines. The Journal of Systems and Software, 106(??):42–58, August 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215000801.

Nasr:2017:AEP

[NBA^{+}17] Sana Ben Nasr, Guillaume Bécan, Mathieu Acher, João Bosco Ferreira Filho, Nicolas Sannier, Benoit Baudry, and Jean-Marc Davril. Automated extraction of product comparison matrices from informal product descriptions. The Journal of Systems and Software, 124(??):82–103, February 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216302242.

Nurdiani:2016:IAL

[NBF16] Indira Nurdiani, Jürgen Börstler, and Samuel A. Fricker. The impacts of agile and lean practices on project constraints: a tertiary study. The Journal of Systems and Software, 119 (??):162–183, September 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300863.

Nurdiani:2019:UOA

[NBF^{+}19] Indira Nurdiani, Jürgen Börstler, Samuel Fricker, Kai Petersen, and Panagiota Chatzipetrou. Understanding the order of agile practice introduction: Comparing agile maturity models and practitioners’ experience. The Journal of Systems and Software, 156(??):1–20, October 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301207.

Nguyen:2019:DRT

[NBH19] Vu Nguyen, Barry Boehm, and LiGuo Huang. Determining relevant training data for effort estimation using Windowbased COCOMO calibration. The Journal of Systems and Software, 147(??):124–146, January 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218302310.

Novielli:2019:ISI

[NBM19] Nicole Novielli, Andrew Begel, and Walid Maalej. Introduction to the special issue on affect awareness in software engineering. The Journal of Systems and Software, 148(??):180–182, February 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0164121218302504.

Neto:2013:DRL

[NBR^{+}13] Alberto Costa Neto, Rodrigo Bonifácio, Márcio Ribeiro, Carlos Eduardo Pontual, Paulo Borba, and Fernando Castor. A design rule language for aspect-oriented programming. The Journal of Systems and Software, 86 (9):2333–2356, September 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000861.

Ntanos:2014:CAF

[NBR^{+}14] C. Ntanos, C. Botsikas, G. Rovis, P. Kakavas, and D. Askounis. A context awareness framework for cross-platform distributed applications. The Journal of Systems and Software, 88 (??):138–146, February 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002562.

Narayan:2010:AAB

[NC10] Sumit Narayan and John A. Chandy. AT-TEST: ATTributes-based Extendable STorage. The Journal of Systems and Software, 83(4):548–556, April 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Naedele:2015:MES

[NCK^{+}15] Martin Naedele, Hong-Mei Chen, Rick Kazman, Yuanfang Cai, Lu Xiao, and Carlos V. A. Silva. Manufacturing execution systems: a vision for managing software development. The Journal of Systems and Software, 101(??):59–68, March 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214002532.

Nasseri:2010:CMR

[NCS10] E. Nasseri, S. Counsell, and M. Shepperd. Class movement and relocation: an empirical study of Java inheritance evolution. The Journal of Systems and Software, 83(2):303–315, February 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ni:2019:ESP

[NCW^{+}19] Chao Ni, Xiang Chen, Fangfang Wu, Yuxiang Shen, and Qing Gu. An empirical study on Pareto based multiobjective feature selection for software defect prediction. The Journal of Systems and Software, 152(??):215–238, June 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300573.

Nakagawa:2018:SLS

[NCWK18] Elisa Yumi Nakagawa, Rafael Capilla, Eoin Woods, and Philippe Kruchten. Sustainability and longevity of systems and architectures. The Journal of Systems and Software, 140(??):1– 2, June 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300335.

Nishi:2018:SCC

[ND18] Manziba Akanda Nishi and Kostadin Damevski. Scalable code clone detection and search based on adaptive prefix filtering. The Journal of Systems and Software, 137(??):130–142, March 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217302790.

Nori:2013:SWB

[NDS13] Fatemeh Nori, Mahmood Deypir, and Mohamad Hadi Sadreddini. A sliding window based algorithm for frequent closed itemset mining over data streams. The Journal of Systems and Software, 86(3):615–623, March 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212002828.

Neugebauer:2017:PAR

[NEM17] Olaf Neugebauer, Michael Engel, and Peter Marwedel. A parallelization approach for resourcerestricted embedded heterogeneous MPSoCs inspired by OpenMP. The Journal of Systems and Software, 125(??):439– 448, March 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216301534.

Naeem:2014:EIC

[NES^{+}14] Ensherah A. Naeem, Mustafa M. Abd Elnaby, Naglaa F. Soliman, Alaa M. Abbas, Osama S. Faragallah, Noura Semary, Mohiy M. Hadhoud, Saleh A. Alshebeili, and Fathi E. Abd El-Samie. Efficient implementation of chaotic image encryption in transform domains. The Journal of Systems and Software, 97(??):118–127, November 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001575.

Nakagawa:2011:AOR

[NFSM11] Elisa Y. Nakagawa, Fabiano C. Ferrari, Mariela M. F. Sasaki, and José C. Maldonado. An aspectoriented reference architecture for Software Engineering Environments. The Journal of Systems and Software, 84 (10):1670–1684, October 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001038.

Nassif:2013:TES

[NHC13] Ali Bou Nassif, Danny Ho, and Luiz Fernando Capretz. Towards an early software estimation using log-linear regression and a multilayer perceptron model. The Journal of Systems and Software, 86(1):144–160, January 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212002221.

Narman:2012:UEA

[NHH^{+}12] Per Närman, Hannes Holm, David Höök, Nicholas Honeth, and Pontus Johnson. Using enterprise architecture and technology adoption models to predict application usage. The Journal of Systems and Software, 85(8):1953–1967, August 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212000556.

Nawahdah:2013:SBV

[NI13] Mamoun Nawahdah and Tomoo Inoue. Setting the best view of a virtual teacher in a mixed reality physical-task learning support system. The Journal of Systems and Software, 86(7):1738– 1750, July 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121200249X.

Noferesti:2017:HBD

[NJ17] Morteza Noferesti and Rasool Jalili. HB^{2}DS: a behavior-driven highbandwidth network mining system. The Journal of Systems and Software, 127(??):266–277, May 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301042.

Nanos:2014:XHP

[NK14] Anastassios Nanos and Nectarios Koziris. Xen2MX: High-performance communication in virtualized environments. The Journal of Systems and Software, 95(??):217–230, September 2014. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121214001058.

Nguyen:2015:CLC

[NK15] Nhan Nguyen and Mohammad Maifi Hasan Khan. A closed-loop context aware data acquisition and resource allocation framework for dynamic data driven applications systems (DDDAS) on the cloud. The Journal of Systems and Software, 109(??):88–105, November 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215001521.

Necasky:2012:ECM

[NKMM12] Martin Necaský, Jakub Klímek, Jakub Malý, and Irena Mlýnková. Evolution and change management of XMLbased systems. The Journal of Systems and Software, 85(3):683–707, March 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211002524.

Niu:2017:AUP

[NKZ17] Haoran Niu, Iman Keivanloo, and Ying Zou. API usage pattern recommendation for software development. The Journal of Systems and Software, 129(??):127–139, July 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301200.

Naseem:2013:CCS

[NMM13] Rashid Naseem, Onaiza Maqbool, and Siraj Muhammad. Cooperative clustering for software modularization. The Journal of Systems and Software, 86 (8):2045–2062, August 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000794.

Nguyen:2017:EEL

[NNVD17] Khanh-Van Nguyen, Phi Le Nguyen, Quoc Huy Vu, and Tien Van Do. An energy efficient and load balanced distributed routing scheme for wireless sensor networks with holes. The Journal of Systems and Software, 123(??):92–105, January 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121216302023.

Neyem:2012:RSD

[NOPF12] Andrés Neyem, Sergio F. Ochoa, José A. Pino, and Rubén Darío Franco. A reusable structural design for mobile collaborative applications. The Journal of Systems and Software, 85(3):511–524, March 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211001403.

Nogueira:2012:FBD

[NPC12] Luís Nogueira, Luís Miguel Pinho, and Jorge Coelho. A feedback-based decentralised coordination model for distributed open real-time systems. The Journal of Systems and Software, 85 (9):2145–2159, September 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212001082.

Olszewska:2016:QML

[nPHW^{+}16] Marta Olszewska (née Plaska), Jeanette Heidenberg, Max Weijola, Kirsi Mikkonen, and Ivan Porres. Quantitatively measuring a large-scale agile transformation. The Journal of Systems and Software, 117(??):258– 273, July 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121600087X.

Qiao:2011:TFM

[nQYD11] Ya nan Qiao, Qi Yong, and Hou Di. Tensor Field Model for higherorder information retrieval. The Journal of Systems and Software, 84 (12):2303–2313, December 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001622.

Nayebi:2010:PAO

[NSAK10] Abbas Nayebi, Hamid Sarbazi-Azad, and Gunnar Karlsson. Performance analysis of opportunistic broadcast for delay-tolerant wireless sensor networks. The Journal of Systems and Software, 83(8):1310– 1317, August 2010. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Nostro:2016:AFN

[NSDI16] Nicola Nostro, Romina Spalazzese, Felicita Di Giandomenico, and Paola Inverardi. Achieving functional and non functional interoperability through synthesized connectors. The Journal of Systems and Software, 111(??):185–199, January 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215002149.

Novais:2017:EAC

[NSM17] Renato Novais, José Amancio Santos, and Manoel Mendonça. Experimentally assessing the combination of multiple visualization strategies for software evolution analysis. The Journal of Systems and Software, 128(??):56–71, June 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217300572.

Nt:2013:BKR

[NTdSX13] Jose A. Rodrigues Nt, Luiz Fernando Cardoso Tomaz, Jano Moreira de Souza, and Geraldo Xexéo. Bringing knowledge into recommender systems. The Journal of Systems and Software, 86(7):1751–1758, July 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002804.

Nguyen:2011:DLS

[NTRN11] Hong-Quang Nguyen, David Taniar, J. Wenny Rahayu, and Kinh Nguyen. Double-layered schema integration of heterogeneous XML sources. The Journal of Systems and Software, 84(1):63–76, January 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Nakagawa:2019:ECR

[NTT19] Hiroyuki Nakagawa, Hiromu Toyama, and Tatsuhiro Tsuchiya. Expression caching for runtime verification based on parameterized probabilistic models. The Journal of Systems and Software, 156(??):300–311, October 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121219301463.

Nunez-Varela:2017:SCM

[NVPGMPSM17] Alberto S. Nuñez-Varela, Héctor G. Pérez-Gonzalez, Francisco E. Martínez-Perez, and Carlos Soubervielle-Montalvo. Source code metrics: a systematic mapping study. The Journal of Systems and Software, 128(??):164– 197, June 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217300663.

Wang:2012:ESS

[nWsCqW12] Wei neng Wang, Fu sheng Chen, and Zhi qiang Wang. An endurance solution for solid state drives with cache. The Journal of Systems and Software, 85(11):2553– 2558, November 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212001574.

Nikooghadam:2010:EUE

[NZM10] Morteza Nikooghadam, Ali Zakerolhosseini, and Mohsen Ebrahimi Moghaddam. Efficient utilization of elliptic curve cryptosystem for hierarchical access control. The Journal of Systems and Software, 83(10):1917–1929, October 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Oliveira:2011:RET

[OAC11] Toacy C. Oliveira, Paulo Alencar, and Don Cowan. ReuseTool — an extensible tool support for object-oriented framework reuse. The Journal of Systems and Software, 84(12):2234–2252, December 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001531.

Ovatman:2013:MBC

[OB13] Tolga Ovatman and Feza Buzluca. Model-based cache-aware dispatching of object-oriented software for multicore systems. The Journal of Systems and Software, 86 (11):2754–2770, November 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001441.

Oliveira:2018:DAD

[OBS^{+}18] Juliana Oliveira, Deise Borges, Thaisa Silva, Nelio Cacho, and Fernando Castor. Do Android developers neglect error handling? A maintenance-centric study on the relationship between Android abstractions and uncaught exceptions. The Journal of Systems and Software, 136(??):1–18, February 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217302558.

Ooi:2012:DSP

[OCC12] Boon-Yaik Ooi, Huah-Yong Chan, and Yu-N. Cheah. Dynamic service placement and replication framework to enhance service availability using team formation algorithm. The Journal of Systems and Software, 85 (9):2048–2062, September 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212000428.

Oyetoyan:2013:SCD

[OCC13] Tosin Daniel Oyetoyan, Daniela S. Cruzes, and Reidar Conradi. A study of cyclic dependencies on defect profile of software components. The Journal of Systems and Software, 86(12):3162–3182, December 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001878.

Okamura:2010:CEA

[OD10] Hiroyuki Okamura and Tadashi Dohi. Comprehensive evaluation of aperiodic checkpointing and rejuvenation schemes in operational software system. The Journal of Systems and Software, 83(9):1591– 1604, September 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Otaduy:2017:UAT

[OD17] I. Otaduy and O. Diaz.

User acceptance testing for Agile-developed web-based applications: Empowering customers through wikis and mind maps. The Journal of Systems and Software, 133(??):212–229, November 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121730002X.

Ottensooser:2012:MSB

[OFR^{+}12] Avner Ottensooser, Alan Fekete, Hajo A. Reijers, Jan Mendling, and Con Menictas. Making sense of business process descriptions: an experimental comparison of graphical and textual notations. The Journal of Systems and Software, 85(3):596–606, March 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211002408.

Orehovacki:2013:EPE

[OGK13] Tihomir Orehovacki, Andrina Granić, and Dragutin Kermek. Evaluating the perceived and estimated quality in use of Web 2.0 applications. The Journal of Systems and Software, 86(12):3039–3059, December 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213001362.

Ochoa:2018:SLR

[OGRJ^{+}18] Lina Ochoa, Oscar González-Rojas, Alves Pereira Juliana, Harold Castro, and Gunter Saake. A systematic literature review on the semi-automatic configuration of extended product lines. The Journal of Systems and Software, 144(??):511–532, October 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301511.

Osman:2015:ACF

[OH15] Rasha Osman and Peter G. Harrison. Approximating closed fork-join queueing networks using product-form stochastic Petri-nets. The Journal of Systems and Software, 110(??):264–278, December 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001855.

Ou:2010:CPA

[OHJ10] Hsia-Hung Ou, Min-Shiang Hwang, and Jinn-Ke Jan. A cocktail protocol with the Authentication and Key Agreement on the UMTS. The Journal of Systems and Software, 83(2):316–325, February 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Oliveto:2017:SCA

[OHL17] Rocco Oliveto, Abram Hindle, and Dawn J. Lawrie. Source code analysis and manipulation. The Journal of Systems and Software, 129(??):58–59, July 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217300638.

Ojala:2016:ASR

[Oja16a]  Arto Ojala.  Adjust-

ing software revenue and pricing strategies in the era of cloud computing. The Journal of Systems and Software, 122(??):40–51, December 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216301546.

Ojala:2016:DCB

[Oja16b] Arto Ojala. Discovering and creating business opportunities for cloud services. The Journal of Systems and Software, 113(??):408–417, March 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002393.

Ohzahata:2011:ESP

[OK11] Satoshi Ohzahata and Konosuke Kawashima. An experimental study of peer behavior in a pure P2P network. The Journal of Systems and Software, 84(1):21–28, January 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ochodek:2018:PIA

[OK18] Miroslaw Ochodek and Sylwia Kopczyńska. Perceived importance of agile requirements engineering practices — a survey. The Journal of Systems and Software, 143(??):29–43, September 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300955.

Ouedraogo:2012:ARS

[OKMD12] Moussa Ouedraogo, Djamel Khadraoui, Haralambos Mouratidis, and Eric Dubois. Appraisal and reporting of security assurance at operational systems level. The Journal of Systems and Software, 85(1):193–208, January 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121100210X.

Ouni:2015:IMO

[OKS^{+}15] Ali Ouni, Marouane Kessentini, Houari Sahraoui, Katsuro Inoue, and Mohamed Salah Hamdi. Improving multi-objective code-smells correction using development history. The Journal of Systems and Software, 105(??):18–39, July 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000631.

Oliveira:2015:ASW

[OLV15] Rui André Oliveira, Nuno Laranjeiro, and Marco Vieira. Assessing the security of web service frameworks against Denial of Service attacks. The Journal of Systems and Software, 109(??):18–31, November 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001454.

Ou:2013:RDH

[OLZN13] Bo Ou, Xiaolong Li, Yao Zhao, and Rongrong Ni. Reversible data hiding based on PDE predictor. The Journal of Systems and Software, 86(10):2700–2709, October 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001374.

Ofuonye:2013:SWC

[OM13] Ejike Ofuonye and James Miller. Securing webclients with instrumented code and dynamic runtime monitoring. The Journal of Systems and Software, 86(6):1689– 1711, June 2013. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121213000514.

Ouadjaout:2016:SAA

[OMLB16] Abdelraouf Ouadjaout, Antoine Miné, Noureddine Lasla, and Nadjib Badache. Static analysis by abstract interpretation of functional properties of device drivers in TinyOS. The Journal of Systems and Software, 120(??):114–132, October 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216301261.

Oquendo:2011:GEI

[OPS11] Flavio Oquendo, Eltjo Poort, and Judith Stafford. Guest Editors’ introduction to the special issue. The Journal of Systems and Software, 84(9):1423, September 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211001634.

Ou:2018:CSR

[OSH^{+}18] Zhonghong Ou, Meina Song, Zhen-Huan Hwang, Antti Ylä-Jääski, Ren Wang, Yong Cui, and Pan Hui. Is cloud storage ready? Performance comparison of representative IP-based storage systems. The Journal of Systems and Software, 138(??):206–221, April 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300025.

Oriol:2017:SUO

[OT17] Xavier Oriol and Ernest Teniente. Simplification of UML/OCL schemas for efficient reasoning. The Journal of Systems and Software, 128(??):130–149, June 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121730064X.

Ozakinci:2018:ESD

[ÖT18] Rana Özakinci and Ayça Tarhan. Early software defect prediction: a systematic map and review. The Journal of Systems and Software, 144(??):216–239, October 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121218301213.

Ovatman:2011:EIP

[OWB11] Tolga Ovatman, Thomas Weigert, and Feza Buzluca. Exploring implicit parallelism in class diagrams. The Journal of Systems and Software, 84 (5):821–834, May 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Olsson:2019:ESD

[OWG19] Thomas Olsson, Krzysztof Wnuk, and Tony Gorschek. An empirical study on decision making for quality requirements. The Journal of Systems and Software, 149(??):217– 233, March 2019. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218302668.

Okutan:2016:NKP

[OY16] Ahmet Okutan and Olcay Taner Yildiz. A novel kernel to predict software defectiveness. The Journal of Systems and Software, 119 (??):109–121, September 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300759.

Omheni:2014:MBA

[OZO^{+}14] Nouri Omheni, Faouzi Zarai, Mohammad S. Obaidat, Ikram Smaoui, and Lotfi Kamoun. A MIH-based approach for best network selection in heterogeneous wireless networks. The Journal of Systems and Software, 92(??):143–156, June 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214000338.

Paschali:2017:ROS

[PAB^{+}17] Maria-Eleni Paschali, Apostolos Ampatzoglou, Stamatia Bibi, Alexander Chatzigeorgiou, and Ioannis Stamelos. Reusability of open source software across domains: a case study. The Journal of Systems and Software, 134(??):211–227, December 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301991.

Prieto:2013:SCB

[PÁC13] Víctor M. Prieto, Manuel Álvarez, and Fidel Cacheda. SAAD, a content based Web Spam Analyzer and Detector. The Journal of Systems and Software, 86(11):2906–2918, November 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213001684.

Poulding:2015:OSG

[PACH15] S. Poulding, R. Alexander, J. A. Clark, and M. J. Hadley. The optimisation of stochastic grammars to enable cost-effective probabilistic structural testing. The Journal of Systems and Software, 103(??):296–310, May 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214002738.

Palmieri:2012:PBR

[Pal12] Francesco Palmieri. Percolationbased routing in the Internet. The Journal of Systems and Software, 85 (11):2559–2573, November 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001689.

Patrick:2015:SBT

[PAOC15] Matthew Patrick, Rob Alexander, Manuel Oriol, and John A. Clark. Subdomain-based test data generation. The Journal of Systems and Software, 103(??):328– 342, May 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002647.

Perez:2014:DCC

[PAR14] Alexandre Perez, Rui Abreu, and André Riboira. A dynamic code coverage approach to maximize fault localization efficiency. The Journal of Systems and Software, 90(??):18–28, April 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214000090.

Prayati:2010:MAT

[PAS^{+}10] A. Prayati, Ch. Antonopoulos, T. Stoyanova, C. Koulamas, and G. Papadopoulos. A modeling approach on the TelosB WSN platform power consumption. The Journal of Systems and Software, 83(8): 1355–1363, August 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). See corrigendum [APS^{+}10].

Park:2011:AAS

[PB11] Seunghun Park and Doo-Hwan Bae. An approach to analyzing the software process change impact using process slicing and simulation. The Journal of Systems and Software, 84(4):528–543, April 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Park:2015:ISR

[PB15] Jinhee Park and Jongmoon Baik. Improving software reliability prediction through multicriteria based dynamic model selection and combination. The Journal of Systems and Software, 101(??):236–244, March 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002878.

Pleuss:2012:MDS

[PBD^{+}12] Andreas Pleuss, Goetz Botterweck, Deepak Dhungana, Andreas Polzer, and Stefan Kowalewski. Model-driven support for product line evolution on feature level. The Journal of Systems and Software, 85(10):2261–2274, October 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002093.

Pelliccione:2019:ISI

[PBM19] Patrizio Pelliccione, Jan Bosch, and Marija Mikic. Introduction to the special issue on architecting autonomous and smart systems. The Journal of Systems and Software, 150(??):1–2, April 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300032.

Pardillo:2010:DSL

[PC10] Jesús Pardillo and Cristina Cachero. Domainspecific language modelling with UML profiles by decoupling abstract and concrete syntaxes. The Journal of Systems and Software, 83 (12):2591–2606, December 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Patel:2015:EHL

[PC15] Pankesh Patel and Damien Cassou. Enabling highlevel application development for the Internet of Things. The Journal of Systems and Software, 103(??):62–84, May 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215000187.

Preda:2011:DDC

[PCCB^{+}11] Stere Preda, Frédéric Cuppens, Nora Cuppens-Boulahia, Joaquin Garcia-Alfaro, and Laurent Toutain. Dynamic deployment of contextaware access control policies for constrained security devices. The Journal of Systems and Software, 84(7):1144–1159, July 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pospieszny:2018:EAS

[PCCK18] Przemysław Pospieszny, Beata Czarnacka-Chrobot, and Andrzej Kobylinski. An effective approach for software project effort and duration estimation with machine learning algorithms. The Journal of Systems and Software, 137(??):184–196, March 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302947.

Perez-Castillo:2012:FCS

[PCCLdGP12] Ricardo Pérez-Castillo, José A. Cruz-Lemus, Ignacio García-Rodríguez de Guzmán, and Mario Piattini. A family of case studies on business process mining using MARBLE. The Journal of Systems and Software, 85(6):1370–1385, June 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000088.

Perez-Castillo:2019:BPM

[PCFRP19] Ricardo Pérez-Castillo, María Fernández-Ropero, and Mario Piattini. Business process model refactoring applying IBUPRO-FEN. An industrial evaluation. The Journal of Systems and Software, 147(??):86–103, January 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121830222X.

Piro:2014:ICS

[PCG^{+}14] G. Piro, I. Cianci, L. A. Grieco, G. Boggia, and P. Camarda. Information centric services in Smart Cities. The Journal of Systems and Software, 88(??):169–188, February 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213002586.

Patikirikorala:2012:EMM

[PCHW12] Tharindu Patikirikorala, Alan Colman, Jun Han, and Liuping Wang. An evaluation of multimodel self-managing control schemes for adaptive performance management of software systems. The Journal of Systems and Software, 85 (12):2678–2696, December 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0164121212001628.

Peng:2012:STS

[PCYZ12] Xin Peng, Bihuan Chen, Yijun Yu, and Wenyun Zhao. Self-tuning of software systems through dynamic quality tradeoff and value-based feedback control loop. The Journal of Systems and Software, 85(12):2707–2719, December 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121200132X.

Pruteanu:2012:LDF

[PD12] Andrei Pruteanu and Stefan Dulman. LossEstimate: Distributed failure estimation in wireless networks. The Journal of Systems and Software, 85(12):2785–2795, December 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002233.

Parthasarathy:2016:AED

[PD16] Sudhaman Parthasarathy and Maya Daneva. An approach to estimation of degree of customization for ERP projects using prioritized requirements. The Journal of Systems and Software, 117(??):471–487, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121630022X.

Prieur-Drevon:2018:RSS

[PDBD18] L. Prieur-Drevon, R. Beamonte, and M. R. Dagenais. R-SHT: a state history tree with R-Tree properties for analysis and visualization of highly parallel system traces. The Journal of Systems and Software, 135(??):55–68, January 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302108.

Palsetia:2016:SNX

[PDK^{+}16] Nushafreen Palsetia, G. Deepa, Furqan Ahmed Khan, P. Santhi Thilagam, and Alwyn R. Pais. Securing native XML database-driven web applications from XQuery injection vulnerabilities. The Journal of Systems and Software, 122 (??):93–109, December 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301571.

Pradhan:2016:ARD

[PDL^{+}16] Subhav Pradhan, Abhishek Dubey, Tihamer Levendovszky, Pranav Srinivas Kumar, William A. Emfinger, Daniel Balasubramanian, William Otte, and Gabor Karsai. Achieving resilience in distributed software systems via selfreconfiguration. The Journal of Systems and Software, 122(??):344– 363, December 2016. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121216300590.

Papamichail:2019:MRS

[PDS19] Michail D. Papamichail, Themistoklis Diamantopoulos, and Andreas L. Symeonidis. Measuring the reusability of software components using static analysis metrics and reuse rate information. The Journal of Systems and Software, 158(??):??, December 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301979.

Peng:2011:ESB

[Pen11] Kun Peng. An efficient shuffling based eVoting scheme. The Journal of Systems and Software, 84 (6):906–922, June 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Palviainen:2011:REP

[PEO11] Marko Palviainen, Antti Evesti, and Eila Ovaska. The reliability estimation, prediction and measuring of componentbased software. The Journal of Systems and Software, 84(6):1054– 1070, June 2011. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pinto:2012:DDD

[PFF12] Mónica Pinto, Lidia Fuentes, and Luis Fernández. Deriving detailed design models from an aspect-oriented ADL using MDD. The Journal of Systems and Software, 85(3):525–545, March 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001269.

Pernstaal:2013:LGR

[PFG13] J. Pernstål, R. Feldt, and T. Gorschek. The lean gap: a review of lean approaches to large-scale software systems development. The Journal of Systems and Software, 86 (11):2797–2821, November 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001477.

Procaccianti:2016:EET

[PFL16] Giuseppe Procaccianti, Héctor Fernández, and Patricia Lago. Empirical evaluation of two best practices for energyefficient software development. The Journal of Systems and Software, 117(??):185–198, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000777.

Pizzoleto:2019:SLR

[PFO^{+}19] Alessandro Viola Pizzoleto, Fabiano Cutigi Ferrari, Jeff Offutt, Leo Fernandes, and Márcio Ribeiro. A systematic literature review of techniques and metrics to reduce the cost of mutation testing. The Journal of Systems and Software, 157(??):??, November 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301554.

Pacheco:2012:SLR

[PG12] Carla Pacheco and Ivan Garcia. A systematic literature review of stakeholder identification methods in requirements elicitation. The Journal of Systems and Software, 85 (9):2171–2181, September 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001288.

Perez:2015:MQP

[PG15] Héctor Pérez and J. Javier Gutiérrez. Modeling the QoS parameters of DDS for event-driven real-time applications. The Journal of Systems and Software, 104(??):126–140, June 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121500059X.

Petrillo:2019:SDC

[PGP^{+}19] Fabio Petrillo, Yann-Gaël Guéhéneuc, Marcelo Pimenta, Carla Dal Sasso Freitas, and Foutse Khomh. Swarm debugging: the collective intelligence on interactive debugging. The Journal of Systems and Software, 153(??):152–174, July 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300780.

Perez:2017:DAD

[PGPC17] Héctor Pérez, J. Javier Gutiérrez, Salva Peiró, and Alfons Crespo. Distributed architecture for developing mixed-criticality systems in multi-core platforms. The Journal of Systems and Software, 123(??):145–159, January 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S016412121630156X.

Pozo:2012:CMD

[PGRQVV12] S. Pozo, R. M. Gasca, A. M. Reina-Quintero, and A. J. Varela-Vaca. CONFIDDENT: a modeldriven consistent and non-redundant layer-3 firewall ACL design, development and maintenance framework. The Journal of Systems and Software, 85(2):425–457, February 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211002354.

Peiris:2013:ASE

[PH13] Manjula Peiris and James H. Hill. Adapting system execution traces to support analysis of software system performance properties. The Journal of Systems and Software, 86 (11):2849–2862, November 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213001635.

Preuveneers:2016:SSA

[PHBJ16] Davy Preuveneers, Thomas Heyman, Yolande Berbers, and Wouter Joosen. Systematic scalability assessment for feature oriented multi-tenant services. The Journal of Systems and Software, 116(??):162–176, June 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002897.

Polancic:2010:EEA

[PHR10] Gregor Polančič, Marjan Heričko, and Ivan Rozman. An empirical examination of application frameworks success based on technology acceptance model. The Journal of Systems and Software, 83(4):574–584, April 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Parolia:2013:PDC

[PJK13] Neeraj Parolia, James J.

Jiang, and Gary Klein. The presence and development of competency in IT programs. The Journal of Systems and Software, 86(12):3140–3150, December 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121213001763.

Pinisetty:2017:PRV

[PJT^{+}17] Srinivas Pinisetty, Thierry Jéron, Stavros Tripakis, Yliès Falcone, Hervé Marchand, and Viorel Preoteasa. Predictive runtime verification of timed properties. The Journal of Systems and Software, 132(??):353–365, October 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217301310.

Papazachos:2010:PEB

[PK10a] Zafeirios C. Papazachos and Helen D. Karatza. Performance evaluation of bag of gangs scheduling in a heterogeneous distributed system. The Journal of Systems and Software, 83(8):1346– 1354, August 2010. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Perez:2010:BRM

[PK10b] Abel Marrero Pérez and Stefan Kaiser. Bottomup reuse for multi-level testing. The Journal of Systems and Software, 83 (12):2392–2415, December 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pramila:2018:ICA

[PKS18] Anu Pramila, Anja Keskinarkaus, and Tapio Seppänen. Increasing the capturing angle in print-cam robust watermarking. The Journal of Systems and Software, 135(??):205–215, January 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302522.

Penichet:2010:RBA

[PLGT10] Victor M. R. Penichet, Maria D. Lozano, José A. Gallud, and Ricardo Tesoriero. Requirementbased approach for groupware environments design. The Journal of Systems and Software, 83(8): 1478–1488, August 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pascual:2015:AME

[PLHP^{+}15] Gustavo G. Pascual, Roberto E. Lopez-Herrejon, Mónica Pinto, Lidia Fuentes, and Alexander Egyed. Applying multiobjective evolutionary algorithms to dynamic software product lines for reconfiguring mobile applications. The Journal of Systems and Software, 103(??):392–411, May 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121400291X.

Palomba:2018:CUR

[PLVB^{+}18] Fabio Palomba, Mario Linares-Vásquez, Gabriele Bavota, Rocco Oliveto, Massimiliano Di Penta, Denys Poshyvanyk, and Andrea De Lucia. Crowdsourcing user reviews to support the evolution of mobile apps. The Journal of Systems and Software, 137(??):143–162, March 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302807.

Psomopoulos:2010:BAD

[PM10] Fotis E. Psomopoulos and Pericles A. Mitkas. Bioinformatics algorithm development for Grid environments. The Journal of Systems and Software, 83(7):1249–1257, July 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pitangueira:2015:SRS

[PMB15] Antônio Mauricio Pitangueira, Rita Suzana P. Maciel, and Márcio Barros. Software requirements selection and prioritization using SBSE approaches: a systematic review and mapping of the literature. The Journal of Systems and Software, 103(??):267– 280, May 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002118.

Pan:2013:LBR

[PMDH13] Zhibin Pan, Xiaoxiao Ma, Xiaoman Deng, and Sen Hu. Low bitrate information hiding method based on search-order-coding technique. The Journal of Systems and Software, 86 (11):2863–2869, November 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001672.

Plaza:2011:MAA

[PMMM11] Inmaculada Plaza, Lourdes Martín, Sergio Martin, and Carlos Medrano. Mobile applications in an aging society: Status and trends. The Journal of Systems and Software, 84(11):1977–1988, November 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121100135X.

Paige:2016:EMM

[PMR16] Richard F. Paige, Nicholas Matragkas, and Louis M. Rose. Evolving models in model-driven engineering: State-of-the-art and future challenges. The Journal of Systems and Software, 111(??):272–280, January 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215001909.

Prudencio:2012:LLQ

[PMWC12] João Gustavo Prudêncio, Leonardo Murta, Cláudia Werner, and Rafael Cepêda. To lock, or not to lock: That is the question. The Journal of Systems and Software, 85(2):277–289, February 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001063.

Papagiannis:2014:HAS

[PN14] Anastasios Papagiannis and Dimitrios S. Nikolopoulos. Hybrid address spaces: a methodology for implementing scalable high-level programming models on non-coherent many-core architectures. The Journal of Systems and Software, 97(??):47–64, November 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001496.

Palanca:2012:DGO

[PNJGF12] Javier Palanca, Martí Navarro, Vicente Julian, and Ana García-Fornes. Distributed goal-oriented computing. The Journal of Systems and Software, 85(7):1540–1557, July 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121212000337.

Park:2014:OFF

[PNY14] Hyunchan Park, Sam H. Noh, and Chuck Yoo. O1FS: Flash file system with O(1) crash recovery time. The Journal of Systems and Software, 97(??):86–96, November 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214001514.

Potena:2013:OAP

[Pot13] Pasqualina Potena. Optimization of adaptation plans for a serviceoriented architecture with cost, reliability, availability and performance tradeoff. The Journal of Systems and Software, 86(3):624–648, March 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212003226.

Pereira:2016:SHB

[PPB16] Geovandro C. C. F. Pereira, Cassius Puodzius, and Paulo S. L. M. Barreto. Shorter hashbased signatures. The Journal of Systems and Software, 116(??):95– 100, June 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001466.

Pascarella:2019:FGJ

[PPB19] Luca Pascarella, Fabio Palomba, and Alberto Bacchelli. Fine-grained just-in-time defect prediction. The Journal of Systems and Software, 150(??):22–36, April 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302656.

Pino:2010:USG

[PPG^{+}10] Francisco J. Pino, Oscar Pedreira, Félix García, Miguel Rodríguez Luaces, and Mario Piattini. Using Scrum to guide the execution of software process improvement in small organizations. The Journal of Systems and Software, 83(10):1662– 1677, October 2010. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pardo:2013:CSH

[PPG^{+}13] César Pardo, Francisco J. Pino, Félix Garcia, Maria Teresa Baldassarre, and Mario Piattini. From chaos to the systematic harmonization of multiple reference models: a harmonization framework applied in two case studies. The Journal of Systems and Software, 86(1):125–143, January 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212002282.

Perez-Palacin:2012:QEM

[PPMM12] Diego Perez-Palacin, Raffaela Mirandola, and José Merseguer. QoS and energy management with Petri nets: a self-adaptive framework. The Journal of Systems and Software, 85 (12):2796–2811, December 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212001306.

Perez-Palacin:2014:RBQ

[PPMM14] Diego Perez-Palacin, Raffaela Mirandola, and José Merseguer. On the relationships between QoS and software adaptability at the architectural level. The Journal of Systems and Software, 87(??):1–17, January 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001982.

Perez-Palacin:2017:AME

[PPMM17] Diego Perez-Palacin, Raffaela Mirandola, and José Merseguer. Accurate modeling and efficient QoS analysis of scalable adaptive systems under bursty workload. The Journal of Systems and Software, 130(??):24–41, August 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217300833.

Paschou:2015:EHP

[PPN^{+}15] Mersini Paschou, Christos Papadimitiriou, Nikolaos Nodarakis, Konstantinos Korezelidis, Evangelos Sakkopoulos, and Athanasios Tsakalidis. Enhanced healthcare personnel rostering solution using mobile technologies. The Journal of Systems and Software, 100(??):44–53, February 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214002271.

Pironti:2012:FBS

[PPS12] Alfredo Pironti, Davide Pozza, and Riccardo Sisto. Formally based semi-automatic implementation of an open security protocol. The Journal of Systems and Software, 85(4):835–849, April 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211002846.

Petroni:2016:LFL

[PQBP16] F. Petroni, L. Querzoni, R. Beraldi, and M. Paolucci. LCBM: a fast and lightweight collaborative filtering algorithm for binary ratings. The Journal of Systems and Software, 117(??):583–594, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0164121216300371.

Prasetya:2018:TAQ

[Pra18] I. S. W. B. Prasetya.

Temporal algebraic query of test sequences. The Journal of Systems and Software, 136(??):223– 236, February 2018. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S016412121730153X.

Pira:2017:DDC

[PRN17] Einollah Pira, Vahid Rafe, and Amin Nikanjam. Deadlock detection in complex software systems specified through graph transformation using Bayesian optimization algorithm. The Journal of Systems and Software, 131 (??):181–200, September 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301061.

Pedrycz:2011:MJS

[PRS11] Witold Pedrycz, Barbara Russo, and Giancarlo Succi. A model of job satisfaction for collaborative development processes. The Journal of Systems and Software, 84 (5):739–752, May 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pachauri:2013:ATD

[PS13] Ankur Pachauri and Gursaran Srivastava. Automated test data generation for branch testing using genetic algorithm: an improved approach using branch ordering, memory and elitism. The Journal of Systems and Software, 86(5):1191– 1208, May 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212003263.

Ploskas:2014:GAP

[PS14] Nikolaos Ploskas and Nikolaos Samaras. GPU accelerated pivoting rules for the simplex algorithm. The Journal of Systems and Software, 96(??):1–9, October 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001174.

Paixao:2015:ROA

[PS15] Matheus Paixao and Jefferson Souza. A robust optimization approach to the next release problem in the presence of uncertainties. The Journal of Systems and Software, 103(??):281–295, May 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121400212X.

Pierantonio:2016:MEI

[PS16] Alfonso Pierantonio and Bernhard Schätz. Models and evolution: an introduction to the special issue. The Journal of Systems and Software, 111(??):270–271, January 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001089.

Pereira:2013:SLC

[PSdO^{+}13] Geovandro C. C. F.

Pereira, Mateus A. S. Santos, Bruno T. de Oliveira, Marcos A. Simplicio, Jr., Paulo S. L. M. Barreto, Cíntia B. Margi, and Wilson V. Ruggiero. SM-SCrypto: a lightweight cryptographic framework for secure SMS transmission. The Journal of Systems and Software, 86(3):698–706, March 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212003056.

Pareto:2012:CPA

[PSEE12] Lars Pareto, Anna Börjesson Sandberg, Peter Eriksson, and Staffan Ehnebom. Collaborative prioritization of architectural concerns. The Journal of Systems and Software, 85 (9):1971–1994, September 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001252.

Papadimitriou:2012:FAL

[PSM12] Alexis Papadimitriou, Panagiotis Symeonidis, and Yannis Manolopoulos. Fast and accurate link prediction in social networking systems. The Journal of Systems and Software, 85 (9):2119–2132, September 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001069.

Pereira:2011:FIF

[PSNB11] Geovandro C. C. F.

Pereira, Marcos A. Simplício Jr., Michael Naehrig, and Paulo S. L. M. Barreto. A family of implementation-friendly BN elliptic curves. The Journal of Systems and Software, 84(8):1319– 1326, August 2011. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Paterno:2011:EAU

[PSS11] Fabio Paternò, Carmen Santoro, and Lucio Davide Spano. Engineering the authoring of usable service front ends. The Journal of Systems and Software, 84 (10):1806–1822, October 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001257.

Parejo:2016:MOT

[PSS^{+}16] José A. Parejo, Ana B.

Sánchez, Sergio Segura, Antonio Ruiz-Cortés, Roberto E. Lopez-Herrejon, and Alexander Egyed. Multi-objective test case prioritization in highly configurable systems: a case study. The Journal of Systems and Software, 122(??):287–310, December 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301935.

Plakidas:2017:ERS

[PSZ17] Konstantinos Plakidas, Daniel Schall, and Uwe Zdun. Evolution of the R software ecosystem: Metrics, relationships, and their impact on qualities. The Journal of Systems and Software, 132(??):119–146, October 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301371.

Pinto:2015:LSS

[PTF^{+}15] Gustavo Pinto, Weslley Torres, Benito Fernandes, Fernando Castor, and Roberto S. M. Barros. A large-scale study on the usage of Java’s concurrent programming constructs. The Journal of Systems and Software, 106(??):59–81, August 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000849.

Prado:2018:TCS

[PV18] Marllos Paiva Prado and Auri Marcelo Rizzo Vincenzi. Towards cognitive support for unit testing: a qualitative study with practitioners. The Journal of Systems and Software, 141(??):66–84, July 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300529.

Poort:2012:RAR

[PvV12] Eltjo R. Poort and Hans van Vliet. RCDA: Architecting as a risk- and cost management discipline. The Journal of Systems and Software, 85 (9):1995–2013, September 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000994.

Petersen:2010:SPI

[PW10] Kai Petersen and Claes Wohlin. Software process improvement through the Lean Measurement (SPI-LEAM) method. The Journal of Systems and Software, 83(7): 1275–1287, July 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pill:2018:AGF

[PW18] Ingo Pill and Franz Wotawa. Automated generation of (F)LTL oracles for testing and debugging. The Journal of Systems and Software, 139(??):124–141, May 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300281.

Pradhan:2019:ERM

[PWA^{+}19] Dipesh Pradhan, Shuai Wang, Shaukat Ali, Tao Yue, and Marius Liaaen. Employing rule mining and multi-objective search for dynamic test case prioritization. The Journal of Systems and Software, 153(??):86– 104, July 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121930072X.

Por:2012:UTB

[PWC12] Lip Yee Por, KokSheik Wong, and Kok Onn Chee. UniSpaCh: a text-based data hiding method using Unicode space characters. The Journal of Systems and Software, 85(5):1075– 1082, May 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211003177.

Pei:2013:ARW

[PWLL13] Qingqi Pei, Xiang Wang, Yuan Li, and Hui Li. Adaptive reversible watermarking with improved embedding capacity. The Journal of Systems and Software, 86 (11):2841–2848, November 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001623.

Phaphoom:2015:SSM

[PWS^{+}15] Nattakarn Phaphoom, Xiaofeng Wang, Sarah Samuel, Sven Helmer, and Pekka Abrahamsson. A survey study on major technical barriers affecting the decision to adopt cloud services. The Journal of Systems and Software, 103(??):167– 181, May 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215000382.

Peng:2010:IWM

[PWW10] Hong Peng, Jun Wang, and Weixing Wang. Image watermarking method in multiwavelet domain based on support vector machines. The Journal of Systems and Software, 83(8):1470–1477, August 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Park:2016:THB

[PWY^{+}16] Sungjin Park, Jong-Jin Won, Jaenam Yoon, Kyong Hoon Kim, and Taisook Han. A tiny hypervisor-based trusted geolocation framework with minimized TPM operations. The Journal of Systems and Software, 122(??):202–214, December 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216301856.

Peng:2013:IFL

[PXT^{+}13] Xin Peng, Zhenchang Xing, Xi Tan, Yijun Yu, and Wenyun Zhao. Improving feature location using structural similarity and iterative graph mapping. The Journal of Systems and Software, 86(3):664–676, March 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212003007.

Psiuk:2015:GDA

[PZ15] Marek Psiuk and Krzysztof Zielinski. Goal-driven adaptive monitoring of SOA systems. The Journal of Systems and Software, 110(??):101–121, December 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001764.

Pazzi:2010:DEN

[PZB10] Richard W. Pazzi, Zhenxia Zhang, and Azzedine Boukerche. Design and evaluation of a novel MAC layer handoff protocol for IEEE 802.11 wireless networks. The Journal of Systems and Software, 83(8):1364– 1372, August 2010. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Qusef:2014:RTC

[QBO^{+}14] Abdallah Qusef, Gabriele Bavota, Rocco Oliveto, Andrea De Lucia, and Dave Binkley. Recovering test-to-code traceability using slicing and textual analysis. The Journal of Systems and Software, 88(??):147– 168, February 2014. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121213002574.

Qu:2015:ECS

[QGZ^{+}15] Yu Qu, Xiaohong Guan, Qinghua Zheng, Ting Liu, Lidan Wang, Yuqiao Hou, and Zijiang Yang. Exploring community structure of software call graph and its applications in class cohesion measurement. The Journal of Systems and Software, 108(??):193–210, October 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001259.

Qiu:2017:USR

[QLBS17] Dong Qiu, Bixin Li, Earl T. Barr, and Zhendong Su. Understanding the syntactic rule usage in Java. The Journal of Systems and Software, 123(??):160–172, January 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216302126.

Quiroga:2016:ORP

[QOLJG16] Jose Quiroga, Francisco Ortin, David Llewellyn-Jones, and Miguel Garcia. Optimizing runtime performance of hybrid dynamically and statically typed languages for the .Net platform. The Journal of Systems and Software, 113(??):114–129, March 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002654.

Qin:2016:SSB

[QXYL16] Yi Qin, Chang Xu, Ping Yu, and Jian Lu. SIT: Sampling-based interactive testing for selfadaptive apps. The Journal of Systems and Software, 120(??):70–88, October 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216301029.

Qian:2012:LDH

[QZ12] Zhenxing Qian and Xinpeng Zhang. Lossless data hiding in JPEG bitstream. The Journal of Systems and Software, 85(2):309–313, February 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211002123.

Qian:2014:IAF

[QZ14] Zhenxing Qian and Xinpeng Zhang. Improved anti-forensics of JPEG compression. The Journal of Systems and Software, 91(??):100–108, May 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214000168.

Ronglong:2016:SOS

[RA16] Suthat Ronglong and Chonlameth Arpnikanondt. Signal: an open-source cross-platform universal messaging system with feedback support. The Journal of Systems and Software, 117(??):30–54, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000510.

Rodriguez:2015:DPP

[RAJ15] Francy D. Rodríguez, Silvia T. Acuña, and Natalia Juristo. Design and programming patterns for implementing usability functionalities in web applications. The Journal of Systems and Software, 105(??):107– 124, July 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000795.

Rashid:2015:TTS

[RAK15] Muhammad Rashid, Muhammad Waseem Anwar, and Aamir M. Khan. Toward the tools selection in model based system engineering for embedded systems — a systematic literature review. The Journal of Systems and Software, 106(??):150–163, August 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121500103X.

Rahmani:2014:ARA

[RAS14] M. Rahmani, A. Azadmanesh, and H. Siy. Architectural reliability analysis of frameworkintensive applications: a web service case study. The Journal of Systems and Software, 94 (??):186–201, August 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121400082X.

Rodrigues:2012:DAA

[RASL12] Genaína Nunes Rodrigues, Vander Alves, Renato Silveira, and Luiz A. Laranjeira. Dependability analysis in the Ambient Assisted Living Domain: an exploratory case study. The Journal of Systems and Software, 85(1):112–131, January 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0164121211002056.

Rogstad:2016:CES

[RB16] Erik Rogstad and Lionel Briand. Costeffective strategies for the regression testing of database applications: Case study and lessons learned. The Journal of Systems and Software, 113(??):257–274, March 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002794.

rezaBazi:2017:CFC

[rBHM17] Hamid reza Bazi, Alireza Hassanzadeh, and Ali Moeini. A comprehensive framework for cloud computing migration using meta-synthesis approach. The Journal of Systems and Software, 128(??):87–105, June 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217300456.

Rocha:2019:UAT

[RBS19] Thaís Rocha, Paulo Borba, and João Pedro Santos. Using acceptance tests to predict files changed by programming tasks. The Journal of Systems and Software, 154(??):176–195, August 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300998.

Rafique:2011:RSC

[RBT11] M. Mustafa Rafique, Ali R. Butt, and Eli Tilevich. Reusable software components for accelerator-based clusters. The Journal of Systems and Software, 84(7): 1071–1081, July 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rasmussen:2018:LFB

[RBW18] Rune Rasmussen, Alistair Barros, and Fuguo Wei. A likelihoodfree Bayesian derivation method for service variants. The Journal of Systems and Software, 143(??):87–99, September 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300943.

Reyes:2011:OSP

[RCCVB11] Francisco Reyes, Narciso Cerpa, Alfredo Candia-Véjar, and Matthew Bardeen. The optimization of success probability for software projects using genetic algorithms. The Journal of Systems and Software, 84(5):775– 785, May 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rezaei:2014:RBI

[RCL14] Reza Rezaei, Thiam Kian Chiew, and Sai Peck Lee. A review on Ebusiness Interoperability Frameworks. The Journal of Systems and Software, 93(??):199–216, July 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121400051X.

Ruan:2019:DRI

[RCPZ19] Hang Ruan, Bihuan Chen, Xin Peng, and Wenyun Zhao. DeepLink: Recovering issue-commit links based on deep learning. The Journal of Systems and Software, 158(??):??, December 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301803.

Rago:2019:DCM

[RDPM19] Alejandro Rago, J. Andres Diaz-Pace, and Claudia Marcos. Do concern mining tools really help requirements analysts? An empirical study of the vetting process. The Journal of Systems and Software, 156(??):181–203, October 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301359.

Russo:2019:SVD

[RDVC19] Ernesto Rosario Russo, Andrea Di Sorbo, Corrado A. Visaggio, and Gerardo Canfora. Summarizing vulnerabilities’ descriptions to support experts during vulnerability assessment activities. The Journal of Systems and Software, 156(??):84–99, October 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S016412121930130X.

Rufiange:2014:VPV

[RF14] Sébastien Rufiange and Christopher P. Fuhrman. Visualizing protected variations in evolving software designs. The Journal of Systems and Software, 88(??):231– 249, February 2014. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S016412121300263X.

Raibulet:2018:CTS

[RF18] Claudia Raibulet and Francesca Arcelli Fontana. Collaborative and teamwork software development in an undergraduate software engineering course. The Journal of Systems and Software, 144(??):409–422, October 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301389.

Rahmani:2010:NRT

[RFM10] Hossein Rahmani, Mohammad Mehdi Faghih, and Mohsen Ebrahimi Moghaddam. A new real time disk-scheduling method based on GSR algorithm. The Journal of Systems and Software, 83 (11):2147–2164, November 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ren:2010:CSH

[RG10] Yanli Ren and Dawu Gu. CCA2 secure (hierarchical) identity-based parallel key-insulated encryption without random oracles. The Journal of Systems and Software, 83(1): 153–162, January 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rivas:2017:SFE

[RGH17] Juan M. Rivas, J. Javier Gutiérrez, and Michael González Harbour. A supercomputing framework for the evaluation of realtime analysis and optimization techniques. The Journal of Systems and Software, 124 (??):120–136, February 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216302163.

Rabiser:2017:CFR

[RGV^{+}17] Rick Rabiser, Sam Guinea, Michael Vierhauser, Luciano Baresi, and Paul Grünbacher. A comparison framework for runtime monitoring approaches. The Journal of Systems and Software, 125(??):309–321, March 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL /

/www.sciencedirect.
com/science/article/
pii/S0164121216302618.

Rehn:2018:ICP

[RHHT18] Adam Rehn, Jason Holdsworth, John Hamilton, and Singwhat Tee. An input-centric performance model for computational offloading of mobile applications. The Journal of Systems and Software, 138(??):1–18, April 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217303035.

Rodriguez:2017:CDS

[RHL^{+}17] Pilar Rodríguez, Alireza Haghighatkhah, Lucy Ellen Lwakatare, Susanna Teppola, Tanja Suomalainen, Juho Eskeli, Teemu Karvonen, Pasi Kuvaja, June M. Verner, and Markku Oivo. Continuous deployment of software intensive products and services: a systematic mapping study. The Journal of Systems and Software, 123 (??):263–291, January 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215002812.

Rupakheti:2018:PPU

[RHM^{+}18] Chandan R. Rupakheti, Mark Hays, Sriram Mohan, Stephen Chenoweth, and Amanda Stouder. On a pursuit for perfecting an undergraduate requirements engineering course. The Journal of Systems and Software, 144(??):366–381, October 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218301390.

Rivero:2013:MTE

[RHRC13] Carlos R. Rivero, Inma Hernández, David Ruiz, and Rafael Corchuelo. MostoDE: a tool to exchange data amongst semantic-web ontologies. The Journal of Systems and Software, 86(6):1517–1529, June 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000174.

Rivero:2015:MTE

[RHRC15] Carlos R. Rivero, Inma Hernández, David Ruiz, and Rafael Corchuelo. MostoDEx: a tool to exchange RDF data using exchange samples. The Journal of Systems and Software, 100(??):67–79, February 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002301.

Robert-Inacio:2011:SAP

[RITF^{+}11] Frédérique Robert-Inacio, Alain Trémeau, Mike Fournigault, Yannick Teglia, and Pierre-Yvan Liardet. Shape analysis for power signal cryptanalysis on secure components. The Journal of Systems and Software, 84 (5):753–762, May 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rola:2016:CMW

[RKK16] Pawel Rola, Dorota Kuchta, and Dominika Kopczyk. Conceptual model of working space for Agile (Scrum) project team. The Journal of Systems and Software, 118(??):49–63, August 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300401.

Ren:2018:BTC

[RLL^{+}18] Jiankang Ren, Chi Lin, Qian Liu, Mohammad S. Obaidat, Guowei Wu, and Guozhen Tan. Broadcast tree construction framework in Tactile Internet via dynamic algorithm. The Journal of Systems and Software, 136(??):59–73, February 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302704.

Ren:2013:DTE

[RLY^{+}13] Yufei Ren, Tan Li, Dantong Yu, Shudong Jin, and Thomas Robertazzi. Design and testbed evaluation of RDMA-based middleware for highperformance data transfer applications. The Journal of Systems and Software, 86(7):1850– 1863, July 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000277.

Rieger:2019:TDE

[RM19a] Christoph Rieger and Tim A. Majchrzak. Towards the definitive evaluation framework for cross-platform app development approaches. The Journal of Systems and Software, 153(??):175– 199, July 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300743.

Rojas:2019:TCP

[RM19b] Luis A. Rojas and José A. Macías. Toward collisions produced in requirements rankings: a qualitative approach and experimental study. The Journal of Systems and Software, 158(??):??, December 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301918.

Risco-Martin:2014:MAO

[RMCH^{+}14] José L. Risco-Martín, J. Manuel Colmenar, J. Ignacio Hidalgo, Juan Lanchares, and Josefa Díaz. A methodology to automatically optimize dynamic memory managers applying grammatical evolution. The Journal of Systems and Software, 91(??):109–123, May 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121400017X.

Russello:2011:PBP

[RMD11] Giovanni Russello, Leonardo Mostarda, and Naranker Dulay. A policy-based publish/subscribe middleware for sense-and-react applications. The Journal of Systems and Software, 84(4):638–654, April 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rafailidis:2014:LHN

[RNC14] Dimitrios Rafailidis, Alexandros Nanopoulos, and Eleni Constantinou. “With a little help from new friends”: Boosting information cascades in social networks based on link injection. The Journal of Systems and Software, 98(??):1–8, December 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001745.

Roumani:2017:ATE

[RNR17] Yaman Roumani, Joseph K.

Nwankpa, and Yazan F. Roumani. Adopters’ trust in enterprise open source vendors: an empirical examination. The Journal of Systems and Software, 125(??):256– 270, March 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

//www.sciencedirect.
com/science/article/
pii/S0164121216302503.

Raspotnig:2013:CRI

[RO13a] Christian Raspotnig and Andreas Opdahl. Comparing risk identification techniques for safety and security requirements. The Journal of Systems and Software, 86(4):1124–1151, April 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121212003305.

Redondo:2013:ESD

[RO13b] Jose Manuel Redondo and Francisco Ortin. Efficient support of dynamic inheritance for class- and prototypebased languages. The Journal of Systems and Software, 86(2):278–301, February 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212002324.

Ruano-Ordas:2013:ESS

[ROFGFRM13] D. Ruano-Ordás, J. Fdez-Glez, F. Fdez-Riverola, and J. R. Méndez. Effective scheduling strategies for boosting performance on rule-based spam filtering frameworks. The Journal of Systems and Software, 86 (12):3151–3161, December 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213001805.

Rabiser:2011:KAP

[ROR11] Rick Rabiser, Pádraig O’Leary, and Ita Richardson. Key activities for product derivation in software product lines. The Journal of Systems and Software, 84(2):285– 300, February 2011. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rosa:2013:CIE

[RPK^{+}13] Wilson Rosa, Travis Packard, Abishek Krupanand, James W. Bilbro, and Max M. Hodal. COTS integration and estimation for ERP. The Journal of Systems and Software, 86(2):538–550, February 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002713.

Rhee:2010:TSS

[RPSL10] Hyun Sook Rhee, Jong Hwan Park, Willy Susilo, and Dong Hoon Lee. Trapdoor security in a searchable public-key encryption scheme with a designated tester. The Journal of Systems and Software, 83(5):763–771, May 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Razavian:2019:ERS

[RPT19] Maryam Razavian, Barbara Paech, and Antony Tang. Empirical research for software architecture decision making: an analysis. The Journal of Systems and Software, 149(??):360–381, March 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121830267X.

Ren:2017:NLN

[RQD^{+}17] Jianbao Ren, Yong Qi, Yuehua Dai, Yu Xuan, and Yi Shi. Nosv: a lightweight nestedvirtualization VMM for hosting high performance computing on cloud. The Journal of Systems and Software, 124 (??):137–152, February 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216302151.

Ruiz:2017:TSD

[RRM17] Fco. Javier Bermúdez Ruiz, Óscar Sánchez Ramón, and Jesús García Molina. A tool to support the definition and enactment of model-driven migration processes. The Journal of Systems and Software, 128(??):106– 129, June 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217300584.

Ramirez:2019:SMO

[RRV19] Aurora Ramírez, José Raúl Romero, and Sebastián Ventura. A survey of many-objective optimisation in searchbased software engineering. The Journal of Systems and Software, 149(??):382–395, March 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302759.

Rana:2014:SSR

[RSB^{+}14] Rakesh Rana, Mirosław Staron, Christian Berger, Jörgen Hansson, Martin Nilsson, Fredrik Törner, Wilhelm Meding, and Christoffer Höglund. Selecting software reliability growth models and improving their predictive accuracy using historical projects data. The Journal of Systems and Software, 98(??):59–78, December 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121400185X.

Rana:2016:ADI

[RSB^{+}16] Rakesh Rana, Mirosław Staron, Christian Berger, Jörgen Hansson, Martin Nilsson, and Wilhelm Meding. Analyzing defect inflow distribution and applying Bayesian inference method for software defect prediction in large software projects. The Journal of Systems and Software, 117(??):229–244, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216000480.

Rezende:2019:SPS

[RSBA19] Allan Vinicius Rezende, Leila Silva, André Britto, and Rodrigo Amaral. Software project scheduling problem in the context of search-based software engineering: a systematic review. The Journal of Systems and Software, 155 (??):43–56, September 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301086.

Rockai:2018:DMC

[RSCB18] Petr Rockai, Vladimír Still, Ivana Cerná, and Jirí Barnat. DiVM: Model checking with LLVM and graph memory. The Journal of Systems and Software, 143(??):1–13, September 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300700.

Rodriguez:2012:EFT

[RSGH12] D. Rodríguez, M. A.

Sicilia, E. García, and R. Harrison. Empirical findings on team size and productivity in software development. The Journal of Systems and Software, 85(3):562–570, March 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002366.

Raatikainen:2019:SPL

[RTM19] Mikko Raatikainen, Juha Tiihonen, and Tomi Männistö. Software product lines and variability modeling: a tertiary study. The Journal of Systems and Software, 149(??):485–510, March 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121830284X.

Roumelis:2017:EQP

[RVCM17] George Roumelis, Michael Vassilakopoulos, Antonio Corral, and Yannis Manolopoulos. Efficient query processing on large spatial databases: a performance study. The Journal of Systems and Software, 132(??):165–185, October 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301383.

Raemaekers:2017:SVI

[RvDV17] S. Raemaekers, A. van Deursen, and J. Visser. Semantic versioning and impact of breaking changes in the Maven repository. The Journal of Systems and Software, 129(??):140–158, July 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300243.

Ren:2019:EAB

[RXY^{+}19] Jiankang Ren, Zichuan Xu, Chao Yu, Chi Lin, Guowei Wu, and Guozhen Tan. Execution allowance based fixed priority scheduling for probabilistic realtime systems. The Journal of Systems and Software, 152(??):120–133, June 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300524.

Ryan:2013:CCS

[Rya13] Mark D. Ryan. Cloud computing security: the scientific challenge, and a survey of solutions. The Journal of Systems and Software, 86 (9):2263–2268, September 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212003378.

Rong:2018:REE

[RZL^{+}18] Guoping Rong, He Zhang, Bohan Liu, Qi Shan, and Dong Shao. A replicated experiment for evaluating the effectiveness of pairing practice in PSP education. The Journal of Systems and Software, 136(??):139–152, February 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301668.

Razo-Zapata:2012:MAB

[RZMPM12] Iván S. Razo-Zapata, Carlos Mex-Perera, and Raúl Monroy. Masquerade attacks based on user’s profile. The Journal of Systems and Software, 85(11):2640–2651, November 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001707.

Shoufan:2011:BEP

[SA11] Abdulhadi Shoufan and Tolga Arul. A benchmarking environment for performance evaluation of tree-based rekeying algorithms. The Journal of Systems and Software, 84(7):1130–1143, July 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Scacchi:2012:URL

[SA12] Walt Scacchi and Thomas A.

Alspaugh. Understanding the role of licenses and evolution in open architecture software ecosystems. The Journal of Systems and Software, 85(7):1479– 1494, July 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000805.

Sagar:2014:CMN

[SA14] Vidhu Bhala R. Vidya Sagar and S. Abirami. Conceptual modeling of natural language functional requirements. The Journal of Systems and Software, 88(??):25–41, February 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002379.

Siad:2016:NFI

[SA16] A. Siad and M. Amara. A new framework for implementing identity-based cryptosystems. The Journal of Systems and Software, 118(??):36–48, August 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300346.

Spanos:2018:MTA

[SA18] Georgios Spanos and Lefteris Angelis. A multi-target approach to estimate software vulnerability characteristics and severity scores. The Journal of Systems and Software, 146 (??):152–166, December 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302061.

Sakkopoulos:2010:WPT

[SAA^{+}10] E. Sakkopoulos, D. Antoniou, P. Adamopoulou, N. Tsirakis, and A. Tsakalidis. A web personalizing technique using adaptive data structures: The case of bursts in web visits. The Journal of Systems and Software, 83 (11):2200–2210, November 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Shang:2012:UPD

[SAH12] Weiyi Shang, Bram Adams, and Ahmed E. Hassan. Using Pig as a data preparation language for large-scale mining software repositories studies: an experience report. The Journal of Systems and Software, 85(10):2195–2204, October 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002007.

Shokripour:2015:TBA

[SAKZ15] Ramin Shokripour, John Anvik, Zarinah M. Kasirun, and Sima Zamani. A time-based approach to automatic bug report assignment. The Journal of Systems and Software, 102(??):109–122, April 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002933.

Salman:2017:IML

[Sal17] Hamzeh Eyal Salman.

Identification multi-level frequent usage patterns from APIs. The Journal of Systems and Software, 130(??):42–56, August 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217300869.

Sanchez:2016:AMD

[SÁM^{+}16] Pedro Sánchez, Bárbara Álvarez, José Miguel Morales, Diego Alonso, and Andrés Iborra. An approach to modeling and developing teleoreactive systems considering timing constraints. The Journal of Systems and Software, 117(??):317–333, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300139.

Sanchez:2017:EST

[SÁMI17] Pedro Sánchez, Bárbara Álvarez, Ramón Martínez, and Andrés Iborra. Embedding statecharts into Teleo-Reactive programs to model interactions between agents. The Journal of Systems and Software, 131(??):78–97, September 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0164121217301024.

Sanchez:2012:TRS

[SAMN12] Pedro Sánchez, Diego Alonso, José Miguel Morales, and Pedro Javier Navarro. From Teleo-Reactive specifications to architectural components: a model-driven approach. The Journal of Systems and Software, 85(11):2504–2518, November 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001537.

Santhanam:2016:QOS

[San16] Ganesh Ram Santhanam.

Qualitative optimization in software engineering: a short survey. The Journal of Systems and Software, 111(??):149–156, January 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002022.

Szoke:2017:ESR

[SAN^{+}17] Gábor Szőke, Gábor Antal, Csaba Nagy, Rudolf Ferenc, and Tibor Gyimóthy. Empirical study on refactoring large-scale industrial systems and its effects on maintainability. The Journal of Systems and Software, 129(??):107– 126, July 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301558.

Savolainen:2015:WDY

[SAR15] Paula Savolainen, Jarmo J. Ahonen, and Ita Richardson. When did your project start? — The software supplier’s perspective. The Journal of Systems and Software, 104(??):32–40, June 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000473.

Srikanth:2012:ITE

[SB12] Hema Srikanth and Sean Banerjee. Improving test efficiency through system test prioritization. The Journal of Systems and Software, 85(5):1176– 1187, May 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0164121212000027.

Staahl:2014:MCI

[SB14] Daniel Ståhl and Jan Bosch. Modeling continuous integration practice differences in industry software development. The Journal of Systems and Software, 87(??):48–59, January 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002276.

Salama:2017:AMR

[SB17a] Maria Salama and Rami Bahsoon. Analysing and modelling runtime architectural stability for self-adaptive software. The Journal of Systems and Software, 133 (??):95–112, November 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301620.

Sevcech:2017:RPS

[SB17b] Jakub Sevcech and Maria Bielikova. Repeating patterns as symbols for long time series representation. The Journal of Systems and Software, 127(??):179–194, May 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300772.

Son:2019:LAV

[SB19] Jungmin Son and Rajkumar Buyya. Latencyaware Virtualized Network Function provisioning for distributed edge clouds. The Journal of Systems and Software, 152(??):24–31, June 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300391.

Steghofer:2017:NSB

[SBAH17] Jan-Philipp Steghöfer, Håkan Burden, Hiva Alahyari, and Dominik Haneberg. No silver brick: Opportunities and limitations of teaching Scrum with Lego workshops. The Journal of Systems and Software, 131(??):230–247, September 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0164121217301206.

Soldani:2016:TMA

[SBB^{+}16] Jacopo Soldani, Tobias Binz, Uwe Breitenbücher, Frank Leymann, and Antonio Brogi. ToscaMart: a method for adapting and reusing cloud applications. The Journal of Systems and Software, 113(??):395–406, March 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002903.

Solanki:2019:TKD

[SBDB19] Monika Solanki, Bojan Bozić, Christian Dirschl, and Rob Brennan. Towards a knowledge driven framework for bridging the gap between software and data engineering. The Journal of Systems and Software, 149(??):476–484, March 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302772.

Sellami:2013:CWS

[SBGT13] Mohamed Sellami, Olfa Bouchaala, Walid Gaaloul, and Samir Tata. Communities of Web service registries: Construction and management. The Journal of Systems and Software, 86(3):835–853, March 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212003123.

Sampaio:2019:PSE

[SBT19] Gabriela Sampaio, Paulo Borba, and Leopoldo Teixeira. Partially safe evolution of software product lines. The Journal of Systems and Software, 155(??):17–42, September 2019. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121219300834.

Shao:2017:CCB

[SBZ^{+}17] Yiyang Shao, Weidong Bao, Xiaomin Zhu, Wenhua Xiao, and Jian Wang. Chord: Checkpointbased scheduling using hybrid waiting list in shared clusters. The Journal of Systems and Software, 131(??):22–34, September 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121217300961.

Saxena:2014:SSS

[SC14] Neetesh Saxena and Narendra S. Chaudhari. SecureSMS: a secure SMS protocol for VAS and other applications. The Journal of Systems and Software, 90(??):138–150, April 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214000041.

Singh:2019:CMC

[SC19] David E. Singh and Jesus Carretero. Combining malleability and I/O control mechanisms to enhance the execution of multiple applications. The Journal of Systems and Software, 148(??):21–36, February 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302425.

Srikanth:2016:TCP

[SCC16] Hema Srikanth, Mikaela Cashman, and Myra B. Cohen. Test case prioritization of build acceptance tests for an enterprise cloud application: an industrial case study. The Journal of Systems and Software, 119 (??):122–135, September 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300851.

Sanchez-Carmona:2018:FML

[SCGL^{+}18] Adrián Sánchez-Carmona, Frédéric Guidec, Pascale Launay, Yves Mahéo, and Sergi Robles. Filling in the missing link between simulation and application in opportunistic networking. The Journal of Systems and Software, 142(??):57–72, August 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300694.

Seo:2013:SGD

[SCL13] Jooyoung Seo, Byoungju Choi, and Sihyun Lee. Software generated device exception for more intensive device-related software testing: an industrial field study. The Journal of Systems and Software, 86 (12):3193–3212, December 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002008.

Sabatucci:2015:ALS

[SCMS15] Luca Sabatucci, Mariano Ceccato, Alessandro Marchetto, and Angelo Susi. Ahab’s legs in scenario-based requirements validation: an experiment to study communication mistakes. The Journal of Systems and Software, 109 (??):124–136, November 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001648.

Silva:2013:CAD

[SCO13] Luís A. Bastião Silva, Carlos Costa, and José Luís Oliveira. A common API for delivering services over multi-vendor cloud resources. The Journal of Systems and Software, 86 (9):2309–2317, September 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213001052.

Sabatucci:2015:GOA

[SCS15] Luca Sabatucci, Massimo Cossentino, and Angelo Susi. A goal-oriented approach for representing and using design patterns. The Journal of Systems and Software, 110(??):136–154, December 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121500165X.

Seo:2012:LES

[SCwY12] Jooyoung Seo, Byoungju Choi, and Sueng wan Yang. Lightweight embedded software performance analysis method by kernel hack and its industrial field study. The Journal of Systems and Software, 85(1):28–42, January 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211000781.

Saito:2016:PSR

[SD16a] Yasuhiro Saito and Tadashi Dohi. Predicting software reliability via completely monotone nonparametric estimator with grouped data. The Journal of Systems and Software, 117(??):296– 306, July 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300073.

Schwartz:2016:CER

[SD16b] Amanda Schwartz and Hyunsook Do. Costeffective regression testing through Adaptive Test Prioritization strategies. The Journal of Systems and Software, 115 (??):61–81, May 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121216000169.

Skersys:2016:MBM

[SDB16] Tomas Skersys, Paulius Danenas, and Rimantas Butleris. Modelbased M2M transformations based on dragand-drop actions: Approach and implementation. The Journal of Systems and Software, 122(??):327–341, December 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216301947.

Skersys:2018:ESB

[SDB18] Tomas Skersys, Paulius Danenas, and Rimantas Butleris. Extracting SBVR business vocabularies and business rules from UML use case diagrams. The Journal of Systems and Software, 141(??):111–130, July 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121830061X.

Senapathi:2017:RMS

[SDG17] Mali Senapathi and Meghann L. Drury-Grogan. Refining a model for sustained usage of agile methodologies. The Journal of Systems and Software, 132(??):298–316, October 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217301498.

Shakiba:2010:IID

[SDM10] Mohsen Shakiba, Mohammad Dakhilalian, and Hamid Mala. An improved impossible differential cryptanalysis of Zodiac. The Journal of Systems and Software, 83 (4):702–709, April 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Souza:2013:ESI

[SdSGdMSN^{+}13] Iuri Santos Souza, Gecynalda Soares da Silva Gomes, Paulo Anselmo da Mota Silveira Neto, Ivan do Carmo Machado, Eduardo Santana de Almeida, and Silvio Romero de Lemos Meira. Evidence of software inspection on feature specification for software product lines. The Journal of Systems and Software, 86(5):1172–1190, May 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212003251.

Santos:2019:EEM

[SdSLS^{+}19] Bruno M. Santos, André de S. Landi, Daniel S. Santibáñez, Rafael S. Durelli, and Valter V. de Camargo. Evaluating the extension mechanisms of the knowledge discovery metamodel for aspect-oriented modernizations. The Journal of Systems and Software, 149(??):285–304, March 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302711.

Stolee:2016:CSI

[SED16] Kathryn T. Stolee, Sebastian Elbaum, and Matthew B. Dwyer. Code search with input/ output queries: Generalizing, ranking, and assessment. The Journal of Systems and Software, 116(??):35–48, June 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000874.

Saoud:2016:FBC

[SFMB16] Zohra Saoud, Noura Faci, Zakaria Maamar, and Djamal Benslimane. A fuzzy-based credibility model to assess Web services trust under uncertainty. The Journal of Systems and Software, 122(??):496–506, December 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002162.

Spinellis:2012:OAO

[SG12] Diomidis Spinellis and Vaggelis Giannikas. Organizational adoption of open source software. The Journal of Systems and Software, 85(3):666–682, March 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002512.

Sampaio:2016:ECS

[SG16] Luciano Sampaio and Alessandro Garcia. Exploring context-sensitive data flow analysis for early vulnerability detection. The Journal of Systems and Software, 113(??):337–361, March 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002873.

Sicari:2012:DDD

[SGBCP12] Sabrina Sicari, Luigi Alfredo Grieco, Gennaro Boggia, and Alberto Coen-Porisini. DyDAP: a dynamic data aggregation scheme for privacy aware wireless sensor networks. The Journal of Systems and Software, 85(1):152–166, January 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002068.

Sljivo:2017:MGR

[SGC^{+}17] Irfan Sljivo, Barbara Gallina, Jan Carlson, Hans Hansson, and Stefano Puri. A method to generate reusable safety case argumentfragments from compositional safety analysis. The Journal of Systems and Software, 131 (??):570–590, September 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301273.

Spinnner:2019:OML

[SGEK19] Simon Spinnner, Johannes Grohmann, Simon Eismann, and Samuel Kounev. Online model learning for self-aware computing infrastructures. The Journal of Systems and Software, 147(??):1–16, January 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302188.

Salamah:2012:VTS

[SGK12] Salamah Salamah, Ann Gates, and Vladik Kreinovich. Validated templates for specification of complex LTL formulas. The Journal of Systems and Software, 85(8):1915–1929, August 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000659.

Soares:2013:CAA

[SGMHJ13] Gustavo Soares, Rohit Gheyi, Emerson Murphy-Hill, and Brittany Johnson. Comparing approaches to analyze refactoring activity on software repositories. The Journal of Systems and Software, 86(4):1006–1022, April 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121200297X.

Struck:2013:EOL

[SGO13] Simon Struck, Matthias Güdemann, and Frank Ortmeier. Efficient optimization of large probabilistic models. The Journal of Systems and Software, 86(10):2488– 2501, October 2013. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213000770.

Salvaneschi:2012:COP

[SGP12] Guido Salvaneschi, Carlo Ghezzi, and Matteo Pradella. Contextoriented programming: a software engineering perspective. The Journal of Systems and Software, 85(8):1801–1817, August 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121200074X.

Song:2015:HHB

[SGW^{+}15] Jie Song, Chaopeng Guo, Zhi Wang, Yichan Zhang, Ge Yu, and Jean-Marc Pierson. HaoLap: a Hadoop based OLAP system for big data. The Journal of Systems and Software, 102(??):167– 181, April 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002076.

Seiffert:2017:ACA

[SH17] Dominic Seiffert and Oliver Hummel. Adapting collections and arrays: Another step towards the automated adaptation of object ensembles. The Journal of Systems and Software, 123(??):79–91, January 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121630200X.

Si:2016:RBE

[SHBA^{+}16] Yain-Whar Si, Kin-Kuan Hoi, Robert P. Biuk-Aghai, Simon Fong, and Defu Zhang. Runbased exception prediction for workflows. The Journal of Systems and Software, 113(??):59–75, March 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002496.

Syed:2019:TGB

[SHBC19] Zahid Syed, Jordan Helmick, Sean Banerjee, and Bojan Cukic. Touch gesture-based authentication on mobile devices: the effects of user posture, device size, configuration, and inter-session variability. The Journal of Systems and Software, 149(??):158–173, March 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218302516.

Song:2011:SRS

[SHC^{+}11] Hui Song, Gang Huang, Franck Chauvel, Yingfei Xiong, Zhenjiang Hu, Yanchun Sun, and Hong Mei. Supporting runtime software architecture: a bidirectionaltransformation-based approach. The Journal of Systems and Software, 84 (5):711–723, May 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Su:2016:UBC

[SHGT16] Moon Ting Su, John Hosking, John Grundy, and Ewan Tempero. Usage-based chunking of software architecture information to assist information finding. The Journal of Systems and Software, 122(??):215–238, December 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301637.

Sun:2015:RSB

[SHH^{+}15] Zhoubao Sun, Lixin Han, Wenliang Huang, Xueting Wang, Xiaoqin Zeng, Min Wang, and Hong Yan. Recommender systems based on social networks. The Journal of Systems and Software, 99(??):109–119, January 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002064.

Strode:2012:CCL

[SHHL12] Diane E. Strode, Sid L.

Huff, Beverley Hope, and Sebastian Link. Coordination in co-located agile software development projects. The Journal of Systems and Software, 85(6):1222–1238, June 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212000465.

Shim:2010:IBA

[Shi10] Kyung-Ah Shim. An ID-based aggregate signature scheme with constant pairing computations. The Journal of Systems and Software, 83 (10):1873–1880, October 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Shirazi:2012:FOS

[Shi12]  Farid Shirazi.  Free

and Open Source Software versus Internet content filtering and censorship: a case study. The Journal of Systems and Software, 85(4):920– 931, April 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211002883.

Shim:2017:PME

[Shi17] Hyotaek Shim. PHash: a memory-efficient, highperformance key-value store for large-scale data-intensive applications. The Journal of Systems and Software, 123(??):33–44, January 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301959.

Sasaki:2014:TKQ

[SHN14] Yuya Sasaki, Takahiro Hara, and Shojiro Nishio. Top-k query processing for replicated data in mobile peer to peer networks. The Journal of Systems and Software, 92(??):45–58, June 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002628.

Sobernig:2016:ERD

[SHS16] Stefan Sobernig, Bernhard Hoisl, and Mark Strembeck. Extracting reusable design decisions for UML-based domainspecific languages: a multi-method study. The Journal of Systems and Software, 113(??):140– 172, March 2016. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121215002617.

Satir:2012:CBT

[SI12] Esra Satir and Hakan Isik. A compressionbased text steganography method. The Journal of Systems and Software, 85(10):2385–2394, October 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001379.

Sinha:2017:RBC

[SJ17] Atish P. Sinha and Hemant Jain. Reusing business components and objects for modeling business systems: the influence of decomposition characteristics and analyst experience. The Journal of Systems and Software, 131(??):550– 569, September 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121216301303.

Shyur:2013:DMA

[SJC13] Huan-Jyh Shyur, Chichang Jou, and Keng Chang. A data mining approach to discovering reliable sequential patterns. The Journal of Systems and Software, 86(8):2196–2203, August 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000873.

Suri:2010:SIA

[SJH^{+}10] Neeraj Suri, Arshad Jhumka, Martin Hiller, András Pataricza, Shariful Islam, and Constantin Sârbu. A software integration approach for designing and assessing dependable embedded systems. The Journal of Systems and Software, 83 (10):1780–1800, October 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sanchez:2011:FDH

[SJR^{+}11] Pedro Sánchez, Manuel Jiménez, Francisca Rosique, Bárbara Álvarez, and Andrés Iborra. A framework for developing home automation systems: From requirements to code. The Journal of Systems and Software, 84(6):1008–1021, June 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Stavrinides:2010:SMT

[SK10] Georgios L. Stavrinides and Helen D. Karatza. Scheduling multiple task graphs with end-to-end deadlines in distributed real-time systems utilizing imprecise computations. The Journal of Systems and Software, 83(6):1004–1014, June 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sasa:2011:EAP

[ŠK11] Ana Šaša and Marjan Krisper. Enterprise architecture patterns for business process support analysis. The Journal of Systems and Software, 84 (9):1480–1506, September 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211000689.

Sudevalayam:2013:AAM

[SK13] Sujesha Sudevalayam and Purushottam Kulkarni. Affinity-aware modeling of CPU usage with communicating virtual machines. The Journal of Systems and Software, 86(10):2627–2638, October 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001246.

Stachtiari:2018:CES

[SK18] Emmanouela Stachtiari and Panagiotis Katsaros. Compositional execution semantics for business process verification. The Journal of Systems and Software, 137(??):217–238, March 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302595.

Shabtai:2010:IDM

[SKE10] Asaf Shabtai, Uri Kanonov, and Yuval Elovici. Intrusion detection for mobile devices using the knowledge-based, temporal abstraction method. The Journal of Systems and Software, 83(8): 1524–1537, August 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Syu:2017:TSF

[SKF17] Yang Syu, Jong-Yih Kuo, and Yong-Yi Fanjiang. Time series forecasting for dynamic quality of web services: an empirical study. The Journal of Systems and Software, 134(??):279–303, December 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217302042.

Skianis:2013:IAP

[Ski13] Charalabos Skianis. Introducing automated procedures in 3G network planning and optimization. The Journal of Systems and Software, 86(6):1596–1602, June 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213000307.

Sauvanaud:2018:ADD

[SKK^{+}18a] Carla Sauvanaud, Mohamed Kaâniche, Karama Kanoun, Kahina Lazri, and Guthemberg Da Silva Silvestre. Anomaly detection and diagnosis for cloud services: Practical experiments and lessons learned. The Journal of Systems and Software, 139(??):84– 106, May 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218300256.

Schneider:2018:PAT

[SKK^{+}18b] Kurt Schneider, Jil Klünder, Fabian Kortum, Lisa Handke, Julia Straube, and Simone Kauffeld. Positive affect through interactions in meetings: the role of proactive and supportive statements. The Journal of Systems and Software, 143(??):59–70, September 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121830089X.

Santos:2010:ACD

[SKL10] André L. Santos, Kai Koskimies, and Antónia Lopes. Automating the construction of domainspecific modeling languages for object-oriented frameworks. The Journal of Systems and Software, 83(7):1078–1093, July 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Skopik:2014:SSG

[Sko14] Florian Skopik. The social smart grid: Dealing with constrained energy resources through social coordination. The Journal of Systems and Software, 89(??):3–18, March 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213001064.

Sievi-Korte:2019:SAD

[SKRB19] Outi Sievi-Korte, Ita Richardson, and Sarah Beecham. Software architecture design in global software development: an empirical study. The Journal of Systems and Software, 158(??):??, December 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301748.

Soualhia:2017:TSB

[SKT17] Mbarka Soualhia, Foutse Khomh, and Sofiène Tahar. Task scheduling in big data platforms: a systematic literature review. The Journal of Systems and Software, 134(??):170–189, December 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217301954.

Salmeron:2010:MAR

[SL10] Jose L. Salmeron and Cristina Lopez. A multicriteria approach for risks assessment in ERP maintenance. The Journal of Systems and Software, 83 (10):1941–1953, October 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Shahin:2014:SRS

[SLB14] Mojtaba Shahin, Peng Liang, and Muhammad Ali Babar. A systematic review of software architecture visualization techniques. The Journal of Systems and Software, 94 (??):161–185, August 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214000831.

Sun:2015:SCI

[SLL^{+}15] Xiaobing Sun, Bixin Li, Hareton Leung, Bin Li, and Junwu Zhu. Static change impact analysis techniques: a comparative study. The Journal of Systems and Software, 109(??):137–149, November 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001661.

Suei:2012:SBG

[SLLL12] Pei-Lun Suei, Yung-Feng Lu, Rong-Jhang Liao, and Shi-Wu Lo. A signature-based Grid index design for mainmemory RFID database applications. The Journal of Systems and Software, 85(5):1205–1212, May 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212000222.

Sun:2014:CIA

[SLLL14] Xiaobing Sun, Hareton Leung, Bin Li, and Bixin Li. Change impact analysis and changeability assessment for a change proposal: an empirical study. The Journal of Systems and Software, 96(??):51–60, October 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121400123X.

Su:2017:CID

[SLLY17] Yihsiung Su, Pin Luarn, Yue-Shi Lee, and Show-Jane Yen. Creating an invalid defect classification model using text mining on server development.

The Journal of Systems and Software, 125(??): 197–206, March 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

//www.sciencedirect.
com/science/article/
pii/S0164121216302527.

Smolander:2016:PSS

[SLR16] Kari Smolander, Casper Lassenius, and Matti Rossi. Preface to the special section on software business. The Journal of Systems and Software, 113(??):407, March 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002356.

Shen:2015:LPS

[SLW^{+}15] Wenfeng Shen, Zhaokai Luo, Daming Wei, Weimin Xu, and Xin Zhu. Loadprediction scheduling algorithm for computer simulation of electrocardiogram in hybrid environments. The Journal of Systems and Software, 102(??):182–191, April 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121215000138.

Shao:2012:AKP

[SLZ12] Jun Shao, Peng Liu, and Yuan Zhou. Achieving key privacy without losing CCA security in proxy re-encryption. The Journal of Systems and Software, 85(3):655–665, March 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211002421.

Staron:2016:MMA

[SM16] Mirosław Staron and Wilhelm Meding. MeS-RAM — a method for assessing robustness of measurement programs in large software development organizations and its industrial evaluation. The Journal of Systems and Software, 113(??):76–100, March 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215002368.

Safabahar:2017:NSA

[SM17a] Babak Safabahar and Meghdad Mirabi. A new structure and access mechanism for secure and efficient XML data broadcast in mobile wireless networks. The Journal of Systems and Software, 125(??):119– 132, March 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

//www.sciencedirect.
com/science/article/
pii/S0164121216302394.

Santos:2017:DAI

[SM17b] André L. Santos and Brad A. Myers. Design annotations to improve API discoverability. The Journal of Systems and Software, 126(??):17–33, April 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121630262X.

Staahl:2017:CCI

[SMB17] Daniel Ståhl, Torvald Mårtensson, and Jan Bosch. The continuity of continuous integration: Correlations and consequences. The Journal of Systems and Software, 127(??):150–167, May 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121217300328.

Stachtiari:2018:EVS

[SMK^{+}18] Emmanouela Stachtiari, Anastasia Mavridou, Panagiotis Katsaros, Simon Bliudze, and Joseph Sifakis. Early validation of system requirements and design through correctness-by-construction. The Journal of Systems and Software, 145(??):52–78, November 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121830150X.

Shchapov:2017:TPI

[SMM17] Vladislav A. Shchapov, Aleksei G. Masich, and Grigorii F. Masich. The technology of processing intensive structured dataflow on a supercomputer. The Journal of Systems and Software, 127(??):258–265, May 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216301054.

Shakshuki:2011:CSS

[SMS11] Elhadi M. Shakshuki, Haroon Malik, and Tarek R. Sheltami. A comparative study on simulation vs. real time deployment in wireless sensor networks. The Journal of Systems and Software, 84(1):45– 54, January 2011. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Santos:2018:HDD

[SMSH18] Eddie Antonio Santos, Carson McLean, Christopher Solinas, and Abram Hindle. How does Docker affect energy consumption? Evaluating workloads in and out of Docker containers. The Journal of Systems and Software, 146(??):14–25, December 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301456.

Salah:2012:MSL

[SMZC12] K. Salah, A. Manea, S. Zeadally, and Jose M. Alcaraz Calero. Mitigating starvation of Linux CPU-bound processes in the presence of network I/O. The Journal of Systems and Software, 85(8):1899–1914, August 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212000660.

Stankovic:2013:SSC

[SNDC13] Dragan Stankovic, Vesna Nikolic, Miodrag Djordjevic, and Dac-Buu Cao. A survey study of critical success factors in agile software projects in former Yugoslavia IT companies. The Journal of Systems and Software, 86(6):1663–1678, June 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213000496.

Stievenart:2019:GMR

[SNDD19] Quentin Stiévenart, Jens Nicolay, Wolfgang De Meuter, and Coen De Roover. A general method for rendering static analyses for diverse concurrency models modular. The Journal of Systems and Software, 147(??):17–45, January 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218302206.

SanchezGuinea:2016:SRE

[SNL16] Alejandro Sánchez Guinea, Grégory Nain, and Yves Le Traon. A systematic review on the engineering of software for ubiquitous systems. The Journal of Systems and Software, 118(??):251–276, August 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300553.

Seyedzadeh:2014:RCI

[SNM14] Seyed Mohammad Seyedzadeh, Benyamin Norouzi, and Sattar Mirzakuchaki. RGB color image encryption based on Choquet fuzzy integral. The Journal of Systems and Software, 97(??):128–139, November 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001563.

Sommerville:2013:TCC

[Som13] Ian Sommerville. Teaching cloud computing: a software engineering perspective. The Journal of Systems and Software, 86 (9):2330–2332, September 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000198.

Siebra:2016:TCT

[SOS^{+}16] Clauirton A. Siebra, Rebeka G. Oliveira, Carolyn B. Seaman, Fabio Q. B. Silva, and Andre L. M. Santos. Theoretical conceptualization of TD: a practical perspective. The Journal of Systems and Software, 120(??):219–237, October 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300644.

Saied:2018:IRS

[SOS^{+}18] Mohamed Aymen Saied, Ali Ouni, Houari Sahraoui, Raula Gaikovina Kula, Katsuro Inoue, and David Lo. Improving reusability of software libraries through usage pattern mining. The Journal of Systems and Software, 145 (??):164–179, November 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0164121218301699.

Sutcliffe:2014:EUD

[SP14] Alistair Sutcliffe and George Papamargaritis. End-user development by application-domain configuration. The Journal of Systems and Software, 91(??):85–99, May 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002872.

Sahin:2016:BRA

[SPC16] Cagri Sahin, Lori Pollock, and James Clause. From benchmarks to real apps: Exploring the energy impacts of performance-directed changes. The Journal of Systems and Software, 117(??):307–316, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000893.

Song:2018:PPE

[SPCT18] Zheng Jason Song, Jing Pu, Junjie Cheng, and Eli Tilevich. Performance and programming effort trade-offs of Android persistence frameworks. The Journal of Systems and Software, 146(??):99–111, December 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301705.

Shao:2017:DSA

[SPLW17] Fei Shao, Rong Peng, Han Lai, and Bangchao Wang. DRank: a semiautomated requirements prioritization method based on preferences and dependencies. The Journal of Systems and Software, 126(??):141–156, April 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301911.

Schwartz:2018:IFM

[SPMG18] Amanda Schwartz, Daniel Puckett, Ying Meng, and Gregory Gay. Investigating faults missed by test suites achieving high code coverage. The Journal of Systems and Software, 144(??):106–120, October 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301201.

Santos:2017:SAU

[SPSR17] André L. Santos, Gonçalo Prendi, Hugo Sousa, and Ricardo Ribeiro. Stepwise API usage assistance using n-gram language models. The Journal of Systems and Software, 131(??):461–474, September 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121216300917.

Santos:2015:SRM

[SPTM15] Heliomar Santos, João Felipe Pimentel, Viviane Torres Da Silva, and Leonardo Murta. Software rejuvenation via a multi-agent approach. The Journal of Systems and Software, 104(??): 41–59, June 2015. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121215000412.

Steinberger:2018:CLV

[SRBT18] Michal Steinberger, Iris Reinhartz-Berger, and Amir Tomer. Cross lifecycle variability analysis:

Utilizing requirements and testing artifacts. The Journal of Systems and Software, 143 (??):208–230, September 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300864.

Santos:2018:SRC

[SRJL^{+}18] José Amancio M. Santos, João B. Rocha-Junior, Luciana Carla Lins Prates, Rogeres Santos do Nascimento, Mydiã Falcão Freitas, and Manoel Gomes de Mendonça. A systematic review on the code smell effect. The Journal of Systems and Software, 144(??):450–477, October 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301444.

Shahid:2015:LBB

[SRS15] Mohammad Shahid, Zahid Raza, and Mohammad Sajid. Level based batch scheduling strategy with idle slot reduction under DAG constraints for computational grid. The Journal of Systems and Software, 108(??):110–133, October 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001260.

Scott:2016:TBS

[SRSC16] Ezequiel Scott, Guillermo Rodríguez, Álvaro Soria, and Marcelo Campo. Towards better Scrum learning using learning styles. The Journal of Systems and Software, 111(??):242–253, January 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002265.

Santos:2012:STD

[SRT^{+}12] A. Santos, J. Romero, J. Taibo, C. Rodriguez, and A. Carballal. Selftuning of disk input– output in operating systems. The Journal of Systems and Software, 85(1):77–86, January 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001993.

Sama:2010:MLF

[SRWE10] Michele Sama, David S. Rosenblum, Zhimin Wang, and Sebastian Elbaum. Multi-layer faults in the architectures of mobile, context-aware adaptive applications. The Journal of Systems and Software, 83(6):906–914, June 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Senapathi:2012:UPA

[SS12] Mali Senapathi and Ananth Srinivasan. Understanding post-adoptive agile usage: an exploratory cross-case analysis. The Journal of Systems and Software, 85(6):1255–1268, June 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212000489.

Sun:2013:HPP

[SS13] Zhe Sun and Jun Shen. A high performance peer to cloud and peer model augmented with hierarchical secure communications. The Journal of Systems and Software, 86(7):1790–1796, July 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002531.

Siqueira:2014:TEM

[SS14a] Fábio Levy Siqueira and Paulo Sérgio Muniz Silva. Transforming an enterprise model into a use case model in business process systems. The Journal of Systems and Software, 96(??):152–171, October 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001447.

Sor:2014:MLD

[SS14b] Vladimir Sor and Satish Narayana Srirama. Memory leak detection in Java: Taxonomy and classification of approaches. The Journal of Systems and Software, 96(??):139–151, October 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001423.

Smith:2015:ISC

[SS15] Jim Smith and Chris Simons. The influence of search components and problem characteristics in early life cycle class modelling. The Journal of Systems and Software, 103(??):440–451, May 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002659.

Schaefer:2017:ISI

[SS17] Ina Schaefer and Ioannis Stamelos. Introduction to the Special Issue on “International Conference on Software Reuse 2015”. The Journal of Systems and Software, 131 (??):323–324, September 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301401.

Sharma:2018:SSS

[SS18] Tushar Sharma and Diomidis Spinellis. A survey on software smells. The Journal of Systems and Software, 138(??):158– 173, April 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0164121217303114.

Suomalainen:2011:SPR

[SSAS11] Tanja Suomalainen, Outi Salo, Pekka Abrahamsson, and Jouni Similä. Software product roadmapping in a volatile business environment. The Journal of Systems and Software, 84(6):958–975, June 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Stray:2016:DSM

[SSD16] Viktoria Stray, Dag I. K.

Sjøberg, and Tore Dybå. The daily stand-up meeting: a grounded theory study. The Journal of Systems and Software, 114(??):101–124, April 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000066.

Santos:2015:USF

[SSF15] Alan R. Santos, Afonso Sales, and Paulo Fernandes. Using SAN formalism to evaluate Follow-The-Sun project scenarios. The Journal of Systems and Software, 100(??):182–194, February 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002386.

Sierra:2019:SSA

[SSK19] Giancarlo Sierra, Emad Shihab, and Yasutaka Kamei. A survey of self-admitted technical debt. The Journal of Systems and Software, 152(??):70–82, June 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300457.

Schloegel:2016:RAS

[SSMvD16] Uta Schloegel, Sebastian Stegmann, Alexander Maedche, and Rolf van Dick. Reducing age stereotypes in software development: the effects of awareness- and cooperation-based diversity interventions. The Journal of Systems and Software, 121(??):1–15, November 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301352.

Sioutas:2015:DPS

[SSP^{+}15] S. Sioutas, E. Sakkopoulos, A. Panaretos, D. Tsoumakos, P. Gerolymatos, G. Tzimas, and Y. Manolopoulos. D-P2P-Sim+: a novel distributed framework for P2P protocols performance testing. The Journal of Systems and Software, 100(??):211–233, February 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002416.

Silhavy:2017:ASR

[SSP17] Radek Silhavy, Petr Silhavy, and Zdenka Prokopova. Analysis and selection of a regression model for the Use Case Points method using a stepwise approach. The Journal of Systems and Software, 125(??):1– 14, March 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

//www.sciencedirect.
com/science/article/
pii/S016412121630231X.

Schroder:2018:AEC

[SSR18] Sandra Schröder, Mohamed Soliman, and Matthias Riebisch. Architecture enforcement concerns and activities

— an expert study. The Journal of Systems and Software, 145(??):??, November 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301614.

Shatnawi:2017:RSP

[SSS17] Anas Shatnawi, Abdelhak-Djamel Seriai, and Houari Sahraoui. Recovering software product line architecture of a family of object-oriented product variants. The Journal of Systems and Software, 131(??):325–346, September 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121216301327.

Shatnawi:2017:RER

[SSSA17] Anas Shatnawi, Abdelhak-Djamel Seriai, Houari Sahraoui, and Zakarea Alshara. Reverse engineering reusable software components from object-oriented APIs. The Journal of Systems and Software, 131 (??):442–460, September 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121630098X.

Schatten:2016:RSA

[SST16] Markus Schatten, Jurica Seva, and Igor Tomicić. A roadmap for scalable agent organizations in the Internet of Everything. The Journal of Systems and Software, 115(??):31–41, May 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000170.

Samaras:2011:ATS

[ST11] Christos V. Samaras and Vassilis Tsaoussidis. Adjusting transport segmentation policy of DTN Bundle Protocol under synergy with lower layers. The Journal of Systems and Software, 84(2): 226–237, February 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sbattella:2013:NSI

[ST13] Licia Sbattella and Roberto Tedesco. A novel semantic information retrieval system based on a threelevel domain model. The Journal of Systems and Software, 86(5):1426– 1452, May 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213000137.

Stamelos:2010:SPM

[Sta10] Ioannis Stamelos. Software project management anti-patterns. The Journal of Systems and Software, 83(1):52–59, January 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Stavru:2014:CER

[Sta14] Stavros Stavru. A critical examination of recent industrial surveys on agile method usage. The Journal of Systems and Software, 94(??):87–97, August 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214000764.

Sari:2019:SLR

[STA19] Asli Sari, Ayse Tosun, and Gülfem Isiklar Alptekin. A systematic literature review on crowdsourcing in software engineering. The Journal of Systems and Software, 153(??):200– 219, July 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300779.

Santos:2019:EST

[STS^{+}19] Joanna C. S. Santos, Katy Tarrit, Adriana Sejfia, Mehdi Mirakhorli, and Matthias Galster. An empirical study of tactical vulnerabilities. The Journal of Systems and Software, 149(??):263–284, March 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302322.

Singh:2012:IBP

[SV12] Harendra Singh and Girraj Kumar Verma. IDbased proxy signature scheme with message recovery. The Journal of Systems and Software, 85(1):209–214, January 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002159.

Sneed:2019:RIL

[SV19] Harry Sneed and Chris Verhoef. Re-implementing a legacy system. The Journal of Systems and Software, 155(??):162– 184, September 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301050.

Silva:2019:CCP

[SVM19] Luciana L. Silva, Marco Tulio Valente, and Marcelo A. Maia. Co-change patterns: a large scale empirical study. The Journal of Systems and Software, 152(??):196–214, June 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219300597.

Salfner:2010:ASA

[SW10] F. Salfner and K. Wolter. Analysis of service availability for time-triggered rejuvenation policies. The Journal of Systems and Software, 83 (9):1579–1590, September 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Saiedian:2019:ASE

[SW19] Hossein Saiedian and Hironori Washizak. Advancing software engineering education: New practices and perspectives. The Journal of Systems and Software, 147(??):104–105, January 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302048.

Smite:2013:OIS

[SWA^{+}13] Darja Smite, Claes Wohlin, Aybüke Aurum, Ronald Jabangwe, and Emil Numminen. Offshore insourcing in software development: Structuring the decision-making process. The Journal of Systems and Software, 86(4):1054–1067, April 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002816.

Sun:2016:RQO

[SWES16] Yu Sun, Jules White, Sean Eade, and Douglas C. Schmidt. ROAR: a QoS-oriented modeling framework for automated cloud resource allocation and optimization. The Journal of Systems and Software, 116(??):146– 161, June 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001715.

Sun:2011:SUP

[SXYM11] Ying Sun, Chunxiang Xu, Yong Yu, and Yi Mu. Strongly unforgeable proxy signature scheme secure in the standard model. The Journal of Systems and Software, 84 (9):1471–1479, September 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211000653.

Si:2014:EMD

[SXYW14] Guannan Si, Jing Xu, Jufeng Yang, and Shuo Wen. An evaluation model for dependability of Internet-scale software on basis of Bayesian Networks and trustworthiness. The Journal of Systems and Software, 89(??):63–75, March 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002306.

Seceleanu:2016:GEF

[SY16a] Cristina Seceleanu and Kenichi Yoshida. Guest editorial foreword. The Journal of Systems and Software, 121(??):125, November 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301492.

Siewe:2016:PPT

[SY16b] François Siewe and Hongji Yang. Privacy protection by typing in ubiquitous computing systems. The Journal of Systems and Software, 120(??):133–153, October 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301315.

Salifu:2012:AMS

[SYBN12] Mohammed Salifu, Yijun Yu, Arosha K. Bandara, and Bashar Nuseibeh. Analysing monitoring and switching problems for adaptive systems. The Journal of Systems and Software, 85 (12):2829–2839, December 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002257.

Su:2017:HSO

[SYT^{+}17] Qianqian Su, Jia Yu, Chengliang Tian, Hanlin Zhang, and Rong Hao. How to securely outsource the inversion modulo a large composite number. The Journal of Systems and Software, 129(??):26–34, July 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217300705.

Sun:2017:EDR

[SYXL17] Xiaobing Sun, Hui Yang, Xin Xia, and Bin Li. Enhancing developer recommendation with supplementary information via mining historical commits. The Journal of Systems and Software, 134(??):355–368, December 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0164121217302091.

Shahriar:2011:TCA

[SZ11] Hossain Shahriar and Mohammad Zulkernine. Taxonomy and classification of automatic monitoring of program security vulnerability exploitations. The Journal of Systems and Software, 84(2):250–269, February 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Stroele:2013:GLA

[SZS13] Victor Ströele, Geraldo Zimbrão, and Jano M. Souza. Group and link analysis of multirelational scientific social networks. The Journal of Systems and Software, 86(7):1819–1830, July 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000472.

Song:2016:MLB

[SZW^{+}16] Qinbao Song, Xiaoyan Zhu, Guangtao Wang, Heli Sun, He Jiang, Chenhao Xue, Baowen Xu, and Wei Song. A machine learning based software process model recommendation method. The Journal of Systems and Software, 118(??):85–100, August 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300425.

Triantafyllidis:2016:PAN

[TAB^{+}16] Konstantinos Triantafyllidis, Waqar Aslam, Egor Bondarev, Johan J. Lukkien, and Peter H. N. de With. ProMARTES: Accurate network and computation delay prediction for componentbased distributed systems. The Journal of Systems and Software, 117(??):450–470, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300188.

Teixeira:2017:MAC

[TAF^{+}17] Sergio Teixeira, Bruno Alves Agrizzi, José Gonçalves Pereira Filho, Silvana Rossetto, and Roquemar de Lima Baldam. Modeling and automatic code generation for wireless sensor network applications using model-driven or business process approaches: a systematic mapping study. The Journal of Systems and Software, 132(??):50–71, October 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301255.

Tang:2010:CSA

[TAJ^{+}10] Antony Tang, Paris Avgeriou, Anton Jansen, Rafael Capilla, and Muhammad Ali Babar. A comparative study of architecture knowledge management tools. The Journal of Systems and Software, 83(3):352–370, March 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tahvili:2018:ETE

[TAS^{+}18] Sahar Tahvili, Wasif Afzal, Mehrdad Saadatmand, Markus Bohlin, and Sharvathul Hasan Ameerjan. ESPRET: a tool for execution time estimation of manual test cases. The Journal of Systems and Software, 146(??):26–41, December 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121218301778.

Tom:2013:ETD

[TAV13] Edith Tom, Aybüke Aurum, and Richard Vidgen. An exploration of technical debt. The Journal of Systems and Software, 86(6):1498–1516, June 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213000022.

Thurimella:2013:MMA

[TB13] Anil Kumar Thurimella and Bernd Brügge. A mixed-method approach for the empirical evaluation of the issue-based variability modeling. The Journal of Systems and Software, 86(7):1831– 1849, July 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213000186.

Totaro:2016:IHP

[TBC^{+}16] G. Totaro, M. Bernaschi, G. Carbone, M. Cianfriglia, and A. Di Marco. ISODAC: a high performance solution for indexing and searching heterogeneous data. The Journal of Systems and Software, 118(??):115–133, August 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002678.

Teixeira:2013:SCC

[TBG13] Leopoldo Teixeira, Paulo Borba, and Rohit Gheyi. Safe composition of configuration knowledgebased software product lines. The Journal of Systems and Software, 86(4):1038–1053, April 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121200307X.

Tichy:2017:E

[TBG17] Matthias Tichy, Jan Bosch, and Michael Goedicke. Editorial. The Journal of Systems and Software, 123 (??):173–175, January 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301741.

Tang:2018:ISD

[TBSvdW18] Antony Tang, Floris Bex, Courtney Schriek, and Jan Martijn E. M. van der Werf. Improving software design reasoning — a reminder card approach. The Journal of Systems and Software, 144(??):22–40, October 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218301043.

Tsantalis:2010:IRO

[TC10] Nikolaos Tsantalis and Alexander Chatzigeorgiou. Identification of refactoring opportunities introducing polymorphism. The Journal of Systems and Software, 83 (3):391–404, March 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tsantalis:2011:IEM

[TC11] Nikolaos Tsantalis and Alexander Chatzigeorgiou. Identification of extract method refactoring opportunities for the decomposition of methods. The Journal of Systems and Software, 84(10):1757–1782, October 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001191.

Tsao:2012:SHL

[TC12] Shiao-Li Tsao and Jian Jhen Chen. SEProf: a highlevel software energy profiling tool for an embedded processor enabling power management functions. The Journal of Systems and Software, 85(8):1757–1769, August 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000775.

Tsai:2016:CIS

[TC16a] Chih-Fong Tsai and Fu-Yu Chang. Combining instance selection for better missing value imputation. The Journal of Systems and Software, 122(??):63–71, December 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301583.

Tsai:2016:TTS

[TC16b] Ting-Hao Tsai and Ya-Shu Chen. Thermalthrottling server: a thermal-aware real-time task scheduling framework for three-dimensional multicore chips. The Journal of Systems and Software, 112(??):11–25, February 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002319.

Tsai:2012:SSE

[TCCH12] Tsai-Hsuan Tsai, Hsien-Tsung Chang, Yi-Ming Chang, and Gung-Shiung Huang. Sharetouch: a system to enrich social network experiences for the elderly. The Journal of Systems and Software, 85(6):1363–1369, June 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121200009X.

Tsai:2014:EIS

[TCK14] Chih-Fong Tsai, Zong-Yao Chen, and Shih-Wen Ke. Evolutionary instance selection for text classification. The Journal of Systems and Software, 90(??):104–113, April 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214000077.

Tuma:2018:TAS

[TCS18] K. Tuma, G. Calikli, and R. Scandariato. Threat analysis of software systems: a systematic literature review. The Journal of Systems and Software, 144(??):275–294, October 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301304.

Teodoro:2016:CSE

[TdCAF16] Silvana Teodoro, Andriele Busatto do Carmo, Daniel Couto Adornes, and Luiz Gustavo Fernandes. A comparative study of energy-aware scheduling algorithms for computational grids. The Journal of Systems and Software, 117(??):153– 165, July 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000509.

Turner:2014:DSP

[TDW^{+}14] Hamilton Turner, Brian Dougherty, Jules White, Russell Kegley, Jonathan Preston, Douglas C. Schmidt, and Aniruddha Gokhale. DRE system performance optimization with the SMACK cache efficiency metric. The Journal of Systems and Software, 98(??):25–43, December 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001836.

Tibermacine:2010:FLA

[TFS10]  Chouki  Tibermacine,

Régis Fleurquin, and Salah Sadou. A family of languages for architecture constraint specification. The Journal of Systems and Software, 83(5): 815–831, May 2010. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Trivedi:2010:MDC

[TG10] Kishor S. Trivedi and Sachin Garg. In memoriam: Dr. Chandra Kintala. The Journal of Systems and Software, 83(9): 1555, September 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tao:2017:BCB

[TG17] Chuanqi Tao and Jerry Gao. On building a cloud-based mobile testing infrastructure service system. The Journal of Systems and Software, 124(??):39–55, February 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216302230.

Tichy:2017:RCS

[TGBF17] Matthias Tichy, Michael Goedicke, Jan Bosch, and Brian Fitzgerald. Rapid continuous software engineering. The Journal of Systems and Software, 133(??):159, November 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301887.

Trubiani:2017:ETU

[TGE17] Catia Trubiani, Achraf Ghabi, and Alexander Egyed. Exploiting traceability uncertainty between software architectural models and extrafunctional results. The Journal of Systems and Software, 125(??):15– 34, March 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216302369.

Tian:2019:GAB

[TGKL19] Tian Tian, Dunwei Gong, Fei-Ching Kuo, and Huai Liu. Genetic algorithm based test data generation for MPI parallel programs with blocking communication. The Journal of Systems and Software, 155 (??):130–144, September 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219300810.

Torres:2011:SMD

[TGP11] Alexandre Torres, Renata Galante, and Marcelo S. Pimenta. A synergistic model-driven approach for persistence modeling with UML. The Journal of Systems and Software, 84(6):942–957, June 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tso:2012:SSC

[THS12] Raylin Tso, Xinyi Huang, and Willy Susilo. Strongly secure certificateless short signatures. The Journal of Systems and Software, 85(6):1409–1417, June 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000039.

Tsai:2010:DSA

[THWC10] Ching-Hong Tsai, Kuo-Chan Huang, Feng-Jian Wang, and Chun-Hao Chen. A distributed server architecture supporting dynamic resource provisioning for BPMoriented workflow management systems. The Journal of Systems and Software, 83(8):1538– 1552, August 2010. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tawosi:2015:ASD

[TJH15] Vali Tawosi, Saeed Jalili, and Seyed Mohammad Hossein Hasheminejad. Automated software design using ant colony optimization with semantic network support. The Journal of Systems and Software, 109(??):1–17, November 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215001442.

Taherizadeh:2018:MSA

[TJT^{+}18] Salman Taherizadeh, Andrew C. Jones, Ian Taylor, Zhiming Zhao, and Vlado Stankovski. Monitoring self-adaptive applications within edge computing frameworks: a state-of-the-art review. The Journal of Systems and Software, 136(??):19–38, February 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121730256X.

Thabit:2014:RRW

[TK14] Rasha Thabit and Bee Ee Khoo. Robust reversible watermarking scheme using Slantlet transform matrix. The Journal of Systems and Software, 88(??):74–86, February 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213002380.

Trubiani:2014:GBH

[TKCR14] Catia Trubiani, Anne Koziolek, Vittorio Cortellessa, and Ralf Reussner. Guilt-based handling of software performance antipatterns in Palladio architectural models. The Journal of Systems and Software, 95(??):141–165, September 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001010.

Tschersich:2011:TPE

[TKH^{+}11] Markus Tschersich, Christian Kahl, Stephan Heim, Stephen Crane, Katja Böttcher, Ioannis Krontiris, and Kai Rannenberg. Towards privacyenhanced mobile communities — architecture, concepts and user trials. The Journal of Systems and Software, 84 (11):1947–1960, November 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001609.

Tchamgoue:2015:PAS

[TKJ15] Guy Martin Tchamgoue, Kyong Hoon Kim, and Yong-Kee Jun. Poweraware scheduling of compositional real-time frameworks. The Journal of Systems and Software, 102(??):58–71, April 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214002891.

Tchamgoue:2016:EBD

[TKJ16] Guy Martin Tchamgoue, Kyong Hoon Kim, and Yong-Kee Jun. EventHealer: Bypassing data races in event-driven programs. The Journal of Systems and Software, 118 (??):208–220, August 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216300012.

Tchamgoue:2013:CRT

[TKJL13] Guy Martin Tchamgoue, Kyong Hoon Kim, Yong-Kee Jun, and Wan Yeon Lee. Compositional real-time scheduling framework for periodic reward-based task model. The Journal of Systems and Software, 86(6):1712–1724, June 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000526.

Thum:2019:FOC

[TKK^{+}19] Thomas Thüm, Alexander Knüppel, Stefan Krüger, Stefanie Bolle, and Ina Schaefer. Featureoriented contract composition. The Journal of Systems and Software, 152(??):83–107, June 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300044.

Tripathi:2018:ARE

[TKP^{+}18] Nirnaya Tripathi, Eriks Klotins, Rafael Prikladnicki, Markku Oivo, Leandro Bento Pompermaier, Arun Sojan Kudakacheril, Michael Unterkalmsteiner, Kari Liukkunen, and Tony Gorschek. An anatomy of requirements engineering in software startups using multi-vocal literature and case survey. The Journal of Systems and Software, 146(??):130–151, December 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121218301729.

Talaei-Khoei:2011:PBA

[TKSRP11] Amir Talaei-Khoei, Terje Solvoll, Pradeep Ray, and Nandan Parameshwaran. Policy-based Awareness Management (PAM): Case study of a wireless communication system at a hospital. The Journal of Systems and Software, 84(10):1791–1805, October 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211001245.

Taba:2017:ESU

[TKZW17] Seyyed Ehsan Salamati Taba, Iman Keivanloo, Ying Zou, and Shaohua Wang. An exploratory study on the usage of common interface elements in Android applications. The Journal of Systems and Software, 131 (??):491–504, September 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216301170.

Tang:2014:SAR

[TL14] Antony Tang and Man F.

Lau. Software architecture review by association. The Journal of Systems and Software, 88(??):87–101, February 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002409.

Treude:2018:UEG

[TLA18] Christoph Treude, Larissa Leite, and Maurício Aniche. Unusual events in GitHub repositories. The Journal of Systems and Software, 142 (??):237–247, August 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300876.

Torre:2018:SIC

[TLGE18] Damiano Torre, Yvan Labiche, Marcela Genero, and Maged Elaasar. A systematic identification of consistency rules for UML diagrams. The Journal of Systems and Software, 144 (??):121–142, October 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301249.

Tajmajer:2016:NPP

[TLK^{+}16a] Tomasz Tajmajer, Spyros Lalis, Manos Koutsoubelias, Aleksander Pruszkowski, Jarosław Domaszewicz, Michele Nati, and Alexander Gluhak. Node/ proxy portability: Designing for the two lives of your next WSAN middleware. The Journal of Systems and Software, 117(??):366–383, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300036.

Tsai:2016:BDM

[TLK16b] Chih-Fong Tsai, Wei-Chao Lin, and Shih-Wen Ke. Big data mining with parallel computing: a comparison of distributed and MapReduce methodologies. The Journal of Systems and Software, 122(??):83–92, December 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0164121216301625.

Tsaur:2012:ESM

[TLL12] Woei-Jiunn Tsaur, Jia-Hong Li, and Wei-Bin Lee. An efficient and secure multi-server authentication scheme with key agreement. The Journal of Systems and Software, 85(4):876–882, April 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002810.

Tsai:2013:ZWS

[TLL13] Hung-Hsu Tsai, Yen-Shou Lai, and Shih-Che Lo. A zero-watermark scheme with geometrical invariants using SVM and PSO against geometrical attacks for image protection. The Journal of Systems and Software, 86(2):335–348, February 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002440.

Tan:2010:CQA

[TLWS10] Zijing Tan, Chengfei Liu, Wei Wang, and Baile Shi. Consistent query answers from virtually integrated XML data. The Journal of Systems and Software, 83(12):2566–2578, December 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tian:2016:ETR

[TLZ^{+}16] Zhenzhou Tian, Ting Liu, Qinghua Zheng, Ming Fan, Eryue Zhuang, and Zijiang Yang. Exploiting thread-related system calls for plagiarism detection of multithreaded programs. The Journal of Systems and Software, 119 (??):136–148, September 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300838.

Tuli:2019:FBB

[TMTB19] Shreshth Tuli, Redowan Mahmud, Shikhar Tuli, and Rajkumar Buyya. FogBus: a blockchainbased lightweight framework for edge and fog computing. The Journal of Systems and Software, 154(??):22–36, August 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300822.

Tantisuwankul:2019:TAC

[TNK^{+}19] Jirateep Tantisuwankul, Yusuf Sulistyo Nugroho, Raula Gaikovina Kula, Hideaki Hata, Arnon Rungsawang, Pattara Leelaprute, and Kenichi Matsumoto. A topological analysis of communication channels for knowledge sharing in contemporary GitHub projects. The Journal of Systems and Software, 158(??):??, December 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301906.

Torrente:2013:SHB

[TPGdS13] M. Carmen Suárez Torrente, A. Belén Martínez Prieto, Darío Alvarez Gutiérrez, and M. Elena Alva de Sagastegui. Sirius: a heuristic-based framework for measuring web usability adapted to the type of website. The Journal of Systems and Software, 86(3):649–663, March 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212002993.

Tsougenis:2012:PEM

[TPKT12] E. D. Tsougenis, G. A. Papakostas, D. E. Koulouriotis, and V. D. Tourassis. Performance evaluation of moment-based watermarking methods: a review. The Journal of Systems and Software, 85(8):1864–1884, August 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212000684.

Tsirakis:2017:LSO

[TPTV17] Nikos Tsirakis, Vasilis Poulopoulos, Panagiotis Tsantilas, and Iraklis Varlamis. Large scale opinion mining for social, news and blog data. The Journal of Systems and Software, 127(??):237– 248, May 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216300814.

Trainer:2018:BGB

[TR18] Erik H. Trainer and David F. Redmiles. Bridging the gap between awareness and trust in globally distributed software teams. The Journal of Systems and Software, 144(??):328–341, October 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301250.

Talebi:2019:PPM

[TS19] Mohammad Mehdi Talebi and Mohsen Sharifi. PARS: a parallel model for scaled processing of complex events. The Journal of Systems and Software, 155(??):1–16, September 2019. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121219301074.

Toosi:2019:EAS

[TSCB19] Adel Nadjaran Toosi, Jungmin Son, Qinghua Chi, and Rajkumar Buyya. ElasticSFC: Auto-scaling techniques for elastic service function chaining in network functions virtualizationbased clouds. The Journal of Systems and Software, 152(??):108–119, June 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219300421.

Tang:2011:MMA

[TSL^{+}11] Xuehai Tang, Bing Sun, Ruilin Li, Chao Li, and Juhua Yin. A meet-in-the-middle attack on reduced-round ARIA. The Journal of Systems and Software, 84(10):1685–1692, October 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121100104X.

Tang:2011:IDC

[TSLL11] Xuehai Tang, Bing Sun, Ruilin Li, and Chao Li. Impossible differential cryptanalysis of 13-round CLEFIA-128. The Journal of Systems and Software, 84(7):1191–1196, July 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Troya:2018:AIL

[TSRC18] Javier Troya, Sergio Segura, and Antonio Ruiz-Cortés. Automated inference of likely metamorphic relations for model transformations.

The Journal of Systems and Software, 136 (??):188–208, February 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217300870.

Tsioliaridou:2010:FCN

[TT10] Ageliki Tsioliaridou and Vassilis Tsaoussidis. Fast convergence to network fairness. The Journal of Systems and Software, 83 (5):745–762, May 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tung:2013:NAC

[TT13] Yuan-Hsin Tung and Shian-Shyong Tseng. A novel approach to collaborative testing in a crowdsourcing environment. The Journal of Systems and Software, 86(8):2143–2153, August 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000782.

See corrigendum [TTT14].

Tibermacine:2015:PIR

[TTC15] Okba Tibermacine, Chouki Tibermacine, and Foudil Cherif. A process to identify relevant substitutes for healing failed WS- * orchestrations. The Journal of Systems and Software, 104(??):1–16, June 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000436.

Tibermacine:2018:ERN

[TTC18] Okba Tibermacine, Chouki Tibermacine, and Foudil Cherif. Estimating the reputation of newcomer web services using a regression-based method. The Journal of Systems and Software, 145 (??):112–124, November 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301626.

Tsai:2010:RLI

[TTL10] H.-H. Tsai, H.-C. Tseng, and Y.-S. Lai. Robust lossless image watermarking based on αtrimmed mean algorithm and support vector machine. The Journal of Systems and Software, 83(6):1015–1028, June 2010. CODEN JSSODM.

ISSN 0164-1212 (print), 1873-1228 (electronic).

Trappey:2013:SLM

[TTL^{+}13] Charles V. Trappey, Amy J. C. Trappey, Gilbert Y. P. Lin, W. T. Lee, and Ta-Hui Yang. SETZ logistics models and system framework for manufacturing and exporting large engineering assets. The Journal of Systems and Software, 86(7):1797–1805, July 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002737.

Tahir:2013:SRF

[TTM13] Abbas Tahir, Davide Tosi, and Sandro Morasca. A systematic review on the functional testing of Semantic Web services. The Journal of Systems and Software, 86 (11):2877–2889, November 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001659.

Tuzun:2019:AIA

[TTMI19] Eray Tüzün, Bedir Tekinerdogan, Yagup Macit, and Kürsat Ince. Adopting integrated application lifecycle management within a largescale software company: an action research approach. The Journal of Systems and Software, 149(??):63–82, March 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302565.

Torchiano:2013:RBP

[TTR^{+}13] Marco Torchiano, Federico Tomassetti, Filippo Ricca, Alessandro Tiso, and Gianna Reggio. Relevance, benefits, and problems of software modelling and model driven techniques — a survey in the Italian industry. The Journal of Systems and Software, 86(8):2110–2126, August 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000824.

Tung:2014:CSN

[TTT14] Yuan-Hsin Tung, Shian-Shyong Tseng, and Wei-Tek Tsai. Corrigendum to “A novel approach to collaborative testing in a crowdsourcing environment” in the Journal of Systems and Software 86 (2013) 2143–2153. The Journal of Systems and Software, 87(??):137, January 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002161.

See [TT13].

Terra:2018:JNH

[TVMS18] Ricardo Terra, Marco Tulio Valente, Sergio Miranda, and Vitor Sales. JMove: a novel heuristic and tool to detect move method refactoring opportunities. The Journal of Systems and Software, 138(??):19–36, April 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302960.

Tu:2019:ASB

[TXCX19] Jingxuan Tu, Xiaoyuan Xie, Tsong Yueh Chen, and Baowen Xu. On the analysis of spectrum based fault localization using hitting sets. The Journal of Systems and Software, 147(??):106–123, January 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218302231.

Tian:2012:LFR

[TXLC12] Wanyong Tian, Chun Jason Xue, Minming Li, and Enhong Chen. Loop fusion and reordering for register file optimization on stream processors. The Journal of Systems and Software, 85(7):1673–1681, July 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212000453.

Tafsiri:2018:CDA

[TY18] Seyedeh Aso Tafsiri and Saleh Yousefi. Combinatorial double auctionbased resource allocation mechanism in cloud computing market. The Journal of Systems and Software, 137(??):322–334, March 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0164121217302819.

Tang:2012:KCU

[TZ12] Xiaoyu Tang and Qingtian Zeng. Keyword clustering for user interest profiling refinement within paper recommender systems. The Journal of Systems and Software, 85(1):87–101, January 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001981.

Tekinerdogan:2019:SIA

[TZB19] Bedir Tekinerdogan, Uwe Zdun, and M. Ali Babar. Special issue on architecting for hyper connectivity and hyper virtualization. The Journal of Systems and Software, 149(??):531–532, March 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302802.

Ullah:2019:ATB

[UB19] Faheem Ullah and Muhammad Ali Babar. Architectural tactics for big data cybersecurity analytics systems: a review. The Journal of Systems and Software, 151(??):81– 118, May 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300172.

Unphon:2010:SAA

[UD10] Hataichanok Unphon and Yvonne Dittrich. Software architecture awareness in long-term software product evolution. The Journal of Systems and Software, 83 (11):2211–2226, November 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Unterkalmsteiner:2015:ARE

[UGFK15] Michael Unterkalmsteiner, Tony Gorschek, Robert Feldt, and Eriks Klotins. Assessing requirements engineering and software test alignment — five case studies. The Journal of Systems and Software, 109(??):62–77, November 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001508.

Usman:2017:PLM

[UIK17] Muhammad Usman, Muhammad Zohaib Iqbal, and Muhammad Uzair Khan. A product-line modeldriven engineering approach for generating feature-based mobile applications. The Journal of Systems and Software, 123(??):1–32, January 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301972.

Umer:2019:SBA

[ULS19] Qasim Umer, Hui Liu, and Yasir Sultan. Sentiment based approval prediction for enhancement reports. The Journal of Systems and Software, 155(??):57–69, September 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301104.

Ullah:2010:DSM

[URG10] Muhammad Irfan Ullah, Günther Ruhe, and Vahid Garousi. Decision support for moving from a single product to a product portfolio in evolving software systems. The Journal of Systems and Software, 83 (12):2496–2512, December 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ulutas:2011:MIS

[UUN11] Mustafa Ulutas, Güzin Ulutas, and Vasif V. Nabiyev. Medical image security and EPR hiding using Shamir’s secret sharing scheme. The Journal of Systems and Software, 84(3):341–353, March 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ulutas:2013:ISI

[UUN13] Mustafa Ulutas, Güzin Ulutas, and Vasif V. Nabiyev. Invertible secret image sharing for gray level and dithered cover images. The Journal of Systems and Software, 86(2):485–500, February 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212002701.

Uzzafer:2013:SMS

[Uzz13] Masood Uzzafer. A simulation model for strategic management process of software projects. The Journal of Systems and Software, 86(1):21–37, January 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121200177X.

Valenca:2017:TPE

[VA17] George Valença and Carina Alves. A theory of power in emerging software ecosystems formed by small-to-medium enterprises. The Journal of Systems and Software, 134(??):76–104, December 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301863.

vanAngeren:2016:CWA

[vAAJ16] Joey van Angeren, Carina Alves, and Slinger Jansen. Can we ask you to collaborate? Analyzing app developer relationships in commercial platform ecosystems. The Journal of Systems and Software, 113(??):430–445, March 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215002502.

Valenca:2018:SMP

[VAJ18] George Valença, Carina Alves, and Slinger Jansen. Strategies for managing power relationships in software ecosystems. The Journal of Systems and Software, 144(??):478–500, October 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218301468.

Vallejo:2010:MAM

[VAM^{+}10] D. Vallejo, J. Albusac, J. A. Mateos, C. Glez-Morcillo, and L. Jimenez. A modern approach to multiagent development. The Journal of Systems and Software, 83(3):467– 484, March 2010. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Verner:2014:FMS

[VBC^{+}14] J. M. Verner, M. A. Babar, N. Cerpa, T. Hall, and S. Beecham. Factors that motivate software engineering teams: a four country empirical study. The Journal of Systems and Software, 92(??):115–127, June 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121400020X.

Venters:2018:SSR

[VCB^{+}18] Colin C. Venters, Rafael Capilla, Stefanie Betz, Birgit Penzenstadler, Tom Crick, Steve Crouch, Elisa Yumi Nakagawa, Christoph Becker, and Carlos Carrillo. Software sustainability: Research and practice from a software architecture viewpoint. The Journal of Systems and Software, 138(??):174–188, April 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217303072.

Vale:2016:TEY

[VCdA^{+}16] Tassio Vale, Ivica Crnkovic, Eduardo Santana de Almeida, Paulo Anselmo da Mota Silveira Neto, Yguaratã Cerqueira Cavalcanti, and Silvio Romero de Lemos Meira. Twenty-eight years of component-based software engineering. The Journal of Systems and Software, 111(??):128– 148, January 2016. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215002095.

Vilela:2017:IBR

[VCMG17] Jéssyka Vilela, Jaelson Castro, Luiz Eduardo G. Martins, and Tony Gorschek. Integration between requirements engineering and safety analysis: a systematic literature review. The Journal of Systems and Software, 125(??): 68–92, March 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216302333.

vandenBerg:2019:HEA

[vdBSvS^{+}19] Martin van den Berg, Raymond Slot, Marlies van Steenbergen, Peter Faasse, and Hans van Vliet. How enterprise architecture improves the quality of IT investment decisions. The Journal of Systems and Software, 152(??):134–150, June 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0164121219300433.

vanderRaadt:2010:RBE

[vdRBSvV10] Bas van der Raadt, Marc Bonnet, Sander Schouten, and Hans van Vliet. The relation between EA effectiveness and stakeholder satisfaction. The Journal of Systems and Software, 83 (10):1954–1969, October 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Vavliakis:2013:RPR

[VGM13] Konstantinos N. Vavliakis, Theofanis K. Grollios, and Pericles A. Mitkas. RDOTE — publishing relational databases into the Semantic Web. The Journal of Systems and Software, 86(1):89– 99, January 2013. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121212002117.

Valdivia-Garcia:2018:CPB

[VGSN18] Harold Valdivia-Garcia, Emad Shihab, and Meiyappan Nagappan. Characterizing and predicting blocking bugs in open source projects. The Journal of Systems and Software, 143 (??):44–58, September 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300530.

vanHeesch:2012:DFA

[vHAH12] U. van Heesch, P. Avgeriou, and R. Hilliard. A documentation framework for architecture decisions. The Journal of Systems and Software, 85(4):795–820, April 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002755.

vanHeesch:2013:DDD

[vHAT13] U. van Heesch, P. Avgeriou, and A. Tang. Does decision documentation help junior designers rationalize their decisions? A comparative multiplecase study. The Journal of Systems and Software, 86(6):1545–1565, June 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000228.

Vogel-Heuser:2017:MAP

[VHFF^{+}17] Birgit Vogel-Heuser, Juliane Fischer, Stefan Feldmann, Sebastian Ulewicz, and Susanne Rösch. Modularity and architecture of PLCbased software for automated production systems: an analysis in industrial companies. The Journal of Systems and Software, 131 (??):35–62, September 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217300985.

Vogel-Heuser:2015:ESA

[VHFST15] Birgit Vogel-Heuser, Alexander Fay, Ina Schaefer, and Matthias Tichy. Evolution of software in automated production systems: Challenges and research directions. The Journal of Systems and Software, 110(??):54–84, December 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001818.

vanHeesch:2017:PDS

[vHJPB^{+}17] U. van Heesch, A. Jansen, H. Pei-Breivold, P. Avgeriou, and C. Manteuffel. Platform design space exploration using architecture decision viewpoints — a longitudinal study. The Journal of Systems and Software, 124(??):56–81, February 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121630214X.

Vilbergsdottir:2014:ARV

[VHL14] Sigurbjorg Groa Vilbergsdottir, Ebba Thora Hvannberg, and Effie Lai-Chong Law. Assessing the reliability, validity and acceptance of a classification scheme of usability problems (CUP). The Journal of Systems and Software, 87(??):18–37, January 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002136.

Vasilecas:2016:RCB

[VKL16] Olegas Vasilecas, Diana Kalibatiene, and Dejan Lavbic. Ruleand context-based dynamic business process modelling and simulation. The Journal of Systems and Software, 122(??):1–15, December 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301509.

Vasconcellos:2017:ASA

[VLC^{+}17] Francisco J. S. Vasconcellos, Geraldo B. Landre, José Adson O. G. Cunha, Juliano L. Oliveira, Ronaldo A. Ferreira, and Auri M. R. Vincenzi. Approaches to strategic alignment of software process improvement: a systematic literature review. The Journal of Systems and Software, 123(??):45–63, January 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301893.

Vanhanen:2018:SEP

[VLL18] Jari Vanhanen, Timo O. A. Lehtinen, and Casper Lassenius. Software engineering problems and their relationship to perceived learning and customer satisfaction on a software capstone project. The Journal of Systems and Software, 137(??):50–66, March 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217302716.

Vara:2012:FMD

[VM12] Juan Manuel Vara and Esperanza Marcos. A framework for modeldriven development of information systems: Technical decisions and lessons learned. The Journal of Systems and Software, 85(10):2368–2384, October 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212001367.

Vidal:2013:TAR

[VM13] Santiago A. Vidal and Claudia A. Marcos. Toward automated refactoring of crosscutting concerns into aspects. The Journal of Systems and Software, 86(6): 1482–1497, June 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212003524.

Viana:2013:DSM

[VPdP13] Matheus C. Viana, Rosângela A. D. Penteado, and Antônio F. do Prado. Domain-Specific Modeling Languages to improve framework instantiation. The Journal of Systems and Software, 86 (12):3123–3139, December 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001775.

Vu:2010:ODH

[VPL^{+}10] Thi Hong Nhan Vu, Namkyu Park, Yang Koo Lee, Yongmi Lee, Jong Yun Lee, and Keun Ho Ryu. Online discovery of Heart Rate Variability patterns in mobile healthcare services. The Journal of Systems and Software, 83 (10):1930–1940, October 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Vazquez-Poletti:2013:SFC

[VPMVM^{+}13] J. L. Vazquez-Poletti, R. Moreno-Vozmediano, R. S. Montero, E. Huedo, and I. M. Llorente. Solidifying the foundations of the cloud for the next generation Software Engineering. The Journal of Systems and Software, 86 (9):2321–2326, September 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001350.

Vierhauser:2016:RFR

[VRG^{+}16] Michael Vierhauser, Rick Rabiser, Paul Grünbacher, Klaus Seyerlehner, Stefan Wallner, and Helmut Zeisel. ReMinds: a flexible runtime monitoring framework for systems of systems. The Journal of Systems and Software, 112(??):123–136, February 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001478.

Varga:2018:AMM

[VRPT18] Jovan Varga, Oscar Romero, Torben Bach Pedersen, and Christian Thomsen. Analytical metadata modeling for next generation BI systems. The Journal of Systems and Software, 144(??):240–254, October 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0164121218301274.

Verbelen:2012:AMI

[VSDD12] Tim Verbelen, Pieter Simoens, Filip De Turck, and Bart Dhoedt. AIO-LOS: Middleware for improving mobile application performance through cyber foraging. The Journal of Systems and Software, 85 (11):2629–2639, November 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001641.

Verbelen:2011:DDQ

[VSS^{+}11] Tim Verbelen, Tim Stevens, Pieter Simoens, Filip De Turck, and Bart Dhoedt. Dynamic deployment and quality adaptation for mobile augmented reality applications. The Journal of Systems and Software, 84(11):1871–1882, November 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121100166X.

Vardalis:2014:EPD

[VT14] Dimitris Vardalis and Vassilis Tsaoussidis. Exploiting the potential of DTN for energy-efficient internetworking. The Journal of Systems and Software, 90(??):91–103, April 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214000089.

Valsamis:2017:ETM

[VTZ^{+}17] Angelos Valsamis, Konstantinos Tserpes, Dimitrios Zissis, Dimosthenis Anagnostopoulos, and Theodora Varvarigou. Employing traditional machine learning algorithms for big data streams analysis: the case of object trajectory prediction. The Journal of Systems and Software, 127(??):249– 257, May 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121630084X.

vanVliet:2010:RSP

[vV10] Hans van Vliet. Reviewers are a sparse and precious resource. The Journal of Systems and Software, 83(3):351, March 2010. CODEN JSSODM.

ISSN 0164-1212 (print), 1873-1228 (electronic).

vanVliet:2013:STJ

[vV13] Hans van Vliet. Signs of a thriving journal. The Journal of Systems and Software, 86(1):1, January 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002944.

Veras:2015:BPA

[VVA^{+}15] Paulo C. Véras, Emilia Villani, Ana Maria Ambrosio, Marco Vieira, and Henrique Madeira. A benchmarking process to assess software requirements documentation for space applications. The Journal of Systems and Software, 100(??):103–116, February 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002404.

Vlietland:2016:ACS

[VvSvV16] Jan Vlietland, Rini van Solingen, and Hans van Vliet. Aligning codependent Scrum teams to enable fast business value delivery: a governance framework and set of intervention actions. The Journal of Systems and Software, 113(??):418–429, March 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002435.

vanVliet:2016:DMS

[vVT16] Hans van Vliet and Antony Tang. Decision making in software architecture. The Journal of Systems and Software, 117(??):638–644, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000157.

vonWangenheim:2013:SEG

[vWSB13] Christiane Gresse von Wangenheim, Rafael Savi, and Adriano Ferreti Borgatto. SCRUMIA — an educational game for teaching SCRUM in computing courses. The Journal of Systems and Software, 86 (10):2675–2687, October 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001295.

Varvarigou:2017:SIS

[VZT17] Theodora Varvarigou, Dimitrios Zissis, and Konstantinos Tserpes. Special issue on “Software architectures and systems for real time data stream analytics”. The Journal of Systems and Software, 127(??):177–178, May 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217300377.

Wang:2015:CET

[WAG15] Shuai Wang, Shaukat Ali, and Arnaud Gotlieb. Cost-effective test suite minimization in product lines using search techniques. The Journal of Systems and Software, 103(??):370–391, May 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001757.

Wautelet:2019:MDI

[Wau19]  Yves Wautelet.  A

model-driven IT governance process based on the strategic impact evaluation of services. The Journal of Systems and Software, 149(??):462–475, March 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302826.

Woungang:2012:CEB

[WAWO12] Isaac Woungang, Felix O. Akinladejo, David W. White, and Mohammad S. Obaidat. Codingerror based defects in enterprise resource planning software: Prevention, discovery, elimination and mitigation. The Journal of Systems and Software, 85(7):1682– 1698, July 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000544.

Wermelinger:2010:CSA

[WB10] Michel Wermelinger and Arosha Bandara. Commentary on ‘Software architectures and mobility: A Roadmap’. The Journal of Systems and Software, 83(6):899–901, June 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Weinreich:2012:TSS

[WB12] Rainer Weinreich and Georg Buchgeher. Towards supporting the software architecture life cycle. The Journal of Systems and Software, 85(3):546–561, March 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001361.

Woods:2015:MLS

[WB15] Eoin Woods and Rabih Bashroush. Modelling large-scale information systems using ADLs — an industrial experience report. The Journal of Systems and Software, 99(??):97–108, January 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002052.

Werner:2019:SII

[WB19] Claudia Werner and Goetz Botterweck. Special issue ICSR 2017. The Journal of Systems and Software, 152(??):

32, June 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121930041X.

Wolfenstetter:2018:ITT

[WBBK18] Thomas Wolfenstetter, Mohammad R. Basirati, Markus Böhm, and Helmut Krcmar. Introducing TRAILS: a tool supporting traceability, integration and visualisation of engineering knowledge for product service systems development. The Journal of Systems and Software, 144(??):342–355, October 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301365.

White:2010:ADF

[WBS^{+}10] J. White, D. Benavides, D. C. Schmidt, P. Trinidad, B. Dougherty, and A. Ruiz-Cortes. Automated diagnosis of feature model configurations. The Journal of Systems and Software, 83(7):1094–1107, July 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wu:2011:EEM

[WC11] Shih-Lin Wu and Shu-Chia Chen. An energyefficient MAC protocol with downlink traffic scheduling strategy in IEEE 802.11 infrastructure WLANs. The Journal of Systems and Software, 84(6):1022–1031, June 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wong:2016:ESI

[WC16] W. Eric Wong and W. K.

Chan. Editorial of the special issue to celebrate the 35th anniversary of JSS. The Journal of Systems and Software, 116(??):1, June 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000856.

Wu:2017:TCS

[WCB^{+}17] Guanlin Wu, Junjie Chen, Weidong Bao, Xiaomin Zhu, Wenhua Xiao, and Ji Wang. Towards collaborative storage scheduling using alternating direction method of multipliers for mobile edge cloud. The Journal of Systems and Software, 134(??):29–43, December 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301826.

Wang:2012:LSD

[WCC12] Xiaofeng Wang, Kieran Conboy, and Oisin Cawley. “leagile” software development: an experience report analysis of the application of lean approaches in agile software development. The Journal of Systems and Software, 85(6):1287–1299, June 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000404.

Wu:2013:CRL

[WCC13] Shih-Lin Wu, Jen-Jee Chen, and Wen-Chiang Chou. Cell-related location area planning for 4G PCS networks with variable-order Markov model. The Journal of Systems and Software, 86(10):2688–2699, October 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0164121213001301.

Wang:2014:HCD

[WCC^{+}14] Chung-Chuan Wang, Ya-Fen Chang, Chin-Chen Chang, Jinn-Ke Jan, and Chia-Chen Lin. A high capacity data hiding scheme for binary images based on block patterns. The Journal of Systems and Software, 93(??):152–162, July 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214000545.

Wang:2010:EMB

[WCCL10] Zhi-Hui Wang, Chin-Chen Chang, Kuo-Nan Chen, and Ming-Chu Li. An encoding method for both image compression and data lossless information hiding. The Journal of Systems and Software, 83(11):2073–2082, November 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wong:2012:SID

[WCTK12] W. Eric Wong, W. K.

Chan, T. H. Tse, and Fei-Ching Kuo. Special issue on dynamic analysis and testing of embedded software. The Journal of Systems and Software, 85(1):1–2, January 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002548.

Wang:2015:DCS

[WCX15] Xiaogang Wang, Jian Cao, and Yang Xiang. Dynamic cloud service selection using an adaptive learning mechanism in multi-cloud computing. The Journal of Systems and Software, 100(??):195–210, February 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002398.

Wong:2010:FCC

[WDC10] W. Eric Wong, Vidroha Debroy, and Byoungju Choi. A family of code coverage-based heuristics for effective fault localization. The Journal of Systems and Software, 83(2):188–208, February 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wang:2012:AMF

[WDC12] Xiaoying Wang, Zhihui Du, and Yinong Chen. An adaptive model-free resource and power management approach for multi-tier cloud environments. The Journal of Systems and Software, 85(5):1135–1146, May 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211003244.

Wen:2016:EAD

[Wen16] Yean-Fu Wen. Energyaware dynamical hosts and tasks assignment for cloud computing. The Journal of Systems and Software, 115(??):144– 156, May 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000236.

Walter:2018:CST

[WFF18] Bartosz Walter, Francesca Arcelli Fontana, and Vincenzo Ferme. Code smells and their collocations: a large-scale experiment on open-source systems. The Journal of Systems and Software, 144(??):1–21, October 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301109.

Wang:2019:MBM

[WFY^{+}19] Hongbing Wang, Huanhuan Fei, Qi Yu, Wei Zhao, Jia Yan, and Tianjing Hong. A motifsbased Maximum Entropy Markov Model for realtime reliability prediction in System of Systems. The Journal of Systems and Software, 151(??):180–193, May 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300330.

Wang:2014:WWA

[WGC^{+}14] Yabin Wang, Ruizhi Gao, Zhenyu Chen, W. Eric Wong, and Bin Luo. WAS: a weighted attribute-based strategy for cluster test selection. The Journal of Systems and Software, 98(??):44–58, December 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0164121214001848.

Wang:2019:CCS

[WGKW19] Yang Wang, Daniel Graziotin, Stefan Kriso, and Stefan Wagner. Communication channels in safety analysis: an industrial exploratory case study. The Journal of Systems and Software, 153(??):135–151, July 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300755.

White:2014:EFM

[WGS^{+}14] Jules White, José A.

Galindo, Tripti Saxena, Brian Dougherty, David Benavides, and Douglas C. Schmidt. Evolving feature model configurations in software product lines. The Journal of Systems and Software, 87(??):119–136, January 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002434.

Wei:2012:NCI

[WGZ^{+}12] Xiaopeng Wei, Ling Guo, Qiang Zhang, Jianxin Zhang, and Shiguo Lian. A novel color image encryption algorithm based on DNA sequence operation and hyper-chaotic system. The Journal of Systems and Software, 85(2):290–299, February 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002147.

Wen:2015:EEH

[WH15] Yean-Fu Wen and Ko-Yu Hung. Energy efficiency heterogeneous wireless access selection for multiple types of applications. The Journal of Systems and Software, 101(??):97–109, March 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002763.

Wang:2012:FOP

[WHY^{+}12] Xu An Wang, Xinyi Huang, Xiaoyuan Yang, Longfei Liu, and Xuguang Wu. Further observation on proxy re-encryption with keyword search. The Journal of Systems and Software, 85(3):643–654, March 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002433.

Wieringa:2014:ERM

[Wie14]  Roel Wieringa.  Em-

pirical research methods for technology validation: Scaling up to practice. The Journal of Systems and Software, 95(??):19–31, September 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002793.

Wale-Kolade:2015:IUW

[WK15] Adeola Yetunde Wale-Kolade. Integrating usability work into a large inter-organisational agile development project: Tactics developed by usability designers. The Journal of Systems and Software, 100(??):54–66, February 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002337.

Wang:2017:IOC

[WKbOS17] Jijie Wang, Mark Keil, Lih bin Oh, and Yide Shen. Impacts of organizational commitment, interpersonal closeness, and Confucian ethics on willingness to report bad news in software projects. The Journal of Systems and Software, 125(??):220– 233, March 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

//www.sciencedirect.
com/science/article/
pii/S0164121216302515.

Wei:2019:MBS

[WKD^{+}19] Ran Wei, Tim P. Kelly, Xiaotian Dai, Shuai Zhao, and Richard Hawkins. Model based system assurance using the structured assurance case metamodel. The Journal of Systems and Software, 154(??):211–233, August 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301062.

Wu:2011:HQI

[WKH11] Chia-Chun Wu, Shang-Juh Kao, and Min-Shiang Hwang. A high quality image sharing with steganography and adaptive authentication scheme. The Journal of Systems and Software, 84 (12):2196–2207, December 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001488.

Weldemariam:2011:FAE

[WKV11] Komminist Weldemariam, Richard A. Kemmerer, and Adolfo Villafiorita. Formal analysis of an electronic voting system: an experience report. The Journal of Systems and Software, 84(10):1618–1637, October 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211000756.

Wang:2010:MCW

[WKZL10] Hongbing Wang, Zuling Kang, Ning Zhou, and Li Li. A model checker for WS-CDL. The Journal of Systems and Software, 83(10):1651–1661, October 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wijayasiriwardhane:2010:CPS

[WL10] Thareendhra Wijayasiriwardhane and Richard Lai. Component Point: a system-level size measure for Component-Based Software Systems. The Journal of Systems and Software, 83(12):2456– 2470, December 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wallshein:2015:SCE

[WL15a] Corinne C. Wallshein and Andrew G. Loerch. Software cost estimating for CMMI Level 5 developers. The Journal of Systems and Software, 105(??):72–78, July 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000710.

Wang:2015:AFL

[WL15b] Xiaoyan Wang and Yongmei Liu. Automated fault localization via hierarchical multiple predicate switching. The Journal of Systems and Software, 104(??):69–81, June 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000448.

Wang:2016:FLU

[WL16] Xiaoyan Wang and Yongmei Liu. Fault localization using disparities of dynamic invariants. The Journal of Systems and Software, 122(??):144–154, December 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301820.

Wang:2017:MCD

[WL17] You-Chiun Wang and Shu-Ju Liu. Minimumcost deployment of adjustable readers to provide complete coverage of tags in RFID systems. The Journal of Systems and Software, 134(??):228–241, December 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302054.

Wang:2013:CIS

[WLC13a] Chun-Sheng Wang, Ying-Ho Liu, and Kuo-Chung Chu. Closed intersequence pattern mining. The Journal of Systems and Software, 86(6):1603–1612, June 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000290.

Wang:2013:HSI

[WLC13b] Zhi-Hui Wang, Chin-Feng Lee, and Ching-Yun Chang. Histogramshifting-imitated reversible data hiding. The Journal of Systems and Software, 86(2):315–323, February 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002415.

Wurfel:2016:GRE

[WLD16] David Würfel, Rainer Lutz, and Stephan Diehl. Grounded requirements engineering: an approach to use case driven requirements engineering. The Journal of Systems and Software, 117(??):645–657, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002277.

Wang:2013:HCL

[WLH13] Kan Wang, Zhe-Ming Lu, and Yong-Jian Hu. A high capacity lossless data hiding scheme for JPEG images. The Journal of Systems and Software, 86(7):1965–1975, July 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000812.

Wu:2013:SPV

[WLL^{+}13] Kuang-Shyr Wu, Jen-Chun Lee, Tsung-Ming Lo, Ko-Chin Chang, and Chien-Ping Chang. A secure palm vein recognition system. The Journal of Systems and Software, 86(11):2870–2876, November 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001660.

Wang:2015:MVW

[WLL15] Xinzhi Wang, Xiangfeng Luo, and Huiming Liu. Measuring the veracity of web event via uncertainty. The Journal of Systems and Software, 102(??):226–236, April 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121400154X.

Wong:2017:MFO

[WLL17] W. Eric Wong, Xuelin Li, and Philip A. Laplante. Be more familiar with our enemies and pave the way forward: a review of the roles bugs played in software failures. The Journal of Systems and Software, 133(??):68–94, November 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301334.

Wang:2019:ETR

[WLL19a] Lulu Wang, Jingyue Li, and Bixin Li. Erratum to “Tracking runtime concurrent dependences in Java threads using thread control profiling” [The Journal of Systems and Software 148 (2019) 116–131]. The Journal of Systems and Software, 153(??): 44, July 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/

pii/S0164121219300615.

See [WLL19b].

Wang:2019:TRC

[WLL19b] Lulu Wang, Jingyue Li, and Bixin Li. Tracking runtime concurrent dependences in Java threads using thread control profiling. The Journal of Systems and Software, 148(??):116–131, February 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302395.

See erratum [WLL19a].

Wang:2017:HSP

[WLZ^{+}17a] Chao Wang, Xi Li, Huizhen Zhang, Aili Wang, and Xuehai Zhou. Hot spots profiling and dataflow analysis in custom dataflow computing SoftProcessors. The Journal of Systems and Software, 125(??):427– 438, March 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

//www.sciencedirect.
com/science/article/
pii/S0164121216301182.

Wang:2017:DIM

[WLZ^{+}17b] Run Wang, Pei Liu, Lei Zhao, Yueqiang Cheng, and Lina Wang. deExploit: Identifying misuses of input data to diagnose memory-corruption exploits at the binary level. The Journal of Systems and Software, 124(??):153–168, February 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121630228X.

Weyns:2012:ISI

[WMAS12] Danny Weyns, Sam Malek, Jesper Andersson, and Bradley Schmerl. Introduction to the special issue on state of the art in engineering self-adaptive systems. The Journal of Systems and Software, 85 (12):2675–2677, December 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002208.

Weyns:2017:ISI

[WMC17] Danny Weyns, Raffaela Mirandola, and Ivica Crnkovic. Introduction to the special issue on “New frontiers in software architecture”. The Journal of Systems and Software, 130(??):57–58, August 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217300341.

Woungang:2010:SAM

[WMD^{+}10] Isaac Woungang, Guangyan Ma, Mieso K. Denko, Sudip Misra, Han-Chieh Chao, and Mohammad S. Obaidat. Survivable ATM mesh networks: Techniques and performance evaluation. The Journal of Systems and Software, 83(3):457–466, March 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wu:2011:MAP

[WMOKY11] Yulei Wu, Geyong Min, Mohamed Ould-Khaoua, and Hao Yin. Modelling and analysis of pipelined circuit switching in interconnection networks with bursty traffic and hot-spot destinations. The Journal of Systems and Software, 84(12):2097–2106, December 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001397.

Weidlich:2012:PCB

[WMW12] Matthias Weidlich, Jan Mendling, and Mathias Weske. Propagating changes between aligned process models. The Journal of Systems and Software, 85 (8):1885–1898, August 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212000672.

Wang:2019:LAA

[WMW^{+}19] Weichao Wang, Zhaopeng Meng, Zan Wang, Shuang Liu, and Jianye Hao. LoopFix: an approach to automatic repair of buggy loops. The Journal of Systems and Software, 156(??):100–112, October 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301384.

Wei:2012:QSF

[WMWZ12] Tongquan Wei, Piyush Mishra, Kaijie Wu, and Junlong Zhou. Quasistatic fault-tolerant scheduling schemes for energyefficient hard real-time systems. The Journal of Systems and Software, 85(6):1386–1399, June 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000064.

White:2017:QSA

[WNC17] Gary White, Vivek Nallur, and Siobhán Clarke. Quality of service approaches in IoT: a systematic mapping. The Journal of Systems and Software, 132(??):186–203, October 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121730105X.

Wu:2015:QSL

[WOC15] Tin-Yu Wu, Mohammad S. Obaidat, and Hung-Lin Chan. QualityScan scheme for load balancing efficiency in vehicular ad hoc networks (VANETs). The Journal of Systems and Software, 104(??):60–68, June 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121215000321.

Wohlin:2016:EGT

[Woh16] Claes Wohlin. Erratum to “A General Theory of Software Engineering: Balancing Human, Social and Organizational Capitals” [The Journal of Systems & Software 109c (2015) 229–242]. The Journal of Systems and Software, 117(??): 129, July 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216000753.

See [WSM15].

Wu:2012:UFS

[WOLS12] Xiaotian Wu, Duanhao Ou, Qiming Liang, and Wei Sun. A userfriendly secret image sharing scheme with reversible steganography based on cellular automata. The Journal of Systems and Software, 85(8):1852–1863, August 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212000696.

Wong:2010:EJT

[Won10] W. Eric Wong. Editorial for the JSS Top Scholar Special Issue. The Journal of Systems and Software, 83(1):1, January 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Woods:2012:IAA

[Woo12]  Eoin Woods.  Indus-

trial architectural assessment using TARA. The Journal of Systems and Software, 85 (9):2034–2047, September 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001264.

Wang:2018:RTT

[WPL^{+}18] Bangchao Wang, Rong Peng, Yuanbang Li, Han Lai, and Zhuo Wang. Requirements traceability technologies and technology transfer decision support: a systematic review. The Journal of Systems and Software, 146(??):59–79, December 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301754.

Wang:2010:HFT

[WQJZ10] Tao Wang, Zhenxing Qin, Zhi Jin, and Shichao Zhang. Handling overfitting in test costsensitive decision tree learning by feature selection, smoothing and pruning. The Journal of Systems and Software, 83(7):1137–1147, July 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Winter:2010:SSF

[WR10] Jeff Winter and Kari Rönkkö. SPI success factors within product usability evaluation. The Journal of Systems and Software, 83(11):2059– 2072, November 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wohlin:2013:RMS

[WRdMSN^{+}13] Claes Wohlin, Per Runeson, Paulo Anselmo da Mota Silveira Neto, Emelie Engström, Ivan do Carmo Machado, and Eduardo Santana de Almeida. On the reliability of mapping studies in software engineering. The Journal of Systems and Software, 86(10):2594–2610, October 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001234.

Winter:2014:IOB

[WRR14] Jeff Winter, Kari Rönkkö, and Mikko Rissanen. Identifying organizational barriers — a case study of usability work when developing software in the automation industry. The Journal of Systems and Software, 88(??):54–73, February 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002343.

Wiese:2017:UCI

[WRS^{+}17] Igor Scaliante Wiese, Reginaldo Ré, Igor Steinmacher, Rodrigo Takashi Kuroda, Gustavo Ansaldi Oliva, Christoph Treude, and Marco Aurélio Gerosa. Using contextual information to predict cochanges. The Journal of Systems and Software, 128(??):220–235, June 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121216301194.

Walker:2013:AOS

[WRTP^{+}13] Martin Walker, Mark-Oliver Reiser, Sara Tucci-Piergiovanni, Yiannis Papadopoulos, Henrik Lönn, Chokri Mraidha, David Parker, DeJiu Chen, and David Servat. Automatic optimisation of system architectures using EAST-ADL. The Journal of Systems and Software, 86(10):2467–2487, October 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213000885.

Wu:2012:RGB

[WS12] Xiaotian Wu and Wei Sun. Random gridbased visual secret sharing for general access structures with cheatpreventing ability. The Journal of Systems and Software, 85(5):1119– 1134, May 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211003220.

Wu:2013:SIS

[WS13] Xiaotian Wu and Wei Sun. Secret image sharing scheme with authentication and remedy abilities based on cellular automata and discrete wavelet transform. The Journal of Systems and Software, 86(4):1068–1088, April 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212003147.

Wuyts:2014:EEP

[WSJ14] Kim Wuyts, Riccardo Scandariato, and Wouter Joosen. Empirical evaluation of a privacy-focused threat modeling methodology. The Journal of Systems and Software, 96(??):122–138, October 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121400137X.

Wohlin:2015:GTS

[WSM15] Claes Wohlin, Darja Smite, and Nils Brede Moe. A general theory of software engineering: balancing human, social and organizational capitals. The Journal of Systems and Software, 109(??):229–242, November 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215001740.

See erratum [Woh16].

Wong:2011:ASS

[WTG^{+}11] W. Eric Wong, T. H. Tse, Robert L. Glass, Victor R. Basili, and T. Y. Chen. An assessment of systems and software engineering scholars and institutions (2003–2007 and 2004–2008). The Journal of Systems and Software, 84(1):162–168, January 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wang:2015:SDA

[WTG^{+}15] Wendong Wang, Ye Tian, Xiangyang Gong, Qinglei Qi, and Yannan Hu. Software defined autonomic QoS model for future Internet. The Journal of Systems and Software, 110(??):122–135, December 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0164121215001776.

Wu:2011:MSF

[Wu11] Wei-Wen Wu. Mining significant factors affecting the adoption of SaaS using the rough set approach. The Journal of Systems and Software, 84 (3):435–441, March 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Weldemariam:2011:PSA

[WV11] Komminist Weldemariam and Adolfo Villafiorita. Procedural security analysis: a methodological approach. The Journal of Systems and Software, 84(7):1114–1129, July 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Walraven:2014:ECM

[WVT^{+}14] Stefan Walraven, Dimitri Van Landuyt, Eddy Truyen, Koen Handekyn, and Wouter Joosen. Efficient customization of multi-tenant Softwareas-a-Service applications with service lines. The Journal of Systems and Software, 91(??):48–62, May 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214000326.

Wang:2010:CAL

[WWL^{+}10] Meng Wang, Yi Wang, Duo Liu, Zhiwei Qin, and Zili Shao. Compilerassisted leakage-aware loop scheduling for embedded VLIW DSP processors. The Journal of Systems and Software, 83 (5):772–785, May 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wang:2013:GBR

[WWLG13] Fangda Wang, Hongzhi Wang, Jianzhong Li, and Hong Gao. Graph-based reference table construction to facilitate entity matching. The Journal of Systems and Software, 86(6):1679–1688, June 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213000484.

WiedermannAgner:2013:BSU

[WWSS13] Luciane Telinski Wiedermann Agner, Inali Wisniewski Soares, Paulo Cézar Stadzisz, and Jean Marcelo Simão. A Brazilian survey on UML and model-driven practices for embedded software development. The Journal of Systems and Software, 86(4):997–1005, April 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212003160.

Wang:2015:ISD

[WWSZ15] Jinyong Wang, Zhibo Wu, Yanjun Shu, and Zhan Zhang. An imperfect software debugging model considering log-logistic distribution fault content function. The Journal of Systems and Software, 100(??):167–181, February 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002374.

Wu:2012:SMD

[WWY^{+}12] Guoquan Wu, Jun Wei, Chunyang Ye, Hua Zhong, Tao Huang, and Hong He. Specification and monitoring of datacentric temporal properties for service-based systems. The Journal of Systems and Software, 85 (12):2738–2754, December 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212001604.

Wang:2011:CHI

[WWYZ11] Xu An Wang, Jian Weng, Xiaoyuan Yang, and Minqing Zhang. Cryptanalysis of an (hierarchical) identity based parallel key-insulated encryption scheme. The Journal of Systems and Software, 84(2):219–225, February 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wang:2014:WAA

[WWZ^{+}14] Tao Wang, Jun Wei, Wenbo Zhang, Hua Zhong, and Tao Huang. Workload-aware anomaly detection for Web applications. The Journal of Systems and Software, 89(??):19–32, March 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213000721.

Wu:2010:TAT

[WX10] Guowei Wu and Zichuan Xu. Temperature-aware task scheduling algorithm for soft real-time multi-core systems. The Journal of Systems and Software, 83(12):2579– 2590, December 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wang:2017:OCB

[WXY^{+}17] Hongda Wang, Jianchun Xing, Qiliang Yang, Ping Wang, Xuewei Zhang, and Deshuai Han. Optimal control based regression test selection for service-oriented workflow applications. The Journal of Systems and Software, 124(??):274–288, February 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300942.

Wang:2017:RES

[WXZ^{+}17] Tao Wang, Jiwei Xu, Wenbo Zhang, Jianhua Zhang, Jun Wei, and Hua Zhong. ReSeer: Efficient search-based replay for multiprocessor virtual machines. The Journal of Systems and Software, 126(??):101– 112, April 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121216301248.

Wang:2013:HPR

[WYCC13] Zhi-Hui Wang, Hai-Rui Yang, Ting-Fang Cheng, and Chin-Chen Chang. A high-performance reversible data-hiding scheme for LZW codes. The Journal of Systems and Software, 86(11):2771– 2778, November 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121300143X.

Wen:2011:DSH

[WZ11] Yamin Wen and Fangguo Zhang. Delegatable secret handshake scheme. The Journal of Systems and Software, 84 (12):2284–2292, December 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211001580.

Zhang:2013:PAS

[wZfG13] Yi wen Zhang and Rui feng Guo. Power-aware scheduling algorithms for sporadic tasks in realtime systems. The Journal of Systems and Software, 86(10):2611–2619, October 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001222.

See corrigendum [wZfG14a].

Zhang:2014:CPA

[wZfG14a] Yi wen Zhang and Rui feng Guo. Corrigendum to “Power-aware scheduling algorithms for sporadic tasks in real-time systems” [J. Syst. Softw. 86 (2013) 2611–2619]. The Journal of Systems and Software, 94(??):202, August 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001046.

See [wZfG13].

Zhang:2014:PAF

[wZfG14b] Yi wen Zhang and Rui feng Guo. Power-aware fixed priority scheduling for sporadic tasks in hard real-time systems. The Journal of Systems and Software, 90(??):128– 137, April 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214000053.

Wu:2014:BBS

[WZJI14] Quanwang Wu, Qingsheng Zhu, Xing Jian, and Fuyuki Ishikawa. Broker-based SLA-aware composite service provisioning. The Journal of Systems and Software, 96(??):194–201, October 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214001459.

Wei:2012:CSO

[WZM12a] Fushan Wei, Zhenfeng Zhang, and Chuangui Ma. Corrigendum to “Gateway-oriented password-authenticated key exchange protocol in the standard model’ [J. Syst. Softw. 85 (March (3)) (2012) 760– 768]. The Journal of Systems and Software, 85(9):2192, September 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212001550.

See [WZM12b].

Wei:2012:GOP

[WZM12b] Fushan Wei, Zhenfeng Zhang, and Chuangui Ma. Gateway-oriented password-authenticated key exchange protocol in the standard model. The Journal of Systems and Software, 85(3):760–768, March 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002597.

See corrigendum [WZM12a].

Wang:2018:URR

[WZY^{+}18] Ying Wang, Zhiliang Zhu, Bo Yang, Fangda Guo, and Hai Yu. Using reliability risk analysis to prioritize test cases. The Journal of Systems and Software, 139(??):14–31, May 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300128.

Xu:2019:SPM

[XB19a] H. Xu and A. Burns. A semi-partitioned model for mixed criticality systems. The Journal of Systems and Software, 150(??):51–63, April 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121219300020.

Xu:2019:BSS

[XB19b] Minxian Xu and Rajkumar Buyya. Brownout-Con: a software system based on brownout and containers for energyefficient cloud computing. The Journal of Systems and Software, 155(??):91–103, September 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301189.

Xu:2012:AID

[XCM^{+}12] Chang Xu, S. C. Cheung, Xiaoxing Ma, Chun Cao, and Jian Lu. Adam: Identifying defects in context-aware adaptation. The Journal of Systems and Software, 85 (12):2812–2828, December 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212001318.

Xie:2011:TVM

[XHM^{+}11] Xiaoyuan Xie, Joshua W. K. Ho, Christian Murphy, Gail Kaiser, Baowen Xu, and Tsong Yueh Chen. Testing and validating machine learning classifiers by metamorphic testing. The Journal of Systems and Software, 84(4):544–558, April 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Xiao:2013:ESF

[Xia13] Lu Xiao. The effects of a shared free form rationale space in collaborative learning activities. The Journal of Systems and Software, 86(7):1727–1737, July 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002178.

Xiao:2015:SDV

[XJZ^{+}15] Zhijiao Xiao, Jianmin Jiang, Yingying Zhu, Zhong Ming, Shenghua Zhong, and Shubin Cai. A solution of dynamic VMs placement problem for energy consumption optimization based on evolutionary game theory. The Journal of Systems and Software, 101(??):260–272, March 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121400288X.

Xu:2019:TTS

[XLL^{+}19] Zhou Xu, Shuai Li, Xiapu Luo, Jin Liu, Tao Zhang, Yutian Tang, Jun Xu, Peipei Yuan, and Jacky Keung. TSTSS: a two-stage training subset selection framework for cross version defect prediction. The Journal of Systems and Software, 154(??):59–78, August 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219300627.

Xu:2015:SBR

[XLM^{+}15] Zheng Xu, Yunhuai Liu, Lin Mei, Chuanping Hu, and Lan Chen. Semantic based representing and organizing surveillance big data using video structural description technology. The Journal of Systems and Software, 102(??):217– 225, April 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214001551.

Xiao:2018:FBI

[XLW18] Peng Xiao, Bin Liu, and Shihai Wang. Feedbackbased integrated prediction: Defect prediction based on feedback from software testing process. The Journal of Systems and Software, 143 (??):159–171, September 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218301067.

Xu:2019:LLD

[XLX^{+}19] Zhou Xu, Shuai Li, Jun Xu, Jin Liu, Xiapu Luo, Yifeng Zhang, Tao Zhang, Jacky Keung, and Yutian Tang. LDFR: Learning deep feature representation for software defect prediction. The Journal of Systems and Software, 158(??):??, December 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219301761.

Xhafa:2011:UGS

[XPBC11] Fatos Xhafa, Claudi Paniagua, Leonard Barolli, and Santi Caballé. Using Grid services to parallelize IBM’s Generic Log Adapter. The Journal of Systems and Software, 84 (1):55–62, January 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Xu:2018:MIA

[XSL^{+}18] Congying Xu, Xiaobing Sun, Bin Li, Xintong Lu, and Hongjing Guo. MULAPI: Improving API method recommendation with API usage location. The Journal of Systems and Software, 142(??):195–205, August 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218300840.

Xie:2018:ISI

[XST18] Xiaoyuan Xie, Markus Stumptner, and T. H. Tse. Introduction to the special issue on program debugging. The Journal of Systems and Software, 140(??):109– 110, June 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218300463.

Xiao:2012:VLM

[XTZX12] Xi Xiao, Xinguang Tian, Qibin Zhai, and Shutao Xia. A variable-length model for masquerade detection. The Journal of Systems and Software, 85(11):2470–2478, November 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001471.

See corrigendum [XTZX13].

Xiao:2013:CSV

[XTZX13] Xi Xiao, Xinguang Tian, Qibin Zhai, and Shutao Xia. Corrigendum to “A variable-length model for masquerade detection” [J. Syst. Softw. 85 (2012) 2470–2478]. The Journal of Systems and Software, 86(3):868, March 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212003093.

See [XTZX12].

Xu:2014:DCF

[XWZC14] Ruzhi Xu, Shuaiqiang Wang, Xuwei Zheng, and Yinong Chen. Distributed collaborative filtering with singular ratings for large scale recommendation. The Journal of Systems and Software, 95(??):231–241, September 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214001150.

Xie:2017:SAA

[XYCL17] Haihua Xie, Jingwei Yang, Carl K. Chang, and Lin Liu. A statistical analysis approach to predict user’s changing requirements for software service evolution. The Journal of Systems and Software, 132 (??):147–164, October 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217301358.

Xu:2019:RSB

[XYZ^{+}19] Yanhong Xu, Beibei Yin, Zheng Zheng, Xiaoyi Zhang, Chenglong Li, and Shunkun Yang. Robustness of spectrumbased fault localisation in environments with labelling perturbations. The Journal of Systems and Software, 147 (??):172–214, January 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302164.

Xu:2010:UTP

[XZP^{+}10] Wenjun Xu, Zude Zhou, D. T. Pham, C. Ji, M. Yang, and Quan Liu. Unreliable transport protocol using congestion control for high-speed networks. The Journal of Systems and Software, 83 (12):2642–2652, December 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Xu:2016:CBA

[XZZ^{+}16] Jiwei Xu, Wenbo Zhang, Zhenyu Zhang, Tao Wang, and Tao Huang. Clustering-based acceleration for virtual machine image deduplication in the cloud environment. The Journal of Systems and Software, 121(??):144–156, November 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000534.

Yazdi:2016:FCS

[YAKK16] Hamed Shariat Yazdi, Lefteris Angelis, Timo Kehrer, and Udo Kelter. A framework for capturing, statistically modeling and analyzing the evolution of software models. The Journal of Systems and Software, 118(??):176–207, August 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300437.

Younas:2011:SII

[YAT11] Muhammad Younas, Irfan Awan, and Makoto Takizawa. Special issue on the information networking and services. The Journal of Systems and Software, 84(1):1, January 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yan:2013:MEA

[YAY13] Dong Yan, Syed Zubair Ahmad, and Dong Yang. Matthew effect, ABC analysis and project management of scale-free information systems. The Journal of Systems and Software, 86(2):247–254, February 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0164121212002294.

Yoo:2017:OSB

[YBE17] Shin Yoo, David Binkley, and Roger Eastman. Observational slicing based on visual semantics. The Journal of Systems and Software, 129(??):60–78, July 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216300255.

Yang:2011:GSS

[YC11] Ching-Nung Yang and Yu-Ying Chu. A general (k, n) scalable secret image sharing scheme with the smooth scalability. The Journal of Systems and Software, 84(10):1726–1733, October 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001105.

Yang:2012:PST

[YC12] Jun-Han Yang and Tian-Jie Cao. Provably secure three-party password authenticated key exchange protocol in the standard model. The Journal of Systems and Software, 85(2):340–350, February 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002172.

Yamashita:2013:CSS

[YC13] Aiko Yamashita and Steve Counsell. Code smells as system-level indicators of maintainability: an empirical study. The Journal of Systems and Software, 86(10):2639–2653, October 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213001258.

Yang:2017:MSC

[YCA17] Hongji Yang, Feng Chen, and Suleiman Aliyu. Modern software cybernetics: new trends. The Journal of Systems and Software, 124 (??):169–186, February 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301595.

Yang:2016:EBB

[YCC16] Ching-Nung Yang, Cheng-Hua Chen, and Song-Ruei Cai. Enhanced Boolean-based multi secret image sharing scheme. The Journal of Systems and Software, 116(??): 22–34, June 2016. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121215000205.

Yang:2013:PFM

[YCF^{+}13] Dingyu Yang, Jian Cao, Jiwen Fu, Jie Wang, and Jianmei Guo. A pattern fusion model for multistep-ahead CPU load prediction. The Journal of Systems and Software, 86(5):1257–1266, May 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212003354.

Yang:2014:GTA

[YCG^{+}14] Weiran Yang, Zhenyu Chen, Zebao Gao, Yunxiao Zou, and Xiaoran Xu. GUI testing assisted by human knowledge: Random vs. functional. The Journal of Systems and Software, 89(??):76–86, March 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002392.

Yang:2017:MDF

[YCLC17] Shengtao Yang, Henry C. B. Chan, Patrick P. Lam, and Peter H. J. Chong. MeshFS: a distributed file system for cloud-based wireless mesh network. The Journal of Systems and Software, 131(??):201–217, September 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S016412121730122X.

Yang:2013:ERD

[YCLY13] Wei-Jen Yang, Kuo-Liang Chung, Hong-Yuan Mark Liao, and Wen-Kuang Yu. Efficient reversible data hiding algorithm based on gradient-based edge direction prediction. The Journal of Systems and Software, 86(2):567–580, February 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002786.

Yang:2015:POA

[YCWW15] Dingyu Yang, Jian Cao, Sai Wu, and Jie Wang. Progressive online aggregation in a distributed stream system. The Journal of Systems and Software, 102(??):146– 157, April 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002556.

Yang:2012:GAQ

[YDGB^{+}12] Yong Yang, Marlon Dumas, Luciano García-Bañuelos, Artem Polyvyanyy, and Liang Zhang. Generalized aggregate Quality of Service computation for composite services. The Journal of Systems and Software, 85(8):1818–1830, August 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000726.

Yang:2015:CCD

[YF15] Hang Yang and Simon Fong. Countering the concept-drift problems in big data by an incrementally optimized stream mining model. The Journal of Systems and Software, 102(??):158–166, April 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001526.

Yu:2015:AAS

[YFT^{+}15] Yijun Yu, Virginia N. L. Franqueira, Thein Than Tun, Roel J. Wieringa, and Bashar Nuseibeh. Automated analysis of security requirements through risk-based argumentation. The Journal of Systems and Software, 106(??):102–116, August 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000850.

Yan:2016:ACS

[YFZ^{+}16] Meng Yan, Ying Fu, Xiaohong Zhang, Dan Yang, Ling Xu, and Jeffrey D. Kymer. Automatically classifying software changes via discriminative topic model: Supporting multi-category and cross-project. The Journal of Systems and Software, 113(??):296– 308, March 2016. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S016412121500285X.

Yousafzai:2016:COM

[YGN^{+}16] Abdullah Yousafzai, Abdullah Gani, Rafidah Md Noor, Anjum Naveed, Raja Wasim Ahmad, and Victor Chang. Computational offloading mechanism for native and Android runtime based mobile applications. The Journal of Systems and Software, 121(??):28–39, November 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301364.

Yoo:2010:UHA

[YH10] Shin Yoo and Mark Harman. Using hybrid algorithm for Pareto efficient multi-objective test suite minimisation. The Journal of Systems and Software, 83(4):689–701, April 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yang:2013:IRS

[YH13] Wan-Shiou Yang and San-Yih Hwang. iTravel: a recommender system in mobile peer-to-peer environment. The Journal of Systems and Software, 86(1):12–20, January 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001768.

Yarinezhad:2019:RAW

[YH19] Ramin Yarinezhad and Seyed Naser Hashemi. A routing algorithm for wireless sensor networks based on clustering and an FPT-approximation algorithm. The Journal of Systems and Software, 155(??):145–161, September 2019. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0164121219301190.

Yang:2014:ATA

[YHM^{+}14] Zhibin Yang, Kai Hu, Dianfu Ma, Jean-Paul Bodeveix, Lei Pi, and Jean-Pierre Talpin. From AADL to Timed Abstract State Machines: a verified model transformation. The Journal of Systems and Software, 93(??):42–68, July 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214000727.

Yli-Huumo:2016:HDS

[YHMS16] Jesse Yli-Huumo, Andrey Maglyas, and Kari Smolander. How do software development teams manage technical debt? — An empirical study. The Journal of Systems and Software, 120(??):195–218, October 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121630053X.

Yu:2017:FMT

[YJZ17] Qiao Yu, Shujuan Jiang, and Yanmei Zhang. A feature matching and transfer approach for cross-company defect prediction. The Journal of Systems and Software, 132(??):366–378, October 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301346.

Yu:2012:IRI

[YKC^{+}12] Jia Yu, Fanyu Kong, Xiangguo Cheng, Rong Hao, and Jianxi Fan. Intrusion-resilient identitybased signature: Security definition and construction. The Journal of Systems and Software, 85(2):382–391, February 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002342.

Yang:2016:MPM

[YL16] Jen-Ho Yang and Pei-Yu Lin. A mobile payment mechanism with anonymity for cloud computing. The Journal of Systems and Software, 116(??):69–74, June 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121500151X.

Yang:2016:SSA

[YLA16a] Chen Yang, Peng Liang, and Paris Avgeriou. A survey on software architectural assumptions. The Journal of Systems and Software, 113(??):362–380, March 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002824.

Yang:2016:SMS

[YLA16b] Chen Yang, Peng Liang, and Paris Avgeriou. A systematic mapping study on the combination of software architecture and agile development. The Journal of Systems and Software, 111(??):157–184, January 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002125.

Yang:2017:ICS

[YLA^{+}17] Chen Yang, Peng Liang, Paris Avgeriou, Ulf Eliasson, Rogardt Heldal, Patrizio Pelliccione, and Tingting Bi. An industrial case study on an architectural assumption documentation framework. The Journal of Systems and Software, 134(??):190–210, December 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301966.

Yang:2018:EJW

[YLC18] Jeong Yang, Young Lee, and Kai H. Chang. Evaluations of JaguarCode: a web-based objectoriented programming environment with static and dynamic visualization. The Journal of Systems and Software, 145(??):147–163, November 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121830147X.

Yu:2012:TAD

[YLCZ12] Kai Yu, Mengxiang Lin, Jin Chen, and Xiangyu Zhang. Towards automated debugging in software evolution: Evaluating delta debugging on real regression bugs from the developers’ perspectives. The Journal of Systems and Software, 85(10):2305–2317, October 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211002743.

Yang:2016:MAR

[YLXZ16] Jianfeng Yang, Yu Liu, Min Xie, and Ming Zhao. Modeling and analysis of reliability of multirelease open source software incorporating both fault detection and correction processes. The Journal of Systems and Software, 115(??):102– 110, May 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000200.

Yu:2017:BNB

[YLYL17] Xiao Yu, Jin Liu, Zijiang Yang, and Xiao Liu. The Bayesian network based program dependence graph and its application to fault localization. The Journal of Systems and Software, 134(??):44–53, December 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301796.

Yu:2016:CBE

[YLZ^{+}16] Qihong Yu, Jiguo Li, Yichen Zhang, Wei Wu, Xinyi Huang, and Yang Xiang. Certificate-based encryption resilient to key leakage. The Journal of Systems and Software, 116(??):101–112, June 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215001326.

Ying:2013:RLA

[YM13] Ming Ying and James Miller. Refactoring legacy AJAX applications to improve the efficiency of the data exchange component. The Journal of Systems and Software, 86(1):72–88, January 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212002129.

Yaman:2017:ICE

[YMM^{+}17] Sezin Gizem Yaman, Myriam Munezero, Jürgen Münch, Fabian Fagerholm, Ossi Syd, Mika Aaltola, Christina Palmu, and Tomi Männistö. Introducing continuous experimentation in large software-intensive product and service organisations. The Journal of Systems and Software, 133(??):195–211, November 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0164121217301474.

See corrigendum [YMM^{+}19].

Yaman:2019:CIC

[YMM^{+}19] Sezin Gizem Yaman, Myriam Munezero, Jürgen Münch, Fabian Fagerholm, Ossi Syd, Mika Aaltola, Christina Palmu, and Tomi Männistö. Corrigendum to “Introducing continuous experimentation in large software-intensive product and service organisations” [The Journal of Systems and Software 133 (2017) 195–211]. The Journal of Systems and Software, 151(??): 119, May 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300329.

See [YMM^{+}17].

Yang:2011:FTF

[YSDT11] Yan Yang, Samia Souissi, Xavier Défago, and Makoto Takizawa. Faulttolerant flocking for a group of autonomous mobile robots. The Journal of Systems and Software, 84(1):29–36, January 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yanes:2017:OBR

[YSG17] Nacim Yanes, Sihem Ben Sassi, and Henda Hajjami Ben Ghezala. Ontologybased recommender system for COTS components. The Journal of Systems and Software, 132(??):283–297, October 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217301590.

Yong:2013:CCT

[YSJ13] Jianming Yong, Weiming Shen, and Anne James. Collaborative computing technologies and systems. The Journal of Systems and Software, 86(7):1725–1726, July 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213000708.

Yeh:2010:TRR

[YSL^{+}10] Kuo-Hui Yeh, Chunhua Su, N. W. Lo, Yingjiu Li, and Yi-Xiang Hung. Two robust remote user authentication protocols using smart cards. The Journal of Systems and Software, 83(12):2556– 2565, December 2010.

CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yu:2014:ATC

[YSSaR14] Tingting Yu, Ahyoung Sung, Witawas Srisa-an, and Gregg Rothermel. An approach to testing commercial embedded systems. The Journal of Systems and Software, 88(??):207–230, February 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213002604.

Yang:2013:ROM

[YTW^{+}13] Yin Yang, Zhihu Tan, Jiguang Wan, Changsheng Xie, Jie Yu, and Jian He. A reliability optimization method for RAID-structured storage systems based on active data migration. The Journal of Systems and Software, 86(2):468–484, February 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002683.

Yan:2013:CSC

[YWEL^{+}13] Xuehu Yan, Shen Wang, Ahmed A. Abd El-Latif, Jianzhi Sang, and Xiamu Niu. Corrigendum to ‘T. Chen, K. Tsao, Threshold visual secret sharing by random grids’ [J. Syst. Softw. 84 (2011) 1197–1208]. The Journal of Systems and Software, 86(5):1462–1463, May 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000253.

See [CT11b].

Yang:2011:HCS

[YWHL11] Cheng-Hsing Yang, Sheng-Chang Wu, Shu-Chien Huang, and Yih-Kai Lin. Huffman-code strategies to improve MFCVQbased reversible data hiding for VQ indexes. The Journal of Systems and Software, 84(3):388–396, March 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wang:2011:RDA

[yWpNyL11] Xiang yang Wang, Pan pan Niu, and Ming yu Lu. A robust digital audio watermarking scheme using wavelet moment invariance. The Journal of Systems and Software, 84(8):1408– 1421, August 2011. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wang:2013:RBC

[yWpWyYpN13] Xiang yang Wang, Chun peng Wang, Hong ying Yang, and Pan pan Niu. A robust blind color image watermarking in quaternion Fourier transform domain. The Journal of Systems and Software, 86(2):255–277, February 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002312.

Yang:2011:DHS

[YWTW11] Cheng-Hsing Yang, Chi-Yao Weng, Hao-Kuan Tso, and Shiuh-Jeng Wang. A data hiding scheme using the varieties of pixel-value differencing in multimedia images. The Journal of Systems and Software, 84 (4):669–678, April 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yang:2010:VPL

[YWWS10] Cheng-Hsing Yang, Chi-Yao Weng, Shiuh-Jeng Wang, and Hung-Min Sun. Varied PVD + LSB evading detection programs to spatial domain in data embedding systems. The Journal of Systems and Software, 83 (10):1635–1643, October 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yan:2018:NDL

[YXH^{+}18] Ruibo Yan, Xi Xiao, Guangwu Hu, Sancheng Peng, and Yong Jiang. New deep learning method to detect code injection attacks on hybrid applications. The Journal of Systems and Software, 137(??):67–77, March 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217302571.

Yang:2018:EVS

[YXP^{+}18] Wenhua Yang, Chang Xu, Minxue Pan, Chun Cao, Xiaoxing Ma, and Jian Lu. Efficient validation of self-adaptive applications by counterexample probability maximization. The Journal of Systems and Software, 138(??):82–99, April 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0164121217303023.

Yang:2016:EPA

[YYS^{+}16] Guangyang Yang, Jia Yu, Wenting Shen, Qianqian Su, Zhangjie Fu, and Rong Hao. Enabling public auditing for shared data in cloud storage supporting identity privacy and traceability. The Journal of Systems and Software, 113(??):130–139, March 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121500268X.

Yu:2015:CAR

[YZC15] Dongjin Yu, Yanyan Zhang, and Zhenli Chen. A comprehensive approach to the recovery of design pattern instances based on sub-patterns and method signatures. The Journal of Systems and Software, 103(??):1– 16, May 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121500014X.

Yang:2013:LQA

[YZG^{+}13] Liang Huai Yang, Jian Zhou, Weihua Gong, Jiakui Zhao, and Lijun Chen. Lifetime and QoS-aware energy-saving buffering schemes. The Journal of Systems and Software, 86(5):1408– 1425, May 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121213000125.

Yin:2014:EDS

[YZL^{+}14] Bo Yin, Siwang Zhou, Yaping Lin, Yonghe Liu, and Yupeng Hu. Efficient distributed skyline computation using dependency-based data partitioning. The Journal of Systems and Software, 93(??):69–83, July 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214000703.

Yu:2018:EDS

[YZY^{+}18] Dongjin Yu, Ping Zhang, Jiazha Yang, Zhenli Chen, Chengfei Liu, and Jie Chen. Efficiently detecting structural design pattern instances based on ordered sequences. The Journal of Systems and Software, 142(??):35–56, August 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300670.

Zimmer:2012:OFC

[ZA12] P. Ann Zimmer and Joanne M. Atlee. Ordering features by category. The Journal of Systems and Software, 85(8):1782–1800, August 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000751.

Zaina:2015:DMU

[ZÁ15] Luciana A. M. Zaina and Alexandre Álvaro. A design methodology for user-centered innovation in the software development area. The Journal of Systems and Software, 110(??):155–177, December 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121500182X.

Zarour:2015:IBP

[ZADA15] Mohammad Zarour, Alain Abran, Jean-Marc Desharnais, and Abdulrahman Alarifi. An investigation into the best practices for the successful design and implementation of lightweight software process assessment methods: a systematic literature review. The Journal of Systems and Software, 101(??):180–192, March 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214002726.

Zhao:2010:GNQ

[ZADM10] Liang Zhao, Ahmed Y. Al-Dubai, and Geyong Min. GLBM: a new QoS aware multicast scheme for wireless mesh networks. The Journal of Systems and Software, 83(8):1318–1326, August 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhang:2019:UWT

[ZAY19] Man Zhang, Shaukat Ali, and Tao Yue. Uncertainty-wise test case generation and minimization for Cyber-Physical Systems. The Journal of Systems and Software, 153(??):1–21, July 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121219300561.

Zhou:2017:RTC

[ZCC^{+}17] Junlong Zhou, Kun Cao, Peijin Cong, Tongquan Wei, Mingsong Chen, Gongxuan Zhang, Jianming Yan, and Yue Ma. Reliability and temperature constrained task scheduling for makespan minimization on heterogeneous multi-core platforms. The Journal of Systems and Software, 133(??):1–16, November 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301607.

Zhang:2019:FLS

[ZCC^{+}19] Weiqiang Zhang, Shing-Chi Cheung, Zhenyu Chen, Yuming Zhou, and Bin Luo. File-level socio-technical congruence and its relationship with bug proneness in OSS projects. The Journal of Systems and Software, 156(??):21–40, October 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301177.

Zhang:2017:PCC

[ZcKS17] Fanlong Zhang, Siau cheng Khoo, and Xiaohong Su. Predicting change consistency in a clone group. The Journal of Systems and Software, 134(??):105–119, December 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217301875.

Zhang:2011:NPS

[ZCT^{+}11] Zhenyu Zhang, W. K. Chan, T. H. Tse, Y. T. Yu, and Peifeng Hu. Non-parametric statistical fault localization. The Journal of Systems and Software, 84(6):885–905, June 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhang:2016:TMA

[ZCY^{+}16] Tao Zhang, Jiachi Chen, Geunseok Yang, Byungjeong Lee, and Xiapu Luo. Towards more accurate severity prediction and fixer recommendation of software bugs. The Journal of Systems and Software, 117(??):166–184, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216000765.

Zhao:2011:EGD

[ZCZZ11] Yanchang Zhao, Jie Cao, Chengqi Zhang, and Shichao Zhang. Enhancing grid-density based clustering for high dimensional data. The Journal of Systems and Software, 84 (9):1524–1539, September 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211000690.

Zhang:2011:TVA

[ZDC^{+}11] Tianle Zhang, Zhihui Du, Yinong Chen, Xiang Ji, and Xiaoying Wang. Typical Virtual Appliances: an optimized mechanism for virtual appliances provisioning and management. The Journal of Systems and Software, 84(3):377– 387, March 2011. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zanoni:2015:AML

[ZFS15] Marco Zanoni, Francesca Arcelli Fontana, and Fabio Stella. On applying machine learning techniques for design pattern detection. The Journal of Systems and Software, 103(??):102–117, May 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121215000217.

Zheng:2019:TUB

[ZFY^{+}19] Wei Zheng, Chen Feng, Tingting Yu, Xibing Yang, and Xiaoxue Wu. Towards understanding bugs in an open source cloud management stack: an empirical study of OpenStack software bugs. The Journal of Systems and Software, 151(??):210–223, May 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219300354.

Zhao:2010:PSA

[ZG10] Jianjie Zhao and Dawu Gu. Provably secure authenticated key exchange protocol under the CDH assumption. The Journal of Systems and Software, 83(11):2297–2304, November 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zou:2010:NGH

[ZGL^{+}10] Dexuan Zou, Liqun Gao, Steven Li, Jianhua Wu, and Xin Wang. A novel global harmony search algorithm for task assignment problem. The Journal of Systems and Software, 83(10):1678–1688, October 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhu:2013:EEE

[ZGSH13] Xiaomin Zhu, Rong Ge, Jinguang Sun, and Chuan He. 3E: Energyefficient elastic scheduling for independent tasks in heterogeneous computing systems. The Journal of Systems and Software, 86(2):302–314, February 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212002336.

Zhi:2015:CBQ

[ZGYS^{+}15] Junji Zhi, Vahid Garousi-Yusifoglu, Bo Sun, Golara Garousi, Shawn Shahnewaz, and Guenther Ruhe. Cost, benefits and quality of software development documentation: a systematic mapping. The Journal of Systems and Software, 99(??):175–198, January 2015. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214002131.

Zhao:2013:EHW

[ZGZ^{+}13] Xinjie Zhao, Shize Guo, Fan Zhang, Tao Wang, Zhijie Shi, Huiying Liu, Keke Ji, and Jing Huang. Efficient Hamming weight-based sidechannel cube attacks on PRESENT. The Journal of Systems and Software, 86(3):728–743, March 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212003081.

Zhang:2012:DTC

[Zha12a] Shichao Zhang. Decision tree classifiers sensitive to heterogeneous costs.

The Journal of Systems and Software, 85(4):771– 779, April 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002603.

Zhang:2012:NNS

[Zha12b] Shichao Zhang. Nearest neighbor selection for iteratively k NN imputation. The Journal of Systems and Software, 85 (11):2541–2552, November 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212001586.

Zhao:2016:POS

[Zha16] Wenbing Zhao. Performance optimization for state machine replication based on application semantics: a review. The Journal of Systems and Software, 112(??):96–109, February 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002411.

Zhu:2012:EAS

[ZHAY12] Yan Zhu, Hongxin Hu, Gail-Joon Ahn, and Stephen S. Yau. Efficient audit service outsourcing for data integrity in clouds. The Journal of Systems and Software, 85(5):1083–1095, May 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211003189.

Zhu:2011:BAF

[ZHGL11] Xiaomin Zhu, Chuan He, Rong Ge, and Peizhong Lu. Boosting adaptivity of fault-tolerant scheduling for real-time tasks with service requirements on clusters. The Journal of Systems and Software, 84(10):1708–1716, October 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211001087.

Zhang:2017:RMB

[ZHH^{+}17] Panfeng Zhang, Ping Huang, Xubin He, Hua Wang, and Ke Zhou. Resemblance and mergence based indexing for high performance data deduplication. The Journal of Systems and Software, 128(??):11–24, June 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217300389.

Zhang:2010:FLT

[ZJC^{+}10] Zhenyu Zhang, Bo Jiang, W. K. Chan, T. H. Tse, and Xinming Wang. Fault localization through evaluation sequences. The Journal of Systems and Software, 83(2):174– 187, February 2010. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhang:2010:SQF

[ZJL10] Shichao Zhang, Zhi Jin, and Jingli Lu. Summary queries for frequent itemsets mining. The Journal of Systems and Software, 83(3):405–411, March 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhang:2011:MDI

[ZJZ11] Shichao Zhang, Zhi Jin, and Xiaofeng Zhu. Missing data imputation by utilizing information within incomplete instances. The Journal of Systems and Software, 84 (3):452–459, March 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhang:2017:MLF

[ZJZ^{+}17] Miao Zhang, Shujuan Jiang, Yanmei Zhang, Xingya Wang, and Qiao Yu. A multi-level feedback approach for the class integration and test order problem. The Journal of Systems and Software, 133(??):54–67, November 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217301802.

Zalewski:2013:BAE

[ZK13] Andrzej Zalewski and Szymon Kijas. Beyond ATAM: Early architecture evaluation method for large-scale distributed systems. The Journal of Systems and Software, 86(3):683–697, March 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212003032.

Zhong:2012:IPA

[ZL12a] Jinmin Zhong and Xuejia Lai. Improved preimage attack on oneblock MD4. The Journal of Systems and Software, 85(4):981–994, April 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002937.

Zhou:2012:CBF

[ZL12b] Qing Zhou and Xiaofeng Liao. Collision-based flexible image encryption algorithm. The Journal of Systems and Software, 85(2):400–407, February 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002329.

Zhang:2017:RPA

[ZL17] Yulai Zhang and Guiming Luo. Recursive prediction algorithm for non-stationary Gaussian Process. The Journal of Systems and Software, 127(??):295–301, May 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121216301455.

Zhang:2014:DFD

[ZLC^{+}14] Dazhi Zhang, Donggang Liu, Christoph Csallner, David Kung, and Yu Lei. A distributed framework for demanddriven software vulnerability detection. The Journal of Systems and Software, 87(??):60–73, January 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121213002288.

Zhang:2013:SSW

[ZLD13] Jinghui Zhang, Junzhou Luo, and Fang Dong. Scheduling of scientific workflow in non-dedicated heterogeneous multicluster platform. The Journal of Systems and Software, 86(7):1806–1818, July 2013. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212002877.

Zhang:2010:TPS

[ZLG10] Pengcheng Zhang, Bixin Li, and Lars Grunske. Timed Property Sequence Chart. The Journal of Systems and Software, 83(3):371–390, March 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhang:2012:STC

[ZLL^{+}12] Dazhi Zhang, Donggang Liu, Yu Lei, David Kung, Christoph Csallner, Nathaniel Nystrom, and Wenhua Wang. Sim-Fuzz: Test case similarity directed deep fuzzing. The Journal of Systems and Software, 85(1):102–111, January 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S016412121100197X.

Zhou:2014:FRB

[ZLmLN14] Lijian Zhou, Wanquan Liu, Zhe ming Lu, and Tingyuan Nie. Face recognition based on curvelets and local binary pattern features via using local property preservation. The Journal of Systems and Software, 95(??):209–216, September 2014. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S016412121400106X.

Zhang:2010:MDC

[ZLT10] Hongyu Zhang, Yuan-Fang Li, and Hee Beng Kuan Tan. Measuring design complexity of Semantic Web ontologies. The Journal of Systems and Software, 83(5):803–814, May 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhang:2012:CCB

[ZLW^{+}12] Leo Yu Zhang, Chengqing Li, Kwok-Wo Wong, Shi Shu, and Guanrong Chen. Cryptanalyzing a chaos-based image encryption algorithm using alternate structure. The Journal of Systems and Software, 85 (9):2077–2085, September 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121200101X.

Zhang:2011:IPM

[ZLZ11] Ruijun Zhang, Jie Lu, and Guangquan Zhang. An information presentation method based on tree-like super entity component. The Journal of Systems and Software, 84(8):1306–1318, August 2011. CODEN JSSODM.

ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhang:2012:ERB

[ZM12] Jianhong Zhang and Jane Mao. An efficient RSA-based certificateless signature scheme. The Journal of Systems and Software, 85(3):638–642, March 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211002445.

Zhong:2018:MRM

[ZM18] Hao Zhong and Hong Mei. Mining repair model for exceptionrelated bug. The Journal of Systems and Software, 141(??):16–31, July 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300505.

Zerrougui:2014:TNA

[ZMB14] Salim Zerrougui, Farid Mokhati, and Mourad Badri. Toward a new aspect-mining approach for multi-agent systems. The Journal of Systems and Software, 98(??):9–24, December 2014. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214001824.

Zimmermann:2012:RAM

[ZMK12] Olaf Zimmermann, Christoph Miksovic, and Jochen M. Küster. Reference architecture, metamodel, and modeling principles for architectural knowledge management in information technology services. The Journal of Systems and Software, 85 (9):2014–2033, September 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121212001343.

Zhang:2010:CCM

[ZML10] Pengcheng Zhang, Henry Muccini, and Bixin Li. A classification and comparison of model checking software architecture techniques. The Journal of Systems and Software, 83(5):723–744, May 2010. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhang:2017:FGA

[ZML17] Rui Zhang, Hui Ma, and Yao Lu. Fine-grained access control system based on fully outsourced attribute-based encryption. The Journal of Systems and Software, 125(??):344–353, March 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL /

/www.sciencedirect.
com/science/article/
pii/S0164121216302606.

Zhu:2017:EFA

[ZP17] Mengmeng Zhu and Hoang Pham. Environmental factors analysis and comparison affecting software reliability in development of multirelease software. The Journal of Systems and Software, 132(??):72–84, October 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121217301036.

Zhang:2016:PSS

[ZS16] Yiji Zhang and Raul Santelices. Prioritized static slicing and its application to fault localization. The Journal of Systems and Software, 114(??):38–53, April 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121500237X.

Zhou:2019:ABI

[ZSB19] Bowen Zhou, Satish Narayana Srirama, and Rajkumar Buyya. An auctionbased incentive mechanism for heterogeneous mobile clouds. The Journal of Systems and Software, 152(??):151–164, June 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219300548.

Zein:2016:SMS

[ZSG16] Samer Zein, Norsaremah Salleh, and John Grundy. A systematic mapping study of mobile application testing techniques. The Journal of Systems and Software, 117(??):334–356, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216300140.

Zhang:2014:NCM

[ZT14] Miao Zhang and Xiaojun Tong. A new chaotic map based image encryption schemes for several image formats. The Journal of Systems and Software, 98(??):140–154, December 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001915.

Zernadji:2016:IQR

[ZTCZ16] Tarek Zernadji, Chouki Tibermacine, Foudil Cherif, and Amina Zouioueche. Integrating quality requirements in engineering web service orchestrations. The Journal of Systems and Software, 122(??):463–483, December 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215002423.

Zhou:2018:ISI

[ZTPT18] Zhi Quan Zhou, Dave Towey, Pak-Lok Poon, and T. H. Tse. Introduction to the special issue on test oracles. The Journal of Systems and Software, 136(??):187, February 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121217301814.

Zhang:2011:CEI

[ZTZ^{+}11] Hongyu Zhang, Hee Beng Kuan Tan, Lu Zhang, Xi Lin, Xiaoyin Wang, Chun Zhang, and Hong Mei. Checking enforcement of integrity constraints in database applications based on code patterns. The Journal of Systems and Software, 84 (12):2253–2264, December 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121211001567.

Zhou:2015:STA

[ZW15] Junlong Zhou and Tongquan Wei. Stochastic thermalaware real-time task scheduling with considerations of soft errors. The Journal of Systems and Software, 102(??):123– 133, April 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121214002817.

Zhang:2019:EAV

[ZWC^{+}19] Xinqian Zhang, Tingming Wu, Mingsong Chen, Tongquan Wei, Junlong Zhou, Shiyan Hu, and Rajkumar Buyya. Energy-aware virtual machine allocation for cloud with resource reservation. The Journal of Systems and Software, 147(??):147–161, January 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218302152.

Zheng:2018:LMS

[ZWF^{+}18] Yan Zheng, Zan Wang, Xiangyu Fan, Xiang Chen, and Zijiang Yang. Localizing multiple software faults based on evolution algorithm. The Journal of Systems and Software, 139(??):107– 123, May 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121218300268.

Zhang:2018:WSD

[ZWM^{+}18] Neng Zhang, Jian Wang, Yutao Ma, Keqing He, Zheng Li, and Xiaoqing (Frank) Liu. Web service discovery based on goal-oriented query expansion. The Journal of Systems and Software, 142(??):73–91, August 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121218300748.

Zhang:2017:HEC

[ZXC^{+}17] Ruiqing Zhang, Chang Xu, S. C. Cheung, Ping Yu, Xiaoxing Ma, and Jian Lu. How effectively can spreadsheet anomalies be detected: an empirical study. The Journal of Systems and Software, 126(??):87–100, April 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216300103.

Zhou:2010:LSL

[ZXG10] Shuigeng Zhou, Ming Xu, and Jihong Guan. LES-SON: a system for lecture notes searching and sharing over Internet. The Journal of Systems and Software, 83(10):1851– 1863, October 2010. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhou:2010:ACM

[ZXL10] Yuming Zhou, Baowen Xu, and Hareton Leung. On the ability of complexity metrics to predict fault-prone classes in object-oriented systems. The Journal of Systems and Software, 83(4):660– 674, April 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhang:2011:PBM

[ZXTT11] Chong Zhang, Weidong Xiao, Daquan Tang, and Jiuyang Tang. P2Pbased multidimensional indexing methods: a survey. The Journal of Systems and Software, 84 (12):2348–2362, December 2011. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121211001968.

Zhang:2018:SUU

[ZYA^{+}18] Man Zhang, Tao Yue, Shaukat Ali, Bran Selic, Oscar Okariz, Roland Norgre, and Karmele Intxausti. Specifying uncertainty in use case models. The Journal of Systems and Software, 144(??):573–603, October 2018. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0164121218301316.

Zhou:2019:AJM

[ZYY^{+}19] Yu Zhou, Xin Yan, Wenhua Yang, Taolue Chen, and Zhiqiu Huang. Augmenting Java method comments generation with context information based on neural networks. The Journal of Systems and Software, 156(??):328–340, October 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121219301529.

Zhang:2017:TAC

[ZYZ^{+}17] Long Zhang, Lanfei Yan, Zhenyu Zhang, Jian Zhang, W. K. Chan, and Zheng Zheng. A theoretical analysis on cloning the failed test cases to improve spectrum-based fault localization. The Journal of Systems and Software, 129(??):35–57, July 2017. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217300808.

Zhang:2012:LRA

[ZYZL12] Xin-Chang Zhang, Mei-Hong Yang, Xiao-Jing Zhu, and Wan-Ming Luo. A loss recovery approach for reliable application layer multicast. The Journal of Systems and Software, 85(5):1198– 1204, May 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121212000076.

Zhang:2014:GCT

[ZYZZ14] Zhiqiang Zhang, Jun Yan, Yong Zhao, and Jian Zhang. Generating combinatorial test suite using combinatorial optimization. The Journal of Systems and Software, 98(??):191–207, December 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121214001939.

Zhao:2012:FCS

[ZZ12] Xingwen Zhao and Fangguo Zhang. Fully CCA2 secure identitybased broadcast encryption with black-box accountable authority. The Journal of Systems and Software, 85(3):708–716, March 2012. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S016412121100255X.

Zhang:2016:HMI

[ZZ16] Hailong Zhang and Yongbin Zhou. How many interesting points should be used in a template attack? The Journal of Systems and Software, 120(??):105–113, October 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121216301224.

Zhang:2018:EUU

[ZZC18] Xiao-Yi Zhang, Zheng Zheng, and Kai-Yuan Cai. Exploring the usefulness of unlabelled test cases in software fault localization. The Journal of Systems and Software, 136(??):278–290, February 2018. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0164121217301589.

Zhu:2015:CAE

[ZZP15] Mengmeng Zhu, Xuemei Zhang, and Hoang Pham. A comparison analysis of environmental factors affecting software reliability. The Journal of Systems and Software, 109(??):150–160, November 2015. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0164121215000977.